A Complete Bibliography of Publications in Statistics and Probability Letters: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

02 August 2023 Version 1.25

Title word cross-reference

(1, 2) [Yan19a]. (α, β) [Kay15]. (∞, ∞) [BO19]. (k_{1}, k_{2}) [UK18]. (n − k + 1) [MF15, ZY10a]. (q, d) [BG16]. (X + Y) [AV13b]. 1 [Hel19, Red17, SC15, Sze10]. 1/2 [Efr10]. {2} [Sma15, Jac13, JKD15a, JKD15b, Kha14, LM17, MK10, Osȩ14a, WL11]. 2^{k} [Lu16a, Lu16b]. 2^{n−p} [Yan13]. 2 × 2 [BB12, MBTC13]. 2 × c [SB16]. 3 [CGLN17, HWB10]. 3x + 1 [DP16]. > [HB18]. > 1/2 [DO11a]. [0, 1] [LP19a]. L^{1}(ℓ^{q}) [Osȩ15a]. _{k} [BML14]. _{r} [VCM14]. A [CDS17, MPA12, PPM18, KDW15]. A_{1} [Osȩ13a]. α [AH12b, hCyP15, GT14, LT19b, MY14, MU13, Mic11, MM14b, PP14b, Sre12, Yan17b, ZZT17, ZYS19, ZZ13b, ZLC18, ZZ19a]. AR(1) [CL10, Mar16, PPN10]. AR(p) [WN14b]. ARH(1) [RMÁL19]. β [EM10, JW16, Kre18, Poi19, RBY10, Su10, XXY17, YZ16]. β → ∞ [JW16]. C[0, 1] [Mor18]. C[_{0}, 1] [KM19]. C^{∞} [LL17c]. D [CDS17, DS15, Jac11, KHN12, Sin19, Sma14, ZP12, Dav12b, ÖA15, Rez15].

D_{s} [KHN12]. D ∩ L [BS12a]. δ [Ery12, PB15a]. d ≥ 3 [DH19]. DS [WZ17]. d × R [BS19b]. E [DS15, HR19a, JKD14, KDW15]. E(f_{NOD}) [CKF17]. ℓ_{1} [AH11]. ℓ_{1}^{N} [Ose19]. ℓ_{p} [Zen14]. ϵ [BCD19]. exp(x) [BM11]. F [MMM13]. G [Che10d, LW15c, RH11, Ren13, Son15, WSMG18, HC10b]. γ < −1/2 [SJ14]. GARCH(p, q) [Roh13]. h [YYY17]. H ∈ (0, 1/2) [SC12b]. I_{1,p} [BCGFR13, JB13a]. I_{2,p} [BCGFR13, JB13a]. INGARCH(p, q) [CW16]. K [Mat12, MM14a, Nag19, DZZ13, DZF18, Ery14, Jas19, JS19, KV13, KLK12, LFM15, MGSL11, MM11, PS18a, RIK16, RA13b, RA13a, TE17, ZK10, Zha18]. L [KO11, LLC^{+}12, LLCG12, Nag13, OK15]. L^{1} [ANRW15, DS11, ZW15a, dMO18]. L^{2} [Osȩ13b, Tap10, LL17d]. L^{2} log L [Osȩ14b]. L^{∞} [Che10b]. L^{p} [Ben17, Izu13, LZ10, LD10, OS15, TJS13]. L_{0} [Ciu15]. L_{1} [CW11, DFKK12, MMF17]. L_{2} [CLQ12, EQ14, Xia17, ZWHQ15]. L_{p} [BBP17, GLM19, LM15a, YH16, ZBL19]. lim sup [Dav13]. M [DSX13, DZX18, EBG16, FG12, GLS13, HWZ14, Lin17b, Mao15a, MYA15, Vir16, WZT12, Ana17, Ber13a, FC10, Raq10, SY10]. MA(p) [WN14b]. R [LYC14, YWC10]. R^{d} [CG13, IMU10, Yan12]. R^{n} [VY12]. R_{+} [MS12b]. Z [CLEAMS16]. Z_{+}^{d} [MZ17]. L_{1} [ZJZ13]. N [CH12, Had11, Had13, HWYW12, PH13a, RR14, WHRY11, WW13b, Bél11, CP11, DZZ13, DS15, LFM15, MM11, MF15, RIK16, Raq10, ZK10, ZY10a, Zha18]. n ≡ 3 mod 4 [JKD14]. ν [Xu12b]. O [DSS18, LW12, YLS10]. P [Llo10, Llo12, AIM19, BMP10b, CS17, Che13b, DT10, Gre12, LX10a, PJ10, RW16b, SL12, SDNS16, XT11]. P (X < Y) [Dat13, TNP17]. p/n → 0 [Bao12, Xie13]. p = 3 [Pin13]. Φ [CW19, PL15a, Sze11, Zho10]. Φ_{p} [Pro13]. π [WX18, Xu15]. Ψ [HS12c]. Q [Il’18, Myr19, Bry14, OT19, Sza15, Wan16]. Q_{Φ_{1}} [Yu11]. Q_{Φ_{2}} [Yu11]. R [Ric10, Wan13c, ÖA15]. r_{W 2} [PRS15]. ρ^{+} [GZ13]. ρ^{−} [XW17]. S [GB16, Yor14, LZZ15b]. σ [LVMS14, Vid18]. Σ^{2} [ZG17a]. T [KM13c, AP13, AN15, BV10, BBP17, BSWS13, Car10, Che18b, Fan17, FS10, HHY11, LMH14, ORO15, PT13, PLN16, TMS11, VM12, WY16]. t_{3} [AN10]. τ [vDLMW19]. θ [CW12]. U [CD18b, DR02, DR15, EQ14, FMPV12, GD15, GD18, Nas12, VB17]. V [Lee12b, AKM19]. W [Pak18]. W_{2} [Bob13]. W × F [WWD15]. X/(X + Y) [AV13b]. y [XF17]. Z^{2} [Mis13]. Zd [Gao19].

-Brownian [Bry14, GT14, ZZ13b, ZLC18, RH11, WSMG18]. -class [Pro13]. -compactness [LVMS14]. -concave [DZF18]. -conformal [AIM19]. -consistent [DFKK12, RR14]. -constraints [AH11]. -convergence [LL17c, OS15]. -convex [Osȩ14a]. -curves [Tap10]. -D [Hel19, SC15, Kha14]. -deficiency [ZJZ13]. -demimartingale [WHRY11]. -demimartingales [CH12, Had11, Had13, HWYW12, WW13b]. -densities [OT19]. -dependent [Ana17]. -dimensional [Dav12b, HWB10]. -discrepancies [CLQ12]. -discrepancy [EQ14, ZWHQ15]. -distribution [Che18b, PT13]. -divergences [GB16]. -entropy [RBY10, CW19]. -error [LZ10]. -estimates [Vir16]. -estimation [DSX13, Mao15a, WZT12]. -estimator [DZX18, FG12, HWZ14]. -estimators [EBG16, Lin17b]. -expectation [Che10d]. -expectations [HC10b]. -exponential [DSS18]. -fields [Vid18]. -folded [XT11]. -functions [AKM19]. -FWER [MGSL11]. -Gaussian [MYA15]. -Hermite [JW16]. -identical [Il’18]. -independence [Il’18]. -independent [Myr19]. -inverses [Sma15]. -Laguerre [JW16]. -level [Jac13, JKD15a, JKD15b, LZZ15b, Mat12]. -Lévy [Ren13]. -martingales [Osȩ13b]. -means [MM14a]. -method [CW12]. -metric [YH16]. -mixing [hCyP15, GZ13, Kre18, Poi19, Sze11, XXY17, XW17, Zho10]. -model [YZ16]. -moment [PL15a]. -moments [Nag13]. -monotone [Rez15]. -monotonicity [Sre12]. -nearest [KLK12]. -norm [GLM19]. -Normal [Sza15, Son15]. -optimal [CDS17, DS15, HR19a, KHN12, LYC14, MPA12, PPM18, Sin19, Sma14, WZ17, ZP12]. -optimality [Jac11, JKD14]. -Ornstein [Wan16]. -out-of- [DZZ13, LFM15, MM11, MF15, RIK16, Raq10, ZY10a, Zha18]. -permutation [Yor14]. -Poincaré [ZW15a]. -principal [MK10]. -quantile [XT11]. -quantiles [BBP17]. -random [BV10, Yan19a]. -record [JS19]. -records [RA13b, RA13a]. -regression [GLS13]. -regularly [YLS10]. -risk [Ben17]. -runs [UK18]. -SDEs [LW15c]. -selfdecomposable [MU13]. -shock [Ery12, PB15a]. -Slepian-processes [BG16]. -solutions [ZBL19]. -square [Wan13c]. -squared [Ric10]. -stability [Ber13a]. -stable [LT19b, MY14, Mic11, MM14b, Yan17b, ZZT17, ZYS19, ZZ19a]. -statistic [Fan17, HHY11]. -statistics [DR15, DR02, FMPV12, GD15, GD18, KO11, LLC^{+}12, LLCG12, Nas12, OK15, VB17]. -Stieltjes [OT19]. -subexponential [LW12]. -symmetric [JB13a, BCGFR13, Nag19]. -TARCH [EM10]. -th [Jas19, PJ10]. -th-order [FC10]. -torus [PH13a]. -transforms [YYY17]. -type [LM17, EQ14]. -uniform [Lee12b]. -urn [CGLN17]. -value [Gre12, LX10a, SL12, Llo12]. -valued [KM19, Mor18, Ose19]. -values [BMP10b, Che13b, RW16b, SDNS16, Llo10]. -Wasserstein [Xia17]. -weak [HS12c]. -wise [PS18a].

/1 [BML14, VCM14].

1 [BML14, Bri15, PSC12, VCM14]. 110 [Ton17]. 12 [HB18]. 123 [BSCQ18]. 137 [Suz20]. 145 [Sko20]. 1queue [Kin12].

27 [Bro11]. 2RV [LH14a].

3 [ZP15].

57 [DR15].

68 [Duc10].

78 [Kim13, SF12].

81 [Kak11a, MN13]. 82 [Abu12a, Bai13, DFL13, DFT13]. 83 [MM14c, RA13b]. 87 [BRBB23].

A-optimal [ASP11]. Aalen [BTT16, Ane17]. ABC [SHL15]. Abelian [Myr19]. aberration [BDM17, ySDM14, YYS18]. absence [AG18, MHH11]. Absolute [Nak13b, BS12a, BK13, Ern17, HR15, NKKY13, PP11, Ush11b, WYL10, Wu14, XT11]. Absolutely [Neu13, dSF12, Mes14]. absorbing [HZZ19, KW14b]. absorption [Mil19b]. Abundance [LZL^{+}19, MHH11]. Accelerated [KFHS11, Hat12, Zha11b]. Accelerating [CH13, PH13a]. acceleration [Sak15]. acceptance [RMPB17]. accessibility [TU17]. according [CK16a]. accumulated [Hew17]. accurate [SLC16]. achieving [SW11]. Acknowledgement [Bro11]. across [TM16]. action [HY13, Ose19]. active [BZV11, MM11, VAZB10, Zha18]. Acyclic [CC16]. ad [Llo12]. adaptation [vWvZ17]. adapted [BH11, QX13, THC17]. Adaptive [KZ14, PB15b, WW14, BB16, BL10b, BB10, BB11c, BCNM15, DFH18, Li15b, PP14b, RW17, RZ10, SS11a, WST14, ZZLH15, ZKG10]. adding [CG17, dB13]. Additive [SA12b, ASP11, BTT16, Dun16b, EM16, HL13a, Kak11a, Kak11b, LV11b, Lia12b, Lia13, LHM14, LXL19, MZZ15, PSS12, PT19, QCZ15, SV19, SZ13b, TSL18, WCD19, YY16]. Adeles [Urb12, Yas13]. ADHD [ASN16]. adiabatic [Kov10]. adjacent [HS10]. adjoint [Min17]. Adjusted [LQ19, Ric10, Bak18, RW16b, SS11b, SDNS16, TPNSN19, XSL^{+}14, ZX18]. Adjusting [BDB^{+}10]. adjustment [Kak12, Llo12, Lu16a, KOR15b]. adjustments [FP13, Kak11a, Kak11b]. Admissibility [Kab11]. Adomian [KV17a]. advantage [HS11a]. advection [Sab18]. adverse [Pac18]. Aeppli [MB14]. affine [GS17a, KL17, Shu18a]. after [KM18, RG18, Yak15]. against [BDM17, DdRS^{+}11, GMA12, GM17b, Han12b, OK15]. age [McC12, SKJ15]. age- [McC12]. Ageing [NV11, BNM13b, GM17a, MF15]. aggregate [BS11, BZ13, JCM15, Kau14, Sha17]. aggregating [BP15]. Aggregation [CP14, CH19, EPSU16, NP16, PS15]. aging [DM11, Nav18]. agreement [GM10b]. AIC [JT14, NMS15, WZ10]. Ait [Dun16a]. Ait-Sahalia [Dun16a]. al [JB13a]. Algebraic [Lan11, Son16, Mat12]. algebras [Fis13]. algorithm [AT14b, CMW18, CD18a, CL15, CP13b, DCR19, Giu15, HH18, HC17, HM10b, KW14a, KN17, LW15b, MM14a, MN16, NHN^{+}11, Tan15a, TZT16, VM12, WMM15, WH10, XSL^{+}14, Yao13b, ZMG10, ZHWH12, ZHWH12]. algorithms [CP17a, CJT10, JH14, KMR13]. alias [ZGZ18]. Allocation [Jon17, Bel16b, BB11c, LTP18, MBTC13, WL15a, Wri17, XZZ19, Zha18]. allocations [MM11]. Almost [CW12, CLM12, Coe15, CP13b, Mou19, Wu11b, Xin12, XW17, BBHH10, GAS13, JWCC19, KM12a, MW14a, PP19, Tan13a, ZZ15c]. almost-exact [MW14a]. almost-sure [BBHH10]. Alpha [HA13]. Alpha-Skew-Laplace [HA13]. alternating [Hew17, Rat15]. alternative [BMD13, BM17, CH11b, GMA12, LP13, Llo10, LDD15, ML17, OO11, WL14b]. alternatives [AB11, AM11, BDvdAW16, Car10, DN16, GM17b, Han12b]. always [Mis14, dB13]. Amato [Arn12]. ambiguity [JS18, ZZZ18]. American [PY18]. among [PS16b, TZ16]. amount [MPA12, PP14b, RSV10, ZW13]. amounts [BNB16]. Ampère [KM13d]. amplitude [Efr14]. amplitude-modulated [Efr14]. analog [HJS19]. analogue [MM18]. analyse [Shi18]. analyses [BCR10, HLR15]. Analysis [WR18, AT14a, AD11, Bak18, BB16, BNM13b, BL10b, BK18, CCS16, CM17, CH17, DMST13, Doe18, FXT12, Gho11a, GDNM18, HZW16, HW13b, Ilm13, IM15, JL16, Kab16, KT10, KM13b, LD16a, LCX12, LG12, LLZ13, Mas11, MNO15, Mih12, Ruk12, SM16, SWH^{+}11, Sta12, SZ13b, TL19, Tan15a, TSL18, TKO12, TM16, TR12, TA14, Vel12, WS12, WK10a, ZY11]. analyst [Won18]. Analytic [Gho11a, VM12, KS16a, Roy12]. Analytical [SGV15, CJN19]. analyzing [SB16, WL11, ZCZ18]. ancillary [NS12]. Andersen [ZY10b]. angular [HdCC17]. anisotropic [Ben17, LX10b, Mis13, Söh10]. annealing [RR11]. annuity [WWY12]. ANOVA [WZT14]. Answer [WC17]. Anticipated [LR13b, TH14, WWR12, THC17, Xu12a, Yu19, ZX19]. Anticipative [WS17]. any [BS15f]. APF [DKT19]. Apollonian [ZM16a]. Application [Alf13, ET19, Kak16, Nas12, Sab18, Abd11, BLWZ11, BRA^{+}12a, BOK17, CC16, Fan17, KOR15b, LBM14, LSS13b, LA13, Oga18, Ost18, Pin12, Pin14, Res11, SKJ15, TY15, Tsi12, Wan14b, WWH17, Xu12a, ZM16a, Zho15]. Applications [Mil19b, AM17, AY17, BE17, BK13, BK10, BK17, BRBB14, BRBB23, Bor16, BM18, CQT12, CS16c, CM17, DD11, Duc04, Duc10, FL19, FNP18, Fer18, FMPV12, GM14a, HT14, IA10, JHF15, KK11, Kre18, Kum15, LVY15, Lar15, LV11b, LZW11, LYW11, Li13b, Lin19a, Lon13, LRH15, Ma15, Mac11, MP15a, MU16, MA14b, MMM13, OQ10, Pam18, PLH13, PSS16, Pin15a, QSS17, RS12a, SA13, Tau15, TYNZ15, VS13, WJY11, WH14, WWD15, XF17, YWY10, ZHL13, ZWB17]. applied [EJ11, JDK^{+}11, dML12]. Applying [Mar11a]. approach [AH13a, AR13, BMD13, Bob13, BCMR13, CTN17, CCS16, Che14d, CJN19, Cho16a, CGH11, DWWG18, Di 11, FCU11, GM17b, GDNM18, HP18a, HH13a, HS11b, HP15, LML15, LCJ10, LP10a, LL14, LX10a, MZZ15, Nag18, NV11, NSS14, QX13, SSN16, Sch14, SR12a, TL12, WW10, WL11, XZ13, ZCM15, ZY11, vDLMW19]. Approaches [VB17, Sha14]. appropriate [BK18]. Approximate [LR12b, MM15, ABC10, CM10, Gov14, NTT18, PR15, VRR13]. approximated [PPN10]. Approximating [NZ19, CSL16]. Approximation [LMR12, PZ14, WY10, XTT18, Ana17, BB14b, BH10a, BH10b, BS19c, BdS13, Che10b, CSY12, Dal19, DFS10, DS12, EK10, GX15, GM18a, Had17, Hwa18, Jou14, JS11, LKK12, Lem11b, Lem11a, LL10, Lin19a, LM15a, Mao15a, Mül19, NT17, NL19, NLF11, PTW10, PS17, RS16, Sas13a, SG15a, SGV15, UK18, Yao14, Zha16]. Approximations [FL10b, SWZ19, BD10, GFA10, GS13a, Gov15, Lee19, LL17c, LBH11, MW16, Mih12, MT17a, MO14, Mna11, OV15, Pin14, SWL10, UV13, WX18, Xu13, Xu15]. arbitrage [HK19, Ost13]. Arbitrary [Hwa13, BSK19, JHW18, Jou14]. ARCH [Duc10, GML10, GLML12, AAEH11, AT15a, BTT18, CW11, Duc04, MS10b, Shi14]. archimax [Wys13]. Archimedean [DM11, DJM11, Mai18, Res11, Rez15, Wan12b, Wys12, XLY17]. arcsine [JHF15]. arctangent [Pap16a]. area [FCU11, Liu13, MW14a]. areas [Bar12]. arising [DER15, FLRS13, JBS10]. arithmetic [Had19]. ARMA [HL11, LLH11, McE17]. arrangement [LL17b, PYK15]. array [HPW15, QN15, YCL14]. array-based [YCL14]. arrays [AKU17, BS19a, CYH16, HW13a, KP15b, QHS13, QG13, Ton16, Ton17, VA16, WL17, ZJL18, dML12]. arrival [CX18, GSW11, LW14a, SZ12]. arrivals [FK10, LR13a]. article [Sug19]. Artifactual [CS16a]. aspects [NS12, Sza19]. Assessing [Haz11, HPP19, LCKK14, Bos14, GM10b]. asset [FS12, KT10, LPS12]. assets [LY18b]. assisted [QCZ15]. associated [BVP10, CDL11, CS14b, DR02, DR15, GD15, GD18, HT15, KV16, MS11, RS16, SK16, TXWW18]. Association [Che13b, PNBW15, CHN14, GSST12, WS10a, YZ13]. associative [HLV15]. assuming [Bow16]. Assumptions [QB14, GSW11, HS10, Maa19, Roz14, Zho10]. Asymmetric [CPH12, QO16, Arv19, CH14, CP17c, Els16, Hay12, HJM19, HS19, Hür13, HBPC10, Hwa18, Lee12b, MM13c, MM14c, RH16, Sri15, ZJL18]. Asymmetry [dSGPM12]. Asymptotic [BS12a, BE17, BM17, FGA11, FW17, Gir16, Gna12, GF13, HCW13, Hew17, HNGS15, HLCH19, Isl16, Jas16, Jou14, KV17b, Kou12b, KW14b, LL19a, LC10, LW14a, LPN13, LW11, LS15, LX12, Lin17b, LBH11, Luo12, MBS13, MP15b, Nad16, PR17, Rah17, RS12c, Sep18, Son12a, Sun17, TXWW18, Uem16, Uno13, WMZW13, Wan14a, WN13a, Yab17, YS12, YYS16, YWZ14, ZL14, AB17, ANV13, AvHZ17, BS11, Bel16a, BS15d, BMN15, BL10c, BMS10, Bou17, CK16a, CFS12, CL19a, CF13, Che18b, CCGPW17, DFH18, EB17, GD15, Gee15, GR10, Han12b, HWB10, HM10b, HB11, Ilm13, Jar13, JLR19, JS11, KS10, KM11, KLW11, KA10b, Kre17, LP19a, LYCK11, Li18, LZ16b, Maa19, MO16, Nad15a, PL19, Rob10, RSC10, RS15a, SWL10, Sko19, Sko20, Tau15, Tud14, VM18, Wit18, WC18, Xia19a]. asymptotic [Xin12, Yan14, Yan15b, Yan17a, YL19, vLT11, Gre12]. Asymptotically [dWGM12, GL13, Gho14, GGV14, Ko13, Li13b, MK16, TXWW18]. Asymptotics [AJ16, GS10, IL11, LS12a, LPH13, PS11b, SK17, Wan19, WQY12, YZ16, YW10, ZZ15b, AD12, ADM15, DMPX17, DK11, For11, FN17, GL13, HK14, KL11, KD16, Kur19, MB15, MB16, Nak17, PL15b, RW16a, Tak12, VB17, WL14b, XY12, Zha14a, Zha17b, ZL16a]. atmospheric [FFM18]. Atomic [ZZ12b, PS18c]. attained [ADM15, Nav14]. attracting [LTX12]. attraction [BR17b, GGS12, Mar12b, NW13, TK10, Vas18, YHW11]. attribute [ST14]. attributes [Cai17]. augmentation [Han12a, KN15, Roy12, Tan15a]. auto [BW15, Rul14, dSGPM12]. auto-cross [BW15]. auto-logistic [Rul14]. auto-regressive [dSGPM12]. autocorrelated [CC10b, Sma14]. autocorrelation [BKR18, FS17a]. autocorrelations [McE12]. autocovariances [McE17]. Automatic [TXX16]. autonomous [HB18]. autoregression [ABD18, AD19, Cha10, Hwa19, LYB13]. autoregressions [DFT12, DFT13]. autoregressive [AB12a, AM14, BB18, Che18d, DPZ16, DR17, EP17, FA14, HS15, KS15a, KM19, KP10, LLHW10, Liu12a, Mas18, NR12, NMUJ16, PO18, RZ10, Wu14, Yeh11, Zha11a, ZZLL17]. auxiliary [AO12, TL12]. availability [HM13, Mat14]. average [AG13, Chi12a, CM11c, DH19, DHL14, DMS19, Dur13, HT13a, LQC15, LMLW15, McE12, MIL16, Pen11, SZW12, STD12, TE12a, Wu14, ZJZ16, Zho10]. averaged [KT10, NM18]. averages [Hom12, RS15a, SR12a]. Averaging [Rad19, Mes10, ZDM17]. avoiding [ÖÇ13]. axially [HZR12]. Azuma [RSdB15]. Azzalini [ML16].

B [BPXS19a]. Backbone [KR12]. Backward [Owo15, AM13a, BBM19, De 11, FRZ10, GLM^{+}13, HLW10, JY15b, KFHS11, LZ10, LL12a, LL12b, LR13b, LRH15, MZ19, MSS18, QX13, TJS13, TH14, THC17, WS17, WWR12, Yan19b, Yu19, ZX19, ZR12]. Bag [RP19]. Bahadur [AFJ11]. balanced [AWL17, EQ15b, KM12b, RMPB17, SD10b, SM13, ZCM15, ZRN14, ZRN17]. Balancing [Yat15, GXH17]. ball [Kob16]. Ballot [Len11]. Banach [ÁLBRM16, CQT12, DT10, LL13, Osȩ12c, Osȩ14a, STD12, ST13b, ZZ12b, ZH12b]. Banach-space-valued [ZH12b]. band [MV15, YXL14, ZZ15d]. band-limited [MV15]. bands [HEM10]. Bandwidth [JdU16, BD13b, BES18, DFH18, EL18, LLSW11, MMF17, OKT13, Oua13]. bandwidths [DG13, HP12b]. bankruptcy [CR16]. barrier [Abu13, Elh14, LL12b, LS16b, SYZ18, Xu12a]. Bartlett [Bar11, CO14, Can16, Kak12]. Bartlett-type [Kak12]. based [AMB10, AT15a, ABI19, BS15a, BB12, BMP10a, BKR18, BES18, CF18, CN16, Cha19b, CWI10, CD18b, CJN19, CCGPW17, DSW13, DZ17, DS18b, FP11b, FA12, FZW12, GO12, GD18, Gau14, GB16, GXH17, GMM10, Gut14, HP14, Hat12, He18, HLV15, HRB19, HS10, HZJ^{+}10, IV19, Ilm13, IKAH17, JP12a, JKP19, JLJ16, KMJ10, KBM15, KS15c, KS19, KS17b, LBG11, LP10a, LW13, LLO14, Li15c, LCLQ16, LW16, LLK^{+}18, LHT15, Lin17b, LLH15, LZL^{+}19, Lop15, Lu16a, Lu16b, LD17b, LG13, LPH13, MJH18, MGA11, MM18, MS12a, MO16, MM13c, MM14c, NV11, NSH17, NMS15, Now16, PB16, PLY18, Par14, PPA16, QLH^{+}16, ROSL17, RS14a, RSdB15, RNB17, RW16b, RG18, SA12a, SNS10, SWW11, SK15, sS10a, SZ12, Sri15, Ste11, Sub12, SJ14, SS12b, SSN13, TE17, TSL18, TE14, VR12, VRR13, Wan13c]. based [WZT14, WYC15, WQD15, Xia17, XKBG15, Xu12b, XZZ19, YCL14, YHMM15, ZP16, ZAV12, ZJS13, Zha19a, ZHWH12, Zhu13b, ZB15, vL16, vSK15]. basic [Len11]. basket [HJS11]. batch [CX18]. Batún [JB13a]. Batún-Cutz [JB13a]. Baum [CS14a, GS11]. Bayes [BM15, HP18b, HWM15, KWL15, KHHD16, MS16a, Ngu16, Rau13, SB19, WW10, WLLZ11, ZS19, ZRAN13, ZRN14, ZRN17]. Bayesian [AT14a, Bal11, BW19, CKP17, CCS16, DD10, DMC17, DN14b, DD14, EPSU16, FMA16, FXT12, GL19, GBR12, Gou15, HH13a, HS12a, JH14, KN11, KD14, LD16a, Li15a, MB15, MPA12, MW14b, Oh14, OPS16, PB15b, PS11a, RBSB16, SM19, SD11, Sri15, Tan15a, TA14, VRR13, VB11, Wal16, WL11, WS12, WR18, WN13b, Won13, XZ13, YJLL16, YW17, ZJS13, ZW18b, vDLMW19]. BDSDEs [ZT19]. be [BT16, Hei18, Nav14, Yor14]. beats [FA18]. behave [ATV10]. behaved [Tyk11]. behavior [ANV13, Aya13, BS12a, BLL10, BLKL13, BE17, BMS10, CS16a, Coe15, Ery11a, FF14, GD15, HH16, Jas19, JLR19, KM11, KLW11, Kou12b, LP19a, LLP11, LYCK11, LW14a, Mat14, RS15a, Roz10, Son12a, Tud14, WWW14, WWZ18, WN13a, YYS16, Yan17a, dlCJB19]. behaviors [AEK19, FGA11, MWZ11, Yan14, ZC19]. behaviour [AB17, EM10, GB13, Leh15]. Behnken [PN14, ZYL11]. bell [JS15, GSST12, WW16]. bell-shape [JS15]. belonging [Vas18]. benchmark [BL10b]. Benford [BSO10, Wój13a]. benthic [JDK^{+}11]. Berkson [YGTT13]. Bernoulli [AMP16a, BMB16, HM15, IJB17, KPO13, Mod11, Mod14, Rob19, Sch13, WQY12, ZK10, ZZ15c]. Bernstein [BS19a, Bel16a, GM16, HdCC17, Lu15, Min17, Zhe11]. Berry [ES13, LWX11, MC15, WJY11, ZJ15]. Bessel [AA19, KMN10]. Best [KM19, Osȩ12a]. beta [BB11b, BV11, CL11, Duo15, HV16, HJM19, Hom12, MBS13, McK17, Mu15, OT15, PPN10]. bets [ED17]. better [Cao18, LMM13]. between [AJ14, AH13c, BM15, BKR18, BVP10, Chi10, Cho18, Dav17, De 16a, Fen14, Fer11, HLR15, HY18, HWM15, JN10, Jon12, LW14a, Lin17a, LM18, Mak11, MW14b, NW10a, NS10, Oga17, Pin17, SA12a, Sch14, Tri18, YH16, YY13, You18, Yu11, Zin13]. bewildered [Won18]. Beyond [EFP19]. bi [BCNM15, LMLW15, ND10, YCQY16]. bi-degree [YCQY16]. bi-dimensional [BCNM15]. bi-exponential [ND10]. bi-infinite [LMLW15]. Bias [BBdWG16, GT14, JDK^{+}11, LLZ13, OKT13, WS16, Yat16b, ACD12, CFS12, GGR11, GI14, HS19, IL11, JD17, KZ14, KZ10, LD16b, MNWA14, MMF17, Sak17, Wes13, WN11, Yao12b]. Bias-corrected [JDK^{+}11, Wes13]. Bias-correction [GT14]. bias-reduced [KZ14]. biased [FGA11, GB19b, Jon19, LZ14b, MZZ15, SCZ15, SMZ18, WCD19]. biasing [Tyu12]. BIC [WZ10]. bidimensional [Li18, ZW12]. bifurcating [PO18, TE12a, Zha11a, dSGPM12]. Big [Azz18, BK18, Bow18, CKK18, Wit18, Won18, BBM18, BvdG18, Cao18, CG18, Cer18, Chu18, DH18, Dun18, FA18, GDNM18, GMDP18, Jam18, Leh15, Qua18, Rei18, Sco18, Sec18, Sha18, Shi18, SIBC18, TR18, San18]. Bihari [QX13]. bilateral [CFBD13, Liu12b]. bilinear [BG15a, BG15b, CPH12, FPZ13]. Bin [Sta16]. binary [AMP16b, Bak18, Bha13, CCM19, CC10b, Di 11, DSW13, ED17, KHN12, KA10a, KK19b, LSCL17, LW17b, Mak11, MBTC13, OPS16, ÖA15, ROSL17, Sha13a]. binary-data [Bak18]. Bingham [Sch16]. Binomial [RBSB16, Zha16, AJ16, BRA^{+}12a, BGM19, BMD13, BMD15, BSL13, BK13, BV11, BRBB14, BRBB23, CL10, DBB13, Doe18, GM14c, Gre11, HD13, LMPR19, PB16, PR16, Sre12, Thu14, UK18, YK11, Zha12]. binomial-type [Zha12]. bioassay [BL10b]. biological [DvH11]. biorthogonal [Che18a]. biostatistics [Wit18]. Birnbaum [FZB16, FLRS13, IK14, LC10, LCX12, TPNSN19, ZB15]. birth [BD17, But15, CSL16, CCSC11, DMS19, EHP13, JLY14, Mar11a, PTW10, WZ13a, Whi12]. birth-and-death [BD17, Whi12]. birth-death [CSL16, CCSC11, Mar11a, PTW10, WZ13a]. birthday [LS15]. Bismut [YY18]. bispectrum [IT14, PP14a]. Bivariate [SK13, BS10, BS14a, DM13, Dom13, EB17, FW17, FS11, GMS18, Han10, HS12a, HD13, HP14, HPP19, JBS10, KK11, KS17b, MU10, Mas11, MB14, MT17b, Nad13, Nad15b, Nad16, NSL11, OT15, sS15b, Sre12, TV17a, Van16, WLP17, WL17, YK11]. Black [DF14, GM16, SS17, Wan19]. Blackwell [Chr12, Had19]. blind [MNOT12, TMN16]. Block [HHF17, BBGH12, CLYW15, Din15a, DPP16, DS15, HM10b, KP17, LLSW11, WW10, WR18, WCZ18]. block-iterative [HM10b]. blocked [DS15, Jac11, JKD14, JKD15a, JKD15b, WCZ18]. blocking [Yan13]. blocks [Jac10, Jac13, JKD15a, JKD15b, PN14]. blowup [Dun19, GLM19]. BLUEs [Ren15, Ren14, TZ16]. blunting [Bak17]. BMO [TMS19]. Board [Ano17k, Ano17l, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano14l, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano15k, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l]. Board [Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l]. bonds [CJN19]. Bonferroni [PS11b]. Boolean [HS17]. boosting [Hay12]. Bootstrap [CP11, DSW13, NdBC12, WR10, BD13b, BD10, DPP16, HS12c, KS15b, KKP15, LSS15, LM15a, Llo10, RS12c, SW13, SKK10, Wan13a, ZP19, dML12]. Bootstrapping [Eck18, EIW14, Shi14, HS13b, SH13]. Borel [Dav13, Fro12, Ste14]. both [CYLL19, CPH12, LZY15, WHB10]. Bound [ASH18, BB11a, CKF17, CC16, DRZ19, Efr10, EQ14, GK12, Gor12, GM14c, HL16, Hil19, JMDW15, Jov14, LP17, McC13, Osȩ10b, PR16, Poi19, SC15, WJY11, Wis11, Yan12, ZWHQ15, Zhe17]. boundaries [AS15]. Boundary [BG16, Den17, AM13a, FKZ15, GV16, KZ10, MN17, Öns13, PKK14, Pen14, Zha10, ZZZ18]. bounded [AA13a, ATV10, Bri15, Deb12, EAB11, FL10b, HS11a, KP15c, Sas13b]. bounding [LP11, YAT16a]. Bounds [Ana17, Chi12a, DJ13, DM17a, Fro12, Fro15, LV11a, Lee12a, Lot19, Mar14b, VAB15, AG13, Bai12, Bai13, Ben18, BMNZ18, Bou16, CLQ12, Chi12b, CSL16, ES13, GAB16, HCW13, Jan18, JW17, KM13c, KSW15, LWX11, Lu11, Lu12, Mak10b, NW10a, Nav14, Puc13, Röl18, RS15b, Sas13a, Sas13b, Sau12, SD16, Ser18, Tyu12, UV13, VA16, Xia19a, ZK10, ZK14, ZK15, ZKSK18, ZJ15]. Box [PN14, ZYL11]. boxes [IJB17]. boxplot [BVV14]. boy [Pol13]. boy-or-girl [Pol13]. brain [Chu18]. branching [BBHH10, BP19c, CYLL19, Chu14, FL19, GW15, GZ15, GW16, Gao19, GZ14, HA15, HMZ11, HB10, Joh14, KT19, KD12, KR12, Li11b, LM17, Ma14a, Ma15, MSPSJ17, ÖÇ13, Öz16, Rah11, Rah17, Shi19b, Sun13a, SZ18, WGL11, WL13, WLLL17, WK10b, Xu14, YR11]. breakdown [RV19b, SÖV14]. breaking [Mil19a]. breaks [HS15]. breeding [BBHH10]. Bridge [WSW10, BD11, CPY15, Fer18, GT14, ZZ13b, ZLC18]. bridges [AS15]. Brillinger [Hei18]. Brillinger-mixing [Hei18]. Brown [EKS11]. Browne [ML16]. Brownian [DF11, ALL12, ANRW15, AS15, AEK19, ATV10, BDP14, BTT18, BL17, BRO14a, BBHH10, BD11, Bon12, BMNZ18, BH12, BH17c, Bry14, BW18, CVA13, CY11, CG13, CPY15, CR13, CP17b, DOT14, DL10, Dav12b, DO11a, Dun06, Fer18, FKZ15, FJW15, FM13, GM14a, Gas10, Gat13, GT14, GS12, Gzy19, HT14, HT13b, HS12b, Hir11, Hob13, HB18, JB13b, JV10, KP15a, KKNH18, KS16b, KS17c, LE19, LX18, LYCK11, Li17a, LY18a, LN19, LW11, Liu13, MP16a, Mar11a, MX17b, MS13, Met10, Muk17, OS15, OSW18, ÖÇ13, Öz16, PH13a, Pot15, RH11, SS18, SC12b, Shi19c, SV17, Sun13a, SG18, VY12, VAB15, Wan14d, WYY14, WX19a, WSMG18, WS17, XTT18, XL18, Yam15, YS10, YR11, YY13, Yu19, ZZ13b, ZLC18]. Brunick [For14b]. Brunk [ST13b]. BSDE [FL10a, Shi19a]. BSDEs [AEFO18, FJ10, FJ12, Fan16, Izu13, JB13b, LS16b, LY17, Rév12, TJD10, XF17, Xu12a, ZF13, ZL15]. budget [Bel16b]. Burgers [NTT18, WWZ18]. burn [MIL16]. burn-in [MIL16]. Burr [KA10b]. business [Jam18]. busy [BML14].

C [LW13]. C-MGARCH [LW13]. CAANA [Ko18]. Cahn [JWW14]. calculate [SLC16]. calculating [LCW17]. calculation [HHY11, MN17]. calculations [VM12]. calibration [Bel13, Ohy13]. call [Sec18]. Can [Nav14, Cao18, Roz16]. cancer [CWI10]. Canonical [HP12a, Bog15, BV15b, Pin12, Suz18, Suz20, ZP12, ZX17a]. Cantelli [Che18d, Dav13, Fro12, Ste14]. Cantelli-like [Dav13]. Cantor [Arn11, BS12b, Hu15]. capacities [Ter15]. capacity [HLV15, MZ19, RIK16]. capital [DZ19b, XZZ19]. capture [LZL^{+}19]. Capturing [LMH14]. Caputo [ADH19]. cardinal [FK10]. Carlo [BBCD^{+}18, DRZ19, Nie16]. CARMA [BL15]. cascade [LLP19]. case [ASN16, BS10, BDvdAW16, BCM14, BCNM15, Che13b, Coh10, DFL12, DFL13, GS16, GM18b, Had19, HJM19, IKM16, KW14a, LMW15, LLP11, OV15, PRS15, Ruk12, SWH^{+}11, Vet14, WL11]. case-control [GS16]. cases [Oga17]. catastrophes [LLY19]. categorical [BM10c, JH11]. Cauchy [ÇKN17, Fak17, GZSZ19, Izu13, KRYAV16, KAVRY17, ST13a, Wal17].

Causal [AHM19, BL14b, Chi12a, Chi12b, Gho12, Giu15, KJS15, Lu16a, LD17b, Lu18]. Causality [PD12, RT15, BLWZ11, PD11, Tri18, VP12]. causation [CCFL18]. cause [ED14]. causes [SA12b]. caution [SYZ18]. Cayley [ABD14, DYB11, SY10]. CDF [BCI14]. CDS [DYW14]. censored [AEAN18, ADP12, AFJ11, ABI19, BD13a, BKL19, BBdWG16, Che19b, Che12b, Fak10, HTA10, IV19, JQZ12, KLM11, KM12a, KBM15, LZ14b, MZZ15, MKK17, MASR14, Mes10, MN11, MN13, PP15, PNC15, SK15, sS15a, SCZ15, SMZ18, TZW12, TZW13, TSZ15, WL11, WSU11, WCD19, XSZ17, ZBZ11, ZJL13, cZzXgL17]. censoring [ACD12, BK19a, BG11, Bou15, BPS17, DL13a, Din10, FXT12, HS12a, LL17b, LW15a, PN12, QB14, QCZ15, Rob10, sS10b, SB10, Wan12b, ZZ11, ZZM15, vBC19]. censorship [AM19b, SY19a]. census [GG13]. Centered [BCR10, CW19, CLQ12, EQ14, Ruz14]. centers [HLR15]. centers-coupled-with-radii [HLR15]. Central [Arb19, CW15, DR02, DR15, McK15, Rok15, WC14, Ane17, BPW18, CK16b, Eks14, Fro15, Gao19, Kam17, Kat19, Mül19, PTT19, RS16, Röl18, Ton16, Ton17, WWS19, XW17, BHJ12, BBH14, Bob13, EM16, MW11, MS11, NW10b, Sha13b, Ter15, Tyu12, VY12, VW14, WGL11]. centred [FK12]. centric [LAG^{+}18, MYA15]. certain [Aza13, CK16b, CC16, HS10, Jon12, LZ10, LPW10, MA14b, Sko19, Sko20]. Cesáro [Tru13]. chain [AR14, BDER14, Dal19, DN14a, FMPV12, Gei17, KL11, LP19a, Liu12a, Sta12, WC18, Xu13]. Chains [LP11, BS14b, CH13, Cho18, DM14, DYB11, ELS13, FC10, Fra15, GFH16, GAS13, JT18, Jia16, KS12, Kau14, KR19, Kov10, Li17c, LMLW15, LPS15, LR13b, MZ17, Mar11a, Mia14, SY10, SHK14, Ste14, Ter17, ZK14, ZK15, ZKSK18, ZW15a, Zho13b]. challenges [Azz18, Chu18, PYS^{+}18]. chance [Bos14]. Change [TL19, CTN17, CTW10, Ciu11, Ery18, Fit14, Huh12, JP11, KS15a, KN15, LQZ13, LZL13, Mar16, Mih12, Roh16, She11, She13, SD15, WK11a, ZG16, ZWT15]. Change-point [TL19, Fit14, LQZ13, Mih12, Roh16, She13, SD15]. change-points [Ciu11]. changed [HS12b, Kob16, KNV11, MN14]. changes [Bre12b, Bre14b, Roh16, Tim15, WA17]. Changing [PP14b]. Chaos [Tud14, HRB19]. chaoses [Gau19, KL15, Mel16]. chaoticity [WC14]. Characterising [LG16, WW18]. Characteristic [Pin15a, Vri16, BH15, BK10, MU16, Nak13a, Shu16, Su10, Tra14]. Characteristics [INO10]. Characterization [Bis10, CKO19, GM10a, MO16, Mse18, Myr19, Nan10, PF17, TE12b, TE14, AN10, BLB16, BM13b, BMS16, DW18, Ejs16, Kru10, KT11, LV13, LS16a, LA13, Pos10, Sre12, SG15b, Ush11a, Wan11b, Wes15, YC16, Yeh11, ZCM15]. Characterizations [FA12, HS10, KI15, NSL11, SI16, AV13a, BR17b, Osȩ12b]. characterize [BNM13a]. Characterizing [QW16, Ber13a, KM13d]. charged [Wan13b]. Charlier [FQZ18, WN14c]. Charlier-like [FQZ18]. chart [KS17b, NMS15]. Chebyshev [Bud14, Nav14, Rao10, Sza19, Yas13, ZH12b]. Checking [WXL19, Duc04, Duc10, EJ11]. checks [DSW13]. Chen [Lê16, SK13].

Chernoff [CD12]. Chesneau [KL16]. Chi [SKRT16, JLR19, RW16a, WWD15]. Chi-square [SKRT16, JLR19, RW16a]. chi-squared [WWD15]. Chinese [Mil19a]. choice [CDS17, MD16, Sin19, WK10a]. Cholesky [HX14, Mad15, PW15]. Choosing [MNP16]. Choquet [Aga15, Alf15]. Chover [hCyP15, LZ16a, Wan14c]. Chover-type [hCyP15, Wan14c]. Chung [FD10, Hil19, Liu13, LX10b]. Chunked [NM18]. Chunked-and-averaged [NM18]. Cifarelli [You18]. CIR [Alf13, MY14, PS18b, XW16, Yan17b]. circle [Bél11, HZR16, IA10, Sch14, TS17]. Circulant [BGHS11, AS18, ACMM13, BSCQ17, BSCQ18]. circular [BJD16]. circularly [CP17b]. circularly-symmetric [CP17b]. circumscribing [Xu15]. CKLS [CW15]. claim [BS12a, BNB16]. claims [BS11, BZ13, FN17, GL13, JCM15, Lu11, Lu12, LZ16b, SXM16, YW10, YL19]. class [BS14a, BJW17, BHJ18, CKO19, CvGL18, CZ17b, Di 11, DZ18, DCR19, Ery16, FL10a, FS17b, GZ15, Gau14, Gra11, IL17, JLR19, JWW14, KM12b, Kob16, KS17c, Li11a, LR11, LFM14, MP15a, MM18, Mel16, MX18, NC16, ÖA15, Pro13, RM11, Sep18, sS10a, ST13a, SdOG12, WC15, XY17, YWY10, ZZ15b, ZM16a]. classes [BS11, Car10, CYW13, Gna12, Nav18, OT19, Rob13, SD16]. classical [KM13a, Lin17a, LXZ15, LZJ17, MB16, OS18, WYL10, YGTT13]. Classification [GBR12, MS18b, DDZ15, LSCL17, MMZ17, YJLL16]. classified [Ski15]. classifier [CCM19, MK16, ZS19]. classifiers [Rau13]. Clayton [SS14]. Clayton-family [SS14]. clear [CLWH17, ZG17b]. climate [CG18, She11]. clinical [BB11c, CWI10]. clinical-genomic [CWI10]. cliques [HLV15]. close [Llo10, NÚF12]. Closed [LW17b, Shu18a]. Closed-form [LW17b]. closeness [AB10, BD13a, BB12, JDB12, MS16b, MS16a, RA13b, RA13a]. Closure [DSS18, YLS14, Bob10, LH14a, LW12, YWY10]. CLT [AL16, CZLT18, KP15a, Spa10a]. cluster [ASN16, AWL17, BN10, BDB^{+}10, Jal16, Sha13a]. clustered [EBG16, HBL11, TLNO11]. clustering [FP11a, Fis11, GMM10, Lop15, MM13c, MM14c]. clustering-based [GMM10]. clusters [CD18a, LTP18]. coboundary [Mor18]. codes [Deb12]. Coefficient [LDLC18, BWW11, CSS14, FA14, FP13, HZJ^{+}10, JQZ12, LML15, LZLW16, LHT15, LYC14, MH12, Mas11, MMC14, NP16, PS15, PRS15, Poi19, SZW12, SS15c, TSL18, TLF12, TXX16, Wan13c, WS13, WZ17, YL10, YGL14, ZW18b, ZZLH15, cZzXgL17]. coefficient-based [TSL18]. coefficients [AEFO18, AM14, AA13b, BIK12, FJ10, FF18, GHH^{+}10, HLW10, LL12b, LT19a, Lon13, Mat12, Nak13b, NT17, NL19, Pam18, PP18b, SG15a, Shi19a, Shu19b, Tah14, WWR12, Xia19c, XWML19, ZK10, ZT19, vWvZ17]. Coherent [WH14, AB12b, BS18c, Ery11a, KHN16, Nav18, PB13, ST19a, Yan14, Yan15a, Yan17a, ZFZ13]. cohort [LTP18, LZYW19]. coin [BB11b, Li13d]. coin-tossing [Li13d]. coincident [MNP16]. Cointegrated [SB19]. cointegrating [Can16, SH13]. cointegration [dFAM19]. collaboration [SWH^{+}11]. collaborative [WRvdL11]. Collapsing [DM14]. Collatz [DP16]. collection [DP19b]. collision [WL13, YS10]. color [BT14b, SM13]. colored [Li17b, XY17]. colors [KM12b]. coloured [MV10]. columns [CKF17]. COM [BRBB23, BRBB14, Pog16]. COM-Poisson [Pog16]. combination [AP13, JBS10, MvdBV^{+}11, QN15]. combinations [GSK12, JP16]. combinatorial [Fro15]. combined [EQ15a, Els16, MK16]. Combining [DN14a, IL11, LLH15, Che13b]. coming [Mir14]. Comment [For18, ML16]. comments [Lau10]. common [BB12, BMS10, BdS13, CS17, GMS18, KZ17, KS19, LLH15, MKK17, MNP16, PS15, Pol13, Roz16, Ruk12]. communication [RR11]. commutative [Qiu14]. commute [HL19]. Comonotonicity [MD15]. compact [AMG17, CQT12, HY13, Myr19, Pei15]. compactly [TNP17]. Compactness [Kha19, Kru17, LVMS14]. company [LH14b]. Comparing [LP10b, PH13b, JP12a, MGSL11, SNS10, SGG10]. Comparison [BPRS13, Had13, NK11, PTW10, PH10a, ADP12, BKR18, CRG17, De 11, KD14, LY17, LW14b, LW15c, MM13b, Oh14, Ruk16, SM19, THC17, Wan15, YS12, YD12]. Comparisons [DZZ13, PSS12, CSK13, DK14, FZ13, FZ15, FZB16, FD18, MM12, MA14b, PYK15, RIK16, RM13, Roy12, ST19a, VAZB10, Yat15, ZB11]. Compatibility [Che10a, DR17]. Compatible [WWW19]. compensated [WZ19]. compensator [Seo17]. competing [BG11, ED14, Gho12, HH18, PS11a, SNS10]. competing-risk [PS11a]. competition [Ma14a, Ma15, Pac18]. Competitive [GHR16, GHR17]. complement [BWW11, SKRT16]. Complementary [HA12, Jon18]. Complete [Che18b, MKK17, Nad15a, QC14, WWH17, Zho10, GZ13, KM12a, Ko13, Ko18, Liu11, Mao15b, PP18a, STD12]. completely [Gri19, PPM18]. completeness [FS16]. completion [CLEA18, CLEAMS16]. complex [BK10, BH11, HYWT17, Ilm13, LNI16, MVI11, WN10c]. complier [AG13, Chi12a]. component [BB14b, BS15f, BCR10, BZV11, CM17, CP19, HLR15, HL13a, Ilm13, KHN16, LZ17a, MPA12, Mar14a, MNO15, MM12, PR17, Rév12, TKO12, WXL19, WQD15]. component-amount [MPA12]. components [AB12b, BB14a, BK17, BMS10, DZZ13, Ery11a, FZ13, FZ15, FZB16, Li15c, Lia12b, MKJ15, MM13b, ST19a, Smu19, Wan15, WN10c, ZB11, Nos19]. componentwise [RMÁL19, ZP16]. Composite [MZZ15, GB16, GB19a, JQZ12, NT14, PSS15, TZW12, TZT16, WMZW13]. composition [Bre14b]. compositional [BJW17]. compositions [Jur18]. Compound [HD13, GX15, GW12a, Han10, HP18b, Hui10, KZZ13, KA10b, MST10, Nie19, PL18, PZ17, Rat15, SLZ13, UV13, Yan17a, YY11b]. compressibility [Dav12a]. comprising [MM13b, Wan15]. compromise [ZG17b]. computable [HL16]. Computation [McE17, ZWLB17, CL16, HM13, Nie16, PJ13, RW16b, SM19]. computational [Qua18]. computations [PP11]. computer [Cer18]. Computing [CGLN17, HP18b, MM13a, Mar14a, BGM19]. concave [Dev12, DZF18, KMS15, LM18, SD10a]. concavity [FD18, Haz11, Mu15].

Concentration [ACP19, GGR11, GI14, BPRS13, Cha10, CRB19, DZ11, Din14, LPH13, Mau10, Ser17]. concept [KI15, WLLL17]. concepts [BS18c, NV11]. concerning [Di 11, JDB12]. concise [WZ15b]. concomitant [ZV15]. Concomitants [SBN16, KS17b, WN10a]. concordance [PSS12]. Condition [EF16, AM13a, Arb19, DM13, Gei17, Hu16, KS12, KMJ10, Osȩ13a, Owo15, QX13, Rul14, Tah14, TH14, WWW10]. condition-based [KMJ10]. Conditional [CZ17b, JB19, PS16b, Pap16b, PB13, PS17, sS15a, SdOG12, YIS15, ART14, AH13a, AM19b, ALS11, CCFL18, CvEZ10, CZCL19, CW16, DNKL12, De 16a, DR17, GX15, GGO15, GW12c, Had13, HX14, HCT16, IA10, JP17, Khm17, KKKP17, Li19, LX12, Lop10, MJH18, MdJ19, Mil19b, Nog13, Oes15, Osȩ12a, Par14, PY17, RNB17, RR13, ST19a, SB16, Son12b, SHL15, Sza15, WW13b, Yab17, YQW10, YH15, ZFZ13, ZPZ18]. conditional-sum-of-squares [Yab17]. Conditionally [YZ13, BS12a, Che10a, Had13, JGW14, Osȩ10c, Sas13b]. conditioned [FL19, Mou19, Pap16b]. conditions [AD19, AH13b, ASVY14, BSO10, Din10, DFT12, DFL12, DFL13, DFT13, GZ13, KM17, LTvdVR11, PL18, RS10a, Sma14, ST10, SG18, Sun13b, Tab18, THC17]. conductance [Bou18]. Conducting [Men18]. cones [CG17]. Confidence [BK19a, FK12, Ruk16, TL15, BCI14, BB16, BFS16, BKL19, GEC17, HMT13, HEM10, Kab11, KM18, KM13a, KS15b, KKMSA13, LSS15, LLH15, NW10a, NdBC12, RH15, SS13b, Sub12, SJ14, Thu14, VM18, Wan12a, Wan13a, WZG16, YXL14, Zha14b, ZZ15d, ZWLB17]. confirmation [NS14]. conflict [EJ11, RS12b]. conformable [ÇKN17]. conformal [AIM19]. confounded [DM12b]. confounding [BDB^{+}10, LZZ15b]. congruence [Gna12, Kin10]. conjecture [BM11, Che16, Shu16]. conjugate [LA13]. conjunctions [DHJT14]. Connecting [CZ15]. connection [Lin17a, Sto12, WT18]. connections [TZ16, VSP13]. connectivity [BS12b]. connectome [PYS^{+}18]. consequences [ACP19, Len11]. conservative [FP11b]. considering [XD15]. Consistency [ÁLBRM16, BB14b, BG15a, BG15b, DL13b, DMPX17, IK10b, IR11, Kre17, SD10a, Smu19, VD18, BSK19, BL10c, BES18, CRC10, CW11, CGH11, Fak10, Gre12, HS12c, KWL15, KZ17, KBM15, KV13, Lia12a, Lu15, Luo10, LG15b, RMÁL19, SL13a, SX13, SV15, Vil12, Wal10, WF15, XJW15, YLHY19, You17]. Consistent [BF10, MHH11, RP19, SV10, ZF16, BH11, CP11, Chi10, DFKK12, GW12b, GW12c, Mis14, RR14, dMO18]. constant [ADH19, BL17, BLB16, DJ13, GL13, KP14, LGW12, PH10b, Pog16, TV17a, WCM11, YW10]. constants [Fer14, NZ19, Osȩ12a, ZJZ16]. Constrained [ASH18, BH16, Kin12, Leh15, LHT15, SW11, TSK13]. constraint [Bel16b, Ser18, WST14]. Constraints [Dav12a, AH11, BCI14, BB10, DSX13, Ern17, HWM15, MS16b, Oh14, ZW13]. Constructing [CQZ15, Els16, LB17, Wys12, ZG17b]. Construction [CLYW15, LLO14, LCLQ16, LDD15, Shi19c, WZYL17, KHN12, LMPR19, McC12, Rez15, ySDM14, TL15, Wan12c, XDWY18]. Constructions [OT15, RR13]. contact [Xue17, Xue18, Xue19]. containing [AMP16b, CLWH17, MPA15]. contamination [CFBD13]. contingency [JKP19, KPK15, PKK14, Zha19a]. contingent [CJN19]. continual [Azr12, Tia16]. continued [AA17, Fan15b]. Continuity [JV10, Alb18, ASVY14, BJQS16, Li17b, Nak13b, SG18]. Continuous [Gus12a, JS18, ATV10, BGT15, Bre12b, CL19a, CXZ15, CL15, CCSC11, DM17a, DL13b, EL18, FJ12, FL19, GFH16, GJ10, Hor16, HLW10, HMZ11, JY15a, KD12, KK19b, KC19, KR12, LL19a, Lee12b, LP11, Li17c, LT19a, Ma14a, Ma15, MNX13, Mes14, MSPSJ17, NKKY13, Neu13, NL19, Osȩ10b, PD12, PNBW15, ST13a, TL15, TJD10, Wei19, WK11a, Wój13a, Xu14, ZK14, ZK15, ZKSK18, dSF12]. continuous-scale [TL15]. continuous-state [FL19, KR12, Ma14a, Ma15, MSPSJ17, Xu14]. Continuous-time [JS18, Bre12b, CCSC11, GFH16, Li17c, Wei19, ZK14, ZK15, ZKSK18]. continuously [HS12b]. Continuum [SH17, CZ15, Yao13a]. contoured [Tsu10]. contract [Dur13]. Contraction [CL13, SKB14, SKB15]. contracts [Pac18]. Contrast [Pin17, ZZ11]. contrasts [CCS16, DM12b, NW10a]. contributions [MMZ17]. control [CSK13, Che13b, Döh14, GS16, GF13, HA14, KS17b, LH14b, MGSL11, NW10a, NMS15, QLH^{+}16, RM13, RBSB16, SGG10, SDNS16, ZYG17]. controlled [BCD19]. Controlling [Mak10a, WX12]. controls [MSS18, YHMM15]. converge [Wój13a]. Convergence [BM18, CMW18, Che13a, Che19a, CL15, Cou19, Gra11, Jov14, KP15a, KD16, MXZZ18, Mar12a, Mas11, Pet11, Sze16, Wan14d, dH15, Alf13, AG11, BB11a, BG13, BT14b, Bao12, BJW17, BP19b, Ben17, BK10, BV15a, BG11, Bor16, Bór18, Bou15, Buc15, BF17, Che10b, CWZ12, CH14, CS14b, CW19, CP13b, CD10b, DD12, DR11, DG13, ET19, FD10, Gao19, Gov14, GZ13, Gut14, He14, Hel19, HMS14, HM10b, IMM16, IKM16, KM12a, Ko13, Ko18, KK13b, KZZ13, Kri18, Li13c, LL17c, LPNW14, Lon13, Mar14b, MS12a, MT17a, Mou19, NT17, OS15, PP19, PD11, QY15, QC14, QHS13, QG13, RM11, SC15, STD12, Son14, SHL15, Tab18, Tru13, UU11, Vid18, WYY14, WMM15, Wan18, Woj13b, WWH17, XY12, XH12, YYY17, ZK15, ZKSK18, Zho10]. convergent [CLM12]. converges [DMS19]. converse [De 11, LY17]. convertible [CJN19]. convex [Bak15, BD15, Cal13, CQT12, Dav12b, DM13, GSK12, HS16b, KL15, MD15, Osȩ14a, QN15, WH14, Yan15a, Yan17a, Yan19b, ZL15]. Convexity [ST18, Alf15, BDK18]. Convolution [Vil12, BLS14, BS15f, FP17, Gna12, Muk10, PL15b, PF17, WYY13, YWY10]. convolutions [Cha17, SWW11, ZB10]. Conway [RBSB16]. coordinate [CMS19, INO10]. coordinate-wise [CMS19]. Copula [LW15a, AJ14, Bou17, BES18, BD10, BCNM15, FS11, HZW16, JT14, PPA16, RNB17, SK16, SA13, TMS11, Wan12b, Wys13, XLY17]. copula-based [PPA16, RNB17]. Copula-graphic [LW15a]. copulas [BDJ17, DM11, DFS16, DFS10, DJM11, DS12, FSNÚF11, JK13, JT14, MS10a, Mai18, Nad15b, NS10, PS18c, Res11, Rez15, Tru13, WLP17, Wys12, ZP16]. Cordeiro [Kak12]. correctability [CO14]. corrected [JDK^{+}11, Wes13, Yao12b]. Correction [DFT13, PS11a, Suz20, Can16, GT14, HS19, Lem11a, PNT15, Sak17, Sri15]. corrections [WS16]. corrector [ZZ13a]. correlated [BS11, CO13, HS16a, MBTC13, Met10, MMC14]. Correlation [Mau19, MN14, BL12, BH14, Cal13, CG17, DR11, DDZ15, DER15, FP13, GdL18, Jia13, KP17, LL12c, Mad15, Mas11, NdBC12, Olk14, PRS15, PW15, SBN16, Su10, WWW19, Wes13, Xia19c, YJLL16, ZK10, ZW18b, Zha19a, ZX17a]. correlations [Liu11, QO15]. corresponding [Arn11]. Corrigendum [Bai13, BSCQ18, DR15, DFL13, Duc10, DF11, Kak11a, Kim13, RA13b, Ren15, Ton17]. cosine [YSZ19]. cost [BBM19, DZ11, DZF18, LLP11, MIL16]. costs [BS14b, LZY15, LY18b, SS17, TLNO11]. count [CZ17b, RBSB16]. countable [DP19a]. counterexample [GF13, Gut14]. Counterexamples [Ter15]. counterpart [D’O11b]. counterparty [DYW14]. counting [BMB16, Bre12a, Bre12b, Bre14b, MP15a, Ste15]. Counts [AMP16a, Spa11, Yao14]. couple [Vid18]. coupled [EK10, HLR15, JY15b, WWS19, dlCJB19]. Coupling [CSL16, JK13, Ouy10]. Couplings [Cri19, GGR11, GI14, RR13]. CoVaR [BDJ17]. Covariance [AM14, EFP19, Jia13, Aga15, Alb18, AB17, BW15, Bao12, BM10a, CZLT18, CHM16, DTV16, HLM16, HZR12, IL17, JP17, Kin10, LP13, LCW17, Lo17, LK13, MD13, Mao16, NHBG19, PL19, PNT15, Pin14, Pos10, RSC10, Sch12, SR12b, Tah12, Wes13, Xie13, XH15, YH17, Zha17b, ZIPP17, vW17]. Covariate [FP13, GXH17, Lu16a, Che19b, Din10, MH12, SS11a, ZX18]. covariate-adjusted [ZX18]. Covariate-balancing-propensity-score-based [GXH17]. covariates [CGH11, DCR19, Fin11, HZJ^{+}10, Khm17, LNC12, LZL^{+}19, MS18b, NT14, SS15c, WZW17]. cover [HL19]. Coverage [HEM10, BM15, HS17, KM18]. covering [GMS18]. covolatility [DS14]. Cox [FS18, GK12, KZZ13, Ruz14, SS10c, TR12, WST14, YS19]. Cramér [ASH18, Onz11, Yan12, Ye16]. Cramér-Rao [Onz11]. crashes [BCM14]. Creating [Bak17, SWH^{+}11]. credibility [BM15]. credible [VRR13]. Credit [BH17a, Elh14, LWD13, WYL10, KOR15b]. crisis [SHBHD11]. criteria [Gre12, Kha19, PJ10, Pro13, Pru19, XL16]. criterion [CKF17, Ciu11, Döh14, HM10a, HP18b, HR11, JT14, LV11b, LZZ15b, LM17, Lia12a, QHS13, WZT14, ZDM17, ZYG17]. critical [EF16, GZ14, Li11b, Sco18, SZ18, WK10b, Xue18]. criticality [WLLL17, YR11]. criticism [MWGT17]. Cross [Ski15, BW15, CFS12, DM11, EFP19, IL17, JT14, KZ17, Kin10, MvdBV^{+}11, Nic17, Roh16, Ros13]. cross-aging [DM11]. Cross-classified [Ski15]. cross-covariance [IL17, Kin10]. cross-sectional [CFS12, KZ17, Roh16]. cross-validation [MvdBV^{+}11]. Crossing [Osu12, BG16, Den17, DM13, MN17]. crossings [Mat12]. crossover [Kab16]. Crucial [Sco18]. Csörgo [WX19a]. CUB [IP14a]. cubic [ASP11, HH13a]. Cumulants [WN14a, WN19b]. Cumulative [PP15, AEAN18, CN16, CCGPW17, ED14, FP17, KRB13, KS17a, KH18, KT11, LM18, PRS12, PK14, SNS10, SWZ19, TE17, XT11, ZYW15]. cure [Che19b, PB16, WZS17]. Curie [Mar14b]. current [AB10, AG18, KFHS11, RA13b, RA13a]. curve [BH15, FGA11, NdBC12]. curves [DNLR17, DP10, HA14, JM14, Pru19, Sak12, Sak15, Tap10]. cusp [Fuj10]. cusps [Bak17]. customer [SZ12]. customer-arrival [SZ12]. customers [KK18]. CUSUMSQ [HS15]. Cutz [JB13a]. cycles [WC18]. cyclic [WX18]. cyclical [HP11].

D [Hel19, SC15, Kha14]. DAGs [Nos19]. Damped [Rat13]. Dantzig [dC13]. Darroch [MPA12]. Data [BTT16, Fen14, TR18, ADP12, AFJ11, ABI19, Ane17, AD11, Azz18, Bak18, BD13a, BL12, BKP15, BOK17, BJW17, BBM18, BCL18, BMN15, BK18, BS18b, BS19c, Bow13, Bow18, BF10, BNSA13, Buc15, BvdG18, Cao18, CDG15, CG18, Cer18, CM10, CP11, Che10b, CM11b, CSG11, Che14b, Che19b, Cho16b, Chu18, CKK18, CZ17b, DMC17, DL13b, DSW13, DH18, DZX18, Dun18, ED14, EBG16, EJ11, FA18, GFA10, GLS13, Gho12, GB19a, GDNM18, Giu18, GdL18, GMDP18, HBL11, Han10, HD13, HP14, HLR15, Har16, Hay12, HX14, HHF17, HH18, HYWT17, IV19, IP14a, Jam18, JQZ12, JH11, Jon19, JDK^{+}11, Kau14, KLM11, KM12a, KZ17, KS17a, KBM15, KS15c, LLL13, LML15, LAG^{+}18, LCJ10, LZT13, LW16, LZ14b, LZYW19, LX12, LWL16, LZ13, LRL19, MZZ15, MKK17, Mao14, Mao15b, Mas11, MASR14]. data [McC12, Men18, MvdBV^{+}11, Mes10, MN11, MN13, MYA15, Nag18, NA13, Now16, OW18, PP15, PTW19, PT19, Qua18, Rei18, RBSB16, Roh16, RR10a, Roy12, Sak12, Sak15, SA12b, Sco18, Sec18, Sha13a, SW13, Sha18, sS10a, sS10b, SX13, sS15a, SCZ15, Shi18, SMZ18, SIBC18, SHBHD11, TZW13, Tan15a, TXX16, TLNO11, TMS11, Tsa13, Tyl10, Vie18, WHB10, WW10, WL11, WMZW13, Wan14a, WMFW15, WZ10, Whi12, Wit18, WSU11, Won18, WW16, WS10b, WCD19, XKBG15, XSZ17, YL10, YXL14, YJLL16, YJ12, You17, ZW18c, ZPZ18, ZS19, ZMG10, ZBZ11, ZJL13, ZG16, ZY11, ZGK13, cZzXgL17, dSGPM12, dB13, San18]. data-centric [LAG^{+}18]. data-dependent [CM10]. Data-driven [BTT16, Fen14, WW10]. Davis [KD16, LQC15, LG15a]. Dawid [LD16b]. death [AdUÁMM11, BD17, But15, CSL16, CCSC11, EHP13, Gou15, Mar11a, PTW10, WZ13a, Whi12, Wit18]. debate [Sec18]. debit [WYL10]. decay [Bou18, CK16a, GA13, WL13, ZYL14]. decaying [Tra14]. deceleration [BG19]. decent [HA15]. Decidable [Dav13]. decision [Kab13, RSV10]. decision-theoretic [Kab13]. decomposability [KP10]. Decomposing [Sha17]. Decomposition [MZ19, Bar11, GR11, HX14, KV17a, KR12, Lee19, Mad15, TV17b, YC16]. decompositions [ZZ12b]. Deconvolution [Ben16, Dat13, TNP17, ZGG16, Ben17, Ben18, PT19, SV10, Son10b, ZW18a]. decoupling [CD12]. decrease [JWCC19, Roz16]. decreasing [BS17b, Roz12]. Dedekind [Urb12]. Default [Elh14, BH17a, BMS16, HJS11, QY15]. deficiency [ZBZ11, ZJZ13]. deficit [YSZ19]. defined [BS19c, GMP19]. defining [WSU11]. Definitive [WA17]. deformations [EM16]. degenerate [QHS13, TZB13]. degenerated [WW18]. degradation [PLY18]. degree [LV11c, SQY18, WZ17, YCQY16]. degrees [GMS18, XKBG15]. delay [BH17c, GA12, HB18, LLK^{+}18, Liu14, MX17a, QLH^{+}16, RSR16, Tie13, WH11, ZYL14, ZX19]. delayed [CF18, DZD12, GKW19, WSMG18, YL19, ZR12]. delays [Che10c, CvGL18, CYS11, LF17, LD17a, ZDD18]. delimitation [Pro13]. demeaned [Ai16]. demimartingale [WHRY11]. demimartingales [CH12, Had11, Had13, Had17, HWYW12, WW13b]. demisubmartingales [Rao12]. Democrats [BH17b]. demonstration [Ger13, Ger18]. Demystifying [LD16b]. dense [LT14]. Densely [Gir16]. densities [Dev12, Efr10, HdCC17, Haz11, HL13a, JP12a, JSA12, Ker16, KMN10, Lan18, MS16b, OT19, PTV16, PH13b, RS12b, SW12, TNP17, XYZ16, YLS10]. Density [BD13b, PT19, ACD12, AM19a, AM19b, ALS11, BV10, Bha13, Bou15, BCNM15, CKP17, CM10, CP11, CP14, CDL11, Che18b, CW11, Che12b, Cou19, CP19, DFH18, DFKK12, DL13b, DG13, Efr14, EL18, Fak10, Gat13, GM14b, GSW11, GW12b, HP11, HP12a, HP12b, HS19, HP15, Jal16, JW17, JWCC19, Jou14, JS11, KLK12, LL19a, LYW11, Li15b, LX12, LM15a, LPS15, Lu15, MS12b, MS18a, MNWA14, MMF17, OKT13, Oua13, Sak17, SWW11, SD10a, SV10, Shu18b, Son10a, Son10b, SS11b, Sri15, Tah14, TS17, WW18, XXY12, YE10, YSZ19, YHY16, ZW18a, Zhe11, Zhu13b]. dependable [BLL14]. Dependence [Fer11, BZ13, Bra11a, DJ13, DFT12, DFL12, DFL13, DFT13, Duc04, Duc10, DJM11, EAB10, FF18, FS12, FS10, GMS18, GF13, GBGPR19, Has10, HPP19, HS12c, JN10, LPS12, LP19b, Mas11, MP16b, McE12, OK15, Roh16, SLZ13, SK16, SdOG12, TWM15, TMS11, Wal10, WHB10, Wis11, YLS12, ZP16]. dependency [SDNS16]. dependency-adjusted [SDNS16]. dependent [Ana17, Ane17, BML14, BT16, BV15b, BF10, Buc15, CM10, CS19, DFS^{+}13, DL13b, EAB11, EL18, Ery18, Fak10, Fen19, FN17, Ger13, GHH^{+}10, GZ14, HP13, HWB10, HSY17, HW13a, HCW13, Isl16, JT11, JGW14, Jir13, KO11, KP14, Ko13, LSCL17, LL19a, Lee19, Li11b, LL19b, LGW12, MKJ15, MR15a, Mat12, Mat14, MASR14, McC12, MMF17, Nak13b, PS11b, Puc13, RR11, SW13, sS10b, SZ13a, SX13, SXM16, Sze15, Tau15, Tru10, WY10, Wan12c, Wan12b, WMFW15, WQY12, WWH17, XWT19, YW10, YHW11, YS13, You17, ZZ15b, ZZ15c, Zha16, cZzXgL17, dS18a]. depends [LE19]. depth [BF17, DMS19, Nag17a, Nag17b, Nag19, PV17, Tsa13, Wan18]. depths [LRL19]. derivation [Mil19a, YHY16]. derivations [CF13]. Derivative [Son14, YYC17, HP12a, HP12b, SY17, YR11]. derivatives [JY15a, LWD13, Liu14, QY15]. Deriving [MGSL11]. Descents [ABBK15, ABB^{+}17, Yak15]. Descriptive [BM10c]. design [BB10, DFKK12, HR19a, HA12, JLJ16, KK13b, KK15, LZ17a, PNC15, SA12a, XD15, YSLL10, ZYL11]. design-based [SA12a]. Designs [CLWH17, ASP11, AWL17, AH13c, AHM19, BB16, BSS13, BS15b, BB11c, BDM17, BH11, CDS17, CKF17, CQZ15, CYLL19, DPZ16, DM12b, EQ14, EQ15a, EQ15b, EQ15c, Els16, Fil12, HP18b, HA12, He18, Jac10, JVVS10, KHN12, Kao14, KP17, KK19b, LCF15, LLO14, LZZ15b, LCLQ16, LTP18, LQ18, LYC14, Lu16a, Lu16b, LD17b, MPA12, MP13, MK14, MDPP16, MD16, MX18, OQ10, PPM18, PN14, Pro13, RS14a, Sin19, Sma14, Sma15, ySDM14, WZT14, WZYL17, WA17, WCZ18, WWS19, WZ17, XDWY18, Yan13, YCL14, YYS18, ZP12, ZW13, ZZ16, ZY11]. Destructive [PB16]. Detecting [BG19, NP10, RV19a, CTN17, JP11, ZG16]. Detection [Zho15, ASN16, HLLK13, MP13, Roh16, Tim15, ZWT15]. determinant [MWGT17]. determinantal [CF11, Olk14]. determinants [Vol14]. determination [Xia17]. determine [Ciu11]. determining [ABC10, CD18a, YDG13]. Deterministic [LOKB11, BBCD^{+}18]. detrended [ALL12]. developments [YH16]. Deviation [Sau12, BT14b, BK13, Bou16, BS19c, CW15, CG13, Che12b, DM12a, DL13a, Fan17, Gas10, Gee15, Hu15, JW16, Kou12b, Kre18, LWYZ17, LLP19, LY16, Lu11, Lu12, NKKY13, PO18, Rao18, Roz12, Roz14, Roz16, SXM16, Shi19b, Sun13a, SZ18, Tao19, VS13, Wu14, XT11, Yab17, YH14, ZLC18, ZZ19b]. Deviations [BBL11, Yan15a, Yan17a, BM13a, BR17a, BZ13, CHM15, DS14, Fan15b, FC15, Fen17, FS17b, FKZ15, GM14a, GW15, GZ15, GW16, GM18b, GMP19, HT13b, HCW13, Hir11, Hui10, JCM15, Li19, LZ13, LLY19, MX17a, Mac11, MP15a, Mac16, MP17, PTW10, SZ12, Son10b, ST10, Vol14, Wan13b, YLS10, Yao18, Ye16, ZG10, ZL12, ZZ13b, ZL14, Zhu13a, Zhu14, Zhu15, BC11]. dexterous [ST14]. Diaconis [Hil19, LP19a]. diagnosis [MvdBV^{+}11]. Diagnostic [LCX12, Duc04, Duc10, Ema15, SdOG12, TL15]. diagonal [Wys12]. diameter [MX17b]. DIC [KD14]. Dickey [Arv17]. Dickman [DF10]. difference [AH13c, AT15b, LG13, SR12a, STD12, TL15, TWM15, Wan12c, cZgL12, ZZ15d]. difference-based [LG13]. differences [CSY12, HWM15, KKMSA13, LSS15, MGG11, RM13, SS13b, ST13b, SDJ15, SLC16, WYY14, XSZ17]. Different [AEK19, JKD15a, JKD15b, WN10b]. differentiability [Suz18, Suz20, WX19a]. differentiable [Efr19]. differential [ADH19, BBM19, BCD19, BH10b, BH12, BH17c, CHM15, Che10c, CW12, CvGL18, CD10a, CYS11, De 11, DO11a, FRZ10, FS17b, GLM^{+}13, HB18, HLW10, JY15b, Kre17, KKT14, LV11b, LZ10, LL12a, LL12b, LL17c, LF17, LFM14, LTX12, LR13b, LRH15, MX17a, MB15, MB16, MSS18, Nak13b, NT17, NL19, OSW18, Owo15, PJ10, QX13, QLH^{+}16, RH11, RSR16, SS13a, TJS15, TJS13, TH14, THC17, VT10, VD18, Wan17a, WS17, WH11, Wu11a, WWR12, WZ13b, WW18, WX19b, XL18, Yan19b, YD12, Yu19, ZYL14, ZYS19, ZX17b, ZR12, Zho15, ZF16, ZC13]. differentially [BO19]. Diffusion [JS15, PTW10, AaM18, CL19b, CP19, DMPX17, Fan15a, Fuj10, GS17a, HMS14, Hui10, KP14, KMV11, LWD13, LC14, Mak10a, MOD18, Mih12, Mis14, MST10, NA13, NL19, Sab17, Sab18, Sak12, Sak15, SL13b, SWH12, TN17, WZT12, XY13, XYZ16, Yun18, ZW15b]. diffusions [Abu12a, Abu12b, BC11, BTT11, Cui14, HZ16a, KOR15b, MN14, Pen14, SY18, SY19b, WJY11, XSL^{+}14, YYY17]. digits [LCJ10]. dihedral [McC13]. dilation [SKB14, SKB15]. Dilatively [KW15]. Dimension [BS19b, MM13c, MM14c, AT15b, CZ15, DZ18, Gho11b, KL17, KSW15, Li11a, MNX13, Nos19, Par17b, Pot15, Vie18, Yoo15, YDG13, ZCZ18]. dimensional [Abu12a, Abu12b, AH11, AZ17, AV14, BS19a, BV11, Bis10, BL11, BCNM15, CZLT18, CJN19, Ciu15, Dav12b, FJ12, Fan16, FS15, FN17, GS17a, GZWW15, HS16a, HP13, HWB10, HZW16, HZ16a, Hua16, Izu13, JL16, KS17a, LSCL17, Li15a, LZ16b, LW15c, MLF13, Mao14, Mao15b, Mao16, Mod11, Nak13b, NHBG19, Osȩ15b, Pan19, PL19, PQW14, Rau13, Rob15, Sab18, SG15a, Sha12c, SZ13a, SY19b, Tom16, WW14, Wan14a, WZW17, Xue17, Xue18, Xue19, YJLL16, YYY17, Yao14, ZZ13a, ZZ15a, ZPZ18, ZX17a, Zhe11, ZGK13, Zhu15]. dimensionality [HL13b]. dimensions [Bes14, DWW17, DMWX15, DTV16, Gov15, Li11b, PRD13]. diminishing [AKV12]. Direct [Len11, Mad15, Chi10, VD18]. Directed [YCQY16, CC16]. direction [AN15, CSK13, HW13b]. direction-mixed [CSK13]. directional [EG11, GZ18, YDG13]. directions [CZ17a, JD17, Tsi12]. Dirichlet [BJW17, HNW19, Jon17, Shi19c, WZ13a, WX18, ZA12, Zha19b]. discontinuous [LL12b, NT17, SG15a, TJS13]. discounted [BS11, DZD12, JYL12, Kat19, LZZ15a, LXZ15, MWH17, WMH11, YSZ19]. discovery [BHJ18, DR11, HA14, Luo10, WLY^{+}14]. discrepancies [CLQ12]. discrepancy [BM15, CQZ15, CS18, CCGPW17, DRZ19, EQ14, EQ15a, EQ15b, Els16, YZ18, ZWHQ15, ZZ16]. Discrete [AAC15, Ery15, PS19a, Ale18, AvHZ17, BD15, BDK18, Bal11, BMP10a, BBL11, CG12, CS15, CF13, CJY14, DMC17, DFL12, DFL13, HMT13, HJS19, HY13, JLY14, KKL18, KZ10, KV16, LLK^{+}18, MZ17, OV15, PZ14, QLH^{+}16, RR14, Sas13b, SL13a, Seo15, Shm13, SvdV17, TN17, WY10, Wan19, Yak15, YHMM15, ZW15a, Zha17a, ZGG16, Zhu15, SI16]. discrete-state-feedback [LLK^{+}18]. discrete-time [CF13, JLY14, QLH^{+}16, Sas13b, YHMM15]. discretely [ES13, HMZ11, KW14b, Ma10a, Yan17b]. discretization [GMM10, ZHWH12]. Discriminant [HZW16, Lop13, Vel12]. discrimination [IV19, JL16, NSL11]. discussions [CYW13]. disease [Bow16]. diseases [CHN14]. disjoint [Yor14]. disparity [KB15, Liu17]. dispersion [AKM19, BLL10, Bre12a, DBB13, Gre11, LD10, Ric10, RSC10, RBSB16, XW10]. dispersive [BNB16, DM11]. distance [BMP10a, Bob13, Dav17, DY10, DER15, GEC17, KV16, LNI16, LM18, MXZZ18, Sas13a, Sch14, SR12b, Wal16, WWW10, Xia17, Zha19a, ZZ19a, Zin13]. distances [JP12a, Mak11, MM14a, Mod14, Mod16, RH16]. distinguish [BD11]. distorted [SGV15]. distortion [CV19, LPH13, WZ15b, YH16]. distributed [BS12b, CM17, DS16, KA10b, Lin17b, LC14, LMS17, Ma13, McK17, PLL13, RS10b, vBC19, Fin11]. Distribution [Bow17, CVA13, Khm17, Mes14, PW15, QL11, Rob15, Rob19, SD10b, WN10a, WHB10, Zen14, ZHWH12, Abd11, AN10, ABI19, AB11, AKP18, Arn11, Aya13, BT14a, BW15, Bai12, Bai13, BRA^{+}12a, BDK18, BD13a, BS15a, BS17a, BGM19, BSL13, BLB16, BP19a, BBGH12, Bel16a, BMN16, BPXS19a, BPXS19b, BBP17, BM10b, BM11, BV11, BSWS13, BRBB14, BRBB23, BV15b, CG12, CRC10, CMT14, CMW18, CQ17, CWZ12, CYW13, CH14, Che18b, Che18d, CR13, Cho16b, Cla14, CLG13, CL11, DBB13, DD10, DN14b, DK15, DF10, DEM14, DM17b, DP19b, DM12c, ER17, Ejs16, EIW14, Ery11b, Ery14, FGA11, FLRS13, FJW15, Fro17, FXT12, FP17, Fur08, GO12, GD18, Gat13, Gau14, Gei17, GF11, GEV16, GSW11, GT16, HV16, Han10, Han12b, HA13, HWB10, HR19b, HZ16a]. distribution [HZZ19, Hew17, HJM19, HJS19, Hür13, JT18, JBS10, JHF15, JW17, Jur10, KRYAV16, KLM17, KI15, Kim08, Kim13, KPO13, Kru10, LO13, LCJ10, LT17, LLP11, Li11a, Li13a, LBS17, LZ17b, LTW18, LQ19, LH13, LLH15, LP17, MU10, MT13, MA10b, Mam15, Mic11, MGSL11, MB14, MA11, Mir14, MM14b, MYA15, NAC13, Nag13, NKKY13, NA13, ND12, NW13, Now16, Oes15, OH15, OT15, Pan17, PT13, PDW10, PLN16, PS18b, PCGV14, Pet11, PTT19, Pin14, RH16, Rob10, Rod13, Roz16, Sab18, Sch14, SV10, SK15, SB16, sS15b, SCS14, SWZ19, Shi19b, Shi12, ST14, SKRT16, SR12a, Sre12, Ste15, SF12, Tak12, TJS15, TZB13, Tsu10, Tud14, Tum15, TW15, TE12b, WF15, WL15b, WH10, WN10b, WN14b, WN15, WN19a, WN19b, Xie13, XT11, Yab17, YC16, YH16]. distribution [YWY10, YZ13, Zen17, Zen19, ZGG16, Zho13b, ZB15, ZRN17, vSK15]. Distribution-free [QL11, SD10b, WHB10, GO12, Gau14]. Distributional [HNW19]. Distributions [BK19b, ZM16a, AH13a, ACP19, ABGM16, Bak17, BS14a, BLKL13, BCGFR13, BS18a, BBdWG16, BPRS13, BBL11, BB11b, BNM13a, BM11, Bis10, BLS14, BN10, Bra11b, BVV14, CK16a, CKO19, CXW13, CL16, CL15, CF11, CLG13, CP17c, DSS18, Dat13, Dom13, EB17, Ery16, FMA16, FA12, FL10b, FS14, Fra11, FS10, GM19, GZSZ19, GGV14, HMT13, HdCC17, Has10, HEM10, HB10, IMU10, INO10, IA10, JB13a, JB19, JP17, Jon12, Jon18, Jur13, KAVRY17, KM13c, Kha19, Khm17, KZ10, KKNH18, KV17b, KP16, KA11, LBG11, LVY15, LSS15, LW12, LN15, LLH15, LMM13, LM15b, MU13, MK10, ML16, MA14b, Mna11, MM13c, MM14c, Mu15, Myr19, Nad13, Nag19, Nak15, Nan10, NS10, NP10, NSH17, Oga18, Ost17, Ost18, Osu12, ORO15, PV17]. distributions [PSS13, PL15b, PZ14, RS10a, Rat15, RBY10, Roh16, RS10b, RA12, Rub15, SAM13, SM10, SD16, Sep18, SKJ15, Sha12a, Shu16, Shu18a, ST13a, Son15, Spa10a, Sre10, Ste19, SW19, Sza15, SI16, TV17a, Ter17, Tra14, TE14, UV13, VM18, VGE15a, VGE15b, Vin11, Wan11b, WX18, WL17, Wes15, WN10b, WN13a, WN14a, WWD15, Wu11b, Yan15b, YIS15, Zha12, dC16]. distributiveness [Il’18, KS16a]. disturbance [WZ16]. diurnal [Fen14]. divergence [Bro14b, CO14, CK14, Din15b, SSN16, TE12b, Din14]. divergence-free [CK14]. divergences [GB16]. diverging [HL13b, LKK12, LL10, LZY14, WSW10]. Diversity [WK10a]. Divided [SR12a]. Dividend [LC14, PLL13, SYZ18, Bra12b, DZD12, JYL12, LH14b, LXZ15, SLZ13, YY11b, ZDD18]. dividends [DZ19b, GW12a]. divisibility [Res11]. divisible [BLL10, Bra11b, DM17b, FS17a, IMU10, Kha19, Nak13a, SW19, Tra14, WK11b]. DMRL [Nav18, SD16]. do [Cao18, HR19a, Shi18]. Dobrushin [Sze12]. Does [dB13, ATV10, BH14]. domain [BK10, BCR10, GGS12, Kur19, Mar12b, NW13, TK10, Vas18, Yan19b, Zha17b]. domains [BR17b, BBCD^{+}18, YHW11]. dominance [Osu12, Pol13, TW15]. dominant [KM12b]. Dominated [Mau10]. dominatedly [YW10]. dominatedly-varying-tailed [YW10]. dominates [BS15f]. domination [CS15, RS12a]. Donsker [ABTV19]. Doob [Jov19]. Dot [NM19]. Double [Che14c, jLL19, RW16a, Abu13, AD19, FCU11, LS16b, Liu12a, MD13, MvdBV^{+}11, NAC13, QHS13, QN15, ZY10a, ZW15b, ZZLL17, Zin13]. double-barrier [Abu13]. Doubly [KRB13, SB10, AM13a, FRZ10, FKZ15, GLM^{+}13, LL12b, MN11, MN13, Owo15, SWW15, Tan19]. down [Sha17, WM18]. Downton [KK11]. draw [WM18]. draw-down [WM18]. Drift [BTT11, Alf13, CVA13, Coq15, Gas10, LE19, Mis14, Pam18, WYY13]. drifted [Gat13]. drifts [SZ18]. driven [Abd15, AC15, BPS14, BTT11, BTT16, BH12, BH17c, CZ17b, DY10, DO11a, FRZ10, FY16, Fen14, HB18, IM15, JB13b, KS16b, LY18a, LT19a, Liu12a, MY14, OSW18, Rad19, RH11, Son12a, Son14, Son16, SG18, Vri16, WW10, Wan11a, WZ19, WSMG18, WS17, Wu11a, XY17, XL18, Yam15, Yu19, ZZT17, ZYS19, ZX17b, ZWB17, dlCJB19, DMST13]. drop [KL17]. dropout [SHBHD11]. dropouts [CM11b]. DTI [Sak15]. Dual [BW18, CM16, BB10, LC14, PLL13, SKB15, WN14c]. Dual-record [CM16]. duality [DP10, MZ17]. dually [BH11]. Dubins [Bak12]. Dunkl [Osȩ15b]. Duration [LW15b, AH12a, RR14]. durations [Fen14, KFHS11]. Dvoretzky [FP11b, WD12]. dyads [HS13a]. Dyck [Kou12b]. Dynamic [Buo18, BS18c, HH16, AAMB10, CN16, DD14, Kay15, KH18, LN15, MNP16, NSL11, Smu19, Vil11, WLP17, Zha19b]. dynamical [Hor16]. dynamics [Bob10, KT10, LLW14, LLY19, ZZT17, ZHL13, Zho15]. Dynkin [Che10d].

E. [Bob13]. each [WN10c]. earnings [Kin12]. easily [Tsi12]. edges [Cai17]. Edgeworth [FD10, WN10c, WN14c]. Editorial [Ano17k, Ano17l, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano14l, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano15k, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l]. Editorial [Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l]. Effect [VT11, ART14, BM10a, CLYW15, Chi10, Chi12a, Chi12b, CPH12, DM11, Hat12, LTvdVR11, LS10, TSZ15, WHB10, WZT14, Zho13a]. effective [Mis13]. effects [Bak18, BKP15, Bre14a, CLWH17, Chi10, De 16a, Din10, DHL14, DS15, Fil12, HBL11, HLB17, HHF17, Jac10, Jac11, Jac13, JKD14, JKD15a, JKD15b, KWL15, LLL13, LZ14b, LZL13, MB15, MB16, MP13, MMPW16, Ros13, Rub15, WCM11, WT18, WSMG18, WS10b, WQD15, YXL14, ZHL13]. efficiencies [MO16]. Efficiency [DM12b, Mar11b, Ros13, XSZ17, BB15, CZ17a, DN16, GR10, Gir16, Han12a, KS10, Luo12, Maa19, PSS15, Uem16, Wan13c, WN13b]. Efficient [ABI19, Cho16b, LZLW16, MW16, MW14a, RW16b, SM19, Gat15, KJS15, Onz11, SW12, Sha13a, SS15c, Tan15a, TMN16, WLS11, WL14a, Wri17, YGL14, ZM17]. Ehrenfest [CGLN17]. eigenfunction [GLM19, LL14]. eigenvalue [WZ13a]. eigenvalues [AS18, BdS13, CQ17, DTV16, HLCH19, JW16, LLSW11, Pac17, PR17, Red17, Rei15, Zen14, Zen17]. EIV [ASH18]. elastic [Zho13a]. elasticity [KP14]. election [KMR13, KW14a, LW15b]. electric [CGLN17]. Electronic [Sha18]. elegant [CF13]. elementary [Aga15, DS11, Doe18, Kir19, Mil19a, Yan12, ZCM15]. elements [Bud14, CC10a, Rao10, ZH12b]. Elicitable [WZ15b]. elitist [CP13b]. ellipsoid [Zen14]. ellipsoidal [MM15]. elliptic [CHW18]. Elliptical [HS16b, Cha19b, Fra11, Has10, JP17, LVY15, Shu16, Shu18a, Shu18b, Shu18c, Shu19b, VB11]. elliptically [Tsu10]. ELR [Kak11a, Kak11b]. EM-based [PB16]. Embedded [Sta12]. embedding [Bak12, BS15d, Jia16]. Empirical [CQ17, Dur13, HZJ^{+}10, Hua12, LGP10, LLHW10, LLH11, LK14, MJH18, PQW14, SL13b, SJ14, VZ18, YL10, ZZ11, Zha11b, ZZM15, ZZ15d, BK10, Bis10, Bou17, BD10, Buc15, CO14, DS18b, EIW14, Gir16, HHF17, IK10a, IL11, LQ19, LRL19, Mar14b, MU16, She13, Shi19b, SK11, SA13, TL12, Tau15, Tsa13, VL19, Wan12a, WW10, WZG16, WN10b, WR10, WT14, Zen17, ZLZ16, ZY11]. empirical-variance [Gir16]. endpoint [BB15, GGS12, NP10]. endpoints [Gho12]. energy [Gir16, RR11, Wan13b]. Engel [Fan15b, Hu15]. Engelbert [Cui14]. engineering [LAG^{+}18]. Enhanced [Vil12, MvdBV^{+}11]. enough [Roz16]. ensemble [CCM19, CQ17]. ensembles [Che18a, GBR12, JW16, NW13, Su10, JW17]. entries [Arb19, BGHS11, HSY17, Löw18, Vol14]. Entropic [Bob13, Bow13, Bow17, Li13c, Yan14, Yan15a, Yan17a]. entropies [Kay15]. Entropy [MT17a, AAMB10, AEAN18, ACMM13, CN16, CW19, Cho18, CD10b, FA12, GB13, KK13a, KRB13, KT11, Kum15, Li13c, LR12b, MASR14, MY11, NSS14, PK14, RBY10, SC12a, SS12b, SSN13, TJE16, TE17, Xia19b, YG13, Yan15b]. envelope [BDvdAW16]. envelopes [Gri19]. environment [ABD14, BP19c, LLP19, MLF13, RV19a, WGL11, WLLL17]. environmental [BOK17, BL10b]. environments [De 17, LMLW15]. Epidemic [Mar16, Cla14, El 13, LO13]. EPPF [MW12a]. epsilon [PK11, Par17a]. epsilon-insensitive [PK11, Par17a]. equal [BL10a, Fil12, Sha13a]. equality [Oh14, RS10a, SL12]. equally [CKF17]. equation [BC14, BN16, BH10a, DO11a, HY13, JLY14, KM13d, Li17b, LWYZ17, LXL19, LHW19, MZZ15, NK17, NA13, Nak13b, NTT18, Pac16, Pac17, RZ12, Rad19, Son12a, Tie13, Vii16, Wu11a, YYS16, YY18, Zho15, ZF16]. equations [AM17, AH13b, ADH19, BBM19, BH10b, BH12, BH17c, BH18, BL14b, BL15, CHM15, Che10c, CW12, CvGL18, CF11, CYS11, D’O11b, DOT14, De 11, DDE17, FRZ10, FS17b, GLM^{+}13, GA13, Gov15, HB18, HLW10, JY15b, KT19, Kre17, KKT14, LV11b, LZ10, LL12a, LL12b, LL17c, LF17, LZ17b, LCKK14, LFM14, LTX12, LR13b, LRH15, Ma14a, MX17a, MB15, MB16, MSS18, NT17, NL19, Owo15, PJ10, QX13, QLH^{+}16, RH11, RSR16, Sab18, SS13a, TJS15, TJS13, TH14, THC17, VT10, VD18, WWW14, WWX17, Wan17a, WWZ18, WS17, WH11, WWR12, WZ13b, WW18, WCD19, WX19b, XL18, Yan19b, Yu19, ZYL14, ZIPP17, ZYS19, ZX19, ZMG10, ZX17b, ZR12, ZBL19, ZC13, aZLW18]. equi [DR11]. equi-correlation [DR11]. equicontinuity [Jar13]. equidistance [HA12]. equidistribution [Gna12]. equilibrium [BCD19, GM18a, YE10, ZZ15a]. equity [QY15, WWY12]. equity-indexed [WWY12]. Equivalence [You18, De 16a, JGW14, LT10, MK18, Ost17, Ost18, Ren14, Ren15, Sre10]. equivalencies [SA12a]. Equivalency [HLR15]. Equivalent [CV19, GZ13, EKS11, Maa19, WYY13, YWY10]. equivariant [BV15a]. era [BBM18, Cao18, Cer18, Dun18, Jam18, Qua18, San18, Sco18, Sec18, Sha18]. Erdos [LT14, GGR11]. Erdos-Rényi [GGR11]. ergodic [AG11, CL19b, Fuj10, GLS13, HZ16a, HZZ19, Hei18, Löw18, Mia14]. Ergodicity [AD19, ABD18, CJ14, Fit14, KLY14, KR19, Lee12b, Li13a, MS10b, MN16, NW10b, RS14b, Son16, ZZT17]. Erlang [HBH10, LMW15, YLHY19, ZB10]. Erratum [Abu12a, BRBB23, GHR17, MN13, MM14c, Sko20, SF12]. Error [JW17, LK13, RS15b, TSL18, AHN10, Bha17, CWZ12, Che19b, Che12b, Che18d, Cho16b, Dat13, Döh14, GM14b, GG12, HS13a, HZJ^{+}10, HS15, Hwa19, KST18, KN11, LZ10, Luo10, Nie16, Nku11, SV10, Shi12, SS11a, SDNS16, TNP17, VB11, WX12, Wan13c, WQD15, cZgL12, vW17]. error-prone [HZJ^{+}10]. errors [AR18, DPZ16, EP17, FA14, Gee15, Hua12, Hwa13, JH14, LWX11, LK14, MR15a, MMC14, PT19, RP19, SCS14, SSL^{+}12, Sma14, Son10b, SY11, SZ13b, TZW13, TXWW18, Wan19, WZ17, WS15, XXY17, XFC17, YSLL10, YGTT13, ZW18a, ZL12, ASH18]. Errors-In [ASH18]. errors-in-function [Hua12]. errors-in-variable [WS15]. errors-in-variables [SCS14, SY11, SZ13b, TZW13, XFC17]. ESD [Bao12]. Esseen [LWX11, MC15, WJY11, ZJ15, ES13]. essential [HL16, KKW11]. ESTAR [Han12b]. estimate [ALS11, BS11, BCL18, BK13, CL19a, DBB13, DL13b, Han18, He12, KM12a, Nic17, Osȩ15a, Ste19, ZW11]. estimated [Abd15, AH13c, ROSL17]. Estimates [Mis13, WYY13, WN11, Aya13, BGM19, BOR17, CHW18, Che14d, Cou19, Dun16b, GLM19, Ilm13, Jov19, KST18, KL15, LZ10, Mel16, Mes10, Osȩ10a, PP11, SHL15, SG18, Tum15, Vir16, Wal10, ZJS13, dB13]. Estimating [BLZ13, BDER14, CDG15, Chi10, FF12, GGS12, GSW11, HS13a, LSS13a, MM14b, MT17b, PY17, Ruk12, SS12a, Tsi12, WCD19, YSZ19, ER17, FP17, Gir16, HB11, JDK^{+}11, LCKK14, LZL^{+}19, MZZ15, MPA12, QCZ15, ST14, YS19, ZIPP17, ZMG10]. Estimation [ADP12, BK17, CHM16, DFS^{+}13, GG13, HX14, HCT16, HH18, HMZ11, KK13a, KKW11, KK15, LW13, LQZ13, LR11, LMM13, MDN18, NT14, PP14a, Pen11, Rah11, TZT16, WS10b, Xia19c, YGTT13, ACD12, AaM18, AF12, AAEH11, AO12, ABI19, AM14, AT14b, AM13b, AM19b, ANRW15, BD15, BDB19, BFS16, BBdWG16, BDS19, BMN16, BTT11, Bha13, BBS11, BY12, BD13b, BF10, BCNM15, CKP17, CFS12, CDG15, CMT14, CK13, CP11, CS17, CM16, CDL11, CSG11, CXW13, Che19b, Cho16b, CP17b, CW16, CZ17b, DNT10, DNKL12, DFH18, DD11, DE13a, DE13b, DFKK12, DY10, DHL14, DSX13, ED14, Efr10, Efr14, Efr19, EL18, FA14, FPZ13, Fen14, Fuj10, GS16, GM14b, GGV14, GGQ17, GHR16, GHR17, Gre11, GZWW15, GW12b, HS12a, HC10a, HP11, HCP19, HP12a, HS19]. estimation [HN10, Huh12, IV19, Ili12, Jac10, JQZ12, Kab13, KJS15, Ker16, KN11, KMJ10, KLNB14, KP15a, KN17, KKW14, KZ10, KKP15, KS17c, KB15, KLK12, LLL13, LML15, LZLW16, LYW11, LP13, LZT13, Li15b, Lia12b, LHT15, LX12, LWL16, LB17, LW17b, Liu17, LZL^{+}19, LK13, MTT11, Mao15a, Mar12a, Mas11, MS16a, MASR14, MS12b, MMC14, MK18, MHH11, MNWA14, MMF17, Nag18, NS12, NSH17, ND10, Now19, OKT13, Oua13, PLY18, Par14, PL19, PPA16, PS11a, RP19, RR10a, Sak12, SW12, Sch12, Ser18, sS10b, SMZ18, Ski15, Son10a, SZW10, SS11b, SB10, SS15c, TL19, Tak12, Tan19, TZW12, TA14, Tsu10, TS17, Uno13, VSP13, VB11, WZ15a, WSW10, WW10, WJY11, WZT12, WW13a, WL14a, WP14, WYC15, Wan17b, WZW17, Wu14, Xu14, XH15, XS18, YGL14, YY16, Yan17b, ZV15, ZYW15, Zha11a, ZZ12a]. estimation [ZIPP17, ZYS19, ZJL13, ZZLH15, ZX17a, ZZ19a, ZKG10, ZJZ16, ZL14, ZP15, ZF16, cZzXgL17, Zhu13b, ZRAN13, ZRN14, ZRN17, dWGM12, vDLMW19]. Estimations [Yan14, LKK12, ZW18a]. estimator [AFJ11, Ana17, AH12b, BB14b, BB12, BSK19, Bel16a, BB18, BG15a, BG15b, BRZ10, BH11, CRC10, Che19a, CW11, Che18d, DL13a, DMPX17, DZX18, ES13, FK12, FP11b, FG12, Gat15, GGO15, GT14, GSST12, Han12a, He18, HP12b, HWZ14, JW17, KLM17, KZ17, KZ14, LL19a, LLHW10, LWX11, LLC^{+}12, LW16, LZ13, LM18, Lu15, Luo10, Luo12, LK13, Ma10a, MM10, Mar11b, MN11, MN13, MWZ11, MNOT12, Mis14, Ohy13, OO11, ÖA15, PSS15, Par14, QO16, Rao18, RR14, SX13, SCZ15, SK11, SA13, TXWW18, TSK13, Uem16, VD18, WMZW13, WF15, WZ16, Wes13, WW16, WS15, XJW15, Yab17, YH14, Yao12b, YWZ14, ZPK14, ZM17, ZJZ13, ZM16b, ZLC18, cZgL12, ZG17a, dMO18, vW17]. Estimators [MSW15, Ane17, BL10c, BMS10, BV15a, Bou15, BS18b, BS19c, Bro14b, BPS17, CP14, CMS19, Che12b, DZ18, DG13, EBG16, Fak10, FS14, GG12, Han18, Hui10, IK14, IL11, KLM11, KBM15, KV16, KKNH18, KV17b, KS16b, KMS15, LLH11, LLCG12, LYB13, Lin17b, LB17, LW15a, LM15a, LPS15, LG13, Mac11, MP17, MS16b, MS12a, NM18, Nie19, Par14, PH10a, RNB17, RSC10, Ros13, Sak15, Sak17, SA12a, SWW11, SD10a, sS15b, Son10b, VR12, WLS11, WW14, WC15, WS16, Wis11, XSZ17, YH17, Zha10, ZBZ11, ZL12, ZJ15, ZGR12, ZC12]. ethics [BB11c]. Euclidean [HNGS15, Tyk11, Zen14]. Euler [Alf13, Kre17, LT19a, NT17, NL19]. EV [LHT15, MWZ11]. evaluating [Kak16, ROSL17]. evaluation [Elh14, PR15]. Evaluations [Raq10]. event [Kao14]. event-related [Kao14]. events [FCP18, Fro12, sS15b]. evidence [DdRS^{+}11, Gre12]. evolution [AH13b, FNP18, GFH16, GA13, Gov15, Khm13, LV11c, Van18]. evolutionary [HS16a, LCKK14]. Ewens [Rod13]. EWMA [NMS15]. Exact [AD12, ADM15, EB17, Gao19, GP17, LH13, Mai18, MKA19, PTT19, Pin13, RH15, Sha14, Wri17, Yan15c, ZK10, AB11, BCI14, Fre16, HA14, MKK17, MW14a, MWGT17, NKKY13, Wan17b, ZWLB17]. example [Khm13, Zha17b]. Examples [Coh10, Li17d, MU13, LW12, Lin19a]. exceedance [EB17, SK13]. Exceedances [PTW19, BN10]. exceeding [GM14c, PR16, ST19b]. exceeds [Doe18]. excellent [Bow18]. excess [LZT13, LH14b]. excess-of-loss [LH14b]. Excessive [SY18]. exchange [Tsi12]. exchangeability [Bow16]. Exchangeable [SW19, BBHV13, Bis10, CP13a, CL13, Di 11, Ery11b, HPP19, Mar14a, PDW10, RR16, ST19a, Sha13a, SL10, VA16, ZK10]. exciting [FP11a, KC19]. exclusion [ZZ15a]. excursion [AP13, BW18, MS11, ST19b]. excursions [SS10c]. exercise [PY18]. Existence [CvGL18, CP19, Fan16, HCP19, LSSR13, Liu14, Öns13, ZT19, FS14, LYB13, LS10, Mah18, MSS18, MM14b, Nog13, Shu18b, TJD10, Tie13, Tum15, Wan17a, WS10b, XY17]. exit [CCS17, GS17a, Gat13, LZ14a, LN19, PTW10, WM18]. exits [Vid17]. exogenous [Goi12]. exotic [Liu17]. expanding [Kur19]. expansion [DHL14, KA10b, LL14, LHM14, Tud14, Yam15, YSZ19, ZZ19b, vLT11]. Expansions [Nad13, Nad15b, ALL12, Ai16, Che18b, FQZ18, FD10, KST18, Li15a, LPN13, Nad15a, Nad16, WLP17, WL17, WN10c, WN14c, ZL16b]. expectation [AH13a, Bha17, Che10d, Doe18, DCR19, Fen17, Fen19, GM14c, Hu16, LWY17, LMM13, Ren13, Son12b, YIS15]. expectation-maximization [Bha17, DCR19]. expectations [BNM13a, DKM19, Had13, HC10b, Lin19b, SGV15, ZL18]. Expected [Kin12, RH16, BMNZ18, CGLN17, CPY15, CC16, Gor12, MX17b, MWH17, Nic17, OK15, Puc13, Sch13, VAB15, WMH11, XZZ19]. Expectile [ZCZ18]. experiment [LZ17a, PNC15, Wal16]. experimental [HP18b, HR19a, Kao14, XD15]. experimentation [RSV10]. experiments [AG13, BK19a, BKL19, HHY11, MD16, QO15, RS14a, SA12a]. explanatory [KHN12, ROSL17]. Explicit [SL15, ZCM15, BIK12, Mak10b]. exploratory [HP15]. explosion [Ger11, WZ19]. Explosive [HBPC10, BDvdAW16, Hwa13, KV17b, XL16]. exponent [CP17b, NS13a, Pen11]. Exponential [AV13a, CYS11, HR19b, HRB19, KLY14, LF17, WHYL10, Woj13b, YD12, ZZT17, AKP18, BRA^{+}12a, BS14a, BD13a, BCI14, BLL10, BSK19, BS15e, BV15b, Bro14b, CS15, Che10c, CW12, Che16, CCS17, CvGL18, Che18c, CD10a, CW19, DSS18, De 16a, DDE17, GM10a, GK12, GA12, HH13a, Hob13, HB18, Hwa18, IK10a, Jan18, KK13a, KI15, KKMSA13, KK11, Lee12a, LT17, LMP18, LSS15, LD17a, Lou13, LA13, MA10b, MKK17, MM10, MGG11, MA11, NSH17, ND10, Now16, Osȩ13a, OS10, PB15b, Rat15, Ric10, RA12, Roz12, SV19, Sas13b, Sha12a, SXM16, SS13b, Son14, ST10, Tum15, WL15b, WC17, YC16, YHMM15, Zaj19, ZW15b, ZWB17]. exponential-tilt [IK10a]. exponential-variance [PB15b]. exponentiality [CN16]. exponentially [LC14, PLL13, XXY17, vBC19]. exponentials [MWA16, vLT11]. exponentiated [BRA^{+}12a, FZ15, KC16]. exponentiation [Nak13a]. exponents [FY16]. exposure [LLZ13, TR12]. expressions [RSC10]. extended [BPXS19a, BPXS19b, Kao13, ND12, OH15, PS18b, Tum15, WT14, YH15]. extending [ZGZ18]. Extension [Che14b, KK18, Fuj10, Hay12, KL16, LLR18, ZL17]. Extensions [KZ10, Oga17, GM16, Min17, PPA16, YJ12]. external [Mar14b]. Extinction [LO13, WK10b]. extrapolated [Mau19, WC15]. extrapolation [HCT16]. extrema [KP16, Tan15b]. Extremal [BPXS19a, BPXS19b, DH19, FF14, PLN16, FF18, HJS19, WN10a]. Extreme [CH19, CH11b, GML15, BOK17, BMN15, CWZ12, CH14, Che14c, FMA16, GGV14, GGQ17, GHR16, GHR17, HdCC17, HCT16, LP10b, Mül19, Nad13, WA17]. Extreme-aggregation [CH19]. Extreme-trimmed [GML15]. Extremes [BS15c, PS16c, CXW13, FC15, Fer11, HW13a, LL19b, LPN13, LPNW14, Nad15a, Nad16, PS19b, Tan13b, WN19b, WW16, ZL16b, dH15]. extropy [JS19, Qiu17, QJ18].

facet [VB17]. facilitated [ZC19]. factor [ABC10, BdS13, Can16, CLYW15, He18, Jac10, KWL15, LD16a, LG12, MNP16, PW15, Smu19, WCZ18]. Factorial [BDM17, DM12b, Jac10, Lu16a, Lu16b, LD17b, OQ10, PPM18, RS14a, Sma15]. factorials [ZWHQ15]. Factorizable [PS18c]. factorization [Jur14, NHN^{+}11]. factors [ABC10, Aza13, CYLL19, DM12b, HWM15, LZ19, MDPP16, PPM18, WLLZ11]. facultative [TZB13]. failed [AB12b]. Failure [Bel13, BLS14, BS15f, Bou15, CTN17, ED14, Hat12, Kay15, KFHS11, Nan10, Rob10, Zha11b]. Failures [Dun18, CF18, Ger18, Mak11]. fairness [BB11b]. faithfulness [LL17a]. Fake [Hob13]. false [BHJ18, DR11, FP11b, Gor12, HA14, MGSL11, WLY^{+}14]. familial [Bow16]. Families [Jon18, AB10, BLKL13, BPRS13, BDJ17, Bro14b, CS15, CSK13, Che16, De 16a, GM10a, HH13a, LPW10, LA13, OT19, Ost18, ZJL18]. family [AMG17, CO14, CQT12, Che18c, Duo15, Fak17, Hwa18, LMP18, SS14, SDNS16, WX12, WYC15, WWD15, XLY17]. family-wise [SDNS16, WX12]. FARIMA [RS12c]. Fast [MN17, PR15, DWWG18, NHN^{+}11, Roz16, WH10, Wój13a]. FDR [BHJ18, GF13]. Feasible [WZT14, WCZ18]. feature [CZCL19, LSCL17]. features [Gho11a]. featuring [CPH12]. feedback [GA12, Kao13, LLK^{+}18, QLH^{+}16, YHMM15]. Fefferman [Osȩ17a]. Feller [DJ14, KSW15, LLR18, QHS13, SL10, XY13, YH15]. Ferrari [Kak12]. Feynman [KOR15b, LT19b, Mak10b]. Fibonacci [Chr12, Neu17]. fiducial [KKMSA13, VM18]. field [AP13, BCD19, CHM15, CK14, Cho19, DH19, DDZ15, EF16, FS17a, GS13b, Hil14, Kar11, Ko13, LL12a, LRH15, Mah18, Mar14b, Mou19, Öz16, Sha13b, Tur14, Xue17, Xue19, ZC13]. fields [BDP14, BDP10, BM12, Che14c, DLO16, DMWX15, DT10, HS16a, HT13a, IL17, Kam17, Li15b, LX10b, Ma13, MW11, MS11, Mor18, NC16, PP18a, PS16c, PS19b, Rob13, Söh10, SG14, STD12, ST13b, Tah12, Tan13b, TT17, Tru10, VS10, Vid18]. fig [BK18]. fig-wood [BK18]. filling [MX18, RS14a]. filter [KS17a, PJ13, Vil12]. Filtering [ELS13, Cou19, Lin13, Ngu16]. filtrations [KN15]. final [Wan13c]. Financial [Giu18, Bre14a, DD14, Fen14]. financing [LH14b]. Finding [Rao13, WRvdL11, WH10]. Finetti [Kir19]. Finite [AC11, DZ19a, FJ10, Jur10, JS11, PL17, AR13, Aza13, BH16, BS14b, BM10a, Bis10, CDG15, CJY14, CCSC11, Dun19, EF16, Fre16, GL13, GF11, HCP19, HL13a, HL16, Hil14, Jia16, Kab16, Li17a, LMR12, Mar12a, MT17a, NKKY13, ORO15, PD12, SÖV14, SSL^{+}12, SS12a, Spa10a, Wei19, YL19, ZK15, ZKSK18, ZF13]. finite-dimensional [Bis10]. finite-horizon [Wei19]. Finite-sample [Jur10, JS11, NKKY13]. finite-time [Dun19, GL13, Li17a, YL19]. finitely [CLEA18, VA16]. Finiteness [Bul14, NP10]. finitized [CLG13]. First [CCS17, Hie14, Iso15, ÖA15, SWW15, Abu12a, Abu12b, Abu13, BB14b, Bap11, BL11, Che18d, CP19, DK11, Fro17, GS17a, Gat13, HS12b, HR15, IMM16, LCJ10, Met10, Pen14, ST19b, Spa11, VT10, WZ13a, WZ17, XY12, Zha11a, ZH12a]. first-degree [WZ17]. First-order [ÖA15, Che18d, Zha11a]. First-passage [Hie14, Abu12a, Abu12b, Abu13, HS12b, IMM16, VT10, ZH12a]. Fisher [BSK19, Ern17, Gat13, HM10b, PNC15, RR10b, Sab18, TE14, Wal16]. fit [CCGPW17, DN16, GEV16, RH16, Rob15, Rob19, SD10b, SY11, VGE15a, VGE15b]. Fitted [Nos19]. Fitting [Whi12, Gir16, Ref14]. Fixed [BB16, IMU10, Lia13, Nie19, BKP15, Gra11, HHF17, LLL13, LZY15, LL17b, LV11c, Ros13, TL15, WZT14]. fixed-effects [Ros13]. Fixed-width [BB16]. flatness [AM19a]. Fleming [dSF12]. flexible [AR13, BOK17, LMPR19, PNC15, Rub15]. flights [De 17]. flips [BB11b]. floating [Sab17]. flows [Pam18, XWML19]. fluctuating [Tzi17]. Fluctuation [Li11b, AS18, GZ14]. Fluctuations [HSY17, MY14]. fMRI [Kao14]. Focused [ZDM17]. folded [BB11b, WW11, XT11]. foldover [EQ15c]. follow [TM16]. follow-up [TM16]. following [ABBK15, ABB^{+}17]. force [BL17, GL13]. forces [dlCJB19]. forecasting [DN14a]. forecasts [EPSU16]. forest [RP19]. forests [IK10b, MFL15]. form [BT13, Gau19, LW17b, Mou19, Pog16, Shi19c]. formal [RR13]. Formation [ASN16, Nav18]. forms [Che14a, EAB10, EAB11, LTW18]. Formula [Mic11, ADH19, BML14, Che10d, GM16, HT14, Kat13, KP14, KK18, Lo17, Maa19, Son14, SG15b, SLC16, WY18, Yam15, YY18]. formulae [Cho18, SL15]. formulas [DF14, Tak19, Zha12]. formulation [IK14, Mad15]. formulations [ABGM16, ML16]. forward [Cha19b, HLW10, JY15b, LG15b, ZX19]. forward-backward [HLW10, JY15b, ZX19]. four [EQ14, MDPP16]. four-level [EQ14]. Fourier [CMT14, FM13, Kat19, Pot15, WY18, YSZ19, Zha17a]. fourth [Gau19]. Fractal [MNX13]. fractals [Pot15, RZ12]. Fractional [DOT14, DLO16, FS18, KKT14, KMV11, SY17, ADH19, ABTV19, AH12b, ATV10, BDP14, BTT18, BC14, BM13a, BM17, Ben16, Ben18, BTT11, BMNZ18, BH12, BH17c, CVA13, ÇKN17, CY11, CHW18, CP17b, CM11c, D’O11b, Dav12b, DO11a, Dun06, DF11, Dun19, ES13, Fan15a, FY16, FM13, GWK17, GT16, HT13b, HB18, HN10, Jac10, JB13b, JV10, Jur13, KV17a, KP15a, KKNH18, KS16b, KS17c, LX18, LST17, LYCK11, Li17b, LY18a, LZ11, MS13, Muk17, NA13, Nol18, OP12, OP13, OSW18, SS18, Sau12, SC12b, Shi19a, SS17, Son12a, SV17, SG15b, SG18, TN17, VAB15, Wan14d, WYY14, Wan19, WX19a, WS17, Wu11a, XTT18, XL18, Yam15, YS10, YYS16, YY18, Yu19, aZLW18]. fractional-diffusion [NA13]. fractionally [Wu14]. fractions [AA17, Fan15b]. fragmentation [Gho11a]. frailty [BSK19, GB19b, GM10b, HS12a, HD13, HP14, LI12, Lin17a]. Frame [EG11]. framework [AHM19, AMP16a, EPSU16, Gho12]. frameworks [Ale18]. Frechet [AT14a, GSK12, BR17b]. free [Cha17, CZCL19, CK14, DH19, GO12, Gau14, Khm17, KS17a, LSCL17, Pei15, QL11, Rob15, Rob19, Sab17, SD10b, WHB10, ZZZ18, Zho13b]. Freedman [LP19a]. freedom [XKBG15]. Freidlin [BC11]. frequencies [SS12a, WK11b]. frequency [BB14a, BCR10, KM11, MS18a, SB16]. frequentist [BM15, CM10, Won13]. Friedman [GM18a]. Friedman-like [GM18a]. front [LLP11]. Full [vWvZ17, PSS15]. Fuller [Arv17]. fully [JY15b]. function [AFJ11, ABI19, AB17, AM13b, ATV10, Bel16a, BNM13b, BRZ10, BS18b, BS19c, Bra12b, CTN17, CO13, CRC10, CMT14, Che18b, Che18c, Cho16b, DZD12, DL13a, DDZ15, DL17, ER17, FS11, GBR12, HP18a, HH18, Hua12, Huh12, JYL12, Jou14, KM11, KLM11, KN11, KH18, KKW11, LL19a, LZ13, LW15a, LXZ15, LZL^{+}19, Lop13, LP17, MM10, MASR14, MU16, Mes10, MO14, MS18a, ML17, Nag18, NV11, NS13a, Nak13a, NSH17, Oga18, Osȩ12a, Osȩ14c, OQ10, Pak18, Pin15a, sS10b, SWZ19, Shu16, STW19, SS12b, SSN13, Tah12, Tsa13, Urb12, Vri16, Wal10, WMH11, WMZW13, XT11, YS19, Yas13, YHY16, ZK10, ZY10b, ZJL13, ZJ15, ZWLB17, ZRN14, ZRN17]. Functional [Ane17, BGT15, CK16b, Lo17, MS11, Rei15, Tor13, TK10, Vas18, AM14, ÁLBRM16, ALS11, BCL18, Ben17, BB18, BV15a, BH10b, BH12, CDG15, CD10a, CHM16, DNKL12, DLZZ18, FJW15, Gee15, GLS13, GDNM18, Goi12, HWZ14, JY15a, JY15b, JN10, KLY14, KM19, Kre18, Kri18, KOR15b, KV13, LSS13b, LCKK14, LX12, LWL16, LZ13, LTX12, MH12, Mak10b, MS18b, MMC14, PZ17, PJ10, RSR16, Sha13b, TJS15, Vie18, WHB10, WZ10, WH11, Xu12a, YD12, ZW18c, ZS19, ZC12]. functionals [BGT15, CG12, Coe15, Dun16b, Jir13, LXL19, Oes15, PP11, SV19, Son12b, WN10b, YY13]. functions [AKM19, Alb18, Bai12, Bai13, BP15, BK10, BW19, BOR17, CDL11, Che16, CM17, Che18c, CGH11, DEM14, EHP13, ED14, Efr19, EFP19, Fro17, FP17, Gir16, GAB16, Had13, HS11b, HZR16, HZRS19, HB11, IL17, JT18, Jal16, Jar13, JDK^{+}11, Khm13, KZ10, Lar15, Li11a, LMLW15, MM13a, MV15, MDR12, McK17, Nag17a, Nan10, Osȩ17b, Ose19, Osȩ17a, Poi19, QN15, RS15b, SNS10, SS14, sS15b, Shu18b, Sub12, Suz18, Suz20, Tra14, UU11, Wes13, YY11a, ZZ11, ZZM15, Zhe11, ZZ15d, ZRAN13, dMO18]. fundamental [Mil19b]. Further [Lau10, XH15, XLL19]. future [AMB10, BABF19, Fre14, MA11, OW18, TW17]. fuzzy [Bon12]. FWER [MGSL11].

G [Bri15]. Galton [BBP18, Che13a, He16]. gambler [AH12a, Kat13]. gambling [Zaj14]. game [AH13a, AH12a, AMR19, DdRS^{+}11, Gut14, Nak17]. game-theoretic [DdRS^{+}11]. games [BCD19, GML15, Nak15, Wei19]. Gamma [HP14, PS16a, SF12, BLS14, BG19, CO13, DD10, DN14b, Dur17, Fur08, HS12a, JS16, KT10, KM13d, LMR12, MU10, MM13b, MA14b, PSS13, PRD13, Tak12, Wes15, XS18, Zha10, ZB11, vWvZ17, GB19b, VGE15b, WR18]. gamma-Gaussian [KM13d]. gamma-Gompertz [BG19]. gap [MW18, SA12b]. gaps [BPW18]. GARCH [Arv19, BL14a, CPH12, HL11, HH16, HBPC10, Hwa19, Kha14, KH18, Lee12b, LLW14, Tao19]. Gauss [KLY14]. Gaussian [AP13, AD12, Arv19, ASVY14, BGT15, BT16, Bal11, Ben16, Ben18, CD18a, CYW12, CSS14, CG17, Che14c, CK14, Cho19, DH19, DHJT14, DR11, DMWX15, DDZ15, EF16, FCU11, GFA10, GP17, Gri11, HL11, HT13a, HW13a, HPW15, HZW16, IK14, JW17, JSA12, KST18, KM13d, KKKP17, KMV11, Lan18, LPS12, LL12c, LLH15, LWY17, LX10b, Mah18, Mat17, MNBO11, Mou19, MYA15, NW13, NC16, PB15b, Pen11, PL17, PTW19, PS16c, PS19b, SAM13, Ser17, Sha13b, SL12, Söh10, SL15, SS10c, Tah12, Tan13a, Tan13b, TT17, Tan15b, Tur14, TW15, Vig12, VGE15a, Vol14, WL17, XWT19, XY17, Yaz15, Zha14b, vWvZ17]. gene [QO15]. general [ABD18, BW19, BMS10, CKO19, CWZ12, Ciu11, Cla14, EJ12, Eks14, Fan16, GZ18, GA13, HH19, HP12b, Hu16, Kat13, KB15, KHN16, LV11a, LS16a, LDZ10, LZZ15b, LY16, LGT15, MW18, Mad15, MK10, Mia14, MNBO11, QX13, RZ12, Rad19, Ren14, Ren15, RM11, Sma15, Sun13b, TZ16, TW17, Tim15, WC15, Wan18, WM18, WCM11, Wu14, XF17, XDWY18, ZYL14, BC11]. generalised [LS15, PT13]. Generalization [AH11, BS15a, Bar11, BRBB14, BRBB23, Bud14, DER15, GR17, KS16a, Kor15a, PL15a, XLY17, XZZ19]. Generalizations [Had19, JHW18, Den15, Fra11, Lo17, Tum15]. Generalized [CXZ15, Chr12, DPP16, Ery12, EGX16, GWK17, GJ10, HLB17, HC10b, JVVS10, JB13b, Kak12, Kum15, LI12, LRL19, NM19, Shi19a, Shu18a, TE17, ZX17a, AA11, AM13a, AB11, AM11, Arn11, BT14a, Bao12, BLB16, Bel12, BBS11, BY12, BCMR13, BVV14, BDB^{+}10, BW18, CO14, CO13, CZLT18, CJY14, DL10, DD10, DN14b, DK15, Döh14, Dun16a, FS12, GM10a, GSK12, GML10, GLML12, HV16, Hel19, HJM19, HH16, HBF14, Huh12, IL11, JH14, KA10a, Kay15, KM13c, KKMSA13, KR13, KA11, KT11, KN10, LT17, Li13a, LX10a, LL17d, MA10b, Mar12b, MASR14, NAC13, NS13a, Nak15, OH15, PJ13, RSdB15, SKB15, SL12, Shu16, Shu18b, Smu19, TJD10, Tie13, Vig12, WSW10, WX12, WW14, XZY13, Xu12a, YZ16, Zen19, ZY10b, ZDM17, ZIPP17]. Generalizing [Spa10b, Ber13a]. generated [Bél11, CLEA18, FC10, HP13, HBPC10, Hwa13, KL15, KZZ13, MS11, Yor14, Zen14]. Generating [BH14, Dev12, HP18a, Jon19, ML17, Shm13, STW19, UU11]. Generation [Ale18, BP19c, Mod11, PS19a, SBN16]. generative [LS12b]. generators [FJ12, Fan16, GLM19, Izu13, Rez15, TJD10, TJS13, XF17, ZF13, ZL15, ZR12]. generic [MDR12, Nag18]. Genericity [KL17]. genes [WRvdL11]. genetic [CP13b]. genetics [PYS^{+}18]. genome [Che13b]. genome-wide [Che13b]. genomic [CWI10]. genomics [SK13]. genotyping [HS13a]. Genton [ML16]. Geo [PSC12]. geodesics [Yao13a]. Geometric [AKM19, ABB^{+}17, Ery14, Fit14, KR19, MN16, ABBK15, BRO14a, BS12b, BP19c, BMB16, Dal19, Di 11, HNGS15, Jan18, KW14a, LW15b, MS10b, MW12a, NR12, RS14b, SC15, Wan15, YE10]. geometrically [Mia14]. geometry [AP13, Osȩ12c]. Gerber [GR17, Bra12b, DZD12, LXZ15]. getting [GM16]. GI [PSC12]. GI/Geo/1 [PSC12]. Gibbs [CJT10, Coe15, Fit14, Gir16, Hel19, RHP14, SC15, VB17, WR18]. Gibbs-energy [Gir16]. Gini [Eis15, FGA11, GR11, WZG16]. girl [Pol13]. Girone [You18]. Girsanov [CJY14, QW16, WW18]. given [Khm17, KN17, ST19b, ZWLB17]. Glivenko [Che18d]. GLMs [WYH14]. Global [CD10a, LTX12, KK13b, MV15]. Gluing [Ouy10]. GMANOVA [YS12]. GMM [KZ17]. Goal [SW11, Mak10a]. goals [JS18]. Godambe [HB11]. gold [BK18]. Gompertz [BG19, HS12a]. good [Gre12]. goodness [CCGPW17, DN16, GEV16, RH16, Rob15, Rob19]. goodness-of-fit [CCGPW17, DN16, GEV16, RH16]. Gordon [LW14b]. Gradient [CHW18, Che14d, Lem13]. gradual [Tim15]. Graham [Hil19]. Gram [FQZ18]. Granger [Tri18]. graph [GGR11, NM19]. graphic [LW15a]. graphical [JH11, KKKP17, PF17]. graphs [BS12b, BH19, HNGS15, HBF14, LT14, Pak13, Rao13, YCQY16, CC16]. group [AWL17, ASK18, AA19, CLEA18, GZWW15, Gus12b, Mal18, XD15, YK11]. grouped [Now16, Now19, PT19, Rob19]. grouping [Zho13a]. groups [AMG17, Fis11, McC13, Myr19]. growing [LLSW11, LV11c]. growth [BBHH10, FJ12, Fan16, Izu13, Nie19, XF17, ZF13, ZGK13, ZL15]. Grushin [TU17]. Grushin-type [TU17]. guesswork [Li19]. guide [SIBC18]. Gumbel [DK15, Tom16]. Gut [LQC15, LG15a]. GWAS [CHN14].

Haar [Osȩ14b]. Hadamard [Cha17, Kao14]. Haenszel [BB12, BH11]. Haezendonck [ZZZ13]. half [SK15]. halfspace [BF17, Nag19, PV17, Wan18]. Hamburger [MT17a]. hard [MP16a]. Hardy [HY18, Wal15, Yu11, YY14]. Harmonic [LLP19, DLM15, MW18, dlCJB19]. Harnack [LY18a, Wan11a, Zha13a, ZC13]. Hartman [LZ16a]. Hastings [HL16]. Hausdorff [KSW15]. Hausman [BL12]. having [Huh12, JKD15a, JKD15b]. Hawkes [Seo15, Seo17, Seo19, Tak19, Yao18, Zhu13a]. Hazard [AM13b, LL19b, AvHZ17, BZV11, BT14c, CCSC11, DL13a, HP14, LDZ10, LQZ13, MM18, MM13b, Pac18, SS14, WYC15, Wan15, ZZ11]. hazards [ASK18, DE13b, HL13b, LZYW19, QCZ15, SA12b, sS15a, TR12, WK11a, ZZ15d]. head [Nov17]. health [Sha18]. healthcare [Sha18]. Heat [RZ12, BC14, BH10a, BH18, Bou18, CHW18, Li17b, Pac16, Pac17, Rad19, Son12a, YYS16, YY18, aZLW18]. heavy [Arv17, BS11, BP19c, BGHS11, BVV14, DS16, DZF18, FMA16, GF11, GGV14, Hwa19, LQ19, LW12, NA13, PY17, SZ13a, ST13a, ST10, Tsi12, YLS14, YIS15, Zha14a, Bra16]. heavy-tailed [BP19c, BVV14, Hwa19, LQ19, LW12, YLS14, YIS15]. Hedging [SS17, Chi13, Wan19]. Hellinger [BMP10a]. help [TE12b]. heredity [WST14]. Hermite [JW16, CXZ15, Kak16, PT10, Su10]. Hessian [CZ17a]. Heston [AE19]. heterogeneity [Han10, Ruk12, SL12, SQY18]. Heterogeneous [WT18, AKP18, BNB16, DZZ13, El 13, FZ13, FZB16, HS15, KC16, LCX12, LL15, LL19b, MKK17, ZB10, ZB11]. Heteroscedastic [LBG11, BRZ10, Cha10, EP17, LZLW16, MMPW16, WHB10, WZ17, ZW18a, dH15]. heteroscedasticity [CSK13, KKMSA13, MGG11, SS13b, ZCZ18]. heteroskedastic [Cho16b]. heteroskedasticity [LW17b, MdJ19]. Heyde [XH12]. Hidden [Far11, ELS13, Lem11b, Lem11a, Liu12a, PJ13, SHBHD11]. Hierarchical [FP12, GBR12, Rub15]. High [JL16, Mod11, AH11, AZ17, Bes14, Ciu15, FS15, GGS12, GZWW15, He14, HZW16, KM11, Mao14, Mao15b, Mao16, NHBG19, Par17b, PL19, PQW14, PS19b, Sab18, SK13, Wan14a, Xue17, Xue18, Xue19, YJLL16, ZX17a, ZCZ18, ZGK13]. High-dimensional [Mod11, Ciu15, FS15, Mao14, Mao16, NHBG19, PL19, Sab18, Xue17, Xue18, Xue19, YJLL16]. high-frequency [KM11]. high-level [SK13]. Higher [KST18, LTvdVR11, TS17, Wei10, ZL16b, CS14a, D’O11b, DHL14, Gei17, HP12a, IL11, Jal16, Kak16, Kau14, MWGT17, PRD13]. Higher-order [Wei10, ZL16b, D’O11b, Gei17, HP12a, Kak16]. Highest [CM10, Liu14]. highly [Men18]. Hilbert [ÁLBRM16, BH12, BH17c, Bud14, HT15, KO17, Kin10, Ko18, Osȩ12b, Ose19, PK11, Par17a, Rao10, SB19]. Hilbert-space-valued [Bud14, Rao10]. Hilbertian [HZ16b]. Hilliard [JWW14]. histogram [KDW15]. Hitting [Gzy19, HL19, NC16, Zho13b, AS15, Bul14, CGLN17, CL16, Cho18, Dal19, Hof13, Iso15, JS15, Lee12a, LT14, PR12, Rao13, SWW15]. HIV [ZHL13, Zho15]. HNBUE [GM17b]. hoc [Llo12]. Hoeffding [Bor14, Bor16, CL19b, Lo17, Mia14, YJ12, Zhe17]. Hölder [BJQS16, ASVY14, BKY19, Fro17, Li17b, LT19a, NL19, Pen11]. holding [Pen14]. holes [SH17]. Hollander [AB11, AM11, GM17a]. homogeneity [CYW12, GMA12, HH19, LX10a, SAM13]. homogeneous [CP13b, Cui14, DM14, FLRS13, GB19a, LST17, MXZZ18, MFL15, MN16, Mis14, SBA12, TT17, ZZ19b]. homophily [SQY18]. Hong [WS15]. Hood [AMR19]. Hopfield [LD17a, Zha13b]. horizon [Coh10, LL14, PD12, Wei19]. Hörmander [Tah14]. Hougaard [Gri11]. hub [KP17]. hull [Dav12b]. humidity [FFM18]. Hurst [CP17b, DO11a, HB18, JV10, KP15a, KS16b, KS17c]. Hybrid [KKP15, SKK10, KS15b, LLK^{+}18, Lka19, SD15, YHMM15, vBC19]. hyperbolic [DOT14, FQZ18, FS12, Hir11, JH14, KR13]. hypercube [CYLL19, MDPP16, RS14a, WZYL17, XDWY18, YCL14]. hypergraphs [HL19]. hyperplanar [AS15]. hyperrectangles [Tzi17]. hypersphere [Gat13, Ski10]. hyperspheres [ZZ16]. Hyperspherical [PW15]. Hypo [Rat15]. Hypo-exponential [Rat15]. hypotheses [GB16, GB19a, LDD15, PSS13]. Hypothesis [BMP10a, XFC17, CSK13, MW14b, RSdB15, SK13, SKJ15, Sma15].

i.i.d [AA13b, CC10a, CS14a, MV10, QC14, Sun13b, XY12, dMO18]. I.I.D. [ABD14, SD14]. ICA [LNI16]. identical [Il’18, KS16a, Xia19a]. identically [DS16, Lin17b, LMS17, RS10b]. Identifiability [Din10, NK17, ORO15, WYH14]. Identification [BM10b, BM11, PKK14, SS14, WS13, AC15, Lia12b, Lu18, Mal18, TMN16, YY16]. identified [BFS16]. identify [Nos19]. Identifying [HZ16b, HH13b, LZ19, WST14]. identities [NS13b]. identity [BKY19, FS16, Muk10, YY13]. idiosyncratic [Bre14a]. IFR [BS15f, BS18c]. II [BK19a, BKL19, FXT12, HTA10, MB14, PN12, PP15, SK15]. iid [Vas18, CD10b, WN10c]. III [BB11c]. illness [AdUÁMM11]. illness-death [AdUÁMM11]. illustration [BCI14]. illustrations [Muk10]. IM [JLJ16]. IM-based [JLJ16]. image [Dav17, Jov14]. images [DMWX15, KSW15]. imaging [Sab17, Sak15]. immigration [BBP18, Chu14, GZ14, HMZ11, KR12, LM17, Ma14a, Rah11, Sun13a, WL13, Xu14]. impact [GL19]. impatient [KK18]. implementation [ZDD18]. implications [SHL15]. implicit [Alf13]. Importance [NLF11, DRZ19]. improve [dB13]. Improved [ABC10, Kak11a, Kak11b, MMPW16, MW12b, Sas13a, UV13, XS18, CM16, CHN14, KS10, KMS15, SS15c, Wes13, ZX17a]. improvement [Wal17]. Improving [BFS16, MNWA14, CMS19, HS11b, Han12a]. impulse [KH18]. impulses [LTX12]. Impulsive [Che10c, CvGL18, CD10a, LF17, LXZ15, PJ10, WSMG18, YD12]. Impulsive-integral [Che10c]. Imputation [HYWT17, KP15c, QCZ15]. imputed [KS15c]. IMRL [LP17, Nav18, SD16]. inaccuracy [TT15]. inactivity [GAB16, ST19a, ZY10a, ZFZ13]. INAR [NR12, NP16, WS16]. INARCH [Wei10, WS16]. incidence [ED14, LO13, SNS10, Sha14]. inclusion [BB15]. income [BPRS13, ZG19]. incomplete [AG13, GV16, JVVS10, JKP19, KPK15, LW16, Mal18, MS18b, Nas12, PP18a, PPM18, RR10a, Tan15a, WQD15]. Inconsistent [KS15b, Zha17b]. incorporated [Ohy13]. Incorporating [BB11c]. increases [LNC12]. increasing [CL16, DWW17, LL17b, Now16, PYK15, SKJ15, VT10, Wój13a, YCQY16, Zha17b]. increment [BB15, Vii16]. incremental [SR12a]. increments [Cri18, Cri19, GS10, Pen11, WYY13, YLS14, ZH14]. Independence [Zha19a, AV13b, CQ18, CC10b, Ejs16, FW17, GW12c, Il’18, KS16a, KKKP17, Lop10, Mao14, Mao15b, MP15b, NÚF12, Ost18, PS16b, Pap16b, QW16, RL14, Vir16, YKY17, ZPZ18]. independent [BS12a, BV10, CX18, CYH16, Cri18, Cri19, GL13, GB19a, Had13, Har16, HX14, Ilm13, Kru17, Li13b, LMS17, MNO15, MW12a, MM13b, Myr19, Nak16, PL19, PS18a, Pin15b, QG13, Roz12, Roz16, Ruz14, SBA12, Sre10, TV17b, Ton16, Ton17, WW18, Xia19a, YYC17, Zhe17, dC16]. independently [CFS12]. indeterminate [OS10]. index [Akr16, Arn12, ALS11, BMN15, BMN16, BF10, Cas10, FGA11, FZW12, GGO15, GHR16, GHR17, Has10, HB18, Ili12, KS16b, KS17c, LLL13, LZLW16, LP10b, LQ19, LNI16, LX12, MJH18, MNP16, ROSL17, Ser18, SJ14, WW13a, WL14a, WZG16, WSU11, YGL14, YXL14]. indexed [Dem10, DYB11, GW15, GZ15, GW16, Gut14, HB10, Ser17, SY10, WWY12]. indicator [OQ10, QCZ15, SB10, Suz18, Suz20]. indicators [BMS16, LW15a, Zha16]. indices [CH11a, MDN18, ZLT11]. Individual [ZIPP17, Pru19, WQD15]. Individual-specific [ZIPP17]. induced [BP19c, MdJ19]. inductive [Wan12c]. Inequalities [BNM13a, Osȩ12b, Osȩ14a, AA13a, BV15b, BH18, CH12, DZ11, Din14, DZF18, DLM15, EAB10, EAB11, EK10, Fan15a, Fan17, Fer14, Fro17, GSK12, GM16, GI14, HWYW12, JP10, JHW18, Jov19, LVY15, LYW11, LY18a, LL12c, LW14b, Mia14, NS13b, Osȩ10c, Osȩ12a, Osȩ12c, Osȩ14b, Osȩ17b, PS18a, PL15a, Pin13, Rao12, RS12a, Ruz14, Tan15b, Ush11b, WHYL10, Wan11a, WHRY11, WW13b, WD12, XZ16, YWJ14]. Inequality [TSK13, Aga15, AK10, Arn12, BBM19, BS19a, BBHV13, BSWS13, BO19, Bud14, Che10c, Che14a, CG17, CW19, CL19b, CD12, CRB19, DS11, Din15b, DZ17, FP11b, FMPV12, Fro17, GR17, Had11, HRB19, Kre18, LZW11, Li13c, Lin19b, Liu12b, LWY17, LW14b, Min17, Nav14, Oga17, Oh14, Olk14, Onz11, Osȩ11, Osȩ13b, Osȩ14c, Osȩ15b, Osȩ17a, Pei15, Qiu14, Rao10, RSdB15, Ser17, Sze12, Sze15, Tru10, Wal15, Wal17, YJ12, Zha13a, ZW15a, ZG19, Zha19b, ZH12b, ZC13]. infection [Xue18]. infectious [Cla14]. infectiousness [Chi12b]. Inference [CL10, GdL18, Kha14, LZ14b, TLF12, Vet14, AHN10, AT15a, AHM19, BBGMPG15, BW19, BG19, Bow16, CG18, CD18b, DD10, DN14b, DP10, EP17, FMA16, FZW12, GO12, GXH17, HS11b, HZJ^{+}10, HB11, KJS15, KS15c, Kur19, LTvdVR11, LZYW19, Lu16a, LD16b, LD17b, MJH18, MB16, MOD18, OS18, PB16, PB15b, Rah17, Roh13, RG18, SK15, She13, SL13b, Tsi17, Wes13, Won13, ZW18b, Zha11b, ZZM15, vL16]. inferences [LLH15, Lu16b, MKK17, ZLZ16]. inferential [ZB15]. inferiority [JLJ16, Sha14]. inferring [McE12]. infill [Kur19, WL14b]. infimum [SD14]. infinite [AP18, AV14, AL16, BT14b, BS14b, CK16b, Coh10, FJ10, Gov15, Gra11, HS16a, HR15, KKT14, LLHW10, LLH11, LL14, Li15a, LMLW15, Nak16, Res11, RSR16, SWZ19, Sta16, TE12a, WH11, ZF13]. infinite-dimensional [AV14, Li15a]. infinitely [BLL10, Bra11b, DM17b, FS17a, IMU10, Kha19, Nak13a, SW19, Tra14]. infinitesimal [Bre12a]. infinity [LNC12, Roz10]. inflated [BSL13, BMB16, DM17a, PNBW15]. inflated-parameter [BSL13, BMB16].

Influence [Ema15, HW13b, LNC12]. influenced [YG13, Yan17a]. influential [Zho15]. information [AO12, BDS19, CFS12, CTN17, CF18, EPSU16, Ern17, FK10, HM10a, IV19, JT14, PRS12, PN12, PP15, PNC15, PP14b, RS12b, RR10b, Sha17, Son10a, TL12, TE14, Wal16, WWY12, YAT16a, ZDM17]. information-based [CF18]. informative [Bow18, CM11b, Din10, SHBHD11]. Ingersoll [FS18]. inhomogeneous [BH19, BS15e, Che14d, CRB19, Sta16, Vri16, WJY11, ZK15, ZKSK18]. initial [Gou15, Hwa13, LB17, Wan17a]. initially [KM12b]. injections [DZ19b]. inner [Zen15]. inner-product [Zen15]. innovation [CW11]. innovations [Arv17, HL11, HBPC10, Mar16, PS15, Zha11a]. input [XTT18]. insensitive [PK11, Par17a]. inspection [Gus12a]. instability [LCW17]. instant [But15]. instrumental [CFS12, LR12b]. insurance [CR16, LH14b]. insurer [ZYG17]. Insuring [DdRS^{+}11]. Integer [KS13, Iso15, MK14, NR12]. integer-valued [NR12]. integers [Kha19]. Integrability [PL18, Cha15, HR11, LV13]. integrable [BM18, FL10a]. Integral [Pog16, Aga15, AP13, Alf15, BDP14, BL15, Che10c, DER15, FCU11, GM17b, JV10, Jur13, Jur18, KWL15, LV11b, LZ11, OP13, Pak18, PT10, Sak12, Sak15, SV19, VD18, WW16]. integrals [AY17, ANRW15, Jur14, Kak16, Kam17, Lan11, MA14b, PS17, Tap10, Xu18, Yas19, Zin13]. Integrated [Nag17a, BL14a, CM16, DN16, GM14b, GGO15, PL17, Wu14, ZPK14, Tao19]. Integration [Tak19, Bry14, Pin12, SC12b, SG15b]. integrative [CWI10]. integro [BB17, Wan17a]. integro-differential [Wan17a]. integro-local [BB17]. integrodifferential [DDE17]. intensity [BNM13b, DMPX17, DYW14, Elh14, PH10a, Poi19]. inter [GSW11]. inter-arrival [GSW11]. interacting [BW11, JWW14]. interaction [BCD19]. interactions [FF12, Jac10, PVY17, WST14, WCM11]. interarrival [LP17, Spa10a]. interdependence [LW14a]. interest [BL17, DMS13, Dun16a, GL13, GW12a, LGW12, MWH17, PH10b, Won13, YW10, ZPK14]. interesting [JHF15]. interests [WYL10]. interference [Fil12, RH15, VT11]. intermediate [GS11, HCT16, LGP10]. Intermittency [BN16, GLT18, BC14]. internal [SX13]. international [Tsi12]. interpoint [Mod14, Mod16]. interpolation [KK13b]. interpretation [WX19b]. intersect [TV17b]. intersection [CY11, Liu11, OSW18, SY17, YYC17]. intersection-union [Liu11]. Interval [FPZ13, LZT13, WYC15, AD12, BB16, BPS17, CM11b, FJ10, HLR15, HM13, Kab11, Kab13, KRB13, Lot19, MY11, NdBC12, PSS13, sS15a, SV15, Sun17, TL15, Uno13, VR12, Wan12a, WL11, WZG16, Wan17b, ZWLB17]. interval-censored [sS15a, WL11]. interval-observed [CM11b]. interval-valued [SV15, Sun17]. intervals [AMB10, BABF19, BCI14, FK12, Fre14, GEC17, KS10, KM18, KM13a, KKMSA13, LSS15, MMPW16, Par17b, RH15, RS12c, SS13b, SV15, Sub12, SJ14, TM16, Thu14, Wan13a, ZF13, Zha14b]. intraclass [ZW18b]. Intrinsic [HZR16, HZRS19, SY19b]. Introducing [Sug19]. Intuitive [MO14, OV15].

Invariance [Mar12b, PD11, Jir13, ZZ15c]. Invariant [DJM11, Jur13, BS18a, Gei17, INO10, LD16a, Liu14, PV17, ZC19]. Inverse [DHL14, KV15, TN17, VGE15a, Abu12a, Abu12b, Abu13, Arn11, AT15b, CP13a, CYW12, CSS14, Han12a, IK14, JSA12, KMV11, LPS12, Li15a, LCW17, LLH15, LM15b, LA13, Mse18, NTT18, Oga18, PZ14, SAM13, Seo19, sS15b, SWL10, SL12, STW19, WHYL10, Zha14b, ZIPP17, Zha12, vWvZ17]. inverses [Sma15]. inversion [Mna11, ML13, WY18]. Invertibility [CD12]. Inverting [Jon19]. invested [Kin12]. investigation [HP15]. investment [BCM14, ZW12]. investments [FN17]. investors [WWY12]. involving [BNM13a, Ery15, Ger13, Ger18, Kak16, MA14b, WM18]. irreducibility [BD17]. irregular [KKNH18]. irrespective [RS10b]. Ising [Hel19, Kar11, PVY17, SC15]. isobars [BDB19]. isolated [GGR11]. isotonic [GJ10, She11, SD15, SZ13b]. Isotonicity [Bel12]. isotropic [De 17]. isotropy [EFP19]. Isserlis [MNBO11, Vig12]. issues [CKK18, FFM18, Sak12, ZB15]. Itô [FCU11, HT14, LSS13b, PS17, Zin13]. items [Gus12b]. Iterated [NP16, hCyP15, Dur13, IJB17, LL13, Lin13, Liu13, LX10b, LG15b, MLF13, McK17, MN11, MN13, MC15, Ngu16, Vas18, Wan14c, XY12, XZY13, Zha17c]. iterates [Gna12, Tru13]. iterative [CMW18, HM10b, Ref14, Tsi17].

Jackknife [CK13, WZG16, ZLZ16, CY15]. Jacobi [FS18, ZG10]. James [GG12]. Jeffreys [Rod13, RL14]. Jensen [Lin19b]. Jirina [LL17d]. Jittered [PRS18]. job [Men18]. John [LZP17, YWJ14]. Joint [CX18, Ery11b, LTR13, PS15, AKP18, BE17, Che14c, CR13, FCP18, FJW15, Kri18, LZ14a, McC12, PP19, RS10b, VT10, WWPY18, WZ10, WN15, XZ13]. jointly [Cai17]. Journeys [DH18]. joy [Men18]. jump [De 16b, DMPX17, GS17a, KK15, KW14b, KOR15b, Leh15, LWD13, LR12a, MOD18, MN16, Rat13, SL13b, SWH12, WZT12, WWX17, WWZ18, XY13, YG13, Yan17a, ZX17b, ZW15b, ZBL19, dSF12]. jump-diffusion [DMPX17, GS17a, LWD13, SL13b, SWH12, WZT12, XY13]. jump-telegraph [LR12a, Rat13]. jump-type [dSF12]. jumping [Pen14]. jumps [BCD19, BL11, BH10b, CHM15, CvGL18, CYS11, De 11, Fro17, Hui10, Lia13, LZ11, MX17a, QW16, Sas13b, Shi19a, TH14, THC17, Wan16, Wei10, Yan19b, ZY10b, ZYL14, ZWB17].

K-distributed [Ma13]. Kac [Bou17, KOR15b, LT19b, LR12a, Mak10b]. Kantorovich [FP14]. Kaplan [Ane17]. Karamata [PLH13]. Karhunen [ALL12, Ai16, Li15a, LHM14]. Kato [Gov15]. Katz [CS14a, GS11, Spa10b]. Kelly [ED17]. Kemperman [Ruz14]. Kendall [DM17a, Lo17, Vin11, vDLMW19]. Kernel [AM19b, BMN16, SS11b, YH17, AFJ11, Bou18, BS18b, BS19c, CDG15, CRC10, CL19a, CSG11, Che18d, Cou19, DFH18, DL13a, DL13b, DG13, EL18, Fak10, Gho14, GGQ17, HP12b, HS19, HLCH19, IK14, JW17, KM12a, KZ10, KV16, Kre18, LL19a, LWL16, LM15a, LPS15, MS12a, MS12b, MK16, MNWA14, PK11, Par17a, QCZ15, Sak17, Son10b, TS17, Wal10, You17, Zen15, Zen19, Zha10, ZJZ13, ZM16b, ZKG10, ZB15]. kernel-assisted [QCZ15]. kernel-type [AFJ11, BS18b, MS12a]. kernels [Ben18, Bow13, CHW18, Duo15, FP17, GM14b, HP12a, HL16, MMF17, Nog13, ST13a, XLL19]. Khasminskii [WH11]. Khasminskii-type [WH11]. Khintchine [Jon19, PS18a]. Kiefer [WD12, FP11b, Pro13]. Kimberling [Res11]. kind [DZ11, KN15, PTT19, SS18]. kinds [LZ10]. King [BS15a]. kink [Wis11]. knowledge [BCNM15]. known [DDZ15, FP13, LP10b, SV10]. Kolmogorov [AK10, EAB11, Fre16, KS17a, SL10, YH15]. Krasulina [Che19a]. kriging [HZR16]. Kronecker [HLM16]. Kullback [CCGPW17, HM10a, IV19, NSL11, PRS12, PP15, SSN16]. Kumaraswamy [Mam15]. Kummer [HV16, Kou12a, Wes15]. Kuramoto [WWW14, WWX17]. kurtosis [KM13c, Oga17].

L [LS12a]. labels [MV10, VS10]. lack [JN10, SD10b, SY11]. lack-of-fit [SD10b, SY11]. LAD [LLHW10, LLH11, ZL14]. ladles [BK18]. lag [CM17, HK19]. Laguerre [JW16]. Lai [LS12a]. lambda [LBH11]. Lambert [Pak18, WWD15]. Lamperti [Ma15]. Langevin [Öns13, Vii16, XSL^{+}14]. Laplace [MM14c, Abd11, BLB16, BS18a, BW18, CL11, GWK17, GS17a, GEV16, HA13, HJM19, HBH10, Hür13, LBS17, Mna11, MM13c, NMUJ16, RH16, Roz10, SCS14, Sri15, YY11a, ZB15]. Laplacian [Din15a]. Large [AMMO13, BC11, BM13a, BR17a, BS19c, CHM15, DM12a, DS14, DYB11, Fan15b, FS17b, For11, FKZ15, GM14a, GW15, GZ15, GW16, GM18b, GMP19, HT13b, Hir11, JCM15, KS19, Li19, LLY19, Mac11, MP15a, Mac16, MP17, Nas12, QHS13, RS10b, ST13b, ST10, SZ18, Sun13b, Ter15, TE12a, Tor13, WHRY11, Wan16, YH14, ZG19, ZG10, Zhu14, Zhu15, AD11, BHJ10, BT14b, BZ13, CQT12, CS16c, CS16b, CZLT18, CS19, DL13a, Din15a, DT10, FC15, Fen17, For14a, Gas10, Gee15, HCW13, HT15, Hor16, Hu16, Hua16, KM17, KD16, Kor15a, Kre18, Li11b, LLR17, LLR18, LLP19, LY16, Lu11, Lu12, MKA19, MVI11, Mou19, Nak16, PTW10, QN15, QSS17, SXM16, Shi19b, SS10c, Sun13a, VS13, WN10a, WZW17, YS12, YLS10, Yao14, YH15, Zen14, Zen15, Zen19, Zha17a, ZL18, Zha19a, ZZ13b, ZLC18, ZZ19b]. large [dS18a, dMO18]. large-dimensional [WZW17, Yao14]. large-maturity [For14a]. Large-time [For11]. larger [CP11]. largest [BH17b]. Lasso [XXY17, RZ10, dC13, WW14, WST14]. last [FJW15]. latency [MP17]. Latent [Jon17, BBM18, DCR19, Sha12c, SdOG12, vDLMW19]. Latin [CYLL19, MDPP16, RS14a, WZYL17, XDWY18, YCL14]. lattice [AIM19, Gao19, HWB10, Iso15, Roz12, ZZ12b, ZZ12a]. lattice-valued [ZZ12b]. lattices [FSNÚF11]. Law [AMMO13, DYB11, Fen19, Nas12, QHS13, RS10b, ST13b, Sun13b, Ter15, Tor13, WHRY11, BSO10, Bis10, BM12, BMS16, hCyP15, CH11a, CYH16, CS16c, CS19, CP19, Cui14, FG12, Hor16, Hu16, IJB17, JV10, JDK^{+}11, KD16, Kor15a, LMP18, LL13, LLR18, Liu13, LQC15, LG15a, Löw18, LX10b, MLF13, Mar12b, MN11, MN13, MC15, Pak18, Pap16a, SV17, SL10, Vas18, Wój13a, XY12, XZY13, YH17, YY13, YH15, Zha17c, Zha17a, ZL18, dML12, dMO18, TE12a]. Laws [Wan14c, AP18, BHJ10, BR17b, CQT12, CM11a, CS16b, DJ14, DT10, FQZ18, GS17b, HW13a, HT15, HBF14, KM17, LLR17, Lin17a, MKA19, Nak16, Nak13b, Nol18, Pak13, PS16b, Pap16b, QN15, QSS17, Sze11, TK10, XYZ16, dS18a]. lay [Won18]. layer [Bra12b, DZD12, JYL12]. lead [HK19]. lead-lag [HK19]. leader [KMR13, KW14a]. leaders [AEK19]. leadership [Ste15]. leaf [AI17]. learning [Bal11, GMM10, TR18, WRvdL11]. Least [KLM11, LL10, Ma10a, Wu14, AAEH11, AT15a, BL10c, DMPX17, ES13, He18, LSS13a, LM18, Mao15a, SZW10, WZ16, YH14, ZK10]. least-squares [DMPX17]. leave [Lot19]. Lebesgue [Kam17, NS13a]. lecture [BHJ18]. Lee [EQ15a, Els16]. left [AEFO18, ADP12, Che19b, FJ12, Fil12, Kol17, LL17b, PCGV14, QB14, sS10b, sS15a, XSZ17, ZBZ11, ZJL13, Fin11]. left- [Fil12]. left-censoring [LL17b]. left-continuous [FJ12]. left-Lipschitz [AEFO18]. Left-Spherically [Fin11]. left-truncated [ADP12, Che19b, PCGV14, sS15a, ZJL13]. Leibler [CCGPW17, IV19, NSL11, PRS12, PP15, SSN16]. lemma [BD15, Fro12, LSS13b, Lo17, Shu18c, Ste14, JMDW15]. lemmas [Dav13]. Length [MPA15, AMP16b, Aza13, BML14, FGA11, GB19b, Jon19, KM13a, Kao13, LZ14b, MZZ15, Nov17, PH10a, SM16, SCZ15, SMZ18, Shu19a, ST19b, WCD19, ZGZ18]. length-biased [FGA11, GB19b, LZ14b, MZZ15, SCZ15, SMZ18, WCD19]. length-intensity [PH10a]. Leone [Mir14]. LePage [DD12]. Lépingle [Qiu14]. Lett [Abu12a, Bai13, BSCQ18, BRBB23, DR15, DFL13, DFT13, Duc10, Kak11a, Kim13, MN13, MM14c, RA13b, Ren15, Sko20, SF12, Suz20, Ton17]. Lett. [GHR17]. Letters [Bro11, Sug19]. level [BLL14, CDS17, CKF17, CR16, EQ14, EQ15a, EQ15c, GB19a, Jac10, Jac13, JKD15a, JKD15b, LZZ15b, MK14, MD16, Mat12, MdJ19, OQ10, PP14b, SK13, Sha13b, Sin19, TL15, WA17, WL14b, Woo15, YSLL10, ZWHQ15]. levels [LCLQ16]. leverage [Bre14a]. Levin [LZ17b]. Lévy [Nie19, Suz20, Abd15, ADM15, BN16, Bra11b, CCS17, CL16, CM11c, CR16, CP17c, DY10, DZ19b, FRZ10, FCU11, FL10b, FS18, GKW19, Gat15, GMP19, Gri11, GLM19, IM15, Jur13, KM11, KM13b, KZZ13, LZ14a, LZZ15a, LT19a, Liu13, Ma10a, MW14a, Mic11, Ren13, Son10a, Son14, Son16, Suz18, VT10, Vet14, Vid17, Vid19, Vri16, Wan11a, Wan14c, Wan14b, WWPY18, WM18, ZDD18, ZLT11, ZWB17, ZYG17, DMST13]. Lévy-driven [Abd15, IM15, Vri16, ZWB17, DMST13]. LIBOR [Ger11]. Lie [AMG17]. life [BCI14, CN16, DK14, GMDP18, KI15, KN10, MKJ15, Nan10, SD16, SKJ15, WC17, WCD19, ZY10a, ZJL13, ZJ15]. life-test [BCI14]. lifetime [BKL19, CF18, Ery16, HTA10, PB16, PLY18, Raq10, ST19a, SSN13]. lifetimes [LP19b, MKJ15, Mar14a, PB13, vBC19]. Lift [DKM19]. light [Bra11b, NP10, PL15b]. light-tailed [NP10]. lightly [VS13]. like [CYW12, Dav13, FQZ18, GM18a, Xu13]. Likelihood [BNB16, BG19, KS15c, LL15, PJ13, WK11a, vL16, AF12, Ana17, BGM19, BB12, BSK19, BH17b, BG15a, BG15b, BY12, Bro14b, BPS17, CO14, CM16, CW16, CZ17b, DBB13, DWWG18, DE13b, DS18b, GT14, HC10a, HHF17, HWM15, HZJ^{+}10, Hua12, IK10a, IL11, JP11, JW12, KLM17, KM11, KLNB14, LGP10, LLHW10, LLH11, LYB13, LQ19, MJH18, MD13, McC12, Mis14, MVI11, OKT13, PSS15, PR15, PB16, Par14, PQW14, Rao18, RSC10, RR10a, RG18, Sch12, SL13a, She13, SK11, SL13b, SHL15, Sri15, Sub12, SJ14, TL12, Tum15, VR12, VZ18, Wan12a, WRvdL11, Wan14a, WP14, WF15, WL15a, WZG16, WZS17, WR10, WT14, XJW15, YL10, YZ14, Yan17b, YWZ14, ZPK14, ZZ12a, ZLZ16, ZZ11, Zha11b, ZZ13b, ZZM15, ZLC18, ZZ19b, ZY11, ZZ15d]. likelihood- [RG18]. Likelihood-based [KS15c]. likelihoods [De 16a]. LIL [GS17b, LZ16a]. Limit [AMR19, ANRW15, BHJ12, BBH14, Bob13, EM16, HW13a, HBF14, KC19, MW11, MS11, MWA16, Nak15, NW10b, Pak13, PS19a, Pan17, Seo15, Seo17, Seo19, Sha13b, Shi12, Ski10, SD14, Ste11, SBN16, Tao19, Ter15, Tyu12, VY12, VW14, WGL11, YR11, Ane17, Arb19, Arv17, Arv19, BGT15, BPW18, CW15, CK16b, Che13a, DL10, DF14, DR02, DR15, Eks14, EJ11, Fro15, Gao19, GFA10, GAS13, GZ14, HSY17, He12, HL13a, HA15, Hwa19, KT19, Kam17, Kat19, KKNH18, LBM14, LL13, LZ16a, LW11, LMLW15, Mas18, McK15, MN11, MN13, MWZ11, MR15b, Mül19, PZ17, Pap16b, Pet11, RS16, Rok15, Röl18, RR16, SY10, Sre10, SS10c, Sze10, Tan13a, Tan13b, Ter17, Ton16, Ton17, TK10, WWS19, Wu11b, WK11b, WC14, XWT19, XW17, Xue17, Xue19, Yan19a, Zha17c]. limited [MV15]. Limiting [BBGH12, HB10, Jas19, Leh15, LLP11, Sch14, Xie13, Zen17, BW15, BLL10, BLS14, BN10, CCFL18, GD18, HJM19, Uem16, VCM14, ZC19]. limits [BT13, CM11a, Li11b, LL17d, McK17, NP16, SW13, YWC10, ZZ13a]. limsup [Pot15]. Lindeberg [DJ14]. Lindley [Cha19a]. line [HA15, XS18]. Linear [AH12b, BB15, CCS16, ZP19, AM17, AP13, AZ17, AD11, BDP10, BCL18, BBGMPG15, BCMR13, BH10a, BDB^{+}10, But15, CTN17, CTW10, Che12b, CC16, CHM16, DNLR17, DLZZ18, DO11a, DSX13, DZX18, Dun19, Eck18, EIW14, FJ12, Fan16, FJW15, GHH^{+}10, Goi12, Gou15, GXH17, HS11b, HP11, HHF17, HB18, HL13b, Hu16, HZJ^{+}10, Hua12, HWZ14, Huh12, Izu13, JBS10, JP16, JH14, Kao13, KST18, KLW11, KZ14, KM19, Ko13, Kri18, LO13, LT17, LNC12, LWX11, LP13, Li15c, LBS17, Lia12a, Lia12b, LHT15, Lop13, LGT15, Luo12, LG13, MA10b, MR15a, MH12, MW11, McK17, MWZ11, Mou19, Oes15, PH10b, PQW14, QO16, Rat14, Rat17, Ren14, Ren15, Rok15, Rub15, RG18, SB19, SY18, SCS14, Sko19, Sko20, Tab18, Tan19, TZW13, TZ16, TZT16, TW17, TK10, WSW10, WS13, WW13a, WW14, WMFW15, WY16]. linear [Wan17a, WZW17, WR18, WL14b, WS10b, XXY17, Xu12b, XFC17, YL10, ZF13, Zha17c, ZW18b, ZC12, ZWT15, dB13]. linear-growth [FJ12, ZF13]. linearity [MR15a]. linearly [Cas10, WC15]. LINEX [Hwa18]. Liouville [Shi19c]. Lipschitz [AEFO18, FJ10, Owo15, PP18b, QW16, Shi19a, SG18, Tah14, TH14, WWR12, XWML19, Yu19]. loads [Tor13]. Local [BCL18, BNSA13, DNLR17, DK11, DN16, Mes10, SZW12, WZT12, WMFW15, YLS10, ZL16a, ZKG10, BRZ10, BB17, CY11, CYW13, DF14, DHL14, FKZ15, Gao19, GAS13, HL13a, Kak16, KM11, KN15, LE19, LBM14, LX18, Li17d, Muk17, OS15, OKT13, PZ17, SY17, TLZ12, VP12, WMZW13, Wan17a, YS10, YYC17, Yan19a, ZG16]. Local- [BNSA13]. Locally [KK19b, KHN12, KKP15, Myr19, NL19, PP14a, Tah14]. location [AB10, BB14a, BKL19, Bow17, DS18b, Gau14, HEM10, KKMSA13, LLC^{+}12, MKK17, MK10, MGG11, SS13b, Tim15, Tyl10, Vig12, Wis11, WN13a, ZPK14]. location-scale [AB10, BKL19, LLC^{+}12]. loci [SKK10, WRvdL11]. Loeve [ALL12, Ai16, Li15a, LHM14]. Log [JP11, KD14, Mu15, BR17b, CSG11, DS11, Dev12, FD18, HK14, Haz11, HJM19, HEM10, HP15, JW12, KN11, LCX12, Roz10, SD10a, SK11, TE12b, ZZ13b, ZZ19b]. log-Birnbaum [LCX12]. log-concave [Dev12, SD10a]. Log-concavity [Mu15, FD18, Haz11]. log-density [HP15]. log-Fréchet [BR17b]. log-Laplace [HJM19]. Log-likelihood [JP11, SK11, ZZ13b, ZZ19b]. log-location-scale [HEM10]. log-normal [HK14, TE12b]. log-Sobolev [DS11]. logarithm [hCyP15, CH11a, CYH16, Fen19, IJB17, LL13, Liu13, LX10b, MLF13, MN11, MN13, MC15, Vas18, Wan14c, XY12, XZY13, Zha17c, dML12]. Logarithmic [Fan15a, Osȩ10a, Pei15, WWD15, Gat15, GM18b, TL19]. logarithmically [KL15]. logistic [BS15a, GS16, KHN12, KO17, LI12, Li13a, LLZ13, ÖA15, Pin14, ROSL17, Rul14, SK15, Tan19, Tie13, WR18, Yeh11, dC16, vSK15]. logit [YWZ14]. loglikelihood [VRR13]. lognormal [LZT13, WF15]. logrank [JdU16]. Long [WWW14, WWX17, WWZ18, AR18, Aza13, BT13, BS14b, Bra16, Buc15, HP11, HS15, LQZ13, LK14, Lu12, Mis13, MIL16, TWM15, WHB10, Wis11, ZZ13a]. long-range [Buc15, Mis13, Wis11, ZZ13a]. long-run [MIL16]. long-tailed [Lu12]. long-term [LQZ13]. longest [LY16, Nov17]. longitudinal [ACW10, CM11b, KJS15, LW16, SHBHD11, Tan15a, TXX16, YL10, YXL14, ZW18c]. look [Ngu16]. Lorentz [HY18]. Lorenz [DP10, FGA11]. loss [BW19, CVA13, DdRS^{+}11, GBR12, Hwa18, KN11, KK18, KMS15, LH14b, PK11, Par17a, VAB15, ZRAN13, ZRN14, ZRN17]. lottery [Pol13]. low [LTvdVR11, Tom16, WN11]. low-dimensional [Tom16]. Lower [ASH18, BB11a, KSW15, Lu11, Lu12, Ser18, YWC10, ZKSK18, Ben18, CLQ12, CC16, Efr10, EQ14, FJ12, GM14c, HCW13, Hil19, LZZ15b, NW10a, PR16, Sas13a, TE17, WYC15, Wis11, Yan12, ZWHQ15]. LPQD [MC15]. LR [Kak11a, Kak11b, NHBG19]. LR/ELR [Kak11a, Kak11b]. LRT [HH19]. LS [MWZ11]. LSL [CC10a]. LSS [CZLT18]. lurking [GBVS11]. Lyapunov [FY16, XL16]. Lynden [GSST12, WW16].

M [Kin12, ML16, BML14, Bri15, Kin12, VCM14]. M/E [BML14, VCM14]. M/G/1 [Bri15]. M/M/1queue [Kin12]. MA [Yab17]. Machine [CGH11, Dun18, Gho14]. Machines [LDLC18]. macrofauna [JDK^{+}11]. macros [SWH^{+}11]. Mahalanobis [GEC17, MM14a]. main [Cha19a, CLYW15, CLWH17, DS15, Jac10, Jac11, Jac13, JKD14, JKD15a, JKD15b, YL19]. main-effect [CLYW15]. Maintaining [TMS11]. maintenance [KMJ10]. majorant [LM18]. Majorization [Bai12, Bai13, DZ17]. making [MKK17]. Malliavin [Suz20, IMPR16, SS13a, Suz18]. Mallows [You18]. Mandelbrot [LLP19]. manifold [Bes14, IKAH17, Pei15]. manifolds [Che14d, TU17]. MANOVA [HH19, ZX12]. Mantel [BB12, BH11]. many [Zha13b]. Mao [CQ18]. map [BHJ10, BHJ12, MHH11, Ngu16, Tom16]. mapping [SKK10]. mappings [IMU10, Jur13, Jur18]. maps [AIM19, Dur13]. Marcenko [YH17, Yao12a]. Marcenko-Pastur [YH17]. Marchenko [HSY17]. Marcinkiewicz [Fer14, Had11, Sze11, Tru10, ZL18]. Marcus [NW10a]. margin [Xu12b]. margin-based [Xu12b]. Marginal [ACW10, HP11, ZHWH12, DHOV13, Jur13, LW15a, Lop10, NS10, PPN10, Roh16, Sha17, TR12]. marginally [WWS19]. marked [Tak19, JMS17]. market [Ger11, ZYG17]. markets [Fen14]. Markov [BSCQ18, AMG17, AR14, ACMM13, BDER14, BS14b, BG15a, BG15b, BBCD^{+}18, BSCQ17, Bre12a, Bre12b, Bre14b, BMS16, CH13, CJ14, Cho18, DMS13, Dal19, DN14a, DM14, DYB11, ELS13, Far11, FC10, FMPV12, Fra15, Gei17, GFH16, GAS13, GSW11, HZZ19, HL13a, HA15, HMS14, JT18, Jar13, Jia16, KS12, Kau14, KL11, KH18, KLY14, KR19, Kov10, KW14b, LP11, Lem11b, Lem11a, Li17c, LM17, LWD13, Liu12a, LMLW15, LXL19, LPS15, LR13b, MZ17, MXZZ18, Mia14, MN16, Nog13, PZ17, PJ13, Rao18, SY10, SG14, SHK14, SHBHD11, Sta12, Ste14, SWH12, TLZ12, Urb12, Woj13b, WK10b, WC18, Xu13, XW16, Yas13, ZK14, ZK15, ZKSK18, ZH12a, ZW15a, Zho13b]. Markov-modulated [HMS14, SWH12, XW16]. Markov-switching [BG15a, BG15b]. Markovian [Cho19, AE19, AMP16b, CX18, FY16, For14b, LR13a, QLH^{+}16, Rév12, Seo19, ZYL14]. Marshall [BD15, BMS16, MS10a, ND12]. Martingale [BdS13, HY18, Lee19, WK11b, Yu11, YY14, AL16, Bou16, CSY12, DT10, ET19, HL12, Jov19, KN15, NS13a, Osȩ10a, Osȩ12a, Osȩ12b, Osȩ14b, Osȩ14c, Osȩ17b, Osȩ17a, PL15a, Qiu14, Röl18, STD12, ST13b, WYY14, YG13, YR11, cZgL12]. martingales [Bór18, BO19, Cas10, JP10, LZW11, Li17d, LZP17, NS13b, Osȩ10b, Osȩ10c, Osȩ11, Osȩ12c, Osȩ13a, Osȩ13b, Osȩ14a, Osȩ15b, Sas13b, Sau12, VP12, YWJ14, Yu11, YY14, ZZ12b]. Maruyama [LT19a, NT17, NL19]. masking [YQW10]. mass [BK12, MvdBV^{+}11, SLM19]. mass-nonlocal [SLM19]. matched [JLJ16, LD17b]. matched-pair [LD17b]. matched-pairs [JLJ16]. Matching [CLG13, CP17a, LN15, WL15b, Yat16b]. Matérn [Gir16, IL17]. mathematical [LOKB11]. matrices [AS18, Arb19, Bao12, BBGH12, BH14, BGHS11, CQ17, Che18a, Din15a, GS13a, HSY17, HR19b, Hil14, HLCH19, JP17, Kao14, Kou12a, LLSW11, LP13, LCW17, LW17a, Löw18, Mao16, Olk14, PW15, Red17, RM11, Sep18, Ser17, Uem16, Vol14, WWW19, WN10c, Xie13, YH17, Yor14, Zen14, Zen15, Zen17, Zen19]. matricial [Duc04, Duc10]. matrix [BW15, BDER14, CCS17, CD12, DTV16, GM19, Ilm13, Jia13, KST18, LK13, Mad15, Mat17, Mil19b, NHN^{+}11, NHBG19, Pac16, Pin14, RSC10, SKRT16, Tsu10, XH15, ZRN17, vW17]. matrix-exponential [CCS17]. matrix-variate [ZRN17]. Matsumoto [Kou12a, Wes15]. maturity [For14a]. MAVE [RW17]. Max [JGW14, BR17b, Hof13, KR19, LT10, Oes15, PS16b, Res11, Rob13, SW19, WS10a, YHW11]. max-domains [YHW11]. max-infinite [Res11]. max-infinitely-divisible [SW19]. max-linear [Oes15]. max-stable [Hof13, KR19, Oes15, PS16b, Rob13, WS10a]. Max-sum [JGW14, LT10]. Maxima [HPW15, ANV13, ABB^{+}17, BS10, BMNZ18, Fer12, Kri18, Lin19a, Nak17, PP18a, PP19, Rob10, Tan13a, TT17, WLP17, WL17, YC16]. Maximal [BD17, CH12, NS13a, WHRY11, AMG17, ABBK15, BOR17, EAB10, He16, JHW18, Kao13, LMS17, MP15b, NW13, Osȩ10b, Osȩ11, Osȩ15b, Osȩ17b, Osȩ17a, PV17, Rao12, Sch14, SZ18, Sze15, Yak15]. Maximin [MP13]. maximization [Bha17, DCR19]. maximizing [AN15, Fer18, MWH17]. Maximum [ACMM13, BCGFR13, BPS17, Che12b, DBB13, JB13a, KLM17, RR10a, Yan17b, ZPK14, Ana17, AKP18, BGM19, BSK19, BH17b, BM11, BG15a, BG15b, BY12, BK19b, Bro14b, CVA13, CR13, CP19, CW16, CZ17b, GT14, HC10a, HK14, HWM15, HCP19, IP14b, KLNB14, Kou12b, LR13a, LYB13, LR12b, Mak11, MPA12, Mes14, Mis14, Nak13b, PSS15, Par14, Rao18, RSC10, RR16, Sch12, SL13a, SHL15, Sre10, Tum15, VAB15, WRvdL11, Wan14d, WP14, WF15, WN13a, WN14b, WN15, XJW15, Xia19c, XZ16, XTT18, Yan15b, YLS14, YWZ14, YK11, ZZ12a, ZLC18, ZWLB17]. maximums [STD12]. Maxwell [RBSB16]. Maxwell-Binomial [RBSB16]. may [BT16]. MBP [LW13]. McKean [WX19b]. MCMC [AG11, Bes14, JH14, RS15b, XKBG15]. MCMC-based [XKBG15]. McMillan [DYB11]. McNicholas [ML16]. Mean [BL11, GG12, LL12a, LD17a, LFM14, LRH15, WSMG18, Xue17, AP18, AO12, AT15b, BWW11, BCI14, BH16, BLB16, BCD19, BK13, Bou15, CHM15, CDG15, DL17, DM12c, Fak17, GLM19, JB19, Kab11, Kar11, KN10, LZT13, LTR13, LLH15, LR12b, Luo10, MKJ15, Mar16, Mil19b, Nak16, Nan10, NHBG19, Now16, Par14, PR16, Rah11, Raq10, Röl18, RR10a, Ruk12, SW11, SD16, SKJ15, TM16, Tsu10, Uno13, WC17, WCD19, XZ13, XT11, YE10, Yu17, ZYW15, ZLZ16, ZJL13, ZYG17, Zhu13b, ZRAN13, ZRN14, ZRN17, ZC13, vBC19, vW17, Xue19]. mean-constrained [BH16]. Mean-field [LL12a, LRH15, BCD19, CHM15, Kar11, ZC13]. Mean-square [WSMG18]. mean-variance [BWW11, JB19, SW11, ZYG17]. meanders [BW18]. means [CD18a, CS17, CMS19, GP17, GSK12, GM18b, GMP19, HH19, LSS15, MM14a, NW10a, PY17, RV19b, SL12, Sun17, Wan18]. measurable [Nag17a, QSS17]. measure [ACP19, BD17, Bis10, EJ12, GGR11, GR11, Gre12, Kol17, KN15, KT11, Kum15, LW11, LXL19, Lu18, LPH13, RZ12, Rad19, Ric10, Sha13b, Sto12, Vet14, YG13, Yan14, ZG19, ZZZ13, dSF12]. measure-valued [LXL19]. measured [BS15b]. measurement [AHN10, Bel16b, Bha17, Che19b, GG12, Nku11, SSL^{+}12, Son10b, VB11, ZW18a, Zho15]. measurements [cZgL12]. measures [AMMO13, AJ14, Bak15, BP19b, Ber13b, BM10c, BM12, BVP10, Bra11a, Bra11b, CH19, CV19, Dav17, DZF18, DLM15, Duc04, Duc10, Gov14, Gri19, GB13, Jur13, KM17, LSS13a, LZZ15a, Liu14, LVMS14, MW18, NSL11, PL18, SC12a, SH11, SY18, SK16, TE12b, WZ15b, WZ19, WH14, Yan15a, Yan17a, YD12, ZC19]. mechanism [FP11a, JKP19, KPK15]. mechanisms [LP10a]. median [AO12, Che14b, GSK12, HCP19, LMP18, NKKY13, RM13, RV19b, SV15]. medians [MGA11]. medical [DvH11]. meet [GFH16]. Meier [Ane17]. Meixner [KM11]. members [KN17]. membrane [MP16a]. memory [AR18, BT13, Gus12a, HP11, HLV15, HS15, LSSR13, LK14]. merits [ABGM16, ML16]. mesh [Sab17]. message [SM16]. meta [Bak18, Ruk12, SWH^{+}11]. meta-analysis [Bak18, Ruk12, SWH^{+}11]. Method [GL19, Abd11, Azr12, BCI14, Bes14, BL10b, BH14, BS18b, BS19c, Cha19b, CW12, FD10, GX15, GR10, HBH10, Hür13, IL11, JLJ16, KV17a, KP15c, LGP10, LW13, Li15b, LS16b, LQ19, MKK17, MWGT17, Nic17, QCZ15, Rol10, Sab17, Sab18, Sak17, Sha17, Shm13, SS15c, TZW12, TMN16, Tia16, Tsi17, WR10, YGL14, YW17, Zaj14, ZZ13a, ZX19]. methodology [MVI11]. methods [ACD12, Bob10, CMT14, Gho11b, MGSL11, MK18, PS19a, PYS^{+}18]. metric [GSK12, Kat19, LVMS14, SY19b, YH16]. metrics [Bow17]. Metropolis [HL16, MDR12, XSL^{+}14]. Metropolis-adjusted [XSL^{+}14]. MGARCH [LW13]. microarray [HHY11]. Mikosch [ZL17]. mild [DDE17, GA12, GA13, Gov14]. mildly [Hwa19]. mimicking [For14b]. min [HLCH19, DS12]. minima [PP18a, PP19]. minimal [Roz14, YG13]. Minimax [AH13c, Ben18, Wis11, Ben16, Ben17, Efr10, Gho14, Tsu10, ZRAN13, ZRN14, ZRN17]. Minimizing [Ern17, LY18b, CKF17, SWH12]. Minimum [Bro14b, DY10, LNI16, Mak11, ND10, SM16, ZZ19a, AMP16b, BM10b, BK19b, CR13, HCP19, KB15, LZZ15b, LL15, MPA15, NS12, ySDM14, WN15, WN19a, Xia19c, YYS18]. Minimum-norm [ND10]. Minkowski [Shu19a]. misclassification [LLZ13, XD15]. Mises [Gat13, Sab18, BV15b]. mislabeled [Hay12]. mismeasured [BNSA13, LW17b]. Missing [PN12, BK12, BBGMPG15, CL19a, Che10b, Cho16b, DSW13, ED14, Efr14, GdL18, GXH17, HX14, Ili12, JKP19, KPK15, KP15c, LW15a, LZL^{+}19, NT14, PL19, PTW19, QCZ15, SB10, SS15c, WXL19, XFC17, ZMG10, dSGPM12]. missing-at-random [KPK15]. missingness [JKP19]. misspecification [Lem13]. misspecified [Hat12, MOD18, QO16, Sri15]. mixability [BP15]. Mixed [HBL11, AEFO18, ABD18, AD19, AR13, AD11, Bal11, BDS19, BDB^{+}10, CSK13, De 17, HLB17, KK19b, LMW15, LP13, LYB13, Li15c, LCLQ16, LD17a, MMPW16, MNBO11, Nag18, NR12, PB15a, PCGV14, PPN10, Rub15, SS13a, SD10b, VPCG13, WR18, WK11b, ZWHQ15, ZHL13]. mixed-effects [MMPW16, ZHL13]. mixed-Gaussian [MNBO11]. mixed-pair [BDS19]. mixing [AFJ11, ACMM13, BT16, BJ14, hCyP15, CMW18, CL15, DP19b, Eks14, GFA10, GZ13, HS11a, Hei18, Jon17, KS12, Kre18, LYW11, Lon13, Poi19, Sze11, Tab18, TY15, XXY17, XW17, Zho10]. Mixture [EQ15b, Lou13, AR13, BH16, BJW17, CKP17, CQZ15, HM10a, HB10, KA11, LDZ10, LZ17a, LBS17, LCKK14, LMH14, LMR12, MPA12, Mu15, NMUJ16, NSH17, ORO15, Pan17, SKRT16, VM12, WP14, WZS17, Yao13b, YZ18, YGTT13, ZP12, ZFZ13, ZW13]. mixtures [BB14a, Ber13b, JB19, KAVRY17, Kim08, Kim13, LBG11, Mar12a, MK10, MM13c, MM14c, Vig12, WYH14, YLHY19, Yu17]. MLE [CMW18, LSSR13, Now16, sS15a, Zha17b]. MLEs [YLHY19, vBC19]. Möbius [MZ17]. mode [DNKL12, MYA15]. mode-centric [MYA15]. Model [Abd15, Ciu15, GZWW15, JT14, LSCL17, Mao15a, SLM19, Sub12, WZ10, AT14a, AE19, AR18, ASP11, ASK18, AH13c, AdUÁMM11, AM19b, Arv19, ALS11, BTT18, BL17, Bak18, Bal11, BL12, BKP15, BH16, BOK17, BJW17, BSK19, BM10a, Ben16, Ben17, BZ13, BGO18, BW11, BMS10, BG19, BMB16, BTT16, Bos14, Bou18, Bra16, Bra12b, CW15, Cha19b, CWI10, CM11b, CGLN17, CZLT18, CH19, Che19b, CZCL19, Che18d, Cho16a, DMS13, DN14a, DMC17, DZD12, DE13b, DSW13, Din10, DWW17, DYW14, DSX13, DS18b, Dun16a, Elh14, Ery12, Fak10, FNP18, FZW12, FPZ13, Fil12, FS12, Fit14, For11, For14a, FF12, FN17, GMS18, GW12a, GS16, Ger11, GFH16, GB19b, GGO15, Goi12, HS12a, HP14, HC10a, Hat12, HX14, HZW16, Hel19, HJS11, HLV15, HH16, HL13b, HWZ14, HB10, HS15]. model [IK10a, JS16, JYL12, JCM15, JL16, Jov14, KM18, KHN12, KWL15, KST18, Kar11, KA10a, KMJ10, KS17a, KH18, KS15c, KM13d, KKKP17, KV17b, Kre18, KD14, LOKB11, LO13, LI12, LLL13, Lee12b, LMPR19, Lem13, LPS12, LWX11, LQZ13, Li17a, Li18, LWD13, LHT15, LSSR13, LZYW19, LX12, LZL13, LGW12, LLZ13, LC14, LXZ15, LW17b, Lop15, LMR12, Lu11, Lu12, LGT15, LZ16b, MY14, MZZ15, MP17, MBS13, MPA12, Mar14b, MS10b, MS12a, MdJ19, Mis14, MM13c, MM14c, MYA15, NA13, Nos19, ND10, OS18, Ost13, Ost18, ÖA15, PB16, PLY18, Pan19, PB15b, PVY17, PH10b, PLL13, PL17, PS18b, PS11a, PPN10, QO16, QCZ15, ROSL17, Ren14, Ren15, RBSB16, Roh13, RG18, Rul14, SY19a, SKJ15, SZ12, SZ13a, sS15a, SXM16, SLZ13, Shi14, SC15, SS17, ST14, SL13b, SHBHD11, SWH12, SQY18]. model [SB10, SdOG12, SS15c, TH16, TZW12, TZW13, TR12, TZ16, TW17, TPNSN19, Tom16, TZB13, Tor13, WY10, WYL10, WL11, WS12, WZT14, WST14, WY16, WXL19, WZ17, Wis11, WL14b, WQD15, WWH17, WCD19, XW10, YZ16, YH14, YJLL16, Yan17b, YL19, YS19, YY11b, ZY10b, ZW12, ZJS13, ZDM17, Zha19b, ZG10, Zha11b, Zha13b, ZC12, cZgL12, ZZ15d, ZP15, ZZLL17, dFAM19, dML12]. Model-based [Sub12, Lop15, MM13c, MM14c]. model-free [CZCL19, KS17a]. Modeling [Han10, ACP19, BOK17, CKP17, Cai17, GDNM18, KHHD16, WZ10, ZZ12a]. modelization [DNLR17]. Modelling [ASN16, BJD16, Gho11b]. models [AKM19, AB12a, AF12, Akr16, ASH18, ABC10, AR13, AI17, AD11, BLL10, BBM18, Bha17, BBGMPG15, BL10c, BG15a, BG15b, BK17, BCMR13, BG11, Bow16, BDB^{+}10, CG18, CF18, Che10a, CY15, CH11b, Cla14, CL10, CW16, CZ17b, DPZ16, DFKK12, DR17, DZX18, Duc04, Duc10, DCR19, Eck18, EP17, EM10, Ema15, Ery15, FA14, Far11, GR17, GS17a, GBR12, Gir16, GM10b, GG12, GZWW15, GXH17, HL11, HS17, HM10a, HD13, HT13a, HLM16, HHF17, He18, Hie14, HZJ^{+}10, Hua12, Huh12, IP14a, JH11, KT19, KS15a, Kau14, Kha14, KS19, KKNH18, KC16, Kur19, LBG11, LML15, LZLW16, Lem11b, Lem11a, LC10, LP10b, LLHW10, LDZ10, LLH11, LCX12, LP13, LW13, LYB13, LLW14, Li15c, LW16, Lia12b, LS12b, LTR13, LS10, LMH14, LYC14, LZJ17, LK14, Lou13, Luo12, LG13, LZY14, Mah18, MR15a]. models [MH12, Mao15a, MMPW16, MWZ11, MMC14, MSW15, NK17, NR12, NSL11, NMUJ16, Oh14, PB15a, PSS12, PO18, PQW14, PJ13, PF17, Ren14, Ren15, RR14, Ric10, RSC10, Ros13, RL14, Rub15, RG18, RS14b, SNS10, SA12b, Sch12, SL13a, Sha12c, SZW12, Smu19, SD10b, SZW10, SY11, SZ13b, TYK18, Tan19, TSL18, TLF12, TXX16, Tsi17, VB11, Vie18, VM12, WZ15a, WSW10, Wan12b, WS13, WW13a, WW14, WZG14, WL14a, WP14, WC17, WZS17, WZW17, WR18, Wan19, WCM11, Whi12, WS10b, Wu14, WS15, XJW15, XKBG15, XWT19, XZ13, XFC17, Yan15b, YW10, YL10, YGL14, YXL14, YY16, Yao13b, YGTT13, YWZ14, YQW10, ZW13, ZW18b, ZZLH15, ZX17a, ZX18, ZHL13, ZWT15, cZzXgL17]. Moderate [CG13, DL13a, Hu15, Hui10, JW16, MX17a, PO18, Rao18, SZ12, Son10b, Wan13b, Yao18, ZL12, Zhu13a, BS19c, CW15, DS14, Fan15b, GZ15, GW16, Kou12b, LWYZ17, LLP19, LZ13, Tao19, Yab17, Ye16, ZL14]. modes [WH10]. modification [IK10a, RMPB17]. modifications [JP12b]. Modified [CTN17, GEC17, LWL16, DZ11, DZF18, Fan15b, MD13, TH16, VR12, VRR13, WST14, WZ10, ZX12]. modulated [DMS13, Efr14, HMS14, PZ17, SWH12, XW16]. moduli [WX19a]. Modulo [Mas18, Alb18, Sze10]. modulus [BBH14]. Moment [Ger11, IMM16, KM17, KL15, LM15b, Mna11, Now19, BT14a, BS15e, CO13, Che14a, CH12, CV19, Ern17, GZ13, HR15, Hu16, Hür13, Jon10, Ko18, LN15, Mel16, MT17a, Mor18, ML17, NS13b, Oga18, OS10, PJ10, PL15a, QC14, Roz14, Ruz14, STW19, SG18, Sun13b, Sze15, Tru10, Uem16, UU11, WHYL10, WS16, WQD15, WWH17, XY12, YYS18, Zho10]. moment-based [WQD15]. moment-indeterminate [OS10]. Moment-recovered [Mna11]. Moment-type [Now19]. Moments [Gee15, Nag13, Sza15, WN19b, Cha15, CS14a, GGS12, GT16, IL11, Lan11, Lee12a, LLP19, LPN13, Mes14, Mir14, ML13, MS18a, MM14b, Nol18, PZ14, SV19, Sch14, SWL10, SL15, Sze16, Ush11b, VT10, Wei10, WZ16, WS10b, Yun18, ZCM15, Zha12]. Monge [FP14, KM13d]. Monitoring [CTW10, NMS15, QL11, ZY10a]. Monotone [WLY^{+}14, AMMO13, Cri18, GM17a, HLW10, KS12, Kar11, LL12b, LHT15, MZ17, Rez15, RR10a, Tan15a, Tzi17, WK11a]. monotonic [BS15b, DSX13, GB13, MM18, ZL15]. Monotonicity [KM13a, MA14b, Nag17b, Ald13, SC12a, Sre12, Xia19b, ZJ13, ZT19, vBC19]. Monte [BBCD^{+}18, DRZ19, Nie16]. Mood [Che14b]. moral [Pac18]. Móri [BBL11]. mortality [BG19, Lin17a]. Mosco [QG13]. most [KKKP17]. motion [ALL12, ANRW15, AH12b, AEK19, ATV10, BTT18, BL17, BRO14a, BBHH10, BD11, Bon12, BMNZ18, BH12, BH17c, Bry14, CVA13, CR13, CP17b, DOT14, DL10, Dav12b, DO11a, Dun06, DF11, FKZ15, FJW15, FM13, GM14a, GWK17, Gas10, Gat13, GS12, Gzy19, HT14, HT13b, HS12b, Hob13, HB18, JB13b, JV10, KP15a, KKNH18, KS16b, KS17c, LE19, LX18, LYCK11, LY18a, MP16a, Mar11a, MX17b, MS13, Met10, Muk17, ÖÇ13, Öz16, PH13a, PRD13, RH11, SC12b, Shi19c, SV17, Sun13a, VAB15, Wan14d, WX19a, WSMG18, WS17, XTT18, XL18, Yam15, YR11, Yu19]. motions [CY11, CG13, Hir11, OSW18, SS18, SG18, VY12, YS10, YYC17]. motivated [AV13a]. move [LLP11]. move-to-front [LLP11]. moving [AS15, CXW13, CK14, CM11c, De 17, Fer12, HT13a, LQC15, McE12, STD12, TE12a, Wu14, Zho10]. MSE [CL19a, DHL14]. MTD [Kau14]. Muckenhoupt [Osȩ13a]. multi [Akr16, Bra12b, BMS16, CKF17, DZD12, Gra11, He18, JYL12, KK19a, LN19, Lu11, Lu12, LW15c, MK14, MDR12, McC12, MST10, Nak13b, Ser18, Sha12c, Vil11, WCZ18, WZ13b, Zhe11]. multi-class [Gra11]. multi-dimensional [LW15c, Nak13b, Sha12c, Zhe11]. multi-factor [He18, WCZ18]. multi-index [Akr16, Ser18]. multi-layer [Bra12b, DZD12, JYL12]. multi-level [CKF17, MK14]. multi-point [MDR12]. multi-risk [Lu11, Lu12]. multi-state [KK19a, McC12]. multi-step [Vil11]. multi-strips [LN19]. multi-threshold [MST10]. multi-valued [WZ13b]. multi-variate [BMS16]. multicategorical [LLZ13]. Multiclass [DDZ15]. multicollinearity [Par17b]. Multidimensional [AS15, FF18, CH11a, GV16, JH11, LY17]. multifractal [SH11]. Multifractional [MR15b, DL10, Men12, Tie13]. multilinear [BT13]. multinomial [AJ16, GM10a, HH13b, Ost17, Ost18, WH10, WN14a, WK11b, YJLL16, YZ13]. Multiple [CSK13, BB12, BK19a, BHJ18, Bre12b, CX18, CCFL18, CHN14, Döh14, FA14, Gor14, HS16b, KZ14, KLNB14, KM17, KP15c, MA14b, Pin12, RM13, RW16b, SA12b, SM16, SK13, TM16, WZG14, WC17, WK10a, Yat15]. multiple-choice [WK10a]. multiple-outlier [WC17]. multiple-output [HS16b]. multiple-run-rules [CCFL18]. multiplicative [ACD12, Dur17, KLY14, Zha13a]. Multiplicity [SDNS16]. Multiplicity- [SDNS16]. multipliers [CL13]. multiply [KS15c]. Multiplying [Mat17]. multiset [Che12a, HC17]. multistable [Aya13, MR15b]. multistate [SM10]. multitype [Chu14, HA15]. Multivariate [BLWZ11, BT13, DFS10, FSNÚF11, Fer12, JSA12, LP10a, Mod16, NMS15, Onz11, AH13a, AR13, AB17, AZ17, AMP16a, AN15, BM17, Bel16a, BPXS19a, BSWS13, Bre12a, Bro07, Bro11, Cha19b, DN14a, DD10, DN14b, DK15, Duo15, Eck18, Eis15, EBG16, Fer11, Fur08, GGQ17, HC10a, HdCC17, HPW15, Haz11, HP12b, Hür13, INO10, IL17, JB19, JMS17, KRYAV16, Kak16, LVY15, LBS17, LTW18, LRL19, Mar12b, MM15, Mna11, Nav14, Oga17, Oga18, PNT15, PPA16, Rod13, Roh16, RR10a, SD10a, Shu18b, SL15, SF12, TW17, Tsa13, Tsi12, TW15, Tyl10, VM12, WN11, WN14c, WN14b, Yao18, Yeh11, YHY16, YKY17, ZW11, ZRAN13, ZRN14]. multivariate- [BSWS13]. mutations [Che13a]. Mutual [OSW18, BDS19, NÚF12].

naive [LM18, ZS19]. Nash [BCD19]. natural [Che16, Che18c, DM12b, HH13a, LCJ10]. Navier [ZBL19]. NBUE [AB11, AM11]. near [BG13, BS10, Dem10, Gir16, Jas16, ZZ15a]. near-efficiency [Gir16]. near-maxima [BS10]. near-symmetric [ZZ15a]. nearest [KLK12, Oua13, Yao14]. nearest-neighbor [Oua13]. Nearly [SL13a, WZYL17]. Necessary [ASVY14, Sma14, WWW10]. necklace [HP18a]. need [Hei18]. negative [AB12a, BGM19, BMD15, BLZ13, BLL14, CMS19, Coq15, DBB13, DJ13, DZ19b, EAB10, Gat15, Gre11, HD13, LMPR19, LZ14a, LZZ15a, PB16, Rao12, Roz10, ST19b, Sre12, Vas18, WYY13, WWPY18, WN14a, YZ13]. negatively [CS14b, Fen19, HT15, Ko13, TXWW18, YW10]. neighbor [Fil12, KLK12, Oua13, Yao14]. neighborhoods [Zha16]. Nelson [Ane17]. nested [CYLL19, DCR19, RZ12, Rez15, Sch13, XDWY18, ZJL18]. net [Zho13a]. network [Cai17, Chu18, GFH16, SQY18, Tor13]. networks [Bal11, CGLN17, DD14, FNP18, LD17a, LV11c, WSMG18, Yan15b, YZ16, ZM16a]. Neumann [AM13a, Che10b]. neural [HLV15, LD17a, WSMG18]. neuronal [MP17]. Neutral [BH12, BH10b, Che13a, CvGL18, CYS11, Gov14, LTX12, MX17a, TJS15]. Newton [FP13, VSP13]. Neyman [CRB19, Jal16, Wri17]. Nirenberg [LZP17, YWJ14]. NM [HL11]. NN [KV13]. No [FK10, HK19, Hat12, Ost13, WHB10]. no-arbitrage [Ost13]. No-information [FK10]. no-treatment [Hat12]. NOD [WHYL10]. node [Cai17]. noise [BN16, BJQS16, Ben16, Ben18, CHM16, DE13a, Dun19, FY16, HJS11, IKM16, KS19, LV11b, LW14a, Li17b, LZ19, MY14, PP18b, Son12a, ST10, Tie13, TZB13, TN17, Wan14d, Wu11a, YY18, Zha13a, ZBL19]. noises [BW11, JWW14, Ma10a, MY14, WWX17, WWZ18, XY17, Yan17b, ZYS19, ZZ19a]. noisy [Sak12, Vet14]. nomenclature [ABGM16, ML16]. nominal [BM10c]. Non [BL14b, GZSZ19, IV19, MASR14, PL19, Ter17, Van18, Xia19a, Yu19, AaM18, AB12a, Arv19, BKP15, BBP18, BCR10, CMS19, Cho19, CP13b, Den17, DE13b, DM14, DL17, FJ10, FLRS13, FS14, GS17a, GB19a, Giu15, HL11, HB18, HPP19, JLJ16, LO13, LST17, Lin17b, MW16, Mac16, MM18, Mar14b, MFL15, MV10, Mou19, NHBG19, OO11, PTT19, PP18b, PS18c, QW16, Qiu14, Rah11, Rah17, Rao12, Roz10, Sha14, SG18, Tsi17, TH14, Tum15, Vas18, WYY13, WP14, WX19a, WZ19, WWR12, XWML19, YE10, ZZ19b]. non-affine [GS17a]. Non-asymptotic [PL19, Xia19a]. non-atomic [PS18c]. non-autonomous [HB18]. Non-causal [BL14b, Giu15]. non-centered [BCR10]. non-central [PTT19]. non-commutative [Qiu14]. non-convolution [WYY13]. non-crossing [Den17]. non-differentiability [WX19a]. non-equilibrium [YE10]. non-exchangeable [HPP19]. non-existence [Tum15]. non-explosion [WZ19]. non-Gaussian [HL11]. non-homogeneous [CP13b, DM14, FLRS13, GB19a, LST17, MFL15, ZZ19b]. non-identical [Xia19a]. non-identically [Lin17b]. non-independent [PL19]. non-inferiority [JLJ16, Sha14]. non-iterative [Tsi17]. non-linear [LO13]. Non-Lipschitz [Yu19, FJ10, PP18b, QW16, SG18, TH14, WWR12, XWML19]. non-Markovian [Cho19]. non-monotonic [MM18]. non-negative [AB12a, CMS19, Rao12, Roz10, Vas18]. non-normality [BKP15, NHBG19, OO11]. Non-parametric [IV19, MASR14, Van18]. non-Poisson [MW16]. non-proportional [DE13b]. non-standard [Mac16]. non-stationary [Arv19, BBP18, MW16, Rah11, WP14]. Non-subjective [GZSZ19]. non-supercritical [Rah17]. non-zero [DL17, Mar14b]. noncentral [SKRT16]. noncommutative [QHS13, TMS19]. noncompliance [AG13, Chi12a]. nonconvex [LKK12]. nondeterminism [LX18]. none [WZ16]. Nonexistence [YS19]. nonhomogeneous [DYB11, SY10]. nonidentical [Abd11]. nonidentifiability [Wan12b]. nonignorable [MK18, PKK14]. nonintegrability [CHR16]. Nonlinear [XL16, ZHL13, AM13a, Bob10, He18, LBG11, LC10, LB17, LZY14, MS10b, PO18, SZW10, Vil11, WXL19, WS15, XJW15, XY17, YH14, ZJ13, ZL18]. nonlinearly [Lee19]. nonlocal [SLM19, WWW14, WWX17]. nonnegative [AM13b, AKU17, CS16c, Had11, HCW13, LMS17, Mel16, Osȩ11, SWL10]. nonnormal [SW13]. nonorthogonal [Jac11]. Nonparametric [ACD12, AMB10, AC15, BS18b, CSG11, DE13a, ED14, Efr14, GLS13, Huh12, Ker16, KKW14, KHHD16, Li15b, MU16, SY19a, TM16, Vil11, ZYS19, Zhu13b, BCL18, BL10b, BBS11, BRZ10, BV15a, BG11, BES18, CMW18, CCM19, Cho16b, Efr10, Efr19, Fre14, GMA12, Gee15, GW12c, Har16, HYWT17, KBM15, KKP15, Kur19, LBH11, LZ13, MGA11, ML17, MMF17, Nag18, Nic17, PSS16, PPA16, PS11a, SX13, SCZ15, SMZ18, Shi14, TY15, WYH14, WXL19, WCM11, Wis11, WWH17, Xia17, XLL19, Yao12b, YGTT13, ZPZ18, ZL12, ZL16a, ZM16b, ZX18, cZgL12]. nonrecursive [NK17]. nonresponse [MK18, PKK14]. nonsingular [Mad15]. nonstationary [AAEH11, DL17, HH16, LLW14, LR11, PP18a, XWT19]. nonsymmetric [LL12c, Osȩ10a]. Nonzero [Wei19]. Nonzero-sum [Wei19]. norm [BOR17, GLM19, Hwa13, ND10, Osȩ12c, SS10c, VL19]. Normal [Had17, HP12b, SW13, Sza15, WP14, Ana17, BPXS19a, BPXS19b, Ber13b, BM10b, BK17, CS17, CQ18, CFBD13, CH14, Ejs16, FP13, FS11, HK14, HX14, HJS19, JBS10, JB19, JSA12, Kab11, KRYAV16, KAVRY17, Kim08, Kim13, KS15c, KMV11, KA11, LBG11, LPS12, LMP18, LTW18, LPN13, LPNW14, LW14b, MKJ15, Mam15, Mar12b, MM15, MT17b, Nad15a, Nad16, OH15, Par14, PS11b, RR10a, Ruk16, Son15, TE12b, Uno13, Wan11b, WW11, WLP17, WN11, YHY16, YKY17, Yu17, ZG16, ZL16b, ZRAN13, ZRN14, ZRN17, vDLMW19, ORO15]. normality [BKP15, BS15d, BMN15, BL10c, Car10, Gee15, Jas16, Kre17, KW14b, LL19a, LX12, Lin17b, NHBG19, OO11, SWW11, Sep18, Sha12b, Sko19, Sko20, TXWW18, WMZW13]. normalization [BR17b, CWZ12, CD18b, FC15, IKM16, WL17]. normalized [CD10b, Fan17, Fen17, FD10, HWM15, Jon10, Pin15b, Xie13, XW17]. normals [BH16, LMR12]. Norms [Zaj19, LM15a]. Note [Bri15, El 13, JS16, Kab13, Ai16, AN15, Arv17, ALS11, Azr12, BW15, BC14, BSL13, BP19b, Bha13, BL14a, BPW18, Bon12, Bor14, BV15b, Bou16, Bra11a, BD10, BCNM15, hCyP15, Cai17, Can16, CDG15, CRC10, CS15, CMS19, CJT10, Che14a, Chi12b, Cri18, CR16, DMC17, De 17, DN14b, Dev12, Din15b, DZ17, DR17, Dur17, FRZ10, FP14, FS15, Fin11, FP17, GB19b, GBGPR19, GA12, GSST12, GM10b, HS11b, Han12a, HT13a, HN15, He16, HS12b, HB18, HR11, HR15, Hwa19, IMPR16, Jac13, JB13a, JP17, JP12a, JMS17, JY15a, KHN12, Kim08, Kim13, Kin10, KDW15, Kov10, Kri18, KS17c, LMW15, LT17, LP11, LT10, LLSW11, LG12, LYB13, Li15a, LL17a, Li17a, LLR17, Lia12a, LL12c, LS10, Lin17a, LZL13, Liu12b, LGW12, LG15a, LZJ17, LW17a, Lka19, Lop10, LVMS14, Ma10a]. note [MW18, MU10, Mak10b, Mao15b, Mao16, MV15, MS10b, MM11, MK18, Nak17, OS15, PSC12, Pen14, PPA16, RH11, RS12a, RS12b, RR13, Rol10, RMÁL19, SAM13, Sch12, SWL10, Shi19b, Shu19b, SGG10, Son15, SHK14, SQY18, Sze11, Tap10, Tia16, TA14, Tyl10, VS10, Vel12, VRR13, Wan13a, WYH14, WL15a, WWPY18, Wan18, Won13, WK10b, XW10, XXY12, Yan15b, YZ14, Yao12a, Yao13b, Yao13a, Yoo15, Zaj14, ZYW15, Zen15, Zha10, ZX12, ZZ12a, ZH14, ZJL13, ZG17b, ZG17a, ZGZ18]. Notes [Li13c]. notion [PSS16, ZP16]. notions [Il’18]. novel [MWGT17, ZHWH12]. Novikov [LHW19]. nuisance [TYNZ15, ZPK14]. null [AB11, GB16, MW14b]. nulls [FP11b]. Number [EF16, ABC10, AB12b, BCM14, BK19b, CD18a, CKF17, Ciu11, Fis11, GGR11, Gor12, Gus12b, IJB17, IA10, KN17, LKK12, LL10, LNC12, Li13d, LZ19, LS15, LZY14, MGSL11, QN15, Sch13, WSW10, YS12, Ye16]. numbering [Sug19]. Numbers [BS10, BHJ10, CQT12, CS16c, CS16b, CS19, Chr12, Dem10, DT10, HT15, Hor16, Hu16, Jas16, KM17, KD16, Kor15a, LLR17, LLR18, MKA19, Nak16, Neu13, QSS17, WA17, YH15, Zha17a, ZL18, dS18a, dMO18, AMMO13, DYB11, Nas12, QHS13, RS10b, ST13b, Sun13b, Ter15, TE12a, Tor13, WHRY11]. Numerical [CL16, DEM14, LCW17, Sak12, BB14b, Gho11a, LZ10, Lin19a, LFM14, ZX19]. Nuttall [Pak18].

Objective [AT14a, FXT12, LS12b, TH16, WS12, ZW18b, KWL15, WY16]. observable [KZ17]. observation [CZ17b, DDZ15, FFM18, LC14, Lka19, NSH17, PLL13, QLH^{+}16]. observation-driven [CZ17b]. observational [Gho11b]. observations [BV15b, But15, Dem10, Efr14, Ili12, Jas16, KO11, Lin17b, OK15, PL19, SW12, TW17, Vet14, WN10b, YHMM15]. observed [BDER14, BS15c, CDG15, CM11b, DMPX17, ES13, HMZ11, KW14b, Ma10a, Rah17, VD18, YS12, Yan17b, vL16]. Observing [Vid19]. Obstacle [AM13a]. obtain [VT10]. obtained [TE12b]. obtaining [HBH10]. occupancy [SvdV17]. Occupation [KMN10, ZW15b, ZWB17, CPY15, LZ14a, LW11, SH11, WWPY18]. occupied [IJB17]. occurrence [FJW15, GV16, PKK14]. odd [DS15, Jac13]. odds [BB12, CSG11, ER17, JLJ16, WZ15a, WL11, Wan17b, ZZM15]. ODE [PP18b]. Offline [AAEH11]. offset [BDvdAW16]. offspring [HS13a, Rah11]. Olkin [BMS16, MS10a, ND12]. OLS [Mar11b, Sha12b, Uem16]. OLSEs [TZ16]. omics [QO15]. on-line [XS18]. One [BABF19, FJ12, HP13, LL12b, Ush11a, Vir16, Abu12a, Abu12b, AP18, Alf13, CCS16, Che16, Che18c, Cui14, Fan16, GS17a, HZ16a, Hua16, Izu13, KWL15, LG12, LT19a, LS10, Lin17b, LB17, MLF13, MW12b, Osȩ15b, Pan19, PN12, PH10b, Rob15, SG15a, SY19b, VZ18, YYY17, ZZ13a, ZZ15a, Zhu15]. One-dimensional [FJ12, Abu12a, Abu12b, Fan16, GS17a, HZ16a, Hua16, Izu13, MLF13, Osȩ15b, Pan19, SG15a, SY19b, YYY17, ZZ13a, ZZ15a, Zhu15]. one-factor [LG12]. one-parameter [Che16, Che18c]. One-sample [BABF19]. one-sided [AP18, LT19a, MW12b, PH10b]. One-step [Vir16, Lin17b, LB17, PN12]. one-way [KWL15, LS10]. ones [AMP16b, BDvdAW16, KPO13]. online [AAEH11]. Only [Spa11]. onto [Son10a]. open [CCS17, WC17]. operating [AB12b, BH15]. operator [AM14, BDP14, CHW18, CHM16, Kin10, LRH15, Wan14c]. operators [BVP10, Gau19, Min17, QSS17, Woj13b]. Opportunities [Azz18, PY18]. Optimal [AEFO18, BS15b, Bel16b, BB11c, CL19a, CY15, DPZ16, EL18, Fer14, GZ18, HA14, He14, He18, KJS15, KK13b, LCF15, LL14, LZY15, LTP18, LH14b, MBTC13, MK14, MD16, MS12a, Pac18, WWY12, Yan13, ZW13, Zha18, ZW18c, ZZZ13, ASP11, AJ14, BSS13, BCM14, BB10, Buo18, But15, CDS17, CKF17, Chi13, CC16, Coh10, Den15, DS15, EQ15c, Els16, Fil12, Gus12b, HMT13, HP18b, HR19a, HW13b, HB11, Hwa18, JS18, KHN12, KP17, KK19b, KLK12, LLC^{+}12, LLCG12, LYC14, LZL^{+}19, Mal18, MPA12, MSS18, MK16, PN12, PPM18, Pro13, RSV10, Sak17, Sin19, Sma14, Tyu12, WZ17, Wri17, Yoo15, ZP12, ZZZ18, PRS18]. Optimality [BS15d, YY11b, AH13b, BB11c, Buo18, Jac11, Jac13, JKD14, JKD15a, Kao13, Pru19]. optimization [Di 11, HS16b]. Option [BGO18, BK10, GM16, KP14, SWH12]. optional [AM17, Ber13a]. options [PY18]. oracle [BB11a]. Order [Cha10, JBS10, LD16a, Abd11, AEAN18, AMB10, Alf13, AA11, AAC15, AV13a, Bak15, BS17a, BNB16, BS17b, BT14c, CS14a, Che18d, Cri18, D’O11b, Dem10, DM13, DFS16, DHL14, Efr10, Ery14, FW17, FC10, Fre14, Gei17, GGS12, GGQ17, HH19, HTA10, He14, He18, HWM15, HP12a, HP12b, HBH10, HS10, Hwa18, IL11, Jal16, Jas16, Kak16, KS12, KST18, KD12, Kau14, Kay15, KKL18, KS17b, KC16, LMW15, Lar15, LV13, LTvdVR11, LZZ15b, LL15, Lin19a, Liu14, MM18, MKA19, MD15, MWA16, MA11, Mir14, MM14b, MW14b, Nag13, NS10, Nav18, Osu12, ÖA15, PLH13, PK14, PL15b, Pin12, Qiu17, QJ18, RS16, RSC10, SKB14, SKB15, SY10, SK17, SL13b, Su10, Tak12, TMN16, TT15, TS17, WN10a, Wan12c, WL15a, WLP17, Wei10, WL17, YZ14, Zha11a]. order [ZJ13, ZL16b, dWGM12]. Order-invariant [LD16a]. order-preservation [ZJ13]. order-restricted [MW14b]. order-theoretic [KS12]. Ordered [ZY10a, Cas10, GO12, Gau14, ML17, SS13b, SR12a]. Ordering [KC16, ZB10, Cal13, KP15b, WW11, Wan15, WC17, YZ14]. orderings [BB11b, BZV11, DD11, LD10]. orders [BNB16, LL19b, ND12, RS10a, WZ16]. ordinal [CG12, LDD15, Lu18]. ordinary [Kre17, KKT14, Son10b, VD18, Zho15, ZF16]. oriented [MDN18]. origin [DK11]. Orlicz [Yu11, YY14]. Ornstein [Ma10a, Abd15, Ai16, ÁLBRM16, BSS13, BS15b, BPS14, BR17a, BY12, DY10, ES13, HN10, JM14, SWW15, SG15b, TLZ12, Vri16, Wan11a, Wan16, Xin12, XXY12, ZZ19b, ZWB17]. orthant [CS19, DM13, Lee12a, MD15]. Orthogonal [MDPP16, CLYW15, Jac10, Jac11, PN14, Rév12, Sep18, WZYL17, YCL14, ZJL18]. orthogonality [AV13b, VP12]. oscillators [dlCJB19]. oscillatory [dlCJB19]. Osgood [LV11b]. our [Shi18]. Out-Of-Bag [RP19]. outcomes [AHM19, AG13, LDD15, Lu18]. Outcrossings [KR13]. outdegrees [He16]. outlier [WC17]. output [HS16b]. outputs [CHM16]. overflow [Kin12]. overloaded [KK18]. overshoot [WYY13].

P [ML16]. Packing [DMWX15, Li11a, ZLT11]. paid [GW12a]. pair [BDS19, CP19, Ejs16, Kru17, LD17b]. pair-wise [Kru17]. paired [CDS17, Sin19, WZ10]. pairs [ED17, JLJ16]. pairwise [Li13b, NÚF12, SS13b, Sun13b, dS18a]. panel [BL12, BKP15, BDvdAW16, HHF17, LLL13, WS10b, YJ12]. panels [WQD15]. parabolic [BH17a]. paradox [Kab13, Pol13]. parallel [FZ15, FZB16, MM12, MM13b, WL15a, Wan15, WC17, ZB11]. Parameter [GS16, HN10, KMJ10, Xu14, Zha11a, AH12b, BSL13, BLKL13, BKL19, BTT11, BMB16, Cha19b, CXW13, Che16, Che18c, DBB13, DY10, DO11a, GBR12, GGQ17, JV10, KLM17, KS15a, KP15a, LSS15, Liu17, MKK17, Men12, Mis14, OH15, Rob15, Sha12a, Tak12, TYNZ15, WJY11, WF15, WL15b, Won13, ZPK14, ZJS13, Zha19b, ZZ19a, ZF16, dWGM12]. parameterization [PW15]. parameterized [Rol10]. parameters [Abd15, BS15b, BFS16, BM10b, BM11, CYW12, DKT19, DWW17, FL10a, FF12, GMA12, Gau14, Gre11, HMT13, HS12a, HLM16, KKMSA13, LKK12, LL10, LZY14, MP17, MGG11, MM14b, MT17b, NM18, PSS13, Rat15, SAM13, SS13b, TZ16, VB11, WSW10, Zha17b]. Parametric [AaM18, CXW13, EP17, LSS15, OS18, HS19, Hua12, IV19, LS16a, MS16b, MASR14, Van18, Won13]. parametrization [MZ19, TR12]. parent [HS13a]. parent-offspring [HS13a]. Pareto [BBdWG16, BMN16, FXT12, GGO15, MJH18, NAC13, RBY10, Tud14]. Pareto-type [BMN16, MJH18, RBY10]. Parisian [BL17, CR16, CP17c, Lka19, PL17, ZDD18]. parking [SvdV17]. PARMA [BL10c]. parsimonious [LMH14]. Part [OV15, Gri19, Hua12, Pin15a, Tsu10, BPXS19a, BPXS19b]. Partial [AA17, SC12a, TW15, Xia19b, BM11, hCyP15, Che10c, CvGL18, CYS11, DFS^{+}13, DDE17, DLZZ18, EPSU16, FS17b, GdL18, GB13, Kri18, LF17, Liu14, LTX12, Lu18, LG13, MD16, MSS18, PR12, RS10a, RR16, Sha17, SK16, SD14, TZ16, Vas18, Wan17a, WWY12, Wu11a, WX19b, XW17, XFC17, YAT16a, Yan19b]. Partially [MH12, BBGMPG15, BCMR13, CDG15, CLWH17, DM12b, DMPX17, DSX13, Ery11b, GO12, HHF17, HL13b, HZJ^{+}10, Hua12, LHT15, PDW10, Rah17, SV10, SZW10, Tan19, VD18, WS13, WW13a, WZW17, WXL19, YL10, vL16].

Particle [Vil12, Cou19, SD11]. particular [Hwa18]. partition [Far11]. partitioned [MVI11]. partitioning [VT11]. partitions [SW19]. parts [SG15b, Tak19]. party [BH17b]. Parzen [He14]. Pascal [GW12a, ZB10]. passage [Abu12a, Abu12b, Abu13, Bap11, BL11, HS12b, Hie14, IMM16, Met10, Pen14, VT10, ZH12a]. Passive [Cho19]. past [KRB13, MASR14, PD12, SSN13]. Pastur [HSY17, YH17, Yao12a]. path [HNGS15, KP14, Kou12b, Par17a, Pei15, QW16, WK10b, WW18]. path-dependent [KP14]. path-independence [QW16]. path-independent [WW18]. paths [CC16, JY15a, Ren13]. pathway [CWI10]. pathway-based [CWI10]. Pathwise [WZ19, ZX17b, Öns13, Yas19]. pattern [ZGZ18]. Patterns [FC10, Fen14, LMH14, SM10, Zha13b]. payments [ZDD18]. payoffs [FK10]. PCS [SÖV14]. PDEs [FY16]. peakedness [JP16]. Pearson [JDK^{+}11, MN14, Oga17]. penalised [WW10]. Penalization [WZ13b, ABC10, LS16b, VR12]. penalizations [LT19b]. Penalized [HC10a, ACW10, DE13b, LKK12, YY16, YLHY19]. penalty [Ciu15, DZD12, JYL12, Lia12a, LXZ15, WMH11, ZY10b]. Peng [HC10b]. Pepys [VSP13]. Per-site [SvdV17]. Percent [Yat16b]. percentile [ZJ15]. percentiles [Xia19a]. percolation [AC11, BH19, Mis13, SH17, Yao13a, ZZ13a]. perfect [Öns13, ZAV12]. Performance [KV16, Kab16, Zha10]. perils [McE12]. period [BML14, Cla14, Kab16]. periodic [ABD18, BL14a, DE13a, DZ19b, PY18]. periodically [HS16a, She11]. periodicity [RV19a]. periodogram [JP12a]. periodogram-based [JP12a]. Permutation [ZAV12, PV17, RM13, Yor14]. Permutation-based [ZAV12]. permutation-invariant [PV17]. permutations [AIP18]. persistence [CTW10]. persistent [WL14b, Woo15]. Persistently [PTV16]. perspective [BvdG18, Cer18, CH11b, GDNM18, LP10a, Qua18]. perspectives [SK13]. Perturb [Tom16]. Perturb-and-MAP [Tom16]. perturbation [ART14, CW15, Li17a, ZK14]. Perturbations [Li17c, PP18b, Tom16]. perturbed [BTT11, IP14b, LR13a, LN19, LZJ17, MST10, XYZ16]. Petersburg [Gut14, GML15, Nak15, Nak17]. Petrov [Kor15a]. PGARCH [BL10c]. phantom [JT18]. phase [Abd11, BB11c, JYL12, LZYW19, ZH12a]. phase-type [Abd11, JYL12]. Phases [SM13]. phenomenon [AM19a]. Phylogenetic [JM14, SHK14]. Picard [LL17c]. Pickands [MS10a]. piece [YW17]. piece-wise [YW17]. Piecewise [BBCD^{+}18, Rat17, CTN17, MM10, Rat14]. Pisot [Neu13]. Pitman [RA13b, AB10, BD13a, BBHV13, JDB12, Maa19, MS16b, MS16a, RA13a]. Pivotal [SK15, CRG17]. pivoting [BCI14]. placed [Dav12a]. placement [KLW11]. plan [PN12]. planar [MX17b, Tor13]. plane [PTW19]. plans [CLYW15, DS15, Jac11, Jac13, JKD14, JKD15a, JKD15b, Yan13, ZG17b]. plants [GG13]. plastic [PVY17]. plot [AHM19]. plug [ÁLBRM16]. plug-in [ÁLBRM16]. pMAX [FF14]. Poincaré [DJ13, ZW15a, Zha19b]. Point [JMS17, KO17, MA11, Abu12a, Abu12b, Abu13, BM13b, Coe15, CRB19, DZ19a, DS11, DWWG18, Ery18, Fer18, Fit14, Gat13, Hei18, HLLK13, Huh12, Jal16, KK15, LQZ13, MW16, MPA12, MDR12, MS16b, Mih12, MW14b, PTW19, Pet11, Poi19, RV19b, Roh16, SÖV14, Ser18, She13, SD15, TL19, TR12, VR12, WK11a, XWT19, Yao14, ZWT15, ZWLB17, Zhu15]. point-null [MW14b]. points [Bél11, Bre12b, CTN17, Ciu11, EF16, HR19a, IMU10, MK10, Osu12, PTW10, PRS18, Rat17, Sch14, YS12]. Pointwise [DG13, GLM19, LZ13, Bór18]. Poisson [BRBB23, DFT13, ASN16, ABD18, AD19, ABTV19, BM13a, BM17, BM11, BRBB14, Bre14b, Cha19a, CDL11, CvGL18, CLG13, CYS11, DKT19, De 17, DS11, DMPX17, DFT12, FK10, FLRS13, Fit14, FF12, GX15, GW15, GMS18, Han10, HLLK13, Hui10, JS16, JLY14, KS15a, KV17a, KP15b, KKT14, Kru10, KNV11, LST17, LW14a, LH13, LXL19, MW16, MFL15, MST10, Mod16, MR15b, MN17, Nie19, OP12, OP13, ÖÇ13, PZ17, PCGV14, Pog16, Rat15, SBA12, Sas13a, SLZ13, ST14, ST10, THC17, VPCG13, WZ19, WS16, Xin12, Yan17a, Yao14, YY11b, ZYL14]. Poisson-gamma [JS16]. Poisson-stopped-sums [VPCG13]. Poissonized [ZM16a]. Polar [Söh10]. policies [Azz18]. policy [Chi13, LH14b, LXZ15, Mal18, WL15a]. Polish [Pet11]. Polya [Sta16, BB14a, BT14b, BD11, Cal13, DW18, FS17a, KM12b, MB14, SM13, WR18, ZCM15]. polygon [Bél11]. polygons [WX18, Xu15]. polymer [Hua16, Wan13b]. Polynomial [LML15, BT13, DZ11, Din14, HdCC17, LFM14, Mat12, ZWLB17]. polynomial-form [BT13]. polynomially [Tra14]. polynomials [Kak16, Lan11, Su10, Sza19, ZP12]. polytopes [Öns13]. pooling [VS10]. poor [KM18]. population [AO12, BBHH10, CDG15, CM16, CSS14, FNP18, HCP19, KN17, LLY19, MT17b, RA13b, RA13a, RR10a, TZB13, WLLZ11, ZLZ16, ZZT17, ZG16]. populations [BMP10a, Chi10, Fre16, KK13a, LLC^{+}12, MKK17, MM15, MSPSJ17, Pin17, SSL^{+}12, SS12a, Zha14b]. Portfolio [Fra11, ZYG17, JS18, LWD13]. portfolios [WH14]. posedness [LHW19]. position [AMP16b]. Positions [Aza13, SZ18]. positive [BC11, BLZ13, CS17, Lar15, Lee12a, Mic11, Pin15a, QSS17, ST19b, Tie13, Tsu10, WN19a, ZP16, ZDD18]. positive-part [Tsu10]. positively [NA13]. positives [MGSL11]. positivity [Arb19, CK14, LL17a]. possible [AEK19, GMS18]. possibly [Gra11, Hat12, Ker16]. Posterior [BJW17, WY16, BB11a, CM10, CJT10, HNW19, MOD18, Rub15, Tsi17]. posteriors [LS10, LSSR13, SHL15]. pot [BK18]. potential [AHM19, AG13, BH17a, BBHH10, LZZ15a, WS10b, Xin12]. Potts [Mar14b]. Poussin [Cha15, HR11]. Power [AA13b, GB19a, LZ11, TLNO11, AAEH11, AT15a, BDvdAW16, BKR18, BR17b, BM11, CO14, CWZ12, CCGPW17, FQZ18, FC15, FP17, GLML12, GMDP18, JDK^{+}11, Lee12b, Liu11, MW12b, Neu13, Tum15, VZ18, Wan14a, WL15b, WL17, YS12]. power-raised [FQZ18]. powered [ZL16b]. powerful [KKKP17]. powers [AJ16, CHW18, CS18, Kak16, Wój13a]. PPMI [Akr16]. practical [Llo12, SIBC18]. pre [LZ14a]. pre-exit [LZ14a]. Precise [BZ13, SXM16, XY12, XZY13, HCW13, KD16, YLS10]. precision [Bel16b, Nie19, vW17]. predator [FF12]. predator-prey [FF12]. predictable [Osȩ15a, YY14]. predicting [Ste15]. Prediction [Dun06, DF11, LG15b, SSL^{+}12, SV17, AMB10, BABF19, BSS13, BH17a, CWI10, Coh10, DNT10, Fre14, Goi12, KS10, KN11, LTR13, Liu17, MMPW16, MS16a, MvdBV^{+}11, MA11, Onz11, Pru19, RS12c, TW17, Vil11, Wan13c]. predictive [CF11, EJ11, Mar12a, MS16b]. predictor [ÁLBRM16, KM19]. predictors [AZ17, DD14, Onz11, RMÁL19]. Preliminary [MTT11]. presence [AHN10, BDM17, CDG15, KZ17, MHH11, NS12, Par17b, QB14, RH15, WLLZ11, ZMG10]. presence-absence [MHH11]. present [JCM15, Pac18]. Preservation [LL17b, Nav18, SBA12, Bak15, ZJ13]. preserve [SS11a]. preserving [Li11a]. presmoothed [JdU16]. Presmoothing [AdUÁMM11]. prewhitening [ZJZ16]. prey [FF12]. price [KT10]. Pricing [CJN19, HJS11, Ber13a, BK10, BGO18, GM16, KP14, LWD13, QY15, SWH12]. principal [BB14b, BMS10, BCR10, CZ17a, Chi10, HLR15, MK10, PR17, Smu19, TKO12, Tsi12, Nos19]. principle [CW15, CG13, CL13, Gee15, Kou12b, Leh15, LWYZ17, LY16, PO18, Rad19, Rao18, SZ18, XZ16, YD12, ZZ15c, ZLC18]. principled [Men18]. Principles [CG18, BS19c, DM12a, DL13a, Hu15, JW16, Jir13, LLP19, Mar12b]. Prior [LNC12, Ohy13, EJ11, HNW19, LD16a, Rod13, RL14, SLM19, WL15b, YJLL16, Zhe11]. prior-data [EJ11]. Priority [Xue19, Bro11]. priors [CM10, GL19, GZSZ19, KWL15, LS12b, LA13, PB15b, TH16, WY16, vWvZ17]. Probab [Abu12a, Bai13, BSCQ18, BRBB23, DR15, DFL13, DFT13, Duc10, GHR17, Kak11a, Kim13, MN13, MM14c, RA13b, Ren15, Sko20, SF12, Suz20, Ton17]. Probabilistic [EK10, Wan17a, WX19b, Che14d, Cho16a, GLML12, Sza19, Wal15, WY18]. probabilities [ART14, AdUÁMM11, AJ16, BG16, CCFL18, CL13, Chu14, Den17, Dun16a, Fro12, FN17, GR17, HEM10, Kob16, Lem11b, Lem11a, Li18, LZ16b, Mac11, MP16b, MN17, NC16, Nie16, Nog13, Roz12, Roz14, Roz16, SW11, SZ13a, Shi19b, VCM14, WY10, Yan15c, ZW12]. Probability [AA13a, Bro11, Red17, Sug19, AJ14, BM15, BS12a, Bak15, CK16a, CKO19, CRG17, DdRS^{+}11, DD11, DHJT14, DZF18, Doe18, DHL14, Dun16b, Dun19, DM12c, GL13, Gat15, Gna12, Gov14, GM14c, Han12a, HR19b, Hof13, Jou14, Kat13, KK18, KN15, Li11a, Li17a, LY18b, Lot19, LMS17, LVMS14, MU16, OS18, ÖÇ13, PR16, PH10b, PTV16, PS19b, QHS13, Sau12, sS15b, Ste19, Sto12, Vid18, Xue17, Xue19, YW10, YHW11, YLS12, YAT16a, YH16, YL19, Yor14, YQW10, ZK10, Zha14a, Zhe17]. probit [LS10, ST18, YJLL16]. problem [AEFO18, Abu12a, Abu12b, Abu13, AM13a, AH12a, Bak12, BS15d, BK12, But15, DN14a, DE13a, Den15, DP16, FP14, Fuj10, Jia16, Lau10, LS15, Liu17, LD16b, MGA11, Met10, NTT18, SD15, SvdV17, Sma14, TN17, VSP13, WC17]. problems [AH11, AV14, Buo18, ÇKN17, Che10b, Di 11, EGX16, FK10, KK19a, KK13b, Li15a, LZ14b, LC14, LZJ17, MT17a, Mil19b, PLL13, RSV10, Sab17, VSP13, WM18, ZZZ18]. procedure [Ger13, HH13b, IKAH17, KZ14, Ref14, Uno13, WX12]. procedures [Gor14, LLR17, LB17, MS16a, NT14, WR10, XD15]. Proceed [SYZ18]. process [Abu13, Ai16, Alf13, AA19, AB17, AM14, ÁLBRM16, AKV12, ABTV19, AD12, ADM15, BJW17, BM17, BR17a, BB18, BS15e, BD11, Bou17, BD10, Buc15, But15, CX18, Cha19a, CDL11, CF13, CSY12, CCS17, CL19b, Coq15, DKT19, DH19, Den17, DY10, DZ19b, DP19b, DMS19, Dur17, EIW14, EKS11, FRZ10, FL19, FLRS13, FF12, GKW19, GW15, GZ15, GW16, Gou15, GZ14, HP18a, HNW19, Hil14, Hil19, HA15, HH13b, Hui10, HB10, Hwa19, Jal16, KD12, KV17a, KLY14, KM19, LST17, LWX11, LR13a, Liu12a, Liu13, LK14, MT13, Mak10a, Mar16, MOD18, Mar12b, Men12, Mic11, Mil19a, MWH17, MST10, MR15b, NMS15, NZ19, Oes15, OP12, Pen11, Rah11, Rat17, Ren13, SBA12, Seo19, Ser17, Sha12c, SZ12, SK17, SS10c, SG15b, Tau15, TE12a, Tim15, Vid17]. process [Vid19, WGL11, WMH11, WZ13a, Wan14d, WLLL17, WM18, WN15, Wu11a, XXY12, XW16, XTT18, XS18, Xue17, Xue18, YG13, Yan17a, Yeh11, Yun18, ZA12, ZZ15a, Zha19b, ZP19, dSF12, HL13a, Xue19, SM13]. processes [Abd15, AaM18, AAEH11, AT15a, AMG17, AA17, ACMM13, ASVY14, BT13, BGT15, BT16, BBP18, BM13a, BPS14, BD17, BL14a, BBCD^{+}18, BG16, BL11, BY12, Bog15, BSCQ17, BSCQ18, BN10, Bre12b, BL15, CX18, CL19a, CJ14, CXZ15, CL16, CJN19, CPH12, Chu14, CF11, Coe15, CRB19, CM11c, CJY14, CSL16, Cri18, Cri19, CCSC11, CR16, CP17c, DNT10, DZ19a, DM12a, De 16b, DHJT14, DFS^{+}13, DS11, DWWG18, DY10, DMPX17, DHOV13, DFL12, DFL13, Dun06, DF11, Dur13, EHP13, EL18, Ery11b, ES13, FF14, FCU11, FL10b, FS18, FP12, Fuj10, Gat15, GFA10, GP17, Gho11a, Gir16, GMP19, Giu15, GML10, GLML12, GLT18, GSW11, Gri11, GS10, HP11, HZZ19, Hei18, Hew17, Hir11, Hof13, HLLK13, HN10, HMZ11, HZR12, HMS14, HBPC10, Hwa13]. processes [IKM16, Iso15, Jas16, JK13, JLR19, JLY14, Jur13, KT10, KM11, KM13b, KW15, KC19, Kin10, KR13, KSW15, Kob16, KZZ13, KP10, KKT14, KW14b, Kri18, KNV11, KR12, LL19a, Lau10, Lee12a, Lee19, LPW10, LW14a, LZ14a, LZZ15a, LM17, LT19b, Lia13, LR11, LZ11, LHM14, LQC15, Liu17, LXL19, LR12a, LL17d, Ma10a, Ma14a, Ma15, MW16, MP15a, Mac16, MXZZ18, MFL15, MN16, Mih12, MC15, MN17, MSPSJ17, NP16, Nie19, Öns13, OP13, Osȩ13a, PR15, PTW10, PZ17, PP14a, PNT15, PH10a, PTW19, Pet11, PS15, PT10, Poi19, QL11, Rah17, Rao18, Rat13, Rat14, Rat15, RZ10, RS12c, SV19, SB19, Seo15, Seo17, Son14, SWW15, Son16, Spa10a, ST10, Sta12, Sta16, Tak19, Tan13a, Tan15b, Tao19, TLZ12, TK10, Urb12, Vas18, VB17, VT10, Vii16, VW14]. processes [Vri16, VL19, WS10a, Wan11a, WZT12, WL13, Wan14c, Wan14b, Wan16, WWPY18, Wei10, WS16, WN14b, WK10b, XY13, XWT19, Xin12, Xu13, Xu14, XYZ16, Yao14, Yao18, Yas13, Yaz15, Zha11a, ZH12a, ZH14, ZZT17, Zha17c, ZDD18, ZZ19b, ZLT11, ZX17b, Zho10, ZW15b, ZWB17, Zhu13a, Zhu15, dSGPM12, vL16, JMS17]. Product [AB12a, NM19, Cha17, CQ17, GFA10, HLM16, Jal16, JT11, LWY17, MN11, MN13, Oga18, OS10, Red17, SH11, SL15, Tru13, WN10c, YHW11, YS13, Zen15, Zhu14]. product-limit [GFA10]. products [Arb19, GS13a, Hu15, Zen17]. profile [BB12, IK10a, MD16, Mou19, VR12, WZW17]. profiles [FFM18]. profinite [CLEA18, CLEAMS16]. prognostic [WSU11]. programming [MK14]. programs [CC16]. progressive [BPS17, FXT12, PN12, PNC15]. progressively [AEAN18, BKL19, HTA10, PP15, SK15]. Projection [Akr16, YZ18, BB11a, BB14b, BMS10, For14b, GM14b, IKAH17, MX18, Osȩ15a, Son10a]. projection-pursuit [BMS10]. projections [PR17, Ski10]. Prokhorov [ST13b]. prone [HZJ^{+}10]. proof [Aga15, Alf15, AG11, AG18, Cui14, DS11, JP16, Kir19, LSS13b, OS10, Shu16, Shu18b, Wal15, WZ15b, WY18, Yan12, Zho13b]. proofs [BPW18]. Propensity [Gho11b, GXH17]. Proper [Vid17, DMC17]. Properties [CP17b, ED17, GSK12, HM13, JH11, SK16, TYNZ15, BP19a, Bel16a, Bel12, BPXS19a, BPXS19b, BH15, BSCQ17, BSCQ18, BRBB14, BRBB23, BJ14, Cha19a, CCM19, DSS18, Din15a, Fra11, HTA10, HNGS15, HC10b, Ilm13, JH14, KC19, KC16, LW11, LW12, LZL13, MS16b, MS16a, MNOT12, MX18, Nag17b, Pan19, PLN16, PR17, QO15, RS12c, SBA12, SY10, TV17a, Tau15, Tom16, TE12b, TE14, Wan13c, WL13, Wan14b, XW10, XWT19, YWZ14, ZY10a, ZG19, ZB10, ZB15]. property [Ald13, Ber13a, Bór18, HJS19, JHF15, Jur14, Kou12a, LH14a, LMM13, Ost13, Pol13, Sch16, Tri18, Vir16, Wan12b, WY16, WW18, XY13, YLS14]. prophet [Osȩ13b]. proportion [DR11, FP11b, HH13b, MBTC13, ZM17]. proportional [ASK18, DE13b, HL13b, LZY15, LY18b, LZYW19, Lin17a, Ref14, sS15a, TR12, TYNZ15, WZ15a, WL11, WYC15, ZZ15d]. proportions [Hom12, SM19, Thu14, Wan12c, Wan13a, YK11]. propriety [Rub15]. Proschan [AB11, AM11, GM17a]. proteomic [MvdBV^{+}11]. proximity [BVP10]. Pseudo [MPA12, MVI11, UK18, AY17, LZ17a, SK11, SHL15, WR10]. Pseudo-Bayesian [MPA12]. Pseudo-binomial [UK18]. pseudo-empirical [SK11]. pseudo-integrals [AY17]. Pseudo-likelihood [MVI11, SHL15]. public [Azz18]. pure [DMS19, Gou15, ZX17b]. Purposely [MOD18]. pursuit [Akr16, BB14b, BMS10]. puzzling [Tri18].

QELE [HL11]. QMLE [Arv19]. quadrant [dS18a]. Quadratic [LKK12, AKM19, BGT15, BBHH10, Che14a, Din15b, EAB10, EAB11, GS12, HS11b, Hui10, Jon10, KLY14, LTW18, Vel12, Wes13, YY13, ZP12, ZRAN13]. quadruple [Ton16, Ton17]. quadruple-wise [Ton16, Ton17]. Qualitative [HWM15, LSS13a]. Quality [GMDP18]. Quantification [GL19]. Quantile [CGH11, PK11, SS12b, SSN13, AFJ11, AT14b, BS15a, Ciu15, DNT10, DLZZ18, GFA10, HS16b, HYWT17, JQZ12, KKW14, LLC^{+}12, LLCG12, Lia12a, MDN18, MS18a, NV11, NSS14, NT14, OPS16, Par17a, RNB17, SNS10, SSN16, SMZ18, Sri15, TL19, TL12, TZW12, TSZ15, TZT16, WMZW13, WMFW15, XT11, XSZ17, ZBZ11, ZJZ13, ZL16a, ZWT15, vSK15]. quantile-based [BS15a, vSK15]. quantiles [Bel12, BBP17, DHOV13, FMA16, GGV14, HCT16, Jur10, JP12b, LGP10, RA13b, RA13a, SW13, SDJ15, TWM15, WN10b]. quantitative [Kat19, Röl18, SKK10, WRvdL11]. Quantizations [Bak15]. Quasi [Cas10, KLNB14, AF12, BH19, BG15a, BG15b, BH10a, CJ14, CP17c, DWWG18, DRZ19, FSNÚF11, HZ16a, HZZ19, Kha19, LYB13, Li13b, Nak13a, WP14, Wan17a, XJW15]. quasi-asymptotically [Li13b]. quasi-copulas [FSNÚF11]. quasi-ergodic [HZ16a, HZZ19]. quasi-ergodicity [CJ14]. quasi-infinitely [Kha19, Nak13a]. quasi-likelihood [AF12, DWWG18, XJW15]. quasi-linear [BH10a, Wan17a]. Quasi-martingales [Cas10]. Quasi-maximum [KLNB14, BG15a, BG15b, LYB13]. quasi-Monte [DRZ19]. quasi-stationary [CP17c]. quasi-transitive [BH19]. quaternion [YH17]. Quenched [JT18, VW14, MLF13]. question [BBL11]. questionnaires [WK10a]. questions [Bow18]. queue [BML14, Bri15, IM15, KK18]. queueing [MW16, PSC12, VCM14, Whi12]. queues [DMST13]. queuing [FL10c].

R [LYCK11, ML16]. R/S [LYCK11]. Rademacher [PS18a, Pin15b]. radial [ACP19, Hir11]. radii [HLR15]. radiosonde [FFM18]. radius [HL16]. Radon [ML13]. raised [FQZ18]. Ramanujan [EHP13]. Random [ABD14, GHH^{+}10, KP10, NM19, Pac17, PRD13, RS16, SKB14, WX18, Xu15, Abd11, Abu12a, Abu12b, Abu13, AS18, AP18, AP13, AA13a, AO12, Ale18, AM14, Ana17, AKU17, ANV13, Arb19, ABBK15, ABB^{+}17, AD12, AKP18, AvHZ17, BS12a, Bak18, BS12b, BDP10, Bap11, Bar12, BDER14, BS15c, BBHV13, BDS19, BCM14, Bél11, BV10, BTT11, BJD16, BK19b, BH14, BM12, Bou18, BM18, BL14b, Bud14, CP13a, CW15, CQT12, CM11a, CHR16, CQ18, CL19a, Che18a, CC10a, CH11a, CS14a, CS14b, CXZ15, CYH16, CS16c, CS16b, CS19, CS18, Che14c, CLM12, CL13, CK14, CD12, CC16, CLEA18, CLEAMS16, CD10b, DLO16, DS16, Dav12a, Dav13, DP19a, De 16a, De 16b, De 17, DFKK12, Dev12, DR02, DK14, DR15, DKM19, Din15a, Doe18, DK11, DMWX15, DDZ15]. random [DT10, DG13, EAB10, EAB11, Ejs16, EB17, EF16, FA14, FNP18, Fen19, FS17a, For18, GW15, GZ15, Gao19, GD15, GD18, GF11, GW12a, GGR11, GM18b, GMP19, GBGPR19, GdL18, Gri19, GZ13, GS13b, Had13, HS16a, HS11a, Han10, HS12a, HT13a, HWB10, HSY17, HK14, HCW13, HL19, HT15, Hil14, Hil19, HNGS15, HBH10, HM15, Hom12, Hor16, HBF14, HR15, HZR16, HZRS19, HLCH19, Hwa13, IP14b, IL17, IK10b, Isl16, JT11, JGW14, Jir13, Joh14, Jou14, JP16, Jur13, Jur14, Jur18, Kam17, KWL15, KI15, KK11, KPK15, KK19a, KS13, KM17, Ko13, Ko18, KK13b, KK15, KL15, KZZ13, KA10b, Kou12a, KP16, Kru17, LBM14, LV11a, LH14a, LLSW11, LYW11, Li13b, LL15, Li15b, LLP19, LL19b, LS10, LZL13, Liu12b, LYC14, LW15a, LMLW15, LT14, LX10b, LMS17, Ma13, Mah18]. random [MB15, MB16, MW11, MLF13, MV15, Mas18, Mat12, McC13, McK17, MNX13, MW12a, MS11, MD15, Mes14, MV10, MNBO11, Mod11, MMM13, Mor18, Mou19, Myr19, NKKY13, Nak16, NTT18, NP16, Neu13, Neu17, NC16, NT14, Nku11, Ohy13, OS10, Pac16, Pak13, PL18, PYK15, PP18a, Pan17, PS18a, PR16, PDW10, PS15, Pin15a, Pin15b, PS16c, PS19b, PW15, Puc13, QC14, QG13, QN15, RP19, Rao10, Red17, RV19a, Rob10, Rob13, RS10b, Roz10, Roz12, Roz16, RM11, Rub15, RR16, Ruz14, Sab17, SY19a, SBA12, Sch14, Sha13b, SWL10, SWZ19, Shi19b, Shm13, Shu18c, Shu19a, Shu19b, SV15, Ski10, Söh10, SL15, STW19, Sre10, SD14, SW19, SL10, Sun13a, SZ18, Sun13b, Sze10, Sze15, Tah12, Tan13b, TT17, Tau15, TY15, TLF12, Ton16]. random [Ton17, Tru10, TN17, Tur14, Tyk11, VS10, Vas18, Vig12, Vol14, WGL11, WW11, WYY13, WLLL17, WZ19, Wei17, WZ17, WN10c, WWD15, Wój13a, WQY12, WWH17, Xia19a, XY12, XX13, Yak15, Yan14, YLS10, YHW11, YLS12, YS13, YXL14, Yan15c, Yan19a, YCQY16, YWY10, YWC10, Zaj19, Zen15, Zen17, Zen19, ZH14, ZZ15b, ZZ15c, ZB10, Zha12, Zhe17, ZH12b, Zhu14, Zhu15, dS18a, dMO18, dC16, dlCJB19, Sab18]. random-coefficient [PS15]. random-effects [Bak18, De 16a]. randomization [LDD15, Lu16a, Lu16b, LD17b, RS14a, SS11a]. randomization-based [Lu16a, Lu16b, LD17b]. randomized [AWL17, Chi12a, DP16, LSS13a, PPM18, SA12a, ST14, Thu14, VT11, XLL19]. Randomly [DS16, Hom12, BRO14a, Dem10, DZX18, GW16, Gut14, JQZ12, Lin19a, RS15a, SR12a, TZW13, Tzi17, WW16, YLS12, YLS14, YIS15, Yor14, vWvZ17]. randomness [Van16]. Range [XX13, BT16, BM10a, Buc15, HL16, Hua16, Mis13, Shm13, TWM15, WHB10, Wis11, ZZ13a]. Range-renewal [XX13]. Rank [DD11, FZW12, DS18b, GO12, Han18, IP14a, KS19, PRS15, sS10a, Sko19, Sko20, Tab18, WWW19, ZDM17, ZW11]. Rank-based [FZW12, KS19, sS10a]. rank-ordered [GO12]. ranked [AWL17, AO12, CXW13, ER17, JDB12, MW17, TJE16, ZV15, ZM17, ZDX14, ZLZ16]. ranking [ZAV12]. ranks [Aza13]. Rao [ASH18, MN16, Onz11]. Raphson [FP13]. rapid [HMS14, ZA12]. rapidly [JT11]. rare [ST14]. Rate [BT14b, KM12a, KBM15, Sak15, STD12, XH12, Yao14, AvHZ17, BB11a, BJW17, BHJ18, BBHH10, BLS14, BS15f, Bor16, Bou15, BZV11, BT14c, CTN17, Che13a, Che19a, CCSC11, DMS13, De 16b, DL13a, Dun16a, Gao19, GGQ17, HA14, HP14, Hel19, KP15a, KD16, LDZ10, LL19b, LGW12, MM18, Mar12a, Nan10, Nie19, OS15, PB16, PL19, PH10b, SC15, SDNS16, WX12, Wan15, Woj13b, WLY^{+}14, Xue18, YW10, ZK15, ZKSK18, ZGK13]. Rates [Bou15, CWZ12, CH14, LPNW14, Ben17, BF17, Cou19, Döh14, FD10, FL10c, HS13a, He14, KK13b, LCKK14, MS12a, MM13b, Sha14, Tsi12, XZY13, ZYL14]. rather [AV13b]. Ratio [AO12, BWW11, BB12, BNB16, BH17b, CO13, CSG11, Eis15, GEV16, JP11, Jas19, JW12, JLJ16, LL15, MD13, OO11, Sub12, VZ18, VGE15b, Wan14a, WL15a, WLLL17, Wan17b, WK11a, YZ14, ZZ13b, ZZM15, ZZ19b]. ratios [MKA19, VRR13]. Rayleigh [Sch16]. Razumikhin [GA12, LLK^{+}18]. RCA [LTR13]. RDEU [CH19]. Re [IK14, Van18]. Re-evolution [Van18].

Re-formulation [IK14]. reach [Mak10a]. Reaching [JS18]. reaction [Sab18]. real [AKM19, ANV13, IT14, Red17]. real-valued [IT14]. realized [DS14, HS13b]. really [BH17b]. reassessment [Azr12, Tia16]. recapture [LZL^{+}19, McC12]. receiver [BH15]. reciprocal [Fak17, IK14, Mai18]. reciprocals [Neu13]. reciprocity [Lou13]. recognition [LSSR13]. recognition-memory [LSSR13]. record [AMB10, BABF19, BE17, CM16, FCP18, JS19, Kum15, KN10, Qiu17, TE17, WYC15]. Records [GS13b, NS14, AB10, Jas19, PS16a, PS19a, RA13b, RA13a, Sha18, SBN16]. recovered [Mna11]. Recovering [GT16, LCJ10, PCGV14]. Recovery [MS18a, McC12]. rectangular [Zen17]. recuperation [Gat15]. Recurrence [Bór18, BL14b, KFHS11, WL13]. recurrent [AaM18, sS15b]. recursion [Mar12a]. recursions [GHH^{+}10]. Recursive [CF11, ZJZ16, AI17, BS18b, BS19c, KP14, KKW14, Zha12]. redistribution [HH18]. reduced [BBdWG16, Han18, KZ14]. reducibility [BSCQ17, BSCQ18]. Reducing [BMD15, CFS12, MMF17, Yat16b, Sch12]. Reduction [Che18c, AT15b, CZ15, DZ18, Gho11b, MM13c, MM14c, Nos19, Shm13, Vie18, Yat16b, Yoo15]. redundancies [BZV11, Zha18]. redundancy [Deb12, MM11, VAZB10]. redundant [LFM15]. reference [HP12b, MDPP16]. refined [BB17, BMNZ18, Zhe17]. Refining [Mül19]. Reflected [HLW10, LE19, Xu12a, Yan19b, ZR12, AM13a, BY12, FRZ10, FKZ15, KM13b, Lee12a, LL12a, LL12b, LS16b, XXY12, Yan19a]. reflecting [SG15a]. reflection [Öns13]. reflections [AEFO18]. Refracted [MSPSJ17, ZW15b]. regarding [QB14]. regenerative [AR14, SK17]. regime [BGO18, DYW14, Hie14, HH16, Hua16, KH18, LWD13]. regime-switching [DYW14, HH16, KH18]. region [BKL19]. regions [BABF19, BK19a, CM10, DM12c, HEM10, KS15b, KR13, MM15, Ruk16]. register [Kao13]. Regression [JP12b, MKJ15, MFL15, ACW10, AR13, AH13c, AH11, AZ17, BL10a, BLB16, BSK19, BCL18, BZ13, BBS11, BRZ10, BV15a, BTT16, BES18, BS19c, CL19a, CZ15, Che12b, Cho16b, Ciu15, CHM16, DN14a, DNLR17, DPZ16, DE13b, DFKK12, DWW17, DLZZ18, DSX13, DZX18, Eck18, FA14, FK12, GBVS11, Gee15, GS16, GLS13, GZ18, GK12, GJ10, GZWW15, GW12b, HM10a, HBL11, HLB17, HS16b, HS12a, HC10a, HCT16, HYWT17, HS10, HWZ14, Huh12, IR11, JM14, JQZ12, JH11, JH14, Jur10, KST18, KA10a, KO17, KLM11, KM12a, KFHS11, KKW11, KK15, Kre18, Kur19, LBG11, LC10, LNC12, LWX11, LCX12, LQZ13, LBS17, Lia12a, LWL16, LB17, LLZ13, LZ13, LYC14, LW17b, LK14, LR12b, Lu16b, LA13, LG15b, MBS13, Mes10, MWZ11, MMC14, MSW15, NT14, OPS16, ÖA15, PK11]. regression [Par17a, Par17b, Pin14, QO16, ROSL17, Rei15, SY19a, SA12a, SZW12, SX13, SCS14, Shi14, SY11, Sri15, SZ13b, TL12, TZW12, TSZ15, TSL18, TY15, TZT16, TSK13, TPNSN19, VB11, Wal10, WMZW13, WMFW15, WY16, WCM11, Wes15, WZ17, Wis11, WL14b, Woo15, WS15, WWH17, XJW15, XXY17, Xu12b, YH14, YJLL16, Yao12b, YW17, YGTT13, YDG13, ZDM17, ZL16a, ZM16b, ZX17a, ZCZ18, ZKG10, ZW11, cZgL12, ZWT15, dML12, dB13, CGH11]. regression-based [Lu16b, SA12a]. regression-type [BZ13]. regressions [BG11, SH13, WLS11]. regressive [LYB13, dSGPM12]. Regressor [WZ16, DNKL12, LW17b]. regressors [KV13, LGT15, ZGK13]. regular [AC11, ET19, LZZ15b, Nog13]. Regularity [LM17, CP19, LTvdVR11, SS13a, aZLW18]. Regularization [PT10]. regularized [HHY11, NTT18, TSL18]. Regularly [BBP18, OV15, Alb18, FS11, JT11, YLS10]. regulating [MSPSJ17]. reinforced [FP12]. reinsurance [LZY15, LY18b, LH14b, ZZZ13]. rejection [BH14, Gri19, HA14]. rejections [Gor12]. related [AEFO18, AI17, Dun06, DF11, Had11, Hil14, HC10b, Jov19, Kao14, KM13c, LS12a, LZ16a, LW12, LW14b, MB16, MR15b, Osȩ12b, Pin13, Pin15a, Ren13, RBY10, Roh13, Sau12, SI16, Vin11, Zha12]. relates [Osu12]. relating [HNW19]. Relation [BVP10, MW14b]. relations [Can16, Kin10]. relationship [Jon19, LCKK14, Liu11]. Relationships [Jon12, HK19, NS10]. Relative [Kru17, MF15, ZBZ11, ABGM16, CZ17a, FD18, GR10, Maa19, ML16, Nie19, SY19a]. relativity [McK15]. Relaxation [Kar11, SHK14]. Reliability [BNM13b, KHN16, CC16, Ery10, MM13a, NSH17]. remainder [Fro15, He12]. remark [AJ14, CC10a, CH11a, CJ14, Lê16, dC13]. Remarks [BDP10, CY11, DP19a, Jur14, Jur18, LTW18, Lon13, NS13b, Rao12, LD10, LGT15, Sau12, Vid18]. REML [LP13]. Removal [HR19a]. Renewal [GF11, Gut11, AB17, BS15c, BS15e, CF13, Chr12, DZD12, FN17, GZ15, GW16, GSW11, GS10, Had19, Hew17, IMM16, IKM16, JCM15, Kol17, Li17a, Li18, LP17, LZ16b, MO14, OV15, PNT15, Rat17, SKJ15, SXM16, Spa10a, Sta12, Ter17, XX13, YL19, vL16]. renewal-reward [AB17, PNT15]. renewals [Ye16]. renormalization [Pog16]. Rényi [CD10b, FA12, GGR11, LT14, NSS14, YC16]. repairable [NK11]. reparameterized [TPNSN19]. reparametrization [RHP14]. repeated [BK19a, CFS12, CM11b, PR17, Rei15, Ros13, cZgL12]. repeated-cross-section [Ros13]. repeats [MPA15]. replacement [CRG17, HRB19, SKJ15]. replicated [CKF17]. reporting [AG13]. Representation [ET19, KK11, LY17, Muk17, Vii16, Yaz15, ZL15, AFJ11, BIK12, Bou17, BES18, BdS13, EK10, EKS11, HZ16b, HZR12, KKNH18, KKT14, Lau10, MS10a, Ren13, SCZ15, TMS19, TE12a, Wan17a, XF17, ZFZ13, ZF13]. representations [BDP14, BW18, Dom13, Fis13, PT10, SKRT16]. reproducibility [DD11]. reproducing [PK11, Par17a]. reproductive [XW10]. required [MIL16]. resample [CP11]. Resampled [Vil12]. resampling [RW16b]. resampling-based [RW16b]. rescaling [LR12a]. research [Sha18]. Residual [JS19, AEAN18, CF18, CN16, DK14, GW12b, HTA10, Has10, KI15, KRB13, KT11, KN10, Li15c, LP19b, MKJ15, Nan10, NSS14, PRS12, PK14, PP15, PB13, PSS16, QJ18, RP19, Raq10, ST19a, SKJ15, TJE16, TT15, WC17, WCD19, ZY10a, ZJL13, ZJ15]. residual-based [Li15c]. residuals [Bro07, Bro11, DFKK12, LK14]. resistance [Mis13]. Resnick [EKS11]. Resolution [Che16]. Resolvable [YCL14]. resolving [GR10]. respect [Bry14, Ery11a, Gna12, HP18b, JV10, SC12b, WL15a, ZJZ13]. respondent [AC15]. respondent-driven [AC15]. response [BB16, BB10, BB11c, CL19a, DSW13, HW13b, KA10a, KH18, LSS13a, MP17, Pru19, SS11a, ST14, SHBHD11, WXL19, XFC17]. response-adaptive [BB16, BB10, BB11c]. responses [AR13, AH13c, AZ17, BBGMPG15, GXH17, KK19b, MBTC13, SK11]. Ressel [Vin11]. rest [HLLK13]. restaurant [Mil19a]. restoration [Jov14]. Restricted [WZW17, BBCD^{+}18, DP10, KJS15, MW14b, Ren14, Ren15, RBSB16, ySDM14, TM16]. restriction [Zhe11]. restrictions [HH19, RS14a]. result [Bor14, BH10a, Bre12b, EJ11, For14b, Lê16, PLY18, TE12a, Wan15, Xue19]. resulting [YZ13]. results [AA11, AS15, BD13a, BM17, CK16a, CF13, DL13b, FCP18, Had11, HNW19, HLCH19, Isl16, JDB12, KT19, KT11, LQ18, LS15, LBH11, LZ16b, MNX13, MM12, Pin13, Pin15a, Res11, Roh16, SXM16, SD14, SBN16, SA13, TJE16, VS13, VM18, XH15, ZB11]. Return [Neu17, DK11]. reverse [CG17]. reversed [BZV11, BT14c, HP14, LDZ10, LL19b, MM13b, WYC15]. reversibility [AG18]. reversible [CH13, Jia16, LPS15, WC18]. reversion [YE10]. Révész [WX19a]. revision [Res11]. revisit [Li18, LD16b]. Revisited [ASH18, FD10, Gzy19, Xia19b]. Revisiting [TV17b]. reward [AB17, Ery14, PNT15, SG14]. rewards [AB17, BS14b, Hew17, KK19a, PNT15]. rho [PPA16]. Richter [DF14]. ridge [BTT16, Ema15, Luo10, Luo12, LK13, Par17b, TSK13, vW17]. ridit [BB16]. Riemann [Yas19]. Riemannian [Pei15]. Riesz [KLM17, LM15b]. right [ADP12, BMN15, BMN16, BG11, Che19b, Fil12, HTA10, LZ14b, MZZ15, MYA15, NP10, QB14, sS10a, SCZ15, SMZ18, XSZ17, ZZ11, ZBZ11, ZJL13, ZZM15]. right-censored [Che19b, LZ14b, MZZ15, SCZ15, SMZ18, ZJL13]. right-neighbor [Fil12]. right-truncated [BMN15, BMN16, sS10a]. Rio [Bob13]. Risk [Ber13b, CS16a, SWH12, BL17, Ben17, BZ13, BG11, BTT16, Bra12b, CH17, CV19, CR16, DZD12, DYW14, DZ19b, Elh14, FS14, FN17, GR17, Gat15, Han18, HP18b, JYL12, JCM15, LR13a, Li17a, Li18, LGW12, LXZ15, LZJ17, Lu11, Lu12, LZ16b, LPH13, MZZ15, Mac11, MWH17, OS18, PLL13, PL17, PP11, PS11a, QY15, QO16, Sha17, SZ12, SZ13a, SXM16, SLZ13, WY10, WYL10, WMH11, WZ15b, WWPY18, WM18, WH14, Wei19, Yan14, Yan15a, Yan17a, YW10, YL19, ZY10b, ZW12, ZDD18, ZYG17, ZZZ13, HM15]. Risk-minimizing [SWH12]. risk-sensitive [Wei19]. risk-switching [GR17]. riskless [LY18b]. risks [ED14, Gho12, HK14, HH18, SNS10, Sha17, SZ13a, WY10]. risky [FS12, FN17, LPS12]. Robin [AMR19]. Robust [AHN10, BBS11, BBGMPG15, DNT10, GGV14, GA13, KS15a, KN11, LBS17, LLK^{+}18, LZYW19, Liu17, LZY14, SCS14, XD15, ZPZ18, ZZLH15, ZM16b, BCL18, BDM17, BRZ10, BCMR13, BV15a, CHN14, KV17b, OO11, SB10, Tan19, YGL14, YY16]. robustified [LRL19]. Robustifying [TKO12]. Robustly [YHMM15]. robustness [Azr12, GB19a, jLL19, Sch16, TR12, ZGR12]. ROC [Wan12a, WZ15a, WSU11]. role [BBM18, Cer18, CM10, LAG^{+}18, LP19b, Qua18, San18, Sco18, Sec18, Sha18]. room [BHJ18]. root [Arv17, BG13, BDvdAW16, CK13, CS16a, CP11, CY15, Han12b, JW12, Mar11b, RR14, Yab17, ZL14, ZP19]. root- [CP11, RR14]. rooted [SY10, SHK14]. roots [AR18, YWY10]. Rosenblatt [BT14a, CSY12, MT13]. Rosenthal [Pin13]. Rosenthal-type [Pin13]. Ross [FS18]. rotationally [LT19b]. rotations [Rau13]. rough [BJQS16, OSW18, Xu12b]. round [CLQ12]. roundoff [Shi12]. rowwise [CYH16, QG13]. Ruin [MP16b, PH10b, SZ13a, ZW12, AH12a, BS12a, BL17, CR16, FN17, GR17, GL13, Gat15, Kat13, LR13a, Li17a, Li18, LY18b, Lin19a, LZJ17, Lka19, LZ16b, OS18, PL17, WY10, WYL10, YW10, YL19, YSZ19]. rule [LLP11, Tsu10, You17]. rules [CCFL18, DMC17, JHW18, TYK18]. run [BS14b, Bra16, CCFL18, Ery11b, MIL16, Nov17]. running [ANV13, CP19]. runs [AMP16b, Ery12, Ery15, Ery18, IA10, KPO13, LY16, MPA15, Ste11, UK18].

S [LYCK11]. safe [KR13]. Sahalia [Dun16a]. Saigo [KV17a]. Sale [Cho16a]. same [AJ14, Fre16]. Sample [HHY11, Xia17, BABF19, BMD15, Bao12, BLB16, BNB16, DHOV13, FS15, Fre14, GO12, Ili12, Jia13, Jur10, JS11, Kab16, KM13a, KHHD16, LSS13a, LL15, LZ14b, LBH11, LG16, MGA11, McE12, MMM13, NKKY13, Nku11, PH13b, SÖV14, SLC16, Thu14, TLNO11, Uem16, WD12, WN19a, WN19b, Wri17, WR10, WT14, Xie13, YH17, YK11, ZG19, ZJZ13]. sampled [Gir16]. sampler [Che12a, Fit14, Hel19, RHP14, SC15, WR18]. samples [ABBK15, AV13a, BK19a, GO12, Gau14, LPNW14, MVI11, Now19, PP18a, Ruk16, SK15, TJE16, TV17b, Yak15, YZ14, ZLZ16, ZL16b]. sampling [AWL17, AO12, AC15, BK18, CRG17, CXW13, DRZ19, ER17, FGA11, Gri19, He14, HCP19, HRB19, JDB12, KM11, LZYW19, MW14a, MW17, NLF11, Ohy13, RMPB17, Ski15, TLNO11, ZV15, ZM17, ZDX14, ZY11, PRS18]. sandwich [Sri15]. Sanov [WWW10]. Sarmanov [MP16b]. SAS(R) [SWH^{+}11]. satisfy [Osȩ13a]. Saunders [FZB16, FLRS13, IK14, LC10, LCX12, TPNSN19, ZB15]. SCAD [FK12, GZWW15, Lia12a]. scalable [Ale18, BBCD^{+}18]. scalar [DDZ15, FS17a, SKRT16, Tah14, Won13]. scalar-valued [FS17a]. Scale [KAVRY17, AB10, BKL19, BSK19, BRZ10, CYW12, CXW13, DKT19, GMA12, GS12, Han10, HEM10, HP15, KLM17, Kim08, Kim13, KV17b, LBG11, LLC^{+}12, LZL13, MGA11, PSS13, RR10b, SAM13, TL15, WN13a, ZPK14]. scale-like [CYW12]. scale-space [HP15]. scaled [AA19, CQ17, SK15]. Scaling [Yan19a, ZZ13a, NP16, vWvZ17]. Scan [ZG16, ASN16, HP13, WZG14]. scanning [ASN16]. Scatter [Nag19, Cha19b, FG12, Ilm13, Tyl10, Vir16]. scenario [CH17]. scheme [Alf13, Bha17, Cal13, CK16b, LT19a, Lka19, MDR12, PN12, WK11b]. schemes [BT14b, Kre17, LZ10, WZ13b, ZW18c]. Schmidt [Cui14]. Schnabel [GG13]. Scholes [DF14, GM16, SS17, Wan19]. Schott [CQ18]. Schur [SKRT16, TV17a]. Schur-constant [TV17a]. Schwarz [Lia12a, Wal17]. science [Bow18, Cer18, Giu18, Men18, OW18, Rei18]. sciences [DvH11]. scientist [Qua18]. score [CM17, Gho11b, GXH17, TYNZ15, YZ16]. scores [KD14, MP15b, Yat15]. Scoring [TYK18, DMC17, HM10b]. Scott [CRB19, Jal16]. scrambled [SK11]. Screening [ASK18, CZCL19, KS17a, LSCL17, WA17, ZPZ18]. SDE [KS16b, Yam15]. SDEs [AM13a, GA12, LY18a, LT19a, LW15c, Pam18, QW16, SG15a, Son16, SG18, Tah14, WZ19, XWML19, ZZ19a]. seamless [Ciu15]. search [Cha19b, KN17, LLP11, Sak17, SD11]. seasonal [DEM14, DPP16]. secant [FQZ18]. Second [Hwa18, KD12, Lin19a, PL15b, RSC10, WLP17, WL17, GGQ17, He18, Osu12, PLH13, PTT19, SS18, SL13b, Su10, Tak12, TMN16, dWGM12]. Second-order [PL15b, RSC10, WLP17, He18, PLH13, SL13b, Su10, Tak12]. secretary [FK10]. section [Ros13]. sectional [CFS12, KZ17, Roh16]. sections [Wys12]. seemingly [WLS11, ZX17a]. segment [AKV12, PH10a]. segmental [MP15b]. selected [MT17b]. selection [AZ17, AV14, BG19, Cha10, Ciu15, DMC17, EL18, FA14, GZWW15, HL13b, JdU16, JBS10, JS18, JQZ12, Jia13, JD17, JHW18, JT14, KM18, LML15, LD16b, LG13, LG15b, LZY14, Mao15a, NMS15, OPS16, Pac18, RW17, RZ10, RG18, SD11, SLM19, TZW12, TXX16, WL14a, WZ10, ZJS13, ZYG17, dML12]. selections [BS19a]. Selective [RG18, LD16b]. selector [BD13b, dC13]. selectors [DFH18]. Self [CP13a, Fan17, Fen17, MSPSJ17, CD18b, FP11a, GM18a, Jon10, KL17, KC19, Kob16, KP10, Min17, PP18b, SY17, Vii16, Wal17, XW17, Yaz15]. self-adjoint [Min17]. self-affine [KL17]. self-decomposability [KP10]. self-equilibrium [GM18a]. self-exciting [FP11a, KC19]. self-improvement [Wal17]. self-intersection [SY17]. Self-inverse [CP13a]. self-normalization [CD18b]. Self-normalized [Fan17, Fen17, Jon10, XW17]. Self-regulating [MSPSJ17]. self-similar [Kob16, PP18b, Vii16, Yaz15]. Selfdecomposability [CM11c, BJ14]. selfdecomposable [DZ19a, MU13]. Semi [DLZZ18, LXL19, MR15a, SG14, AA19, DMS13, Dun19, ELS13, FJ12, Fan16, Gho12, HH18, ST10, Vas18, Yeh11, ZH12a, ZC12]. Semi-additive [LXL19]. semi-competing [Gho12, HH18]. semi-continuous [FJ12]. semi-exponential [ST10]. Semi-functional [DLZZ18, ZC12]. semi-group [AA19]. semi-linear [Dun19, Fan16]. semi-logistic [Yeh11]. Semi-Markov [SG14, DMS13, ELS13, LXL19, ZH12a]. semi-stable [Vas18]. Semi-strong [MR15a]. Semicircle [FG12, Löw18]. semicontinuous [QN15]. semifoldover [Yan13]. semigroup [HY13]. semigroups [Wan14b, ZJ13]. semilinear [FS17b, Gov15, GLM19, Son14, Zha13a]. semimartingales [AM17, LSS13b]. Semiparametric [Che19b, LLL13, sS10b, SZ13b, ACW10, Ema15, HZJ^{+}10, JL16, KS17b, Luo12, MK18, RR14, SZW12, sS10a, SB10, WZ15a, WL11, WS13, WYH14, WCM11, XZ13, YL10, YY16, ZP15]. semisimple [AMG17]. semistable [KW15]. semivarying [ZZLH15]. sensitive [ST14, Wei19]. Sensitivity [KT10, MDN18, ROSL17, TL15, TR12]. separability [MD13, RT15]. separation [Bow17, Fra11, KO17, MNOT12, TMN16]. sequence [ANV13, AMP16b, Bis10, CHR16, EGX16, Ery18, GZ13, HPW15, Izu13, Jou14, KPO13, KK19a, Liu12b, MPA15, Osȩ15a, PB15b, SM10, Sch13, VY12, WHYL10, YCQY16]. sequences [Bha13, BN10, hCyP15, Dav12a, Dav13, Eks14, EB17, Gut14, GS17b, HP13, Kao13, Mak11, Mat14, MW12a, PP19, PTW19, Rob10, Sze11, Sze16, XW17]. Sequential [BY12, KN17, Tim15, ZWT15, Buc15, BT14c, Gou15, Mih12, RSdB15, RR13, RSV10, Uno13, YK11]. serial [BL12, McE12]. series [ANV13, AA13b, BM11, BG15a, BG15b, BMB16, BM18, BZV11, CP11, Che10b, CTW10, CM17, CC10b, CZ17b, DD12, DL17, Efr14, FZ13, FZB16, Goi12, HZ16b, Hu15, KKP15, Kre18, LML15, LS12a, LX12, LWL16, MNO15, McE12, MNOT12, MM12, MMC14, Neu13, PS11a, PPN10, PH13b, Roz10, SM16, Spa11, TMN16, Tri18, VAZB10, WZG14, WL15a, Woo15, Wu14, XKBG15, Yan15c, YSZ19, Yao12a, vWvZ17]. service [Bri15]. Set [CH17, Sto12, AWL17, AO12, BFS16, Ber13a, Bon12, Cas10, CXW13, CCS17, ER17, ET19, FM13, GO12, JDB12, Khm13, KB15, LLC^{+}12, LTX12, Spa10a, TJE16, ZV15, ZM17, ZDX14, ZLZ16, ZWLB17]. Set-valued [CH17, Bon12, Khm13]. Sethuraman [Mil19a]. sets [AP13, AD11, BS15b, Che14c, DKM19, HMT13, KL17, LBM14, LCJ10, LL12c, LVMS14, MS11, Nag17a, QG13, Söh10, Tyk11, Tyl10, VRR13]. setting [BG13, CDG15, Rév12]. setup [Ery10, MBTC13]. seven [LSS13a]. several [BPRS13, KHN12, KK13a, LSS15, LLH15, MKK17, SAM13, SGG10, Zha14b]. severity [LR13a]. Shannon [DYB11, KK13a]. Shape [DP10, Zhe11, BH15, CCSC11, JS15, LS12b, Tak12, WL15b]. Shape-restricted [DP10]. shared [HS12a, HD13, HP14]. sharing [AJ14]. Sharp [AG13, Aya13, BOR17, BO19, Osȩ10b, Osȩ11, Osȩ14b, Osȩ15a, Puc13, SD16, VA16, ZZ13b, BK13, Efr10, Efr19, Gor12, LZW11, OS15, Osȩ15b, Ste19]. Sharpe [BWW11]. Sharpened [AK10]. Sharpness [BH19]. sheet [LW11, WYY14]. sheets [BSS13, BS15b]. Shepp [HT13b, NZ19]. shift [DKT19, Kao13]. shifted [BS15b, KK13a, MM13c, MM14c]. shifts [HH13b, KLNB14, WL14b, Woo15]. Shiu [Bra12b, DZD12, LXZ15]. shock [CH11b, Ery12, Ery15, GMS18, KZ17, LW14a, PB15a]. Short [BT16, GBGPR19, KS17c, PSC12, SM16]. Short-range [BT16]. Shorter [Fre14]. shortest [CC16, HNGS15]. shortest-path [HNGS15]. shortfall [Puc13, XZZ19]. shot [HJS11, IKM16, LW14a, ST10, Wan14d]. shot-noise [LW14a]. Shreve [For14b]. Shrinkage [AF12, FA14, Lia12b, Nku11, Par14, Tsu10, WW13a, CMS19, HS11b, KMS15, MS16a, TA14, WW10]. shuffles [DFS10, DS12]. shuffling [TMS11]. side [WM18]. sided [AP18, LZZ15a, LT19a, LP17, Mel16, MW12b, PH10b, Vid17, ZK15, ZY10b, ZWB17]. Sieve [SZW10, IJB17, LHT15, RS12c, YW17]. Sieves [BB18]. sigma [Fis13]. sigma-algebras [Fis13]. Sign [ZDX14, DTV16, LP10b, TKO12]. signal [DE13a]. signatures [MM13a]. Signed [DS18b, JW12]. significance [MWGT17, PP14b]. significant [LCJ10, ZGK13]. Simes [BSWS13, Llo12]. similar [Fra15, Ker16, Kob16, PP18b, Vii16, Yaz15]. Simple [CF13, MW14b, AN10, AO12, Alf15, Bap11, BH14, CZCL19, Chi13, Cho16a, FPZ13, Fis13, FMPV12, Hay12, HH13b, Joh14, LSS13b, LT14, MD13, MWZ11, Nic17, Pet11, RR14, Wan11b, WH10, XX13, ZZ15a, Zho13b]. simplex [BH17b, PV17, Pin12]. Simplicial [Van16]. simplified [HZR12]. simulated [RR11]. Simulating [DF10, KL11]. Simulation [PS16a, Tur14, BLWZ11, BMS16, Giu15, LW16, MW16, Mai18, SdOG12, Tau15, ZA12]. simulation-based [LW16]. Simultaneous [AZ17, CD18a, CS17, Ger18, Gre11, KKMSA13, MGG11, RM13, SS13b, YXL14, Zha14b, Ben18, HEM10, LSS15, NW10a, NHBG19, TW17, Yat15]. single [ALS11, BM10a, CYH16, FZW12, JLY14, KHN16, LLL13, LZLW16, LSS13a, Leh15, LX12, MGSL11, NSH17, WW13a, WL14a, YGL14, YXL14, dML12]. single-birth [JLY14]. single-index [FZW12, LLL13, LZLW16, WW13a, WL14a, YGL14, YXL14]. singly [KS15c]. Singular [Xu18, LTW18, Shu18b, Uem16, Xu13, XWML19, YHY16, YKY17]. singularity [JS16, Yoo15]. SIR [Cla14, El 13]. SIRS [LO13]. site [Li11b, SvdV17]. site-dependent [Li11b]. Sivashinsky [WWW14, WWX17]. size [AC11, AV13a, BMD15, BZ13, BN10, CP11, CM16, GGR11, Gou15, HHY11, Jac10, Jac13, KM13a, SXM16, TLNO11, VSP13, Xia17, YK11]. size-dependence [BZ13]. size-dependent [SXM16]. size-estimation [VSP13]. sizes [AWL17, BS12a, BP19c, DS15, Gra11, JKD15a, JKD15b, MMM13, PH13b, Sha13a, Sta16, WCZ18, ZDX14]. sketches [XLL19]. Skew [HA13, AA19, AN15, ABGM16, BS15a, BPXS19a, BPXS19b, Ber13b, Car10, DN16, Fra11, FS10, JSA12, KRYAV16, KAVRY17, Kim08, Kim13, KA11, LBG11, LTW18, LPN13, LPNW14, Mah18, Mam15, ML16, OH15, PLY18, PLN16, RL14, Shu16, Shu18a, Shu18b, VM12, XW16, YHY16, YKY17, vSK15, ORO15]. skew- [AN15, Car10, PLN16, VM12, ORO15]. skew-elliptical [Fra11, Shu16, Shu18a, Shu18b]. skew-Gaussian [Mah18]. skew-normal [BPXS19a, BPXS19b, JSA12, LBG11, LTW18, LPN13, LPNW14, Mam15, OH15, YHY16, YKY17, ORO15]. skew-normal-Cauchy [KRYAV16, KAVRY17]. skew-symmetric [RL14]. skew-Wiener [PLY18]. skewed [BCGFR13, BVV14, GM19, JB13a, KM13c, MYA15, NA13, RS14b, SWW15]. skewed-SVCHARME [RS14b]. skewing [LP10a]. Skewness [KM13c, Lop13, AN15, EJ12, LC10, Lop15, MBS13, Oga17]. Skewness-kurtosis [KM13c]. skip [Zho13b]. skip-free [Zho13b]. Skorohod [BS15d, PS17]. Skorokhod [Bak12, DD12]. slab [IR11]. Slash [dC16]. SLE [MZ19]. Slepian [BG16, Den17, DZ17, LHM14, NZ19]. Sliced [CYLL19, AT15b, HA12, WZYL17, YCL14, ZJL18]. SLLN [BDP10, QG13]. Slow [Jon17]. Slower [BP19c]. slowly [Alb18, IKM16, LLSW11]. Small [BLKL13, Chu14, Gas10, Kob16, MY14, PN14, Roz14, Roz16, Vol14, ZYL11, CW15, FA18, JS11, LBH11, Ma10a, PP18b, Ste11, SZ18, Yam15, Yan17b, ZYS19, ZZ19a]. smallest [BNB16, BKL19, BK17, HLCH19, MKA19]. smile [For14a]. Smirnov [Fre16]. smoking [LOKB11]. Smooth [CDL11, Tah14, Wan14b, XYZ16, CMT14, PS16c, PS19b, Son10b, Tan13a, YW17]. smoothed [LLHW10]. smoothers [CDG15]. Smoothing [Gor14, AT14b, Bow13, CSG11, Lem11b, Lem11a, PJ13, ZKG10]. smoothness [SY17, Son12b, XLL19, vWvZ17]. Sobolev [CW19, DS11, DLM15, Efr10, Fan15a, IMPR16, Pam18, Pei15, RSR16, Wan17a, WX19b]. Sobolev-type [RSR16]. sojourn [BM12]. Solution [Par17a, Bak12, BJQS16, Jia16, KS16b, NTT18, QX13, Son12a, Tah14, Tie13, Wu11a, YYS16]. Solutions [JY15b, BL14b, BL15, ÇKN17, DDE17, Dun16a, Fan16, GV16, GA12, GA13, Gov14, HY13, LL17c, LS15, LFM14, PKK14, SG15a, SS13a, SG18, TJD10, TJS13, THC17, WX19b, XL16, Xu12a, ZT19, ZBL19, aZLW18]. Solving [LS16b, LZ10, Sab17, Sab18]. Some [AA11, BR17b, CLQ12, CYW13, CKK18, EAB10, FCP18, HTA10, HWYW12, Jov19, Kak16, KT11, LVY15, LYW11, LQ18, Liu11, LD10, LW14b, LGT15, LZ16b, MWZ11, NR12, NS12, OQ10, Pan19, PJ10, Pin14, QN15, Rob13, Roh16, SY10, Ski15, SA13, TV17a, TJE16, Tan15b, Tyu12, Ush11b, VAZB10, VM18, WW13b, WMM15, ZJL18, Zha12, BLL10, Bor16, Buc15, CF13, CJT10, CF11, DDE17, Din15a, Döh14, Dun06, DF11, Gho11a, GM18b, Gri19, GB13, HP13, Jur14, Jur18, KT19, KR19, KV17b, KS16b, LV11a, LV11c, Lin17a, MP15a, Mac16, Mah18, Mat14, Min17, OT19, Pam18, RS10a, Res11, Spa11, Tom16, Tru10, VS13, WJY11, Woj13b, YW10, ZL12, ZLT11, ZB15]. Sooner [KK19a, EGX16]. soup [BK18]. source [HH13b, MNOT12, TN17]. sources [KM18]. Space [RS14a, BJQS16, BH12, BH17c, Bud14, CQT12, DD12, HP15, KD12, KV17a, KO17, Kin10, KR19, LL13, Ma13, Mia14, MX18, OP12, PK11, Par17a, Pei15, Rao10, Rob15, SB19, Sha12c, ST13b, TYNZ15, Tyk11, Vil11, Wu11a, ZKSK18, ZH12b, ZHL13, aZLW18]. Space-filling [RS14a, MX18]. space-fractional [OP12]. space-time [KV17a, KR19, Wu11a, aZLW18]. space-valued [KO17]. spaces [ÁLBRM16, CV19, DT10, GSK12, HY18, HT15, Hir11, HY13, IMPR16, Ko18, Lan18, LVMS14, NS13a, Osȩ12b, Osȩ12c, Osȩ14a, Pet11, QN15, STD12, SHK14, Suz18, Suz20, TMS19, Vid18, Yu11, YY14]. spacing [Ste11]. spacings [AA11, PS19a]. Sparre [ZY10b]. Sparse [AD11, YJLL16, CSG11, Ger18, Tyl10, XXY17, ZIPP17, Zha19a]. Spatial [AR18, ASN16, BL10a, BM10a, BD17, BK18, BS18b, BS19c, DNT10, DNKL12, DWWG18, DDZ15, DTV16, EFP19, GDNM18, Kur19, LYB13, LCW17, Mar11b, Nag17b, SV15, TKO12, ZZ12a, Zha17b]. spatial-temporal [ZZ12a]. spatial-type [SV15]. spatially [DFH18, Tzi17, You17]. spatio [CG18, IL17]. spatio-temporal [CG18, IL17]. SPDE [AM13a, Dun19, GLM^{+}13, Zha13a]. SPDEs [BJQS16, Gov14, GLM19, Son14, XY17]. Spearman [PPA16, WWW19]. special [MDPP16, McK15]. species [MHH11]. specific [Wan17b, ZIPP17]. specification [LD16a]. specificities [TL15]. specified [Che10a, SD16]. spectra [HS16a, MTT11]. Spectral [JH14, Roy12, RM11, Zen19, BW15, BBGH12, BSCQ17, BSCQ18, BVP10, CKP17, CP14, DWWG18, Din15a, Efr14, HL16, HZ16b, JP12a, JW17, MW18, MvdBV^{+}11, PH13b, Xie13]. spectrally [Coq15, DZ19b, Gat15, LZ14a, LZZ15a, LT19a, Mic11, WWPY18, ZDD18]. Spectrum [HS16a]. speed [Hua16, Lon13]. sphere [CG13, HZR12, HZRS19, Rau13, VY12]. spheres [DLM15, EFP19, Gzy19, MW18, TYK18, Sab18]. Spherical [ZZ16, CQ17, DLO16, LX18, MV15, Shu19a]. Spherically [Duo15, FS14, MK10, Fin11]. sphericity [BKP15]. spike [BJD16, IR11]. spiked [CZLT18, LW17a]. Spitzer [CS19]. Spitzer-type [CS19]. Spline [LHT15, MMC14, ZC12, LML15, ZP12]. Spline-based [LHT15]. splines [ACW10, ZL16a]. split [Thu14]. splitting [PF17]. spread [Bow17, Pin17]. Spurious [BL10a, GBVS11]. square [Bou15, JLR19, LD17a, LFM14, Luo10, Osȩ12a, Osȩ14c, Osȩ17b, Osȩ17a, PRS18, RW16a, SKRT16, Wan13c, WSMG18]. Squared [VL19, GM14b, GG12, KN11, Ric10, SW12, WWD15, vW17]. Squared-norm [VL19]. squares [AAEH11, AT15a, BL10c, DMPX17, ES13, For18, He18, KLM11, LL10, Ma10a, Mao15a, SZW10, Wei17, WZ16, Yab17, YH14]. squares-based [AT15a]. St. [Gut14, GML15, Nak15, Nak17]. Stability [KO11, LZ17b, OK15, Ost13, TJS15, ZYL14, Ber13a, Che10c, CW12, CvGL18, CHW18, CD10a, CYS11, DEM14, DDE17, HB18, LOKB11, LF17, LD17a, LFM14, LTX12, PJ10, WWX17, WSMG18, XW16, YD12]. stabilizability [GA12]. Stabilization [QLH^{+}16, GA13, LLK^{+}18, YHMM15]. Stable [Arv19, GKW19, AT15a, AH12b, Bes14, CM11a, Coq15, DP19a, DJ14, Hof13, Iso15, KS13, KR19, KV15, LLP11, LT19b, MY14, Mic11, MM14b, Mse18, Nol18, Oes15, PR15, Pan19, PS16b, Rob13, SS18, SH11, Tau15, TK10, Vas18, WS10a, Wan14c, Wu11b, Xu13, Xu18, YYC17, Yan17b, ZZT17, ZYS19, ZZ19a]. stable-like [Xu13]. Staden [BS15a]. stage [BB16, Kab16, Uno13, VT11, Xue18, ZGR12, ZY11]. Standard [Mis14, Chi10, KA11, Mac16]. standardized [Fer18]. Standardizing [Lan18, JW12]. standby [Ery11a, HN15, KHN16, LFM15]. star [LS12b, Tru13]. star-shape [LS12b]. start [Ger13, HS19]. start-up [Ger13]. starting [Abu12a, Abu12b, Abu13, Pen11, Rat17]. state [CJY14, FL19, GA12, HMZ11, KD12, KK19a, KW14b, KR12, LLK^{+}18, Ma14a, Ma15, McC12, Mia14, MSPSJ17, QLH^{+}16, Vil11, Xu14, ZKSK18, ZZ15a, ZHL13]. states [SHBHD11]. stationarity [EFP19, Har16]. Stationary [HS13b, SH13, Ai16, Ane17, AD12, Arv19, BGT15, BBP18, BDvdAW16, BN10, BL14b, CDL11, Coe15, CP17c, Dal19, DHJT14, DL13b, DHOV13, HP13, HW13a, HS12c, Hwa13, Hwa19, IM15, Jas16, KKP15, LO13, LL19a, Lau10, Li13a, Liu17, MW16, MZ17, Mat14, MNOT12, PP19, PP14a, Pen11, PTW19, PH13b, Rah11, Rob10, Rob13, RS15b, Sze11, Sze16, Tah12, Tan13a, Tan15b, TMN16, TZB13, Vii16, VW14, WP14, ZH14]. Statist [Abu12a, Bai13, BSCQ18, BRBB23, DR15, DFL13, DFT13, Duc10, Kak11a, Kim13, MN13, MM14c, RA13b, Ren15, Sko20, SF12, Suz20, Ton17]. Statist. [GHR17]. statistic [DD11, Fan17, Fre14, GZ18, HHY11, JW12, Lem13, LYCK11, McC12, PTT19, RW16a, WZG14]. Statistical [BS15d, Bow16, Chu18, FFM18, GDNM18, MNOT12, PYS^{+}18, QO15, Rei18, VP12, ACP19, CG18, CCM19, CKK18, Gho11a, HH13b, KKNH18, MP15a, NMS15, PP14b, PD11, Roh13, TYK18]. statistician [Men18]. Statisticians [Cao18, Shi18]. Statistics [Bro11, BvdG18, DvH11, Dun18, Jam18, San18, Sco18, Abd11, AEAN18, ASN16, AMB10, AA11, AAC15, AI17, AV13a, BS17a, BNB16, BV15b, BT14c, Cer18, CQ18, CK16b, CH17, CD18b, Dem10, DR02, DR15, DFS16, DH18, EG11, EB17, Ery11b, FW17, Fer18, FMPV12, GD15, GD18, GL19, GGQ17, GMDP18, HP13, HTA10, HT13b, HBH10, HS10, JBS10, Jas16, Kak11a, Kak11b, KO11, KLW11, KKL18, KS17b, KS17c, KC16, LMW15, LAG^{+}18, LLC^{+}12, LLCG12, MKA19, MWA16, MP15b, MA11, Mir14, MM14b, Nag13, Nas12, NS10, Nav18, NS12, OK15, OW18, PK14, Pin12, PS11b, Qiu17, QJ18, Qua18, Sec18, SKB14, SKB15, Sha18, Sko19, Sko20, Sma15, SK17, Tab18, TT15, Tyl10, VB17, WN10a, WK11b, YZ14, ZG16, Sug19]. Steck [MWGT17]. Stein [GL19, For14a, FS16, GX15, Gau19, GG12, Han18, LVY15, LS16a, Osȩ17a, PS17, QO16, SK13, Shu18c]. Stein-type [LVY15]. Step [WX12, BK19a, JP12b, KZ14, Lin17b, LB17, MGSL11, MSW15, PN12, Vil11, Vir16, YWZ14, ZP15]. step-stress [BK19a]. Step-up [WX12, KZ14]. stepdown [RW16b]. stepped [LCF15, LTP18]. stepping [BW11, JWW14]. stepping-stone [BW11, JWW14]. steps [PRD13]. stepwise [Gor14]. stick [Mil19a]. stick-breaking [Mil19a]. Stiefel [IKAH17]. Stieltjes [Fak17, MT17a, OT19, Yas19]. Stigler [LCJ10]. Stochastic [BB11b, BH17c, ÇKN17, FZ13, FZ15, FZB16, KT19, LV13, Ma14a, ND12, PYK15, PVY17, RIK16, RSR16, ST19a, SC12b, TU17, WW11, XL18, XWML19, ZJ13, vBC19, AM17, AY17, AH13b, AT14b, ADH19, BBM19, BCD19, BGO18, BW11, Bob10, BS17b, Bos14, BH10a, BH10b, BH12, BH18, BS19c, CHM15, Cal13, CS15, Che10c, CW12, CvGL18, CD10a, CJN19, CJY14, Cri18, CYS11, DD11, De 11, DK14, DDE17, DO11a, FRZ10, FD18, FS17b, For11, GM18a, GP17, GLM^{+}13, GLML12, GA13, Gov15, HC10a, HB18, HLW10, JY15b, JWW14, JCM15, KW15, KZ17, Kin10, Kre17, LOKB11, LO13, Lan11, LV11b, LZ10, LPW10, LL12a, LL12b, Li13a, LL17c, LF17, Li17b, LWYZ17, LD17a, LL17b, LZ17b, LLK^{+}18, LZL13, Liu14, LFM14, LTX12, LR13b, LRH15, MX17a, MB15, MB16, MSS18, MM12, Nak13b, NT17]. stochastic [NL19, Osu12, Owo15, Pac16, Pac17, PJ10, Pol13, QX13, QLH^{+}16, RZ12, Rad19, RS10a, RH11, SS13a, SD11, SWH12, TJS15, Tap10, TJS13, Tim15, Tsi17, TH14, THC17, TW15, VAZB10, WWW14, Wan15, WWX17, WWZ18, Wei19, WSMG18, WS17, WH11, Wu11a, WWR12, WZ13b, WW18, XJW15, XL16, XZ16, YY18, Yan19b, YD12, Yu19, ZYL14, ZYS19, ZX19, ZX17b, ZR12, ZBL19, ZC13, aZLW18, dlCJB19, LHW19]. stochastically [Now16, Ren14, Ren15]. stochasticity [Bre14a]. Stokes [FS16, ZBL19]. Stone [BB17, BW11, JWW14]. stopped [BRO14a, DS16, GS10, Mak10a, Mak10b, SBA12, VPCG13]. stopping [Buo18, Den15, Fis13, LL14, Vid19, ZZZ18]. storage [XTT18]. story [Shi18]. Strassen [LZ16a]. strategies [Kin12, SW11, WWY12]. strategy [AHN10, AF12, Bra12b, DZD12, EQ15c, JYL12, Nku11, SYZ18, SLZ13, YY11b]. stratification [WLLZ11]. stratified [MH12, Nku11, Ohy13]. stratum [Chi10]. strength [AT14a, Ery10]. stress [AT14a, BK19a, Ery10]. stress-strength [AT14a, Ery10]. Strict [Li17d, KN15]. strictly [BL14b, Sze11, Sze16, Tur14]. strings [AMP16a]. strip [AHM19]. strip-plot [AHM19]. strips [LN19]. Strong [Alf13, Bao12, BL10c, BV15a, DZF18, DT10, Fak10, GFA10, GLM^{+}13, GS17b, HY13, Hor16, HS12c, KKL18, LX18, MZ17, Mih12, NT17, SX13, Wal10, XJW15, XY17, dS18a, AP18, AFJ11, AAC15, BT16, BJ14, CS14b, CYH16, CS16c, CS16b, CW11, Dal19, Eks14, Had13, Had17, Hu16, KBM15, KM17, Kor15a, LYW11, LLR18, LMLW15, MR15a, MKA19, QN15, QSS17, RMÁL19, SCZ15, Tab18, VS13, XY13, ZL18, AMMO13, DYB11, Nas12, RS10b, ST13b, Sun13b, Tor13, WHRY11]. strong-consistency [RMÁL19]. Strongly [GW12b, GW12c, BM18, LZ16b, Mis14, Tau15, TY15, XWT19]. structural [GSW11, HS15, NK17, TR12, YDG13]. Structure [BT14a, Liu12a, AE19, AV13a, BBGH12, Cal13, CPH12, FM13, GML10, GLML12, GBGPR19, HLM16, HZR12, KP17, MD13, MP16b, MN14, SH11, SK17, Vel12, WL14a, XX13, YLS12, cZgL12]. structures [DJM11, PPM18]. Student [AP13, AN10, BV10, BBP17, Mar12b, PT13, Wan11b, WY16]. Student- [WY16]. studies [BL10b, Che13b, Gho11b, LZL^{+}19]. study [ASN16, ADP12, Ben16, Cha10, GS16, KW14a, KHN16, MW16, Mar11a, Ruk12, SWH^{+}11, SdOG12, YSLL10]. Studying [BP15, MMZ17]. sub [Gau14, Hu16, SC12b, SXM16, Tur14, YS10, YKY17, Zaj19]. sub-exponential [SXM16, Zaj19]. sub-fractional [SC12b, YS10]. sub-Gaussian [Tur14]. sub-linear [Hu16]. sub-samples [Gau14]. sub-vectors [YKY17]. Subadditivity [HM15]. subcriticality [WLLL17]. subdifferential [LRH15]. subexponential [JWCC19, LW12, Lu11, LZ16b, RS12b, YLS12, YL19]. Subexponentiality [YS13]. subgroups [AMG17]. subject [Nku11, Rob10, sS10b, Wan17b, WN13a]. subject-specific [Wan17b]. subjective [GZSZ19]. sublinear [Fen17, Fen19, Ren13]. submatrices [CD12]. subordinate [BO19]. subordinated [BT16, FJW15, GM14a, Gri11, Pan19]. subordination [JMS17]. subordinator [GKW19, LY18a, MR15b, Mse18, Xu18]. subordinators [KV15]. subsample [MGA11]. subsamples [ZJS13]. Subsampling [BMP10b, CD18b, DL17]. subsequences [SD14]. Subset [RZ10, KM12b]. subsets [GS13b, WN10a]. subsignatures [Mar14a]. subspace [MK10, Vel12]. subspaces [Son10a]. success [AMP16a, Ery18, IA10, MPA15]. successes [Sch13, ZK10]. Successive [BH10b, KKMSA13, LL17c, MGG11, SDJ15, SLC16]. Sudoku [LLO14, LCLQ16]. Sudoku-based [LLO14, LCLQ16]. Sufficient [AH13b, BSO10, McC12, RS10a, AT15b, ASVY14, CZ15, DM13, Döh14, DZ18, Gei17, Nos19, Sma14, WWW10, Yoo15]. suitable [GMP19, KN15]. suite [SWH^{+}11]. Sukhatme [YC16]. Sum [For18, Wei17, AKP18, AA17, BS11, BV10, CLM12, HK14, JGW14, Kri18, LH14a, LT10, Puc13, Roz16, SWZ19, Sre10, Vas18, WS10a, Wei19, Yab17, Zha16, Zhe17, dC16, DD11]. sum- [WS10a]. summands [CLM12]. Sums [Lar15, AvHZ17, BS19a, BBH14, BM11, hCyP15, CC10a, CH11a, CS14b, CS16b, CL13, CD10b, Dal19, DS16, DFS^{+}13, Fen19, FD10, GZ13, Had13, HT15, Isl16, Jan18, Jir13, Jon10, KM17, Ko13, KA10b, Kru17, LV11a, Leh15, Lin19a, Lu11, Lu12, LMS17, Mas18, MD15, PR12, PYK15, Pan17, Pin15b, Roz12, Roz14, RR16, SW12, SXM16, STD12, SD14, Sze10, TT17, TY15, VPCG13, VS13, VW14, WWH17, XW17, Yan14, YLS10, YLS12, YLS14, YIS15, YWY10, YWC10, Zha14a, Zhu14, dML12, vLT11]. Super [Zha19b]. superadditive [EAB10]. superconcentration [Tan15b]. supercritical [Chu14, HA15, Rah11, Rah17, WGL11, Yao13a]. superiority [LH13, Sha14]. Superlarge [Roz12]. superlinear [Lin19b]. superlinearly [Zha13b]. Supermodular [KP15b, BP15]. superposition [Sta12]. superprocesses [HL12]. supersaturated [MK14]. supersmooth [Dat13]. supervised [GMM10]. supplement [CZLT18]. Support [LDLC18, CP17a, HR19a, Joh14, Pro13, Xu12b, CGH11]. supported [TNP17, YWC10]. Suprema [Bog15, MV15]. supremum [AD12, ADM15, Coq15, Fer18, FJW15, JLR19, KKW11, LPW10, Mic11, Pin15b]. sure [BBHH10, CW12, Coe15, CP13b, GAS13, Mou19, PP19, Tan13a, Wu11b, Xin12, XW17, ZZ15c]. surely [CLM12]. surface [Wan12a, WZ15a]. surrender [WWY12]. surrogate [BB15, Gho12]. survey [BLL14, BH11, KHHD16]. surveys [BLZ13, WR10]. Survival [Öz16, AAMB10, BMS16, CF18, CDL11, CWI10, Che14b, Che19b, CZCL19, Din10, GB19b, Han10, HD13, HP14, HH18, IK10b, Kay15, LW15a, MS10a, MM10, ND10, Oga18, SS14, sS10b, STW19, Sub12, Xue17, Xue19, ZZM15, ZZ15d]. Survivors [KMR13, LQZ13]. SVCHARME [RS14b]. swaps [HJS11]. Sweden [BH17b]. switch [LFM15]. switches [Li13d]. switching [AEFO18, BG15a, BG15b, BGO18, DYW14, FY16, GR17, Hie14, HH16, HMS14, KH18, LWD13, QLH^{+}16, XY13, ZYL14]. Sylvester [Hu15]. symbols [BS18a]. Symmetric [Dom13, BCGFR13, CP17b, DS18b, Duo15, EQ15a, EQ15b, FA12, FS14, GLM19, HR19b, HZR12, Iso15, JB13a, KL15, LT19b, MK10, Nag13, Nag19, Osȩ10c, RL14, Sas13b, TWM15, YYC17, ZZ15a]. symmetries [Jon12, Rau13]. symmetrization [LLR17]. symmetrized [BW15]. Symmetry [BS17a, BM13b, MO16, Ush11a]. synergistic [MP13]. synthetic [KS15c]. System [CM16, Mat14, AEFO18, AB12b, Ery10, Ery11a, FL10c, GLM19, HN15, HM13, INO10, KHN16, Li13a, MKJ15, MM13a, Osȩ14b, PSC12, PS11a, Raq10, VCM14, ZFZ13]. systems [Bre12a, Bre14b, BZV11, BS18c, CD10a, DZZ13, EK10, FZ13, FZ15, FZB16, Gra11, Hor16, JWW14, KKT14, Li11b, LFM15, LLK^{+}18, Liu14, Mar14a, MM11, MM12, MM13b, MF15, NK11, Nav18, OSW18, PB13, RIK16, ST19a, VAZB10, Vil11, VD18, WL15a, Wan15, WC17, YD12, YHMM15, ZY10a, Zha18, ZB11].

T. [LS12a]. table [JKP19, PKK14, Wan17b]. tables [BB12, GV16, KPK15, SB16, Zha19a]. taboo [Bul14]. Tail [AH13a, Dun16b, Dun16a, EM10, FS10, HK14, Jan18, JLR19, ÖÇ13, YLS12, ADM15, Aya13, BBdWG16, BMN16, BS15e, BF10, CKO19, CL13, CD12, GL13, GGO15, HPP19, Ili12, KL11, Kol17, KA10b, LW14a, LQ19, Lin19a, LMS17, LPH13, MJH18, SZ13a, ST13a, ST10, SJ14, TMS11, YHW11, YIS15, Yan15c, Zha14a, Zhe17, dWGM12]. tail-decoupling [CD12]. tailed [Arv17, BS12a, BP19c, BGHS11, BVV14, DS16, FMA16, GF11, GGV14, Hwa19, LQ19, LW12, Lu12, NP10, YW10, YLS14, YIS15]. tails [BS11, BV15b, Bra11b, CK16a, CD18b, DZF18, GM18b, JT11, KL15, NA13, PL15b, PY17, Pin15b, Roz12, Tsi12, YWC10, Zha14a]. Takács [BP19a]. taking [HY13, Osȩ14a, ST13b]. Talpaz [NW10a]. Tamed [NL19]. Tamer [WS15]. tandem [DMST13]. tapered [DPP16]. TARCH [EM10]. target [MBTC13]. targeted [WRvdL11]. TASEP [ZC19]. tau [DM17a, Lo17]. tax [WMH11, WWPY18]. taxation [MWH17]. taxed [WM18]. Taylor [BL12, JDK^{+}11]. team [Zaj14]. tears [BR17a]. technique [LSS13a, LLK^{+}18, LG15b]. techniques [SBN16]. Teh [MN16]. telegraph [DOT14, DM12a, LR12a, Mac16, Rat13]. telegraph-type [DOT14]. temperature [FFM18, Kar11, RR11]. temperature-dependent [RR11]. Tempered [CM11a, MS13, SS18, GKW19, KV15]. temporal [CG18, EFP19, IL17, ZZ12a]. tenable [DW18, KM12b, SM13]. tend [Red17]. tensor [Sak12, Sak15]. tent [BHJ10, BHJ12]. term [Han12a, LQZ13, PNT15, Spa11, ZX19]. Termination [Ref14]. terms [GMP19, PTW10, Poi19, YD12]. ternary [LMLW15]. Test [BD11, AR18, AB11, AM11, Arv17, BWW11, BCI14, BMD13, BMD15, BH17b, CQ18, Che14b, CHN14, CM17, Chi13, CCGPW17, DKT19, DD11, FP11a, FS15, Fin11, Fre16, GM17a, GB19a, GEV16, Han12b, Har16, HH13b, HS15, IKAH17, JdU16, JP11, Kak11a, Kak11b, KS15a, KKKP17, Li15c, LZ19, MTT11, MGA11, Mao14, ML17, NT14, NHBG19, PQW14, PS11b, RSdB15, RG18, SNS10, Sha12b, SB16, SD15, SY11, TSZ15, TYNZ15, VGE15b, WQD15, Xia17, YZ16, YK11, Zha19a, ZX18, ZP19]. test-based [RG18]. Testing [BDK18, BCMR13, CN16, Cha19b, CYW12, CFBD13, CC10b, GB16, HLM16, Hat12, HWM15, JN10, Ost17, Ost18, PSS13, Tah12, TZW13, WCM11, BKP15, BMP10a, CSS14, CM17, DD11, Döh14, GMA12, Gau14, GM17b, Gor14, Gus12b, JLJ16, Khm17, KZ14, LX10a, MR15a, Mal18, Mao15b, MGG11, MIL16, Rob15, Rob19, RW16b, SAM13, SK13, SKJ15, Sha14, SL12, Sma15, XD15, ZG10]. Tests [Car10, SWW11, SDJ15, VGE15a, BLWZ11, BDvdAW16, BKR18, BHJ18, Can16, Che13b, DEM14, DN16, GMA12, Gau14, Ger13, Ger18, GP17, Gou15, GW12c, Liu11, LH13, LDD15, LG16, MD13, Mao16, Mar16, MW12b, MO16, MMM13, MW14b, RM13, Sch16, Sha13a, sS10a, SS11a, SD10b, Sun17, TL15, Van16, VZ18, WHB10, Wan14a, WK11a, XFC17, YS12, You18, ZAV12, ZX12, ZDX14, dSGPM12]. TGARCH [WP14]. th [FC10, Jas19, PJ10, SY10, TE17]. th-order [SY10]. their [AM17, AA11, BVP10, CK16a, Dav12a, DFS16, GM16, JP12b, KKT14, Li13b, Liu11, Mas18, MO16, MA14b, MX18, PS19a, Poi19, RS12a, RS10b, Wan14b, WWD15]. them [Osu12]. theorem [AMR19, ABTV19, Arb19, AG11, AG18, BM13b, Bor16, BB17, CW15, Cha15, Che13a, CS14a, Chr12, DL10, De 11, DF14, DR02, DR15, DJ14, DYB11, Eks14, El 13, ET19, FS16, Fro15, Gao19, GAS13, GA12, GZ14, GS11, Had19, He12, HL13a, IP14a, Kam17, Kat19, Kir19, KL16, Kov10, LBM14, LZ16a, LMLW15, LZP17, LW15c, MK10, MNBO11, Mor18, Mül19, Osu12, PLH13, RS16, Ren13, Rok15, Röl18, RR16, SK13, Spa10b, SS10c, Sun13a, Sze10, Tan13a, Ton16, Ton17, THC17, Vig12, WWW10, WLLL17, WWS19, WC14, XF17, XH12, XW17, YR11, Yao12a, ZF13, ZL17, BHJ12, BBH14, Bob13, Che18d, Len11, MW11, MS11, NW10b, Sha13b, Ter15, Tyu12, VY12, Yan12]. theorems [Ane17, ANRW15, BGT15, BPW18, CCFL18, CK16b, LY17, Mas18, McK15, MWA16, MR15b, MC15, PS19a, PZ17, Pan17, Seo15, Seo17, Seo19, Ski10, Ste11, Tan13b, TK10, WMM15, Wu11b, WH11, WK11b, ZL15, EM16, VW14, WGL11]. theoretic [DdRS^{+}11, Kab13, KS12]. Theoretical [HP15, Yoo15]. Theory [MMM13, AH13a, Arv17, Arv19, BH17a, Che12a, DP10, DER15, EG11, GF11, Gut11, HA12, Hwa19, IMM16, Lin19a, Mac11, Mül19, OV15, Shi19c, Ski15, Tao19, Uno13, ZL14]. theta [EHP13]. thick [CD18b]. thin [GM18b]. thinning [DZ19a]. thinning-selfdecomposable [DZ19a]. third [BT14a, Gau19]. Three [CDS17, GM19, BS19a, EQ15a, MKJ15, MP15b, OS10, PPM18, Rau13, Sin19, WF15, YS12, ZWHQ15]. three-dimensional [Rau13]. Three-level [CDS17, EQ15a, Sin19]. Threshold [TSZ15, AT15a, BS12b, CPH12, El 13, GML10, GLML12, HBPC10, MST10, Nic17, SLZ13, YY11b]. threshold-GARCH [HBPC10]. thresholding [Jia13, LG13]. ties [GR10, PRS15]. Tight [GM14c]. Tightened [Sas13b]. tightness [Cha15, LV13]. tilt [IK10a]. Time [Bre12b, KNV11, Woo15, dFAM19, AS15, ACMM13, BML14, BS12a, BGT15, BCI14, Bap11, BG15a, BG15b, BMB16, Bra16, Bre14b, Bri15, CP11, CL19a, CF13, CTW10, CY11, CXZ15, CGLN17, CM17, Che14d, CPY15, CL16, Cho18, CC10b, CJY14, CCSC11, Cui14, CZ17b, DE13a, DK14, DL13b, DK11, DL17, Dun19, DD14, Efr14, EL18, Ery15, EGX16, EFP19, FJ10, Fis13, FL10c, For11, FKZ15, GL13, Gat13, GFH16, Goi12, GAB16, GSW11, GS12, Hat12, HS12b, HY13, HZ16b, Iso15, JS18, JLY14, Kar11, KV17a, KI15, KFHS11, KPO13, KK19a, Kob16, KR19, KKP15, Kre18, LE19, LML15, LL19a, LBM14, Lee12b, LP11, LW14a, Li17a, Li17c, LL17b, LLK^{+}18, LX12, LWL16, Liu13, LC14, Ma13, MZ17, MNO15, McC12, MNX13, MNOT12, MN14, MMC14, Nic17, OS15, PLL13, Pen14, PL17, PD12]. time [PPN10, PH13b, QLH^{+}16, Roh13, SBA12, ST19a, SM10, SA12b, Sas13b, SM16, SHK14, Sta16, Ste15, TMN16, TLZ12, Tri18, TN17, Vid19, Vri16, WY10, WJY11, WWW14, WZG14, WWX17, WWZ18, WM18, Wan19, Wei19, WS10b, Wu11a, Wu14, XKBG15, XWML19, Yam15, YS10, YYC17, YL19, Yan19a, Yao12a, YHMM15, ZK14, ZK15, ZKSK18, ZY10a, ZFZ13, ZF13, ZW15a, Zha11b, ZJZ16, ZR12, Zho13b, aZLW18]. time-average [ZJZ16]. Time-changed [KNV11, HS12b, Kob16, MN14]. time-dependent [BML14, McC12]. time-fractional [TN17]. time-homogeneous [Cui14]. time-inhomogeneous [Che14d, Sta16, Vri16]. time-periodic [DE13a]. Time-varying [dFAM19, Bra16, DD14, LLK^{+}18]. times [BDER14, BS15c, BL11, Bul14, Dal19, Din10, Fis13, GS17a, HL19, HS12b, Hie14, IMM16, JS15, KFHS11, KMN10, Lee12a, LZ14a, LN19, LT14, Muk17, PR12, Rao13, SY17, SWW15, VT10, WWPY18, YG13, Yan17a, ZH12a, ZW15b, ZWB17]. timing [LFM15]. Tobit [DWW17, SY11, ZP15]. Toeplitz [BGT15, BBGH12, GS13a, LLSW11]. Tolerance [Par17b, MM15]. tools [MNO15]. top [Sha17]. top-down [Sha17]. Topp [Mir14]. torus [PH13a]. tossing [Li13d]. total [ART14, Chi10, Dav17, Jov14, KV16, LL17a, MXZZ18, RIK16, Sas13a, VA16, YK11, Zin13]. totally [Lar15]. tournament [RV19b]. trace [GS13a, Sep18]. tracer [CK14, Cho19]. tractable [HJS11]. Tracy [DM17b, NW13]. trades [Fen14]. trait [SKK10, WRvdL11]. trajectories [CDG15]. Trajectory [Bre14b, RV19a]. transaction [LY18b, SS17]. transfer [AG18]. transfer-current [AG18]. transform [Abd11, KLY14, Mna11, ML13, Ose19, Roz10, YY11a, Zha17a]. transformation [Bro07, Bro11, CJY14, HL12, LP10a, LZ17a, Ma15, QW16, ZF16]. transformations [Rok15, Shu18a]. transforms [AL16, BW18, GS17a, HY18, HBH10, Osȩ10a, Osȩ12b, Osȩ14b, YYY17, Yu11, YY14]. Transient [DMST13, IM15, KM13b, DK11, JP11, LW11, Ter17, XX13]. transition [AdUÁMM11, BDER14, GSW11, Jar13, Nie16, XXY12]. transitive [BH19, LG16]. translated [GT16, dMO18]. Translation [SKB15, SKB14]. Transmuted [KP16]. Transport [De 16b, AJ14, Bob13]. Transportation [BBM19, BH18, DZ11, Din14, Din15b, DZF18]. trap [Öz16]. traps [ÖÇ13]. trawl [GLT18]. treatment [DHL14, Hat12, LTvdVR11, LZ14b, LD16b, NW10a, WHB10]. treatments [Yat15]. tree [AI17, Bap11, BD11, Che13a, DYB11, MV10, SY10, SHK14]. trees [AC11, ABD14, HBL11, HLB17, He16, HNGS15, MFL15, WZ13a]. trend [Gut11, KKP15, PS19b, ZYS19]. trends [WN13a]. tri [BM10b]. tri-variate [BM10b]. trial [Kab16]. trials [BB11c, Chi12a, Chi12b, EGX16, Ery18, KK19a, Rob19, SM10, Sch13, VT11, ZK10]. triangles [Bar12]. Triangular [Che18a, AIM19, HPW15, Hil14, KZ10, Nag13, QG13, WL17, ZCM15]. trigonometric [AH13c]. trimmed [BBH14, GML15, VS13, Wan18]. trinary [EGX16]. Triple [LQ18]. trivariate [CR13]. trivial [Ter17, Tsi17]. Trojan [JVVS10]. Trojan-type [JVVS10]. Trotter [Gov15]. Truncated [KW14a, Nol18, ADP12, BMN15, BMN16, Che19b, DZX18, GFA10, KRB13, LN15, LW15b, PCGV14, sS10a, sS15a, Shu18c, WS12, WMZW13, WMFW15, WW16, XSZ17, ZBZ11, ZJL13, Zhu13b, vWvZ17]. Truncation [YY11a, CSL16, QB14, sS10b]. Tsirel’son [HY13]. Tuned [Lin13]. twice [Bou15, KLM11, KM12a, KBM15, Mes10]. Two [GO12, Gri19, Jac10, KM18, LZZ15a, LY18b, Mel16, WM18, WD12, WT14, ZK15, ZP15, AJ14, AH13c, AV13a, ABGM16, BS11, BB16, BMP10a, BV10, BV11, BK18, BL11, Bra11a, BZV11, Chi10, CJN19, CC10b, DTV16, EQ15c, EPSU16, FS15, Fer11, Fra15, Fre16, FN17, FS10, HV16, JT11, JP12b, Kab16, KT19, KPK15, Leh15, LSS15, LZ14b, LZYW19, LP17, LZ16b, LG16, Ma14a, MGA11, MD16, ML16, Men12, Mes14, MM12, MM13b, MK18, NK11, NSH17, OH15, OQ10, PRS18, PJ13, RS10a, Ruk12, Ruk16, SM19, Sha12a, Sha14, SH11, SZ13a, SHK14, SM13, TL15, Uno13, VT11, Vid17, WLS11, Wan12c, WL15a, Wan15, WN10c, Xu14, Xue18, YZ14, YD12, YWZ14, Yor14, ZY10b, ZX12, ZWHQ15, Zha19b, ZZ11, ZZM15, ZGR12, ZY11, ZWB17, Zin13]. two-color [SM13]. two-component [BZV11, MM12]. two-dimensional [BV11, BL11, CJN19, FN17, LZ16b, SZ13a]. two-factor [Jac10]. two-filter [PJ13]. Two-level [Jac10, EQ15c, MD16, OQ10]. two-parallel-series [WL15a]. two-parameter [LSS15, Men12, OH15, Sha12a, Zha19b]. two-period [Kab16]. two-phase [LZYW19]. Two-sample [WD12, WT14, FS15, LG16, MGA11]. two-series-parallel [WL15a]. Two-side [WM18]. Two-sided [LZZ15a, Mel16, ZK15, LP17, Vid17, ZY10b, ZWB17]. two-stage [BB16, Kab16, Uno13, VT11, Xue18, ZGR12, ZY11]. Two-step [ZP15, JP12b, YWZ14]. two-type [KT19, Ma14a, Xu14]. two-way [KPK15, ZX12]. Tyler [FG12]. Type [PP15, Abd11, AFJ11, AB11, AM11, ABTV19, Bak12, BS19a, BK19a, BBdWG16, BMN16, BZ13, BRBB14, BRBB23, BGHS11, BS18b, BO19, hCyP15, CS19, Cui14, DOT14, DF14, DM13, DT10, EQ14, GM17a, GGO15, GA12, HV16, JKD15a, JVVS10, JYL12, JWW14, JP10, KT19, Kak12, Kao14, LVY15, LM17, LZP17, Ma14a, MJH18, MS12a, MC15, Now19, Osȩ10c, Osȩ12a, Osȩ14c, PS18a, Pin13, RSR16, RBY10, Ric10, SS11a, SV15, Sma15, ST13a, TLZ12, Tru10, TU17, Wan14c, Wes15, WH11, Xu14, Yan17b, Yao12a, Ye16, ZJ13, Zha12, dSF12, BD13a, BPS17, FXT12, HTA10, MB14, PN12, SK15, ZP15, BKL19]. Type-3 [ZP15]. Type-I [BD13a, BPS17]. Type-II [PP15, BK19a, FXT12, HTA10, SK15, BKL19]. types [HA15]. typical [Bel13].

UGROW [Pak13]. Uhlenbeck [Ma10a, Abd15, Ai16, ÁLBRM16, BSS13, BS15b, BPS14, BR17a, BY12, DY10, ES13, HN10, JM14, SWW15, SG15b, TLZ12, Vri16, Wan11a, Wan16, Xin12, XXY12, ZZ19b, ZWB17]. ultimate [CR16]. ultracontractive [CJ14]. ultrahigh [KS17a, LSCL17, WW14, ZPZ18]. ultrahigh-dimensional [ZPZ18]. umbrella [GMA12]. UMPU [BMD13, BMD15]. unbalanced [AWL17]. Unbiased [MD13, NSH17, FS14, GGV14, Han18, HCP19, KKKP17, LLCG12, Li13d, ML17, NS12, YS19, ZG17a, dWGM12]. Unbiasedness [HH19]. unbounded [He14, Pam18, PTV16, RS15b]. uncertain [Rok15, YHMM15]. Uncertainty [CvEZ10, Ost13, Pac18]. unconditional [IA10, Sha13a]. uncorrelated [For11, MR15a, TMN16]. under-valued [Sco18]. unequal [CRG17, DS15, PH13b, WCZ18, ZDX14]. unfair [AIP18]. Unified [MW17, LP10a]. Uniform [BDB19, BK10, BG11, BES18, BF17, Cal13, DP19b, FN17, GL13, KV13, Oua13, Pen11, YYS18, Zha14a, BHJ10, BHJ12, BM12, Cha15, CHR16, CQZ15, DG13, Fak10, For18, HR11, Lee12b, LV13, LLO14, LCLQ16, LZ13, Lu15, MC15, Mou19, Öz16, SR12a, Wei17, WN19b, YCL14, ZGG16, dMO18, vLT11]. Uniform-in-bandwidth [Oua13]. uniformity [IKAH17, Sch16, YSLL10, YZ18]. Uniformly [KKKP17, CL19b, EAB11, TJD10]. Unilateral [DYW14, CFBD13]. unimodality [AA11, AAC15, KKL18, ZP16]. union [Liu11, YAT16a]. unions [Fro12]. Unique [HS17, Gei17, Öns13]. Uniqueness [Rul14, WL13, GLM^{+}13, LYB13, SR12b, WZ19, XY17, ZX17b, ZT19]. Unit [AR18, Arv17, BG13, BDvdAW16, BVP10, CK13, CS16a, CG13, CY15, Han12b, JKP19, Mar11b, PRS18, VY12, Yab17, ZL14, ZP19]. Unitary [JW17, NW13]. univariate [BPXS19b, QL11, ZWLB17, ZHWH12]. Universal [Bro07, Bro11, KP17, CRC10, Deb12, HZR16, SL13a]. Universality [AS18]. Universally [Fil12]. unknown [Ben18, CS17, Mac11, WK11a, ZG16, ZF16]. unrelated [WLS11, ZX17a]. unstable [AT15a]. unstructured [Wes13]. upcrossings [OS15, XWT19]. updating [SLC16]. Upper [CKF17, GL13, GAB16, Gor12, Hil14, JMDW15, Lu11, McC13, QN15, Yan15c, YWC10, Zhe17]. urn [BT14b, Cal13, CK16b, CGLN17, FP12, GM18a]. urns [DW18, KM12b, ZCM15, ZM16a]. use [FP17, JKD15b, MM14a, MvdBV^{+}11, Ste14, WSU11]. used [LMM13]. Using [AT15b, LG13, Nos19, Sha12b, SS11a, VT10, WZ15a, ZZ13a, AWL17, AO12, ANRW15, BD11, BNM13b, BW19, BD13b, Bow17, BDB^{+}10, CFS12, CDL11, CSY12, CXW13, CD18b, DMC17, DP10, DYW14, DG13, ER17, Gho11b, HS11b, Han12a, HH13b, Ili12, JK13, JP12a, Kat19, LSS13a, Li15b, LZ19, LK13, Mao15a, ML13, MS18a, MM14b, NSL11, NMS15, PB15b, RM13, SLM19, SK11, ST14, Sma15, TL19, TZT16, TMS11, Tyu12, WYY14, WY16, WZ10, Wes13, YAT16a, ZV15, ZDX14, dFAM19, vDLMW19, vWvZ17]. usual [Kab11]. utility [Bos14, DFH18].

vaccine [Chi12b, VT11]. vague [BP19b]. Valhalla [Wit18]. valid [IL17]. validation [JT14, MvdBV^{+}11]. validity [CM10, WSU11]. Vallée [Cha15, HR11]. valuation [DYW14]. Value [KOR15b, BOK17, BMN15, CC16, Chu14, GGQ17, GHR16, GHR17, Gre12, HdCC17, JCM15, LP10b, LX10a, Llo12, Mül19, Nad13, Ohy13, SL12, VAB15, Wan17a, CS16a, HM15]. Value-at-Risk [HM15]. valued [Bon12, Bud14, CQT12, CH17, DFL12, DFL13, ET19, FS17a, IT14, Ilm13, JP10, KO17, Khm13, Kin10, KS13, KM19, LNI16, LXL19, Lop15, Mor18, NR12, Ose19, Rao10, Sco18, SV15, Sto12, Sun17, Vas18, WZ13b, ZZ12b, ZH12b]. values [ABBK15, BE17, BMP10b, Che13b, Fro17, HY13, Hwa13, JS19, Kum15, KN10, KP15c, Llo10, Osȩ14a, Pen11, Qiu17, RW16b, ST13b, SDNS16, TE17, WYC15, Yak15, YSLL10]. VaR [CS16a]. Varadhan [JMDW15]. VARFIMA [PR15]. Variable [AV14, HL13b, JKP19, JQZ12, RW17, Yat16b, BBM18, CD18a, Doe18, FA14, LML15, LSCL17, LR12b, LG13, LZY14, NS13a, NMS15, OPS16, PR16, Pin15a, SD11, TZW12, TXX16, WL14a, Wój13a, WS15, ZV15]. Variable-based [JKP19]. Variables [ASH18, Abd11, AB12a, AP18, ASK18, Ale18, Ana17, AKU17, ABBK15, ABB^{+}17, AKP18, AvHZ17, BS15c, BDS19, BM10c, BJD16, BK19b, BM18, CP13a, CO13, CQT12, CG12, CHR16, CH11a, CS14a, CS14b, CYH16, CS16c, CS16b, CS19, CS18, CD10b, DS16, DNLR17, DM17a, Dev12, DR02, DR15, Di 11, EAB10, EAB11, EP17, Fen19, For18, GBVS11, GD15, GD18, Gau19, GF11, GM18b, Goi12, GZ13, Had13, HS11a, HWB10, HCW13, HBH10, HM15, HR15, Isl16, Jan18, JT11, JGW14, JN10, Jou14, JP16, KHN12, KO17, KS13, KL15, KA10b, Kru17, KV13, LV11a, LSS13a, Leh15, LYW11, Li13b, LL15, LL19b, LCKK14, Liu12b, LMS17, Mas18, MMZ17, MW12a, MD15, MNBO11, Myr19, NKKY13, Nak16, OS10, PYK15, Pan17, PS18a, PDW10, PNBW15, Pin15b, Puc13]. variables [QC14, RS16, RS10b, Roz10, Roz12, Roz16, RR16, Ruz14, SWL10, SCS14, SWZ19, Shm13, Shu19b, SY11, SL15, Sre10, SD14, SL10, SZ13b, Sun13b, Sze15, TZW13, Tau15, TY15, Ton16, Ton17, Vas18, WW11, Wei17, WWD15, WQY12, WWH17, Xia19a, XY12, XFC17, Yak15, Yan14, YLS10, YHW11, YLS12, YS13, ZZ15b, ZZ15c, ZB10, ZL12, Zha12, ZX17a, Zhe17, Zhu14, dS18a, dMO18, dC16]. Variance [AH12a, PP11, ZV15, AL16, BWW11, BK17, Buo18, CCS16, CDG15, CS17, CvEZ10, Che16, CM17, Che18c, CK14, DFS^{+}13, DMS19, EF16, FMPV12, GF11, GM10a, Gir16, GAB16, HLM16, HS15, JB19, KST18, KP14, LP11, LLHW10, LLH11, Li15c, LMR12, MD13, NS12, PB15b, PS16c, Pos10, QO16, RP19, RS15b, Ruk12, SA12a, SW11, SGG10, SLC16, TV17b, VGE15b, WC15, WC18, XZ13, Yu17, ZG16, ZJZ16, ZYG17]. variance-covariance [MD13]. variance-mean [Yu17]. variances [Ald13, FP13, LX10a, SGG10]. variate [BM10b, BMS16, GM19, ZRN17]. variation [ART14, ADH19, ATV10, BWW11, BP19c, CSS14, Dav17, FL10b, GS12, Jov14, KV16, LZ11, MXZZ18, Sas13a, VSP13, VA16, Zin13]. Variational [QX13, AD11, EK10, Hui10, NLF11, XZ16]. Various [Pru19, Rao13]. VARMA [Giu15]. Varying [LDLC18, MS12b, Alb18, BBP18, Bra16, DD14, FL10c, FS11, HZJ^{+}10, IKM16, JT11, LML15, LZLW16, LLK^{+}18, LHT15, MH12, OV15, Roh13, SZW12, SS15c, TXX16, WS13, YW10, YL10, YLS10, YGL14, cZzXgL17, dFAM19]. varying-coefficient [LHT15, SZW12, SS15c, YGL14]. Vector [CGH11, Lop15, LDLC18, CMS19, Duc04, Duc10, Gna12, JP10, KK11, Ma13, Mar14b, Mat17, McE17, Mes14, NM18, NHBG19, Par14, RZ10, Shu19a, SS10c, Xu12b, ZZLL17, ZRAN13, ZRN14]. Vector-valued [Lop15, JP10]. vectors [AA13a, AN15, BCGFR13, BBHV13, BV10, CQ18, CL13, Ejs16, GMP19, GBGPR19, HT15, JB13a, Ko18, Mes14, Mod11, Mod14, Shu18c, STW19, TKO12, Tri18, Vig12, YKY17, Zaj19]. Velocity [Cho18, Cho19]. Verblunsky [BIK12]. verification [Abd15]. version [BM17, BB17, DP16, LBH11, PLH13, PP19, PCGV14, Qiu14, YH15]. versions [AK10]. versus [CFBD13, DP19b, JT14, Pin17]. vertex [HBF14, LV11c]. vertices [BS12b, GGR11, HLR15]. very [NHN^{+}11]. via [ACD12, ACW10, AM13a, AG11, BS17a, CMS19, CGLN17, CD12, DS12, Els16, Ery12, GM18a, GL19, GI14, GLM^{+}13, JQZ12, JL16, Kao14, KV17a, LS16b, LMH14, MWGT17, MM13c, MM14c, Mül19, MMF17, PJ13, RZ10, SD11, SCS14, Shi19c, Son10a, STW19, Sre12, Vii16, YD12, YDG13]. vicinity [GS17b, Mar11b]. view [LN15, MMF17]. viewpoint [Bou17]. Viot [dSF12]. visibility [Tyk11]. Vlasov [WX19b]. volatilities [HBPC10, LLW14]. volatility [BGO18, For11, HH16, HS13b, KLNB14, LTR13, MOD18, SWH12, TLF12, Tsi17]. Volterra [AE19, LWYZ17, LZ17b, Men12]. volume [MS11, Wan12a]. votes [Ste15].

W. [Bor14]. w.r.t [YR11]. wagers [ED17]. Wagner [LS15]. waiting [EGX16, KPO13, KK19a, SM10]. Wald [Sma15]. Wald-type [Sma15]. Walk [Sab18, Bap11, CM11a, CLEA18, CLEAMS16, DK11, Gao19, Hil14, IP14b, Joh14, LT14, RV19a, Sab17, Ski10, WYY13, XX13, Yan19a]. walks [ABD14, BTT11, BS19b, CXZ15, HL19, KZZ13, MLF13, McC13, MNX13, Neu17, Shi19b, SZ18, Zhu15]. Waller [MPA12]. Walsh [HT14, SZW12]. Walsh-average [SZW12]. warm [Ery11a, HN15]. Wasserstein [Din14, WWW10, Xia17]. Wasserstein-Divergence [Din14]. Watson [BBP18, Che13a, He16]. wave [BN16, BM10a]. Wavelet [ZW18a, cZzXgL17, ACD12, AH12b, AEK19, Ben18, GM14b, LWX11, Li15b, Son10a, TXWW18, TA14, WW10, cZgL12]. wavelets [AM13b, dFAM19]. way [CCS16, KWL15, KPK15, LS10, ZX12]. Weak [AP18, BG13, BS17b, Buc15, CGH11, Gov14, HMS14, IKM16, JP10, Nak16, Osȩ10c, Osȩ12c, QY15, SHL15, Tab18, WYY14, YYY17, BOR17, BO19, CFS12, DL10, DFT12, DFL12, DFL13, DFT13, GF13, HT15, HS12c, Jir13, Kru17, Li13c, LL17b, LZP17, OK15, Osȩ12a, Osȩ14c, THC17, WHB10, Wan18, YY14, YH15, ZT19, aZLW18, QHS13, Ter15]. weak-norm [BOR17]. weak-type [Osȩ14c]. weakened [Maa19]. weakly [BFS16, CQT12, KO11, Tru10, WWH17]. Weber [Den15]. wedge [LCF15, LTP18]. Weibull [BNB16, Bog15, FZ13, FZ15, GGS12, GB19b, GGO15, KC16, LL15, MS12a, PB16, SI16, Zha11a]. Weibull-type [GGO15]. weighing [Sma14]. weight [MDR12]. Weighted [AT15a, DM12c, EBG16, GSK12, Osȩ17b, TZW12, TY15, AAEH11, BOR17, BO19, Buc15, hCyP15, CD18a, CC10a, CH11a, CS14b, CS16b, DRZ19, DHL14, DZX18, Fen19, GM17b, GMP19, GZ13, Han12a, Hom12, Ko13, LW16, Lin19a, LM15a, MU16, MY11, Nak16, Nas12, PYK15, PRS15, RIK16, RBY10, RS15a, RA12, Roz10, Roz14, Roz16, Sha12a, sS15b, SWZ19, SR12a, SLC16, TE14, WN10b, WWH17, Xia17, YLS12, YLS14, YIS15, YAT16a, YCQY16, Zha14a, Zha18, ZW11, dML12, vLT11]. weighted- [RIK16]. Weighted-mean [DM12c]. weighting [CD18a, jLL19, ZW18c]. weights [BL10a, CDL11, HBF14, KHHD16, ST18]. Weiss [Mar14b]. Weissman [For18]. Well [LHW19, Tyk11]. well-behaved [Tyk11]. Well-posedness [LHW19]. Wentzell [BC11]. whether [Fre16, MIL16]. which [BK18]. white [BN16, LZ19, Son12a, Wu11a]. whole [TZ16]. whose [Osȩ13a]. wide [Che13b]. widely [CS19]. Widom [DM17b, NW13]. width [BB16]. Wiener [Abu13, FCU11, Gau19, JK13, JV10, PLY18, Zin13]. Wilcoxon [DD11, ZW11]. wild [KKP15]. Wilks [LBH11, PTT19]. window [ASN16, WZG14]. windows [He14]. winners [Ste15]. Wintner [LZ16a]. wise [CMS19, Kru17, PS18a, SDNS16, Ton16, Ton17, WX12, YW17]. Wishart [Kou12a, LW17a, SKRT16, XKBG15]. Wishing [Van18]. within [Jam18, Pin17]. without [Arb19, BR17a, HRB19, jLL19, WC18]. WLLN [AKU17]. Wolfowitz [FP11b, WD12]. wood [BK18]. words [Aza13]. work [BH14]. workload [Bri15, Kin12]. world [Rei18]. Worst [BCM14]. Worst-case [BCM14]. wrap [CLQ12, ZWHQ15]. wrap-around [ZWHQ15]. wrap-round [CLQ12]. Wrapped [RA12, GZSZ19].

Xia [Lê16].

Yard [Cho16a]. Yard-Sale [Cho16a]. Yor [Kou12a, Qiu14, Wes15]. Yosida [Gov14]. Young [Yam15].

Zero [TPNSN19, BJD16, Cui14, DH19, DM17a, DL17, FM13, GI14, KM12b, Mar14b, PNBW15, Roz16, SM13, TH16, Tyu12]. Zero-adjusted [TPNSN19]. zero-average [DH19]. zero-balanced [KM12b, SM13]. zero-inflated [DM17a, PNBW15]. zero-modified [TH16]. zero-one [Cui14]. zeros [KPO13, LZT13]. zeta [Urb12]. Zipf [DP19b]. zonotopes [DP19a]. Zygmund [Fer14, Had11, Tru10].

References

Alimohammadi:2011:SNR

[AA11] Mahdi Alimohammadi and Mohammad Hossein Alamatsaz. Some new results on unimodality of generalized order statistics and their spacings. Statistics & Probability Letters, 81 (11):1677–1682, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002306.

Ahmad:2013:PIB

[AA13a] I. A. Ahmad and M. Amezziane. Probability inequalities for bounded random vectors. Statistics & Probability Letters, 83(4):1136–1142, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004397.

Arnold:2013:PSD

[AA13b] Barry C. Arnold and Krishna B. Athreya. Power series with i.i.d. coefficients. Statistics & Probability Letters, 83(3):923– 929, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004476.

Athreya:2017:PSP

[AA17] Jayadev S. Athreya and Krishna B. Athreya. Partial sum processes and continued fractions. Statistics & Probability Letters, 130(??):57–62, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302432.

Alili:2019:SGS

[AA19] Larbi Alili and Andrew Aylwin. On the semi-group of a scaled skew Bessel process. Statistics & Probability Letters, 145(??):96– 102, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521830289X.

Alimohammadi:2015:DSU

[AAC15] Mahdi Alimohammadi, Mohammad Hossein Alamatsaz, and Erhard Cramer. Discrete strong unimodality of order statistics. Statistics & Probability Letters, 103(??):176–185, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001303.

Aknouche:2011:OOW

[AAEH11] Abdelhakim Aknouche, Eid M. Al-Eid, and Aboubakry M. Hmeid. Offline and online weighted least squares estimation of nonstationary power ARCH processes. Statistics & Probability Letters, 81(10):1535–1540, October 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001714.

Abi-ayad:2018:PEN

[AaM18] Ilham Abi-ayad and Tahar Mourid. Parametric estimation for non recurrent diffusion processes. Statistics & Probability Letters, 141(??):96–102, October 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302116.

Abbasnejad:2010:DSE

[AAMB10] M. Abbasnejad, N. R. Arghami, S. Morgenthaler, and G. R. Mohtashami Borzadaran. On the dynamic survival entropy. Statistics & Probability Letters, 80(23–24):1962–1971, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521000249X.

Ahmadi:2010:PCC

[AB10] Jafar Ahmadi and N. Balakrishnan. Pitman closeness of current records for location-scale families. Statistics & Probability Letters, 80(21–22):1577–1583, November 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001719.

Anis:2011:END

[AB11] M. Z. Anis and Kinjal Basu. The exact null distribution of the generalized Hollander–Proschan type test for NBUE alternatives. Statistics & Probability Letters, 81(11):1733–1737, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002069.

Abraham:2012:PAM

[AB12a] B. Abraham and N. Balakrishna. Product autoregressive models for non-negative variables. Statistics & Probability Letters, 82(8):1530–1537, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001654.

Asadi:2012:NFC

[AB12b] Majid Asadi and Alexandre Berred. On the number of failed components in a coherent operating system. Statistics & Probability Letters, 82(12):2156–2163, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002829.

Aliyev:2017:ABC

[AB17] Rovshan Aliyev and Veli Bayramov. On the asymptotic behaviour of the covariance function of the rewards of a multivariate renewal-reward process. Statistics & Probability Letters, 127(??):138–149, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301505.

Archibald:2017:GRV

[ABB^{+}17] Margaret Archibald, Aubrey Blecher, Charlotte Brennan, Arnold Knopfmacher, and Helmut Prodinger. Geometric random variables: Descents following maxima. Statistics & Probability Letters, 124(??):140–147, May 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300341.

Archibald:2015:DFM

[ABBK15] Margaret Archibald, Aubrey Blecher, Charlotte Brennan, and Arnold Knopfmacher. Descents following maximal values in samples of geometric random variables. Statistics & Probability Letters, 97(??):229–240, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400399X.

Alessi:2010:IPD

[ABC10] Lucia Alessi, Matteo Barigozzi, and Marco Capasso. Improved penalization for determining the number of factors in approximate factor models. Statistics & Probability Letters, 80(23–24):1806–1813, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002282.

Abdelkader:2011:LTM

[Abd11] Yousry H. Abdelkader. A Laplace transform method for order statistics from nonidentical random variables and its application in phase-type distribution. Statistics & Probability Letters, 81(8):1143–1149, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000927.

Athreya:2014:RWD

[ABD14] Siva Athreya, Antar Bandyopadhyay, and Amites Dasgupta. Random walks in I.I.D. random environment on Cayley trees. Statistics & Probability Letters, 92(??):39–44, September 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001667.

Abdelrazeq:2015:MVL

[Abd15] Ibrahim Abdelrazeq. Model verification for Lévy-driven Ornstein–Uhlenbeck processes with estimated parameters. Statistics & Probability Letters, 104(??):26–35, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001248.

Aknouche:2018:PEG

[ABD18] Abdelhakim Aknouche, Wissam Bentarzi, and Nacer Demouche. On periodic ergodicity of a general periodic mixed Poisson autoregression. Statistics & Probability Letters, 134(??):15–21, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303280.

Azzalini:2016:NRM

[ABGM16] Adelchi Azzalini, Ryan P. Browne, Marc G. Genton, and Paul D. McNicholas. On nomenclature for, and the relative merits of, two formulations of skew distributions. Statistics & Probability Letters, 110(??):201–206, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215303813. See comment [ML16].

Alevizos:2019:EED

[ABI19] Filippos Alevizos, Dimitrios Bagkavos, and Dimitrios Ioannides. Efficient estimation of a distribution function based on censored data. Statistics & Probability Letters, 145(??):359–364, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302967.

Araya:2019:DTT

[ABTV19] Héctor Araya, Natalia Bahamonde, Soledad Torres, and Frederi Viens. Donsker type theorem for fractional Poisson process. Statistics & Probability Letters, 150(??):1–8, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300422.

Abundo:2012:EIF

[Abu12a] Mario Abundo. Erratum to: “An inverse first-passage problem for one-dimensional diffusions with random starting point” [Statist. Probab. Lett. 82 (2012) 7–14]. Statistics & Probability Letters, 82(3):705, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003749. See [Abu12b].

Abundo:2012:IFP

[Abu12b] Mario Abundo. An inverse first-passage problem for onedimensional diffusions with random starting point. Statistics & Probability Letters, 82(1):7–14, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002926. See erratum [Abu12a].

Abundo:2013:DBI

[Abu13] Mario Abundo. The double-barrier inverse first-passage problem for Wiener process with random starting point. Statistics & Probability Letters, 83(1):168–176, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003434.

Arias-Castro:2011:FSP

[AC11] Ery Arias-Castro. Finite size percolation in regular trees. Statistics & Probability Letters, 81(2):302–309, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003032.

Aronow:2015:NIR

[AC15] Peter M. Aronow and Forrest W. Crawford. Nonparametric identification for respondent-driven sampling. Statistics & Probability Letters, 106(??):100–102, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002357.

Abbaszadeh:2012:NED

[ACD12] Mohammad Abbaszadeh, Christophe Chesneau, and Hassan Doosti. Nonparametric estimation of density under bias and multiplicative censoring via wavelet methods. Statistics & Probability Letters, 82(5):932–941, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000260.

Avrachenkov:2013:MEM

[ACMM13] Konstantin Avrachenkov, Laura Cottatellucci, Lorenzo Maggi, and Yong-Hua Mao. Maximum entropy mixing time of circulant Markov processes. Statistics & Probability Letters, 83(3):768– 773, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004385.

Arias-Castro:2019:CMR

[ACP19] Ery Arias-Castro and Xiao Pu. Concentration of measure for radial distributions and consequences for statistical modeling. Statistics & Probability Letters, 145(??):216–223, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303158.

AlKadiri:2010:MLS

[ACW10] M. Al Kadiri, R. J. Carroll, and M. P. Wand. Marginal longitudinal semiparametric regression via penalized splines. Statistics & Probability Letters, 80(15–16):1242–1252, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001094.

Armagan:2011:SVA

[AD11] Artin Armagan and David Dunson. Sparse variational analysis of linear mixed models for large data sets. Statistics & Probability Letters, 81(8):1056–1062, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000757.

Arendarczyk:2012:EAS

[AD12] Marek Arendarczyk and Krzysztof Dȩbicki. Exact asymptotics of supremum of a stationary Gaussian process over a random interval. Statistics & Probability Letters, 82(3):645–652, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003683.

Aknouche:2019:ECD

[AD19] Abdelhakim Aknouche and Nacer Demmouche. Ergodicity conditions for a double mixed Poisson autoregression. Statistics & Probability Letters, 147(??):6–11, April 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303857.

Anh:2019:VCF

[ADH19] P. T. Anh, T. S. Doan, and P. T. Huong. A variation of constant formula for Caputo fractional stochastic differential equations. Statistics & Probability Letters, 145(??):351–358, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303262.

Asghari:2015:ETA

[ADM15] N. M. Asghari, K. Dȩbicki, and M. Mandjes. Exact tail asymptotics of the supremum attained by a Lévy process. Statistics & Probability Letters, 96(??):180–184, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003162.

Ahmadi:2012:ELT

[ADP12] Jafar Ahmadi, Mahdi Doostparast, and Ahmad Parsian. Estimation with left-truncated and right censored data: a comparison study. Statistics & Probability Letters, 82(7):1391–1400, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001101.

Amorim:2011:PTP

[AdUÁMM11] Ana Paula Amorim, Jacobo de Uña-Álvarez, and Luís Meira-Machado. Presmoothing the transition probabilities in the illness-death model. Statistics & Probability Letters, 81(7):797– 806, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000630.

AbiJaber:2019:MSV

[AE19] Eduardo Abi Jaber and Omar El Euch. Markovian structure of the Volterra Heston model. Statistics & Probability Letters, 149(??):63–72, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300306.

Abo-Eleneen:2018:CRE

[AEAN18] Z. A. Abo-Eleneen, B. Almohaimeed, and H. K. T. Ng. On cumulative residual entropy of progressively censored order statistics. Statistics & Probability Letters, 139(??):47–52, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301226.

Aazizi:2018:OSP

[AEFO18] Soufiane Aazizi, Tarik El Mellali, Imade Fakhouri, and Youssef Ouknine. Optimal switching problem and related system of BSDEs with left-Lipschitz coefficients and mixed reflections. Statistics & Probability Letters, 137(??):70–78, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300087.

Ayache:2019:DPB

[AEK19] Antoine Ayache, Céline Esser, and Thomas Kleyntssens. Different possible behaviors of wavelet leaders of the Brownian motion. Statistics & Probability Letters, 150(??):54–60, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300550.

Ahmed:2012:SES

[AF12] S. Ejaz Ahmed and Saber Fallahpour. Shrinkage estimation strategy in quasi-likelihood models. Statistics & Probability Letters, 82(12):2170–2179, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002969.

Ajami:2011:BRK

[AFJ11] M. Ajami, V. Fakoor, and S. Jomhoori. The Bahadur representation for kernel-type estimator of the quantile function under strong mixing and censored data. Statistics & Probability Letters, 81(8):1306–1310, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001246.

Asmussen:2011:NPC

[AG11] Søren Asmussen and Peter W. Glynn. A new proof of convergence of MCMC via the ergodic theorem. Statistics & Probability Letters, 81(10):1482–1485, October 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001751.

Aronow:2013:SBC

[AG13] Peter M. Aronow and Donald P. Green. Sharp bounds for complier average potential outcomes in experiments with noncompliance and incomplete reporting. Statistics & Probability Letters, 83(3):677–679, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004282.

Avena:2018:PTC

[AG18] L. Avena and A. Gaudillière. A proof of the transfercurrent theorem in absence of reversibility. Statistics & Probability Letters, 142(??):17–22, November 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521830230X.

Agahi:2015:EPC

[Aga15] Hamzeh Agahi. An elementary proof of the covariance inequality for Choquet integral. Statistics & Probability Letters, 106(??):173–175, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002448.

Alquier:2011:GCH

[AH11] Pierre Alquier and Mohamed Hebiri. Generalization of ℓ_{1}- constraints for high dimensional regression problems. Statistics & Probability Letters, 81(12):1760–1765, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002422.

Andel:2012:VGD

[AH12a] Jirí Andel and Sárka Hudecová. Variance of the game duration in the gambler’s ruin problem. Statistics & Probability Letters, 82(9):1750–1754, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001988.

Ayache:2012:LFS

[AH12b] Antoine Ayache and Julien Hamonier. Linear fractional stable motion: a wavelet estimator of the α parameter. Statistics & Probability Letters, 82(8):1569–1575, August 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001484.

Abbasi:2013:TCE

[AH13a] Babak Abbasi and S. Zahra Hosseinifard. Tail conditional expectation for multivariate distributions: a game theory approach. Statistics & Probability Letters, 83(10):2228–2235, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002174.

Al-Hussein:2013:SCO

[AH13b] AbdulRahman Al-Hussein. Sufficient conditions for optimality for stochastic evolution equations. Statistics & Probability Letters, 83(9):2103–2107, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001946.

Alqallaf:2013:MDD

[AH13c] Fatemah Alqallaf and S. Huda. Minimax designs for the difference between two estimated responses in a trigonometric regression model. Statistics & Probability Letters, 83(3):909– 915, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004518.

Alqallaf:2019:CIS

[AHM19] Fatemah A. Alqallaf, S. Huda, and Rahul Mukerjee. Causal inference from strip-plot designs in a potential outcomes framework. Statistics & Probability Letters, 149(??):55–62, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300331.

Ahmed:2010:RIS

[AHN10] S. Ejaz Ahmed, Abdulkadir Hussein, and Sévérien Nkurunziza. Robust inference strategy in the presence of measurement error. Statistics & Probability Letters, 80(7–8):726–732, April 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000076.

Ai:2016:NKL

[Ai16] Xiaohui Ai. A note on Karhunen–Loève expansions for the demeaned stationary Ornstein–Uhlenbeck process. Statistics & Probability Letters, 117(??):113–117, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300419.

Altok:2017:LRS

[AI17] S. Altok and Ü. Islak. On leaf related statistics in recursive tree models. Statistics & Probability Letters, 121(??):61–69, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215304119.

Akahori:2019:CMT

[AIM19] Jirô Akahori, Yuuki Ida, and Greg Markowsky. p-conformal maps on the triangular lattice. Statistics & Probability Letters, 151(??):42–48, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300860.

Arslan:2018:UP

[AIP18] Ilker Arslan, Ümit Islak, and Cihan Pehlivan. On unfair permutations. Statistics & Probability Letters, 141(??):31–40, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301925.

Alfonsi:2014:ROT

[AJ14] A. Alfonsi and B. Jourdain. A remark on the optimal transport between two probability measures sharing the same copula. Statistics & Probability Letters, 84(??):131–134, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003337.

Athreya:2016:APB

[AJ16] K. B. Athreya and R. Janicki. Asymptotics of powers of binomial and multinomial probabilities. Statistics & Probability Letters, 112(??):58–62, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215302601.

Antonov:2010:SVK

[AK10] Sergei N. Antonov and Victor M. Kruglov. Sharpened versions of a Kolmogorov’s inequality. Statistics & Probability Letters, 80(3–4):155–160, February 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771520900371X.

Abid:2019:GDM

[AKM19] Rahma Abid, Célestin C. Kokonendji, and Afif Masmoudi. Geometric dispersion models with real quadratic v-functions. Statistics & Probability Letters, 145(??):197–204, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303092.

Arendarczyk:2018:JDS

[AKP18] Marek Arendarczyk, Tomasz. J. Kozubowski, and Anna K. Panorska. The joint distribution of the sum and the maximum of heterogeneous exponential random variables. Statistics & Probability Letters, 139(??):10–19, August 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301330.

Akritas:2016:PPM

[Akr16] Michael G. Akritas. Projection pursuit multi-index (PPMI) models. Statistics & Probability Letters, 114(??):99–103, July 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303023.

Ankirchner:2017:WAN

[AKU17] Stefan Ankirchner, Thomas Kruse, and Mikhail Urusov. WLLN for arrays of nonnegative random variables. Statistics & Probability Letters, 122(??):73–78, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302371.

Ambrus:2012:DSP

[AKV12] Gergely Ambrus, Péter Kevei, and Viktor Vígh. The diminishing segment process. Statistics & Probability Letters, 82(1):191– 195, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003105.

Arvanitis:2016:CMT

[AL16] Stelios Arvanitis and Alexandros Louka. A CLT for martingale transforms with infinite variance. Statistics & Probability Letters, 119(??):116–123, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215303990.

Albin:2018:CFS

[Alb18] J. M. P. Albin. On covariance functions with slowly or regularly varying modulo of continuity. Statistics & Probability Letters, 138(??):177–182, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301196.

Alvarez-Liebana:2016:CPF

[ÁLBRM16] Javier Álvarez-Liébana, Denis Bosq, and María D. Ruiz-Medina. Consistency of the plug-in functional predictor of the Ornstein–Uhlenbeck process in Hilbert and Banach spaces. Statistics & Probability Letters, 117(??):12–22, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630044X.

Aldaz:2013:MPV

[Ald13] J. M. Aldaz. A monotonicity property of variances. Statistics & Probability Letters, 83(5):1416–1419, May 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000436.

Aletti:2018:GDR

[Ale18] Giacomo Aletti. Generation of discrete random variables in scalable frameworks. Statistics & Probability Letters, 132(??):99– 106, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217302900.

Alfonsi:2013:SOO

[Alf13] Aurélien Alfonsi. Strong order one convergence of a drift implicit Euler scheme: Application to the CIR process. Statistics & Probability Letters, 83(2):602–607, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004063.

Alfonsi:2015:SPC

[Alf15] Aurélien Alfonsi. A simple proof for the convexity of the Choquet integral. Statistics & Probability Letters, 104(??):22–25, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001340.

Ai:2012:KLE

[ALL12] Xiaohui Ai, Wenbo V. Li, and Guoqing Liu. Karhunen– Loeve expansions for the detrended Brownian motion. Statistics & Probability Letters, 82(7):1235–1241, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200079X.

Attaoui:2011:NCD

[ALS11] Said Attaoui, Ali Laksaci, and Elias Ould Said. A note on the conditional density estimate in the single functional index model. Statistics & Probability Letters, 81(1):45–53, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002671.

Anis:2011:GHP

[AM11] M. Z. Anis and M. Mitra. A generalized Hollander– Proschan type test for NBUE alternatives. Statistics & Probability Letters, 81(1):126–132, January 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002786.

Aman:2013:OPS

[AM13a] Auguste Aman and Naoul Mrhardy. Obstacle problem for SPDE with nonlinear Neumann boundary condition via reflected generalized backward doubly SDEs. Statistics & Probability Letters, 83(3):863–874, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004117.

Angers:2013:HFE

[AM13b]  Jean-François Angers and Brenda MacGibbon.  Hazard

function estimation with nonnegative “wavelets”. Statistics & Probability Letters, 83(4):969–978, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004865.

Allam:2014:COE

[AM14] Abdelaziz Allam and Tahar Mourid. Covariance operator estimation of a functional autoregressive process with random coefficients. Statistics & Probability Letters, 84(??):1–8, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003167.

Abdelghani:2017:LSE

[AM17] Mohamed N. Abdelghani and Alexander V. Melnikov. On linear stochastic equations of optional semimartingales and their applications. Statistics & Probability Letters, 125(??):207–214, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215301486.

Alhakim:2019:DFP

[AM19a] Abbas Alhakim and S. Molchanov. The density flatness phenomenon. Statistics & Probability Letters, 152(??):156–161, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521930135X.

Aouicha:2019:KEC

[AM19b] Lamia Aouicha and Fatiha Messaci. Kernel estimation of the conditional density under a censorship model. Statistics & Probability Letters, 145(??):173–180, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303080.

Ahmadi:2010:NPI

[AMB10] Jafar Ahmadi, S. M. T. K. MirMostafaee, and N. Balakrishnan. Nonparametric prediction intervals for future record intervals based on order statistics. Statistics & Probability Letters, 80(21–22):1663–1672, November 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002038.

Arafat:2017:FMP

[AMG17] Ahmed Arafat, Jorge Mateu, and Pablo Gregori. A family of Markov processes in maximal compact subgroups of a semisimple Lie groups. Statistics & Probability Letters, 126(??):132– 138, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215303758.

Agahi:2013:SLL

[AMMO13] Hamzeh Agahi, Adel Mohammadpour, Radko Mesiar, and Yao Ouyang. On a Strong Law of Large Numbers for monotone measures. Statistics & Probability Letters, 83(4):1213–1218, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000229.

Aoudia:2016:CBS

[AMP16a] Djilali Ait Aoudia, Éric Marchand, and François Perron. Counts of Bernoulli success strings in a multivariate framework. Statistics & Probability Letters, 119(??):1–10, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303096.

Arapis:2016:LPM

[AMP16b] Anastasios N. Arapis, Frosso S. Makri, and Zaharias M. Psillakis. On the length and the position of the minimum sequence containing all runs of ones in a Markovian binary sequence. Statistics & Probability Letters, 116(??):45–54, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302285.

Angel:2019:LTR

[AMR19] Omer Angel, Anastasios Matzavinos, and Alexander Roitershtein. Limit theorem for the Robin Hood game. Statistics & Probability Letters, 149(??):9–15, June 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300070.

Akhundov:2010:SCS

[AN10] I. Akhundov and V. B. Nevzorov. A simple characterization of Student’s t_{3} distribution. Statistics & Probability Letters, 80(5–6):293–295, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004209.

Arevalillo:2015:NDM

[AN15] Jorge M. Arevalillo and Hilario Navarro. A note on the direction maximizing skewness in multivariate skew-t vectors. Statistics & Probability Letters, 96(??):328–332, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003630.

Anastasiou:2017:BNA

[Ana17] Andreas Anastasiou. Bounds for the normal approximation of the maximum likelihood estimator from m-dependent random variables. Statistics & Probability Letters, 129(??):171– 181, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521730161X.

Anevski:2017:FCL

[Ane17] Dragi Anevski. Functional central limit theorems for the Nelson– Aalen and Kaplan–Meier estimators for dependent stationary data. Statistics & Probability Letters, 124(??):83–91, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300123.

Anonymous:2010:EBa

[Ano10a] Anonymous. Editorial Board. Statistics & Probability Letters, 80 (1):ii, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209004131.

Anonymous:2010:EBb

[Ano10b] Anonymous. Editorial Board. Statistics & Probability Letters, 80(1):ifc, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209004118.

Anonymous:2010:EBc

[Ano10c] Anonymous. Editorial Board. Statistics & Probability Letters, 80(2):ifc, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771520900426X.

Anonymous:2010:EBd

[Ano10d] Anonymous. Editorial Board. Statistics & Probability Letters, 80(3–4):ifc, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209004623.

Anonymous:2010:EBe

[Ano10e] Anonymous. Editorial Board. Statistics & Probability Letters, 80(5–6):ifc, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000209.

Anonymous:2010:EBf

[Ano10f] Anonymous. Editorial Board. Statistics & Probability Letters, 80(7–8):ifc, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000544.

Anonymous:2010:EBg

[Ano10g] Anonymous. Editorial Board. Statistics & Probability Letters, 80(9–10):ifc, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000817.

Anonymous:2010:EBh

[Ano10h] Anonymous. Editorial Board. Statistics & Probability Letters, 80(11–12):ifc, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001021.

Anonymous:2010:EBi

[Ano10i] Anonymous. Editorial Board. Statistics & Probability Letters, 80(13–14):ifc, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001306.

Anonymous:2010:EBj

[Ano10j] Anonymous. Editorial Board. Statistics & Probability Letters, 80(15–16):ifc, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001501.

Anonymous:2010:EBk

[Ano10k] Anonymous. Editorial Board. Statistics & Probability Letters, 80(17–18):ifc, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001793.

Anonymous:2010:EBl

[Ano10l] Anonymous. Editorial Board. Statistics & Probability Letters, 80(19–20):ifc, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001926.

Anonymous:2010:EBm

[Ano10m] Anonymous. Editorial Board. Statistics & Probability Letters, 80(21–22):ifc, November 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002178.

Anonymous:2010:EBn

[Ano10n] Anonymous. Editorial Board. Statistics & Probability Letters, 80(23–24):ifc, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002877.

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. Statistics & Probability Letters, 81 (1):ii, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210003147.

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. Statistics & Probability Letters, 81(1):ifc, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210003123.

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board. Statistics & Probability Letters, 81(2):ifc, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521000355X.

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Statistics & Probability Letters, 81 (3):ifc, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521100006X.

Anonymous:2011:EBe

[Ano11e] Anonymous. Editorial Board. Statistics & Probability Letters, 81 (4):ifc, April 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521100023X.

Anonymous:2011:EBf

[Ano11f] Anonymous. Editorial Board. Statistics & Probability Letters, 81(5):ifc, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000502.

Anonymous:2011:EBg

[Ano11g] Anonymous. Editorial Board. Statistics & Probability Letters, 81 (6):ifc, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521100112X.

Anonymous:2011:EBh

[Ano11h] Anonymous. Editorial Board. Statistics & Probability Letters, 81(7):ifc, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001404.

Anonymous:2011:EBi

[Ano11i] Anonymous. Editorial Board. Statistics & Probability Letters, 81 (8):ifc, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001647.

Anonymous:2011:EBj

[Ano11j] Anonymous. Editorial Board. Statistics & Probability Letters, 81(9):ifc, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001933.

Anonymous:2011:EBk

[Ano11k] Anonymous. Editorial Board. Statistics & Probability Letters, 81(10):ifc, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002203.

Anonymous:2011:EBl

[Ano11l] Anonymous. Editorial Board. Statistics & Probability Letters, 81(11):ifc, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002574.

Anonymous:2011:EBm

[Ano11m] Anonymous. Editorial Board. Statistics & Probability Letters, 81(12):ifc, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003038.

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Statistics & Probability Letters, 82(1):ifc, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003476.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. Statistics & Probability Letters, 82(2):ifc, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003968.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. Statistics & Probability Letters, 82 (3):ifc, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000090.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Statistics & Probability Letters, 82 (4):ifc, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000417.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board. Statistics & Probability Letters, 82(5):ifc, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001009.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board. Statistics & Probability Letters, 82 (6):ifc, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001393.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board. Statistics & Probability Letters, 82(7):ifc, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001745.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board. Statistics & Probability Letters, 82 (8):ifc, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002027.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Statistics & Probability Letters, 82(9):ifc, September 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002428.

Anonymous:2012:EBj

[Ano12j] Anonymous. Editorial Board. Statistics & Probability Letters, 82(10):ifc, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002714.

Anonymous:2012:EBk

[Ano12k] Anonymous. Editorial Board. Statistics & Probability Letters, 82(11):ifc, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003033.

Anonymous:2012:EBl

[Ano12l] Anonymous. Editorial Board. Statistics & Probability Letters, 82(12):ifc, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003513.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Statistics & Probability Letters, 83(1):ifc, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004130.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board. Statistics & Probability Letters, 83(2):ifc, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004622.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board. Statistics & Probability Letters, 83 (3):ifc, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521300031X.

Anonymous:2013:EBd

[Ano13d] Anonymous. Editorial Board. Statistics & Probability Letters, 83 (4):ifc, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521300062X.

Anonymous:2013:EBe

[Ano13e] Anonymous. Editorial Board. Statistics & Probability Letters, 83(5):ifc, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000837.

Anonymous:2013:EBf

[Ano13f] Anonymous. Editorial Board. Statistics & Probability Letters, 83 (6):ifc, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001168.

Anonymous:2013:EBg

[Ano13g] Anonymous. Editorial Board. Statistics & Probability Letters, 83(7):ifc, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001582.

Anonymous:2013:EBh

[Ano13h] Anonymous. Editorial Board. Statistics & Probability Letters, 83 (8):ifc, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001879.

Anonymous:2013:EBi

[Ano13i] Anonymous. Editorial Board. Statistics & Probability Letters, 83(9):ifc, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002265.

Anonymous:2013:EBj

[Ano13j] Anonymous. Editorial Board. Statistics & Probability Letters, 83(10):ifc, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521300271X.

Anonymous:2013:EBk

[Ano13k] Anonymous. Editorial Board. Statistics & Probability Letters, 83(11):ifc, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521300299X.

Anonymous:2013:EBl

[Ano13l] Anonymous. Editorial Board. Statistics & Probability Letters, 83(12):ifc, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003428.

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. Statistics & Probability Letters, 84(??):ifc, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003763.

Anonymous:2014:EBb

[Ano14b] Anonymous. Editorial Board. Statistics & Probability Letters, 85(??):ifc, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521400008X.

Anonymous:2014:EBc

[Ano14c] Anonymous. Editorial Board. Statistics & Probability Letters, 86 (??):ifc, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000388.

Anonymous:2014:EBd

[Ano14d] Anonymous. Editorial Board. Statistics & Probability Letters, 87 (??):ifc, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000601.

Anonymous:2014:EBe

[Ano14e] Anonymous. Editorial Board. Statistics & Probability Letters, 88 (??):ifc, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521400090X.

Anonymous:2014:EBf

[Ano14f] Anonymous. Editorial Board. Statistics & Probability Letters, 89 (??):ifc, June 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001242.

Anonymous:2014:EBg

[Ano14g] Anonymous. Editorial Board. Statistics & Probability Letters, 90 (??):ifc, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001497.

Anonymous:2014:EBh

[Ano14h] Anonymous. Editorial Board. Statistics & Probability Letters, 91(??):ifc, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214001710.

Anonymous:2014:EBi

[Ano14i] Anonymous. Editorial Board. Statistics & Probability Letters, 92(??):ifc, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214002259.

Anonymous:2014:EBj

[Ano14j] Anonymous. Editorial Board. Statistics & Probability Letters, 93(??):ifc, October 1–142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214002661.

Anonymous:2014:EBk

[Ano14k] Anonymous. Editorial Board. Statistics & Probability Letters, 94(??):ifc, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003022.

Anonymous:2014:EB

[Ano14l] Anonymous. Editorial Board. Statistics & Probability Letters, 95(??):ifc, December 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003319.

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board. Statistics & Probability Letters, 96(??):ifc, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521400385X.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. Statistics & Probability Letters, 97(??):ifc, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214004295.

Anonymous:2015:EBc

[Ano15c] Anonymous. Editorial Board. Statistics & Probability Letters, 98 (??):ifc, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000395.

Anonymous:2015:EBd

[Ano15d] Anonymous. Editorial Board. Statistics & Probability Letters, 100(??):ifc, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000863.

Anonymous:2015:EBe

[Ano15e] Anonymous. Editorial Board. Statistics & Probability Letters, 101(??):ifc, June 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001030.

Anonymous:2015:EBf

[Ano15f] Anonymous. Editorial Board. Statistics & Probability Letters, 102(??):ifc, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001431.

Anonymous:2015:EBg

[Ano15g] Anonymous. Editorial Board. Statistics & Probability Letters, 103(??):ifc, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001662.

Anonymous:2015:EBh

[Ano15h] Anonymous. Editorial Board. Statistics & Probability Letters, 104(??):ifc, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002229.

Anonymous:2015:EBi

[Ano15i] Anonymous. Editorial Board. Statistics & Probability Letters, 105(??):ifc, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002539.

Anonymous:2015:EBj

[Ano15j] Anonymous. Editorial Board. Statistics & Probability Letters, 106(??):ifc, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002941.

Anonymous:2015:EB

[Ano15k] Anonymous. Editorial Board. Statistics & Probability Letters, 109:ifc, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521500348X.

Anonymous:2016:EBa

[Ano16a] Anonymous. Editorial Board. Statistics & Probability Letters, 108:ifc, January 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003673.

Anonymous:2016:EBb

[Ano16b] Anonymous. Editorial Board. Statistics & Probability Letters, 109:ifc, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521500396X.

Anonymous:2016:EBc

[Ano16c] Anonymous. Editorial Board. Statistics & Probability Letters, 110(??):ifc, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000171.

Anonymous:2016:EBd

[Ano16d] Anonymous. Editorial Board. Statistics & Probability Letters, 111(??):ifc, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000286.

Anonymous:2016:EBe

[Ano16e] Anonymous. Editorial Board. Statistics & Probability Letters, 112(??):ifc, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000468.

Anonymous:2016:EBf

[Ano16f] Anonymous. Editorial Board. Statistics & Probability Letters, 113(??):ifc, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300098.

Anonymous:2016:EBg

[Ano16g] Anonymous. Editorial Board. Statistics & Probability Letters, 114(??):ifc, July 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300359.

Anonymous:2016:EBh

[Ano16h] Anonymous. Editorial Board. Statistics & Probability Letters, 115(??):ifc, August 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300487.

Anonymous:2016:EBi

[Ano16i] Anonymous. Editorial Board. Statistics & Probability Letters, 116(??):ifc, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300736.

Anonymous:2016:EBj

[Ano16j] Anonymous. Editorial Board. Statistics & Probability Letters, 117(??):ifc, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301092.

Anonymous:2016:EBk

[Ano16k] Anonymous. Editorial Board. Statistics & Probability Letters, 118(??):ifc, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301389.

Anonymous:2016:EBl

[Ano16l] Anonymous. Editorial Board. Statistics & Probability Letters, 119(??):ifc, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301961.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board. Statistics & Probability Letters, 120(??):ifc, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302292.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board. Statistics & Probability Letters, 121(??):ifc, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302632.

Anonymous:2017:EBc

[Ano17c] Anonymous. Editorial Board. Statistics & Probability Letters, 122(??):ifc, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302978.

Anonymous:2017:EBd

[Ano17d] Anonymous. Editorial Board. Statistics & Probability Letters, 123(??):ifc, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300226.

Anonymous:2017:EBe

[Ano17e] Anonymous. Editorial Board. Statistics & Probability Letters, 124(??):ifc, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300627.

Anonymous:2017:EBf

[Ano17f] Anonymous. Editorial Board. Statistics & Probability Letters, 125(??):ifc, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301050.

Anonymous:2017:EBg

[Ano17g] Anonymous. Editorial Board. Statistics & Probability Letters, 126(??):ifc, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301402.

Anonymous:2017:EBh

[Ano17h] Anonymous. Editorial Board. Statistics & Probability Letters, 127(??):ifc, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301682.

Anonymous:2017:EBi

[Ano17i] Anonymous. Editorial Board. Statistics & Probability Letters, 128(??):ifc, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521730192X.

Anonymous:2017:EBj

[Ano17j] Anonymous. Editorial Board. Statistics & Probability Letters, 129(??):ifc, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302511.

Anonymous:2017:EBk

[Ano17k] Anonymous. Editorial board. Statistics & Probability Letters, 130(??):ifc, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302766.

Anonymous:2017:EBl

[Ano17l] Anonymous. Editorial board. Statistics & Probability Letters, 131(??):ifc, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302997.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Statistics & Probability Letters, 132(??):ifc, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303310.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Statistics & Probability Letters, 133(??):ifc, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521730353X.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Statistics & Probability Letters, 134(??):ii, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303802.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Statistics & Probability Letters, 135(??):ii, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300221.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Statistics & Probability Letters, 136(??):ii, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301597.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. Statistics & Probability Letters, 137(??):ii, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301275.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial Board. Statistics & Probability Letters, 138(??):ii, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301524.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board. Statistics & Probability Letters, 139(??):ii, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301962.

Anonymous:2018:EBi

[Ano18i] Anonymous. Editorial Board. Statistics & Probability Letters, 140(??):ii, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302219.

Anonymous:2018:EBj

[Ano18j] Anonymous. Editorial Board. Statistics & Probability Letters, 141(??):ii, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302438.

Anonymous:2018:EBk

[Ano18k] Anonymous. Editorial Board. Statistics & Probability Letters, 142(??):ii, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302736.

Anonymous:2018:EBl

[Ano18l] Anonymous. Editorial Board. Statistics & Probability Letters, 143(??):ii, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303006.

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Statistics & Probability Letters, 145(??):ii, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303675.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Statistics & Probability Letters, 146(??):ii, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303973.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Statistics & Probability Letters, 147(??):ii, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300045.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Statistics & Probability Letters, 148(??):ii, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300483.

Anonymous:2019:EBe

[Ano19e] Anonymous. Editorial Board. Statistics & Probability Letters, 149(??):ii, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300744.

Anonymous:2019:EBf

[Ano19f] Anonymous. Editorial Board. Statistics & Probability Letters, 149(??):iii, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301737.

Anonymous:2019:EBg

[Ano19g] Anonymous. Editorial Board. Statistics & Probability Letters, 150(??):ii, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300999.

Anonymous:2019:EBh

[Ano19h] Anonymous. Editorial Board. Statistics & Probability Letters, 151(??):ii, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301270.

Anonymous:2019:EBi

[Ano19i] Anonymous. Editorial Board. Statistics & Probability Letters, 152(??):ii, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301592.

Anonymous:2019:EBj

[Ano19j] Anonymous. Editorial Board. Statistics & Probability Letters, 153(??):ii, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301981.

Anonymous:2019:EBk

[Ano19k] Anonymous. Editorial Board. Statistics & Probability Letters, 154(??):Article 108572, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0167715219302184.

Anonymous:2019:EBl

[Ano19l] Anonymous. Editorial Board. Statistics & Probability Letters, 154(??):Article 108604, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0167715219302500.

Athreya:2015:LTE

[ANRW15] Krishna B. Athreya, Raoul Normand, Vivekananda Roy, and Sheng-Jhih Wu. Limit theorems for the estimation of L^{1} integrals using the Brownian motion. Statistics & Probability Letters, 100(??):42–47, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000498.

Antonini:2013:ABS

[ANV13] Rita Giuliano Antonini, Thuntida Ngamkham, and Andrei Volodin. On the asymptotic behavior of the sequence and series of running maxima from a real random sequence. Statistics & Probability Letters, 83(2):534–542, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003823.

Al-Omari:2012:REP

[AO12] Amer Ibrahim Al-Omari. Ratio estimation of the population mean using auxiliary information in simple random sampling and median ranked set sampling. Statistics & Probability Letters, 82(11):1883–1890, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002672.

Ahmad:2013:LCG

[AP13] Ola Ahmad and Jean-Charles Pinoli. On the linear combination of the Gaussian and Student’s t random field and the integral geometry of its excursion sets. Statistics & Probability Letters, 83(2):559–567, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200394X.

Adler:2018:WOS

[AP18] André Adler and Anthony G. Pakes. Weak and one-sided strong laws for random variables with infinite mean. Statistics & Probability Letters, 142(??):8–16, November 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302311.

Alfo:2013:FAF

[AR13] Marco Alfò and Irene Rocchetti. A flexible approach to finite mixture regression models for multivariate mixed responses. Statistics & Probability Letters, 83(7):1754–1758, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001259.

Athreya:2014:WMC

[AR14] Krishna B. Athreya and Vivekananda Roy. When is a Markov chain regenerative? Statistics & Probability Letters, 84(??):22– 26, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003192.

Adu:2018:URT

[AR18] N. Adu and G. Richardson. Unit roots test: Spatial model with long memory errors. Statistics & Probability Letters, 140(??):126–131, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301846.

Arbues:2019:CLT

[Arb19] Ignacio Arbués. Central limit theorem for the entries of products of random matrices without the positivity condition. Statistics & Probability Letters, 145(??):254–259, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303134.

Arnold:2011:GCD

[Arn11] Barry C. Arnold. The generalized Cantor distribution and its corresponding inverse distribution. Statistics & Probability Letters, 81(8):1098–1103, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000861.

Arnold:2012:AII

[Arn12] Barry C. Arnold. On the Amato inequality index. Statistics & Probability Letters, 82(8):1504–1506, August 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001630.

Abate:2014:EPC

[ART14] Alessandro Abate, Frank Redig, and Ilya Tkachev. On the effect of perturbation of conditional probabilities in total variation. Statistics & Probability Letters, 88(??):1–8, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000200.

Arvanitis:2017:NLT

[Arv17] Stelios Arvanitis. A note on the limit theory of a Dickey– Fuller unit root test with heavy tailed innovations. Statistics & Probability Letters, 126(??):198–204, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300883.

Arvanitis:2019:SLT

[Arv19] Stelios Arvanitis. Stable limit theory for the Gaussian QMLE in a non-stationary asymmetric GARCH model. Statistics & Probability Letters, 145(??):166–172, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303067.

Atkinson:2015:MHT

[AS15] Michael P. Atkinson and Dashi I. Singham. Multidimensional hitting time results for Brownian bridges with moving hyperplanar boundaries. Statistics & Probability Letters, 100(??):85– 92, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000565.

Adhikari:2018:UFE

[AS18] Kartick Adhikari and Koushik Saha. Universality in the fluctuation of eigenvalues of random circulant matrices. Statistics & Probability Letters, 138(??):1–8, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300567.

Al-Sharadqah:2018:CCR

[ASH18] A. Al-Sharadqah and K. C. Ho. Constrained Cramér–Rao lower bound in Errors-In Variables (EIV) models: Revisited. Statistics & Probability Letters, 135(??):118–126, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303231.

Ahn:2018:SGV

[ASK18] Kwang Woo Ahn, Natasha Sahr, and Soyoung Kim. Screening group variables in the proportional hazards model. Statistics & Probability Letters, 135(??):20–25, April 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303681.

Aboukhamseen:2016:MCD

[ASN16] S. M. Aboukhamseen, A. R. Soltani, and M. Najafi. Modelling cluster detection in spatial scan statistics: Formation of a spatial Poisson scanning window and an ADHD case study. Statistics & Probability Letters, 111(??):26–31, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215004083.

Aggrawal:2011:ODA

[ASP11] Manohar Aggrawal, Poonam Singh, and Mahesh Kumar Panda. A-optimal designs for an additive cubic model. Statistics & Probability Letters, 81(2):259–266, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002841.

Azmoodeh:2014:NSC

[ASVY14] Ehsan Azmoodeh, Tommi Sottinen, Lauri Viitasaari, and Adil Yazigi. Necessary and sufficient conditions for Hölder continuity of Gaussian processes. Statistics & Probability Letters, 94(??):230–235, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002776.

Abbas:2014:OBA

[AT14a] Kamran Abbas and Yincai Tang. Objective Bayesian analysis of the Frechet stress-strength model. Statistics & Probability Letters, 84(??):169–175, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300312X.

Amiri:2014:SSA

[AT14b] Aboubacar Amiri and Baba Thiam. A smoothing stochastic algorithm for quantile estimation. Statistics & Probability Letters, 93(??):116–125, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002181.

Aknouche:2015:WLS

[AT15a] Abdelhakim Aknouche and Nassim Touche. Weighted least squares-based inference for stable and unstable threshold power ARCH processes. Statistics & Probability Letters, 97(??):108– 115, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003800.

Artemiou:2015:USI

[AT15b] Andreas Artemiou and Lipu Tian. Using sliced inverse mean difference for sufficient dimension reduction. Statistics & Probability Letters, 106(??):184–190, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002667.

Azmoodeh:2010:WDF

[ATV10] Ehsan Azmoodeh, Heikki Tikanmäki, and Esko Valkeila. When does fractional Brownian motion not behave as a continuous function with bounded variation? Statistics & Probability Letters, 80(19–20):1543–1550, October 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001690.

Arnold:2013:ECM

[AV13a] Barry C. Arnold and Jose A. Villasenor. Exponential characterizations motivated by the structure of order statistics in samples of size two. Statistics & Probability Letters, 83(2):596– 601, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004002.

Arnold:2013:ORT

[AV13b] Barry C. Arnold and Jose A. Villasenor. On orthogonality of (X + Y) and X/(X + Y) rather than independence. Statistics & Probability Letters, 83(2):584–587, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004075.

Aneiros:2014:VSI

[AV14] Germán Aneiros and Philippe Vieu. Variable selection in infinite-dimensional problems. Statistics & Probability Letters, 94(??):12–20, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002363.

Arts:2017:AHR

[AvHZ17] Joachim Arts, Geert-Jan van Houtum, and Bert Zwart. The asymptotic hazard rate of sums of discrete random variables. Statistics & Probability Letters, 125(??):171–173, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521730055X.

Ahn:2017:UGS

[AWL17] Soohyun Ahn, Xinlei Wang, and Johan Lim. On unbalanced group sizes in cluster randomized designs using balanced ranked set sampling. Statistics & Probability Letters, 123(??):210– 217, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216302917.

Agahi:2017:SPI

[AY17] Hamzeh Agahi and Milad Yadollahzadeh. On stochastic pseudo-integrals with applications. Statistics & Probability Letters, 124(??):41–48, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300081.

Ayache:2013:SET

[Aya13] Antoine Ayache. Sharp estimates on the tail behavior of a multistable distribution. Statistics & Probability Letters, 83(3):680– 688, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004324.

An:2017:SSP

[AZ17] Baiguo An and Beibei Zhang. Simultaneous selection of predictors and responses for high dimensional multivariate linear regression. Statistics & Probability Letters, 127(??):173–177, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301475.

Azad:2013:PRF

[Aza13] Elahe Zohoorian Azad. Positions of the ranks of factors in certain finite long length words. Statistics & Probability Letters, 83(3):836–840, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004543.

Zou:2018:RWS

[aZLW18] Guang an Zou, Guangying Lv, and Jiang-Lun Wu. On the regularity of weak solutions to space-time fractional stochastic heat equations. Statistics & Probability Letters, 139(??):84– 89, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301469.

Azriel:2012:NRC

[Azr12] David Azriel. A note on the robustness of the continual reassessment method. Statistics & Probability Letters, 82(5):902– 906, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000363.

Azzone:2018:BDP

[Azz18] Giovanni Azzone. Big data and public policies: Opportunities and challenges. Statistics & Probability Letters, 136(??):116– 120, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300671.

Bagheri:2019:OSP

[BABF19] S. F. Bagheri, A. Asgharzadeh, E. Basiri, and A. J. Fernández. One-sample prediction regions for future record intervals. Statistics & Probability Letters, 153(??):15–20, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301348.

Bairamov:2012:MBD

[Bai12] Ismihan Bairamov. Majorization bounds for distribution functions. Statistics & Probability Letters, 82(10):1799–1806, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002167. See corrigendum [Bai13].

Bairamov:2013:CMB

[Bai13] Ismihan Bairamov. Corrigendum to “Majorization bounds for distribution functions” [Statist. Probab. Lett. 82 (2012) 1799– 1806]. Statistics & Probability Letters, 83(1):7–8, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003197. See [Bai12].

Baker:2012:DTS

[Bak12] David M. Baker. A Dubins type solution to the Skorokhod embedding problem. Statistics & Probability Letters, 82(6):1054– 1058, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000788.

Baker:2015:QPM

[Bak15] David M. Baker. Quantizations of probability measures and preservation of the convex order. Statistics & Probability Letters, 107:280–285, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003211.

Baker:2017:CND

[Bak17] Rose Baker. Creating new distributions by blunting cusps. Statistics & Probability Letters, 124(??):55–63, May 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730010X.

Baker:2018:ARE

[Bak18] Rose Baker. An adjusted random-effects model for binarydata meta-analysis. Statistics & Probability Letters, 134(??): 93–97, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303401.

Balov:2011:GMM

[Bal11] Nikolay Balov. A Gaussian mixed model for learning discrete Bayesian networks. Statistics & Probability Letters, 81(2):220– 230, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210003287.

Bao:2012:SCE

[Bao12] Zhigang Bao. Strong convergence of ESD for the generalized sample covariance matrices when p/n → 0. Statistics & Probability Letters, 82(5):894–901, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000223.

Bapat:2011:FPT

[Bap11] R. B. Bapat. On the first passage time of a simple random walk on a tree. Statistics & Probability Letters, 81(10):1552– 1558, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002021.

Barone:2011:GBD

[Bar11] P. Barone. A generalization of Bartlett’s decomposition. Statistics & Probability Letters, 81(3):371–381, March 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003342.

Baringhaus:2012:ART

[Bar12] Ludwig Baringhaus. On areas of random triangles. Statistics & Probability Letters, 82(4):733–738, April 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211004093.

Biswas:2010:ORA

[BB10] Atanu Biswas and Rahul Bhattacharya. An optimal responseadaptive design with dual constraints. Statistics & Probability Letters, 80(3–4):177–185, February 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209003903.

Babenko:2011:LBO

[BB11a] Alexandra Babenko and Eduard Belitser. Lower bound for the oracle projection posterior convergence rate. Statistics & Probability Letters, 81(2):175–180, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003214.

Berenhaut:2011:SOF

[BB11b] Kenneth S. Berenhaut and Lauren D. Bergen. Stochastic orderings, folded beta distributions and fairness in coin flips. Statistics & Probability Letters, 81(6):632–638, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000034.

Biswas:2011:ORA

[BB11c] Atanu Biswas and Rahul Bhattacharya. Optimal responseadaptive allocation designs in phase III clinical trials: Incorporating ethics in optimality. Statistics & Probability Letters, 81(8):1155–1160, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000940.

Banerjee:2012:CMH

[BB12] Buddhananda Banerjee and Atanu Biswas. On closeness of the Mantel–Haenszel estimator and the profile likelihood based estimator of the common odds ratio from multiple 2 × 2 tables. Statistics & Probability Letters, 82(11):1990–1993, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002246.

Balabdaoui:2014:LMP

[BB14a] Fadoua Balabdaoui and Cristina Butucea. On location mixtures with Pólya frequency components. Statistics & Probability Letters, 95(??):144–149, December 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002922.

Bali:2014:CNA

[BB14b] Juan Lucas Bali and Graciela Boente. Consistency of a numerical approximation to the first principal component projection pursuit estimator. Statistics & Probability Letters, 94(??):181–191, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002570.

Banerjee:2015:LIE

[BB15] Buddhananda Banerjee and Atanu Biswas. Linear increment in efficiency with the inclusion of surrogate endpoint. Statistics & Probability Letters, 96(??):102–108, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003265.

Bandyopadhyay:2016:FWC

[BB16] Uttam Bandyopadhyay and Atanu Biswas. Fixed-width confidence interval for two-stage response-adaptive designs in ridit analysis. Statistics & Probability Letters, 109:45–51, January 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003387.

Borovkov:2017:RVI

[BB17] Alexander A. Borovkov and Konstantin A. Borovkov. A refined version of the integro-local stone theorem. Statistics & Probability Letters, 123(??):153–159, April 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302887.

Berhoune:2018:SEF

[BB18] Kamila Berhoune and Nawel Bensmain. Sieves estimator of functional autoregressive process. Statistics & Probability Letters, 135(??):60–69, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303620.

Bierkens:2018:PDM

[BBCD^{+}18] Joris Bierkens, Alexandre Bouchard-Côté, Arnaud Doucet, Andrew B. Duncan, Paul Fearnhead, Thibaut Lienart, Gareth Roberts, and Sebastian J. Vollmer. Piecewise deterministic Markov processes for scalable Monte Carlo on restricted domains. Statistics & Probability Letters, 136(??):148–154, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830066X.

Beirlant:2016:BRT

[BBdWG16] J. Beirlant, A. Bardoutsos, T. de Wet, and I. Gijbels. Bias reduced tail estimation for censored Pareto type distributions. Statistics & Probability Letters, 109:78–88, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521530119X.

Basu:2012:LSD

[BBGH12] Riddhipratim Basu, Arup Bose, Shirshendu Ganguly, and Rajat Subhra Hazra. Limiting spectral distribution of block matrices with Toeplitz block structure. Statistics & Probability Letters, 82(7):1430–1438, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200137X.

Bianco:2015:RIP

[BBGMPG15] Ana M. Bianco, Graciela Boente, Wenceslao González-Manteiga, and Ana Pérez-González. Robust inference in partially linear models with missing responses. Statistics & Probability Letters, 97(??):88–98, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003733.

Bazarova:2014:CLT

[BBH14] Alina Bazarova, István Berkes, and Lajos Horváth. On the Central Limit Theorem for modulus trimmed sums. Statistics & Probability Letters, 86(??):61–67, March 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213004057.

Berestycki:2010:ASP

[BBHH10] J. Berestycki, É. Brunet, J. W. Harris, and S. C. Harris. The almost-sure population growth rate in branching Brownian motion with a quadratic breeding potential. Statistics & Probability Letters, 80(17–18):1442–1446, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001483.

Behboodian:2013:PIE

[BBHV13] J. Behboodian, Naveen Bansal, G. G. Hamedani, and Hans Volkmer. The Pitman inequality for exchangeable random vectors. Statistics & Probability Letters, 83(8):1825–1829, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001314.

Berenhaut:2011:DDD

[BBL11] Kenneth S. Berenhaut, John V. Baxley, and Robert G. Lyday. Deviations of discrete distributions and a question of Móri. Statistics & Probability Letters, 81(12):1940–1944, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100229X.

Bartolucci:2018:RLV

[BBM18] Francesco Bartolucci, Silvia Bacci, and Antonietta Mira. On the role of latent variable models in the era of big data. Statistics & Probability Letters, 136(??):165–169, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300683.

Bahlali:2019:TCI

[BBM19] Khaled Bahlali, Brahim Boufoussi, and Soufiane Mouchtabih. Transportation cost inequality for backward stochastic differential equations. Statistics & Probability Letters, 154(??): Article 108586, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219302329.

Bernardi:2017:IQS

[BBP17] Mauro Bernardi, Valeria Bignozzi, and Lea Petrella. On the L_{p}-quantiles for the Student t distribution. Statistics & Probability Letters, 128(??):77–83, September 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301566.

Barczy:2018:RVN

[BBP18] Mátyás Barczy, Zsuzsanna Bősze, and Gyula Pap. Regularly varying non-stationary Galton–Watson processes with immigration. Statistics & Probability Letters, 140(??):106–114, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301913.

Bianco:2011:REN

[BBS11] Ana M. Bianco, Graciela Boente, and Susana Sombielle. Robust estimation for nonparametric generalized regression.

Statistics & Probability Letters, 81(12):1986–1994, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002719.

Baldi:2011:GFW

[BC11] P. Baldi and L. Caramellino. General Freidlin–Wentzell Large Deviations and positive diffusions. Statistics & Probability Letters, 81(8):1218–1229, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001039.

Balan:2014:NIF

[BC14] Raluca M. Balan and Daniel Conus. A note on intermittency for the fractional heat equation. Statistics & Probability Letters, 95(??):6–14, ???? 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214002806.

Benazzoli:2019:NES

[BCD19] Chiara Benazzoli, Luciano Campi, and Luca Di Persio. ϵ– Nash equilibrium in stochastic differential games with meanfield interaction and controlled jumps. Statistics & Probability Letters, 154(??):Article 108522, November 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301506.

Batun-Cutz:2013:MDS

[BCGFR13] J. Batún-Cutz, G. González-Farías, and W.-D. Richter. Maximum distributions for I_{2,p}-symmetric vectors are skewed I_{1,p}- symmetric distributions. Statistics & Probability Letters, 83 (10):2260–2268, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002344.

Balakrishnan:2014:MPC

[BCI14] N. Balakrishnan, E. Cramer, and G. Iliopoulos. On the method of pivoting the CDF for exact confidence intervals with illustration for exponential mean under life-test with time constraints. Statistics & Probability Letters, 89(??):124–130, June 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000868.

Belarbi:2018:LLE

[BCL18] Faiza Belarbi, Souheyla Chemikh, and Ali Laksaci. Local linear estimate of the nonparametric robust regression in functional data. Statistics & Probability Letters, 134(??):128–133, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303498.

Belak:2014:WCO

[BCM14] Christoph Belak, Sören Christensen, and Olaf Menkens. Worstcase optimal investment with a random number of crashes. Statistics & Probability Letters, 90(??):140–148, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400114X.

Boente:2013:TGP

[BCMR13] Graciela Boente, Ricardo Cao, Wenceslao González Manteiga, and Daniela Rodriguez. Testing in generalized partially linear models: a robust approach. Statistics & Probability Letters, 83(1):203–212, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003367.

Bulla:2015:NAE

[BCNM15] Ingo Bulla, Christophe Chesneau, Fabien Navarro, and Tanya Mark. A note on the adaptive estimation of a bi-dimensional density in the case of knowledge of the copula density. Statistics & Probability Letters, 105(??):6–13, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001765.

Boudou:2010:CNC

[BCR10] A. Boudou, E. N. Cabral, and Y. Romain. Centered and non-centered principal component analyses in the frequency domain. Statistics & Probability Letters, 80(2):96–103, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003587.

Bucher:2010:NBA

[BD10] Axel Bücher and Holger Dette. A note on bootstrap approximations for the empirical copula process. Statistics & Probability Letters, 80(23–24):1925–1932, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002440.

Bharath:2011:TDB

[BD11] Karthik Bharath and Dipak K. Dey. Test to distinguish a Brownian motion from a Brownian bridge using Pólya tree process. Statistics & Probability Letters, 81(1):140–145, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002828.

Balakrishnan:2013:PCR

[BD13a] N. Balakrishnan and Katherine F. Davies. Pitman closeness results for Type-I censored data from exponential distribution. Statistics & Probability Letters, 83(12):2693–2698, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002964.

Bose:2013:DEU

[BD13b] Arup Bose and Santanu Dutta. Density estimation using bootstrap bandwidth selector. Statistics & Probability Letters, 83(1):245–256, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200332X.

Balabdaoui:2015:MLD

[BD15] Fadoua Balabdaoui and Cécile Durot. Marshall lemma in discrete convex estimation. Statistics & Probability Letters, 99(??): 143–148, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500022X.

Bezborodov:2017:MIM

[BD17] Viktor Bezborodov and Luca Di Persio. Maximal irreducibility measure for spatial birth-and-death processes. Statistics & Probability Letters, 125(??):25–32, June 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730041X.

Brumback:2010:ACC

[BDB^{+}10] Babette A. Brumback, Amy B. Dailey, Lyndia C. Brumback, Melvin D. Livingston, and Zhulin He. Adjusting for confounding by cluster using generalized linear mixed models. Statistics & Probability Letters, 80(21–22):1650–1654, November 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002002.

Barme-Delcroix:2019:UEI

[BDB19] Marie-Françoise Barme-Delcroix and Margarida Brito. Uniform estimation of isobars. Statistics & Probability Letters, 148(??): 94–100, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300185.

Barsotti:2014:ETM

[BDER14] Flavia Barsotti, Yohann De Castro, Thibault Espinasse, and Paul Rochet. Estimating the transition matrix of a Markov chain observed at random times. Statistics & Probability Letters, 94(??):98–105, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002478.

Bernardi:2017:CFC

[BDJ17] M. Bernardi, F. Durante, and P. Jaworski. CoVaR of families of copulas. Statistics & Probability Letters, 120(??):8–17, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301705.

Balabdaoui:2018:TCD

[BDK18] Fadoua Balabdaoui, Cécile Durot, and Babagnidé François Koladjo. Testing convexity of a discrete distribution. Statistics & Probability Letters, 137(??):8–13, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730384X.

Biswas:2017:FDR

[BDM17] A. Biswas, P. Das, and N. K. Mandal. Factorial designs robust against the presence of an aberration. Statistics & Probability Letters, 129(??):326–334, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302092.

Banys:2010:RSL

[BDP10] Povilas Banys, Youri Davydov, and Vygantas Paulauskas. Remarks on the SLLN for linear random fields. Statistics & Probability Letters, 80(5–6):489–496, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004519.

Baek:2014:IRO

[BDP14] Changryong Baek, Gustavo Didier, and Vladas Pipiras. On integral representations of operator fractional Brownian fields. Statistics & Probability Letters, 92(??):190–198, September 1– 262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001916.

Bystrov:2013:MAE

[BdS13] Victor Bystrov and Antonietta di Salvatore. Martingale approximation of eigenvalues for common factor representation. Statistics & Probability Letters, 83(1):233–237, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521200346X.

Beknazaryan:2019:MIE

[BDS19] Aleksandr Beknazaryan, Xin Dang, and Hailin Sang. On mutual information estimation for mixed-pair random variables. Statistics & Probability Letters, 148(??):9–16, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830405X.

Becheri:2016:PEP

[BDvdAW16] I. Gaia Becheri, Feike C. Drost, Ramon van den Akker, and Oliver Wichert. The power envelope of panel unit root tests in case stationary alternatives offset explosive ones. Statistics & Probability Letters, 109:1–8, January 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300754.

Barakat:2017:ABJ

[BE17] H. M. Barakat and M. A. Abd Elgawad. Asymptotic behavior of the joint record values, with applications. Statistics & Probability Letters, 124(??):13–21, May 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300068.

Belisle:2011:PGR

[Bél11] Claude Bélisle. On the polygon generated by n random points on a circle. Statistics & Probability Letters, 81(2):236–242, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003317.

Bellini:2012:IPG

[Bel12] Fabio Bellini. Isotonicity properties of generalized quantiles. Statistics & Probability Letters, 82(11):2017–2024, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002696.

Belot:2013:FCT

[Bel13] Gordon Belot. Failure of calibration is typical. Statistics & Probability Letters, 83(10):2316–2318, October 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002368.

Belalia:2016:APB

[Bel16a] Mohamed Belalia. On the asymptotic properties of the Bernstein estimator of the multivariate distribution function. Statistics & Probability Letters, 110(??):249–256, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003569.

Belitser:2016:OMA

[Bel16b] Eduard Belitser. Optimal measurement allocation under precision budget constraint. Statistics & Probability Letters, 117(??):46–53, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215303734.

Benhaddou:2016:DMF

[Ben16] Rida Benhaddou. Deconvolution model with fractional Gaussian noise: A minimax study. Statistics & Probability Letters, 117(??):201–208, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300815.

Benhaddou:2017:MCR

[Ben17] Rida Benhaddou. On minimax convergence rates under L^{p}-risk for the anisotropic functional deconvolution model. Statistics & Probability Letters, 130(??):120–125, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302419.

Benhaddou:2018:MLB

[Ben18] Rida Benhaddou. Minimax lower bounds for the simultaneous wavelet deconvolution with fractional Gaussian noise and unknown kernels. Statistics & Probability Letters, 140(??):91–95, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301834.

Berkaoui:2013:CGO

[Ber13a] Abdelkarem Berkaoui. On characterizing and generalizing the optional m-stability property for pricing set. Statistics & Probability Letters, 83(3):856–862, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004567.

Bernardi:2013:RMS

[Ber13b] Mauro Bernardi. Risk measures for skew normal mixtures. Statistics & Probability Letters, 83(8):1819–1824, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001375.

Beskos:2014:SMM

[Bes14] Alexandros Beskos. A stable manifold MCMC method for high dimensions. Statistics & Probability Letters, 90(??):46–52, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001096.

Bouzebda:2018:UBC

[BES18] Salim Bouzebda, Issam Elhattab, and Cheikh Tidiane Seck. Uniform in bandwidth consistency of nonparametric regression based on copula representation. Statistics & Probability Letters, 137(??):173–182, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300324.

Brito:2010:CET

[BF10] Margarida Brito and Ana Cristina Moreira Freitas. Consistent estimation of the tail index for dependent data. Statistics & Probability Letters, 80(23–24):1835–1843, December 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002324.

Burr:2017:UCR

[BF17] Michael A. Burr and Robert J. Fabrizio. Uniform convergence rates for halfspace depth. Statistics & Probability Letters, 124(??):33–40, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300093.

Battey:2016:ICS

[BFS16] Heather Battey, Qiang Feng, and Richard J. Smith. Improving confidence set estimation when parameters are weakly identified. Statistics & Probability Letters, 118(??):117–123, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303631.

Bordes:2011:UCN

[BG11] Laurent Bordes and Kossi Essona Gneyou. Uniform convergence of nonparametric regressions in competing risk models with right censoring. Statistics & Probability Letters, 81(11):1654–1663, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002100.

Bailey:2013:WCN

[BG13] N. Bailey and L. Giraitis. Weak convergence in the near unit root setting. Statistics & Probability Letters, 83(5):1411–1415, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000370.

Bibi:2015:CQMa

[BG15a] Abdelouahab Bibi and Ahmed Ghezal. Consistency of quasimaximum likelihood estimator for Markov-switching bilinear time series models. Statistics & Probability Letters, 99(??):192– 202, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000607.

Bibi:2015:CQMb

[BG15b] Abdelouahab Bibi and Ahmed Ghezal. Consistency of quasimaximum likelihood estimator for Markov-switching bilinear time series models. Statistics & Probability Letters, 100(??): 192–202, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000607.

Bischoff:2016:BCP

[BG16] Wolfgang Bischoff and Andreas Gegg. Boundary crossing probabilities for (q, d)-Slepian-processes. Statistics & Probability Letters, 118(??):139–144, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301043.

Bohnstedt:2019:DMD

[BG19] Marie Böhnstedt and Jutta Gampe. Detecting mortality deceleration: Likelihood inference and model selection in the gamma-Gompertz model. Statistics & Probability Letters, 150(??):68– 73, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300653.

Bose:2011:CTM

[BGHS11] Arup Bose, Suman Guha, Rajat Subhra Hazra, and Koushik Saha. Circulant type matrices with heavy tailed entries. Statistics & Probability Letters, 81(11):1706–1716, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100232X.

Bandara:2019:CML

[BGM19] Udika Bandara, Ryan Gill, and Riten Mitra. On computing maximum likelihood estimates for the negative binomial distribution. Statistics & Probability Letters, 148(??):54–58, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521930015X.

Biswas:2018:OPR

[BGO18] Arunangshu Biswas, Anindya Goswami, and Ludger Overbeck. Option pricing in a regime switching stochastic volatility model. Statistics & Probability Letters, 138(??):116–126, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301019.

Bai:2015:FLT

[BGT15] Shuyang Bai, Mamikon S. Ginovyan, and Murad S. Taqqu. Functional limit theorems for Toeplitz quadratic functionals of continuous time Gaussian stationary processes. Statistics & Probability Letters, 104(??):58–67, September 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001534.

Boufoussi:2010:ARQ

[BH10a] Brahim Boufoussi and Salah Hajji. An approximation result for a quasi-linear stochastic heat equation. Statistics & Probability Letters, 80(17–18):1369–1377, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001379.

Boufoussi:2010:SAN

[BH10b] Brahim Boufoussi and Salah Hajji. Successive approximation of neutral functional stochastic differential equations with jumps. Statistics & Probability Letters, 80(5–6):324–332, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004246.

Brumback:2011:MHE

[BH11] Babette A. Brumback and Zhulin He. The Mantel–Haenszel estimator adapted for complex survey designs is not dually consistent. Statistics & Probability Letters, 81(9):1465–1470, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001568.

Boufoussi:2012:NSF

[BH12] Brahim Boufoussi and Salah Hajji. Neutral stochastic functional differential equations driven by a fractional Brownian motion in a Hilbert space. Statistics & Probability Letters, 82(8):1549– 1558, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001563.

Bohm:2014:GRC

[BH14] Walter Böhm and Kurt Hornik. Generating random correlation matrices by the simple rejection method: Why it does not work. Statistics & Probability Letters, 87(??):27–30, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213004112.

Bhattacharya:2015:SPR

[BH15] Bhaskar Bhattacharya and Gareth Hughes. On shape properties of the receiver operating characteristic curve. Statistics & Probability Letters, 103(??):73–79, August 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001133.

Bao:2016:MCF

[BH16] Junshu Bao and Timothy E. Hanson. A mean-constrained finite mixture of normals model. Statistics & Probability Letters, 117(??):93–99, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215302492.

Bedini:2017:CDP

[BH17a] Matteo L. Bedini and Michael Hinz. Credit default prediction and parabolic potential theory. Statistics & Probability Letters, 124(??):121–125, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300329.

Bergman:2017:SDR

[BH17b] J. Bergman and B. Holmquist. Are the Sweden Democrats really Sweden’s largest party? A maximum likelihood ratio test on the simplex. Statistics & Probability Letters, 124(??):64–68, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300135.

Boufoussi:2017:SDD

[BH17c] Brahim Boufoussi and Salah Hajji. Stochastic delay differential equations in a Hilbert space driven by fractional Brownian motion. Statistics & Probability Letters, 129(??):222–229, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302109.

Boufoussi:2018:TIS

[BH18] Brahim Boufoussi and Salah Hajji. Transportation inequalities for stochastic heat equations. Statistics & Probability Letters, 139(??):75–83, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301329.

Beekenkamp:2019:SIP

[BH19] Thomas Beekenkamp and Tim Hulshof. Sharpness for inhomogeneous percolation on quasi-transitive graphs. Statistics & Probability Letters, 152(??):28–34, September 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300896.

Bharath:2013:NDE

[Bha13] Karthik Bharath. A note on density estimation for binary sequences. Statistics & Probability Letters, 83(12):2735–2742, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003118.

Bhadra:2017:EMS

[Bha17] Anindya Bhadra. An expectation-maximization scheme for measurement error models. Statistics & Probability Letters, 120(??):61–68, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215302790.

Bae:2010:ULL

[BHJ10] Jongsig Bae, Changha Hwang, and Doobae Jun. The uniform laws of large numbers for the tent map. Statistics & Probability Letters, 80(17–18):1437–1441, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001471.

Bae:2012:UCL

[BHJ12] Jongsig Bae, Changha Hwang, and Doobae Jun. The uniform Central Limit Theorem for the tent map. Statistics & Probability Letters, 82(5):1021–1027, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000508.

Benditkis:2018:FDR

[BHJ18] Julia Benditkis, Philipp Heesen, and Arnold Janssen. The false discovery rate (FDR) of multiple tests in a class room lecture. Statistics & Probability Letters, 134(??):29–35, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303115.

Bingham:2012:ERV

[BIK12] N. H. Bingham, Akihiko Inoue, and Yukio Kasahara. An explicit representation of Verblunsky coefficients. Statistics & Probability Letters, 82(2):403–410, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003579.

Bissiri:2010:CLF

[Bis10] Pier Giovanni Bissiri. Characterization of the law of a finite exchangeable sequence through the finite-dimensional distributions of the empirical measure. Statistics & Probability Letters, 80(17–18):1306–1312, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001197.

Bradley:2014:SMS

[BJ14] Richard C. Bradley and Zbigniew J. Jurek. The strong mixing and the selfdecomposability properties. Statistics & Probability Letters, 84(??):67–71, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003131.

Biswas:2016:MCR

[BJD16] Atanu Biswas, Jayant Jha, and Somak Dutta. Modelling circular random variables with a spike at zero. Statistics & Probability Letters, 109:194–201, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301504.

Balan:2016:SRN

[BJQS16] Raluca M. Balan, Maria Jolis, and Lluís Quer-Sardanyons. SPDEs with rough noise in space: Hölder continuity of the solution. Statistics & Probability Letters, 119(??):310–316, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301687.

Barrientos:2017:PCR

[BJW17] Andrés F. Barrientos, Alejandro Jara, and Claudia Wehrhahn. Posterior convergence rate of a class of Dirichlet process mixture model for compositional data. Statistics & Probability Letters, 120(??):45–51, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215300730.

Binkowski:2010:UCE

[BK10] Karol Binkowski and Andrzej Kozek. Uniform convergence of empirical characteristic functions in a complex domain with applications to option pricing. Statistics & Probability Letters, 80(5–6):270–276, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771520900409X.

Berend:2012:MMP

[BK12] Daniel Berend and Aryeh Kontorovich. The missing mass problem. Statistics & Probability Letters, 82(6):1102–1110, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000612.

Berend:2013:SEB

[BK13] Daniel Berend and Aryeh Kontorovich. A sharp estimate of the binomial mean absolute deviation with applications. Statistics & Probability Letters, 83(4):1254–1259, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000242.

Bobotas:2017:ESN

[BK17] Panayiotis Bobotas and Stavros Kourouklis. Estimation of the smallest normal variance with applications to variance components models. Statistics & Probability Letters, 131(??):38–45, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217302651.

Bivand:2018:BDS

[BK18] Roger Bivand and Konstantin Krivoruchko. Big data sampling and spatial analysis: “which of the two ladles, of fig-wood or gold, is appropriate to the soup and the pot?”. Statistics & Probability Letters, 136(??):87–91, ???? 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300579.

Bedbur:2019:CRS

[BK19a] S. Bedbur and U. Kamps. Confidence regions in step-stress experiments with multiple samples under repeated type-II censoring. Statistics & Probability Letters, 146(??):181–186, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303754.

Bobotas:2019:DMM

[BK19b] Panayiotis Bobotas and Markos V. Koutras. Distributions of the minimum and the maximum of a random number of random variables. Statistics & Probability Letters, 146(??):57–64, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303481.

Bedbur:2019:SCR

[BKL19] S. Bedbur, U. Kamps, and J. M. Lennartz. On a smallest confidence region for a location-scale parameter in progressively type-II censored lifetime experiments. Statistics & Probability Letters, 154(??):Article 108545, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301828.

Baltagi:2015:TSN

[BKP15] Badi H. Baltagi, Chihwa Kao, and Bin Peng. On testing for sphericity with non-normality in a fixed effects panel data model. Statistics & Probability Letters, 98(??):123–130, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214004209.

BenHajria:2018:PCB

[BKR18] Raja Ben Hajria, Salah Khardani, and Hamdi Raïssi. A power comparison between autocorrelation based tests. Statistics & Probability Letters, 143(??):1–6, December 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302591.

Brinda:2019:HLI

[BKY19] W. D. Brinda, Jason M. Klusowski, and Dana Yang. Hölder’s identity. Statistics & Probability Letters, 148(??):150–154, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300148.

Baltagi:2010:SSR

[BL10a] Badi H. Baltagi and Long Liu. Spurious spatial regression with equal weights. Statistics & Probability Letters, 80(21–22):1640–1642, November 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001975.

Bhattacharya:2010:ANM

[BL10b] Rabi Bhattacharya and Lizhen Lin. An adaptive nonparametric method in benchmark analysis for bioassay and environmental studies. Statistics & Probability Letters, 80(23–24): 1947–1953, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002476.

Bibi:2010:SCA

[BL10c] Abdelouahab Bibi and Ines Lescheb. Strong consistency and asymptotic normality of least squares estimators for PGARCH and PARMA–PGARCH models. Statistics & Probability Letters, 80(19–20):1532–1542, October 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001689.

Bo:2011:MFP

[BL11] Lijun Bo and Mario Lefebvre. Mean first passage times of two-dimensional processes with jumps. Statistics & Probability Letters, 81(8):1183–1189, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100099X.

Baltagi:2012:HTP

[BL12] Badi H. Baltagi and Long Liu. The Hausman–Taylor panel data model with serial correlation. Statistics & Probability Letters, 82(7):1401–1406, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001095.

Bibi:2014:NIP

[BL14a] Abdelouahab Bibi and Ines Lescheb. A note on integrated periodic GARCH processes. Statistics & Probability Letters, 87(??): 121–124, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000182.

Brandes:2014:NCS

[BL14b] Dirk-Philip Brandes and Alexander Lindner. Non-causal strictly stationary solutions of random recurrence equations. Statistics & Probability Letters, 94(??):113–118, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002442.

Brockwell:2015:CPS

[BL15] Peter J. Brockwell and Alexander Lindner. CARMA processes as solutions of integral equations. Statistics & Probability Letters, 107:221–227, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300171.

Bai:2017:PRB

[BL17] Long Bai and Li Luo. Parisian ruin of the Brownian motion risk model with constant force of interest. Statistics & Probability Letters, 120(??):34–44, January 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301754.

Bar-Lev:2016:CGL

[BLB16] Shaul K. Bar-Lev and Daoud Bshouty. A characterization of the generalized Laplace distribution by constant regression on the sample mean. Statistics & Probability Letters, 113(??):79– 83, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215300651.

Bar-Lev:2013:SPB

[BLKL13] Shaul K. Bar-Lev, Offer Kella, and Andreas Löpker. Small parameter behavior of families of distributions. Statistics & Probability Letters, 83(3):783–789, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004439.

Bar-Lev:2010:LBS

[BLL10] Shaul K. Bar-Lev and Gérard Letac. The limiting behavior of some infinitely divisible exponential dispersion models. Statistics & Probability Letters, 80(23–24):1870–1874, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002361.

Bao:2014:DLN

[BLL14] Yafei Bao, Wenjian Luo, and Yihui Lu. On the dependable level of the negative survey. Statistics & Probability Letters, 89(??): 31–40, June 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000728.

Block:2014:LFR

[BLS14] Henry W. Block, Naftali A. Langberg, and Thomas H. Savits. The limiting failure rate for a convolution of gamma distributions. Statistics & Probability Letters, 94(??):176–180, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521400220X.

Bai:2011:MCT

[BLWZ11] Zhidong Bai, Heng Li, Wing-Keung Wong, and Bingzhi Zhang. Multivariate causality tests with simulation and application. Statistics & Probability Letters, 81(8):1063–1071, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000770.

Bao:2013:EPS

[BLZ13] Yafei Bao, Wenjian Luo, and Xin Zhang. Estimating positive surveys from negative surveys. Statistics & Probability Letters, 83(2):551–558, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200404X.

Bellier:2010:SCM

[BM10a] Edwige Bellier and Pascal Monestiez. A spatial covariance model with a single wave effect and a finite range. Statistics & Probability Letters, 80(17–18):1343–1347, September 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001240.

Bi:2010:IPD

[BM10b] L. Bi and A. Mukherjea. Identification of parameters and the distribution of the minimum of the tri-variate normal. Statistics & Probability Letters, 80(23–24):1819–1826, December 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002300.

Biswas:2010:DMN

[BM10c]  Atanu Biswas and Saumen Mandal.  Descriptive mea-

sures for nominal categorical variables. Statistics & Probability Letters, 80(11–12):982–989, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000520.

Bi:2011:PDI

[BM11] L. Bi and A. Mukherjea. Poisson distributions: Identification of parameters from the distribution of the maximum and a conjecture on the partial sums of the power series for exp(x). Statistics & Probability Letters, 81(5):611–613, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003226.

Borovkov:2012:ULS

[BM12] Konstantin Borovkov and Shaun McKinlay. The uniform law for sojourn measures of random fields. Statistics & Probability Letters, 82(9):1745–1749, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001861.

Beghin:2013:LDF

[BM13a] Luisa Beghin and Claudio Macci. Large deviations for fractional Poisson processes. Statistics & Probability Letters, 83(4):1193– 1202, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000187.

Behboodian:2013:CTS

[BM13b] Javad Behboodian and Reza Modarres. On a characterization theorem of symmetry about a point. Statistics & Probability Letters, 83(9):2057–2059, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300179X.

Bahamyirou:2015:DBB

[BM15] Asma Bahamyirou and Éric Marchand. On the discrepancy between Bayes credibility and frequentist probability of coverage. Statistics & Probability Letters, 97(??):63–68, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003691.

Beghin:2017:ARM

[BM17] Luisa Beghin and Claudio Macci. Asymptotic results for a multivariate version of the alternative fractional Poisson process. Statistics & Probability Letters, 129(??):260–268, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302134.

Boukhari:2018:CSS

[BM18] Fakhreddine Boukhari and Dounyazed Malti. Convergence of series of strongly integrable random variables and applications. Statistics & Probability Letters, 137(??):191–200, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300403.

Borges:2016:GTS

[BMB16] Patrick Borges, Fabio Fajardo Molinares, and Marcelo Bourguignon. A geometric time series model with inflated-parameter Bernoulli counting series. Statistics & Probability Letters, 119(??):264–272, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215304399.

Bandyopadhyay:2013:AAB

[BMD13] Uttam Bandyopadhyay, Shirsendu Mukherjee, and Dhiman Dutta. An approach alternative to the binomial UMPU test. Statistics & Probability Letters, 83(1):1–6, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003148.

Bandyopadhyay:2015:RSS

[BMD15] Uttam Bandyopadhyay, Shirsendu Mukherjee, and Dhiman Dutta. Reducing sample size in negative binomial UMPU test. Statistics & Probability Letters, 103(??):57–61, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001327.

Baek:2014:TDB

[BML14] Jung Woo Baek, Seung Ki Moon, and Ho Woo Lee. A timedependent busy period queue length formula for the M/E_{k} /1 queue. Statistics & Probability Letters, 87(??):98–104, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000157.

Benchaira:2015:ANE

[BMN15] Souad Benchaira, Djamel Meraghni, and Abdelhakim Necir. On the asymptotic normality of the extreme value index for righttruncated data. Statistics & Probability Letters, 107:378–384, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521530081X.

Benchaira:2016:KET

[BMN16] Souad Benchaira, Djamel Meraghni, and Abdelhakim Necir. Kernel estimation of the tail index of a right-truncated Paretotype distribution. Statistics & Probability Letters, 119(??):186– 193, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301493.

Borovkov:2018:NRB

[BMNZ18] Konstantin Borovkov, Yuliya Mishura, Alexander Novikov, and Mikhail Zhitlukhin. New and refined bounds for expected maxima of fractional Brownian motion. Statistics & Probability Letters, 137(??):142–147, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300361.

Basu:2010:HTT

[BMP10a] A. Basu, A. Mandal, and L. Pardo. Hypothesis testing for two discrete populations based on the Hellinger distance. Statistics & Probability Letters, 80(3–4):206–214, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003940.

Berg:2010:SV

[BMP10b] Arthur Berg, Timothy L. McMurry, and Dimitris N. Politis. Subsampling p-values. Statistics & Probability Letters, 80(17–18):1358–1364, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001276.

Boente:2010:ABG

[BMS10] Graciela Boente, Julieta Molina, and Mariela Sued. On the asymptotic behavior of general projection-pursuit estimators under the common principal components model. Statistics & Probability Letters, 80(3–4):228–235, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003976.

Brigo:2016:MMV

[BMS16] Damiano Brigo, Jan-Frederik Mai, and Matthias Scherer. Markov multi-variate survival indicators for default simulation as a new characterization of the Marshall–Olkin law. Statistics & Probability Letters, 114(??):60–66, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521530167X.

Borovkov:2010:LCS

[BN10] Konstantin Borovkov and Serguei Novak. On limiting cluster size distributions for processes of exceedances for stationary sequences. Statistics & Probability Letters, 80(23–24): 1814–1818, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002294.

Balan:2016:IWE

[BN16] Raluca M. Balan and Cheikh B. Ndongo. Intermittency for the wave equation with Lévy white noise. Statistics & Probability Letters, 109:214–223, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003880.

Barmalzan:2016:LRD

[BNB16] Ghobad Barmalzan, Amir T. Payandeh Najafabadi, and Narayanaswamy Balakrishnan. Likelihood ratio and dispersive orders for smallest order statistics and smallest claim amounts from heterogeneous Weibull sample. Statistics & Probability Letters, 110(??):1–7, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003776.

Bhattacharjee:2013:IIE

[BNM13a] Subarna Bhattacharjee, Asok K. Nanda, and Satya Kr. Misra. Inequalities involving expectations to characterize distributions. Statistics & Probability Letters, 83(9):2113–2118, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001831.

Bhattacharjee:2013:RAU

[BNM13b] Subarna Bhattacharjee, Asok K. Nanda, and Satya Kr. Misra. Reliability analysis using ageing intensity function. Statistics & Probability Letters, 83(5):1364–1371, May 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000175.

Brown:2013:LEM

[BNSA13] Patrick E. Brown, Paul Nguyen, Jamie Stafford, and Shiva Ashta. Local- EM and mismeasured data. Statistics & Probability Letters, 83(1):135–140, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003355.

Brzozowski:2019:SWW

[BO19] Michał Brzozowski and Adam Osekowski. Sharp weighted weak type (∞, ∞) inequality for differentially subordinate martingales. Statistics & Probability Letters, 154(??):Article 108561, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521930207X.

Bobryk:2010:CMN

[Bob10] Roman V. Bobryk. On closure methods in nonlinear stochastic dynamics. Statistics & Probability Letters, 80(23–24): 1733–1738, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002117.

Bobkov:2013:EAR

[Bob13] Sergey G. Bobkov. Entropic approach to E. Rio’s Central Limit Theorem for W_{2} transport distance. Statistics & Probability Letters, 83(7):1644–1648, July 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300103X.

Bogucki:2015:SCW

[Bog15] Robert Bogucki. Suprema of canonical Weibull processes. Statistics & Probability Letters, 107:253–263, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003223.

Barakat:2017:NFE

[BOK17] H. M. Barakat, A. R. Omar, and O. M. Khaled. A new flexible extreme value model for modeling the extreme value data, with an application to environmental data. Statistics & Probability Letters, 130(??):25–31, November 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302353.

Bongiorno:2012:NFS

[Bon12] Enea G. Bongiorno. A note on fuzzy set-valued Brownian motion. Statistics & Probability Letters, 82(4):827–832, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000211.

Borisov:2014:NRW

[Bor14] I. S. Borisov. A note on a result by W. Hoeffding. Statistics & Probability Letters, 87(??):7–11, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213004161.

Borisov:2016:RCH

[Bor16] I. S. Borisov. The rate of convergence in Hoeffding’s theorem and some applications. Statistics & Probability Letters, 109:99– 105, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215300134.

Brzozowski:2017:SWW

[BOR17] Michał Brzozowski, Adam Osȩkowski, and Mateusz Rapicki. Sharp weighted weak-norm estimates for maximal functions. Statistics & Probability Letters, 131(??):93–101, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302717.

Borquez:2018:RPP

[Bór18] R. Bórquez. Recurrence property and pointwise convergence of martingales. Statistics & Probability Letters, 135(??):51– 53, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303735.

Bose:2014:SMA

[Bos14] Sudip Bose. A stochastic model for assessing the utility of chance. Statistics & Probability Letters, 93(??):72–77, October 1–142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001886.

Boukeloua:2015:RMS

[Bou15] Mohamed Boukeloua. Rates of mean square convergence of density and failure rate estimators under twice censoring. Statistics & Probability Letters, 106(??):121–128, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002503.

Boucher:2016:NMD

[Bou16] Thomas R. Boucher. A note on martingale deviation bounds. Statistics & Probability Letters, 111(??):8–11, April 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301073.

Bouzebda:2017:KRE

[Bou17] Salim Bouzebda. Kac’s representation for empirical copula process from an asymptotic viewpoint. Statistics & Probability Letters, 123(??):107–113, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302905.

Boukhadra:2018:HKD

[Bou18] Omar Boukhadra. On heat kernel decay for the random conductance model. Statistics & Probability Letters, 133(??):23– 27, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303139.

Bowden:2013:EKD

[Bow13] Roger Bowden. Entropic kernels for data smoothing. Statistics & Probability Letters, 83(3):916–922, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004555.

Bowman:2016:SIF

[Bow16] Dale Bowman. Statistical inference for familial disease models assuming exchangeability. Statistics & Probability Letters, 119(??):220–225, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301511.

Bowden:2017:DSL

[Bow17] Roger J. Bowden. Distribution spread and location metrics using entropic separation. Statistics & Probability Letters, 124(??): 148–153, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300342.

Bowman:2018:BQI

[Bow18] Adrian W. Bowman. Big questions, informative data, excellent science. Statistics & Probability Letters, 136(??):34–36, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300622.

Bignozzi:2015:SMS

[BP15] Valeria Bignozzi and Giovanni Puccetti. Studying mixability with supermodular aggregating functions. Statistics & Probability Letters, 100(??):48–55, May 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000474.

Barabesi:2019:PTD

[BP19a] Lucio Barabesi and Luca Pratelli. On the properties of a Takács distribution. Statistics & Probability Letters, 148(??):66–73, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300112.

Basrak:2019:NVC

[BP19b] Bojan Basrak and Hrvoje Planinić. A note on vague convergence of measures. Statistics & Probability Letters, 153(??):180– 186, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301622.

Bhattacharya:2019:SVG

[BP19c] Ayan Bhattacharya and Zbigniew Palmowski. Slower variation of the generation sizes induced by heavy-tailed environment for geometric branching. Statistics & Probability Letters, 154(??):Article 108550, November 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301877.

Belzunce:2013:CCS

[BPRS13] Félix Belzunce, José F. Pinar, José M. Ruiz, and Miguel A. Sordo. Comparison of concentration for several families of income distributions. Statistics & Probability Letters, 83(4):1036– 1045, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004841.

Bercu:2014:OUD

[BPS14] Bernard Bercu, Frédéric Proïa, and Nicolas Savy. On Ornstein– Uhlenbeck driven by Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 85(??):36–44, February 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003635.

Budhiraja:2017:MLE

[BPS17] Sonal Budhiraja, Biswabrata Pradhan, and Debasis Sengupta. Maximum likelihood estimators under progressive Type-I interval censoring. Statistics & Probability Letters, 123(??):202– 209, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216303078.

Biscio:2018:NGP

[BPW18] Christophe Ange Napoléon Biscio, Arnaud Poinas, and Rasmus Waagepetersen. A note on gaps in proofs of central limit theorems. Statistics & Probability Letters, 135(??):7–10, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303632.

Beranger:2019:EPM

[BPXS19a] B. Beranger, S. A. Padoan, Y. Xu, and S. A. Sisson. Extremal properties of the multivariate extended skew-normal distribution, Part B. Statistics & Probability Letters, 147(??):105– 114, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303870.

Beranger:2019:EPU

[BPXS19b] B. Beranger, S. A. Padoan, Y. Xu, and S. A. Sisson. Extremal properties of the univariate extended skew-normal distribution, Part A. Statistics & Probability Letters, 147(??):73–82, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303869.

Bercu:2017:LDO

[BR17a] Bernard Bercu and Adrien Richou. Large deviations for the Ornstein–Uhlenbeck process without tears. Statistics & Probability Letters, 123(??):45–55, April 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302838.

Bhati:2017:SNC

[BR17b] Deepesh Bhati and Sreenivasan Ravi. Some new characterizations of max domains of attraction of Fréchet and log-Fréchet laws under power normalization. Statistics & Probability Letters, 130(??):92–99, November 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302407.

Bradley:2011:NTM

[Bra11a] Richard C. Bradley. A note on two measures of dependence. Statistics & Probability Letters, 81(12):1823–1826, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002458.

Braverman:2011:IDD

[Bra11b] Michael Braverman. On infinitely divisible distributions with light tails of Lévy measures. Statistics & Probability Letters, 81(11):1648–1653, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002148.

Bakouch:2012:EEB

[BRA^{+}12a] Hassan S. Bakouch, Miroslav M. Ristić, A. Asgharzadeh, L. Esmaily, and Bander M. Al-Zahrani. An exponentiated exponential binomial distribution with application. Statistics & Probability Letters, 82(6):1067–1081, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000764.

Bratiichuk:2012:GSF

[Bra12b] Mykola Bratiichuk. On the Gerber–Shiu function for a risk model with multi-layer dividend strategy. Statistics & Probability Letters, 82(3):496–504, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003555.

Braione:2016:TVL

[Bra16] Manuela Braione. A time-varying long run HEAVY model. Statistics & Probability Letters, 119(??):36–44, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301225.

Borges:2014:CPT

[BRBB14] Patrick Borges, Josemar Rodrigues, Narayanaswamy Balakrishnan, and Jorge Bazán. A COM–Poisson type generalization of the binomial distribution and its properties and applications. Statistics & Probability Letters, 87(??):158–166, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000303. See erratum [BRBB23].

Borges:2023:ECP

[BRBB23] P. Borges, J. Rodrigues, N. Balakrishnan, and J. Bazán. Erratum to “A COM–Poisson type generalization of the binomial distribution and its properties and applications” [Statist. Probab. Lett. 87 (2014) 158–166]. Statistics & Probability Letters, 201(??):??, October 2023. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715223001074.

See [BRBB14].

Breto:2012:IDM

[Bre12a] Carles Bretó. On the infinitesimal dispersion of multivariate Markov counting systems. Statistics & Probability Letters, 82(4): 720–725, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211004111.

Breto:2012:TCR

[Bre12b] Carles Bretó. Time changes that result in multiple points in continuous-time Markov counting processes. Statistics & Probability Letters, 82(12):2229–2234, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200301X.

Breto:2014:ISF

[Bre14a] Carles Bretó. On idiosyncratic stochasticity of financial leverage effects. Statistics & Probability Letters, 91(??):20–26, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001345.

Breto:2014:TCP

[Bre14b] Carles Bretó. Trajectory composition of Poisson time changes and Markov counting systems. Statistics & Probability Letters, 88(??):91–98, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214000492.

Brill:2015:NST

[Bri15] Percy H. Brill. Note on the service time in an M/G/1 queue with bounded workload. Statistics & Probability Letters, 96(??):162– 169, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003393.

Brockwell:2007:URM

[Bro07] A. E. Brockwell. Universal residuals: a multivariate transformation. Statistics & Probability Letters, 77(14):1473–1478, August 2007. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715207000715. See acknowledgement of

priority [Bro11].

Brockwell:2011:APU

[Bro11] A. E. Brockwell. Acknowledgement of priority to: “Universal residuals: a multivariate transformation. Statistics & Probability Letters 27, 2007, 1473–1478”. Statistics & Probability Letters, 81(12):1822, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001477. See [Bro07].

Balanzario:2014:RSG

[BRO14a] Eugenio P. Balanzario, Rosalva Mendoza Ramírez, and Jorge Sánchez Ortiz. The randomly stopped geometric Brownian motion. Statistics & Probability Letters, 90(??):85–92, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001060.

Broniatowski:2014:MDE

[Bro14b] Michel Broniatowski. Minimum divergence estimators, maximum likelihood and exponential families. Statistics & Probability Letters, 93(??):27–33, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002168.

Bryc:2014:IRB

[Bry14] Włodek Bryc. On integration with respect to the q-Brownian motion. Statistics & Probability Letters, 94(??):257–266, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002764.

Boente:2010:RLE

[BRZ10] Graciela Boente, Marcelo Ruiz, and Ruben H. Zamar. On a robust local estimator for the scale function in heteroscedastic nonparametric regression. Statistics & Probability Letters, 80(15–16):1185–1195, August 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000957.

Bairamov:2010:NNM

[BS10] I. Bairamov and A. Stepanov. Numbers of near-maxima for the bivariate case. Statistics & Probability Letters, 80(3–4): 196–205, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003939.

Bai:2011:AES

[BS11] Xiaodong Bai and Lixin Song. The asymptotic estimate for the sum of two correlated classes of discounted aggregate claims with heavy tails. Statistics & Probability Letters, 81(12):1891–1898, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002525.

Bai:2012:ABR

[BS12a] Xiaodong Bai and Lixin Song. Asymptotic behavior of random time absolute ruin probability with D ∩ L tailed and conditionally independent claim sizes. Statistics & Probability Letters, 82(9):1718–1726, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200185X.

Bandyopadhyay:2012:CTR

[BS12b] Antar Bandyopadhyay and Farkhondeh Sajadi. The connectivity threshold of random geometric graphs with Cantor distributed vertices. Statistics & Probability Letters, 82(12):2103–2107, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002878.

Balakrishna:2014:CBE

[BS14a] N. Balakrishna and K. Shiji. On a class of bivariate exponential distributions. Statistics & Probability Letters, 85(??):153–160, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003908.

Baumann:2014:FLR

[BS14b] Hendrik Baumann and Werner Sandmann. On finite long run costs and rewards in infinite Markov chains. Statistics & Probability Letters, 91(??):41–46, August 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001369.

Balakrishnan:2015:GQB

[BS15a] N. Balakrishnan and H. Y. So. A generalization of quantilebased skew logistic distribution of van Staden and King. Statistics & Probability Letters, 107:44–51, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002813.

Baran:2015:ODP

[BS15b] S. Baran and M. Stehlík. Optimal designs for parameters of shifted Ornstein–Uhlenbeck sheets measured on monotonic sets. Statistics & Probability Letters, 99(??):114–124, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000127.

Basrak:2015:ERV

[BS15c] Bojan Basrak and Drago Spoljarić. Extremes of random variables observed in renewal times. Statistics & Probability Letters, 97(??):216–221, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004015.

Belomestny:2015:SSE

[BS15d] Denis Belomestny and John Schoenmakers. Statistical Skorohod embedding problem: Optimality and asymptotic normality. Statistics & Probability Letters, 104(??):169–180, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001728.

Bernackaite:2015:EMT

[BS15e] Emilija Bernackaite and Jonas Siaulys. The exponential moment tail of inhomogeneous renewal process. Statistics & Probability Letters, 97(??):9–15, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003678.

Block:2015:FRC

[BS15f] Henry W. Block and Thomas H. Savits. The failure rate of a convolution dominates the failure rate of any IFR component. Statistics & Probability Letters, 107:142–144, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002758.

Balakrishnan:2017:SDS

[BS17a] Narayanaswamy Balakrishnan and Alessandro Selvitella. Symmetry of a distribution via symmetry of order statistics. Statistics & Probability Letters, 129(??):367–372, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302274.

Bogso:2017:WDS

[BS17b] Antoine-Marie Bogso and Patrice Takam Soh. Weak decreasing stochastic order. Statistics & Probability Letters, 126(??):

49–58, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300767.

Behme:2018:LSI

[BS18a] Anita Behme and Alexander Schnurr. Laplace symbols and invariant distributions. Statistics & Probability Letters, 137(??): 217–223, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300348.

Bouzebda:2018:NRM

[BS18b] Salim Bouzebda and Yousri Slaoui. Nonparametric recursive method for kernel-type function estimators for spatial data. Statistics & Probability Letters, 139(??):103–114, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301378.

Burkschat:2018:DIC

[BS18c] M. Burkschat and F. J. Samaniego. Dynamic IFR concepts for coherent systems. Statistics & Probability Letters, 142(??):1–7, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302281.

Banerjee:2019:BTI

[BS19a] Debapratim Banerjee and Matteo Sordello. A Bernstein type inequality for sums of selections from three dimensional arrays. Statistics & Probability Letters, 154(??):Article 108534, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301683.

Bingham:2019:DW

[BS19b] N. H. Bingham and Tasmin L. Symons. Dimension walks on d × R. Statistics & Probability Letters, 147(??):12–17, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303614.

Bouzebda:2019:LMD

[BS19c] Salim Bouzebda and Yousri Slaoui. Large and moderate deviation principles for recursive kernel estimators of a regression function for spatial data defined by stochastic approximation method. Statistics & Probability Letters, 151(??):17–28, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300835.

Bolanos-Servin:2017:RSP

[BSCQ17] Jorge R. Bolaños-Servin, Raffaella Carbone, and Roberto Quezada. On reducibility and spectral properties of circulant Markov processes. Statistics & Probability Letters, 123(??):27– 33, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216302784. See corrigendum

[BSCQ18].

Bolanos-Servin:2018:CRS

[BSCQ18] Jorge R. Bolaños-Servin, Raffaella Carbone, and Roberto Quezada. Corrigendum to “On reducibility and spectral properties of circulant Markov processes” [Statist. Probab. Lett. 123 (2017) 27–33]. Statistics & Probability Letters, 135(??):140, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303504. See [BSCQ17].

Bednarski:2019:SFC

[BSK19] Tadeusz Bednarski and Magdalena Skolimowska-Kulig. On scale Fisher consistency of maximum likelihood estimator for the exponential regression model under arbitrary frailty. Statistics & Probability Letters, 150(??):9–12, July 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300549.

Bao:2013:NIP

[BSL13] Zhenhua Bao, Lixin Song, and He Liu. A note on the inflatedparameter binomial distribution. Statistics & Probability Letters, 83(8):1911–1914, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001478.

Balanzario:2010:SCB

[BSO10] Eugenio P. Balanzario and Jorge Sánchez-Ortiz. Sufficient conditions for Benford’s law. Statistics & Probability Letters, 80(23–24):1713–1719, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002087.

Baran:2013:ODP

[BSS13] Sándor Baran, Kinga Sikolya, and Milan Stehlík. On the optimal designs for the prediction of Ornstein–Uhlenbeck sheets. Statistics & Probability Letters, 83(6):1580–1587, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521300076X.

Block:2013:MDS

[BSWS13] Henry W. Block, Thomas H. Savits, Jie Wang, and Sanat K. Sarkar. The multivariate-t distribution and the Simes inequality. Statistics & Probability Letters, 83(1):227–232, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003185.

Bai:2013:MLM

[BT13] Shuyang Bai and Murad S. Taqqu. Multivariate limits of multilinear polynomial-form processes with long memory. Statistics & Probability Letters, 83(11):2473–2485, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002393.

Bai:2014:STM

[BT14a] Shuyang Bai and Murad S. Taqqu. Structure of the third moment of the generalized Rosenblatt distribution. Statistics & Probability Letters, 94(??):144–152, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002508.

Bandyopadhyay:2014:RCL

[BT14b] Antar Bandyopadhyay and Debleena Thacker. Rate of convergence and large deviation for the infinite color Pólya urn schemes. Statistics & Probability Letters, 92(??):232– 240, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214001928.

Burkschat:2014:RHR

[BT14c] Marco Burkschat and Nuria Torrado. On the reversed hazard rate of sequential order statistics. Statistics & Probability Letters, 85(??):106–113, February 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003969.

Bai:2016:SRD

[BT16] Shuyang Bai and Murad S. Taqqu. Short-range dependent processes subordinated to the Gaussian may not be strong mixing. Statistics & Probability Letters, 110(??):198–200, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303059.

Bertin:2011:DPE

[BTT11] Karine Bertin, Soledad Torres, and Ciprian A. Tudor. Drift parameter estimation in fractional diffusions driven by perturbed random walks. Statistics & Probability Letters, 81(2):243–249, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002798.

Boruvka:2016:DDR

[BTT16] Audrey Boruvka, Glen Takahara, and Dongsheng Tu. Datadriven ridge regression for aalen’s additive risk model. Statistics & Probability Letters, 109:189–193, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003788.

Bahamonde:2018:AMF

[BTT18] Natalia Bahamonde, Soledad Torres, and Ciprian A. Tudor. ARCH model and fractional Brownian motion. Statistics & Probability Letters, 134(??):70–78, March 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303176.

Buchsteiner:2015:WCW

[Buc15] Jannis Buchsteiner. Weak convergence of the weighted sequential empirical process of some long-range dependent data. Statistics & Probability Letters, 96(??):170–179, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003423.

Budny:2014:GCI

[Bud14] Katarzyna Budny. A generalization of Chebyshev’s inequality for Hilbert-space-valued random elements. Statistics & Probability Letters, 88(??):62–65, May 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000327.

Bulinskaya:2014:FHT

[Bul14] Ekaterina Vladimirovna Bulinskaya. Finiteness of hitting times under taboo. Statistics & Probability Letters, 85(??):15–19, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003593.

Buonaguidi:2018:DOO

[Buo18] B. Buonaguidi. Dynamic optimality in optimal variance stopping problems. Statistics & Probability Letters, 141(??):103– 108, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302177.

Butov:2015:POI

[But15] Alexander A. Butov. On the problem of optimal instant observations of the linear birth and death process. Statistics & Probability Letters, 101(??):49–53, June 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000784.

Berg:2010:DST

[BV10] C. Berg and C. Vignat. On the density of the sum of two independent Student t-random vectors. Statistics & Probability Letters, 80(13–14):1043–1055, July 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000660.

Bibby:2011:TDB

[BV11] Bo Martin Bibby and Michael Væth. The two-dimensional beta binomial distribution. Statistics & Probability Letters, 81(7): 884–891, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521000369X.

Boente:2015:SCR

[BV15a] Graciela Boente and Alejandra Vahnovan. Strong convergence of robust equivariant nonparametric functional regression estimators. Statistics & Probability Letters, 100(??):1–11, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000346.

Borisov:2015:NEI

[BV15b] I. S. Borisov and N. V. Volodko. A note on exponential inequalities for the distribution tails of canonical von Mises’ statistics of dependent observations. Statistics & Probability Letters, 96(??):287–291, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003575.

Buhlmann:2018:SBD

[BvdG18] Peter Bühlmann and Sara van de Geer. Statistics for big data: a perspective. Statistics & Probability Letters, 136(??):37–41, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300610.

Boudou:2010:RBU

[BVP10] Alain Boudou and Sylvie Viguier-Pla. Relation between unit operators proximity and their associated spectral measures. Statistics & Probability Letters, 80(23–24):1724–1732, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002105.

Bruffaerts:2014:GBS

[BVV14] Christopher Bruffaerts, Vincenzo Verardi, and Catherine Vermandele. A generalized boxplot for skewed and heavy-tailed distributions. Statistics & Probability Letters, 95(??):110–117, December 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002971.

Bo:2011:SIM

[BW11] Lijun Bo and Yongjin Wang. On a stochastic interacting model with stepping-stone noises. Statistics & Probability Letters, 81(8):1300–1305, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001234.

Bai:2015:NLS

[BW15] Zhidong Bai and Chen Wang. A note on the limiting spectral distribution of a symmetrized auto-cross covariance matrix. Statistics & Probability Letters, 96(??):333–340, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003514.

Bryc:2018:DRL

[BW18] Wlodzimierz Bryc and Yizao Wang. Dual representations of Laplace transforms of Brownian excursion and generalized meanders. Statistics & Probability Letters, 140(??):77–83, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301780.

Bissiri:2019:GBI

[BW19] Pier Giovanni Bissiri and Stephen G. Walker. On general Bayesian inference using loss functions. Statistics & Probability Letters, 152(??):89–91, September 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301087.

Bai:2011:MVR

[BWW11] Zhidong Bai, Keyan Wang, and Wing-Keung Wong. The mean-variance ratio test — a complement to the coefficient of variation test and the Sharpe ratio test. Statistics & Probability Letters, 81(8):1078–1085, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000812.

Bo:2012:SML

[BY12] Lijun Bo and Xuewei Yang. Sequential maximum likelihood estimation for reflected generalized Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 82(7):1374–1382, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001113.

Bi:2013:PLD

[BZ13] Xiuchun Bi and Shuguang Zhang. Precise large deviations of aggregate claims in a risk model with regression-type sizedependence. Statistics & Probability Letters, 83(10):2248–2255, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002149.

Brito:2011:HRR

[BZV11] Gerandy Brito, Romulo I. Zequeira, and José E. Valdés. On the hazard rate and reversed hazard rate orderings in twocomponent series systems with active redundancies. Statistics & Probability Letters, 81(2):201–206, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003238.

Cai:2017:NJM

[Cai17] Haiyan Cai. A note on jointly modeling edges and node attributes of a network. Statistics & Probability Letters, 121(??):54–60, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215303771.

Calabrese:2013:UCS

[Cal13] Raffaella Calabrese. Uniform correlation structure and convex stochastic ordering in the Pólya urn scheme. Statistics & Probability Letters, 83(1):272–277, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003495.

Canepa:2016:NBC

[Can16] Alessandra Canepa. A note on Bartlett correction factor for tests on cointegrating relations. Statistics & Probability Letters, 110(??):296–304, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003429.

Cao:2018:SCD

[Cao18] Jiguo Cao. Statisticians can do better in the big data era. Statistics & Probability Letters, 136(??):146–147, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300634.

Carota:2010:TNC

[Car10] Cinzia Carota. Tests for normality in classes of skew-t alternatives. Statistics & Probability Letters, 80(1):1–8, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003460.

Cassese:2010:QML

[Cas10] Gianluca Cassese. Quasi-martingales with a linearly ordered index set. Statistics & Probability Letters, 80(5–6):421–426, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520900443X.

Chen:2010:RLW

[CC10a] Pingyan Chen and Ran Chen. A remark on LSL for weighted sums of i.i.d random elements. Statistics & Probability Letters, 80(17–18):1329–1334, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001227.

Chou:2010:TIT

[CC10b] Cheng Chou and Chia-Shang J. Chu. Testing independence of two autocorrelated binary time series. Statistics & Probability Letters, 80(1):69–75, January 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209003551.

Chretien:2016:LBE

[CC16] Stéphane Chrétien and Franck Corset. A lower bound on the expected optimal value of certain random linear programs and application to shortest paths in Directed Acyclic Graphs and reliability. Statistics & Probability Letters, 117(??):221–230, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300827.

Chang:2018:LTC

[CCFL18] Hsing-Ming Chang, Yung-Ming Chang, Winnie H. W. Fu, and Wan-Chen Lee. On limiting theorems for conditional causation probabilities of multiple-run-rules. Statistics & Probability Letters, 138(??):151–156, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301147.

Contreras-Cristan:2017:APG

[CCGPW17] A. Contreras-Cristán, E. Gutiérrez-Peña, and S. G. Walker. On the asymptotic power of a goodness-of-fit test based on a cumulative Kullback–Leibler discrepancy. Statistics & Probability Letters, 120(??):118–125, January 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301991.

Chakraborty:2019:NEB

[CCM19] Tanujit Chakraborty, Ashis Kumar Chakraborty, and C. A. Murthy. A nonparametric ensemble binary classifier and its statistical properties. Statistics & Probability Letters, 149(??): 16–23, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300276.

Cano:2016:LCO

[CCS16] J. A. Cano, C. Carazo, and D. Salmerón. Linear contrasts for the one way analysis of variance: A Bayesian approach. Statistics & Probability Letters, 109:54–62, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003740.

Chen:2017:FEO

[CCS17] Yu-Ting Chen, Yu-Tzu Chen, and Yuan-Chung Sheu. First exit from an open set for a matrix-exponential Lévy process. Statistics & Probability Letters, 127(??):104–110, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301189.

Crossman:2011:SHR

[CCSC11] Richard J. Crossman, Pauline Coolen-Schrijner, and Frank P. A. Coolen. The shape of the hazard rate for finite continuous-time birth-death processes. Statistics & Probability Letters, 81(2):181–187, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003044.

Cheng:2010:GES

[CD10a] Pei Cheng and Feiqi Deng. Global exponential stability of impulsive stochastic functional differential systems. Statistics & Probability Letters, 80(23–24):1854–1862, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002348.

Cui:2010:CRE

[CD10b] Hongfei Cui and Yiming Ding. The convergence of the Rényi entropy of the normalized sums of IID random variables. Statistics & Probability Letters, 80(15–16):1167–1173, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521000091X.

Chretien:2012:IRS

[CD12] Stéphane Chrétien and Sébastien Darses. Invertibility of random submatrices via tail-decoupling and a matrix Chernoff inequality. Statistics & Probability Letters, 82(7):1479–1487, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001319.

Chakraborty:2018:SVW

[CD18a] Saptarshi Chakraborty and Swagatam Das. Simultaneous variable weighting and determining the number of clusters — a weighted Gaussian means algorithm. Statistics & Probability Letters, 137(??):148–156, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521830018X.

Chen:2018:SBI

[CD18b] Willa W. Chen and Rohit S. Deo. Subsampling based inference for U statistics under thick tails using self-normalization. Statistics & Probability Letters, 138(??):95–103, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301020.

Cardot:2015:EKS

[CDG15] Hervé Cardot, Anne De Moliner, and Camelia Goga. Estimating with kernel smoothers the mean of functional data in a finite population setting. A note on variance estimation in presence of partially observed trajectories. Statistics & Probability Letters, 99(??):156–166, April 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000310.

Chaubey:2011:SES

[CDL11] Yogendra P. Chaubey, Isha Dewan, and Jun Li. Smooth estimation of survival and density functions for a stationary associated process using Poisson weights. Statistics & Probability Letters, 81(2):267–276, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002920.

Chai:2017:TLO

[CDS17] Feng-Shun Chai, Ashish Das, and Rakhi Singh. Threelevel A- and D-optimal paired choice designs. Statistics & Probability Letters, 122(??):211–217, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302711.

Ceri:2018:RSE

[Cer18] Stefano Ceri. On the role of statistics in the era of big data: a computer science perspective. Statistics & Probability Letters, 136(??):68–72, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300646.

Cifarelli:2011:REP

[CF11] D. M. Cifarelli and S. Fortini. Recursive equations for the predictive distributions of some determinantal processes. Statistics & Probability Letters, 81(1):8–15, January 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002622.

Chaudhry:2013:SED

[CF13] Mohan Chaudhry and Brent Fisher. Simple and elegant derivations for some asymptotic results in the discrete-time renewal process. Statistics & Probability Letters, 83(1):315–319, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003677.

Cha:2018:IBR

[CF18] Ji Hwan Cha and Maxim Finkelstein. On informationbased residual lifetime in survival models with delayed failures. Statistics & Probability Letters, 137(??):209–216, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300397.

Charnigo:2013:TUV

[CFBD13] Richard Charnigo, Qian Fan, Douglas Bittel, and Hongying Dai. Testing unilateral versus bilateral normal contamination. Statistics & Probability Letters, 83(1):163–167, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003616.

Cai:2012:RAB

[CFS12] Zongwu Cai, Ying Fang, and Jia Su. Reducing asymptotic bias of weak instrumental estimation using independently repeated cross-sectional information. Statistics & Probability Letters, 82(1):180–185, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003142.

Cerchiello:2012:DFD

[CG12] Paola Cerchiello and Paolo Giudici. On the distribution of functionals of discrete ordinal variables. Statistics & Probability Letters, 82(11):2044–2049, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002684.

Chen:2013:MDP

[CG13] Lei Chen and Fuqing Gao. Moderate deviation principle for Brownian motions on the unit sphere in R^{d}. Statistics & Probability Letters, 83(11):2486–2491, November 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002599.

Chen:2017:RGC

[CG17] Xiaohong Chen and Fuchang Gao. A reverse Gaussian correlation inequality by adding cones. Statistics & Probability Letters, 123(??):84–87, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521630284X.

Castruccio:2018:PSI

[CG18] Stefano Castruccio and Marc G. Genton. Principles for statistical inference on big spatio-temporal data from climate models. Statistics & Probability Letters, 136(??):92–96, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300713.

Crambes:2011:WCS

[CGH11] Christophe Crambes, Ali Gannoun, and Yousri Henchiri. Weak consistency of the Support Vector Machine Quantile Regression approach when covariates are functions. Statistics & Probability Letters, 81(12):1847–1858, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002392.

Chen:2017:CEH

[CGLN17] Yung-Pin Chen, Isaac H. Goldstein, Eve D. Lathrop, and Roger B. Nelsen. Computing an expected hitting time for the 3-urn Ehrenfest model via electric networks. Statistics & Probability Letters, 127(??):42–48, August 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730113X.

Chen:2011:RLL

[CH11a] Pingyan Chen and Chunyan Hao. A remark on the law of the logarithm for weighted sums of random variables with multidimensional indices. Statistics & Probability Letters, 81(12):1808–1812, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002380.

Cirillo:2011:ESM

[CH11b] Pasquale Cirillo and Jürg Hüsler. Extreme shock models: An alternative perspective. Statistics & Probability Letters, 81(1):25– 30, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002646.

Christofides:2012:MMI

[CH12] Tasos C. Christofides and Milto Hadjikyriakou. Maximal and moment inequalities for demimartingales and Ndemimartingales. Statistics & Probability Letters, 82(3):683– 691, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003919.

Chen:2013:ARM

[CH13] Ting-Li Chen and Chii-Ruey Hwang. Accelerating reversible Markov chains. Statistics & Probability Letters, 83(9):1956–1962, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001533.

Chen:2014:RCE

[CH14] Shouquan Chen and Jianwen Huang. Rates of convergence of extreme for asymmetric normal distribution. Statistics & Probability Letters, 84(??):158–168, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003374.

Chen:2017:SVR

[CH17] Yanhong Chen and Yijun Hu. Set-valued risk statistics with scenario analysis. Statistics & Probability Letters, 131(??):25– 37, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521730264X.

Chen:2019:EAM

[CH19] Ouxiang Chen and Taizhong Hu. Extreme-aggregation measures in the RDEU model. Statistics & Probability Letters, 148(??): 155–163, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521930029X.

Chandler:2010:OSH

[Cha10] Gabriel Chandler. Order selection for heteroscedastic autoregression: a study on concentration. Statistics & Probability Letters, 80(23–24):1904–1910, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002415.

Chandra:2015:VPT

[Cha15] Tapas Kumar Chandra. de La Vallée Poussin’s theorem, uniform integrability, tightness and moments. Statistics & Probability Letters, 107:136–141, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002904.

Chakrabarty:2017:HPF

[Cha17] Arijit Chakrabarty. The Hadamard product and the free convolutions. Statistics & Probability Letters, 127(??):150–157, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301487.

Cha:2019:PLP

[Cha19a] Ji Hwan Cha. Poisson Lindley process and its main properties. Statistics & Probability Letters, 152(??):74–81, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301117.

Chakraborty:2019:TMS

[Cha19b] Chitradipa Chakraborty. Testing multivariate scatter parameter in elliptical model based on forward search method. Statistics & Probability Letters, 147(??):66–72, April 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303833.

Chen:2010:CCS

[Che10a] Hua Yun Chen. Compatibility of conditionally specified models. Statistics & Probability Letters, 80(7–8):670–677, April 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004830.

Chen:2010:CNS

[Che10b] Hua Yun Chen. On L^{∞} convergence of Neumann series approximation in missing data problems. Statistics & Probability Letters, 80(9–10):864–873, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000337.

Chen:2010:III

[Che10c] Huabin Chen. Impulsive-integral inequality and exponential stability for stochastic partial differential equations with delays. Statistics & Probability Letters, 80(1):50–56, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003526.

Chen:2010:DFU

[Che10d] Xiaoyan Chen. Dynkin’s formula under the G-expectation. Statistics & Probability Letters, 80(5–6):519–526, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004556.

Chen:2012:TMS

[Che12a] Yuguo Chen. A theory for the multiset sampler. Statistics & Probability Letters, 82(3):473–477, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003695.

Cheng:2012:MDE

[Che12b] Fuxia Cheng. Maximum deviation of error density estimators in censored linear regression. Statistics & Probability Letters, 82(9):1657–1664, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001666.

Chen:2013:CRL

[Che13a] Xinxin Chen. Convergence rate of the limit theorem of a Galton–Watson tree with neutral mutations. Statistics & Probability Letters, 83(2):588–595, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004014.

Chen:2013:ATT

[Che13b] Zhongxue Chen. Association tests through combining p-values for case control genome-wide association studies. Statistics & Probability Letters, 83(8):1854–1862, August 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001429.

Chen:2014:NMI

[Che14a] Xiaohui Chen. A note on moment inequality for quadratic forms. Statistics & Probability Letters, 92(??):83–88, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001801.

Chen:2014:EMM

[Che14b] Zhongxue Chen. Extension of Mood’s median test for survival data. Statistics & Probability Letters, 95(??):77–84, ???? 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002855.

Cheng:2014:DEJ

[Che14c] Dan Cheng. Double extreme on joint sets for Gaussian random fields. Statistics & Probability Letters, 92(??):79–82, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001771.

Cheng:2014:PAG

[Che14d] Li-Juan Cheng. A probabilistic approach for gradient estimates on time-inhomogeneous manifolds. Statistics & Probability Letters, 88(??):174–183, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214000571.

Chen:2016:RCV

[Che16] Xiongzhi Chen. Resolution of a conjecture on variance functions for one-parameter natural exponential families. Statistics & Probability Letters, 118(??):107–109, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300979.

Cheliotis:2018:TRM

[Che18a] Dimitris Cheliotis. Triangular random matrices and biorthogonal ensembles. Statistics & Probability Letters, 134(??):36– 44, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303243.

Chen:2018:CAE

[Che18b] Chao-Ping Chen. Complete asymptotic expansions for the density function of t-distribution. Statistics & Probability Letters, 141(??):1–6, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302013.

Chen:2018:RFV

[Che18c] Xiongzhi Chen. Reduction functions for the variance function of one-parameter natural exponential family. Statistics & Probability Letters, 137(??):319–325, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300555.

Cheng:2018:GCT

[Che18d] Fuxia Cheng. Glivenko–Cantelli Theorem for the kernel error distribution estimator in the first-order autoregressive model. Statistics & Probability Letters, 139(??):95–102, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830138X.

Chen:2019:CRK

[Che19a] Jiangning Chen. Convergence rate of Krasulina estimator. Statistics & Probability Letters, 154(??):Article 108562, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302081.

Chen:2019:SEC

[Che19b] Li-Pang Chen. Semiparametric estimation for cure survival model with left-truncated and right-censored data and covariate measurement error. Statistics & Probability Letters, 154(??):Article 108547, November 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301841.

Chiba:2010:EPS

[Chi10] Yasutaka Chiba. Estimating the principal stratum direct effect when the total effects are consistent between two standard populations. Statistics & Probability Letters, 80(11–12):958–961, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000489.

Chiba:2012:BCA

[Chi12a] Yasutaka Chiba. Bounds on the complier average causal effect in randomized trials with noncompliance. Statistics & Probability Letters, 82(7):1352–1357, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001332.

Chiba:2012:NBC

[Chi12b] Yasutaka Chiba. A note on bounds for the causal infectiousness effect in vaccine trials. Statistics & Probability Letters, 82(7):1422–1429, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001356.

Chiu:2013:STO

[Chi13] Wan-Yi Chiu. A simple test of optimal hedging policy. Statistics & Probability Letters, 83(4):1062–1070, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004737.

Cai:2015:LDM

[CHM15] Yujie Cai, Jianhui Huang, and Vasileios Maroulas. Large deviations of mean-field stochastic differential equations with jumps. Statistics & Probability Letters, 96(??):1–9, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002892.

Crambes:2016:ENC

[CHM16] Christophe Crambes, Nadine Hilgert, and Tito Manrique. Estimation of the noise covariance operator in functional linear regression with functional outputs. Statistics & Probability Letters, 113(??):7–15, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000353.

Chen:2014:IRA

[CHN14] Zhongxue Chen, Hanwen Huang, and Hon Keung Tony Ng. An improved robust association test for GWAS with multiple diseases. Statistics & Probability Letters, 91(??):153–161, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001461.

Chorro:2016:SPA

[Cho16a] Christophe Chorro. A simple probabilistic approach of the yardsale model. Statistics & Probability Letters, 112(??):35–40, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000080.

Chown:2016:EEE

[Cho16b] Justin Chown. Efficient estimation of the error distribution function in heteroskedastic nonparametric regression with missing data. Statistics & Probability Letters, 117(??):31–39, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300281.

Choi:2018:VFB

[Cho18] Michael C. H. Choi. Velocity formulae between entropy and hitting time for Markov chains. Statistics & Probability Letters, 141(??):62–67, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521830213X.

Chojecki:2019:PTN

[Cho19] Tymoteusz Chojecki. Passive tracer in non-markovian, Gaussian velocity field. Statistics & Probability Letters, 145(??):21–27, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302773.

Christensen:2012:GFN

[Chr12] Sören Christensen. Generalized Fibonacci numbers and Blackwell’s renewal theorem. Statistics & Probability Letters, 82 (9):1665–1668, September 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200168X.

Chandra:2016:UNS

[CHR16] Tapas K. Chandra, Tien-Chung Hu, and Andrew Rosalsky. On uniform nonintegrability for a sequence of random variables. Statistics & Probability Letters, 116(??):27–37, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215300882.

Chu:2014:SVP

[Chu14] Weijuan Chu. Small value probabilities for supercritical multitype branching processes with immigration. Statistics & Probability Letters, 93(??):87–95, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002193.

Chung:2018:SCB

[Chu18] Moo K. Chung. Statistical challenges of big brain network data. Statistics & Probability Letters, 136(??):78–82, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300658.

Chen:2018:GSE

[CHW18] Yong Chen, Yaozhong Hu, and Zhi Wang. Gradient and stability estimates of heat kernels for fractional powers of elliptic operator. Statistics & Probability Letters, 142(??):44–49, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302463.

Ciuperca:2011:GCD

[Ciu11] Gabriela Ciuperca. A general criterion to determine the number of change-points. Statistics & Probability Letters, 81(8):1267– 1275, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001106.

Ciuperca:2015:MSH

[Ciu15] Gabriela Ciuperca. Model selection in high-dimensional quantile regression with seamless L_{0} penalty. Statistics & Probability Letters, 107:313–323, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003302.

Chen:2014:RQE

[CJ14] Jinwen Chen and Siqi Jian. A remark on quasi-ergodicity of ultracontractive Markov processes. Statistics & Probability Letters, 87(??):184–190, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214000170.

Choe:2019:PCC

[CJN19] Geon Ho Choe, Hyun Jin Jang, and Young Hoon Na. Pricing contingent convertible bonds: an analytical approach based on two-dimensional stochastic processes. Statistics & Probability Letters, 148(??):43–53, May 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218304036.

Chen:2010:NSA

[CJT10] Kun Chen, Wenxin Jiang, and Martin A. Tanner. A note on some algorithms for the Gibbs posterior. Statistics & Probability Letters, 80(15–16):1234–1241, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001082.

Cohen:2014:GGT

[CJY14] Samuel N. Cohen, Shaolin Ji, and Shuzhen Yang. A generalized Girsanov transformation of finite state stochastic processes in discrete time. Statistics & Probability Letters, 84(??):33– 39, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003234.

Chambers:2013:JEU

[CK13] Marcus J. Chambers and Maria Kyriacou. Jackknife estimation with a unit root. Statistics & Probability Letters, 83(7):1677– 1682, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000990.

Chojecki:2014:PVT

[CK14] Tymoteusz Chojecki and Tomasz Komorowski. On positivity of the variance of a tracer moving in a divergence-free Gaussian random field. Statistics & Probability Letters, 91(??):98– 106, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001412.

Cadena:2016:NRT

[CK16a] Meitner Cadena and Marie Kratz. New results for tails of probability distributions according to their asymptotic decay. Statistics & Probability Letters, 109:178–183, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215300237.

Chebunin:2016:FCL

[CK16b] Mikhail Chebunin and Artyom Kovalevskii. Functional central limit theorems for certain statistics in an infinite urn scheme. Statistics & Probability Letters, 119(??):344–348, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521630164X.

Chasiotis:2017:UBN

[CKF17] Vasilis Chasiotis, Stratis Kounias, and Nikolaos Farmakis. Upper bound on the number of multi-level columns in equally replicated optimal designs minimizing the E(f_{NOD}) criterion. Statistics & Probability Letters, 129(??):269–274, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302122.

Cox:2018:BDS

[CKK18] D. R. Cox, Christiana Kartsonaki, and Ruth H. Keogh. Big data: Some statistical issues. Statistics & Probability Letters, 136(??): 111–115, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300609.

Cenesiz:2017:SSC

[ÇKN17] Yücel Çenesiz, Ali Kurt, and Erkan Nane. Stochastic solutions of conformable fractional Cauchy problems. Statistics & Probability Letters, 124(??):126–131, May 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300354.

Cadena:2019:CGC

[CKO19] Meitner Cadena, Marie Kratz, and Edward Omey. Characterization of a general class of tail probability distributions. Statistics & Probability Letters, 154(??):Article 108553, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301907.

Cadonna:2017:BMM

[CKP17] Annalisa Cadonna, Athanasios Kottas, and Raquel Prado. Bayesian mixture modeling for spectral density estimation. Statistics & Probability Letters, 125(??):189–195, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300573.

Cui:2010:IBM

[CL10] Yunwei Cui and Robert Lund. Inference in binomial AR(1) models. Statistics & Probability Letters, 80(23–24):1985–1990, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002531.

Cordeiro:2011:BLD

[CL11] Gauss M. Cordeiro and Artur J. Lemonte. The beta Laplace distribution. Statistics & Probability Letters, 81(8):973–982, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000320.

Chobanyan:2013:CPT

[CL13] S. Chobanyan and S. Levental. Contraction principle for tail probabilities of sums of exchangeable random vectors with multipliers. Statistics & Probability Letters, 83(7):1720–1724, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000813.

Chung:2015:CEA

[CL15] Yeojin Chung and Bruce G. Lindsay. Convergence of the EM algorithm for continuous mixing distributions. Statistics & Probability Letters, 96(??):190–195, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003411.

Choe:2016:NCH

[CL16] Geon Ho Choe and Dong Min Lee. Numerical computation of hitting time distributions of increasing Lévy processes. Statistics & Probability Letters, 119(??):289–294, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301584.

Chaouch:2019:OAM

[CL19a] Mohamed Chaouch and Naâmane Laïb. Optimal asymptotic MSE of kernel regression estimate for continuous time processes with missing at random response. Statistics & Probability Letters, 154(??):Article 108532, November 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521930166X.

Choi:2019:HIU

[CL19b] Michael C. H. Choi and Evelyn Li. A Hoeffding’s inequality for uniformly ergodic diffusion process. Statistics & Probability Letters, 150(??):23–28, July 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300641.

Clancy:2014:SEM

[Cla14] Damian Clancy. SIR epidemic models with general infectious period distribution. Statistics & Probability Letters, 85(??):1– 5, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521300360X.

Cruz-Lopez:2018:RWP

[CLEA18] Manuel Cruz-López and Samuel Estala-Arias. A random walk on the profinite completion of a finitely generated group. Statistics & Probability Letters, 143(??):7–16, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302608.

Cruz-Lxfpez:2016:RWP

[CLEAMS16] Manuel Cruz-López, Samuel Estala-Arias, and Antonio Murillo-Salas. A random walk on the profinite completion of Z. Statistics & Probability Letters, 109:130–138, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003831.

Cochran:2013:MFP

[CLG13] James J. Cochran, Martin S. Levy, and Saeed Golnabi. Matching the finitized Poisson distribution to the matching distributions. Statistics & Probability Letters, 83(6):1484–1489, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000564.

Chobanyan:2012:ASC

[CLM12] S. Chobanyan, S. Levental, and V. Mandrekar. Almost surely convergent summands of a random sum. Statistics & Probability Letters, 82(1):212–216, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003117.

Chatterjee:2012:SNL

[CLQ12] Kashinath Chatterjee, Zhaohai Li, and Hong Qin. Some new lower bounds to centered and wrap-round L_{2}-discrepancies. Statistics & Probability Letters, 82(7):1367–1373, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000831.

Chen:2017:DCP

[CLWH17] Xue-Ping Chen, Jin-Guan Lin, Hong-Xia Wang, and Xing-Fang Huang. Designs containing partially clear main effects. Statistics & Probability Letters, 121(??):12–17, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301766.

Chen:2015:CME

[CLYW15] Xue-Ping Chen, Jin-Guan Lin, Jian-Feng Yang, and Hong-Xia Wang. Construction of main-effect plans orthogonal through the block factor. Statistics & Probability Letters, 106(??):58–64, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002370.

Chang:2010:HPD

[CM10] In Hong Chang and Rahul Mukerjee. Highest posterior density regions with approximate frequentist validity: The role of data-dependent priors. Statistics & Probability Letters, 80(23–24):1791–1797, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002269.

Chakrabarty:2011:TSL

[CM11a] Arijit Chakrabarty and Mark M. Meerschaert. Tempered stable laws as random walk limits. Statistics & Probability Letters, 81(8):989–997, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000344.

Chen:2011:LMR

[CM11b] Huichao Chen and Amita K. Manatunga. A longitudinal model for repeated interval-observed data with informative dropouts. Statistics & Probability Letters, 81(2):292–297, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002956.

Cohen:2011:SMA

[CM11c] Serge Cohen and Makoto Maejima. Selfdecomposability of moving average fractional Lévy processes. Statistics & Probability Letters, 81(11):1664–1669, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002173.

Chatterjee:2016:IIL

[CM16] Kiranmoy Chatterjee and Diganta Mukherjee. An improved integrated likelihood population size estimation in dual-record system. Statistics & Probability Letters, 110(??):146–154, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215300699.

Chen:2017:NVC

[CM17] Yin-Hsiu Chen and Bhramar Mukherjee. A new variance component score test for testing distributed lag functions with applications in time series analysis. Statistics & Probability Letters, 123(??):122–127, April 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302875.

Chang:2019:NIV

[CMS19] Yuan-Tsung Chang, Takeru Matsuda, and William E. Strawderman. A note on improving on a vector of coordinatewise estimators of non-negative means via shrinkage. Statistics & Probability Letters, 153(??):143–150, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301634.

Chacon:2014:FMS

[CMT14] José E. Chacón, Pablo Monfort, and Carlos Tenreiro. Fourier methods for smooth distribution function estimation. Statistics & Probability Letters, 84(??):223–230, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003532.

Chae:2018:CIA

[CMW18] Minwoo Chae, Ryan Martin, and Stephen G. Walker. Convergence of an iterative algorithm to the nonparametric MLE of a mixing distribution. Statistics & Probability Letters, 140 (??):142–146, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301998.

Chahkandi:2016:TER

[CN16] M. Chahkandi and H. Alizadeh Noughabi. Testing exponentiality of the residual life, based on dynamic cumulative residual entropy. Statistics & Probability Letters, 117(??):1–11, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300451.

Candan:2013:MFR

[CO13]  Çagatay Candan and Umut Orguner. The moment func-

tion for the ratio of correlated generalized gamma variables. Statistics & Probability Letters, 83(10):2353–2356, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002381.

Camponovo:2014:BCE

[CO14] Lorenzo Camponovo and Taisuke Otsu. On Bartlett correctability of empirical likelihood in generalized power divergence family. Statistics & Probability Letters, 86(??):38–43, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213004070.

Coeurjolly:2015:ASB

[Coe15] Jean-François Coeurjolly. Almost sure behavior of functionals of stationary Gibbs point processes. Statistics & Probability Letters, 97(??):241–246, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003836.

Cohen:2010:EOP

[Coh10] Albert Cohen. Examples of optimal prediction in the infinite horizon case. Statistics & Probability Letters, 80(11–12):950–957, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000465.

Coqueret:2015:SSN

[Coq15] Guillaume Coqueret. On the supremum of the spectrally negative stable process with drift. Statistics & Probability Letters, 107:333–340, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003314.

Coufal:2019:CRK

[Cou19] David Coufal. Convergence rates of kernel density estimates in particle filtering. Statistics & Probability Letters, 153(??):164– 170, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301713.

Chang:2011:BLR

[CP11] Christopher C. Chang and Dimitris N. Politis. Bootstrap with larger resample size for root-n consistent density estimation with time series data. Statistics & Probability Letters, 81(6):652– 661, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000460.

Cacoullos:2013:SIE

[CP13a] Theophilos Cacoullos and Nickos Papadatos. Self-inverse and exchangeable random variables. Statistics & Probability Letters, 83(1):9–12, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002660.

Cruz:2013:ENH

[CP13b] J. A. Rojas Cruz and A. G. C. Pereira. The elitist nonhomogeneous genetic algorithm: Almost sure convergence. Statistics & Probability Letters, 83(10):2179–2185, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001934.

Chang:2014:ASD

[CP14] Christopher Chang and Dimitris Politis. Aggregation of spectral density estimators. Statistics & Probability Letters, 94(??): 204–213, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214002557.

Cannas:2017:SMA

[CP17a] Massimo Cannas and Gavino Puggioni. On the support of matching algorithms. Statistics & Probability Letters, 131(??):72–77, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302699.

Coeurjolly:2017:PHE

[CP17b] Jean-François Coeurjolly and Emilio Porcu. Properties and Hurst exponent estimation of the circularly-symmetric fractional Brownian motion. Statistics & Probability Letters, 128(??):21– 27, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301360.

Czarna:2017:PQS

[CP17c] Irmina Czarna and Zbigniew Palmowski. Parisian quasistationary distributions for asymmetric Lévy processes. Statistics & Probability Letters, 127(??):75–84, August 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301025.

Coutin:2019:ERL

[CP19] Laure Coutin and Monique Pontier. Existence and regularity of law density of a pair (diffusion, first component running maximum). Statistics & Probability Letters, 153(??):130–138, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301427.

Choi:2012:AGP

[CPH12] M. S. Choi, J. A. Park, and S. Y. Hwang. Asymmetric GARCH processes featuring both threshold effect and bilinear structure. Statistics & Probability Letters, 82(3):419–426, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003774.

Chi:2015:EOT

[CPY15] Zhiyi Chi, Vladimir Pozdnyakov, and Jun Yan. On expected occupation time of Brownian bridge. Statistics & Probability Letters, 97(??):83–87, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003782.

Chang:2017:EDS

[CQ17] Shuhua Chang and Yongcheng Qi. Empirical distribution of scaled eigenvalues for product of matrices from the spherical ensemble. Statistics & Probability Letters, 128(??):8–13, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301335.

Chang:2018:SMT

[CQ18] Shuhua Chang and Yongcheng Qi. On Schott’s and Mao’s test statistics for independence of normal random vectors. Statistics & Probability Letters, 140(??):132–141, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301901.

Castaing:2012:NFC

[CQT12] Charles Castaing, Nguyen Van Quang, and Nguyen Tran Thuan. A new family of convex weakly compact valued random variables in Banach space and applications to laws of large numbers. Statistics & Probability Letters, 82(1):84–95, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002768.

Chen:2015:CUD

[CQZ15] Wen Chen, Zong-Feng Qi, and Yong-Dao Zhou. Constructing uniform designs under mixture discrepancy. Statistics & Probability Letters, 97(??):76–82, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003769.

Choi:2013:TJD

[CR13] ByoungSeon Choi and JeongHo Roh. On the trivariate joint distribution of Brownian motion and its maximum and minimum. Statistics & Probability Letters, 83(4):1046–1053, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004749.

Czarna:2016:NPR

[CR16] Irmina Czarna and Jean-François Renaud. A note on Parisian ruin with an ultimate bankruptcy level for Lévy insurance risk processes. Statistics & Probability Letters, 113(??):54–61, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000420.

Coeurjolly:2019:CII

[CRB19] Jean-François Coeurjolly and Patricia Reynaud-Bouret. A concentration inequality for inhomogeneous Neyman–Scott point processes. Statistics & Probability Letters, 148(??):30–34, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303912.

Chacon:2010:NUC

[CRC10] José E. Chacón and Alberto Rodríguez-Casal. A note on the universal consistency of the kernel distribution function estimator. Statistics & Probability Letters, 80(17–18):1414–1419, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210001434.

Chauvet:2017:CPS

[CRG17] Guillaume Chauvet and Anne Ruiz-Gazen. A comparison of pivotal sampling and unequal probability sampling with replacement. Statistics & Probability Letters, 121(??):1–5, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302024.

Criens:2018:NMS

[Cri18] David Criens. A note on the monotone stochastic order for processes with independent increments. Statistics & Probability Letters, 135(??):127–131, April 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303462.

Criens:2019:CPI

[Cri19] David Criens. Couplings for processes with independent increments. Statistics & Probability Letters, 146(??):161–167, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303638.

Chen:2014:BKT

[CS14a] Pingyan Chen and Soo Hak Sung. A Baum–Katz theorem for i.i.d. random variables with higher order moments. Statistics & Probability Letters, 94(??):63–68, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002417.

Chen:2014:SCW

[CS14b] Pingyan Chen and Soo Hak Sung. On the strong convergence for weighted sums of negatively associated random variables. Statistics & Probability Letters, 92(??):45–52, September 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001680.

Chang:2015:NSD

[CS15] Yuan-Tsung Chang and William E. Strawderman. A note on stochastic domination for discrete exponential families. Statistics & Probability Letters, 98(??):131–135, March 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004234.

Chan:2016:AUR

[CS16a] Ngai Hang Chan and Tony Sit. Artifactual unit root behavior of Value at Risk (VaR). Statistics & Probability Letters, 116(??):88– 93, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215300262.

Chen:2016:SLLc

[CS16b] Pingyan Chen and Soo Hak Sung. On the strong laws of large numbers for weighted sums of random variables. Statistics & Probability Letters, 118(??):87–93, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301018.

Chen:2016:SLLb

[CS16c] Pingyan Chen and Soo Hak Sung. A strong law of large numbers for nonnegative random variables and applications. Statistics & Probability Letters, 118(??):80–86, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300980.

Chang:2017:SEI

[CS17] Yuan-Tsung Chang and William E. Strawderman. Simultaneous estimation of p positive normal means with common unknown variance. Statistics & Probability Letters, 121(??):83–89, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215301243.

Chenavier:2018:DPR

[CS18] Nicolas Chenavier and Dominique Schneider. On the discrepancy of powers of random variables. Statistics & Probability Letters, 134(??):5–14, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303206.

Chen:2019:STL

[CS19] Pingyan Chen and Soo Hak Sung. A Spitzer-type law of large numbers for widely orthant dependent random variables. Statistics & Probability Letters, 154(??):Article 108544, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301816.

Chen:2011:NEL

[CSG11] Ziqi Chen, Ning-Zhong Shi, and Wei Gao. Nonparametric estimation of the log odds ratio for sparse data by kernel smoothing. Statistics & Probability Letters, 81(12):1802–1807, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002276.

Chauhan:2013:MCC

[CSK13] Rajvir Singh Chauhan, Parminder Singh, and Narinder Kumar. Multiple comparisons with a control in direction-mixed families of hypothesis under heteroscedasticity. Statistics & Probability Letters, 83(12):2679–2687, December 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002915.

Crawford:2016:CBA

[CSL16] Forrest W. Crawford, Timothy C. Stutz, and Kenneth Lange. Coupling bounds for approximating birth-death processes by truncation. Statistics & Probability Letters, 109:30–38, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003648.

Chaubey:2014:TCV

[CSS14] Yogendra P. Chaubey, Debaraj Sen, and Krishna K. Saha. On testing the coefficient of variation in an inverse Gaussian population. Statistics & Probability Letters, 90(??):121–128, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001163.

Chen:2012:ARP

[CSY12] Chao Chen, Liya Sun, and Litan Yan. An approximation to the Rosenblatt process using martingale differences. Statistics & Probability Letters, 82(4):748–757, April 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003889.

Cai:2017:MIA

[CTN17] Xia Cai, Yubin Tian, and Wei Ning. Modified information approach for detecting change points in piecewise linear failure rate function. Statistics & Probability Letters, 125(??):130–140, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300548.

Chen:2010:MCP

[CTW10] Zhanshou Chen, Zheng Tian, and Yuesong Wei. Monitoring change in persistence in linear time series. Statistics & Probability Letters, 80(19–20):1520–1527, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001653.

Cui:2014:NPE

[Cui14] Zhenyu Cui. A new proof of an Engelbert–Schmidt type zero-one law for time-homogeneous diffusions. Statistics & Probability Letters, 89(??):118–123, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000960.

Cornilly:2019:EDR

[CV19] Dries Cornilly and Steven Vanduffel. Equivalent distortion risk measures on moment spaces. Statistics & Probability Letters, 146(??):187–192, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303766.

Caglar:2013:DML

[CVA13] Mine Caglar and Ceren Vardar-Acar. Distribution of maximum loss of fractional Brownian motion with drift. Statistics & Probability Letters, 83(12):2729–2734, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002976.

Chen:2010:UCV

[CvEZ10] Jiahua Chen, Constance van Eeden, and James Zidek. Uncertainty and the conditional variance. Statistics & Probability Letters, 80(23–24):1764–1770, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002154.

Chen:2018:EES

[CvGL18] Guiling Chen, Onno van Gaans, and Sjoerd Verduyn Lunel. Existence and exponential stability of a class of impulsive neutral stochastic partial differential equations with delays and Poisson jumps. Statistics & Probability Letters, 141(??):7–18, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302049.

Cheng:2011:SCA

[CW11] Fuxia Cheng and Miin-Jye Wen. The L_{1} strong consistency of ARCH innovation density estimator. Statistics & Probability Letters, 81(5):548–551, May 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000022.

Chen:2012:ASE

[CW12] Lin Chen and Fuke Wu. Almost sure exponential stability of the θ-method for stochastic differential equations. Statistics & Probability Letters, 82(9):1669–1676, September 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001691.

Cai:2015:CLT

[CW15] Yujie Cai and Shaochen Wang. Central limit theorem and moderate deviation principle for CKLS model with small random perturbation. Statistics & Probability Letters, 98(??):6– 11, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003939.

Cui:2016:CML

[CW16] Yunwei Cui and Rongning Wu. On conditional maximum likelihood estimation for INGARCH(p, q) models. Statistics & Probability Letters, 118(??):1–7, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215302200.

Cheng:2019:CSI

[CW19] Lingyan Cheng and Liming Wu. Centered Sobolev inequality and exponential convergence in Φ-entropy. Statistics & Probability Letters, 148(??):101–111, May 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300082.

Chen:2010:IPB

[CWI10] Xi Chen, Lily Wang, and Hemant Ishwaran. An integrative pathway-based clinical-genomic model for cancer survival prediction. Statistics & Probability Letters, 80(17–18): 1313–1319, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001203.

Chen:2012:RCE

[CWZ12] Shouquan Chen, Chao Wang, and Geng Zhang. Rates of convergence of extreme for general error distribution under power normalization. Statistics & Probability Letters, 82(2):385–395, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003403.

Cao:2018:JAP

[CX18] Jianyu Cao and Weixin Xie. Joint arrival process of multiple independent batch Markovian arrival processes. Statistics & Probability Letters, 133(??):42–49, February 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303061.

Chen:2013:PES

[CXW13] Wangxue Chen, Minyu Xie, and Ming Wu. Parametric estimation for the scale parameter for scale distributions using moving extremes ranked set sampling. Statistics & Probability Letters, 83(9):2060–2066, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001764.

Chen:2015:GCT

[CXZ15] Zhenlong Chen, Lin Xu, and Dongjin Zhu. Generalized continuous time random walks and Hermite processes. Statistics & Probability Letters, 99(??):44–53, April 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521500005X.

Chen:2011:RIL

[CY11] Chao Chen and Litan Yan. Remarks on the intersection local time of fractional Brownian motions. Statistics & Probability Letters, 81(8):1003–1012, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000368.

Chen:2015:OJU

[CY15] Ye Chen and Jun Yu. Optimal jackknife for unit root models. Statistics & Probability Letters, 99(??):135–142, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214004179.

Chen:2016:SLLa

[CYH16] Pingyan Chen, Xiaoqin Ye, and Tien-Chung Hu. A strong law and a law of the single logarithm for arrays of rowwise independent random variables. Statistics & Probability Letters, 110(??):169–174, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215303618.

Chen:2019:SLH

[CYLL19] Hao Chen, Jinyu Yang, Dennis K. J. Lin, and Min-Qian Liu. Sliced Latin hypercube designs with both branching and nested factors. Statistics & Probability Letters, 146(??):124–131, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303547.

Cui:2011:ESN

[CYS11] Jing Cui, Litan Yan, and Xichao Sun. Exponential stability for neutral stochastic partial differential equations with delays and Poisson jumps. Statistics & Probability Letters, 81 (12):1970–1977, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002744.

Chang:2012:THI

[CYW12] Ming Chang, Xuqun You, and Muqing Wen. Testing the homogeneity of inverse Gaussian scale-like parameters. Statistics & Probability Letters, 82(10):1755–1760, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001885.

Chen:2013:SDL

[CYW13] Wei Chen, Changjun Yu, and Yuebao Wang. Some discussions on the local distribution classes. Statistics & Probability Letters, 83(7):1654–1661, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000965.

Chen:2015:CCR

[CZ15] Xin Chen and Li-Ping Zhu. Connecting continuum regression with sufficient dimension reduction. Statistics & Probability Letters, 98(??):44–49, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214004106.

Cheng:2017:REP

[CZ17a] Qing Cheng and Liping Zhu. On relative efficiency of principal Hessian directions. Statistics & Probability Letters, 126(??):108– 113, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521730086X.

Cui:2017:CML

[CZ17b] Yunwei Cui and Qi Zheng. Conditional maximum likelihood estimation for a class of observation-driven time series models for count data. Statistics & Probability Letters, 123(??):193– 201, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216302425.

Chen:2019:SMF

[CZCL19] Xiaolin Chen, Yahui Zhang, Xiaojing Chen, and Yi Liu. A simple model-free survival conditional feature screening. Statistics & Probability Letters, 146(??):156–160, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303729.

Zhou:2012:WEN

[cZgL12] Xing cai Zhou and Jin guan Lin. A wavelet estimator in a nonparametric regression model with repeated measurements under martingale difference error’s structure. Statistics & Probability Letters, 82(11):1914–1922, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002623.

Chen:2018:SCL

[CZLT18] Jiaqi Chen, Yangchun Zhang, Weiming Li, and Boping Tian. A supplement on CLT for LSS under a large dimensional generalized spiked covariance model. Statistics & Probability Letters, 138(??):57–65, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301068.

Zhou:2017:WEV

[cZzXgL17] Xing cai Zhou, Ying zhi Xu, and Jin guan Lin. Wavelet estimation in varying coefficient models for censored dependent data. Statistics & Probability Letters, 122(??):179–189, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302498.

Daly:2019:SST

[Dal19] Fraser Daly. On strong stationary times and approximation of Markov chain hitting times by geometric sums. Statistics & Probability Letters, 150(??):74–80, July 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300690.

Dattner:2013:DSE

[Dat13] I. Dattner. Deconvolution of P (X < Y) with supersmooth error distributions. Statistics & Probability Letters, 83(8):1880– 1887, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001454.

Davie:2012:CPR

[Dav12a] George Davie. Constraints placed on random sequences by their compressibility. Statistics & Probability Letters, 82(7):1474– 1478, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001526.

Davydov:2012:CHD

[Dav12b] Yu. Davydov. On convex hull of d-dimensional fractional Brownian motion. Statistics & Probability Letters, 82(1):37–39, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002914.

Davie:2013:DBC

[Dav13] George Davie. Decidable lim sup and Borel–Cantelli-like lemmas for random sequences. Statistics & Probability Letters, 83(1):278–285, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003471.

Davydov:2017:DTV

[Dav17] Youri Davydov. On distance in total variation between image measures. Statistics & Probability Letters, 129(??):393–400, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217302262.

denBoer:2013:DAD

[dB13] A. V. den Boer. Does adding data always improve linear regression estimates? Statistics & Probability Letters, 83(3):829– 835, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004506.

Dai:2013:MLE

[DBB13] Hongsheng Dai, Yanchun Bao, and Mingtang Bao. Maximum likelihood estimate for the dispersion parameter of the negative binomial distribution. Statistics & Probability Letters, 83(1):21– 27, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003227.

deCastro:2013:RLD

[dC13] Yohann de Castro. A remark on the lasso and the Dantzig selector. Statistics & Probability Letters, 83(1):304–314, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003665.

delCastillo:2016:SDS

[dC16] J. M. del Castillo. Slash distributions of the sum of independent logistic random variables. Statistics & Probability Letters, 110(??):111–118, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215302108.

Durante:2019:NEM

[DCR19] Daniele Durante, Antonio Canale, and Tommaso Rigon. A nested expectation-maximization algorithm for latent class models with covariates. Statistics & Probability Letters, 146(??):97– 103, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303390.

Das:2010:BIG

[DD10] Sourish Das and Dipak K. Dey. On Bayesian inference for generalized multivariate gamma distribution. Statistics & Probability Letters, 80(19–20):1492–1499, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001616.

DeCapitani:2011:SOW

[DD11] L. De Capitani and D. De Martini. On stochastic orderings of the Wilcoxon Rank Sum test statistic — with applications to reproducibility probability estimation testing. Statistics & Probability Letters, 81(8):937–946, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001325.

Davydov:2012:CLS

[DD12] Youri Davydov and Clément Dombry. On the convergence of LePage series in Skorokhod space. Statistics & Probability Letters, 82(1):145–150, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002999.

Durante:2014:BDF

[DD14] Daniele Durante and David B. Dunson. Bayesian dynamic financial networks with time-varying predictors. Statistics & Probability Letters, 93(??):19–26, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400217X.

Dieye:2017:ESM

[DDE17] Moustapha Dieye, Mamadou Abdoul Diop, and Khalil Ezzinbi. On exponential stability of mild solutions for some stochastic partial integrodifferential equations. Statistics & Probability Letters, 123(??):61–76, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302383.

Dawid:2011:IAL

[DdRS^{+}11] A. Philip Dawid, Steven de Rooij, Glenn Shafer, Alexander Shen, Nikolai Vereshchagin, and Vladimir Vovk. Insuring against loss of evidence in game-theoretic probability. Statistics & Probability Letters, 81(1):157–162, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002968.

Ducinskas:2015:MCS

[DDZ15] Kestutis Ducinskas, Lina Dreiziene, and Egle Zikariene. Multiclass classification of the scalar Gaussian random field observation with known spatial correlation function. Statistics & Probability Letters, 98(??):107–114, March 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004118.

DeScheemaekere:2011:CCT

[De 11] Xavier De Scheemaekere. A converse comparison theorem for backward stochastic differential equations with jumps. Statistics & Probability Letters, 81(2):298–301, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003020.

DeBin:2016:EBC

[De 16a] Riccardo De Bin. On the equivalence between conditional and random-effects likelihoods in exponential families. Statistics & Probability Letters, 110(??):34–38, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003909.

DeGregorio:2016:TPR

[De 16b] Alessandro De Gregorio. Transport processes with random jump rate. Statistics & Probability Letters, 118(??):127–134, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301031.

DeGregorio:2017:NIR

[De 17] Alessandro De Gregorio. A note on isotropic random flights moving in mixed Poisson environments. Statistics & Probability Letters, 129(??):311–317, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302250.

Dehay:2013:NEP

[DE13a] D. Dehay and K. El Waled. Nonparametric estimation problem for a time-periodic signal in a periodic noise. Statistics & Probability Letters, 83(2):608–615, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004245.

Devarajan:2013:PLE

[DE13b] Karthik Devarajan and Nader Ebrahimi. On penalized likelihood estimation for a non-proportional hazards regression model. Statistics & Probability Letters, 83(7):1703–1710, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000801.

Debowski:2012:BRU

[Deb12] Lukasz Debowski. On bounded redundancy of universal codes. Statistics & Probability Letters, 82(11):2068–2071, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002799.

Dembinska:2010:NON

[Dem10] Anna Dembińska. On numbers of observations near randomly indexed order statistics. Statistics & Probability Letters, 80 (5–6):309–317, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209004222.

Diaz-Emparanza:2014:NDF

[DEM14] Ignacio Díaz-Emparanza and M. Paz Moral. Numerical distribution functions for seasonal stability tests. Statistics & Probability Letters, 86(??):44–49, March 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300401X.

Dendievel:2015:WOS

[Den15] Rémi Dendievel. Weber’s optimal stopping problem and generalizations. Statistics & Probability Letters, 97(??):176–184, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521400371X.

Deng:2017:BNC

[Den17] Pingjin Deng. Boundary non-crossing probabilities for Slepian process. Statistics & Probability Letters, 122(??):28–35, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302255.

Dueck:2015:GIA

[DER15] Johannes Dueck, Dominic Edelmann, and Donald Richards. A generalization of an integral arising in the theory of distance correlation. Statistics & Probability Letters, 97(??):116–119, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003812.

Devroye:2012:NGR

[Dev12] Luc Devroye. A note on generating random variables with logconcave densities. Statistics & Probability Letters, 82(5):1035– 1039, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000326.

Devroye:2010:SDD

[DF10] Luc Devroye and Omar Fawzi. Simulating the Dickman distribution. Statistics & Probability Letters, 80(3–4):242–247, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209004052.

Duncan:2011:CPS

[DF11] Tyrone E. Duncan and Holger Fink. Corrigendum to “Prediction for some processes related to a fractional Brownian motion”. Statistics & Probability Letters, 81(8):1336–1337, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001192. See [Dun06].

Denker:2014:RLL

[DF14] Manfred Denker and Souha Fares. Richter’s local limit theorem and Black–Scholes type formulas. Statistics & Probability Letters, 92(??):241–248, September 1–262, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002053.

daFonseca:2019:TVC

[dFAM19] Eder Lucio da Fonseca, Airlane Pereira Alencar, and Pedro Alberto Morettin. Time-varying cointegration model using wavelets. Statistics & Probability Letters, 145(??):260–267, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521830316X.

Davies:2018:UAB

[DFH18] Tilman M. Davies, Claire R. Flynn, and Martin L. Hazelton. On the utility of asymptotic bandwidth selectors for spatially adaptive kernel density estimation. Statistics & Probability Letters, 138(??):75–81, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301123.

Devroye:2012:CED

[DFKK12] Luc Devroye, Tina Felber, Michael Kohler, and Adam Krzyzak. L_{1}-consistent estimation of the density of residuals in random design regression models. Statistics & Probability Letters, 82(1):173–179, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003208.

Doukhan:2012:WDCb

[DFL12] Paul Doukhan, Konstantinos Fokianos, and Xiaoyin Li. On weak dependence conditions: The case of discrete valued processes. Statistics & Probability Letters, 82(11):1941–1948, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002544. See corrigendum [DFL13].

Doukhan:2013:CWDa

[DFL13] Paul Doukhan, Konstantinos Fokianos, and Xiaoyin Li. Corrigendum to “On weak dependence conditions: The case of discrete valued processes” [Statist. Probab. Lett. 82 (2012) 1941– 1948]. Statistics & Probability Letters, 83(2):674–675, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003902. See [DFL12].

Durante:2010:MSA

[DFS10] Fabrizio Durante and Juan Fernández-Sánchez. Multivariate shuffles and approximation of copulas. Statistics & Probability Letters, 80(23–24):1827–1834, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002312.

Dehling:2013:EVP

[DFS^{+}13] Herold Dehling, Roland Fried, Olimjon Sh. Sharipov, Daniel Vogel, and Max Wornowizki. Estimation of the variance of partial sums of dependent processes. Statistics & Probability Letters, 83(1):141–147, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003173.

Dietz:2016:OST

[DFS16] Markus Dietz, Sebastian Fuchs, and Klaus D. Schmidt. On order statistics and their copulas. Statistics & Probability Letters, 117(??):165–172, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300712.

Doukhan:2012:WDCa

[DFT12] Paul Doukhan, Konstantinos Fokianos, and Dag Tjøstheim. On weak dependence conditions for Poisson autoregressions. Statistics & Probability Letters, 82(5):942–948, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000259. See correction [DFT13].

Doukhan:2013:CWDb

[DFT13] Paul Doukhan, Konstantinos Fokianos, and Dag Tjøstheim. Correction to “On weak dependence conditions for Poisson autoregressions” [Statist. Probab. Lett. 82 (2012) 942– 948]. Statistics & Probability Letters, 83(8):1926–1927, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521300134X. See [DFT12].

Dutta:2013:PUC

[DG13] Santanu Dutta and Alok Goswami. Pointwise and uniform convergence of kernel density estimators using random bandwidths. Statistics & Probability Letters, 83(12):2711–2720, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003088.

deHaan:2015:CHE

[dH15] Laurens de Haan. Convergence of heteroscedastic extremes. Statistics & Probability Letters, 101(??):38–39, June 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000693.

Dryden:2018:JBD

[DH18] Ian L. Dryden and David J. Hodge. Journeys in big data statistics. Statistics & Probability Letters, 136(??):121–125, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300580.

Das:2019:EPZ

[DH19] Sayan Das and Rajat Subhra Hazra. Extremal process of the zero-average Gaussian free field for d ≥ 3. Statistics & Probability Letters, 146(??):42–49, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303444.

Debicki:2014:PCS

[DHJT14] Krzysztof Dȩbicki, Enkelejd Hashorva, Lanpeng Ji, and Kamil Tabiś. On the probability of conjunctions of stationary Gaussian processes. Statistics & Probability Letters, 88(??):141–148, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000534.

Donald:2014:IPW

[DHL14] Stephen G. Donald, Yu-Chin Hsu, and Robert P. Lieli. Inverse probability weighted estimation of local average treatment effects: A higher order MSE expansion. Statistics & Probability Letters, 95(??):132–138, December 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400296X.

Dominicy:2013:SMQ

[DHOV13] Yves Dominicy, Siegfried Hörmann, Hiroaki Ogata, and David Veredas. On sample marginal quantiles for stationary processes. Statistics & Probability Letters, 83(1):28–36, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521200288X.

DiCecco:2011:GAC

[Di 11] Davide Di Cecco. A geometric approach to a class of optimization problems concerning exchangeable binary variables. Statistics & Probability Letters, 81(3):411–416, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003354.

Ding:2010:ICC

[Din10] A. Adam Ding. Identifiability conditions for covariate effects model on survival times under informative censoring. Statistics & Probability Letters, 80(11–12):911–915, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000398.

Ding:2014:WDT

[Din14] Ying Ding. Wasserstein-Divergence transportation inequalities and polynomial concentration inequalities. Statistics & Probability Letters, 94(??):77–85, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400251X.

Ding:2015:SSP

[Din15a] Xue Ding. On some spectral properties of large block Laplacian random matrices. Statistics & Probability Letters, 99(??):61– 69, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000115.

Ding:2015:NQT

[Din15b] Ying Ding. A note on quadratic transportation and divergence inequality. Statistics & Probability Letters, 100(??):115– 123, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000590.

Daly:2013:BPC

[DJ13] Fraser Daly and Oliver Johnson. Bounds on the Poincaré constant under negative dependence. Statistics & Probability Letters, 83(2):511–518, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004087.

Dombry:2014:LFT

[DJ14] Clément Dombry and Paul Jung. A Lindeberg–Feller theorem for stable laws. Statistics & Probability Letters, 84(??):198– 203, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003520.

Durante:2011:IDS

[DJM11] Fabrizio Durante, Piotr Jaworski, and Radko Mesiar. Invariant dependence structures and Archimedean copulas. Statistics & Probability Letters, 81(12):1995–2003, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002823.

Doney:2011:LAT

[DK11] R. A. Doney and D. A. Korshunov. Local asymptotics for the time of first return to the origin of transient random walk. Statistics & Probability Letters, 81(9):1419–1424, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100157X.

Dewan:2014:SCR

[DK14] Isha Dewan and Baha-Eldin Khaledi. On stochastic comparisons of residual life time at random time. Statistics & Probability Letters, 88(??):73–79, May 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000467.

Demirhan:2015:GMG

[DK15] Haydar Demirhan and Zeynep Kalaylioglu. On the generalized multivariate Gumbel distribution. Statistics & Probability Letters, 103(??):93–99, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001352.

Diaye:2019:LER

[DKM19] Marc-Arthur Diaye, Gleb A. Koshevoy, and Ilya Molchanov. Lift expectations of random sets. Statistics & Probability Letters, 145(??):110–117, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302906.

Dabye:2019:ATP

[DKT19] A. S. Dabye, Yu. A. Kutoyants, and E. D. Tanguep. On APF test for Poisson process with shift and scale parameters. Statistics & Probability Letters, 145(??):28–36, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302761.

Dai:2010:WLT

[DL10] Hongshuai Dai and Yuqiang Li. A weak limit theorem for generalized multifractional Brownian motion. Statistics & Probability Letters, 80(5–6):348–356, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004337.

Diallo:2013:MLD

[DL13a] Amadou Oury Korbe Diallo and Djamal Louani. Moderate and large deviation principles for the hazard rate function kernel estimator under censoring. Statistics & Probability Letters, 83(3):735–743, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004269.

Didi:2013:CRK

[DL13b] Sultana Didi and Djamal Louani. Consistency results for the kernel density estimate on continuous time stationary and dependent data. Statistics & Probability Letters, 83(4):1262– 1270, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000254.

Dudek:2017:SNT

[DL17] Anna E. Dudek and Lukasz Lenart. Subsampling for nonstationary time series with non-zero mean function. Statistics & Probability Letters, 129(??):252–259, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302067.

delaCruz:2019:OBC

[dlCJB19] H. de la Cruz, J. C. Jimenez, and R. J. Biscay. On the oscillatory behavior of coupled stochastic harmonic oscillators driven by random forces. Statistics & Probability Letters, 146(??):85– 89, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521830347X.

Du:2015:SIH

[DLM15] Jing Du, Liangzhen Lei, and Yutao Ma. Sobolev inequalities for harmonic measures on spheres. Statistics & Probability Letters, 100(??):104–114, May 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000589.

DOvidio:2016:FSR

[DLO16] Mirko D’Ovidio, Nikolai Leonenko, and Enzo Orsingher. Fractional spherical random fields. Statistics & Probability Letters, 116(??):146–156, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521630030X.

Ding:2018:SFP

[DLZZ18] Hui Ding, Zhiping Lu, Jian Zhang, and Riquan Zhang. Semifunctional partial linear quantile regression. Statistics & Probability Letters, 142(??):92–101, November 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302505.

Denuit:2011:DEC

[DM11] Michel M. Denuit and Mhamed Mesfioui. The dispersive effect of cross-aging with Archimedean copulas. Statistics & Probability Letters, 81(9):1407–1418, September 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001507.

DeGregorio:2012:LDP

[DM12a] Alessandro De Gregorio and Claudio Macci. Large deviation principles for telegraph processes. Statistics & Probability Letters, 82(11):1874–1882, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200257X.

Dey:2012:EFN

[DM12b] Aloke Dey and Rahul Mukerjee. Efficiency factors for natural contrasts in partially confounded factorial designs. Statistics & Probability Letters, 82(12):2180–2188, December 2012.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002982.

Dyckerhoff:2012:WMR

[DM12c] Rainer Dyckerhoff and Karl Mosler. Weighted-mean regions of a probability distribution. Statistics & Probability Letters, 82(2):318–325, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003324.

Denuit:2013:SCC

[DM13] Michel Denuit and Mhamed Mesfioui. A sufficient condition of crossing type for the bivariate orthant convex order. Statistics & Probability Letters, 83(1):157–162, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002866.

Dey:2014:CNH

[DM14] Agnish Dey and Arunava Mukherjea. Collapsing of nonhomogeneous Markov chains. Statistics & Probability Letters, 84(??):140–148, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003362.

Denuit:2017:BKT

[DM17a] Michel M. Denuit and Mhamed Mesfioui. Bounds on Kendall’s tau for zero-inflated continuous variables. Statistics & Probability Letters, 126(??):173–178, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300962.

Dominguez-Molina:2017:TWD

[DM17b] J. Armando Domínguez-Molina. The Tracy–Widom distribution is not infinitely divisible. Statistics & Probability Letters, 123(??):56–60, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302826.

Dawid:2017:NBM

[DMC17] A. Philip Dawid, Monica Musio, and Silvia Columbu. A note on Bayesian model selection for discrete data using proper scoring rules. Statistics & Probability Letters, 129(??):101–106, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301852.

deMicheaux:2012:LSL

[dML12] Pierre Lafaye de Micheaux and Christian Léger. A law of the single logarithm for weighted sums of arrays applied to bootstrap model selection in regression. Statistics & Probability Letters, 82(5):965–971, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000284.

deMicheaux:2018:ULL

[dMO18] Pierre Lafaye de Micheaux and Frédéric Ouimet. A uniform L^{1} law of large numbers for functions of i.i.d. random variables that are translated by a consistent estimator. Statistics & Probability Letters, 142(??):109–117, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302293.

Djouadi:2017:CAP

[DMPX17] Seddik M. Djouadi, Vasileios Maroulas, Xiaoyang Pan, and Jie Xiong. Consistency and asymptotics of a Poisson intensity least-squares estimator for partially observed jump-diffusion processes. Statistics & Probability Letters, 123(??):8–16, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630270X.

DAmico:2013:SMM

[DMS13] Guglielmo D’Amico, Raimondo Manca, and Giovanni Salvi. A semi-Markov modulated interest rate model. Statistics & Probability Letters, 83(9):2094–2102, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001855.

Duffy:2019:VAD

[DMS19] Ken R. Duffy, Gianfelice Meli, and Seva Shneer. The variance of the average depth of a pure birth process converges to 7. Statistics & Probability Letters, 150(??):88–93, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300677.

Debicki:2013:TAL

[DMST13] Krzysztof Dȩbicki, Michel Mandjes, and Iwona Sierpińska-Tulacz. Transient analysis of Lévy-driven tandem queues. Statistics & Probability Letters, 83(7):1776–1781, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001119.

Du:2015:PDI

[DMWX15] Yali Du, Junjie Miao, Dongsheng Wu, and Yimin Xiao. Packing dimensions of the images of Gaussian random fields. Statistics & Probability Letters, 106(??):209–217, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002618.

Damasio:2014:CRM

[DN14a] Bruno Damásio and João Nicolau. Combining a regression model with a multivariate Markov chain in a forecasting problem. Statistics & Probability Letters, 90(??):108–113, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001199.

Dehbozorgi:2014:NBI

[DN14b] N. Dehbozorgi and A. R. Nematollahi. A note on the Bayesian inference for generalized multivariate gamma distribution. Statistics & Probability Letters, 92(??):95–98, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001795.

Durio:2016:LEI

[DN16] A. Durio and Ya. Yu. Nikitin. Local efficiency of integrated goodness-of-fit tests under skew alternatives. Statistics & Probability Letters, 117(??):136–143, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300682.

Dabo-Niang:2012:SCM

[DNKL12] Sophie Dabo-Niang, Zoulikha Kaid, and Ali Laksaci. On spatial conditional mode estimation for a functional regressor. Statistics & Probability Letters, 82(7):1413–1421, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001228.

Demongeot:2017:LLR

[DNLR17] Jacques Demongeot, Amina Naceri, Ali Laksaci, and Mustapha Rachdi. Local linear regression modelization when all variables are curves. Statistics & Probability Letters, 121(??):37– 44, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301845.

Dabo-Niang:2010:RQE

[DNT10] Sophie Dabo-Niang and Baba Thiam. Robust quantile estimation and prediction for spatial processes. Statistics & Probability Letters, 80(17–18):1447–1458, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001550.

Diop:2011:LSD

[DO11a] Mamadou Abdoul Diop and Youssef Ouknine. A linear stochastic differential equation driven by a fractional Brownian motion with Hurst parameter > 1/2. Statistics & Probability Letters, 81(8):1013–1020, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000599.

DOvidio:2011:FCH

[D’O11b] Mirko D’Ovidio. On the fractional counterpart of the higherorder equations. Statistics & Probability Letters, 81(12): 1929–1939, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002689.

Doerr:2018:EAP

[Doe18] Benjamin Doerr. An elementary analysis of the probability that a binomial random variable exceeds its expectation. Statistics & Probability Letters, 139(??):67–74, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301366.

Dohler:2014:SCC

[Döh14] Sebastian Döhler. A sufficient criterion for control of some generalized error rates in multiple testing. Statistics & Probability Letters, 92(??):114–120, September 1–262, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001850.

Domansky:2013:SRB

[Dom13] Victor Domansky. Symmetric representations of bivariate distributions. Statistics & Probability Letters, 83(4):1054–1061, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004853.

DOvidio:2014:FTT

[DOT14] Mirko D’Ovidio, Enzo Orsingher, and Bruno Toaldo. Fractional telegraph-type equations and hyperbolic Brownian motion. Statistics & Probability Letters, 89(??):131–137, June 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000856.

Dentcheva:2010:SRI

[DP10] Darinka Dentcheva and Spiridon Penev. Shape-restricted inference for Lorenz curves using duality theory. Statistics & Probability Letters, 80(5–6):403–412, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004416.

Doumas:2016:RVC

[DP16] Aristides V. Doumas and Vassilis G. Papanicolaou. A randomized version of the Collatz 3x + 1 problem. Statistics & Probability Letters, 109:39–44, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300924.

Davydov:2019:RRC

[DP19a] Youri Davydov and Vygantas Paulauskas. Remarks on random countable stable zonotopes. Statistics & Probability Letters, 153(??):187–191, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301798.

Doumas:2019:UVZ

[DP19b] Aristides V. Doumas and Vassilis G. Papanicolaou. Uniform versus Zipf distribution in a mixing collection process. Statistics & Probability Letters, 154(??):Article 108559, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302056.

Dudek:2016:GST

[DPP16] Anna E. Dudek, Efstathios Paparoditis, and Dimitris N. Politis. Generalized seasonal tapered block bootstrap. Statistics & Probability Letters, 115(??):27–35, August 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215304326.

Dette:2016:ODR

[DPZ16] Holger Dette, Andrey Pepelyshev, and Anatoly Zhigljavsky. Optimal designs for regression models with autoregressive errors. Statistics & Probability Letters, 116(??):107–115, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630027X.

Dewan:2002:CLT

[DR02] Isha Dewan and B. L. S. Prakasa Rao. Central limit theorem for U-statistics of associated random variables. Statistics & Probability Letters, 57(1):9–15, March 1, 2002. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715201001948. See corrigendum [DR15].

Delattre:2011:FDP

[DR11] S. Delattre and E. Roquain. On the false discovery proportion convergence under Gaussian equi-correlation. Statistics & Probability Letters, 81(1):111–115, January 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002750.

Dewan:2015:CT

[DR15] Isha Dewan and B. L. S. Prakasa Rao. Corrigendum to “Central limit theorem for U-statistics of associated random variables [Statist. Probab. Lett. 57 (1) (2002) 9–15]. Statistics & Probability Letters, 106(??):147–148, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002412. See [DR02].

Dreassi:2017:NCC

[DR17] Emanuela Dreassi and Pietro Rigo. A note on compatibility of conditional autoregressive models. Statistics & Probability Letters, 125(??):9–16, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300317.

Dick:2019:WDB

[DRZ19] Josef Dick, Daniel Rudolf, and Houying Zhu. A weighted discrepancy bound of quasi-Monte Carlo importance sampling. Statistics & Probability Letters, 149(??):100–106, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300203.

Deng:2011:EPL

[DS11] Chang-Song Deng and Yan-Hong Song. An elementary proof of the L^{1} log-Sobolev inequality for Poisson point processes. Statistics & Probability Letters, 81(9):1458–1462, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100160X.

Durante:2012:ACS

[DS12] Fabrizio Durante and Juan Fernández Sánchez. On the approximation of copulas via shuffles of Min. Statistics & Probability Letters, 82(10):1761–1767, October 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002192.

Djellout:2014:LMD

[DS14] Hacène Djellout and Yacouba Samoura. Large and moderate deviations of realized covolatility. Statistics & Probability Letters, 86(??):30–37, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213004021.

Dutta:2015:DOB

[DS15] Ganesh Dutta and Rita SahaRay. D- and E-optimal blocked main effects plans with unequal block sizes when n is odd. Statistics & Probability Letters, 107:37–43, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002801.

Danilenko:2016:RSS

[DS16] Svetlana Danilenko and Jonas Siaulys. Randomly stopped sums of not identically distributed heavy tailed random variables. Statistics & Probability Letters, 113(??):84–93, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302248.

daSilva:2018:SLL

[dS18a] João Lita da Silva. Strong laws of large numbers for pairwise quadrant dependent random variables. Statistics & Probability Letters, 137(??):349–358, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300427.

Du:2018:SRB

[DS18b] Xiaojie Du and Anton Schick. Signed rank based empirical likelihood for the symmetric location model. Statistics & Probability Letters, 137(??):40–45, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300063.

daSilva:2012:ACM

[dSF12] Telles Timóteo da Silva and Marcelo Dutra Fragoso. Absolutely continuous measure for a jump-type Fleming–Viot process. Statistics & Probability Letters, 82(3):557–564, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003786.

deSaporta:2012:ATB

[dSGPM12] Benoîte de Saporta, Anne Gégout-Petit, and Laurence Marsalle. Asymmetry tests for bifurcating auto-regressive processes with missing data. Statistics & Probability Letters, 82(7):1439– 1444, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001368.

Danilenko:2018:CPE

[DSS18] Svetlana Danilenko, Jonas Siaulys, and Gediminas Stepanauskas. Closure properties of O-exponential distributions. Statistics & Probability Letters, 140(??):63–70, September 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301664.

Dikta:2013:BBM

[DSW13] Gerhard Dikta, Sundarraman Subramanian, and Thorsten Winkler. Bootstrap based model checks with missing binary response data. Statistics & Probability Letters, 83(1):219–226, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003604.

Du:2013:EPL

[DSX13] Jiang Du, Zhimeng Sun, and Tianfa Xie. M-estimation for the partially linear regression model under monotonic constraints. Statistics & Probability Letters, 83(5):1353–1363, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000072.

Dung:2010:SLL

[DT10] Le Van Dung and Nguyen Duy Tien. Strong laws of large numbers for random fields in martingale type p Banach spaces. Statistics & Probability Letters, 80(9–10):756–763, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521000012X.

Durre:2016:ESS

[DTV16] Alexander Dürre, David E. Tyler, and Daniel Vogel. On the eigenvalues of the spatial sign covariance matrix in more than two dimensions. Statistics & Probability Letters, 111(??):80– 85, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215304065.

Duchesne:2004:MMD

[Duc04] Pierre Duchesne. On matricial measures of dependence in vector ARCH models with applications to diagnostic checking. Statistics & Probability Letters, 68(2):149–160, June 15, 2004. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715204000768. See corrigendum [Duc10].

Duchesne:2010:CMM

[Duc10] Pierre Duchesne. Corrigendum to: “On matricial measures of dependence in vector ARCH models with applications to diagnostic checking” [Statist. Probab. Lett. 68 (2004) 149– 160]. Statistics & Probability Letters, 80(9–10):910, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000477. See [Duc04].

Duncan:2006:PSP

[Dun06] T. E. Duncan. Prediction for some processes related to a fractional Brownian motion. Statistics & Probability Letters, 76(2):128–134, January 15, 2006. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715205002506.

See [DF11].

Dung:2016:TPS

[Dun16a]  Nguyen Tien Dung.  Tail probabilities of solutions to

a generalized ait-sahalia interest rate model. Statistics & Probability Letters, 112(??):98–104, May 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301607.

Dung:2016:TPE

[Dun16b] Nguyen Tien Dung. Tail probability estimates for additive functionals. Statistics & Probability Letters, 119(??):349–356, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301675.

Dunson:2018:SBD

[Dun18] David B. Dunson. Statistics in the big data era: Failures of the machine. Statistics & Probability Letters, 136(??):4–9, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300737.

Dung:2019:PFT

[Dun19] Nguyen Tien Dung. The probability of finite-time blowup of a semi-linear SPDE with fractional noise. Statistics & Probability Letters, 149(??):86–92, June 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300367.

Duong:2015:SSM

[Duo15] Tarn Duong. Spherically symmetric multivariate beta family kernels. Statistics & Probability Letters, 104(??):141–145, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001637.

Durieu:2013:EPI

[Dur13] Olivier Durieu. Empirical processes of iterated maps that contract on average. Statistics & Probability Letters, 83 (11):2454–2458, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002526.

Durante:2017:NMG

[Dur17] Daniele Durante. A note on the multiplicative gamma process. Statistics & Probability Letters, 122(??):198–204, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630253X.

Datta:2011:SBM

[DvH11] Somnath Datta and Hans C. van Houwelingen. Statistics in biological and medical sciences. Statistics & Probability Letters, 81(7):715–716, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000988.

Davidson:2018:CTP

[DW18] Allison Davidson and Mark D. Ward. The characterization of tenable Pólya urns. Statistics & Probability Letters, 135(??): 38–43, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521730367X.

deWet:2012:AUE

[dWGM12] Tertius de Wet, Yuri Goegebeur, and Maria Reimert Munch. Asymptotically unbiased estimation of the second order tail parameter. Statistics & Probability Letters, 82(3):565–573, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003701.

Ding:2017:TRM

[DWW17] Hao Ding, Zhanfeng Wang, and Yaohua Wu. Tobit regression model with parameters of increasing dimensions. Statistics & Probability Letters, 120(??):1–7, January 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301717.

Deng:2018:FSQ

[DWWG18] C. Deng, R. P. Waagepetersen, M. Wang, and Y. Guan. A fast spectral quasi-likelihood approach for spatial point processes. Statistics & Probability Letters, 133(??):59–64, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303103.

Diop:2010:MDP

[DY10] Aliou Diop and Armel Fabrice Yode. Minimum distance parameter estimation for Ornstein–Uhlenbeck processes driven by Lévy process. Statistics & Probability Letters, 80(2):122–127, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003678.

Dong:2011:SLL

[DYB11] Yan Dong, Weiguo Yang, and Jianfang Bai. The Strong Law of Large Numbers and the Shannon–McMillan theorem for nonhomogeneous Markov chains indexed by a Cayley tree. Statistics & Probability Letters, 81(12):1883–1890, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002318.

Dong:2014:UCR

[DYW14] Yinghui Dong, Kam C. Yuen, and Chongfeng Wu. Unilateral counterparty risk valuation of CDS using a regime-switching intensity model. Statistics & Probability Letters, 85(??):25–35, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003623.

Ding:2011:NKM

[DZ11] Ying Ding and Xinsheng Zhang. A new kind of modified transportation cost inequalities and polynomial concentration inequalities. Statistics & Probability Letters, 81(10):1524–1534, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001775.

Ding:2017:NSI

[DZ17] Ying Ding and Xinsheng Zhang. A note on Slepian’s inequality based on majorization. Statistics & Probability Letters, 124(??): 132–139, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300330.

Dong:2018:NCS

[DZ18] Yuexiao Dong and Yongxu Zhang. On a new class of sufficient dimension reduction estimators. Statistics & Probability Letters, 139(??):90–94, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301391.

Davydov:2019:FTS

[DZ19a] Michel Davydov and Sergei Zuyev. Finite thinning-selfdecomposable point processes. Statistics & Probability Letters, 146(??):132– 138, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303572.

Dong:2019:SNL

[DZ19b] Hua Dong and Xiaowen Zhou. On a spectrally negative Lévy risk process with periodic dividends and capital injections. Statistics & Probability Letters, 154(??):Article 108589, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302354.

Deng:2012:GSD

[DZD12] Chao Deng, Jieming Zhou, and Yingchun Deng. The Gerber–Shiu discounted penalty function in a delayed renewal risk model with multi-layer dividend strategy. Statistics & Probability Letters, 82(9):1648–1656, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001678.

Ding:2018:SMT

[DZF18] Ying Ding, Xinsheng Zhang, and Longxiang Fang. Strong modified transportation cost inequalities on k-concave probability measures with heavy tails. Statistics & Probability Letters, 142(??):30–38, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302396.

Du:2018:WEL

[DZX18] Jiang Du, Zhongzhan Zhang, and Tianfa Xie. A weighted M-estimator for linear regression models with randomly truncated data. Statistics & Probability Letters, 138(??):90–94, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301007.

Ding:2013:CSH

[DZZ13] Weiyong Ding, Yiying Zhang, and Peng Zhao. Comparisons of k-out-of-n systems with heterogeneous components. Statistics & Probability Letters, 83(2):493–502, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003847.

Eghbal:2010:SMI

[EAB10] N. Eghbal, M. Amini, and A. Bozorgnia. Some maximal inequalities for quadratic forms of negative superadditive dependence random variables. Statistics & Probability Letters, 80(7–8):587–591, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004726.

Eghbal:2011:KIQ

[EAB11] N. Eghbal, M. Amini, and A. Bozorgnia. On the Kolmogorov inequalities for quadratic forms of dependent uniformly bounded random variables. Statistics & Probability Letters, 81(8):1112– 1120, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000885.

Erem:2017:EAD

[EB17] Aysegul Erem and Ismihan Bayramoglu. Exact and asymptotic distributions of exceedance statistics for bivariate random sequences. Statistics & Probability Letters, 125(??):181–188, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300706.

ElAsri:2016:WEM

[EBG16] M. El Asri, D. Blanke, and E. Gabriel. Weighted M-estimators for multivariate clustered data. Statistics & Probability Letters, 112(??):26–34, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000122.

Eck:2018:BML

[Eck18] Daniel J. Eck. Bootstrapping for multivariate linear regression models. Statistics & Probability Letters, 134(??):141–149, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303450.

Effraimidis:2014:NEC

[ED14] Georgios Effraimidis and Christian M. Dahl. Nonparametric estimation of cumulative incidence functions for competing risks data with missing cause of failure. Statistics & Probability Letters, 89(??):1–7, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000509.

Eisenberg:2017:PKB

[ED17] Bennett Eisenberg and Shuotao Diao. Properties of the Kelly bets for pairs of binary wagers. Statistics & Probability Letters, 125(??):215–219, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300755.

Estrade:2016:NCP

[EF16] Anne Estrade and Julie Fournier. Number of critical points of a Gaussian random field: Condition for a finite variance. Statistics & Probability Letters, 118(??):94–99, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300992.

Estrade:2019:CFS

[EFP19] Anne Estrade, Alessandra Fariñas, and Emilio Porcu. Covariance functions on spheres cross time: Beyond spatial isotropy and temporal stationarity. Statistics & Probability Letters, 151(??):1–7, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300872.

Efromovich:2010:SML

[Efr10] Sam Efromovich. A sharp minimax lower bound for the nonparametric estimation of Sobolev densities of order 1/2. Statistics & Probability Letters, 80(2):77–81, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003198.

Efromovich:2014:NES

[Efr14] Sam Efromovich. Nonparametric estimation of the spectral density of amplitude-modulated time series with missing observations. Statistics & Probability Letters, 93(??):7–13, October 1– 142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002156.

Efromovich:2019:SNE

[Efr19] Sam Efromovich. On sharp nonparametric estimation of differentiable functions. Statistics & Probability Letters, 152(??):9–14, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301105.

Ehler:2011:FTD

[EG11] Martin Ehler and Jennifer Galanis. Frame theory in directional statistics. Statistics & Probability Letters, 81(8):1046– 1051, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000733.

Eryilmaz:2016:GSW

[EGX16] Serkan Eryilmaz, Min Gong, and Min Xie. Generalized sooner waiting time problems in a sequence of trinary trials. Statistics & Probability Letters, 115(??):70–78, August 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215300705.

Ebner:2013:RTF

[EHP13] Bruno Ebner, Norbert Henze, and Panamalai R. Parthasarathy. Ramanujan theta functions and birth and death processes. Statistics & Probability Letters, 83(12):2647–2655, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002800.

Eisenberg:2015:MGR

[Eis15] Bennett Eisenberg. The multivariate Gini ratio. Statistics & Probability Letters, 96(??):292–298, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003587.

ElKtaibi:2014:BED

[EIW14] Farid El Ktaibi, B. Gail Ivanoff, and Neville C. Weber. Bootstrapping the empirical distribution of a linear process. Statistics & Probability Letters, 93(??):134–142, October 1–142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002211.

Evans:2011:LRP

[EJ11] Michael Evans and Gun Ho Jang. A limit result for the prior predictive applied to checking for prior-data conflict. Statistics & Probability Letters, 81(8):1034–1038, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100071X.

Ekstrom:2012:GMS

[EJ12] Magnus Ekström and Sreenivasa Rao Jammalamadaka. A general measure of skewness. Statistics & Probability Letters, 82(8):1559–1568, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200154X.

Ejsmont:2016:CND

[Ejs16] Wiktor Ejsmont. A characterization of the normal distribution by the independence of a pair of random vectors.

Statistics & Probability Letters, 114(??):1–5, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215302637.

Elie:2010:PRA

[EK10] Romuald Elie and Idris Kharroubi. Probabilistic representation and approximation for coupled systems of variational inequalities. Statistics & Probability Letters, 80(17–18):1388– 1396, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001392.

Engelke:2011:ERB

[EKS11] S. Engelke, Z. Kabluchko, and M. Schlather. An equivalent representation of the Brown–Resnick process. Statistics & Probability Letters, 81(8):1150–1154, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000939.

Ekstrom:2014:GCL

[Eks14] Magnus Ekström. A general central limit theorem for strong mixing sequences. Statistics & Probability Letters, 94(??): 236–238, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214002624.

ElMaroufy:2013:NTT

[El 13] Hamid El Maroufy. Note on the threshold theorem of a heterogeneous SIR epidemic. Statistics & Probability Letters, 83(1):292– 296, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003598.

ElHeda:2018:OBS

[EL18] Khadijetou El Heda and Djamal Louani. Optimal bandwidth selection in kernel density estimation for continuous time dependent processes. Statistics & Probability Letters, 138(??):9– 19, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300452.

Elhiwi:2014:DBI

[Elh14] Majdi Elhiwi. Default barrier intensity model for credit risk evaluation. Statistics & Probability Letters, 95(??):125–131, December 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002879.

Elliott:2013:FHS

[ELS13] Robert Elliott, Nikolaos Limnios, and Anatoliy Swishchuk. Filtering hidden semi-Markov chains. Statistics & Probability Letters, 83(9):2007–2014, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001685.

Elsawah:2016:COA

[Els16] A. M. Elsawah. Constructing optimal asymmetric combined designs via Lee discrepancy. Statistics & Probability Letters, 118(??):24–31, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300906.

Elek:2010:TBT

[EM10] Péter Elek and László Márkus. Tail behaviour of —β-TARCH models. Statistics & Probability Letters, 80(23–24): 1758–1763, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002142.

Eck:2016:CLT

[EM16] Daniel J. Eck and Ian W. McKeague. Central Limit Theorems under additive deformations. Statistics & Probability Letters, 118(??):156–162, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521630092X.

Emami:2015:IDR

[Ema15] Hadi Emami. Influence diagnostic in ridge semiparametric models. Statistics & Probability Letters, 105(??):106–113, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001960.

ElKolei:2017:PIA

[EP17] Salima El Kolei and Florian Pelgrin. Parametric inference of autoregressive heteroscedastic models with errors in variables. Statistics & Probability Letters, 130(??):63–70, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302444.

Ernst:2016:BAT

[EPSU16] Philip Ernst, Robin Pemantle, Ville Satopää, and Lyle Ungar. Bayesian aggregation of two forecasts in the partial information framework. Statistics & Probability Letters, 119(??):170–180, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301316.

Elsawah:2014:NLB

[EQ14] A. M. Elsawah and Hong Qin. New lower bound for centered L_{2}-discrepancy of four-level U-type designs. Statistics & Probability Letters, 93(??):65–71, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002107.

Elsawah:2015:LDS

[EQ15a] A. M. Elsawah and Hong Qin. Lee discrepancy on symmetric three-level combined designs. Statistics & Probability Letters, 96(??):273–280, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003472.

Elsawah:2015:MDS

[EQ15b] A. M. Elsawah and Hong Qin. Mixture discrepancy on symmetric balanced designs. Statistics & Probability Letters, 104 (??):123–132, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001583.

Elsawah:2015:NSO

[EQ15c] A. M. Elsawah and Hong Qin. A new strategy for optimal foldover two-level designs. Statistics & Probability Letters, 103(??):116–126, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001315.

Eftekharian:2017:EDF

[ER17] A. Eftekharian and M. Razmkhah. On estimating the distribution function and odds using ranked set sampling. Statistics & Probability Letters, 122(??):1–10, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302218.

Ernst:2017:MFI

[Ern17] Philip A. Ernst. Minimizing Fisher information with absolute moment constraints. Statistics & Probability Letters, 129(??):167–170, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301906.

Eryilmaz:2010:SRS

[Ery10] Serkan Eryilmaz. On system reliability in stress-strength setup. Statistics & Probability Letters, 80(9–10):834–839, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000295.

Eryilmaz:2011:BWS

[Ery11a] Serkan Eryilmaz. The behavior of warm standby components with respect to a coherent system. Statistics & Probability Letters, 81(8):1319–1325, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100126X.

Eryilmaz:2011:JDR

[Ery11b] Serkan Eryilmaz. Joint distribution of run statistics in partially exchangeable processes. Statistics & Probability Letters, 81(1):163–168, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521000297X.

Eryilmaz:2012:GSM

[Ery12] Serkan Eryilmaz. Generalized δ-shock model via runs. Statistics & Probability Letters, 82(2):326–331, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003439.

Eryilmaz:2014:GDO

[Ery14] Serkan Eryilmaz. Geometric distribution of order k with a reward. Statistics & Probability Letters, 92(??):53–58, September 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001825.

Eryilmaz:2015:DTS

[Ery15] Serkan Eryilmaz. Discrete time shock models involving runs. Statistics & Probability Letters, 107:93–100, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002862.

Eryilmaz:2016:NCL

[Ery16] Serkan Eryilmaz. A new class of lifetime distributions. Statistics & Probability Letters, 112(??):63–71, May 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215303746.

Eryilmaz:2018:SRS

[Ery18] Serkan Eryilmaz. On success runs in a sequence of dependent trials with a change point. Statistics & Probability Letters, 132(??):91–98, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302936.

Es-Sebaiy:2013:BEB

[ES13] Khalifa Es-Sebaiy. Berry–Esséen bounds for the least squares estimator for discretely observed fractional Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 83(10):2372–2385, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002447.

Ezzaki:2019:RTS

[ET19] Fatima Ezzaki and Khalid Tahri. Representation theorem of set valued regular martingale: Application to the convergence of set valued martingale. Statistics & Probability Letters, 154(??):Article 108548, November 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301853.

Fashandi:2012:CSD

[FA12] M. Fashandi and Jafar Ahmadi. Characterizations of symmetric distributions based on Rényi entropy. Statistics & Probability Letters, 82(4):798–804, April 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000053.

Fallahpour:2014:SEV

[FA14] Saber Fallahpour and S. Ejaz Ahmed. Shrinkage estimation and variable selection in multiple regression models with random coefficient autoregressive errors. Statistics & Probability Letters, 92(??):199–208, September 1–262, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001965.

Faraway:2018:WSD

[FA18] Julian J. Faraway and Nicole H. Augustin. When small data beats big data. Statistics & Probability Letters, 136(??):142– 145, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300762.

Fakoor:2010:SUC

[Fak10] V. Fakoor. Strong uniform consistency of kernel density estimators under a censored dependent model. Statistics & Probability Letters, 80(5–6):318–323, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004234.

Fakhfakh:2017:MRC

[Fak17] Raouf Fakhfakh. The mean of the reciprocal in a Cauchy– Stieltjes family. Statistics & Probability Letters, 129(??):1–11, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301608.

Fan:2015:LSI

[Fan15a] XiLiang Fan. Logarithmic Sobolev inequalities for fractional diffusion. Statistics & Probability Letters, 106(??):165–172, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521500262X.

Fang:2015:LMD

[Fan15b] Lulu Fang. Large and moderate deviations for modified Engel continued fractions. Statistics & Probability Letters, 98(??):98– 106, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214004180.

Fan:2016:ESO

[Fan16] ShengJun Fan. Existence of solutions to one-dimensional BSDEs with semi-linear growth and general growth generators. Statistics & Probability Letters, 109:7–15, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003636.

Fan:2017:SND

[Fan17] Xiequan Fan. Self-normalized deviation inequalities with application to t-statistic. Statistics & Probability Letters, 127(??):158–164, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301372.

Farcomeni:2011:HMP

[Far11] Alessio Farcomeni. Hidden Markov partition models. Statistics & Probability Letters, 81(12):1766–1770, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002434.

Fisher:2010:PGO

[FC10] Evan Fisher and Shiliang Cui. Patterns generated by m-th-order Markov chains. Statistics & Probability Letters, 80(15– 16):1157–1166, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000908.

Feng:2015:LDE

[FC15] Bo Feng and Shouquan Chen. On large deviations of extremes under power normalization. Statistics & Probability Letters, 99(??):27–35, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000097.

Falk:2018:SRJ

[FCP18] M. Falk, A. Khorrami Chokami, and S. A. Padoan. Some results on joint record events. Statistics & Probability Letters, 135(??): 11–19, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303656.

Ferreiro-Castilla:2011:LAG

[FCU11] Albert Ferreiro-Castilla and Frederic Utzet. Lévy area for Gaussian processes: a double Wiener–Itô integral approach. Statistics & Probability Letters, 81(9):1380–1391, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001556.

Finner:2010:EER

[FD10] Helmut Finner and Thorsten Dickhaus. Edgeworth expansions and rates of convergence for normalized sums: Chung’s 1946 method revisited. Statistics & Probability Letters, 80(23–24):1875–1880, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002373.

Fang:2018:RLC

[FD18] Rui Fang and Weiyong Ding. On relative log-concavity and stochastic comparisons. Statistics & Probability Letters, 137(??): 91–98, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300099.

Feng:2014:DDE

[Fen14] Yuanhua Feng. Data-driven estimation of diurnal patterns of durations between trades on financial markets. Statistics & Probability Letters, 92(??):109–113, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001874.

Feng:2017:SNL

[Fen17] Xinwei Feng. Self-normalized large deviations under sublinear expectation. Statistics & Probability Letters, 123(??):77– 83, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216302899.

Feng:2019:LLW

[Fen19] Xinwei Feng. Law of the logarithm for weighted sums of negatively dependent random variables under sublinear expectation. Statistics & Probability Letters, 149(??):132–141, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300392.

Ferreira:2011:DBT

[Fer11] H. Ferreira. Dependence between two multivariate extremes. Statistics & Probability Letters, 81(5):586–591, May 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000216.

Ferreira:2012:MMM

[Fer12] Helena Ferreira. Multivariate maxima of moving multivariate maxima. Statistics & Probability Letters, 82(8):1489–1496, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001587.

Ferger:2014:OCM

[Fer14] Dietmar Ferger. Optimal constants in the Marcinkiewicz– Zygmund inequalities. Statistics & Probability Letters, 84 (??):96–101, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003271.

Ferger:2018:SBB

[Fer18] Dietmar Ferger. On the supremum of a Brownian bridge standardized by its maximizing point with applications to statistics. Statistics & Probability Letters, 134(??):63–69, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521730322X.

Froda:2012:EPP

[FF12] Sorana Froda and René Ferland. Estimating the parameters of a Poisson process model for predator-prey interactions. Statistics & Probability Letters, 82(12):2252–2259, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002891.

Ferreira:2014:EBP

[FF14] Helena Ferreira and Marta Ferreira. Extremal behavior of pMAX processes. Statistics & Probability Letters, 93(??):46–57, October 1–142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002119.

Ferreira:2018:MED

[FF18] Helena Ferreira and Marta Ferreira. Multidimensional extremal dependence coefficients. Statistics & Probability Letters, 133(??): 1–8, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303127.

Fasso:2018:SIR

[FFM18] A. Fassò, F. Finazzi, and F. Madonna. Statistical issues in radiosonde observation of atmospheric temperature and humidity profiles. Statistics & Probability Letters, 136(??):97–100, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300725.

Frahm:2012:SLT

[FG12] Gabriel Frahm and Konstantin Glombek. Semicircle law of Tyler’s M-estimator for scatter. Statistics & Probability Letters, 82(5):959–964, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000272.

Fakoor:2011:ABL

[FGA11] V. Fakoor, M. Bolbolian Ghalibaf, and H. A. Azarnoosh. Asymptotic behaviors of the Lorenz curve and Gini index in sampling from a length-biased distribution. Statistics & Probability Letters, 81(9):1425–1435, September 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001532.

Filipiak:2012:UOD

[Fil12] Katarzyna Filipiak. Universally optimal designs under an interference model with equal left- and right-neighbor effects. Statistics & Probability Letters, 82(3):592–598, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100366X.

Finos:2011:NLS

[Fin11] Livio Finos. A note on Left-Spherically Distributed test with covariates. Statistics & Probability Letters, 81(6):639–641, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000381.

Fischer:2011:NGC

[Fis11] Aurélie Fischer. On the number of groups in clustering. Statistics & Probability Letters, 81(12):1771–1781, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002367.

Fischer:2013:SRS

[Fis13] Tom Fischer. On simple representations of stopping times and stopping time sigma-algebras. Statistics & Probability Letters, 83(1):345–349, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003707.

Fitzpatrick:2014:GEG

[Fit14] Matthew Fitzpatrick. Geometric ergodicity of the Gibbs sampler for the Poisson change-point model. Statistics & Probability Letters, 91(??):55–61, August 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001394.

Fan:2010:FIT

[FJ10] ShengJun Fan and Long Jiang. Finite and infinite time interval BSDEs with non-Lipschitz coefficients. Statistics & Probability Letters, 80(11–12):962–968, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000490.

Fan:2012:ODB

[FJ12] ShengJun Fan and Long Jiang. One-dimensional BSDEs with left-continuous, lower semi-continuous and linear-growth generators. Statistics & Probability Letters, 82(10):1792–1798, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002155.

Fotopoulos:2015:JDS

[FJW15] Stergios Fotopoulos, Venkata Jandhyala, and Jun Wang. On the joint distribution of the supremum functional and its last occurrence for subordinated linear Brownian motion. Statistics & Probability Letters, 106(??):149–156, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002515.

Ferenstein:2010:NIS

[FK10] Elzbieta Z. Ferenstein and Anna Krasnosielska. No-information secretary problems with cardinal payoffs and Poisson arrivals. Statistics & Probability Letters, 80(3–4):221–227, February 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003964.

Farchione:2012:CIR

[FK12] Davide Farchione and Paul Kabaila. Confidence intervals in regression centred on the SCAD estimator. Statistics & Probability Letters, 82(11):1953–1960, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002611.

Forde:2015:LDB

[FKZ15] Martin Forde, Rohini Kumar, and Hongzhong Zhang. Large deviations for the boundary local time of doubly reflected Brownian motion. Statistics & Probability Letters, 96(??):262–268, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003150.

Fan:2010:CBI

[FL10a] ShengJun Fan and DeQun Liu. A class of BSDE with integrable parameters. Statistics & Probability Letters, 80(23–24): 2024–2031, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002592.

Figueroa-Lopez:2010:ADB

[FL10b] José E. Figueroa-López. Approximations for the distributions of bounded variation Lévy processes. Statistics & Probability Letters, 80(23–24):1744–1757, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002130.

Flick:2010:QST

[FL10c] Allen Flick and Ming Liao. A queuing system with time varying rates. Statistics & Probability Letters, 80(5–6):386–389, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209004398.

Fang:2019:CCS

[FL19] Rongjuan Fang and Zenghu Li. A conditioned continuousstate branching process with applications. Statistics & Probability Letters, 152(??):43–49, September 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301166.

Fierro:2013:BSD

[FLRS13] Raúl Fierro, Víctor Leiva, Fabrizio Ruggeri, and Antonio Sanhueza. On a Birnbaum–Saunders distribution arising from a non-homogeneous Poisson process. Statistics & Probability Letters, 83(4):1233–1239, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004774.

Fouche:2013:FSZ

[FM13] Willem L. Fouché and Safari Mukeru. On the Fourier structure of the zero set of fractional Brownian motion. Statistics & Probability Letters, 83(2):459–466, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003872.

Farias:2016:BIE

[FMA16] Rafael B. A. Farias, Michel H. Montoril, and José A. A. Andrade. Bayesian inference for extreme quantiles of heavy tailed distributions. Statistics & Probability Letters, 113(??):103–107, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215301942.

Fort:2012:SVI

[FMPV12] G. Fort, E. Moulines, P. Priouret, and P. Vandekerkhove. A simple variance inequality for U-statistics of a Markov chain with applications. Statistics & Probability Letters, 82(6):1193– 1201, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000387.

Fu:2017:UAR

[FN17] Ke-Ang Fu and Cheuk Yin Andrew Ng. Uniform asymptotics for the ruin probabilities of a two-dimensional renewal risk model with dependent claims and risky investments. Statistics & Probability Letters, 125(??):227–235, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302960.

Fazekas:2018:PEM

[FNP18] I. Fazekas, Cs. Noszály, and A. Perecsényi. A population evolution model and its applications to random networks. Statistics & Probability Letters, 143(??):17–27, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302566.

Forde:2011:LTA

[For11] Martin Forde. Large-time asymptotics for an uncorrelated stochastic volatility model. Statistics & Probability Letters, 81(8):1230–1232, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001040.

Forde:2014:LMS

[For14a] Martin Forde. The large-maturity smile for the Stein–Stein model. Statistics & Probability Letters, 91(??):145–152, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001400.

Forde:2014:MPB

[For14b] Martin Forde. On the Markovian projection in the Brunick– Shreve mimicking result. Statistics & Probability Letters, 85(??):98–105, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003866.

Forrester:2018:CSS

[For18] Peter J. Forrester. Comment on “Sum of squares of uniform random variables” by I. Weissman. Statistics & Probability Letters, 142(??):118–122, November 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301779. See [Wei17].

Fama:2011:TSE

[FP11a] Yuchen Fama and Vladimir Pozdnyakov. A test for selfexciting clustering mechanism. Statistics & Probability Letters, 81(10):1541–1546, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001702.

Farcomeni:2011:CEP

[FP11b] Alessio Farcomeni and Simona Pacillo. A conservative estimator for the proportion of false nulls based on Dvoretzky, Kiefer and Wolfowitz inequality. Statistics & Probability Letters, 81(12):1867–1870, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002483.

Fortini:2012:HRU

[FP12] S. Fortini and S. Petrone. Hierarchical reinforced urn processes. Statistics & Probability Letters, 82(8):1521–1529, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001551.

Fosdick:2013:CNR

[FP13] Bailey K. Fosdick and Michael D. Perlman. Covariate and Newton–Raphson adjustments for a normal correlation coefficient when the variances are known. Statistics & Probability Letters, 83(12):2627–2633, December 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002824.

Feng:2014:NMK

[FP14] Pengbin Feng and Xuhui Peng. A note on Monge–Kantorovich problem. Statistics & Probability Letters, 84(??):204–211, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003544.

Funke:2017:NEC

[FP17] Benedikt Funke and Christian Palmes. A note on estimating cumulative distribution functions by the use of convolution power kernels. Statistics & Probability Letters, 121(??):90–98, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302115.

Feng:2013:IES

[FPZ13] Huijun Feng, Liang Peng, and Fukang Zhu. Interval estimation for a simple bilinear model. Statistics & Probability Letters, 83(10):2152–2159, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002058.

Faliva:2018:GCL

[FQZ18] Mario Faliva, Piero Quatto, and Maria Grazia Zoia. Gram– Charlier-like expansions of power-raised hyperbolic secant laws. Statistics & Probability Letters, 137(??):229–234, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300385.

Framstad:2011:PSP

[Fra11] N. C. Framstad. Portfolio separation properties of the skewelliptical distributions, with generalizations. Statistics & Probability Letters, 81(12):1862–1866, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002379.

Fralix:2015:WTM

[Fra15] Brian Fralix. When are two Markov chains similar? Statistics & Probability Letters, 107:199–203, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003193.

Frey:2014:SNP

[Fre14] Jesse Frey. Shorter nonparametric prediction intervals for an order statistic from a future sample. Statistics & Probability Letters, 91(??):69–75, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214001424.

Frey:2016:EKS

[Fre16] Jesse Frey. An exact Kolmogorov–Smirnov test for whether two finite populations are the same. Statistics & Probability Letters, 116(??):65–71, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300311.

Frolov:2012:BPU

[Fro12] Andrei N. Frolov. Bounds for probabilities of unions of events and the Borel–Cantelli lemma. Statistics & Probability Letters, 82(12):2189–2197, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002970.

Frolov:2015:BRC

[Fro15] Andrei N. Frolov. Bounds of the remainder in a combinatorial central limit theorem. Statistics & Probability Letters, 105(??):37–46, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001832.

Frolov:2017:IVF

[Fro17] Andrei N. Frolov. On inequalities for values of first jumps of distribution functions and Hölder’s inequality. Statistics & Probability Letters, 126(??):150–156, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300950.

Fan:2010:NDR

[FRZ10] Xiliang Fan, Yong Ren, and Dongjin Zhu. A note on the doubly reflected backward stochastic differential equations driven by a Lévy process. Statistics & Probability Letters, 80(7–8):690–696, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000027.

Fung:2010:TDT

[FS10] Thomas Fung and Eugene Seneta. Tail dependence for two skew t distributions. Statistics & Probability Letters, 80(9–10):784–791, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000167.

Fung:2011:BNC

[FS11] Thomas Fung and Eugene Seneta. The bivariate normal copula function is regularly varying. Statistics & Probability Letters, 81(11):1670–1676, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002057.

Finlay:2012:GHM

[FS12] Richard Finlay and Eugene Seneta. A generalized hyperbolic model for a risky asset with dependence. Statistics & Probability Letters, 82(12):2164–2169, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002787.

Fourdrinier:2014:NEU

[FS14] Dominique Fourdrinier and William Strawderman. On the non existence of unbiased estimators of risk for spherically symmetric distributions. Statistics & Probability Letters, 91(??):6– 13, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001151.

Feng:2015:NHD

[FS15] Long Feng and Fasheng Sun. A note on high-dimensional twosample test. Statistics & Probability Letters, 105(??):29–36, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001789.

Fourdrinier:2016:STS

[FS16] Dominique Fourdrinier and William E. Strawderman. Stokes’ theorem, Stein’s identity and completeness. Statistics & Probability Letters, 109:224–231, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0167715215302613.

Finlay:2017:SVI

[FS17a] Richard Finlay and Eugene Seneta. A scalar-valued infinitely divisible random field with Pólya autocorrelation. Statistics & Probability Letters, 122(??):141–146, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302462.

Foondun:2017:LDC

[FS17b] Mohammud Foondun and Leila Setayeshgar. Large deviations for a class of semilinear stochastic partial differential equations. Statistics & Probability Letters, 121(??):143–151, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630219X.

Fink:2018:FLC

[FS18] Holger Fink and Georg Schlüchtermann. Fractional Lévy Cox– Ingersoll–Ross and Jacobi processes. Statistics & Probability Letters, 142(??):84–91, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302475.

Fernandez-Sanchez:2011:MCQ

[FSNÚF11] Juan Fernández-Sánchez, Roger B. Nelsen, and Manuel Úbeda-Flores. Multivariate copulas, quasi-copulas and lattices. Statistics & Probability Letters, 81(9):1365–1369, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001350.

Fujii:2010:ECE

[Fuj10] Takayuki Fujii. An extension of cusp estimation problem in ergodic diffusion processes. Statistics & Probability Letters, 80(9–10):779–783, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000155.

Furman:2008:MGD

[Fur08] Edward Furman. On a multivariate gamma distribution. Statistics & Probability Letters, 78(15):2353–2360, October 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715208001338. See erratum [SF12].

Falk:2017:AIB

[FW17] Michael Falk and Florian Wisheckel. Asymptotic independence of bivariate order statistics. Statistics & Probability Letters, 125(??):91–98, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300421.

Fu:2012:OBA

[FXT12] Jiayu Fu, Ancha Xu, and Yincai Tang. Objective Bayesian analysis of Pareto distribution under progressive Type-II censoring. Statistics & Probability Letters, 82(10):1829–1836, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002180.

Fan:2016:LEP

[FY16] Xiliang Fan and Chenggui Yuan. Lyapunov exponents of PDEs driven by fractional noise with Markovian switching. Statistics & Probability Letters, 110(??):39–50, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003910.

Fang:2013:SCS

[FZ13] Longxiang Fang and Xinsheng Zhang. Stochastic comparisons of series systems with heterogeneous Weibull components. Statistics & Probability Letters, 83(7):1649–1653, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000953.

Fang:2015:SCP

[FZ15] Longxiang Fang and Xinsheng Zhang. Stochastic comparisons of parallel systems with exponentiated Weibull components. Statistics & Probability Letters, 97(??):25–31, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003666.

Fang:2016:SCP

[FZB16] Longxiang Fang, Xiaojun Zhu, and N. Balakrishnan. Stochastic comparisons of parallel and series systems with heterogeneous Birnbaum–Saunders components. Statistics & Probability Letters, 112(??):131–136, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000146.

Feng:2012:RBI

[FZW12] Long Feng, Changliang Zou, and Zhaojun Wang. Rank-based inference for the single-index model. Statistics & Probability Letters, 82(3):535–541, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003798.

Govindan:2012:NES

[GA12] T. E. Govindan and N. U. Ahmed. A note on exponential state feedback stabilizability by a Razumikhin type theorem of mild solutions of SDEs with delay. Statistics & Probability Letters, 82(7):1303–1309, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001204.

Govindan:2013:RSG

[GA13] T. E. Govindan and N. U. Ahmed. Robust stabilization with a general decay of mild solutions of stochastic evolution equations. Statistics & Probability Letters, 83(1):115–122, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003240.

Goodarzi:2016:UBV

[GAB16] F. Goodarzi, M. Amini, and G. R. Mohtashami Borzadaran. On upper bounds for the variance of functions of the inactivity time. Statistics & Probability Letters, 117(??):62–71, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302364.

Gao:2019:ECR

[Gao19] Zhi-Qiang Gao. Exact convergence rate in the local central limit theorem for a lattice branching random walk on Zd. Statistics & Probability Letters, 151(??):58–66, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300926.

Gasanenko:2010:SDB

[Gas10] Vitalii Gasanenko. Small deviation of Brownian motion with large drift. Statistics & Probability Letters, 80(21–22):1618– 1622, November 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521000177X.

Giuliano-Antonini:2013:ASL

[GAS13] Rita Giuliano-Antonini and Zbigniew S. Szewczak. An almost sure local limit theorem for Markov chains. Statistics & Probability Letters, 83(2):573–579, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003835.

Gatto:2013:MFD

[Gat13] Riccardo Gatto. The von Mises–Fisher distribution of the first exit point from the hypersphere of the drifted Brownian motion and the density of the first exit time. Statistics & Probability Letters, 83(7):1669–1676, July 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300093X.

Gatto:2015:LEE

[Gat15] Riccardo Gatto. A logarithmic efficient estimator of the probability of ruin with recuperation for spectrally negative Lévy risk processes. Statistics & Probability Letters, 99(??):177–184, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000255.

Gaur:2014:NCD

[Gau14] Anil Gaur. A new class of distribution-free tests for testing ordered location parameters based on sub-samples. Statistics & Probability Letters, 90(??):53–59, July 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001047.

Gaunt:2019:SOV

[Gau19] Robert E. Gaunt. Stein operators for variables form the third and fourth Wiener chaoses. Statistics & Probability Letters, 145(??):118–126, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302943.

Gupta:2013:PMB

[GB13] Nitin Gupta and Rakesh Kumar Bajaj. On partial monotonic behaviour of some entropy measures. Statistics & Probability Letters, 83(5):1330–1338, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000400.

Ghosh:2016:TCN

[GB16] Abhik Ghosh and Ayanendranath Basu. Testing composite null hypotheses based on S-divergences. Statistics & Probability Letters, 114(??):38–47, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216000365.

Ghosh:2019:PLR

[GB19a] Abhik Ghosh and Ayanendranath Basu. Power and level robustness of a test for composite hypotheses under independent non-homogeneous data. Statistics & Probability Letters, 148(??): 35–42, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303900.

Giussani:2019:NLB

[GB19b] A. Giussani and M. Bonetti. A note on the lengthbiased Weibull–Gamma frailty survival model. Statistics & Probability Letters, 153(??):32–36, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301385.

Gonzalez-Barrios:2019:SND

[GBGPR19] José M. González-Barrios, Eduardo Gutiérrez-Peña, and Raúl Rueda. A short note on the dependence structure of random vectors. Statistics & Probability Letters, 146(??):200–205, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830378X.

Ginestet:2012:CLF

[GBR12] Cedric E. Ginestet, Nicky G. Best, and Sylvia Richardson. Classification loss function for parameter ensembles in Bayesian hierarchical models. Statistics & Probability Letters, 82(4):859– 863, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521100407X.

Garcia-Belmonte:2011:SRL

[GBVS11] Lizeth García-Belmonte and Daniel Ventosa-Santaulària. Spurious regression and lurking variables. Statistics & Probability Letters, 81(12):2004–2010, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002793.

Garg:2015:ABS

[GD15] Mansi Garg and Isha Dewan. On asymptotic behavior of U-statistics for associated random variables. Statistics & Probability Letters, 105(??):209–220, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002059.

Garg:2018:LDS

[GD18] Mansi Garg and Isha Dewan. On limiting distribution of U-statistics based on associated random variables. Statistics & Probability Letters, 132(??):7–16, January 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302857.

Gorbach:2018:IPC

[GdL18] Tetiana Gorbach and Xavier de Luna. Inference for partial correlation when data are missing not at random. Statistics & Probability Letters, 141(??):82–89, October 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302141.

Giraldo:2018:SMS

[GDNM18] Ramón Giraldo, Sophie Dabo-Niang, and Sergio Martínez. Statistical modeling of spatial big data: an approach from a functional data analysis perspective. Statistics & Probability Letters, 136(??):126–129, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300701.

Garthwaite:2017:MCI

[GEC17] Paul H. Garthwaite, Fadlalla G. Elfadaly, and John R. Crawford. Modified confidence intervals for the Mahalanobis distance. Statistics & Probability Letters, 127(??):131–137, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301293.

Geenens:2015:MEA

[Gee15] Gery Geenens. Moments, errors, asymptotic normality and large deviation principle in nonparametric functional regression. Statistics & Probability Letters, 107:369–377, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521500334X.

Geiger:2017:SCU

[Gei17] Bernhard C. Geiger. A sufficient condition for a unique invariant distribution of a higher-order Markov chain. Statistics & Probability Letters, 130(??):49–56, November 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302390.

Gerhold:2011:MEL

[Ger11] Stefan Gerhold. Moment explosion in the LIBOR market model. Statistics & Probability Letters, 81(5):560–562, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000162.

Gera:2013:SDP

[Ger13] Amos E. Gera. A start-up demonstration procedure involving dependent tests. Statistics & Probability Letters, 83(10):2191– 2196, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002046.

Gera:2018:SDT

[Ger18] Amos E. Gera. Simultaneous demonstration tests involving sparse failures. Statistics & Probability Letters, 135(??):26– 31, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303723.

Gonzalez-Estrada:2016:RGF

[GEV16] Elizabeth González-Estrada and José A. Villaseñor. A ratio goodness-of-fit test for the Laplace distribution. Statistics & Probability Letters, 119(??):30–35, December 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301195.

Geluk:2011:RTR

[GF11] J. L. Geluk and J. B. G. Frenk. Renewal theory for random variables with a heavy tailed distribution and finite variance. Statistics & Probability Letters, 81(1):77–82, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002713.

Gontscharuk:2013:AFC

[GF13] Veronika Gontscharuk and Helmut Finner. Asymptotic FDR control under weak dependence: a counterexample. Statistics & Probability Letters, 83(8):1888–1893, August 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001466.

Ghalibaf:2010:SGA

[GFA10] M. Bolbolian Ghalibaf, V. Fakoor, and H. A. Azarnoosh. Strong Gaussian approximations of product-limit and quantile processes for truncated data under strong mixing. Statistics & Probability Letters, 80(7–8):581–586, April 1–15, 2010.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004714.

Gilboa-Freedman:2016:WMC

[GFH16] Gail Gilboa-Freedman and Refael Hassin. When Markov chains meet: A continuous-time model of network evolution. Statistics & Probability Letters, 116(??):131–138, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215304132.

Guo:2012:MSE

[GG12] Meixi Guo and Malay Ghosh. Mean squared error of James– Stein estimators for measurement error models. Statistics & Probability Letters, 82(11):2033–2043, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002532.

Gormley:2013:ESC

[GG13] R. Gormley and I. B. J. Goudie. Estimation for the Schnabel census with plants. Statistics & Probability Letters, 83(7):1740– 1744, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001132.

Goegebeur:2015:ETI

[GGO15] Yuri Goegebeur, Armelle Guillou, and Michael Osmann. An estimator for the tail index of an integrated conditional Pareto– Weibull-type model. Statistics & Probability Letters, 103(??): 8–16, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001182.

Goegebeur:2017:KES

[GGQ17] Yuri Goegebeur, Armelle Guillou, and Jing Qin. On kernel estimation of the second order rate parameter in multivariate extreme value statistics. Statistics & Probability Letters, 128(??):35–43, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301542.

Ghosh:2011:CMN

[GGR11] Subhankar Ghosh, Larry Goldstein, and Martin Raic. Concentration of measure for the number of isolated vertices in the Erdős-Rényi random graph by size bias couplings. Statistics & Probability Letters, 81(11):1565–1570, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002045.

Girard:2012:EEH

[GGS12] Stéphane Girard, Armelle Guillou, and Gilles Stupfler. Estimating an endpoint with high order moments in the Weibull domain of attraction. Statistics & Probability Letters, 82(12):2136–2144, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002775.

Goegebeur:2014:RAU

[GGV14] Yuri Goegebeur, Armelle Guillou, and Andréhette Verster. Robust and asymptotically unbiased estimation of extreme quantiles for heavy tailed distributions. Statistics & Probability Letters, 87(??):108–114, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000212.

Ghosh:2010:RLR

[GHH^{+}10] Arka P. Ghosh, Diana Hay, Vivek Hirpara, Reza Rastegar, Alexander Roitershtein, Ashley Schulteis, and Jiyeon Suh. Random linear recursions with dependent coefficients. Statistics & Probability Letters, 80(21–22):1597–1605, November 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001744.

Ghorbel:2011:ANA

[Gho11a] M. Ghorbel. Analytic and numerical analysis of some statistical features of fragmentation processes. Statistics & Probability Letters, 81(12):1953–1960, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002550.

Ghosh:2011:PSM

[Gho11b] Debashis Ghosh. Propensity score modelling in observational studies using dimension reduction methods. Statistics & Probability Letters, 81(7):813–820, July 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100085X.

Ghosh:2012:CFS

[Gho12] Debashis Ghosh. A causal framework for surrogate endpoints with semi-competing risks data. Statistics & Probability Letters, 82(11):1898–1902, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002210.

Ghosh:2014:AMK

[Gho14] Debashis Ghosh. An asymptotically minimax kernel machine. Statistics & Probability Letters, 95(??):33–38, ???? 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002843.

Gomes:2016:CEE

[GHR16] M. Ivette Gomes and Lígia Henriques-Rodrigues. Competitive estimation of the extreme value index. Statistics & Probability Letters, 117(??):128–135, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300621.

Gomes:2017:ECE

[GHR17] M. Ivette Gomes and Lígia Henriques-Rodrigues. Erratum to “Competitive estimation of the extreme value index” [Statist. Probab. Lett. 117 (2016) 128–135]. Statistics & Probability Letters, 130(??):40–41, November 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302171.

Goldstein:2014:CIZ

[GI14] Larry Goldstein and Ümit Islak. Concentration inequalities via zero bias couplings. Statistics & Probability Letters, 86(??):17– 23, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213004008.

Girard:2016:ANE

[Gir16] Didier A. Girard. Asymptotic near-efficiency of the “Gibbsenergy and empirical-variance” estimating functions for fitting Matérn models - I: Densely sampled processes. Statistics & Probability Letters, 110(??):191–197, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215302935.

Giurcanu:2015:SAN

[Giu15] Mihai C. Giurcanu. A simulation algorithm for non-causal VARMA processes. Statistics & Probability Letters, 98(??):65– 72, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214004088.

Giudici:2018:FDS

[Giu18] Paolo Giudici. Financial data science. Statistics & Probability Letters, 136(??):160–164, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300695.

Groeneboom:2010:GCI

[GJ10] Piet Groeneboom and Geurt Jongbloed. Generalized continuous isotonic regression. Statistics & Probability Letters, 80 (3–4):248–253, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209004064.

Goldberg:2012:EBC

[GK12] Y. Goldberg and M. R. Kosorok. An exponential bound for Cox regression. Statistics & Probability Letters, 82(7):1267– 1272, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001162.

Gajda:2019:SLP

[GKW19] J. Gajda, A. Kumar, and A. Wyloma’nska. Stable Lévy process delayed by tempered stable subordinator. Statistics & Probability Letters, 145(??):284–292, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303079.

Gao:2013:UAF

[GL13] Qingwu Gao and Xijun Liu. Uniform asymptotics for the finite-time ruin probability with upper tail asymptotically independent claims and constant force of interest. Statistics & Probability Letters, 83(6):1527–1538, June 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000680.

Ghaderinezhad:2019:QIP

[GL19] Fatemeh Ghaderinezhad and Christophe Ley. Quantification of the impact of priors in Bayesian statistics via stein’s method. Statistics & Probability Letters, 146(??):206–212, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303596.

Gomez:2013:SUS

[GLM^{+}13] Alejandro Gomez, Kijung Lee, Carl Mueller, Ang Wei, and Jie Xiong. Strong uniqueness for an SPDE via backward doubly stochastic differential equations. Statistics & Probability Letters, 83(10):2186–2190, October 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002150.

Guerrero:2019:PEE

[GLM19] Eugenio Guerrero and José Alfredo López-Mimbela. Pointwise eigenfunction estimates and mean L_{p}-norm blowup of a system of semilinear SPDEs with symmetric Lévy generators. Statistics & Probability Letters, 149(??):47–54, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300215.

Goncalves:2012:PSP

[GLML12] E. Gonçalves, J. Leite, and N. Mendes-Lopes. On the probabilistic structure of power threshold generalized arch stochastic processes. Statistics & Probability Letters, 82(8):1597–1609, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001575.

Gheriballah:2013:NRF

[GLS13] Abdelkader Gheriballah, Ali Laksaci, and Soumeya Sekkal. Nonparametric M-regression for functional ergodic data. Statistics & Probability Letters, 83(3):902–908, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004531.

Grahovac:2018:ITP

[GLT18] Danijel Grahovac, Nikolai N. Leonenko, and Murad S. Taqqu. Intermittency of trawl processes. Statistics & Probability Letters, 137(??):235–242, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300415.

Ghribi:2010:CME

[GM10a] Abdelaziz Ghribi and Afif Masmoudi. Characterization of multinomial exponential families by generalized variance. Statistics & Probability Letters, 80(11–12):939–944, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000441.

Guo:2010:NAA

[GM10b] Ying Guo and Amita K. Manatunga. A note on assessing agreement for frailty models. Statistics & Probability Letters, 80(7–8):527–533, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004568.

Gajda:2014:LDS

[GM14a] Janusz Gajda and Marcin Magdziarz. Large deviations for subordinated Brownian motion and applications. Statistics & Probability Letters, 88(??):149–156, May 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000522.

Gine:2014:WPK

[GM14b] Evarist Giné and W. R. Madych. On wavelet projection kernels and the integrated squared error in density estimation. Statistics & Probability Letters, 91(??):32–40, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001229.

Greenberg:2014:TLB

[GM14c] Spencer Greenberg and Mehryar Mohri. Tight lower bound on the probability of a binomial exceeding its expectation. Statistics & Probability Letters, 86(??):91–98, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213004082.

Glazyrina:2016:BIT

[GM16] Anna Glazyrina and Alexander Melnikov. Bernstein’s inequalities and their extensions for getting the Black–Scholes option pricing formula. Statistics & Probability Letters, 111(??):86– 92, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215304430.

Ghosh:2017:HPT

[GM17a] Shyamal Ghosh and Murari Mitra. A Hollander–Proschan type test when ageing is not monotone. Statistics & Probability Letters, 121(??):119–127, February 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302140.

Ghosh:2017:WIA

[GM17b] Shyamal Ghosh and Murari Mitra. A weighted integral approach to testing against HNBUE alternatives. Statistics & Probability Letters, 129(??):58–64, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301785.

Gao:2018:SEF

[GM18a] Shuyang Gao and Hosam M. Mahmoud. A self-equilibrium Friedman-like urn via stochastic approximation. Statistics & Probability Letters, 142(??):77–83, November 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302499.

Giuliano:2018:LDS

[GM18b] Rita Giuliano and Claudio Macci. Large deviations for some logarithmic means in the case of random variables with thin tails. Statistics & Probability Letters, 138(??):47–56, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301111.

Gallaugher:2019:TSM

[GM19] Michael P. B. Gallaugher and Paul D. McNicholas. Three skewed matrix variate distributions. Statistics & Probability Letters, 145(??):103–109, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302876.

Gaur:2012:NNT

[GMA12] Anil Gaur, Kalpana K. Mahajan, and Sangeeta Arora. New nonparametric tests for testing homogeneity of scale parameters against umbrella alternative. Statistics & Probability Letters, 82(9):1681–1689, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001848.

Gupta:2018:QLB

[GMDP18] Shivam Gupta, Jorge Mateu, Auriol Degbelo, and Edzer Pebesma. Quality of life, big data and the power of statistics. Statistics & Probability Letters, 136(??):101–104, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300750.

Goncalves:2010:SGT

[GML10] E. Gonçalves and N. Mendes-Lopes. On the structure of generalized threshold arch processes. Statistics & Probability Letters, 80(7–8):573–580, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004702.

Gut:2015:ETS

[GML15] Allan Gut and Anders Martin-Löf. Extreme-trimmed St. Petersburg games. Statistics & Probability Letters, 96(??):341– 345, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003174.

Gupta:2010:CBD

[GMM10] Ankit Gupta, Kishan G. Mehrotra, and Chilukuri Mohan. A clustering-based discretization for supervised learning. Statistics & Probability Letters, 80(9–10):816–824, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000271.

Giuliano:2019:LDW

[GMP19] Rita Giuliano, Claudio Macci, and Barbara Pacchiarotti. Large deviations for weighted means of random vectors defined in terms of suitable Lévy processes. Statistics & Probability Letters, 150(??):13–22, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300616.

Genest:2018:NBP

[GMS18] Christian Genest, Mhamed Mesfioui, and Juliana Schulz. A new bivariate Poisson common shock model covering all possible degrees of dependence. Statistics & Probability Letters, 140(??):202–209, September 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301676.

Gnacadja:2012:AEC

[Gna12] Gilles Gnacadja. Asymptotic equidistribution of congruence classes with respect to the convolution iterates of a probability vector. Statistics & Probability Letters, 82(10):1849–1852, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002003.

Gao:2012:TSD

[GO12] Jinguo Gao and Omer Ozturk. Two sample distribution-free inference based on partially rank-ordered set samples. Statistics & Probability Letters, 82(5):876–884, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000314.

Goia:2012:FLM

[Goi12] Aldo Goia. A functional linear model for time series prediction with exogenous variables. Statistics & Probability Letters, 82(5):1005–1011, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000569.

Gordon:2012:SUB

[Gor12] Alexander Y. Gordon. A sharp upper bound for the expected number of false rejections. Statistics & Probability Letters, 82(8):1507–1514, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000806.

Gordon:2014:SSM

[Gor14] Alexander Y. Gordon. Smoothing of stepwise multiple testing procedures. Statistics & Probability Letters, 87(??):149– 157, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213004173.

Goudie:2015:BST

[Gou15] I. B. J. Goudie. Bayesian sequential tests of the initial size of a linear pure death process. Statistics & Probability Letters, 100(??):176–181, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521500067X.

Govindan:2014:WCP

[Gov14] T. E. Govindan. Weak convergence of probability measures of Yosida approximate mild solutions of neutral SPDEs. Statistics & Probability Letters, 95(??):26–32, ???? 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002612.

Govindan:2015:TKA

[Gov15] T. E. Govindan. On Trotter–Kato approximations of semilinear stochastic evolution equations in infinite dimensions. Statistics & Probability Letters, 96(??):299–306, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003563.

Ghiglietti:2017:ETM

[GP17] Andrea Ghiglietti and Anna Maria Paganoni. Exact tests for the means of Gaussian stochastic processes. Statistics & Probability Letters, 131(??):102–107, December 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302614.

Gerke:2010:MRT

[GR10] Travis A. Gerke and Ronald H. Randles. A method for resolving ties in asymptotic relative efficiency. Statistics & Probability Letters, 80(13–14):1065–1069, July 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000684.

Giudici:2011:GMD

[GR11] P. Giudici and E. Raffinetti. On the Gini measure decomposition. Statistics & Probability Letters, 81(1):133–139, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002816.

Gajek:2017:GGI

[GR17] Leslaw Gajek and Marcin Rud’z. A generalization of gerber’s inequality for ruin probabilities in risk-switching models. Statistics & Probability Letters, 129(??):236–240, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302055.

Graham:2011:CMC

[Gra11] Carl Graham. Convergence of multi-class systems of fixed possibly infinite sizes. Statistics & Probability Letters, 81(1):31– 35, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002658.

Grevstad:2011:SEN

[Gre11] Nels Grevstad. Simultaneous estimation of negative binomial dispersion parameters. Statistics & Probability Letters, 81(12):1751–1755, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002513.

Grendar:2012:VGM

[Gre12] M. Grendár. Is the p-value a good measure of evidence? Asymptotic consistency criteria. Statistics & Probability Letters, 82(6):1116–1119, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200065X.

Grigelionis:2011:HSG

[Gri11] Bronius Grigelionis. On the Hougaard subordinated Gaussian Lévy processes. Statistics & Probability Letters, 81(8):998–1002, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000356.

Griffin:2019:TPE

[Gri19] Jim Griffin. Two part envelopes for rejection sampling of some completely random measures. Statistics & Probability Letters, 151(??):36–41, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521930080X.

Gut:2010:AIS

[GS10] Allan Gut and Josef Steinebach. Asymptotics for increments of stopped renewal processes. Statistics & Probability Letters, 80(7–8):558–565, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771520900460X.

Gut:2011:IBK

[GS11] Allan Gut and Ulrich Stadtmüller. An intermediate Baum– Katz theorem. Statistics & Probability Letters, 81(10):1486– 1492, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001830.

Grow:2012:QVB

[GS12] David Grow and Suman Sanyal. The quadratic variation of Brownian motion on a time scale. Statistics & Probability Letters, 82(9):1677–1680, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001836.

Ginovyan:2013:TAP

[GS13a] Mamikon S. Ginovyan and Artur A. Sahakyan. On the trace approximations of products of Toeplitz matrices. Statistics & Probability Letters, 83(3):753–760, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200435X.

Gut:2013:RSR

[GS13b] Allan Gut and Ulrich Stadtmüller. Records in subsets of a random field. Statistics & Probability Letters, 83(3):689–699, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004312.

Geng:2016:PEL

[GS16] Pei Geng and Lyudmila Sakhanenko. Parameter estimation for the logistic regression model under case-control study. Statistics & Probability Letters, 109:168–177, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003867.

Gapeev:2017:LTF

[GS17a] Pavel V. Gapeev and Yavor I. Stoev. On the Laplace transforms of the first exit times in one-dimensional non-affine jumpdiffusion models. Statistics & Probability Letters, 121(??):152– 162, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215301048.

Gut:2017:SLS

[GS17b] Allan Gut and Ulrich Stadtmüller. Strong laws for sequences in the vicinity of the LIL. Statistics & Probability Letters, 122(??): 63–72, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216302267.

Ginestet:2012:WFM

[GSK12] Cedric E. Ginestet, Andrew Simmons, and Eric D. Kolaczyk. Weighted Frechet means as convex combinations in metric spaces: Properties and generalized median inequalities. Statistics & Probability Letters, 82(10):1859–1863, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002106.

Guessoum:2012:NLB

[GSST12] Zohra Guessoum, Elias Ould Saïd, Ourida Sadki, and Abdelkader Tatachak. A note on the Lynden–Bell estimator under association. Statistics & Probability Letters, 82(11):1994–2000, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002271.

Greenwood:2011:EIA

[GSW11] Priscilla E. Greenwood, Anton Schick, and Wolfgang Wefelmeyer. Estimating the inter-arrival time density of Markov renewal processes under structural assumptions on the transition distribution. Statistics & Probability Letters, 81(2):277– 282, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002932.

Gorgens:2014:BCM

[GT14] Maik Görgens and Måns Thulin. Bias-correction of the maximum likelihood estimator for the α-Brownian bridge. Statistics & Probability Letters, 93(??):78–86, October 1–142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002223.

Gzyl:2016:RDT

[GT16] H. Gzyl and A. Tagliani. Recovering a distribution from its translated fractional moments. Statistics & Probability Letters, 118(??):171–176, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301055.

Gusev:2012:CIM

[Gus12a] Andrey L. Gusev. Continuous inspection with memory. Statistics & Probability Letters, 82(2):303–307, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003269.

Gusev:2012:ONI

[Gus12b] Andrey L. Gusev. The optimal number of items in a group for group testing. Statistics & Probability Letters, 82(12):2083–2085, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002817.

Gut:2011:RTT

[Gut11] Allan Gut. Renewal theory with a trend. Statistics & Probability Letters, 81(8):1292–1299, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001222.

Gut:2014:CRI

[Gut14] Allan Gut. On convergence of randomly indexed sequences; a counterexample based on the St. Petersburg game. Statistics & Probability Letters, 87(??):105–107, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000224.

Ghosh:2016:OBS

[GV16] S. Ghosh and P. Vellaisamy. On the occurrence of boundary solutions in multidimensional incomplete tables. Statistics & Probability Letters, 119(??):63–75, December 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301262.

Geng:2012:CPM

[GW12a] Xianmin Geng and Ying Wang. The compound Pascal model with dividends paid under random interest. Statistics & Probability Letters, 82(7):1331–1336, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001307.

Gyorfi:2012:SCD

[GW12b] László Györfi and Harro Walk. Strongly consistent density estimation of the regression residual. Statistics & Probability Letters, 82(11):1923–1929, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002556.

Gyorfi:2012:SCN

[GW12c] László Györfi and Harro Walk. Strongly consistent nonparametric tests of conditional independence. Statistics & Probability Letters, 82(6):1145–1150, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000703.

Gao:2015:LDP

[GW15] Zhenlong Gao and Weigang Wang. Large deviations for a Poisson random indexed branching process. Statistics & Probability Letters, 105(??):143–148, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001996.

Gao:2016:LMD

[GW16] Zhenlong Gao and Weigang Wang. Large and moderate deviations for a renewal randomly indexed branching process.

Statistics & Probability Letters, 116(??):139–145, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302686.

Gajda:2017:GFL

[GWK17] J. Gajda, A. Wylomańska, and A. Kumar. Generalized fractional Laplace motion. Statistics & Probability Letters, 124(??):101– 109, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300366.

Gan:2015:SMC

[GX15] H. L. Gan and A. Xia. Stein’s method for conditional compound Poisson approximation. Statistics & Probability Letters, 100(??): 19–26, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000486.

Guo:2017:CBP

[GXH17] Donglin Guo, Liugen Xue, and Yuqin Hu. Covariate-balancing-propensity-score-based inference for linear models with missing responses. Statistics & Probability Letters, 123(??):139– 145, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216302851.

Guo:2013:ECC

[GZ13] Mingle Guo and Dongjin Zhu. Equivalent conditions of complete moment convergence of weighted sums for ρ^{+}-mixing sequence of random variables. Statistics & Probability Letters, 83(1):13– 20, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003203.

Guo:2014:FLT

[GZ14] Hongsong Guo and Mei Zhang. A fluctuation limit theorem for a critical branching process with dependent immigration. Statistics & Probability Letters, 94(??):29–38, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002430.

Gao:2015:LMD

[GZ15] Zhenlong Gao and Yanhua Zhang. Large and moderate deviations for a class of renewal random indexed branching process. Statistics & Probability Letters, 103(??):1–5, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001121.

Gillard:2018:ODS

[GZ18] Jonathan Gillard and Anatoly Zhigljavsky. Optimal directional statistic for general regression. Statistics & Probability Letters, 143(??):74–80, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302694.

Ghosh:2019:NSP

[GZSZ19] Malay Ghosh, Xiaolong Zhong, Ashis SenGupta, and Ruoyang Zhang. Non-subjective priors for wrapped Cauchy distributions. Statistics & Probability Letters, 153(??):90–97, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301452.

Guo:2015:MSE

[GZWW15] Xiao Guo, Hai Zhang, Yao Wang, and Jiang-Lun Wu. Model selection and estimation in high dimensional regression models with group SCAD. Statistics & Probability Letters, 103(??):86– 92, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001285.

Gzyl:2019:HSB

[Gzy19] Henryk Gzyl. Hitting spheres with Brownian motion revisited. Statistics & Probability Letters, 154(??):Article 108565, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302111.

He:2012:CDT

[HA12] Yuanzhen He and Mingyao Ai. Complementary design theory for sliced equidistance designs. Statistics & Probability Letters, 82(3):542–547, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003610.

Harandi:2013:ASL

[HA13] S. Shams Harandi and M. H. Alamatsaz. Alpha-skew-Laplace distribution. Statistics & Probability Letters, 83(3):774–782, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004403.

Habiger:2014:ORC

[HA14] Joshua D. Habiger and Akim Adekpedjou. Optimal rejection curves for exact false discovery rate control. Statistics & Probability Letters, 94(??):21–28, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521400248X.

Hong:2015:MLL

[HA15] Jyy-I Hong and K. B. Athreya. Markov limit of line of decent types in a multitype supercritical branching process. Statistics & Probability Letters, 98(??):54–58, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003824.

Hadjikyriakou:2011:MZI

[Had11] Milto Hadjikyriakou. Marcinkiewicz–Zygmund inequality for nonnegative N-demimartingales and related results. Statistics & Probability Letters, 81(6):678–684, June 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000605.

Hadjikyriakou:2013:CCE

[Had13] Milto Hadjikyriakou. Comparison of conditional expectations of functions of strong N-demimartingales and functions of sums of conditionally independent random variables. Statistics & Probability Letters, 83(4):1282–1286, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004907.

Hadjikyriakou:2017:NAS

[Had17] Milto Hadjikyriakou. Normal approximation for strong demimartingales. Statistics & Probability Letters, 122(??):104–108, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521630236X.

Hadjicostas:2019:GAC

[Had19] Petros Hadjicostas. Generalizations of the arithmetic case of Blackwell’s renewal theorem. Statistics & Probability Letters, 149(??):124–131, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300379.

Hanagal:2010:MHB

[Han10] David D. Hanagal. Modeling heterogeneity for bivariate survival data by the compound Poisson distribution with random scale. Statistics & Probability Letters, 80(23–24):1781–1790, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002257.

Han:2012:NIE

[Han12a] Peisong Han. A note on improving the efficiency of inverse probability weighted estimator using the augmentation term. Statistics & Probability Letters, 82(12):2221–2228, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003008.

Hanck:2012:ADU

[Han12b] Christoph Hanck. On the asymptotic distribution of a unit root test against ESTAR alternatives. Statistics & Probability Letters, 82(2):360–364, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003543.

Hansen:2018:SUR

[Han18] Niels Richard Hansen. On Stein’s unbiased risk estimate for reduced rank estimators. Statistics & Probability Letters, 135(??): 76–82, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303607.

Hart:2016:NTS

[Har16] Jeffrey D. Hart. A nonparametric test of stationarity for independent data. Statistics & Probability Letters, 108:40–44, January 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215301292.

Hashorva:2010:RDI

[Has10] Enkelejd Hashorva. On the residual dependence index of elliptical distributions. Statistics & Probability Letters, 80(13– 14):1070–1078, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000702.

Hattori:2012:TNT

[Hat12] Satoshi Hattori. Testing the no-treatment effect based on a possibly misspecified accelerated failure time model. Statistics & Probability Letters, 82(2):371–377, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003373.

Hayashi:2012:SEB

[Hay12] Kenichi Hayashi. A simple extension of boosting for asymmetric mislabeled data. Statistics & Probability Letters, 82(2):348–356, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521100335X.

Hazelton:2011:ALC

[Haz11] Martin L. Hazelton. Assessing log-concavity of multivariate densities. Statistics & Probability Letters, 81(1):121–125, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002774.

Hwang:2010:LMD

[HB10] S. Y. Hwang and J. S. Baek. Limiting mixture distributions for AR(1) model indexed by a branching process. Statistics & Probability Letters, 80(23–24):2003–2008, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002567.

Hwang:2011:GEF

[HB11] S. Y. Hwang and I. V. Basawa. Godambe estimating functions and asymptotic optimal inference. Statistics & Probability Letters, 81(8):1121–1127, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000897.

Hong:2018:NES

[HB18] Phan Thanh Hong and Cao Tan Binh. A note on exponential stability of non-autonomous linear stochastic differential delay equations driven by a fractional Brownian motion with Hurst index >12. Statistics & Probability Letters, 138(??):127–136, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301093.

Hu:2014:LLG

[HBF14] Zhishui Hu, Wei Bi, and Qunqiang Feng. Limit laws in the generalized random graphs with random vertex weights. Statistics & Probability Letters, 89(??):65–76, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000844.

Hlynka:2010:MOL

[HBH10] M. Hlynka, P. H. Brill, and W. Horn. A method for obtaining Laplace transforms of order statistics of Erlang random variables. Statistics & Probability Letters, 80(1):9–18, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003472.

Hajjem:2011:MER

[HBL11] Ahlem Hajjem, François Bellavance, and Denis Larocque. Mixed effects regression trees for clustered data. Statistics & Probability Letters, 81(4):451–459, April 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003433.

Hwang:2010:EVT

[HBPC10] S. Y. Hwang, J. S. Baek, J. A. Park, and M. S. Choi. Explosive volatilities for threshold-GARCH processes generated by asymmetric innovations. Statistics & Probability Letters, 80(1):26– 33, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003496.

Hansen:2010:PML

[HC10a] Elizabeth Hansen and Kung-Sik Chan. Penalized maximum likelihood estimation of a stochastic multivariate regression model. Statistics & Probability Letters, 80(21–22):1643–1649, November 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210001999.

Hu:2010:GPE

[HC10b] Feng Hu and Zengjing Chen. Generalized Peng’s g-expectations and related properties. Statistics & Probability Letters, 80(3– 4):191–195, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003927.

Huang:2017:MEA

[HC17] Weihong Huang and Yuguo Chen. The multiset EM algorithm. Statistics & Probability Letters, 126(??):41–48, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300779.

Hedayat:2019:EUE

[HCP19] A. S. Hedayat, Hansheng Cheng, and Jennifer Pajda-De La O. Existence of unbiased estimation for the minimum, maximum, and median in finite population sampling. Statistics & Probability Letters, 153(??):192–195, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301403.

He:2016:EEC

[HCT16] Fengyang He, Yebin Cheng, and Tiejun Tong. Estimation of extreme conditional quantiles through an extrapolation of intermediate regression quantiles. Statistics & Probability Letters, 113(??):30–37, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000249.

He:2013:ALB

[HCW13] Wei He, Dongya Cheng, and Yuebao Wang. Asymptotic lower bounds of precise large deviations with nonnegative and dependent random variables. Statistics & Probability Letters, 83(1):331–338, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003653.

Cai:2015:NCT

[hCyP15] Guang hui Cai and Xue yan Pan. A note on the Chover-type law of the iterated logarithm for the weighted partial sums of αmixing sequences. Statistics & Probability Letters, 107:150–156, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002928.

Hanagal:2013:CNB

[HD13] David D. Hanagal and Alok D. Dabade. Compound negative binomial shared frailty models for bivariate survival data. Statistics & Probability Letters, 83(11):2507–2515, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521300254X.

Hanson:2017:BPA

[HdCC17] Timothy E. Hanson, Miguel de Carvalho, and Yuhui Chen. Bernstein polynomial angular densities of multivariate extreme value distributions. Statistics & Probability Letters, 128(??):60–66, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301311.

He:2012:ERL

[He12] Jianjun He. An estimate of the remainder of a limit theorem. Statistics & Probability Letters, 82(3):478–487, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003622.

He:2014:OCR

[He14] Fangchao He. Optimal convergence rates of high order Parzen windows with unbounded sampling. Statistics & Probability Letters, 92(??):26–32, September 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001631.

He:2016:NMO

[He16] Xin He. A note on the maximal outdegrees of Galton– Watson trees. Statistics & Probability Letters, 109:1–6, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521500365X.

He:2018:ODM

[He18] Lei He. Optimal designs for multi-factor nonlinear models based on the second-order least squares estimator. Statistics & Probability Letters, 137(??):201–208, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300075.

Heinrich:2018:BMP

[Hei18] Lothar Heinrich. Brillinger-mixing point processes need not to be ergodic. Statistics & Probability Letters, 138(??):31–35, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300749.

Helali:2019:CRG

[Hel19] Amine Helali. The convergence rate of the Gibbs sampler for generalized 1-D Ising model. Statistics & Probability Letters, 155(??):Article 108555, December 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301932.

Hong:2010:CPS

[HEM10] Yili Hong, Luis A. Escobar, and William Q. Meeker. Coverage probabilities of simultaneous confidence bands and regions for log-location-scale distributions. Statistics & Probability Letters, 80(7–8):733–738, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000088.

Hew:2017:ADR

[Hew17] Patrick Chisan Hew. Asymptotic distribution of rewards accumulated by alternating renewal processes. Statistics & Probability Letters, 129(??):355–359, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302316.

Hamza:2013:BAC

[HH13a] Marwa Hamza and Abdelhamid Hassairi. Bayesian approach to cubic natural exponential families. Statistics & Probability Letters, 83(9):1946–1955, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001417.

Hou:2013:ISP

[HH13b] Chia-Ding Hou and Sheng Huang. Identifying the source of proportion shifts in a multinomial process using a simple statistical test procedure. Statistics & Probability Letters, 83(4):1100– 1105, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000047.

Hong:2016:DBV

[HH16] Won-Tak Hong and Eunju Hwang. Dynamic behavior of volatility in a nonstationary generalized regime-switching GARCH model. Statistics & Probability Letters, 115(??):36–44, August 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303503.

Hsieh:2018:ESF

[HH18] Jin-Jian Hsieh and Chia-Hao Hsu. Estimation of the survival function with redistribution algorithm under semi-competing risks data. Statistics & Probability Letters, 132(??):1–6, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302894.

Hamdan:2019:ULH

[HH19] Mustafa Hamdan and Xiaomi Hu. Unbiasedness of LRT on the homogeneity of means in MANOVA under general order restrictions. Statistics & Probability Letters, 148(??):59–65, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300100.

He:2017:BEL

[HHF17] Bang-Qiang He, Xing-Jian Hong, and Guo-Liang Fan. Block empirical likelihood for partially linear panel data models with fixed effects. Statistics & Probability Letters, 123(??):128–138, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302747.

Hirakawa:2011:SSC

[HHY11] Akihiro Hirakawa, Chikuma Hamada, and Isao Yoshimura. Sample size calculation for a regularized t-statistic in microarray experiments. Statistics & Probability Letters, 81(7):870–875, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000642.

Hieber:2014:FPT

[Hie14] Peter Hieber. First-passage times of regime switching models. Statistics & Probability Letters, 92(??):148–157, September 1– 262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001941.

Hildebrand:2014:RPR

[Hil14] Martin Hildebrand. A random process related to a random walk on upper triangular matrices over a finite field. Statistics & Probability Letters, 89(??):77–80, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000832.

Hildebrand:2019:LBC

[Hil19] Martin Hildebrand. On a lower bound for the Chung–Diaconis– Graham random process. Statistics & Probability Letters, 152 (??):121–125, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301233.

Hirao:2011:LDR

[Hir11] Masatake Hirao. Large deviations for the radial processes of the Brownian motions on hyperbolic spaces. Statistics & Probability Letters, 81(11):1561–1564, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002070.

Higbee:2019:ALL

[HJM19] Joshua D. Higbee, Jonathan E. Jensen, and James B. Mc-Donald. The asymmetric log-Laplace distribution as a limiting case of the generalized beta distribution. Statistics & Probability Letters, 151(??):73–78, August 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521930094X.

Herbertsson:2011:PBD

[HJS11] Alexander Herbertsson, Jiwook Jang, and Thorsten Schmidt. Pricing basket default swaps in a tractable shot noise model. Statistics & Probability Letters, 81(8):1196–1207, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001015.

Hillion:2019:EPN

[HJS19] Erwan Hillion, Oliver Johnson, and Adrien Saumard. An extremal property of the normal distribution, with a discrete analog. Statistics & Probability Letters, 145(??):181–186, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302931.

Hashorva:2014:TAR

[HK14] Enkelejd Hashorva and Dominik Kortschak. Tail asymptotics of random sum and maximum of log-normal risks. Statistics & Probability Letters, 87(??):167–174, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000297.

Hayashi:2019:NAL

[HK19] Takaki Hayashi and Yuta Koike. No arbitrage and leadlag relationships. Statistics & Probability Letters, 154(??): Article 108530, November 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301646.

Ha:2011:NQA

[HL11] Jeongcheol Ha and Taewook Lee. NM–QELE for ARMA– GARCH models with non-Gaussian innovations. Statistics & Probability Letters, 81(6):694–703, June 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000423.

He:2012:MTS

[HL12] Hui He and Nana Luan. A martingale transformation for superprocesses. Statistics & Probability Letters, 82(6):1082– 1087, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000739.

Herve:2013:LLT

[HL13a] Loïc Hervé and James Ledoux. A local limit theorem for densities of the additive component of a finite Markov Additive Process. Statistics & Probability Letters, 83(9):2119–2128, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002009.

Hu:2013:VSP

[HL13b] Yuao Hu and Heng Lian. Variable selection in a partially linear proportional hazards model with a diverging dimensionality. Statistics & Probability Letters, 83(1):61–69, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521200329X.

Herve:2016:CBE

[HL16] Loïc Hervé and James Ledoux. A computable bound of the essential spectral radius of finite range Metropolis–Hastings kernels. Statistics & Probability Letters, 117(??):72–79, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300608.

Helali:2019:HTC

[HL19] Amine Helali and Matthias Löwe. Hitting times, commute times, and cover times for random walks on random hypergraphs. Statistics & Probability Letters, 155(??):Article 108535, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301695.

Hajjem:2017:GME

[HLB17] Ahlem Hajjem, Denis Larocque, and François Bellavance. Generalized mixed effects regression trees. Statistics & Probability Letters, 126(??):114–118, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300895.

Huang:2019:SER

[HLCH19] Lu-Jing Huang, Yin-Ting Liao, Lo-Bin Chang, and Chii-Ruey Hwang. The smallest eigenvalues of random kernel matrices: Asymptotic results on the min kernel. Statistics & Probability Letters, 148(??):23–29, May 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218304024.

Hong:2013:PPP

[HLLK13] Yicheng Hong, Chaehun Lee, Sungchul Lee, and Hyungsoo Kim. Poisson point processes with detection and rest. Statistics & Probability Letters, 83(7):1805–1811, July 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001272.

Hao:2016:TVP

[HLM16] Chengcheng Hao, Yuli Liang, and Thomas Mathew. Testing variance parameters in models with a Kronecker product covariance structure. Statistics & Probability Letters, 118(??):182–189, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215300389.

Hao:2015:EBV

[HLR15] Chengcheng Hao, Yuli Liang, and Anuradha Roy. Equivalency between vertices and centers-coupled-with-radii principal component analyses for interval data. Statistics & Probability Letters, 106(??):113–120, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002382.

Heusel:2015:CAM

[HLV15] Judith Heusel, Matthias Löwe, and Franck Vermet. On the capacity of an associative memory model based on neural cliques. Statistics & Probability Letters, 106(??):256–261, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002679.

Huang:2010:RFB

[HLW10] Zongyuan Huang, Jean-Pierre Lepeltier, and Zhen Wu. Reflected forward-backward stochastic differential equations with continuous monotone coefficients. Statistics & Probability Letters, 80(21–22):1569–1576, November 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001665.

Hafidi:2010:KIC

[HM10a] Bezza Hafidi and Abdallah Mkhadri. The Kullback information criterion for mixture regression models. Statistics & Probability Letters, 80(9–10):807–815, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521000026X.

Hudson:2010:ACB

[HM10b] Malcolm Hudson and Jun Ma. On asymptotic convergence of the block-iterative Fisher scoring algorithm. Statistics & Probability Letters, 80(11–12):922–925, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000416.

Huang:2013:PCI

[HM13] Kai Huang and Jie Mi. Properties and computation of interval availability of system. Statistics & Probability Letters, 83(5):1388–1396, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000199.

Hofert:2015:SVR

[HM15] Marius Hofert and Alexander J. McNeil. Subadditivity of Value-at-Risk for Bernoulli random variables. Statistics & Probability Letters, 98(??):79–88, March 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004192.

Huang:2014:WCM

[HMS14] Gang Huang, Michel Mandjes, and Peter Spreij. Weak convergence of Markov-modulated diffusion processes with rapid switching. Statistics & Probability Letters, 86(??):74–79, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213004124.

Habiger:2013:OCS

[HMT13] Joshua D. Habiger, Melinda H. McCann, and Joshua M. Tebbs. On optimal confidence sets for parameters in discrete distributions. Statistics & Probability Letters, 83(1):297–303, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003641.

Huang:2011:EDO

[HMZ11] Jianhui Huang, Chunhua Ma, and Cai Zhu. Estimation for discretely observed continuous state branching processes with immigration. Statistics & Probability Letters, 81(8):1104–1111, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000873.

Hu:2010:PEF

[HN10] Yaozhong Hu and David Nualart. Parameter estimation for fractional Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 80(11–12):1030–1038, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000659.

Hazra:2015:NWS

[HN15] Nil Kamal Hazra and Asok K. Nanda. A note on warm standby system. Statistics & Probability Letters, 106(??):30–38, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002369.

Hirsch:2015:APE

[HNGS15] Christian Hirsch, David Neuhäuser, Catherine Gloaguen, and Volker Schmidt. Asymptotic properties of Euclidean shortest-path trees in random geometric graphs. Statistics & Probability Letters, 107:122–130, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002916.

Hatjispyros:2019:DRR

[HNW19] Spyridon J. Hatjispyros, Theodoros Nicoleris, and Stephen G. Walker. Distributional results relating to the posterior of a Dirichlet process prior. Statistics & Probability Letters, 149(??): 146–152, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300446.

Hobson:2013:FEB

[Hob13] David G. Hobson. Fake exponential Brownian motion. Statistics & Probability Letters, 83(10):2386–2390, October 2013.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002423.

Hofmann:2013:HPM

[Hof13] Martin Hofmann. On the hitting probability of max-stable processes. Statistics & Probability Letters, 83(11):2516–2521, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002617.

Homei:2012:RWA

[Hom12] H. Homei. Randomly weighted averages with beta random proportions. Statistics & Probability Letters, 82(8):1515–1520, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001605.

Horbacz:2016:SLL

[Hor16] Katarzyna Horbacz. Strong law of large numbers for continuous random dynamical systems. Statistics & Probability Letters, 118(??):70–79, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300955.

Haye:2011:MDE

[HP11] Mohamedou Ould Haye and Anne Philippe. Marginal density estimation for linear processes with cyclical long memory. Statistics & Probability Letters, 81(9):1354–1364, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001490.

Henderson:2012:CHO

[HP12a] Daniel J. Henderson and Christopher F. Parmeter. Canonical higher-order kernels for density derivative estimation. Statistics & Probability Letters, 82(7):1383–1387, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000855.

Henderson:2012:NRB

[HP12b] Daniel J. Henderson and Christopher F. Parmeter. Normal reference bandwidths for the general order, multivariate kernel density derivative estimator. Statistics & Probability Letters, 82(12):2198–2205, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002921.

Haiman:2013:ODS

[HP13] George Haiman and Cristian Preda. One dimensional scan statistics generated by some dependent stationary sequences. Statistics & Probability Letters, 83(5):1457–1463, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000540.

Hanagal:2014:GSF

[HP14] David D. Hanagal and Arvind Pandey. Gamma shared frailty model based on reversed hazard rate for bivariate survival data. Statistics & Probability Letters, 88(??):190–196, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000583.

Huh:2015:TIE

[HP15] Jib Huh and Cheolwoo Park. Theoretical investigation of an exploratory approach for log-density in scale-space. Statistics & Probability Letters, 107:272–279, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003235.

Hackl:2018:NPG

[HP18a]  Benjamin Hackl and Helmut Prodinger.  The necklace

process: a generating function approach. Statistics & Probability Letters, 142(??):57–61, November 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302335.

Harman:2018:COE

[HP18b] Radoslav Harman and Maryna Prus. Computing optimal experimental designs with respect to a compound Bayes risk criterion. Statistics & Probability Letters, 137(??):135–141, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300282.

Hua:2019:ABT

[HPP19] Lei Hua, Alan Polansky, and Paramahansa Pramanik. Assessing bivariate tail non-exchangeable dependence. Statistics & Probability Letters, 155(??):Article 108556, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301944.

Hashorva:2015:MTA

[HPW15] Enkelejd Hashorva, Liang Peng, and Zhichao Weng. Maxima of a triangular array of multivariate Gaussian sequence. Statistics & Probability Letters, 103(??):62–72, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001170.

Hu:2011:NVP

[HR11] Tien-Chung Hu and Andrew Rosalsky. A note on the de La Vallée Poussin criterion for uniform integrability. Statistics & Probability Letters, 81(1):169–174, January 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002981.

Hu:2015:NRV

[HR15] Tien-Chung Hu and Andrew Rosalsky. A note on random variables with an infinite absolute first moment. Statistics & Probability Letters, 97(??):212–215, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004003.

Harman:2019:RPD

[HR19a] Radoslav Harman and Samuel Rosa. Removal of the points that do not support an E-optimal experimental design. Statistics & Probability Letters, 147(??):83–89, April 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303936.

Hassairi:2019:EPD

[HR19b] A. Hassairi and A. Roula. Exponential probability distribution on symmetric matrices. Statistics & Probability Letters, 145(??):37–42, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302888.

Hodara:2019:EIC

[HRB19] P. Hodara and P. Reynaud-Bouret. Exponential inequality for chaos based on sampling without replacement. Statistics & Probability Letters, 146(??):65–69, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521830350X.

Huang:2010:CBC

[HS10] Wen-Jang Huang and Nan-Cheng Su. Characterizations based on certain regression assumptions of adjacent order statistics. Statistics & Probability Letters, 80(23–24):1700–1704, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521000204X.

Hamza:2011:MAB

[HS11a] K. Hamza and A. W. Sudbury. The mixing advantage for bounded random variables. Statistics & Probability Letters, 81(8):1190–1195, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001003.

Han:2011:NIQ

[HS11b] Peisong Han and Peter X.-K. Song. A note on improving quadratic inference functions using a linear shrinkage approach. Statistics & Probability Letters, 81(3):438–445, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003500.

Hanagal:2012:BEP

[HS12a] David D. Hanagal and Richa Sharma. Bayesian estimation of parameters for the bivariate Gompertz regression model with shared gamma frailty under random censoring. Statistics & Probability Letters, 82(7):1310–1317, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001216.

Hieber:2012:NFP

[HS12b] Peter Hieber and Matthias Scherer. A note on first-passage times of continuously time-changed Brownian motion. Statistics & Probability Letters, 82(1):165–172, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003129.

Hwang:2012:SCS

[HS12c] Eunju Hwang and Dong Wan Shin. Strong consistency of the stationary bootstrap under Ψ-weak dependence. Statistics & Probability Letters, 82(3):488–495, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100383X.

Haaland:2013:EGE

[HS13a] Øystein A. Haaland and Hans J. Skaug. Estimating genotyping error rates from parent-offspring dyads. Statistics & Probability Letters, 83(3):812–819, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004257.

Hwang:2013:SBR

[HS13b] Eunju Hwang and Dong Wan Shin. Stationary bootstrapping realized volatility. Statistics & Probability Letters, 83 (9):2045–2051, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001569.

Hwang:2015:CTS

[HS15] Eunju Hwang and Dong Wan Shin. A CUSUMSQ test for structural breaks in error variance for a long memory heterogeneous autoregressive model. Statistics & Probability Letters, 99(??): 167–176, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500019X.

Haghbin:2016:IDP

[HS16a] H. Haghbin and Z. Shishebor. On infinite dimensional periodically correlated random fields: Spectrum and evolutionary spectra. Statistics & Probability Letters, 110(??):257–267, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003466.

Hallin:2016:EMO

[HS16b] Marc Hallin and Miroslav Siman. Elliptical multiple-output quantile regression and convex optimization. Statistics & Probability Letters, 109:232–237, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521530359X.

Haenggi:2017:UCB

[HS17] M. Haenggi and A. Sarkar. Unique coverage in Boolean models. Statistics & Probability Letters, 123(??):1–7, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302735.

Hirukawa:2019:ABC

[HS19] Masayuki Hirukawa and Mari Sakudo. Another bias correction for asymmetric kernel density estimation with a parametric start. Statistics & Probability Letters, 145(??):158–165, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302955.

Hasegawa:2017:FMP

[HSY17] Ayako Hasegawa, Noriyoshi Sakuma, and Hiroaki Yoshida. Fluctuations of Marchenko–Pastur limit of random matrices with dependent entries. Statistics & Probability Letters, 127(??):85– 96, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301220.

Hansen:2013:NMA

[HT13a] Linda V. Hansen and Thordis L. Thorarinsdottir. A note on moving average models for Gaussian random fields. Statistics & Probability Letters, 83(3):850–855, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004580.

Hashorva:2013:LDS

[HT13b] Enkelejd Hashorva and Zhongquan Tan. Large deviations of Shepp statistics for fractional Brownian motion. Statistics & Probability Letters, 83(10):2242–2247, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002186.

Hajri:2014:IFW

[HT14] Hatem Hajri and Wajdi Touhami. Itô’s formula for Walsh’s Brownian motion and applications. Statistics & Probability Letters, 87(??):48–53, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214000029.

Hien:2015:WLL

[HT15] N. T. T. Hien and L. V. Thanh. On the weak laws of large numbers for sums of negatively associated random vectors in Hilbert spaces. Statistics & Probability Letters, 107:236–245, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500320X.

Hashemi:2010:SPR

[HTA10] Marzieh Hashemi, Mahdi Tavangar, and Majid Asadi. Some properties of the residual lifetime of progressively Type-II right censored order statistics. Statistics & Probability Letters, 80(9–10):848–859, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000313.

Hu:2015:MDP

[Hu15] Wei Hu. Moderate deviation principles for Engel’s, Sylvester’s series and Cantor’s products. Statistics & Probability Letters, 96(??):247–254, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003551.

Hu:2016:SLL

[Hu16] Cheng Hu. A strong law of large numbers for sub-linear expectation under a general moment condition. Statistics & Probability Letters, 119(??):248–258, December 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301602.

Huang:2012:ELP

[Hua12] Zhensheng Huang. Empirical likelihood for the parametric part in partially linear errors-in-function models. Statistics & Probability Letters, 82(1):63–66, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002847.

Huang:2016:SOD

[Hua16] Chien-Hao Huang. On the speed of the one-dimensional polymer in the large range regime. Statistics & Probability Letters, 110(??):8–17, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215301620.

Huh:2012:NER

[Huh12] Jib Huh. Nonparametric estimation of the regression function having a change point in generalized linear models. Statistics & Probability Letters, 82(4):843–851, April 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000193.

Hui:2010:MDE

[Hui10] Jiang Hui. Moderate deviations for estimators of quadratic variational process of diffusion with compound Poisson jumps. Statistics & Probability Letters, 80(17–18):1297–1305, September 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001185.

Hurlimann:2013:MMM

[Hür13] Werner Hürlimann. A moment method for the multivariate asymmetric Laplace distribution. Statistics & Probability Letters, 83(4):1247–1253, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000278.

Hamza:2016:KDT

[HV16] Marwa Hamza and Pierre Vallois. On Kummer’s distribution of type two and a generalized beta distribution. Statistics & Probability Letters, 118(??):60–69, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301267.

Hashorva:2013:LLE

[HW13a] Enkelejd Hashorva and Zhichao Weng. Limit laws for extremes of dependent stationary Gaussian arrays. Statistics & Probability Letters, 83(1):320–330, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200363X.

Huang:2013:IAD

[HW13b] Yufen Huang and Sheng-Wen Wang. Influence analysis on the direction of optimal response. Statistics & Probability Letters, 83(4):1287–1299, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004440.

Hwang:2013:AIV

[Hwa13] S. Y. Hwang. Arbitrary initial values and random norm for explosive AR(1) processes generated by stationary errors. Statistics & Probability Letters, 83(1):127–134, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003276.

Hwang:2018:SOO

[Hwa18] Leng-Cheng Hwang. Second order optimal approximation in a particular exponential family under asymmetric LINEX loss. Statistics & Probability Letters, 137(??):283–291, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300439.

Hwang:2019:NLT

[Hwa19] Eunju Hwang. A note on limit theory for mildly stationary autoregression with a heavy-tailed GARCH error process. Statistics & Probability Letters, 152(??):59–68, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301129.

Harvey:2010:ADD

[HWB10] Danielle J. Harvey, Qian Weng, and Laurel A. Beckett. On an asymptotic distribution of dependent random variables on a 3-dimensional lattice. Statistics & Probability Letters, 80(11–12):1015–1021, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000635.

Heck:2015:TOC

[HWM15] Daniel W. Heck, Eric-Jan Wagenmakers, and Richard D. Morey. Testing order constraints: Qualitative differences between Bayes factors and normalized maximum likelihood. Statistics & Probability Letters, 105(??):157–162, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521500200X.

Hu:2012:SID

[HWYW12] Shuhe Hu, Xinghui Wang, Wenzhi Yang, and Xuejun Wang. Some inequalities for demimartingales and N-demimartingales. Statistics & Probability Letters, 82(2):232–239, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003427.

Huang:2014:EFL

[HWZ14] Lele Huang, Huiwen Wang, and Andi Zheng. The Mestimator for functional linear regression model. Statistics & Probability Letters, 88(??):165–173, May 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000273.

He:2014:ECD

[HX14] Daojiang He and Kai Xu. Estimation of the Cholesky decomposition in a conditional independent normal model with missing data. Statistics & Probability Letters, 88(??):27–39, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000455.

Hirayama:2013:SST

[HY13] Takao Hirayama and Kouji Yano. Strong solutions of Tsirel’son’s equation in discrete time taking values in compact spaces with semigroup action. Statistics & Probability Letters, 83(3):824– 828, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521200449X.

He:2018:MTB

[HY18] Min He and Lin Yu. Martingale transforms between Hardy– Lorentz spaces. Statistics & Probability Letters, 137(??):46– 53, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300142.

Hu:2017:INQ

[HYWT17] Yanan Hu, Yaqi Yang, Chunyu Wang, and Maozai Tian. Imputation in nonparametric quantile regression with complex data. Statistics & Probability Letters, 127(??):120–130, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302164.

He:2016:QED

[HZ16a] Guoman He and Hanjun Zhang. On quasi-ergodic distribution for one-dimensional diffusions. Statistics & Probability Letters, 110(??):175–180, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215301802.

Horta:2016:ISR

[HZ16b] Eduardo Horta and Flavio Ziegelmann. Identifying the spectral representation of Hilbertian time series. Statistics & Probability Letters, 118(??):45–49, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300967.

Huang:2010:ELB

[HZJ^{+}10] Zhensheng Huang, Zhangong Zhou, Rong Jiang, Weimin Qian, and Riquan Zhang. Empirical likelihood based inference for semiparametric varying coefficient partially linear models with error-prone linear covariates. Statistics & Probability Letters, 80(5–6):497–504, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004520.

Huang:2012:SRC

[HZR12] Chunfeng Huang, Haimeng Zhang, and Scott M. Robeson. A simplified representation of the covariance structure of axially symmetric processes on the sphere. Statistics & Probability Letters, 82(7):1346–1351, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001083.

Huang:2016:IRF

[HZR16] Chunfeng Huang, Haimeng Zhang, and Scott M. Robeson. Intrinsic random functions and universal kriging on the circle. Statistics & Probability Letters, 108:33–39, January 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003417.

Huang:2019:IRF

[HZRS19] Chunfeng Huang, Haimeng Zhang, Scott M. Robeson, and Jacob Shields. Intrinsic random functions on the sphere. Statistics & Probability Letters, 146(??):7–14, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303407.

He:2016:DAH

[HZW16] Yong He, Xinsheng Zhang, and Pingping Wang. Discriminant analysis on high dimensional Gaussian copula model. Statistics & Probability Letters, 117(??):100–112, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300694.

He:2019:QED

[HZZ19] Guoman He, Hanjun Zhang, and Yixia Zhu. On the quasiergodic distribution of absorbing Markov processes. Statistics & Probability Letters, 149(??):116–123, June 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300513.

Inoue:2010:CUD

[IA10] Kiyoshi Inoue and Sigeo Aki. On the conditional and unconditional distributions of the number of success runs on a circle with applications. Statistics & Probability Letters, 80(9–10):874–885, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000349.

Iksanov:2017:LIL

[IJB17] Alexander Iksanov, Wissem Jedidi, and Fethi Bouzeffour. A law of the iterated logarithm for the number of occupied boxes in the Bernoulli sieve. Statistics & Probability Letters, 126(??): 244–252, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301177.

Inagaki:2010:MPE

[IK10a] Kazuhisa Inagaki and Fumiyasu Komaki. A modification of profile empirical likelihood for the exponential-tilt model. Statistics & Probability Letters, 80(11–12):997–1004, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000611.

Ishwaran:2010:CRS

[IK10b] Hemant Ishwaran and Udaya B. Kogalur. Consistency of random survival forests. Statistics & Probability Letters, 80(13– 14):1056–1064, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000672.

Igarashi:2014:RFI

[IK14] Gaku Igarashi and Yoshihide Kakizawa. Re-formulation of inverse Gaussian, reciprocal inverse Gaussian, and Birnbaum– Saunders kernel estimators. Statistics & Probability Letters, 84(??):235–246, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003568.

Iwashita:2017:TPU

[IKAH17] Toshiya Iwashita, Bernhard Klar, Moe Amagai, and Hiroki Hashiguchi. A test procedure for uniformity on the Stiefel manifold based on projection. Statistics & Probability Letters, 128(??):89–96, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301669.

Iksanov:2016:WCR

[IKM16] Alexander Iksanov, Zakhar Kabluchko, and Alexander Marynych. Weak convergence of renewal shot noise processes in the case of slowly varying normalization. Statistics & Probability Letters, 114(??):67–77, July 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215300122.

Israelov:2011:CEL

[IL11] Roni Israelov and Steven Lugauer. Combining empirical likelihood and generalized method of moments estimators: Asymptotics and higher order bias. Statistics & Probability Letters, 81(9):1339–1347, September 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001623.

Ip:2017:CVM

[IL17] Ryan H. L. Ip and W. K. Li. A class of valid Matérn crosscovariance functions for multivariate spatio-temporal random fields. Statistics & Probability Letters, 130(??):115–119, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302602.

Ilinskii:2018:NII

[Il’18] Alexander Il’inskii. On notions of Q-independence and Qidentical distributiveness. Statistics & Probability Letters, 140(??):33–36, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301688.

Ilic:2012:TIE

[Ili12] Ivana Ilić. On tail index estimation using a sample with missing observations. Statistics & Probability Letters, 82(5):949– 958, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000247.

Ilmonen:2013:APS

[Ilm13] Pauliina Ilmonen. On asymptotic properties of the scatter matrix based estimates for complex valued independent component analysis. Statistics & Probability Letters, 83(4):1219–1226, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000217.

Ivanovs:2015:TAS

[IM15] Jevgenijs Ivanovs and Michel Mandjes. Transient analysis of a stationary Lévy-driven queue. Statistics & Probability Letters, 107:341–347, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003296.

Iksanov:2016:MCF

[IMM16] Alexander Iksanov, Alexander Marynych, and Matthias Meiners. Moment convergence of first-passage times in renewal theory. Statistics & Probability Letters, 119(??):134–143, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630133X.

Imkeller:2016:NMS

[IMPR16] Peter Imkeller, Thibaut Mastrolia, Dylan Possamaï, and Anthony Réveillac. A note on the Malliavin–Sobolev spaces. Statistics & Probability Letters, 109:45–53, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003119.

Ichifuji:2010:FPM

[IMU10] Ken Ichifuji, Makoto Maejima, and Yohei Ueda. Fixed points of mappings of infinitely divisible distributions on R^{d}. Statistics & Probability Letters, 80(17–18):1320–1328, September 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001215.

Ilmonen:2010:CMD

[INO10] Pauliina Ilmonen, Jaakko Nevalainen, and Hannu Oja. Characteristics of multivariate distributions and the invariant coordinate system. Statistics & Probability Letters, 80(23–24): 1844–1853, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002336.

Iannario:2014:TCM

[IP14a] Maria Iannario and Domenico Piccolo. A theorem on CUB models for rank data. Statistics & Probability Letters, 91(??):27– 31, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001357.

Iksanov:2014:MPR

[IP14b] Alexander Iksanov and Andrey Pilipenko. On the maximum of a perturbed random walk. Statistics & Probability Letters, 92(??):168–172, September 1–262, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002028.

Ishwaran:2011:CSS

[IR11] Hemant Ishwaran and J. Sunil Rao. Consistency of spike and slab regression. Statistics & Probability Letters, 81(12):1920–1928, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002690.

Islak:2016:ARR

[Isl16] Ümit Islak. Asymptotic results for random sums of dependent random variables. Statistics & Probability Letters, 109:22–29, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215300845.

Isozaki:2015:FHT

[Iso15] Yasuki Isozaki. First hitting time of the integer lattice by symmetric stable processes. Statistics & Probability Letters, 98(??): 50–53, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214004167.

Igloi:2014:WBR

[IT14] E. Iglói and Gy. Terdik. When the bispectrum is realvalued. Statistics & Probability Letters, 95(??):1–5, ???? 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002831.

I:2019:NPE

[IV19] Abdul Sathar E. I. and Viswakala K. V. Non-parametric estimation of Kullback–Leibler discrimination information based on censored data. Statistics & Probability Letters, 155(??): Article 108531, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301658.

Izumi:2013:CSO

[Izu13] Yuki Izumi. The L^{p} Cauchy sequence for one-dimensional BS-DEs with linear growth generators. Statistics & Probability Letters, 83(6):1588–1594, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000795.

Jacroux:2010:TLF

[Jac10] Mike Jacroux. Two-level fractional factorial designs in blocks of size two for the orthogonal estimation of all main effects and two-factor interactions. Statistics & Probability Letters, 80(11–12):926–931, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000428.

Jacroux:2011:OON

[Jac11] Mike Jacroux. On the D-optimality of orthogonal and nonorthogonal blocked main effects plans. Statistics & Probability Letters, 81(1):116–120, January 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002762.

Jacroux:2013:NOL

[Jac13] Mike Jacroux. A note on the optimality of 2-level main effects plans in blocks of odd size. Statistics & Probability Letters, 83(4):1163–1166, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000060.

Jalilian:2016:HOP

[Jal16] Abdollah Jalilian. On the higher order product density functions of a Neyman–Scott cluster point process. Statistics & Probability Letters, 117(??):144–150, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300566.

James:2018:SWB

[Jam18] Gareth M. James. Statistics within business in the era of big data. Statistics & Probability Letters, 136(??):155–159, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300798.

Janson:2018:TBS

[Jan18] Svante Janson. Tail bounds for sums of geometric and exponential variables. Statistics & Probability Letters, 135(??):1–6, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303711.

Jaroszewska:2013:AEM

[Jar13] Joanna Jaroszewska. On asymptotic equicontinuity of Markov transition functions. Statistics & Probability Letters, 83(3):943– 951, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004051.

Jasinski:2016:ANN

[Jas16] Krzysztof Jasiński. Asymptotic normality of numbers of observations near order statistics from stationary processes. Statistics & Probability Letters, 119(??):259–263, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301596.

Jasinski:2019:LBR

[Jas19] Krzysztof Jasiński. Limiting behavior of the ratio of k-th records. Statistics & Probability Letters, 150(??):29–34, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300562.

Jamalizadeh:2013:NMD

[JB13a] Ahad Jamalizadeh and N. Balakrishnan. A note on “Maximum distributions for I_{2,p}-symmetric vectors are skewed I_{1,p}-symmetric distributions” by Batún-Cutz et al. (2013). Statistics & Probability Letters, 83(11):2522–2523, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002642.

Janczak-Borkowska:2013:GBD

[JB13b] Katarzyna Jańczak-Borkowska. Generalized BSDEs driven by fractional Brownian motion. Statistics & Probability Letters, 83(3):805–811, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004452.

Jamalizadeh:2019:CDM

[JB19] Ahad Jamalizadeh and Narayanaswamy Balakrishnan. Conditional distributions of multivariate normal mean-variance mixtures. Statistics & Probability Letters, 145(??):312–316, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303213.

Jamalizadeh:2010:OSL

[JBS10] A. Jamalizadeh, N. Balakrishnan, and Mehdi Salehi. Order statistics and linear combination of order statistics arising from a bivariate selection normal distribution. Statistics & Probability Letters, 80(5–6):445–451, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004465.

Jiang:2015:LDS

[JCM15] Tao Jiang, Sheng Cui, and Ruixing Ming. Large deviations for the stochastic present value of aggregate claims in the renewal risk model. Statistics & Probability Letters, 101(??):83–91, June 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521500070X.

Jiang:2017:DSB

[JD17] Zhichao Jiang and Peng Ding. The directions of selection bias. Statistics & Probability Letters, 125(??):104–109, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300445.

Jozani:2012:PCR

[JDB12] Mohammad Jafari Jozani, Katherine F. Davies, and Narayanaswamy Balakrishnan. Pitman closeness results concerning ranked set sampling. Statistics & Probability Letters, 82(12):2260–2269, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003161.

Jorgensen:2011:BCP

[JDK^{+}11] Bent Jørgensen, Clarice G. B. Demétrio, Erik Kristensen, Gary T. Banta, Hans Christian Petersen, and Matthieu Delefosse. Bias-corrected Pearson estimating functions for Taylor’s power law applied to benthic macrofauna data. Statistics & Probability Letters, 81(7):749–758, July 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000125.

Jacome:2016:BSP

[JdU16] M. A. Jácome and I. López de Ullibarri. Bandwidth selection for the presmoothed logrank test. Statistics & Probability Letters, 117(??):151–157, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300670.

Jiang:2014:MSE

[JGW14] Tao Jiang, Qingwu Gao, and Yuebao Wang. Max-sum equivalence of conditionally dependent random variables. Statistics & Probability Letters, 84(??):60–66, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003295.

Johnson:2011:PGR

[JH11] Devin S. Johnson and Jennifer A. Hoeting. Properties of graphical regression models for multidimensional categorical data. Statistics & Probability Letters, 81(10):1471–1475, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001726.

Jung:2014:SPM

[JH14] Yeun Ji Jung and James P. Hobert. Spectral properties of MCMC algorithms for Bayesian linear regression with generalized hyperbolic errors. Statistics & Probability Letters, 95(??):92–100, December 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214002946.

Jiang:2015:IPA

[JHF15] Jia-Jian Jiang, Ping He, and Kai-Tai Fang. An interesting property of the arcsine distribution and its applications. Statistics & Probability Letters, 105(??):88–95, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001820.

Jiao:2018:GMI

[JHW18] Jiantao Jiao, Yanjun Han, and Tsachy Weissman. Generalizations of maximal inequalities to arbitrary selection rules. Statistics & Probability Letters, 137(??):19–25, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830004X.

Jiang:2013:CST

[Jia13] Binyan Jiang. Covariance selection by thresholding the sample correlation matrix. Statistics & Probability Letters, 83 (11):2492–2498, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002575.

Jia:2016:SRE

[Jia16] Chen Jia. A solution to the reversible embedding problem for finite Markov chains. Statistics & Probability Letters, 116 (??):122–130, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215301231.

Jirak:2013:WIP

[Jir13] Moritz Jirak. On weak invariance principles for sums of dependent random functionals. Statistics & Probability Letters, 83(10):2291–2296, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002198.

Jaworski:2013:CWP

[JK13] Piotr Jaworski and Marcin Krzywda. Coupling of Wiener processes by using copulas. Statistics & Probability Letters, 83 (9):2027–2033, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001727.

Jacroux:2014:OBM

[JKD14] Mike Jacroux and Bonni Kealy-Dichone. On the E-optimality of blocked main effects plans when n ≡ 3 mod 4. Statistics & Probability Letters, 87(??):143–148, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000169.

Jacroux:2015:TOB

[JKD15a] Mike Jacroux and Bonni Kealy-Dichone. On the type I optimality of blocked 2-level main effects plans having blocks of different sizes. Statistics & Probability Letters, 98(??):39–43, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521400412X.

Jacroux:2015:UBL

[JKD15b] Mike Jacroux and Bonni Kealy-Dichone. On the use of blocked 2-level main effects plans having blocks of different sizes. Statistics & Probability Letters, 107:362–368, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003363.

Jeon:2019:VBM

[JKP19] Saebom Jeon, Tae Yeon Kwon, and Yousung Park. Variablebased missing mechanism for an incomplete contingency table with unit missingness. Statistics & Probability Letters, 146(??): 90–96, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303535.

Jiang:2016:HDD

[JL16] Binyan Jiang and Chenlei Leng. High dimensional discrimination analysis via a semiparametric model. Statistics & Probability Letters, 110(??):103–110, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003806.

Jin:2016:IBM

[JLJ16] Hua Jin, Song Li, and Yaolan Jin. The IM-based method for testing the non-inferiority of odds ratio in matched-pairs design. Statistics & Probability Letters, 109:145–151, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003843.

Lee:2019:DRW

[jLL19] Myoung jae Lee and Sanghyeok Lee. Double robustness without weighting. Statistics & Probability Letters, 146(??):175– 180, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303705.

Ji:2019:TAB

[JLR19] Lanpeng Ji, Peng Liu, and Stephan Robert. Tail asymptotic behavior of the supremum of a class of chi-square processes. Statistics & Probability Letters, 155(??):Article 108551, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301889.

Jiang:2014:PED

[JLY14] Shuxia Jiang, Yuanyuan Liu, and Shuai Yao. Poisson’s equation for discrete-time single-birth processes. Statistics & Probability Letters, 85(??):78–83, February 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003891.

Jhwueng:2014:POU

[JM14] Dwueng-Chwuan Jhwueng and Vasileios Maroulas. Phylogenetic Ornstein–Uhlenbeck regression curves. Statistics & Probability Letters, 89(??):110–117, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000881.

Jansen:2015:UBV

[JMDW15] H. M. Jansen, M. R. H. Mandjes, K. De Turck, and S. Wittevrongel. On the upper bound in Varadhan’s Lemma. Statistics & Probability Letters, 103(??):24–29, August 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001157.

Jevtic:2017:NMP

[JMS17] Petar Jevtić, Marina Marena, and Patrizia Semeraro. A note on Marked Point Processes and multivariate subordination. Statistics & Probability Letters, 122(??):162–167, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302486.

Jiofack:2010:TLD

[JN10] Jean Gérard Aghoukeng Jiofack and Guy Martial Nkiet. Testing for lack of dependence between functional variables. Statistics & Probability Letters, 80(15–16):1210–1217, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000982.

Johnson:2014:SSB

[Joh14] Torrey Johnson. On the support of the simple branching random walk. Statistics & Probability Letters, 91(??):107–109, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001473.

Jonsson:2010:QMS

[Jon10] Fredrik Jonsson. On the quadratic moment of self-normalized sums. Statistics & Probability Letters, 80(17–18):1289–1296, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001173.

Jones:2012:RBD

[Jon12] M. C. Jones. Relationships between distributions with certain symmetries. Statistics & Probability Letters, 82(9):1737–1744, September 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001897.

Jonasson:2017:SML

[Jon17] Johan Jonasson. Slow mixing for latent Dirichlet allocation. Statistics & Probability Letters, 129(??):96–100, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301864.

Jones:2018:FCD

[Jon18] M. C. Jones. Families of complementary distributions. Statistics & Probability Letters, 141(??):74–81, October 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302086.

Jones:2019:IKR

[Jon19] M. C. Jones. Inverting Khintchine’s relationship and generating length biased data. Statistics & Probability Letters, 155 (??):Article 108539, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301762.

Joutard:2014:AAP

[Jou14] Cyrille Joutard. Asymptotic approximation for the probability density function of an arbitrary sequence of random variables. Statistics & Probability Letters, 90(??):100–107, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001114.

Jovanovski:2014:CBT

[Jov14] Oliver Jovanovski. Convergence bound in total variation for an image restoration model. Statistics & Probability Letters, 90(??): 11–16, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521400100X.

Jovalekic:2019:SER

[Jov19] Milica Jovalekić. Some estimates related to the Doob’s martingale inequalities. Statistics & Probability Letters, 153(??):124– 129, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521930152X.

Jiao:2010:WTI

[JP10] Yong Jiao and Lihua Peng. Weak type inequalities for vectorvalued martingales. Statistics & Probability Letters, 80(13– 14):1128–1135, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000775.

Jaruskova:2011:LLR

[JP11] Daniela Jarusková and Vladimir I. Piterbarg. Log-likelihood ratio test for detecting transient change. Statistics & Probability Letters, 81(5):552–559, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000137.

Jentsch:2012:NUP

[JP12a] Carsten Jentsch and Markus Pauly. A note on using periodogram-based distances for comparing spectral densities. Statistics & Probability Letters, 82(1):158–164, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003087.

Jureckova:2012:RQT

[JP12b] Jana Jurecková and Jan Picek. Regression quantiles and their two-step modifications. Statistics & Probability Letters, 82(6):1111–1115, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000636.

Ju:2016:NPP

[JP16] Shan Ju and Xiaoqing Pan. A new proof for the peakedness of linear combinations of random variables. Statistics & Probability Letters, 114(??):93–98, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215302169.

Jaworski:2017:NCC

[JP17] Piotr Jaworski and Marcin Pitera. A note on conditional covariance matrices for elliptical distributions. Statistics & Probability Letters, 129(??):230–235, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302079.

Jiang:2012:VSC

[JQZ12] Rong Jiang, Weimin Qian, and Zhangong Zhou. Variable selection and coefficient estimation via composite quantile regression with randomly censored data. Statistics & Probability Letters, 82(2):308–317, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003385.

Jureckova:2011:FSD

[JS11] Jana Jurecková and Radka Sabolová. Finite-sample density and its small sample asymptotic approximation. Statistics & Probability Letters, 81(8):1311–1318, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001258.

Jedidi:2015:DHT

[JS15] Wissem Jedidi and Thomas Simon. Diffusion hitting times and the bell-shape. Statistics & Probability Letters, 102(??): 38–41, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000978.

Jakimauskas:2016:NSP

[JS16] Gintautas Jakimauskas and Leonidas Sakalauskas. Note on the singularity of the Poisson-gamma model. Statistics & Probability Letters, 114(??):86–92, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215303825.

Ji:2018:RGU

[JS18] Shaolin Ji and Xiaomin Shi. Reaching goals under ambiguity: Continuous-time optimal portfolio selection. Statistics & Probability Letters, 137(??):63–69, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300129.

Jose:2019:RER

[JS19] Jitto Jose and E. I. Abdul Sathar. Residual extropy of k-record values. Statistics & Probability Letters, 146(??):1–6, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303432.

Joe:2012:MIG

[JSA12] Harry Joe, Vanamamalai Seshadri, and Barry C. Arnold. Multivariate inverse Gaussian and skew-normal densities. Statistics & Probability Letters, 82(12):2244–2251, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002994.

Jiang:2011:PTD

[JT11] Jun Jiang and Qihe Tang. The product of two dependent random variables with regularly varying or rapidly varying tails. Statistics & Probability Letters, 81(8):957–961, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000307.

Jordanger:2014:MSC

[JT14] Lars Arne Jordanger and Dag Tjøstheim. Model selection of copulas: AIC versus a cross validation copula information criterion. Statistics & Probability Letters, 92(??):249–255, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002089.

Jakubowski:2018:QPD

[JT18] Adam Jakubowski and Patryk Truszczyński. Quenched phantom distribution functions for Markov chains. Statistics & Probability Letters, 137(??):79–83, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300117.

Jureckova:2010:FSD

[Jur10] Jana Jurecková. Finite-sample distribution of regression quantiles. Statistics & Probability Letters, 80(23–24):1940–1946, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002464.

Jurek:2013:IMU

[Jur13] Zbigniew J. Jurek. Invariant measures under random integral mappings and marginal distributions of fractional Lévy processes. Statistics & Probability Letters, 83(1):177–183, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003410.

Jurek:2014:RFP

[Jur14] Zbigniew J. Jurek. Remarks on the factorization property of some random integrals. Statistics & Probability Letters, 94(??):192–195, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002594.

Jurek:2018:RCS

[Jur18] Zbigniew J. Jurek. Remarks on compositions of some random integral mappings. Statistics & Probability Letters, 137(??):277– 282, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300373.

Jolis:2010:CHP

[JV10] Maria Jolis and Noèlia Viles. Continuity in the Hurst parameter of the law of the Wiener integral with respect to the fractional Brownian motion. Statistics & Probability Letters, 80(7–8):566–572, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004696.

Jaggi:2010:GIT

[JVVS10] Seema Jaggi, Cini Varghese, Eldho Varghese, and V. K. Sharma. Generalized incomplete Trojan-type designs. Statistics & Probability Letters, 80(7–8):706–710, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000040.

Jiang:2012:SSR

[JW12] L. Jiang and A. C. M. Wong. On standardizing the signed root log likelihood ratio statistic. Statistics & Probability Letters, 82(4):833–839, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200020X.

Jiang:2016:MDP

[JW16] Hui Jiang and Shaochen Wang. Moderate deviation principles for eigenvalues of β-hermite and β-Laguerre ensembles with β → ∞. Statistics & Probability Letters, 118(??):50–59, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300918.

Jiang:2017:EBK

[JW17] Hui Jiang and Shaochen Wang. Error bounds for kernel density estimator of spectral distribution for Gaussian Unitary Ensembles. Statistics & Probability Letters, 126(??):179–184, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301141.

Jiang:2019:ADS

[JWCC19] Tao Jiang, Yuebao Wang, Zhaolei Cui, and Yuxin Chen. On the almost decrease of a subexponential density. Statistics & Probability Letters, 153(??):71–79, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521930149X.

Jiang:2014:CCH

[JWW14] Yiming Jiang, Suxin Wang, and Yongjin Wang. On a class of Cahn–Hilliard type stochastic interacting systems with steppingstone noises. Statistics & Probability Letters, 84(??):9–16, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003180.

Ji:2015:NFD

[JY15a] Shaolin Ji and Shuzhen Yang. A note on functional derivatives on continuous paths. Statistics & Probability Letters, 106(??):176–183, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002655.

Ji:2015:SFF

[JY15b] Shaolin Ji and Shuzhen Yang. Solutions for functional fully coupled forward-backward stochastic differential equations. Statistics & Probability Letters, 99(??):70–76, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000152.

Jiang:2012:DPF

[JYL12] Wuyuan Jiang, Zhaojun Yang, and Xinping Li. The discounted penalty function with multi-layer dividend strategy in the phasetype risk model. Statistics & Probability Letters, 82(7):1358– 1366, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000843.

Kateri:2010:GRM

[KA10a] Maria Kateri and Alan Agresti. A generalized regression model for a binary response. Statistics & Probability Letters, 80(2):89–95, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209003575.

Kortschak:2010:AET

[KA10b] Dominik Kortschak and Hansjörg Albrecher. An asymptotic expansion for the tail of compound sums of Burr distributed random variables. Statistics & Probability Letters, 80(7–8):612–620, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209004763.

Kumar:2011:GMS

[KA11] C. Satheesh Kumar and M. R. Anusree. On a generalized mixture of standard normal and skew normal distributions. Statistics & Probability Letters, 81(12):1813–1821, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002409.

Kabaila:2011:AUC

[Kab11] Paul Kabaila. Admissibility of the usual confidence interval for the normal mean. Statistics & Probability Letters, 81(3):352– 359, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521000338X.

Kabaila:2013:NPD

[Kab13] Paul Kabaila. Note on a paradox in decision-theoretic interval estimation. Statistics & Probability Letters, 83(1):123–126, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003264.

Kabaila:2016:FSP

[Kab16] Paul Kabaila. The finite sample performance of the twostage analysis of a two-period crossover trial. Statistics & Probability Letters, 117(??):118–127, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521630061X.

Kakizawa:2011:CIA

[Kak11a] Yoshihide Kakizawa. Corrigendum to: “Improved additive adjustments for the LR/ELR test statistics” [Statist. Probab. Lett. 81 (2011) 1245–1255]. Statistics & Probability Letters, 81(10): 1559, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001386. See [Kak11b].

Kakizawa:2011:IAA

[Kak11b] Yoshihide Kakizawa. Improved additive adjustments for the LR/ ELR test statistics. Statistics & Probability Letters, 81(8):1245– 1255, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001076. See corrigendum

[Kak11a].

Kakizawa:2012:GCF

[Kak12] Yoshihide Kakizawa. Generalized Cordeiro–Ferrari Bartlett-type adjustment. Statistics & Probability Letters, 82(11):2008–2016, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002568.

Kakizawa:2016:SII

[Kak16] Yoshihide Kakizawa. Some integrals involving multivariate Hermite polynomials: Application to evaluating higher-order local powers. Statistics & Probability Letters, 110(??):162–168, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302091.

Kampf:2017:CLT

[Kam17] Jürgen Kampf. A central limit theorem for Lebesgue integrals of random fields. Statistics & Probability Letters, 124(??):5– 12, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300032.

Kao:2013:OEM

[Kao13] Ming-Hung Kao. On the optimality of extended maximal length linear feedback shift register sequences. Statistics & Probability Letters, 83(6):1479–1483, June 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000552.

Kao:2014:NTE

[Kao14] Ming-Hung Kao. A new type of experimental designs for event-related fMRI via Hadamard matrices. Statistics & Probability Letters, 84(??):108–112, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003222.

Kargin:2011:RTM

[Kar11] Vladislav Kargin. Relaxation time is monotone in temperature in the mean-field Ising model. Statistics & Probability Letters, 81(8):1094–1097, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000848.

Katriel:2013:GRP

[Kat13] Guy Katriel. Gambler’s ruin probability — a general formula. Statistics & Probability Letters, 83(10):2205–2210, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002101.

Katriel:2019:QDC

[Kat19] Guy Katriel. A quantitative discounted central limit theorem using the Fourier metric. Statistics & Probability Letters, 145(??):321–326, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303377.

Kaur:2014:MMA

[Kau14] Inderdeep Kaur. MTD models for aggregate data from higher order Markov chains. Statistics & Probability Letters, 88(??):

157–164, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000510.

Kahrari:2017:SMS

[KAVRY17] F. Kahrari, R. B. Arellano-Valle, M. Rezaei, and F. Yousefzadeh. Scale mixtures of skew-normal-Cauchy distributions. Statistics & Probability Letters, 126(??):1–6, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300834.

Kayal:2015:GDS

[Kay15] Suchandan Kayal. On generalized dynamic survival and failure entropies of order (α, β). Statistics & Probability Letters, 96(??):123–132, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003289.

Kuchibhotla:2015:GSM

[KB15] Arun Kumar Kuchibhotla and Ayanendranath Basu. A general set up for minimum disparity estimation. Statistics & Probability Letters, 96(??):68–74, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003083.

Kitouni:2015:RSC

[KBM15] Abderrahim Kitouni, Mohamed Boukeloua, and Fatiha Messaci. Rate of strong consistency for nonparametric estimators based on twice censored data. Statistics & Probability Letters, 96(??):255–261, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521400354X.

Kundu:2016:OPO

[KC16] Amarjit Kundu and Shovan Chowdhury. Ordering properties of order statistics from heterogeneous exponentiated Weibull models. Statistics & Probability Letters, 114(??):119–127, July 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000596.

Kim:2019:LPC

[KC19] Gunhee Kim and Geon Ho Choe. Limit properties of continuous self-exciting processes. Statistics & Probability Letters, 155(??):Article 108558, December 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219302044.

Kashikar:2012:SOB

[KD12] Akanksha S. Kashikar and S. R. Deshmukh. Second order branching process with continuous state space. Statistics & Probability Letters, 82(11):1930–1934, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002209.

Krnjajic:2014:BMC

[KD14] Milovan Krnjajić and David Draper. Bayesian model comparison: Log scores and DIC. Statistics & Probability Letters, 88(??): 9–14, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000340.

Kong:2016:CRP

[KD16] Lingtao Kong and Hongshuai Dai. Convergence rate in precise asymptotics for Davis law of large numbers. Statistics & Probability Letters, 119(??):295–300, December 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301638.

Kirschenmann:2015:NH

[KDW15] T. H. Kirschenmann, P. Damien, and S. G. Walker. A note on the e-a histogram. Statistics & Probability Letters, 103(??):105– 109, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001339.

Ker:2016:NEP

[Ker16] Alan P. Ker. Nonparametric estimation of possibly similar densities. Statistics & Probability Letters, 117(??):23–30, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302509.

Keiding:2011:AFT

[KFHS11] Niels Keiding, Jason P. Fine, Oluf H. Hansen, and Rémy Slama. Accelerated failure time regression for backward recurrence times and current durations. Statistics & Probability Letters, 81(7):724–729, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000472.

Kim:2018:DMR

[KH18] Yujin Kim and Eunju Hwang. A dynamic Markov regimeswitching GARCH model and its cumulative impulse response function. Statistics & Probability Letters, 139(??):20–30, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301044.

Kharfouchi:2014:IDG

[Kha14] Soumia Kharfouchi. Inference for 2-D GARCH models. Statistics & Probability Letters, 92(??):99–108, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001837.

Khartov:2019:CCQ

[Kha19] A. A. Khartov. Compactness criteria for quasi-infinitely divisible distributions on the integers. Statistics & Probability Letters, 153(??):1–6, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301373.

Kunihama:2016:NBM

[KHHD16] T. Kunihama, A. H. Herring, C. T. Halpern, and D. B. Dunson. Nonparametric Bayes modeling with sample survey weights. Statistics & Probability Letters, 113(??):41–48, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300037.

Khmaladze:2013:ESV

[Khm13] Estáte V. Khmaladze. On the evolution of set-valued functions: An example. Statistics & Probability Letters, 83(3):898– 901, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003884.

Khmaladze:2017:DFT

[Khm17] Estate Khmaladze. Distribution free testing for conditional distributions given covariates. Statistics & Probability Letters, 129(??):348–354, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302304.

Kabera:2012:NCL

[KHN12] M. Gaëtan Kabera, Linda M. Haines, and Principal Ndlovu. A note on the construction of locally D- and D_{s}-optimal designs for the binary logistic model with several explanatory variables. Statistics & Probability Letters, 82(5):865–870, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000338.

Kundu:2016:RSC

[KHN16] Pradip Kundu, Nil Kamal Hazra, and Asok K. Nanda. Reliability study of a coherent system with single general standby component. Statistics & Probability Letters, 110(??):25–33, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215301851.

Kayid:2015:CED

[KI15] M. Kayid and S. Izadkhah. Characterizations of the exponential distribution by the concept of residual life at random time. Statistics & Probability Letters, 107:164–169, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003132.

Kim:2008:NSM

[Kim08] Hyoung-Moon Kim. A note on scale mixtures of skew normal distribution. Statistics & Probability Letters, 78(13):1694–1701, September 15, 2008. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715208000278. See corrigendum

[Kim13].

Kim:2013:CNS

[Kim13] Hyoung-Moon Kim. Corrigendum to “A note on scale mixtures of skew normal distribution” [Statist. Probab. Lett. 78 (2008) 1694–1701]. Statistics & Probability Letters, 83(8):1937, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001491. See [Kim08].

King:2010:NCC

[Kin10] David King. A note on the cross-covariance operator and on congruence relations for Hilbert space valued stochastic processes. Statistics & Probability Letters, 80(5–6):361–365, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004350.

Kinateder:2012:EEI

[Kin12] Kimberly K. J. Kinateder. Expected earnings of invested overflow strategies for M/M/1queue with constrained workload. Statistics & Probability Letters, 82(11):2072–2081, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002283.

Kirsch:2019:EPF

[Kir19] Werner Kirsch. An elementary proof of de Finetti’s theorem. Statistics & Probability Letters, 151(??):84–88, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300902.

Kennedy:2015:ORE

[KJS15] Edward H. Kennedy, Marshall M. Joffe, and Dylan S. Small. Optimal restricted estimation for more efficient longitudinal causal inference. Statistics & Probability Letters, 97(??):185– 191, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003988.

Kim:2011:RDB

[KK11] Bara Kim and Jeongsim Kim. Representation of Downton’s bivariate exponential random vector and its applications. Statistics & Probability Letters, 81(12):1743–1750, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002446.

Kayal:2013:ESE

[KK13a] Suchandan Kayal and Somesh Kumar. Estimation of the Shannon’s entropy of several shifted exponential populations. Statistics & Probability Letters, 83(4):1127–1135, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000138.

Kohler:2013:OGR

[KK13b] Michael Kohler and Adam Krzyzak. Optimal global rates of convergence for interpolation problems with random design. Statistics & Probability Letters, 83(8):1871–1879, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001399.

Kohler:2015:EJP

[KK15]  Michael Kohler and Adam Krzyzak.  Estimation of a

jump point in random design regression. Statistics & Probability Letters, 106(??):247–255, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002424.

Kim:2018:ELP

[KK18] Bara Kim and Jeongsim Kim. Extension of the loss probability formula to an overloaded queue with impatient customers. Statistics & Probability Letters, 134(??):54–62, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303218.

Kim:2019:SWT

[KK19a] Bara Kim and Jeongsim Kim. Sooner waiting time problems in a sequence of multi-state trials with random rewards. Statistics & Probability Letters, 153(??):171–179, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301786.

Kim:2019:LOD

[KK19b] Soohyun Kim and Ming-Hung Kao. Locally optimal designs for mixed binary and continuous responses. Statistics & Probability Letters, 148(??):112–117, May 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300094.

Koldanov:2017:UMP

[KKKP17] Petr Koldanov, Alexander Koldanov, Valeriy Kalyagin, and Panos Pardalos. Uniformly most powerful unbiased test for conditional independence in Gaussian graphical model. Statistics & Probability Letters, 122(??):90–95, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302437.

Kim:2018:SUD

[KKL18] Bara Kim, Jeongsim Kim, and Sungji Lee. Strong unimodality of discrete order statistics. Statistics & Probability Letters, 140(??):48–52, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301755.

Kharrati-Kopaei:2013:SFG

[KKMSA13] Mahmood Kharrati-Kopaei, Ahad Malekzadeh, and Mohammad Sadooghi-Alvandi. Simultaneous fiducial generalized confidence intervals for the successive differences of exponential location parameters under heteroscedasticity. Statistics & Probability Letters, 83(6):1547–1552, June 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000746.

Kordzakhia:2018:LDE

[KKNH18] Nino E. Kordzakhia, Yury A. Kutoyants, Alexander A. Novikov, and Lin-Yee Hin. On limit distributions of estimators in irregular statistical models and a new representation of fractional Brownian motion. Statistics & Probability Letters, 139(??):141– 151, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301445.

Krampe:2015:HWB

[KKP15] J. Krampe, J.-P. Kreiss, and E. Paparoditis. Hybrid wild bootstrap for nonparametric trend estimation in locally stationary time series. Statistics & Probability Letters, 101(??):54–63, June 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000838.

Kreer:2014:FPP

[KKT14] Markus Kreer, Ayse Kizilersü, and Anthony W. Thomas. Fractional Poisson processes and their representation by infinite systems of ordinary differential equations. Statistics & Probability Letters, 84(??):27–32, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300326X.

Kohler:2011:EES

[KKW11] Michael Kohler, Adam Krzyzak, and Harro Walk. Estimation of the essential supremum of a regression function. Statistics & Probability Letters, 81(6):685–693, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000575.

Kohler:2014:NRQ

[KKW14] Michael Kohler, Adam Krzyzak, and Harro Walk. Nonparametric recursive quantile estimation. Statistics & Probability Letters, 93(??):102–107, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002090.

Khanchi:2011:STA

[KL11] Aziz Khanchi and Gilles Lamothe. Simulating tail asymptotics of a Markov chain. Statistics & Probability Letters, 81 (9):1392–1397, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001374.

Kolesko:2015:MEC

[KL15] Konrad Kolesko and Rafal Latala. Moment estimates for chaoses generated by symmetric random variables with logarithmically convex tails. Statistics & Probability Letters, 107:210–214, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215003107.

Kou:2016:ECT

[KL16] Junke Kou and Youming Liu. An extension of Chesneau’s theorem. Statistics & Probability Letters, 108:23–32, January 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003375.

Kaenmaki:2017:GDD

[KL17] Antti Käenmäki and Bing Li. Genericity of dimension drop on self-affine sets. Statistics & Probability Letters, 126(??):18– 25, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300846.

Kung:2012:ONN

[KLK12] Yi-Hung Kung, Pei-Sheng Lin, and Cheng-Hsiung Kao. An optimal k-nearest neighbor for density estimation. Statistics & Probability Letters, 82(10):1786–1791, October 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001927.

Kebabi:2011:LSE

[KLM11] K. Kebabi, I. Laroussi, and F. Messaci. Least squares estimators of the regression function with twice censored data. Statistics & Probability Letters, 81(11):1588–1593, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002124.

Kammoun:2017:MLE

[KLM17] Kaouthar Kammoun, Mahdi Louati, and Afif Masmoudi. Maximum likelihood estimator of the scale parameter for the Riesz distribution. Statistics & Probability Letters, 126(??):127– 131, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300871.

Kim:2014:QML

[KLNB14] Moosup Kim, Taewook Lee, Jungsik Noh, and Changryong Baek. Quasi-maximum likelihood estimation for multiple volatility shifts. Statistics & Probability Letters, 86(??):50–60, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213004069.

Kim:2011:ABL

[KLW11] Dongjae Kim, Sungchul Lee, and Wensheng Wang. The asymptotic behavior of linear placement statistics. Statistics & Probability Letters, 81(2):326–336, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521000307X.

Kleptsyna:2014:ETQ

[KLY14] Marina Kleptsyna, Alain Le Breton, and Bernard Ycart. Exponential transform of quadratic functional and multiplicative ergodicity of a Gauss–Markov process. Statistics & Probability Letters, 87(??):70–75, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000042.

Kawai:2011:LAB

[KM11] Reiichiro Kawai and Hiroki Masuda. On the local asymptotic behavior of the likelihood function for Meixner Lévy processes under high-frequency sampling. Statistics & Probability Letters, 81(4):460–469, April 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210003512.

Kebabi:2012:RAC

[KM12a] Khedidja Kebabi and Fatiha Messaci. Rate of the almost complete convergence of a kernel regression estimate with twice censored data. Statistics & Probability Letters, 82(11):1908–1913, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521200260X.

Kholfi:2012:CTZ

[KM12b] Sanaa Kholfi and Hosam M. Mahmoud. The class of tenable zero-balanced Pólya urns with an initially dominant subset of colors. Statistics & Probability Letters, 82(1):49–57, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002707.

Kagan:2013:MSS

[KM13a] Abram M. Kagan and Yaakov Malinovsky. Monotonicity in the sample size of the length of classical confidence intervals. Statistics & Probability Letters, 83(1):78–82, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003306.

Kella:2013:TAR

[KM13b] Offer Kella and Michel Mandjes. Transient analysis of reflected Lévy processes. Statistics & Probability Letters, 83(10):2308– 2315, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002137.

Kerman:2013:SKB

[KM13c] Sean C. Kerman and James B. McDonald. Skewness-kurtosis bounds for the skewed generalized T and related distributions. Statistics & Probability Letters, 83(9):2129–2134, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521300196X.

Kokonendji:2013:MAE

[KM13d] Célestin C. Kokonendji and Afif Masmoudi. On the Monge– Ampère equation for characterizing gamma-Gaussian model.

Statistics & Probability Letters, 83(7):1692–1698, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001065.

Klesov:2017:MCS

[KM17] Oleg Klesov and Ilya Molchanov. Moment conditions in strong laws of large numbers for multiple sums and random measures. Statistics & Probability Letters, 131(??):56–63, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302675.

Kabaila:2018:TSP

[KM18] Paul Kabaila and Rheanna Mainzer. Two sources of poor coverage of confidence intervals after model selection. Statistics & Probability Letters, 140(??):185–190, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301822.

Kloucha:2019:BLP

[KM19] Meryem Kada Kloucha and Tahar Mourid. Best linear predictor of a C[_{0}, 1]-valued functional autoregressive process. Statistics & Probability Letters, 150(??):114–120, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300793.

Kim:2010:PEC

[KMJ10] Michael Jong Kim, Viliam Makis, and Rui Jiang. Parameter estimation in a condition-based maintenance model. Statistics & Probability Letters, 80(21–22):1633–1639, November 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001896.

Kovchegov:2010:OTB

[KMN10] Yevgeniy Kovchegov, Nick Meredith, and Eyal Nir. Occupation times and Bessel densities. Statistics & Probability Letters, 80(2):104–110, January 15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209003654.

Kalpathy:2013:SLE

[KMR13] Ravi Kalpathy, Hosam M. Mahmoud, and Walter Rosenkrantz. Survivors in leader election algorithms. Statistics & Probability Letters, 83(12):2743–2749, December 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300309X.

Kubokawa:2015:ISE

[KMS15] Tatsuya Kubokawa, Éric Marchand, and William E. Strawderman. On improved shrinkage estimators for concave loss. Statistics & Probability Letters, 96(??):241–246, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003447.

Kumar:2011:FNI

[KMV11] A. Kumar, Mark M. Meerschaert, and P. Vellaisamy. Fractional normal inverse Gaussian diffusion. Statistics & Probability Letters, 81(1):146–152, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521000283X.

Kundu:2010:GMR

[KN10] Chanchal Kundu and Asok K. Nanda. On generalized mean residual life of record values. Statistics & Probability Letters, 80(9–10):797–806, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000180.

Kiapour:2011:RBP

[KN11] A. Kiapour and N. Nematollahi. Robust Bayesian prediction and estimation under a squared log error loss function. Statistics & Probability Letters, 81(11):1717–1724, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002331.

Kreher:2015:NKA

[KN15] Dörte Kreher and Ashkan Nikeghbali. A new kind of augmentation of filtrations suitable for a change of probability measure by a strict local martingale. Statistics & Probability Letters, 104(??):94–101, September 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001595.

Klass:2017:SSA

[KN17] Michael J. Klass and Krzysztof Nowicki. Sequential search algorithm for estimation of the number of members of a given population. Statistics & Probability Letters, 121(??):101–108, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301870.

Kumar:2011:TCP

[KNV11] A. Kumar, Erkan Nane, and P. Vellaisamy. Time-changed Poisson processes. Statistics & Probability Letters, 81(12):1899–1910, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002665.

Kaluszka:2011:SSW

[KO11] Marek Kaluszka and Andrzej Okolewski. Stability of L-statistics from weakly dependent observations. Statistics & Probability Letters, 81(5):618–625, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210003615.

Ko:2013:CCW

[Ko13] Mi-Hwa Ko. On complete convergence for weighted sums of asymptotically linear negatively dependent random field. Statistics & Probability Letters, 83(12):2615–2620, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002848.

Kazakeviciute:2017:PSL

[KO17] Agne Kazakeviciute and Malini Olivo. Point separation in logistic regression on Hilbert space-valued variables. Statistics & Probability Letters, 128(??):84–88, September 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730158X.

Ko:2018:CMC

[Ko18] Mi-Hwa Ko. On complete moment convergence for CAANA random vectors in Hilbert spaces. Statistics & Probability Letters, 138(??):104–110, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301135.

Kobayashi:2016:SBP

[Kob16] Kei Kobayashi. Small ball probabilities for a class of timechanged self-similar processes. Statistics & Probability Letters, 110(??):155–161, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215301218.

Kolodziejek:2017:LTR

[Kol17] Bartosz Kolodziejek. The left tail of renewal measure. Statistics & Probability Letters, 129(??):306–310, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302225.

Korchevsky:2015:GPS

[Kor15a] Valery Korchevsky. A generalization of the Petrov strong law of large numbers. Statistics & Probability Letters, 104(??):102–108, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001613.

Kromer:2015:FKF

[KOR15b] E. Kromer, L. Overbeck, and J. A. L. Röder. Feynman–Kac for functional jump diffusions with an application to Credit Value Adjustment. Statistics & Probability Letters, 105(??):120–129, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001935.

Koudou:2012:MYP

[Kou12a] Angelo Efoevi Koudou. A Matsumoto–Yor property for Kummer and Wishart random matrices. Statistics & Probability Letters, 82(11):1903–1907, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002581.

Kousha:2012:ABM

[Kou12b] Termeh Kousha. Asymptotic behavior and the moderate deviation principle for the maximum of a Dyck path. Statistics & Probability Letters, 82(2):340–347, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003361.

Kovchegov:2010:NAT

[Kov10] Yevgeniy Kovchegov. A note on adiabatic theorem for Markov chains. Statistics & Probability Letters, 80(3–4):186–190, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003915.

Kozubowski:2010:RSD

[KP10] Tomasz J. Kozubowski and Krzysztof Podgórski. Random self-decomposability and autoregressive processes. Statistics & Probability Letters, 80(21–22):1606–1611, November 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001756.

Kim:2014:RPF

[KP14] Jeong-Hoon Kim and Sang-Hyeon Park. A recursive pricing formula for a path-dependent option under the constant elasticity of variance diffusion. Statistics & Probability Letters, 94(??):39–47, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002405.

Kim:2015:CRC

[KP15a] Yoon Tae Kim and Hyun Suk Park. Convergence rate of CLT for the estimation of Hurst parameter of fractional Brownian motion. Statistics & Probability Letters, 105(??):181–188, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001844.

Kizildemir:2015:SOP

[KP15b] Bünyamin Kizildemir and Nicolas Privault. Supermodular ordering of Poisson arrays. Statistics & Probability Letters, 98(??):136–143, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214004258.

Kwon:2015:NMI

[KP15c] Tae Yeon Kwon and Yousung Park. A new multiple imputation method for bounded missing values. Statistics & Probability Letters, 107:204–209, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003181.

Kozubowski:2016:TDR

[KP16] Tomasz J. Kozubowski and Krzysztof Podgórski. Transmuted distributions and random extrema. Statistics & Probability Letters, 116(??):6–8, September 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300219.

Khodsiani:2017:UOB

[KP17] R. Khodsiani and S. Pooladsaz. Universal optimal block designs under hub correlation structure. Statistics & Probability Letters, 129(??):387–392, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302286.

Kim:2015:MRM

[KPK15] Seongyong Kim, Yousung Park, and Daeyoung Kim. On missingat-random mechanism in two-way incomplete contingency tables. Statistics & Probability Letters, 96(??):196–203, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003460.

Kim:2013:WTD

[KPO13] Sungsu Kim, Chongjin Park, and Jungtaek Oh. On waiting time distribution of runs of ones or zeros in a Bernoulli sequence. Statistics & Probability Letters, 83(1):339–344, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003732.

Kyprianou:2012:BDC

[KR12] A. E. Kyprianou and Y.-X. Ren. Backbone decomposition for continuous-state branching processes with immigration. Statistics & Probability Letters, 82(1):139–144, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003014.

Kluppelberg:2013:OSR

[KR13] Claudia Klüppelberg and Morten Grud Rasmussen. Outcrossings of safe regions by generalized hyperbolic processes. Statistics & Probability Letters, 83(10):2197–2204, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002113.

Koch:2019:GES

[KR19] Erwan Koch and Christian Y. Robert. Geometric ergodicity for some space-time max-stable Markov chains. Statistics & Probability Letters, 145(??):43–49, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302372.

Khorashadizadeh:2013:DTI

[KRB13] M. Khorashadizadeh, A. H. Rezaei Roknabadi, and G. R. Mohtashami Borzadaran. Doubly truncated (interval) cumulative residual and past entropy. Statistics & Probability Letters, 83(5):1464–1471, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521300045X.

Krebs:2017:CAN

[Kre17] Johannes T. N. Krebs. Consistency and asymptotic normality of stochastic Euler schemes for ordinary differential equations. Statistics & Probability Letters, 125(??):1–8, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303886.

Krebs:2018:LDI

[Kre18] Johannes T. N. Krebs. A large deviation inequality for β-mixing time series and its applications to the functional kernel regression model. Statistics & Probability Letters, 133(??):50–58, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303073.

Krizmanic:2018:NJF

[Kri18] Danijel Krizmanić. A note on joint functional convergence of partial sum and maxima for linear processes. Statistics & Probability Letters, 138(??):42–46, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301081.

Kruglov:2010:CPD

[Kru10] Victor M. Kruglov. A characterization of the Poisson distribution. Statistics & Probability Letters, 80(23–24):2032–2034, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002609.

Kruglov:2017:RWC

[Kru17] Victor M. Kruglov. Relative weak compactness of sums of pairwise independent random variables. Statistics & Probability Letters, 122(??):36–41, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302243.

Kahrari:2016:MSN

[KRYAV16] F. Kahrari, M. Rezaei, F. Yousefzadeh, and R. B. Arellano-Valle. On the multivariate skew-normal-Cauchy distribution. Statistics & Probability Letters, 117(??):80–88, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630058X.

Kabaila:2010:AEI

[KS10] Paul Kabaila and Khreshna Syuhada. The asymptotic efficiency of improved prediction intervals. Statistics & Probability Letters, 80(17–18):1348–1353, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001252.

Kamihigashi:2012:OTM

[KS12] Takashi Kamihigashi and John Stachurski. An order-theoretic mixing condition for monotone Markov chains. Statistics & Probability Letters, 82(2):262–267, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100321X.

Klebanov:2013:IVS

[KS13] Lev B. Klebanov and Lenka Slámová. Integer valued stable random variables. Statistics & Probability Letters, 83(6):1513– 1519, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521300059X.

Kang:2015:RPC

[KS15a] Jiwon Kang and Junmo Song. Robust parameter change test for Poisson autoregressive models. Statistics & Probability Letters, 104(??):14–21, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521500139X.

Keener:2015:IHB

[KS15b] Robert W. Keener and Hokeun Sun. Inconsistent hybrid bootstrap confidence regions. Statistics & Probability Letters, 107:115–121, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003053.

Klein:2015:LBI

[KS15c] Martin Klein and Bimal Sinha. Likelihood-based inference for singly and multiply imputed synthetic data under a normal model. Statistics & Probability Letters, 105(??):168–175, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001893.

Kagan:2016:AGI

[KS16a] Abram M. Kagan and Gábor J. Székely. An analytic generalization of independence and identical distributiveness. Statistics & Probability Letters, 110(??):244–248, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003570.

Kubilius:2016:SEH

[KS16b] K. Kubilius and V. Skorniakov. On some estimators of the Hurst index of the solution of SDE driven by a fractional Brownian motion. Statistics & Probability Letters, 109:159–167, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003818.

Kim:2017:CKF

[KS17a] Arlene Kyoung Hee Kim and Seung Jun Shin. The cumulative Kolmogorov filter for model-free screening in ultrahigh dimensional data. Statistics & Probability Letters, 126(??):238– 243, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301037.

Koutras:2017:NBS

[KS17b] M. V. Koutras and E. M. Sofikitou. A new bivariate semiparametric control chart based on order statistics and concomitants. Statistics & Probability Letters, 129(??):340–347, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302195.

Kubilius:2017:SNC

[KS17c] K. Kubilius and V. Skorniakov. A short note on a class of statistics for estimation of the Hurst index of fractional Brownian motion. Statistics & Probability Letters, 121(??):78–82, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302127.

Kolli:2019:LRB

[KS19] Praveen Kolli and Andrey Sarantsev. Large rank-based models with common noise. Statistics & Probability Letters, 151(??):29– 35, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300811.

Karavias:2018:HOE

[KST18] Yiannis Karavias, Spyridon D. Symeonides, and Elias Tzavalis. Higher order expansions for error variance matrix estimates in the Gaussian AR(1) linear regression model. Statistics & Probability Letters, 135(??):54–59, April 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730370X.

Knopova:2015:LBH

[KSW15] V. Knopova, R. L. Schilling, and J. Wang. Lower bounds of the Hausdorff dimension for the images of Feller processes. Statistics & Probability Letters, 97(??):222–228, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214004039.

Kawai:2010:SAA

[KT10] Reiichiro Kawai and Atsushi Takeuchi. Sensitivity analysis for averaged asset price dynamics with gamma processes. Statistics & Probability Letters, 80(1):42–49, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003514.

Kumar:2011:SCR

[KT11] Vikas Kumar and H. C. Taneja. Some characterization results on generalized cumulative residual entropy measure. Statistics & Probability Letters, 81(8):1072–1077, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000794.

Kaj:2019:SEL

[KT19] Ingemar Kaj and Daniah Tahir. Stochastic equations and limit results for some two-type branching models. Statistics & Probability Letters, 150(??):35–46, July 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521930063X.

Kumar:2015:GEM

[Kum15] Vikas Kumar. Generalized entropy measure in record values and its applications. Statistics & Probability Letters, 106(??):46–51, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002278.

Kurisu:2019:NIS

[Kur19] Daisuke Kurisu. On nonparametric inference for spatial regression models under domain expanding and infill asymptotics. Statistics & Probability Letters, 155(??):Article 108543, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301804.

Kudraszow:2013:UCR

[KV13] Nadia L. Kudraszow and Philippe Vieu. Uniform consistency of kNN regressors for functional variables. Statistics & Probability Letters, 83(8):1863–1870, August 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001387.

Kumar:2015:ITS

[KV15] A. Kumar and P. Vellaisamy. Inverse tempered stable subordinators. Statistics & Probability Letters, 103(??):134–141, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001200.

Kokonendji:2016:PDA

[KV16] Célestin C. Kokonendji and Davit Varron. Performance of discrete associated kernel estimators through the total variation distance. Statistics & Probability Letters, 110(??):225–235, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003600.

Kataria:2017:SST

[KV17a] K. K. Kataria and P. Vellaisamy. Saigo space-time fractional Poisson process via Adomian decomposition method. Statistics & Probability Letters, 129(??):69–80, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301827.

Koul:2017:ADS

[KV17b] Hira L. Koul and P. Vellaisamy. Asymptotic distributions of some robust scale estimators in explosive AR(1) model. Statistics & Probability Letters, 126(??):157–163, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300986.

Kalpathy:2014:LEA

[KW14a] Ravi Kalpathy and Mark Daniel Ward. On a leader election algorithm: Truncated geometric case study. Statistics & Probability Letters, 87(??):40–47, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213004197.

Kremer:2014:AND

[KW14b] Alexander Kremer and Rafael Weißbach. Asymptotic normality for discretely observed Markov jump processes with an absorbing state. Statistics & Probability Letters, 90(??):136–139, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001035.

Kern:2015:DSS

[KW15] Peter Kern and Lina Wedrich. Dilatively semistable stochastic processes. Statistics & Probability Letters, 99(??):101–108, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000140.

Kang:2015:COB

[KWL15] Shuaimin Kang, Min Wang, and Tao Lu. On the consistency of the objective Bayes factor for the integral priors in the one-way random effects model. Statistics & Probability Letters, 103(??): 17–23, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500111X.

Kokonendji:2010:EDT

[KZ10] Célestin C. Kokonendji and Silvio S. Zocchi. Extensions of discrete triangular distributions and boundary bias in kernel estimation for discrete functions. Statistics & Probability Letters, 80(21–22):1655–1662, November 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002026.

Kim:2014:ALS

[KZ14] Donggyu Kim and Chunming Zhang. Adaptive linear stepup multiple testing procedure with the bias-reduced estimator. Statistics & Probability Letters, 87(??):31–39, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000054.

Khovansky:2017:CCS

[KZ17] Serguey Khovansky and Oleksandr Zhylyevskyy. On the consistency of a cross-sectional GMM estimator in the presence of an observable stochastic common data shock. Statistics & Probability Letters, 129(??):196–202, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302006.

Korolev:2013:CRW

[KZZ13] V. Yu. Korolev, L. M. Zaks, and A. I. Zeifman. On convergence of random walks generated by compound Cox processes to Lévy processes. Statistics & Probability Letters, 83(10):2432– 2438, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002502.

Luo:2013:CCP

[LA13] Wei Luo and Naomi S. Altman. A characterization of conjugate priors in exponential families with application to inverse regression. Statistics & Probability Letters, 83(2):650–654, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004026.

Lau:2018:RSD

[LAG^{+}18] F. Din-Houn Lau, Niall M. Adams, Mark A. Girolami, Liam J. Butler, and Mohammed Z. E. B. Elshafie. The role of statistics in data-centric engineering. Statistics & Probability Letters, 136(??):58–62, ???? 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300804.

Langovoy:2011:APM

[Lan11] Mikhail Langovoy. Algebraic polynomials and moments of stochastic integrals. Statistics & Probability Letters, 81(6):627– 631, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521100037X.

Lanconelli:2018:SDG

[Lan18] Alberto Lanconelli. Standardizing densities on Gaussian spaces. Statistics & Probability Letters, 137(??):243–250, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300440.

Laradji:2015:STP

[Lar15] A. Laradji. Sums of totally positive functions of order 2 and applications. Statistics & Probability Letters, 105(??):176–180, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001376.

Laurent:2010:FCR

[Lau10] Stéphane Laurent. Further comments on the representation problem for stationary processes. Statistics & Probability Letters, 80(7–8):592–596, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004738.

Linke:2017:CIE

[LB17] Yu. Yu. Linke and I. S. Borisov. Constructing initial estimators in one-step estimation procedures of nonlinear regression. Statistics & Probability Letters, 120(??):87–94, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301857.

Lachos:2011:HNR

[LBG11] Victor H. Lachos, Dipankar Bandyopadhyay, and Aldo M. Garay. Heteroscedastic nonlinear regression models based on scale mixtures of skew-normal distributions. Statistics & Probability Letters, 81(8):1208–1217, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001027.

Liu:2011:NVW

[LBH11] Chunxu Liu, Arne C. Bathke, and Solomon W. Harrar. A nonparametric version of Wilks’ lambda — asymptotic results and small sample approximations. Statistics & Probability Letters, 81(10):1502–1506, October 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001519.

Laissaoui:2014:LTL

[LBM14] Diffalah Laissaoui and Abdelatif Benchérif-Madani. A limit theorem for local time and application to random sets. Statistics & Probability Letters, 88(??):107–117, May 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000364.

Li:2017:RMM

[LBS17] Xiongya Li, Xiuqin Bai, and Weixing Song. Robust mixture multivariate linear regression by multivariate Laplace distribution. Statistics & Probability Letters, 130(??):32–39, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302328.

Lemonte:2010:ASB

[LC10] Artur J. Lemonte and Gauss M. Cordeiro. Asymptotic skewness in Birnbaum–Saunders nonlinear regression models. Statistics & Probability Letters, 80(9–10):892–898, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000362.

Liu:2014:DPD

[LC14] Xiao Liu and Zhenlong Chen. Dividend problems in the dual model with diffusion and exponentially distributed observation time. Statistics & Probability Letters, 87(??):175–183, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000285.

Lawrie:2015:OSW

[LCF15] Jock Lawrie, John B. Carlin, and Andrew B. Forbes. Optimal stepped wedge designs. Statistics & Probability Letters, 99(??): 210–214, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000309.

Lee:2010:SAR

[LCJ10] Joanne Lee, Wendy K. Tam Cho, and George G. Judge. Stigler’s approach to recovering the distribution of first significant digits in natural data sets. Statistics & Probability Letters, 80(2):82–88, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003563.

Lin:2014:ARE

[LCKK14] Pei-Sheng Lin, Feng-Chi Chen, Shu-Fu Kuo, and Yi-Hung Kung. Assessing the relationship of evolutionary rates and functional variables by mixture estimating equations. Statistics & Probability Letters, 94(??):248–256, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002788.

Li:2016:CSB

[LCLQ16] Hongyi Li, Kashinath Chatterjee, Bo Li, and Hong Qin. Construction of Sudoku-based uniform designs with mixed levels. Statistics & Probability Letters, 114(??):111–118, July 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000584.

Lim:2017:NIC

[LCW17] Chae Young Lim, Chien-Hung Chen, and Wei-Ying Wu. Numerical instability of calculating inverse of spatial covariance matrices. Statistics & Probability Letters, 129(??):182–188, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521730202X.

Li:2012:DAH

[LCX12] Ai-Ping Li, Zhao-Xia Chen, and Feng-Chang Xie. Diagnostic analysis for heterogeneous log-Birnbaum–Saunders regression models. Statistics & Probability Letters, 82(9):1690–1698, September 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001964.

Lopez-Diaz:2010:SRD

[LD10] Miguel López-Díaz. Some remarks on L^{p} dispersion orderings. Statistics & Probability Letters, 80(5–6):413–420, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004428.

Leung:2016:OIP

[LD16a] Dennis Leung and Mathias Drton. Order-invariant prior specification in Bayesian factor analysis. Statistics & Probability Letters, 111(??):60–66, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000055.

Lu:2016:DBS

[LD16b] Jiannan Lu and Alex Deng. Demystifying the bias from selective inference: A revisit to Dawid’s treatment selection problem. Statistics & Probability Letters, 118(??):8–15, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300888.

Li:2017:MSE

[LD17a] Xiaofei Li and Deng Ding. Mean square exponential stability of stochastic Hopfield neural networks with mixed delays. Statistics & Probability Letters, 126(??):88–96, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300858.

Lu:2017:RBC

[LD17b] Jiannan Lu and Alex Deng. On randomization-based causal inference for matched-pair factorial designs. Statistics & Probability Letters, 125(??):99–103, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300561.

Lu:2015:CAH

[LDD15] Jiannan Lu, Peng Ding, and Tirthankar Dasgupta. Construction of alternative hypotheses for randomization tests with ordinal outcomes. Statistics & Probability Letters, 107:348–355, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003326.

Lu:2018:VCS

[LDLC18] Xiaoling Lu, Fengchi Dong, Xiexin Liu, and Xiangyu Chang. Varying coefficient support vector machines. Statistics & Probability Letters, 132(??):107–115, January 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302924.

Li:2010:RHR

[LDZ10] Xiaohu Li, Gaofeng Da, and Peng Zhao. On reversed hazard rate in general mixture models. Statistics & Probability Letters, 80(7–8):654–661, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004817.

Le:2016:RRX

[Lê16] Khoa Lê. A remark on a result of Xia Chen. Statistics & Probability Letters, 118(??):124–126, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300852.

LE:2019:RBM

[LE19] Vi LE. Reflected Brownian motion with a drift that depends on its local time. Statistics & Probability Letters, 155(??): Article 108546, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521930183X.

Lee:2012:BEM

[Lee12a] Chihoon Lee. Bounds on exponential moments of hitting times for reflected processes on the positive orthant. Statistics & Probability Letters, 82(6):1120–1128, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000697.

Lee:2012:UEC

[Lee12b] Oesook Lee. V -uniform ergodicity of a continuous time asymmetric power GARCH(1,1) model. Statistics & Probability Letters, 82(4):812–817, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000077.

Lee:2019:MDA

[Lee19] Ji Hyung Lee. Martingale decomposition and approximations for nonlinearly dependent processes. Statistics & Probability Letters, 152(??):35–42, September 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301154.

Lehtomaa:2015:LBC

[Leh15] Jaakko Lehtomaa. Limiting behaviour of constrained sums of two variables and the principle of a single big jump. Statistics & Probability Letters, 107:157–163, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003089.

Lember:2011:CAS

[Lem11a] Jüri Lember. A correction on approximation of smoothing probabilities for hidden Markov models. Statistics & Probability Letters, 81(9):1463–1464, September 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001489.

Lember:2011:ASP

[Lem11b] Jüri Lember. On approximation of smoothing probabilities for hidden Markov models. Statistics & Probability Letters, 81(2):310–316, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210003056.

Lemonte:2013:GSU

[Lem13] Artur J. Lemonte. On the gradient statistic under model misspecification. Statistics & Probability Letters, 83(1):390–398, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521200380X.

Lengyel:2011:DCB

[Len11] Tamás Lengyel. Direct consequences of the basic Ballot Theorem. Statistics & Probability Letters, 81(10):1476–1481, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001866.

Li:2017:ESI

[LF17] Dingshi Li and Xiaoming Fan. Exponential stability of impulsive stochastic partial differential equations with delays. Statistics & Probability Letters, 126(??):185–192, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301165.

Liu:2014:MSP

[LFM14] Wei Liu, Mohammud Foondun, and Xuerong Mao. Mean square polynomial stability of numerical solutions to a class of stochastic differential equations. Statistics & Probability Letters, 92(??):173–182, September 1–262, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002041.

Li:2015:TSS

[LFM15] Xiaohu Li, Rui Fang, and Jie Mi. On the timing to switch on the standby in k-out-of-n redundant systems. Statistics & Probability Letters, 96(??):10–20, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003101.

Li:2012:NOF

[LG12] Benchong Li and Jianhua Guo. A note on one-factor analysis. Statistics & Probability Letters, 82(11):1949–1952, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002222.

Luo:2013:UTD

[LG13] June Luo and Patrick Gerard. Using thresholding differencebased estimators for variable selection in partial linear models. Statistics & Probability Letters, 83(12):2601–2606, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002861.

Liu:2015:NDG

[LG15a] Xiangdong Liu and Hui Guo. A note on the Davis–Gut law. Statistics & Probability Letters, 105(??):163–167, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002047.

Luo:2015:PCF

[LG15b] Shikai Luo and Subhashis Ghosal. Prediction consistency of forward iterated regression and selection technique. Statistics & Probability Letters, 107:79–83, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002849.

Lumley:2016:CTT

[LG16] Thomas Lumley and Daniel L. Gillen. Characterising transitive two-sample tests. Statistics & Probability Letters, 109:118–123, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215302996.

Li:2010:ELM

[LGP10] Zhouping Li, Yun Gong, and Liang Peng. Empirical likelihood method for intermediate quantiles. Statistics & Probability Letters, 80(11–12):1022–1029, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000647.

Lu:2015:SRG

[LGT15] Changli Lu, Shengjun Gan, and Yongge Tian. Some remarks on general linear model with new regressors. Statistics & Probability Letters, 97(??):16–24, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003642.

Liu:2012:NDR

[LGW12] Xijun Liu, Qingwu Gao, and Yuebao Wang. A note on a dependent risk model with constant interest rate. Statistics & Probability Letters, 82(4):707–712, April 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211004081.

Liu:2013:ETS

[LH13] MingTe Liu and Huey-Miin Hsueh. Exact tests of the superiority under the Poisson distribution. Statistics & Probability Letters, 83(5):1339–1345, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000382.

Leng:2014:CPU

[LH14a] Xuan Leng and Taizhong Hu. The closure property of 2RV under random sum. Statistics & Probability Letters, 92(??): 158–167, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214001990.

Liu:2014:OFD

[LH14b] Wei Liu and Yijun Hu. Optimal financing and dividend control of the insurance company with excess-of-loss reinsurance policy. Statistics & Probability Letters, 84(??):121–130, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003325.

Liu:2014:KLE

[LHM14] Jin V. Liu, Zongfu Huang, and Hongjun Mao. Karhunen– Loève expansion for additive Slepian processes. Statistics & Probability Letters, 90(??):93–99, July 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001102.

Liang:2015:SBS

[LHT15] Baosheng Liang, Tao Hu, and Xingwei Tong. Spline-based sieve estimation in monotone constrained varying-coefficient partially linear EV model. Statistics & Probability Letters, 103(??):169– 175, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001212.

Lv:2019:WPS

[LHW19] Wujun Lv, Ping He, and Qinghua Wang. Well-posedness for the Stochastic Novikov equation. Statistics & Probability Letters, 153(??):157–163, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301488.

Li:2011:CPD

[Li11a] Jinjun Li. A class of probability distribution functions preserving the packing dimension. Statistics & Probability Letters, 81(12):1782–1791, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002410.

Li:2011:FLS

[Li11b] Yuqiang Li. Fluctuation limits of site-dependent branching systems in critical and large dimensions. Statistics & Probability Letters, 81(11):1604–1611, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001908.

Lai:2012:GLF

[LI12] Chin-Diew Lai and Muhyiddin Izadi. Generalized logistic frailty model. Statistics & Probability Letters, 82(11):1969–1977, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002234.

Li:2013:SDE

[Li13a] Dingshi Li. The stationary distribution and ergodicity of a stochastic generalized logistic system. Statistics & Probability Letters, 83(2):580–583, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004221.

Li:2013:PQA

[Li13b] Jinzhu Li. On pairwise quasi-asymptotically independent random variables and their applications. Statistics & Probability Letters, 83(9):2081–2087, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001843.

Li:2013:NEC

[Li13c] Lina Li. Notes on entropic convergence and the weak entropy inequality. Statistics & Probability Letters, 83(4):1106–1110, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000084.

Li:2013:NSU

[Li13d] Wenbo Li. On the number of switches in unbiased cointossing. Statistics & Probability Letters, 83(7):1613–1618, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001004.

Li:2015:NKV

[Li15a] Jinglai Li. A note on the Karhunen–Loève expansions for infinite-dimensional Bayesian inverse problems. Statistics & Probability Letters, 106(??):1–4, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002291.

Li:2015:NAD

[Li15b] Linyuan Li. Nonparametric adaptive density estimation on random fields using wavelet method. Statistics & Probability Letters, 96(??):346–355, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003617.

Li:2015:RBT

[Li15c] Zaixing Li. A residual-based test for variance components in linear mixed models. Statistics & Probability Letters, 98(??):73– 78, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214004131.

Li:2017:NFTb

[Li17a] Jinzhu Li. A note on the finite-time ruin probability of a renewal risk model with Brownian perturbation. Statistics & Probability Letters, 127(??):49–55, August 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301281.

Li:2017:HLC

[Li17b] Kexue Li. Hölder continuity for stochastic fractional heat equation with colored noise. Statistics & Probability Letters, 129(??):34–41, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301591.

Li:2017:PCTa

[Li17c] Pei-Sen Li. Perturbations of continuous-time Markov chains. Statistics & Probability Letters, 125(??):17–24, June 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300408.

Li:2017:SLM

[Li17d] Xue-Mei Li. Strict local martingales: Examples. Statistics & Probability Letters, 129(??):65–68, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301645.

Li:2018:RAR

[Li18] Jinzhu Li. A revisit to asymptotic ruin probabilities for a bidimensional renewal risk model. Statistics & Probability Letters, 140(??):23–32, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301433.

Li:2019:LDC

[Li19] Jiange Li. Large deviations for conditional guesswork. Statistics & Probability Letters, 153(??):7–14, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301361.

Lian:2012:NCS

[Lia12a] Heng Lian. A note on the consistency of Schwarz’s criterion in linear quantile regression with the SCAD penalty. Statistics & Probability Letters, 82(7):1224–1228, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001320.

Lian:2012:SEI

[Lia12b] Heng Lian. Shrinkage estimation for identification of linear components in additive models. Statistics & Probability Letters, 82(2):225–231, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003300.

Liao:2013:FJA

[Lia13] Ming Liao. Fixed jumps of additive processes. Statistics & Probability Letters, 83(3):820–823, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200452X.

Lindstrom:2013:TIF

[Lin13] Erik Lindström. Tuned iterated filtering. Statistics & Probability Letters, 83(9):2077–2080, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001806.

Lindholm:2017:NCB

[Lin17a] Mathias Lindholm. A note on the connection between some classical mortality laws and proportional frailty. Statistics & Probability Letters, 126(??):76–82, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300718.

Linke:2017:ANO

[Lin17b] Yuliana Yu. Linke. Asymptotic normality of one-step Mestimators based on non-identically distributed observations. Statistics & Probability Letters, 129(??):216–221, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302031.

Lin:2019:SOT

[Lin19a] Jianxi Lin. Second order tail approximation for the maxima of randomly weighted sums with applications to ruin theory and numerical examples. Statistics & Probability Letters, 153(??):37– 47, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301440.

Lin:2019:JIS

[Lin19b] Qian Lin. Jensen inequality for superlinear expectations. Statistics & Probability Letters, 151(??):79–83, August 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300823.

Liu:2011:SCI

[Liu11] Fang Liu. Some correlations in intersection-union tests and their relationship with complete power. Statistics & Probability Letters, 81(4):518–523, April 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003445.

Liu:2012:SDA

[Liu12a] Ji-Chun Liu. Structure of a double autoregressive process driven by a hidden Markov chain. Statistics & Probability Letters, 82(7):1468–1473, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001344.

Liu:2012:NBI

[Liu12b] Jicheng Liu. A note on the bilateral inequality for a sequence of random variables. Statistics & Probability Letters, 82(5):871– 875, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521200051X.

Liu:2013:CLI

[Liu13] Jin V. Liu. On Chung’s law of the iterated logarithm for the Brownian time Lévy’s area process. Statistics & Probability Letters, 83(5):1404–1410, May 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000412.

Liu:2014:EIM

[Liu14] Kai Liu. Existence of invariant measures of stochastic systems with delay in the highest order partial derivatives. Statistics & Probability Letters, 94(??):267–272, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002752.

Liu:2017:RPE

[Liu17] Yan Liu. Robust parameter estimation for stationary processes by an exotic disparity from prediction problem. Statistics & Probability Letters, 129(??):120–130, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301803.

Luo:2013:ECM

[LK13] June Luo and K. B. Kulasekera. Error covariance matrix estimation using ridge estimator. Statistics & Probability Letters, 83(1):257–264, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003483.

Lorek:2014:EPR

[LK14] Paweł Lorek and Rafal Kulik. Empirical process of residuals for regression models with long memory errors. Statistics & Probability Letters, 86(??):7–16, March 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003994.

Lkabous:2019:NPR

[Lka19] Mohamed Amine Lkabous. A note on Parisian ruin under a hybrid observation scheme. Statistics & Probability Letters, 145(??):147–157, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303122.

Lee:2012:QAN

[LKK12] Sangin Lee, Yongdai Kim, and Sunghoon Kwon. Quadratic approximation for nonconvex penalized estimations with a diverging number of parameters. Statistics & Probability Letters, 82(9):1710–1717, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001873.

Leng:2010:LSA

[LL10] Chenlei Leng and Bo Li. Least squares approximation with a diverging number of parameters. Statistics & Probability Letters, 80(3–4):254–261, February 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004076.

Li:2012:MFR

[LL12a] Zhi Li and Jiaowan Luo. Mean-field reflected backward stochastic differential equations. Statistics & Probability Letters, 82 (11):1961–1968, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002520.

Li:2012:OBR

[LL12b] Zhi Li and Jiaowan Luo. One barrier reflected backward doubly stochastic differential equations with discontinuous monotone coefficients. Statistics & Probability Letters, 82(10):1841– 1848, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001952.

Lim:2012:NGC

[LL12c] Adrian P. C. Lim and Dejun Luo. A note on Gaussian correlation inequalities for nonsymmetric sets. Statistics & Probability Letters, 82(1):196–202, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003221.

Li:2013:LLI

[LL13] Deli Li and Han-Ying Liang. The limit law of the iterated logarithm in Banach space. Statistics & Probability Letters, 83(7):1800–1804, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001284.

Li:2014:OSI

[LL14] Lingfei Li and Vadim Linetsky. Optimal stopping in infinite horizon: An eigenfunction expansion approach. Statistics & Probability Letters, 85(??):122–128, February 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003982.

Li:2015:LRO

[LL15] Chen Li and Xiaohu Li. Likelihood ratio order of sample minimum from heterogeneous Weibull random variables. Statistics & Probability Letters, 97(??):46–53, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400368X.

Li:2017:NFTa

[LL17a] Benchong Li and Yang Li. A note on faithfulness and total positivity. Statistics & Probability Letters, 122(??):168–172, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302528.

Li:2017:PWS

[LL17b] Chen Li and Xiaohu Li. Preservation of weak stochastic arrangement increasing under fixed time left-censoring. Statistics & Probability Letters, 129(??):42–49, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301578.

Li:2017:CPS

[LL17c] Zhen Li and Jicheng Liu. C^{∞}-convergence of Picard’s successive approximations to solutions of stochastic differential equations. Statistics & Probability Letters, 129(??):203–209, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301992.

Lv:2017:LGJ

[LL17d] You Lv and Yuqiang Li. L^{2} limits of generalized Jiřina processes. Statistics & Probability Letters, 129(??):288–296, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521730216X.

Laib:2019:ANK

[LL19a] Naâmane Laïb and Djamal Louani. Asymptotic normality of kernel density function estimator from continuous time stationary and dependent processes. Statistics & Probability Letters, 145(??):187–196, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303109.

Li:2019:HRR

[LL19b] Chen Li and Xiaohu Li. Hazard rate and reversed hazard rate orders on extremes of heterogeneous and dependent random variables. Statistics & Probability Letters, 146(??):104–111, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303523.

Li:2012:OSQ

[LLC^{+}12] Ling-Wei Li, Loo-Hay Lee, Chun-Hung Chen, Bo Guo, and Ya-Jie Liu. An optimal L-statistics quantile estimator for a set of location-scale populations. Statistics & Probability Letters, 82(10):1853–1858, October 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001903.

Li:2012:UOS

[LLCG12] Ling-Wei Li, Loo-Hay Lee, Chun-Hung Chen, and Bo Guo. On unbiased optimal L-statistics quantile estimators. Statistics & Probability Letters, 82(11):1891–1897, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200212X.

Li:2011:ELL

[LLH11] Jinyu Li, Wei Liang, and Shuyuan He. Empirical likelihood for LAD estimators in infinite variance ARMA models. Statistics & Probability Letters, 81(2):212–219, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003299.

Liu:2015:CIC

[LLH15] Xuhua Liu, Na Li, and Yuqin Hu. Combining inferences on the common mean of several inverse Gaussian distributions based on confidence distribution. Statistics & Probability Letters, 105(??):136–142, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002023.

Li:2010:ELS

[LLHW10] Jinyu Li, Wei Liang, Shuyuan He, and Xianbin Wu. Empirical likelihood for the smoothed LAD estimator in infinite variance autoregressive models. Statistics & Probability Letters, 80(17–18):1420–1430, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001458.

Li:2018:RDS

[LLK^{+}18] Yuyuan Li, Jianqiu Lu, Chunhai Kou, Xuerong Mao, and Jiafeng Pan. Robust discrete-state-feedback stabilization of hybrid stochastic systems with time-varying delay based on Razumikhin technique. Statistics & Probability Letters, 139(??):152– 161, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301032.

Lai:2013:SEF

[LLL13] Peng Lai, Gaorong Li, and Heng Lian. Semiparametric estimation of fixed effects panel data single-index model. Statistics & Probability Letters, 83(6):1595–1602, June 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000783.

Lloyd:2010:HCA

[Llo10] Chris J. Lloyd. How close are alternative bootstrap P -values? Statistics & Probability Letters, 80(23–24):1972–1976, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002506.

Lloyd:2012:PAH

[Llo12] Chris J. Lloyd. A practical ad hoc adjustment to the Simes P -value. Statistics & Probability Letters, 82(7):1297–1302, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000818.

Li:2014:CSD

[LLO14] Hongyi Li, Qisheng Li, and Zujun Ou. Construction of Sudoku designs and Sudoku-based uniform designs. Statistics & Probability Letters, 89(??):51–57, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000790.

Leisen:2011:LBS

[LLP11] Fabrizio Leisen, Antonio Lijoi, and Christian Paroissin. Limiting behavior of the search cost distribution for the move-to-front rule in the stable case. Statistics & Probability Letters, 81(12):1827–1832, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002471.

Li:2019:HML

[LLP19] Yingqiu Li, Quansheng Liu, and Xuelian Peng. Harmonic moments, large and moderate deviation principles for Mandelbrot’s cascade in a random environment. Statistics & Probability Letters, 147(??):57–65, April 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303183.

Li:2017:NSP

[LLR17] Deli Li, Han-Ying Liang, and Andrew Rosalsky. A note on symmetrization procedures for the laws of large numbers. Statistics & Probability Letters, 121(??):136–142, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302188.

Li:2018:EFS

[LLR18] Deli Li, Han-Ying Liang, and Andrew Rosalsky. An extension of Feller’s strong law of large numbers. Statistics & Probability Letters, 132(??):83–90, January 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302973.

Li:2011:NER

[LLSW11] Yi-Ting Li, Dang-Zheng Liu, Xin Sun, and Zheng-Dong Wang. A note on eigenvalues of random block Toeplitz matrices with slowly growing bandwidth. Statistics & Probability Letters, 81(12):2026–2029, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100280X.

Li:2014:DVN

[LLW14] Dong Li, Muyi Li, and Wuqing Wu. On dynamics of volatilities in nonstationary GARCH models. Statistics & Probability Letters, 94(??):86–90, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002399.

Logachov:2019:LDP

[LLY19] A. Logachov, O. Logachova, and A. Yambartsev. Large deviations in a population dynamics with catastrophes. Statistics & Probability Letters, 149(??):29–37, June 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300355.

Liu:2013:BAM

[LLZ13] Yaqing Liu, Juxin Liu, and Fuxi Zhang. Bias analysis for misclassification in a multicategorical exposure in a logistic regression model. Statistics & Probability Letters, 83(12):2621–2626, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002897.

Liu:2015:WBA

[LM15a] Bo Liu and Majid Mojirsheibani. On a weighted bootstrap approximation of the L_{p} norms of kernel density estimators. Statistics & Probability Letters, 105(??):65–73, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001911.

Louati:2015:MIR

[LM15b] Mahdi Louati and Afif Masmoudi. Moment for the inverse Riesz distributions. Statistics & Probability Letters, 102(??): 30–37, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000991.

Li:2017:RCT

[LM17] Junping Li and Weiwei Meng. Regularity criterion for 2type Markov branching processes with immigration. Statistics & Probability Letters, 121(??):109–118, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301577.

Lopuhaa:2018:DBN

[LM18] Hendrik P. Lopuhaä and Eni Musta. The distance between a naive cumulative estimator and its least concave majorant. Statistics & Probability Letters, 139(??):119–128, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830141X.

Lin:2014:CPP

[LMH14] Tsung-I Lin, Paul D. McNicholas, and Hsiu J. Ho. Capturing patterns via parsimonious t mixture models. Statistics & Probability Letters, 88(??):80–87, May 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000261.

Lai:2015:PSA

[LML15] Peng Lai, Jie Meng, and Heng Lian. Polynomial spline approach for variable selection and estimation in varying coefficient models for time series data. Statistics & Probability Letters, 96(??):21– 27, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003198.

Liu:2015:SLT

[LMLW15] Wei Liu, Chaoqun Ma, Yingqiu Li, and Suming Wang. A strong limit theorem for the average of ternary functions of Markov chains in bi-infinite random environments. Statistics & Probability Letters, 100(??):12–18, May 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000358.

Lorenzo:2013:EDN

[LMM13] Edgardo Lorenzo, Ganesh Malla, and Hari Mukerjee. Estimation of distributions with the new better than used in expectation property. Statistics & Probability Letters, 83(5):1346–1352, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000291.

Letac:2018:MEF

[LMP18] Gérard Letac, Lutz Mattner, and Mauro Piccioni. The median of an exponential family and the normal law. Statistics & Probability Letters, 133(??):38–41, February 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303164.

Leisen:2019:FCN

[LMPR19] Fabrizio Leisen, Ramsés H. Mena, Freddy Palma, and Luca Rossini. On a flexible construction of a negative binomial model. Statistics & Probability Letters, 152(??):1–8, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301075.

Loregian:2012:AVG

[LMR12] Angela Loregian, Lorenzo Mercuri, and Edit Rroji. Approximation of the variance gamma model with a finite mixture of normals. Statistics & Probability Letters, 82(2):217–224, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003257.

Luczak:2017:MTP

[LMS17] Tomasz Luczak, Katarzyna Mieczkowska, and Matas Sileikis. On maximal tail probability of sums of nonnegative, independent and identically distributed random variables. Statistics & Probability Letters, 129(??):12–16, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301633.

Landriault:2015:NOS

[LMW15] David Landriault, Khouzeima Moutanabbir, and Gordon E. Willmot. A note on order statistics in the mixed Erlang case. Statistics & Probability Letters, 106(??):13–18, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001923.

Liquet:2015:DVM

[LN15] Benoit Liquet and Yoni Nazarathy. A dynamic view to moment matching of truncated distributions. Statistics & Probability Letters, 104(??):87–93, September 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001571.

Lifshits:2019:BET

[LN19] M. Lifshits and A. Nazarov. On Brownian exit times from perturbed multi-strips. Statistics & Probability Letters, 147(??): 1–5, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521830381X.

Leon-Novelo:2012:PIL

[LNC12] Luis Leon-Novelo and George Casella. Prior influence in linear regression when the number of covariates increases to infinity. Statistics & Probability Letters, 82(3):438–445, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003397.

Lietzen:2016:MDI

[LNI16] Niko Lietzén, Klaus Nordhausen, and Pauliina Ilmonen. Minimum distance index for complex valued ICA. Statistics & Probability Letters, 118(??):100–106, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301006.

Lahrouz:2013:ESD

[LO13] Aadil Lahrouz and Lahcen Omari. Extinction and stationary distribution of a stochastic SIRS epidemic model with non-linear incidence. Statistics & Probability Letters, 83(4):960–968, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004804.

Lo:2017:FGH

[Lo17] Ambrose Lo. Functional generalizations of Hoeffding’s covariance lemma and a formula for Kendall’s tau. Statistics & Probability Letters, 122(??):218–226, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302693.

Lahrouz:2011:DSS

[LOKB11] A. Lahrouz, L. Omari, D. Kiouach, and A. Belmaâti. Deterministic and stochastic stability of a mathematical model of smoking. Statistics & Probability Letters, 81(8):1276–1284, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001209.

Longla:2013:RSC

[Lon13] Martial Longla. Remarks on the speed of convergence of mixing coefficients and applications. Statistics & Probability Letters, 83(10):2439–2445, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001144.

Loperfido:2010:NMC

[Lop10] Nicola Loperfido. A note on marginal and conditional independence. Statistics & Probability Letters, 80(23–24):1695–1699, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002014.

Loperfido:2013:SLD

[Lop13] Nicola Loperfido. Skewness and the linear discriminant function. Statistics & Probability Letters, 83(1):93–99, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003379.

Loperfido:2015:VVS

[Lop15] Nicola Loperfido. Vector-valued skewness for model-based clustering. Statistics & Probability Letters, 99(??):230–237, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000243.

Lotov:2019:BPL

[Lot19] V. I. Lotov. Bounds for the probability to leave the interval. Statistics & Probability Letters, 145(??):141–146, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303110.

Louati:2013:MRE

[Lou13] Mahdi Louati. Mixture and reciprocity of exponential models. Statistics & Probability Letters, 83(2):452–458, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003859.

Lowe:2018:SLM

[Löw18] Matthias Löwe. The semicircle law for matrices with ergodic entries. Statistics & Probability Letters, 141(??):90–95, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302128.

Ley:2010:MSM

[LP10a] Christophe Ley and Davy Paindaveine. Multivariate skewing mechanisms: a unified perspective based on the transformation approach. Statistics & Probability Letters, 80(23–24): 1685–1694, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001987.

Li:2010:CEM

[LP10b] Deyuan Li and Liang Peng. Comparing extreme models when the sign of the extreme value index is known. Statistics & Probability Letters, 80(7–8):739–746, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521000009X.

Leisen:2011:NVB

[LP11] Fabrizio Leisen and Cecilia Prosdocimi. A note on variance bounding for continuous time Markov chains. Statistics & Probability Letters, 81(1):153–156, January 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002853.

Li:2013:ARE

[LP13] Erning Li and Mohsen Pourahmadi. An alternative REML estimation of covariance matrices in linear mixed models. Statistics & Probability Letters, 83(4):1071–1077, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004877.

Losidis:2017:TSB

[LP17] Sotirios Losidis and Konstadinos Politis. A two-sided bound for the renewal function when the interarrival distribution is IMRL. Statistics & Probability Letters, 125(??):164–170, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300500.

Ladjimi:2019:ABD

[LP19a] Fetima Ladjimi and Marc Peigné. On the asymptotic behavior of the Diaconis–Freedman chain on [0, 1]. Statistics & Probability Letters, 145(??):1–11, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302062.

Longobardi:2019:RDR

[LP19b] Maria Longobardi and Franco Pellerey. On the role of dependence in residual lifetimes. Statistics & Probability Letters, 153(??):56–64, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301397.

Lv:2013:ARC

[LPH13] Wenhua Lv, Xiaoqing Pan, and Taizhong Hu. Asymptotics of the risk concentration based on the tail distortion risk measure. Statistics & Probability Letters, 83(12):2703–2710, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002952.

Liao:2013:AEM

[LPN13] Xin Liao, Zuoxiang Peng, and Saralees Nadarajah. Asymptotic expansions for moments of skew-normal extremes. Statistics & Probability Letters, 83(5):1321–1329, May 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000539.

Liao:2014:RCE

[LPNW14] Xin Liao, Zuoxiang Peng, Saralees Nadarajah, and Xiaoqian Wang. Rates of convergence of extremes from skew-normal samples. Statistics & Probability Letters, 84(??):40–47, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003258.

Leonenko:2012:NIG

[LPS12] N. N. Leonenko, S. Petherick, and A. Sikorskii. A normal inverse Gaussian model for a risky asset with dependence. Statistics & Probability Letters, 82(1):109–115, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100294X.

Longla:2015:KED

[LPS15] Martial Longla, Magda Peligrad, and Hailin Sang. On kernel estimators of density for reversible Markov chains. Statistics & Probability Letters, 100(??):149–157, May 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000632.

Li:2010:SCF

[LPW10] Wenbo V. Li, Natesh S. Pillai, and Robert L. Wolpert. On the supremum of certain families of stochastic processes. Statistics & Probability Letters, 80(11–12):916–921, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000404.

Li:2018:SNR

[LQ18] Hongyi Li and Hong Qin. Some new results on triple designs. Statistics & Probability Letters, 139(??):1–9, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301172.

Li:2019:AEL

[LQ19] Yizeng Li and Yongcheng Qi. Adjusted empirical likelihood method for the tail index of a heavy-tailed distribution. Statistics & Probability Letters, 152(??):50–58, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521930118X.

Liu:2015:DGL

[LQC15] Xiangdong Liu, Hangyong Qian, and Linqiu Cao. The Davis–Gut law for moving average processes. Statistics & Probability Letters, 104(??):1–6, September 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001364.

Li:2013:ECP

[LQZ13] Yunxia Li, Lianfen Qian, and Wei Zhang. Estimation in a change-point hazard regression model with long-term survivors. Statistics & Probability Letters, 83(7):1683–1691, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001090.

Lii:2011:ECN

[LR11] Keh-Shin Lii and Murray Rosenblatt. Estimation for a class of nonstationary processes. Statistics & Probability Letters, 81(11):1612–1622, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002112.

Lopez:2012:KRJ

[LR12a] Oscar López and Nikita Ratanov. Kac’s rescaling for jumptelegraph processes. Statistics & Probability Letters, 82(10): 1768–1776, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200199X.

Loubes:2012:AME

[LR12b] Jean-Michel Loubes and Paul Rochet. Approximate maximum entropy on the mean for instrumental variable regression. Statistics & Probability Letters, 82(5):972–978, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000533.

Li:2013:MSR

[LR13a] Shuanming Li and Jiandong Ren. The maximum severity of ruin in a perturbed risk process with Markovian arrivals. Statistics & Probability Letters, 83(4):993–998, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004786.

Lu:2013:ABS

[LR13b] Wen Lu and Yong Ren. Anticipated backward stochastic differential equations on Markov chains. Statistics & Probability Letters, 83(7):1711–1719, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001053.

Lu:2015:MFB

[LRH15] Wen Lu, Yong Ren, and Lanying Hu. Mean-field backward stochastic differential equations with subdifferential operator and its applications. Statistics & Probability Letters, 106(??):73– 81, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002187.

Liu:2019:GRE

[LRL19] Xiaohui Liu, Jafer Rahman, and Shihua Luo. Generalized and robustified empirical depths for multivariate data. Statistics & Probability Letters, 146(??):70–79, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303420.

Lin:2010:NEP

[LS10] Xiaoyan Lin and Dongchu Sun. A note on the existence of the posteriors for one-way random effect probit models. Statistics & Probability Letters, 80(1):57–62, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003538.

Li:2012:ARS

[LS12a] Deli Li and Aurel Spataru. Asymptotics related to a series of T. L. Lai. Statistics & Probability Letters, 82(8):1538–1548, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001496.

Liang:2012:OPG

[LS12b] Ye Liang and Dongchu Sun. Objective priors for generative star-shape models. Statistics & Probability Letters, 82(5):991– 997, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000557.

Lindo:2015:ARN

[LS15] Alexey Lindo and Serik Sagitov. Asymptotic results for the number of Wagner’s solutions to a generalised birthday problem. Statistics & Probability Letters, 107:356–361, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003338.

Ley:2016:GPS

[LS16a] Christophe Ley and Yvik Swan. A general parametric Stein characterization. Statistics & Probability Letters, 111(??):67– 71, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216000079.

Li:2016:SDB

[LS16b] Min Li and Yufeng Shi. Solving the double barrier reflected BSDEs via penalization method. Statistics & Probability Letters, 110(??):74–83, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003922.

Lai:2017:MFF

[LSCL17] Peng Lai, Fengli Song, Kaiwen Chen, and Zhi Liu. Model free feature screening with dependent variable in ultrahigh dimensional binary classification. Statistics & Probability Letters, 125(??):141–148, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300603.

Lee:2013:ELS

[LSS13a] Cheon-Sig Lee, Stephen A. Sedory, and Sarjinder Singh. Estimating at least seven measures of qualitative variables from a single sample using randomized response technique. Statistics & Probability Letters, 83(1):399–409, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003768.

Levental:2013:SPF

[LSS13b] Shlomo Levental, Mark Schroder, and Sumit Sinha. A simple proof of functional Itô’s lemma for semimartingales with an application. Statistics & Probability Letters, 83(9):2019–2026, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001715.

Li:2015:PBS

[LSS15] Juan Li, Weixing Song, and Jianhong Shi. Parametric bootstrap simultaneous confidence intervals for differences of means from several two-parameter exponential distributions. Statistics & Probability Letters, 106(??):39–45, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002345.

Lin:2013:EMP

[LSSR13] Xiaoyan Lin, Dongchu Sun, Paul L. Speckman, and Jeffery N. Rouder. Existence of MLE and posteriors for a recognitionmemory model. Statistics & Probability Letters, 83(10):2415– 2421, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002459.

Leonenko:2017:FNH

[LST17] Nikolai Leonenko, Enrico Scalas, and Mailan Trinh. The fractional non-homogeneous Poisson process. Statistics & Probability Letters, 120(??):147–156, January 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301869.

Li:2010:NMS

[LT10] Jinzhu Li and Qihe Tang. A note on max-sum equivalence. Statistics & Probability Letters, 80(23–24):1720–1723, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002099.

Lowe:2014:HTS

[LT14] Matthias Löwe and Felipe Torres. On hitting times for a simple random walk on dense Erdős–Rényi random graphs. Statistics & Probability Letters, 89(??):81–88, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000807.

Lee:2017:NGL

[LT17] Chiang-Sheng Lee and Hsine-Jen Tsai. A note on the generalized linear exponential distribution. Statistics & Probability Letters, 124(??):49–54, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300020.

Li:2019:EMS

[LT19a] Libo Li and Dai Taguchi. On the Euler–Maruyama scheme for spectrally one-sided Lévy driven SDEs with Hölder continuous coefficients. Statistics & Probability Letters, 146(??):15– 26, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303419.

Li:2019:FKP

[LT19b] Yunke Li and Masayoshi Takeda. Feynman–Kac penalizations of rotationally symmetric α-stable processes. Statistics & Probability Letters, 148(??):82–87, May 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300124.

Li:2018:OAC

[LTP18] Fan Li, Elizabeth L. Turner, and John S. Preisser. Optimal allocation of clusters in cohort stepped wedge designs. Statistics & Probability Letters, 137(??):257–263, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300464.

Liang:2013:RMJ

[LTR13] Y. Liang, A. Thavaneswaran, and N. Ravishanker. RCA models: Joint prediction of mean and volatility. Statistics & Probability Letters, 83(2):527–533, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004038.

Li:2011:HOI

[LTvdVR11] Lingling Li, Eric Tchetgen Tchetgen, Aad van der Vaart, and James M. Robins. Higher order inference on a treatment effect under low regularity conditions. Statistics & Probability Letters, 81(7):821–828, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000769.

Li:2018:RSM

[LTW18] Baokun Li, Weizhong Tian, and Tonghui Wang. Remarks for the singular multivariate skew-normal distribution and its quadratic forms. Statistics & Probability Letters, 137(??):105–112, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300105.

Long:2012:GAS

[LTX12] Shujun Long, Lingying Teng, and Daoyi Xu. Global attracting set and stability of stochastic neutral partial functional differential equations with impulses. Statistics & Probability Letters, 82(9):1699–1709, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001939.

Lu:2011:LUB

[Lu11] Dawei Lu. Lower and upper bounds of large deviation for sums of subexponential claims in a multi-risk model. Statistics & Probability Letters, 81(12):1911–1919, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002677.

Lu:2012:LBL

[Lu12] Dawei Lu. Lower bounds of large deviation for sums of long-tailed claims in a multi-risk model. Statistics & Probability Letters, 82(7):1242–1250, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001137.

Lu:2015:UCB

[Lu15] Lu Lu. On the uniform consistency of the Bernstein density estimator. Statistics & Probability Letters, 107:52–61, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002837.

Lu:2016:CAR

[Lu16a] Jiannan Lu. Covariate adjustment in randomization-based causal inference for 2^{k} factorial designs. Statistics & Probability Letters, 119(??):11–20, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301250.

Lu:2016:RBR

[Lu16b] Jiannan Lu. On randomization-based and regression-based inferences for 2^{k} factorial designs. Statistics & Probability Letters, 112(??):72–78, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215304144.

Lu:2018:PIN

[Lu18] Jiannan Lu. On the partial identification of a new causal measure for ordinal outcomes. Statistics & Probability Letters, 137(??): 1–7, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300014.

Luo:2010:DMS

[Luo10] June Luo. The discovery of mean square error consistency of a ridge estimator. Statistics & Probability Letters, 80(5– 6):343–347, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209004325.

Luo:2012:AER

[Luo12] June Luo. Asymptotic efficiency of ridge estimator in linear and semiparametric linear models. Statistics & Probability Letters, 82(1):58–62, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002835.

Landsman:2011:BSG

[LV11a] Zinoviy Landsman and Steven Vanduffel. Bounds for some general sums of random variables. Statistics & Probability Letters, 81(3):382–391, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210003378.

Leon:2011:OCI

[LV11b] Jorge A. León and José Villa. An Osgood criterion for integral equations with applications to stochastic differential equations with an additive noise. Statistics & Probability Letters, 81(4): 470–477, April 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521000341X.

Lindholm:2011:DEF

[LV11c] Mathias Lindholm and Thomas Vallier. On the degree evolution of a fixed vertex in some growing networks. Statistics & Probability Letters, 81(6):673–677, June 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000617.

Leskela:2013:SOC

[LV13] Lasse Leskelä and Matti Vihola. Stochastic order characterization of uniform integrability and tightness. Statistics & Probability Letters, 83(1):382–389, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003690.

Luque-Vasquez:2014:NCS

[LVMS14] Fernando Luque-Vásquez and J. Adolfo Minjárez-Sosa. A note on the σ-compactness of sets of probability measures on metric spaces. Statistics & Probability Letters, 84(??):212–214, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003581.

Landsman:2015:SST

[LVY15] Zinoviy Landsman, Steven Vanduffel, and Jing Yao. Some Steintype inequalities for multivariate elliptical distributions and applications. Statistics & Probability Letters, 97(??):54–62, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003745.

Lin:2011:ALP

[LW11] Huonan Lin and Jian Wang. Asymptotic limit properties of the occupation measure for a transient Brownian sheet. Statistics & Probability Letters, 81(12):1978–1985, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100277X.

Lin:2012:NEH

[LW12] Jianxi Lin and Yuebao Wang. New examples of heavytailed O-subexponential distributions and related closure properties. Statistics & Probability Letters, 82(3):427–432, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003932.

Li:2013:ECM

[LW13] Lihui Li and Tao Wen. Estimation of C-MGARCH models based on the MBP method. Statistics & Probability Letters, 83(2):665– 673, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003811.

Li:2014:ATB

[LW14a] Xiaohu Li and Jintang Wu. Asymptotic tail behavior of Poisson shot-noise processes with interdependence between shock and arrival time. Statistics & Probability Letters, 88(??):15–26, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000431.

Lu:2014:SNN

[LW14b] Dawei Lu and Xiaoguang Wang. Some new normal comparison inequalities related to Gordon’s inequality. Statistics & Probability Letters, 88(??):133–140, May 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400056X.

Liu:2015:CGE

[LW15a] Yi Liu and Qihua Wang. Copula-graphic estimators for the marginal survival function with censoring indicators missing at random. Statistics & Probability Letters, 107:101–110, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002898.

Louchard:2015:TGE

[LW15b] Guy Louchard and Mark Daniel Ward. The truncated geometric election algorithm: Duration of the election. Statistics & Probability Letters, 101(??):40–48, June 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000681.

Luo:2015:CTM

[LW15c] Peng Luo and Falei Wang. On the comparison theorem for multidimensional G-SDEs. Statistics & Probability Letters, 96(??):38– 44, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003216.

Li:2016:WSB

[LW16] Daniel H. Li and Liqun Wang. A weighted simulation-based estimator for incomplete longitudinal data models. Statistics & Probability Letters, 113(??):16–22, June 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521530273X.

Liu:2017:NSW

[LW17a] Dang-Zheng Liu and Yanhui Wang. A note on spiked Wishart matrices. Statistics & Probability Letters, 127(??):1–6, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301190.

Liu:2017:CFE

[LW17b] Yibin Liu and Wenbin Wu. Closed-form estimation of a regression model with a mismeasured binary regressor and heteroskedasticity. Statistics & Probability Letters, 125(??):202– 206, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521730072X.

Liang:2013:MRS

[LWD13] Xue Liang, Guojing Wang, and Yinghui Dong. A Markov regime switching jump-diffusion model for the pricing of portfolio credit derivatives. Statistics & Probability Letters, 83(1):373– 381, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003756.

Ling:2016:MKR

[LWL16] Nengxiang Ling, Chao Wang, and Jin Ling. Modified kernel regression estimation with functional time series data. Statistics & Probability Letters, 114(??):78–85, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215302716.

Li:2011:BEB

[LWX11] Yongming Li, Chengdong Wei, and Guodong Xing. Berry– Esseen bounds for wavelet estimator in a regression model with linear process errors. Statistics & Probability Letters, 81(1):103– 110, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002749.

Liu:2017:GEP

[LWY17] Zhenxia Liu, Zhi Wang, and Xiangfeng Yang. A Gaussian expectation product inequality. Statistics & Probability Letters, 124 (??):1–4, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300044.

Li:2017:MDP

[LWYZ17] Yumeng Li, Ran Wang, Nian Yao, and Shuguang Zhang. A moderate deviation principle for stochastic Volterra equation. Statistics & Probability Letters, 122(??):79–85, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302413.

Liu:2010:NGV

[LX10a] Xuhua Liu and Xingzhong Xu. A new generalized p-value approach for testing the homogeneity of variances. Statistics & Probability Letters, 80(19–20):1486–1491, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001604.

Luan:2010:CLI

[LX10b] Nana Luan and Yimin Xiao. Chung’s law of the iterated logarithm for anisotropic Gaussian random fields. Statistics & Probability Letters, 80(23–24):1886–1895, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002397.

Ling:2012:ANC

[LX12] Nengxiang Ling and Qian Xu. Asymptotic normality of conditional density estimation in the single index model for functional time series data. Statistics & Probability Letters, 82 (12):2235–2243, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003239.

Lan:2018:SLN

[LX18] Xiaohong Lan and Yimin Xiao. Strong local nondeterminism of spherical fractional Brownian motion. Statistics & Probability Letters, 135(??):44–50, April 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303619.

Liu:2019:SAF

[LXL19] Shanshan Liu, Xiaoyu Xing, and Guoxin Liu. Semi-additive functionals of semi-Markov processes and measure-valued Poisson equation. Statistics & Probability Letters, 147(??):45– 56, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303511.

Liu:2015:GSD

[LXZ15] Xiangdong Liu, Jie Xiong, and Shuaiqi Zhang. The Gerber–Shiu discounted penalty function in the classical risk model with impulsive dividend policy. Statistics & Probability Letters, 107:183– 190, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215003168.

Liu:2016:GLD

[LY16] Zhenxia Liu and Xiangfeng Yang. A general large deviation principle for longest runs. Statistics & Probability Letters, 110(??):128–132, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215301462.

Liu:2017:RCC

[LY17] Haodong Liu and Shuzhen Yang. Representation and converse comparison theorems for multidimensional BSDEs. Statistics & Probability Letters, 127(??):67–74, August 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301256.

Li:2018:HIS

[LY18a] Zhi Li and Litan Yan. Harnack inequalities for SDEs driven by subordinator fractional Brownian motion. Statistics & Probability Letters, 134(??):45–53, March 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303292.

Liang:2018:MPR

[LY18b] Xiaoqing Liang and Virginia R. Young. Minimizing the probability of ruin: Two riskless assets with transaction costs and proportional reinsurance. Statistics & Probability Letters, 140(??):167–175, September 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521830186X.

Li:2013:NEU

[LYB13] Mengyuan Li, Dalei Yu, and Peng Bai. A note on the existence and uniqueness of quasi-maximum likelihood estimators for mixed regressive, spatial autoregression models. Statistics & Probability Letters, 83(2):568–572, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004099.

Liu:2014:ODR

[LYC14] Xin Liu, Rong-Xian Yue, and Kashinath Chatterjee. Roptimal designs in random coefficient regression models. Statistics & Probability Letters, 88(??):127–132, May 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000558.

Li:2011:ABR

[LYCK11] Wen Li, Cindy Yu, Alicia Carriquiry, and Wolfgang Kliemann. The asymptotic behavior of the r/s statistic for fractional Brownian motion. Statistics & Probability Letters, 81(1):83–91, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002725.

Li:2011:SIS

[LYW11] Yongming Li, Shanchao Yang, and Chengdong Wei. Some inequalities for strong mixing random variables with applications to density estimation. Statistics & Probability Letters, 81(2):250– 258, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002804.

Li:2010:EEN

[LZ10] Yang Li and Weidong Zhao. L^{p}-error estimates for numerical schemes for solving certain kinds of backward stochastic differential equations. Statistics & Probability Letters, 80(21–22): 1612–1617, November 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001768.

Liu:2011:PVF

[LZ11] Guangying Liu and Xinsheng Zhang. Power variation of fractional integral processes with jumps. Statistics & Probability Letters, 81(8):962–972, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000319.

Liu:2013:PUM

[LZ13] Qiaojing Liu and Shoujiang Zhao. Pointwise and uniform moderate deviations for nonparametric regression function estimator on functional data. Statistics & Probability Letters, 83(5):1372– 1381, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521300028X.

Li:2014:PEJ

[LZ14a] Yingqiu Li and Xiaowen Zhou. On pre-exit joint occupation times for spectrally negative Lévy processes. Statistics & Probability Letters, 94(??):48–55, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521400234X.

Lin:2014:ITE

[LZ14b] Cunjie Lin and Yong Zhou. Inference for the treatment effects in two sample problems with right-censored and lengthbiased data. Statistics & Probability Letters, 90(??):17–24, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001023.

Li:2016:LTR

[LZ16a] Deli Li and Shuhua Zhang. A limit theorem related to the Hartman–Wintner–Strassen LIL and the Chover LIL. Statistics & Probability Letters, 109:16–21, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003594.

Lu:2016:SAR

[LZ16b] Dawei Lu and Bin Zhang. Some asymptotic results of the ruin probabilities in a two-dimensional renewal risk model with some strongly subexponential claims. Statistics & Probability Letters, 114(??):20–29, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216000572.

Li:2017:PCTb

[LZ17a] Guanghui Li and Chongqi Zhang. The pseudo component transformation design for experiment with mixture. Statistics & Probability Letters, 131(??):19–24, December 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302584.

Li:2017:SDS

[LZ17b] Zhi Li and Wei Zhang. Stability in distribution of stochastic Volterra–Levin equations. Statistics & Probability Letters, 122(??):20–27, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521630222X.

Li:2019:INF

[LZ19] Shuangbo Li and Li-Xin Zhang. Identifying the number of factors using a white noise test. Statistics & Probability Letters, 152(??):92–99, September 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301142.

Liu:2017:NRP

[LZJ17] Peng Liu, Chunsheng Zhang, and Lanpeng Ji. A note on ruin problems in perturbed classical risk models. Statistics & Probability Letters, 120(??):28–33, January 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301778.

Ling:2013:NSP

[LZL13] Xiaoliang Ling, Peng Zhao, and Ping Li. A note on the stochastic properties of a scale change random effects model. Statistics & Probability Letters, 83(10):2407–2414, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002460.

Liu:2019:AEB

[LZL^{+}19] Yang Liu, Xiuzhen Zhang, Mengke Li, Guanfu Liu, and Lin Zhu. Abundance estimation based on optimal estimating function with missing covariates in capture–recapture studies. Statistics & Probability Letters, 152(??):15–20, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301063.

Lai:2016:EEH

[LZLW16] Peng Lai, Qingzhao Zhang, Heng Lian, and Qihua Wang. Efficient estimation for the heteroscedastic single-index varying coefficient models. Statistics & Probability Letters, 110(??):84– 93, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215003946.

Liu:2017:WTJ

[LZP17]  Kaituo Liu, Dejian Zhou, and Lihua Peng.  A weak

type John–Nirenberg theorem for martingales. Statistics & Probability Letters, 122(??):190–197, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302759.

Li:2013:IEM

[LZT13] Xinmin Li, Xiaohua Zhou, and Lili Tian. Interval estimation for the mean of lognormal data with excess zeros. Statistics & Probability Letters, 83(11):2447–2453, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002538.

Li:2011:SIM

[LZW11] Bainian Li, Kongsheng Zhang, and Libin Wu. A sharp inequality for martingales and its applications. Statistics & Probability Letters, 81(8):1260–1266, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100109X.

Lv:2014:RVS

[LZY14] Zhike Lv, Huiming Zhu, and Keming Yu. Robust variable selection for nonlinear models with diverging number of parameters. Statistics & Probability Letters, 91(??):90–97, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001448.

Li:2015:ORB

[LZY15] Peng Li, Ming Zhou, and Chuancun Yin. Optimal reinsurance with both proportional and fixed costs. Statistics & Probability Letters, 106(??):134–141, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521500228X.

Lin:2019:RIP

[LZYW19] Chanjuan Lin, Ming Zheng, Wen Yu, and Mingzhe Wu. Robust inference for the proportional hazards model with two-phase cohort sampling data. Statistics & Probability Letters, 153(??):98– 103, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301439.

Li:2015:TSD

[LZZ15a] Yingqiu Li, Xiaowen Zhou, and Na Zhu. Two-sided discounted potential measures for spectrally negative Lévy processes. Statistics & Probability Letters, 100(??):67–76, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000280.

Li:2015:GML

[LZZ15b] Zhiming Li, Shengli Zhao, and Runchu Zhang. On general minimum lower order confounding criterion for s-level regular designs. Statistics & Probability Letters, 99(??):202–209, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000322.

Ma:2010:NLS

[Ma10a] Chunhua Ma. A note on “Least squares estimator for discretely observed Ornstein–Uhlenbeck processes with small Lévy noises”. Statistics & Probability Letters, 80(19–20):1528–1531, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210001677.

Mahmoud:2010:GLE

[MA10b] M. A. W. Mahmoud and Farouq Mohammad A. Alam. The generalized linear exponential distribution. Statistics & Probability Letters, 80(11–12):1005–1014, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000623.

MirMostafaee:2011:PPF

[MA11] S. M. T. K. MirMostafaee and Jafar Ahmadi. Point prediction of future order statistics from an exponential distribution. Statistics & Probability Letters, 81(3):360–370, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003366.

Ma:2013:KDV

[Ma13] Chunsheng Ma. K-distributed vector random fields in space and time. Statistics & Probability Letters, 83(4):1143–1150, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000059.

Ma:2014:SET

[Ma14a] Rugang Ma. Stochastic equations for two-type continuousstate branching processes with immigration and competition. Statistics & Probability Letters, 91(??):83–89, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001588.

Misra:2014:MCI

[MA14b] Neeraj Misra and Mohd. Arshad. Monotonicity of certain integrals involving gamma distributions and their applications in multiple comparisons. Statistics & Probability Letters, 85(??):144–152, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003945.

Ma:2015:LTC

[Ma15] Rugang Ma. Lamperti transformation for continuous-state branching processes with competition and applications. Statistics & Probability Letters, 107:11–17, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002795.

Maaroufi:2019:EFP

[Maa19] Nadir Maaroufi. An equivalent formula for Pitman’s asymptotic relative efficiency under weakened assumptions. Statistics & Probability Letters, 149(??):93–99, June 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300409.

Macci:2011:LDE

[Mac11] Claudio Macci. Large deviations for estimators of unknown probabilities, with applications in risk theory. Statistics & Probability Letters, 81(1):16–24, January 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002634.

Macci:2016:LDS

[Mac16] Claudio Macci. Large deviations for some non-standard telegraph processes. Statistics & Probability Letters, 110(??):119– 127, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215301437.

Madar:2015:DFC

[Mad15] Vered Madar. Direct formulation to Cholesky decomposition of a general nonsingular correlation matrix. Statistics & Probability Letters, 103(??):142–147, August 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001273.

Mahmoudian:2018:ESS

[Mah18] Behzad Mahmoudian. On the existence of some skew-Gaussian random field models. Statistics & Probability Letters, 137(??): 331–335, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521830097X.

Mai:2018:ESR

[Mai18] Jan-Frederik Mai. Exact simulation of reciprocal Archimedean copulas. Statistics & Probability Letters, 141(??):68–73, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302074.

Makasu:2010:CSD

[Mak10a] Cloud Makasu. Controlling a stopped diffusion process to reach a goal. Statistics & Probability Letters, 80(15–16):1218–1222, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000994.

Makasu:2010:NEB

[Mak10b] Cloud Makasu. A note on explicit bounds for a stopped Feynman–Kac functional. Statistics & Probability Letters, 80 (23–24):1977–1979, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0167715210002518.

Makri:2011:MMD

[Mak11] Frosso S. Makri. Minimum and maximum distances between failures in binary sequences. Statistics & Probability Letters, 81(3):402–410, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210003391.

Malinovsky:2018:OPG

[Mal18] Yaakov Malinovsky. On optimal policy in the group testing with incomplete identification. Statistics & Probability Letters, 140(??):44–47, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521830169X.

Mameli:2015:KSN

[Mam15] Valentina Mameli. The Kumaraswamy skew-normal distribution. Statistics & Probability Letters, 104(??):75–81, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001546.

Mao:2014:NTI

[Mao14] Guangyu Mao. A new test of independence for high-dimensional data. Statistics & Probability Letters, 93(??):14–18, October 1– 142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002004.

Mao:2015:MSE

[Mao15a] Guangyu Mao. Model selection of M-estimation models using least squares approximation. Statistics & Probability Letters, 99(??):238–243, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000334.

Mao:2015:NTC

[Mao15b] Guangyu Mao. A note on testing complete independence for high dimensional data. Statistics & Probability Letters, 106(??):82– 85, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002333.

Mao:2016:NTH

[Mao16] Guangyu Mao. A note on tests for high-dimensional covariance matrices. Statistics & Probability Letters, 117(??):89–92, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300554.

Markowsky:2011:ABM

[Mar11a] Greg Markowsky. Applying Brownian motion to the study of birth-death chains. Statistics & Probability Letters, 81(8):1173– 1178, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000964.

Martellosio:2011:EOE

[Mar11b] Federico Martellosio. Efficiency of the OLS estimator in the vicinity of a spatial unit root. Statistics & Probability Letters, 81(8):1285–1291, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001210.

Martin:2012:CRP

[Mar12a] Ryan Martin. Convergence rate for predictive recursion estimation of finite mixtures. Statistics & Probability Letters, 82(2):378–384, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003440.

Martsynyuk:2012:IPM

[Mar12b] Yuliya V. Martsynyuk. Invariance principles for a multivariate Student process in the generalized domain of attraction of the multivariate normal law. Statistics & Probability Letters, 82(12):2270–2277, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003136.

Marichal:2014:CSS

[Mar14a] Jean-Luc Marichal. Computing subsignatures of systems with exchangeable component lifetimes. Statistics & Probability Letters, 94(??):128–134, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002533.

Martschink:2014:BCE

[Mar14b] Bastian Martschink. Bounds on convergence for the empirical vector of the Curie–Weiss–Potts model with a non-zero external field vector. Statistics & Probability Letters, 88(??):118– 126, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000546.

Markeviciute:2016:ECT

[Mar16] J. Markevičiūtė. Epidemic change tests for the mean of innovations of an AR(1) process. Statistics & Probability Letters, 112(??):79–91, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000262.

Masry:2011:ECC

[Mas11] Elias Masry. The estimation of the correlation coefficient of bivariate data under dependence: Convergence analysis. Statistics & Probability Letters, 81(8):1039–1045, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000721.

Masse:2018:MLT

[Mas18] Bruno Massé. Modulo 1 limit theorems for autoregressive random variables and their sums. Statistics & Probability Letters, 135(??):70–75, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303693.

Maya:2014:NPE

[MASR14] R. Maya, E. I. Abdul-Sathar, and G. Rajesh. Non-parametric estimation of the generalized past entropy function with censored dependent data. Statistics & Probability Letters, 90(??):129– 135, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001059.

Matayoshi:2012:LCR

[Mat12] Jeffrey Matayoshi. The K-level crossings of a random algebraic polynomial with dependent coefficients. Statistics & Probability Letters, 82(1):203–211, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003130.

Mathew:2014:SAB

[Mat14] Angel Mathew. System availability behavior of some stationary dependent sequences. Statistics & Probability Letters, 84(??):17– 21, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003210.

Mattei:2017:MGM

[Mat17] Pierre-Alexandre Mattei. Multiplying a Gaussian matrix by a Gaussian vector. Statistics & Probability Letters, 128(??):67–70, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301359.

Maurer:2010:DC

[Mau10] Andreas Maurer. Dominated concentration. Statistics & Probability Letters, 80(7–8):683–689, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004854.

Maugis:2019:CE

[Mau19] Pierre-André G. Maugis. Correlation extrapolated. Statistics & Probability Letters, 145(??):89–95, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302827.

Minkova:2014:TIB

[MB14] Leda D. Minkova and N. Balakrishnan. Type II bivariate Pólya– Aeppli distribution. Statistics & Probability Letters, 88(??):40– 49, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000443.

Maitra:2015:BAS

[MB15] Trisha Maitra and Sourabh Bhattacharya. On Bayesian asymptotics in stochastic differential equations with random effects. Statistics & Probability Letters, 103(??):148–159, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001194.

Maitra:2016:ARC

[MB16] Trisha Maitra and Sourabh Bhattacharya. On asymptotics related to classical inference in stochastic differential equations with random effects. Statistics & Probability Letters, 110(??):278–288, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003442.

Magalhaes:2013:ASB

[MBS13] Tiago M. Magalhães, Denise A. Botter, and Mônica C. Sandoval. Asymptotic skewness for the beta regression model. Statistics & Probability Letters, 83(10):2236–2241, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002162.

Mandal:2013:OTA

[MBTC13] Saumen Mandal, Atanu Biswas, Paula Camelia Trandafir, and Mohammad Ziaul Islam Chowdhury. Optimal target allocation proportion for correlated binary responses in a 2 × 2 setup. Statistics & Probability Letters, 83(9):1991–1997, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001545.

Moon:2015:BET

[MC15] Hee-Jin Moon and Yong-Kab Choi. Berry–Esseen type theorems and the uniform law of the iterated logarithm for LPQD processes. Statistics & Probability Letters, 106(??):191–198, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521500245X.

McCrea:2012:SSL

[McC12] R. S. McCrea. Sufficient statistic likelihood construction for age- and time-dependent multi-state joint recapture and recovery data. Statistics & Probability Letters, 82(2):357–359, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003415.

McCollum:2013:UBR

[McC13] Joseph McCollum. An upper bound on random walks on dihedral groups. Statistics & Probability Letters, 83(8):1894–1900, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001430.

McElroy:2012:PIS

[McE12] Tucker McElroy. The perils of inferring serial dependence from sample autocorrelations of moving average series. Statistics & Probability Letters, 82(9):1632–1636, September 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001721.

McElroy:2017:CVA

[McE17] Tucker McElroy. Computation of vector ARMA autocovariances. Statistics & Probability Letters, 124(??):92–96, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300019.

McKeague:2015:CLT

[McK15] Ian W. McKeague. Central limit theorems under special relativity. Statistics & Probability Letters, 99(??):149–155, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000061.

McKinlay:2017:BDL

[McK17] Shaun McKinlay. On beta distributed limits of iterated linear random functions. Statistics & Probability Letters, 127(??):33– 41, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301219.

Manceur:2013:UML

[MD13] A. M. Manceur and P. Dutilleul. Unbiased modified likelihood ratio tests for simple and double separability of a variancecovariance structure. Statistics & Probability Letters, 83(2):631– 636, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003926.

Mesfioui:2015:COC

[MD15] Mhamed Mesfioui and Michel M. Denuit. Comonotonicity, orthant convex order and sums of random variables. Statistics & Probability Letters, 96(??):356–364, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003538.

Manna:2016:OTL

[MD16] Soumen Manna and Ashish Das. Optimal two-level designs for partial profile choice experiments. Statistics & Probability Letters, 116(??):80–87, September 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300869.

Michel:2019:MLI

[MdJ19] Jon Michel and Robert M. de Jong. A model for level induced conditional heteroskedasticity. Statistics & Probability Letters, 145(??):293–300, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303274.

Maume-Deschamps:2018:EQO

[MDN18] Véronique Maume-Deschamps and Ibrahima Niang. Estimation of quantile oriented sensitivity indices. Statistics & Probability Letters, 134(??):122–127, March 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303413.

Mandal:2016:OLH

[MDPP16] B. N. Mandal, Sukanta Dash, Shyamsundar Parui, and Rajender Parsad. Orthogonal Latin hypercube designs with special reference to four factors. Statistics & Probability Letters, 119(??):181– 185, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301468.

Martino:2012:MPM

[MDR12] Luca Martino, Victor Pascual Del Olmo, and Jesse Read. A multi-point Metropolis scheme with generic weight functions. Statistics & Probability Letters, 82(7):1445–1453, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001514.

Meller:2016:TSM

[Mel16] Rafal Meller. Two-sided moment estimates for a class of nonnegative chaoses. Statistics & Probability Letters, 119(??):213–219, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521630150X.

Mendy:2012:TPV

[Men12] Ibrahima Mendy. The two-parameter Volterra multifractional process. Statistics & Probability Letters, 82(12):2115–2124, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002933.

Meng:2018:CHP

[Men18] Xiao-Li Meng. Conducting highly principled data science: a statistician’s job and joy. Statistics & Probability Letters, 136(??):51–57, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300981.

Messaci:2010:LAE

[Mes10] Fatiha Messaci. Local averaging estimates of the regression function with twice censored data. Statistics & Probability Letters, 80(19–20):1508–1511, October 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521000163X.

Messaci:2014:DMV

[Mes14] Rabah Messaci. Distribution and moments of the vector of the maximum of two absolutely continuous random vectors. Statistics & Probability Letters, 92(??):256–262, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002065.

Metzler:2010:FPP

[Met10] Adam Metzler. On the first passage problem for correlated Brownian motion. Statistics & Probability Letters, 80(5–6):277–284, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209004180.

Misra:2015:RAS

[MF15] Neeraj Misra and Jisha Francis. Relative ageing of (n−k+1)-out-of-n systems. Statistics & Probability Letters, 106(??):272–280, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002461.

Mathlouthi:2015:RTF

[MFL15] Walid Mathlouthi, Marc Fredette, and Denis Larocque. Regression trees and forests for non-homogeneous Poisson processes. Statistics & Probability Letters, 96(??):204–211, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003459.

Mahajan:2011:NTT

[MGA11] Kalpana K. Mahajan, Anil Gaur, and Sangeeta Arora. A nonparametric test for a two-sample scale problem based on subsample medians. Statistics & Probability Letters, 81(8):983– 988, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000332.

Maurya:2011:STS

[MGG11] Vishal Maurya, Anju Goyal, and Amar Nath Gill. Simultaneous testing for the successive differences of exponential location parameters under heteroscedasticity. Statistics & Probability Letters, 81(10):1507–1517, October 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001854.

Miecznikowski:2011:DCD

[MGSL11] Jeffrey C. Miecznikowski, David Gold, Lori Shepherd, and Song Liu. Deriving and comparing the distribution for the number of false positives in single step methods to control k-FWER. Statistics & Probability Letters, 81(11):1695–1705, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002136.

Maity:2012:PLV

[MH12] Arnab Maity and Jianhua Z. Huang. Partially linear varying coefficient models stratified by a functional covariate. Statistics & Probability Letters, 82(10):1807–1814, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002118.

Muller:2011:CES

[MHH11] Christine H. Müller, Richard Huggins, and Wen-Han Hwang. Consistent estimation of species abundance from a presenceabsence map. Statistics & Probability Letters, 81(9):1449–1457, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001362.

Miasojedow:2014:HIG

[Mia14] Blazej Miasojedow. Hoeffding’s inequalities for geometrically ergodic Markov chains on general state space. Statistics & Probability Letters, 87(??):115–120, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000248.

Michna:2011:FSD

[Mic11] Zbigniew Michna. Formula for the supremum distribution of a spectrally positive α-stable Lévy process. Statistics & Probability Letters, 81(2):231–235, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003330.

Mihalache:2012:SAS

[Mih12] Stefan Mihalache. Strong approximations and sequential change-point analysis for diffusion processes. Statistics & Probability Letters, 82(3):464–472, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003804.

Mohammadi:2016:TWB

[MIL16] Faezeh Mohammadi, Muhyiddin Izadi, and Chin-Diew Lai. On testing whether burn-in is required under the long-run average cost. Statistics & Probability Letters, 110(??):217–224, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003612.

Miller:2019:EDC

[Mil19a] Jeffrey W. Miller. An elementary derivation of the Chinese restaurant process from Sethuraman’s stick-breaking process. Statistics & Probability Letters, 146(??):112–117, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303560.

Milliken:2019:AFM

[Mil19b] Evan Milliken. Applications of the fundamental matrix to mean absorption and conditional mean absorption problems. Statistics & Probability Letters, 151(??):106–115, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301026.

Minsker:2017:SEB

[Min17] Stanislav Minsker. On some extensions of Bernstein’s inequality for self-adjoint operators. Statistics & Probability Letters, 127(??):111–119, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301207.

MirMostafaee:2014:MOS

[Mir14] S. M. T. K. MirMostafaee. On the moments of order statistics coming from the Topp–Leone distribution. Statistics & Probability Letters, 95(??):85–91, December 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002958.

Misumi:2013:EER

[Mis13] Jun Misumi. Estimates on the effective resistance in anisotropic long-range percolation on Z^{2}. Statistics & Probability Letters, 83(4):953–959, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000114.

Mishura:2014:SML

[Mis14] Yuliya Mishura. Standard maximum likelihood drift parameter estimator in the homogeneous diffusion model is always strongly consistent. Statistics & Probability Letters, 86(??):24–29, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213004033.

Ma:2018:ELB

[MJH18] Yaolan Ma, Yuexiang Jiang, and Wei Huang. Empirical likelihood based inference for conditional Pareto-type tail index. Statistics & Probability Letters, 134(??):114–121, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303437.

Matsuura:2010:PST

[MK10] Shun Matsuura and Hiroshi Kurata. A principal subspace theorem for 2-principal points of general location mixtures of spherically symmetric distributions. Statistics & Probability Letters, 80(23–24):1863–1869, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521000235X.

Mandal:2014:OML

[MK14] B. N. Mandal and C. Koukouvinos. Optimal multi-level supersaturated designs through integer programming. Statistics & Probability Letters, 84(??):183–191, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003507.

Mojirsheibani:2016:AOK

[MK16] Majid Mojirsheibani and Jiajie Kong. An asymptotically optimal kernel combined classifier. Statistics & Probability Letters, 119(??):91–100, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301304.

Morikawa:2018:NET

[MK18] Kosuke Morikawa and Jae Kwang Kim. A note on the equivalence of two semiparametric estimation methods for nonignorable nonresponse. Statistics & Probability Letters, 140(??):1–6, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301408.

Matula:2019:ESL

[MKA19] Przemysław Matuła, Paweł Kurasiński, and André Adler. Exact strong laws of large numbers for ratios of the smallest order statistics. Statistics & Probability Letters, 152(??):69–73, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301178.

Madadi:2015:RMR

[MKJ15] Mohsen Madadi, Parisa Khalilpoor, and Ahad Jamalizadeh. Regression mean residual life of a system with three dependent components with normal lifetimes. Statistics & Probability Letters, 100(??):182–191, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521500053X.

Malekzadeh:2017:EMM

[MKK17] Ahad Malekzadeh and Mahmood Kharrati-Kopaei. An exact method for making inferences on the common location parameter of several heterogeneous exponential populations: Complete and censored data. Statistics & Probability Letters, 129(??):210– 215, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217302043.

Mnatsakanov:2013:RTI

[ML13] Robert M. Mnatsakanov and Shengqiao Li. The Radon transform inversion using moments. Statistics & Probability Letters, 83(3):936–942, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004427.

McLachlan:2016:CNR

[ML16] Geoffrey J. McLachlan and Sharon X. Lee. Comment on “On nomenclature, and the relative merits of two formulations of skew distributions” by A. Azzalini, R. Browne, M. Genton, and P. McNicholas. Statistics & Probability Letters, 116(??):1–5, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216300220. See [ABGM16].

Murakami:2017:UNT

[ML17] Hidetoshi Murakami and Seong Keon Lee. The unbiased nonparametric test and its moment generating function for the ordered alternative. Statistics & Probability Letters, 126(??):193– 197, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301001.

Mao:2013:QLI

[MLF13] Mingzhi Mao, Ting Liu, and Urszula Foryś. The quenched law of the iterated logarithm for one-dimensional random walks in a random environment. Statistics & Probability Letters, 83(1):52– 60, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003215.

Malla:2010:NPE

[MM10] Ganesh Malla and Hari Mukerjee. A new piecewise exponential estimator of a survival function. Statistics & Probability Letters, 80(23–24):1911–1917, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002427.

Misra:2011:NAR

[MM11] Amit Kumar Misra and Neeraj Misra. A note on active redundancy allocations in k-out-of-n systems. Statistics & Probability Letters, 81(10):1518–1523, October 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001829.

Misra:2012:NRS

[MM12] Neeraj Misra and Amit Kumar Misra. New results on stochastic comparisons of two-component series and parallel systems. Statistics & Probability Letters, 82(2):283–290, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003312.

Marichal:2013:CSS

[MM13a] Jean-Luc Marichal and Pierre Mathonet. Computing system signatures through reliability functions. Statistics & Probability Letters, 83(3):710–717, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004348.

Misra:2013:CRH

[MM13b] Neeraj Misra and Amit Kumar Misra. On comparison of reversed hazard rates of two parallel systems comprising of independent gamma components. Statistics & Probability Letters, 83(6):1567– 1570, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000758.

Morris:2013:DRM

[MM13c] Katherine Morris and Paul D. McNicholas. Dimension reduction for model-based clustering via mixtures of shifted asymmetric Laplace distributions. Statistics & Probability Letters, 83(9):2088–2093, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001326.

See erratum [MM14c].

Melnykov:2014:MAU

[MM14a] Igor Melnykov and Volodymyr Melnykov. On K-means algorithm with the use of Mahalanobis distances. Statistics & Probability Letters, 84(??):88–95, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003246.

Mohammadi:2014:EPS

[MM14b] Mohammad Mohammadi and Adel Mohammadpour. Estimating the parameters of an α-stable distribution using the existence of moments of order statistics. Statistics & Probability Letters, 90(??):78–84, July 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001011.

Morris:2014:EDR

[MM14c] Katherine Morris and Paul D. McNicholas. Erratum to “Dimension reduction for model-based clustering via mixtures of shifted asymmetric Laplace distributions” [Statist. Probab. Lett. 83 (9) (2013) 2088–2093]. Statistics & Probability Letters, 85(??):168, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003106. See [MM13c].

Mbodj:2015:AET

[MM15] Malick Mbodj and Thomas Mathew. Approximate ellipsoidal tolerance regions for multivariate normal populations. Statistics & Probability Letters, 97(??):41–45, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003708.

Majumder:2018:NMA

[MM18] Priyanka Majumder and Murari Mitra. On the non-monotonic analogue of a class based on the hazard rate order. Statistics & Probability Letters, 139(??):135–140, August 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301421.

Montoril:2014:SEF

[MMC14] Michel H. Montoril, Pedro A. Morettin, and Chang Chiann. Spline estimation of functional coefficient regression models for time series with correlated errors. Statistics & Probability Letters, 92(??):226–231, September 1–262, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001977.

Mynbaev:2017:RBN

[MMF17] Kairat Mynbaev and Carlos Martins-Filho. Reducing bias in nonparametric density estimation via bandwidth dependent kernels: L_{1} view. Statistics & Probability Letters, 123(??):17– 22, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216302723.

Moreira:2013:TRS

[MMM13] Elsa E. Moreira, João T. Mexia, and Christoph E. Minder. F tests with random sample sizes. theory and applications. Statistics & Probability Letters, 83(6):1520–1526, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000709.

Mathew:2016:IPI

[MMPW16] Thomas Mathew, Sandeep Menon, Inna Perevozskaya, and Samaradasa Weerahandi. Improved prediction intervals in heteroscedastic mixed-effects models. Statistics & Probability Letters, 114(??):48–53, July 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000560.

Melnykov:2017:SCV

[MMZ17] Yana Melnykov, Volodymyr Melnykov, and Xuwen Zhu. Studying contributions of variables to classification. Statistics & Probability Letters, 129(??):318–325, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302080.

Messaci:2011:LIL

[MN11] Fatiha Messaci and Nahima Nemouchi. A law of the iterated logarithm for the product limit estimator with doubly censored data. Statistics & Probability Letters, 81(8):1241–1244, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001064. See erratum [MN13].

Messaci:2013:ELI

[MN13] Fatiha Messaci and Nahima Nemouchi. Erratum to “A law of the iterated logarithm for the product limit estimator with doubly censored data” [Statist. Probab. Lett. 81 (2011) 1241– 1244]. Statistics & Probability Letters, 83(9):2142, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001673. See [MN11].

Mijena:2014:CST

[MN14] Jebessa B. Mijena and Erkan Nane. Correlation structure of time-changed Pearson diffusions. Statistics & Probability Letters, 90(??):68–77, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214001138.

Miasojedow:2016:GER

[MN16] Blazej Miasojedow and Wojciech Niemiro. Geometric ergodicity of Rao and Teh’s algorithm for homogeneous Markov jump processes. Statistics & Probability Letters, 113(??):1–6, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303928.

Moscovich:2017:FCB

[MN17] Amit Moscovich and Boaz Nadler. Fast calculation of boundary crossing probabilities for Poisson processes. Statistics & Probability Letters, 123(??):177–182, April 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302802.

Mnatsakanov:2011:MRA

[Mna11] Robert M. Mnatsakanov. Moment-recovered approximations of multivariate distributions: The Laplace transform inversion. Statistics & Probability Letters, 81(1):1–7, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002610.

Michalowicz:2011:GIT

[MNBO11] J. V. Michalowicz, J. M. Nichols, F. Bucholtz, and C. C. Olson. A general Isserlis theorem for mixed-Gaussian random variables. Statistics & Probability Letters, 81(8):1233–1240, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001052.

Matilainen:2015:NIC

[MNO15] Markus Matilainen, Klaus Nordhausen, and Hannu Oja. New independent component analysis tools for time series. Statistics & Probability Letters, 105(??):80–87, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001868.

Miettinen:2012:SPB

[MNOT12] Jari Miettinen, Klaus Nordhausen, Hannu Oja, and Sara Taskinen. Statistical properties of a blind source separation estimator for stationary time series. Statistics & Probability Letters, 82(11):1865–1873, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002593.

Martxednez:2016:CDC

[MNP16] Wilmer Martínez, Fabio H. Nieto, and Pilar Poncela. Choosing a dynamic common factor as a coincident index. Statistics & Probability Letters, 109:89–98, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003764.

Mynbaev:2014:IBK

[MNWA14] Kairat T. Mynbaev, Saralees Nadarajah, Christopher S. Withers, and Aziza S. Aipenova. Improving bias in kernel density estimation. Statistics & Probability Letters, 94(??):106–112, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002521.

Meerschaert:2013:FDR

[MNX13] Mark M. Meerschaert, Erkan Nane, and Yimin Xiao. Fractal dimension results for continuous time random walks. Statistics & Probability Letters, 83(4):1083–1093, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000023.

Mitov:2014:IAR

[MO14] Kosto V. Mitov and Edward Omey. Intuitive approximations for the renewal function. Statistics & Probability Letters, 84(??):72– 80, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003283.

Milosevic:2016:CBS

[MO16] B. Milosević and M. Obradović. Characterization based symmetry tests and their asymptotic efficiencies. Statistics & Probability Letters, 119(??):155–162, December 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301237.

Modarres:2011:HDG

[Mod11] Reza Modarres. High-dimensional generation of Bernoulli random vectors. Statistics & Probability Letters, 81(8):1136–1142, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000915.

Modarres:2014:IDB

[Mod14] Reza Modarres. On the interpoint distances of Bernoulli vectors. Statistics & Probability Letters, 84(??):215–222, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003556.

Modarres:2016:MPI

[Mod16] Reza Modarres. Multivariate Poisson interpoint distances. Statistics & Probability Letters, 112(??):113–123, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216000250.

Martin:2018:PMP

[MOD18] Ryan Martin, Cheng Ouyang, and François Domagni. ‘Purposely misspecified’ posterior inference on the volatility of a jump diffusion process. Statistics & Probability Letters, 134(??):106– 113, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303279.

Morrow:2018:MCT

[Mor18] Steven T. Morrow. A moment coboundary theorem for C[0, 1]-valued random fields. Statistics & Probability Letters, 142(??):62–70, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302360.

Mounaix:2019:ASU

[Mou19] Philippe Mounaix. Almost sure uniform convergence of a random Gaussian field conditioned on a large linear form to a non random profile. Statistics & Probability Letters, 148(??):164–168, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300240.

Mandal:2013:MDD

[MP13] Nripes Kumar Mandal and Manisha Pal. Maximin designs for the detection of synergistic effects. Statistics & Probability Letters, 83(7):1632–1637, July 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000928.

Macci:2015:LDC

[MP15a] Claudio Macci and Barbara Pacchiarotti. Large deviations for a class of counting processes and some statistical applications. Statistics & Probability Letters, 104(??):36–48, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001406.

Mijatovic:2015:AIT

[MP15b] Aleksandar Mijatović and Martijn Pistorius. Asymptotic independence of three statistics of maximal segmental scores. Statistics & Probability Letters, 99(??):185–191, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000218.

Mandrekar:2016:BMH

[MP16a] Vidyadhar Mandrekar and Andrey Pilipenko. On a Brownian motion with a hard membrane. Statistics & Probability Letters, 113(??):62–70, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215301279.

Maulik:2016:RPU

[MP16b] Krishanu Maulik and Moumanti Podder. Ruin probabilities under Sarmanov dependence structure. Statistics & Probability Letters, 117(??):173–182, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300803.

Macci:2017:LDE

[MP17] Claudio Macci and Barbara Pacchiarotti. Large deviations for estimators of the parameters of a neuronal response latency model. Statistics & Probability Letters, 126(??):65–75, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300822.

Mandal:2012:PBO

[MPA12] N. K. Mandal, Manisha Pal, and M. L. Aggarwal. Pseudo-Bayesian A-optimal designs for estimating the point of maximum in component-amount Darroch–Waller mixture model. Statistics & Probability Letters, 82(6):1088–1094, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000582.

Makri:2015:LMS

[MPA15] Frosso S. Makri, Zaharias M. Psillakis, and Anastasios N. Arapis. Length of the minimum sequence containing repeats of success runs. Statistics & Probability Letters, 96(??):28–37, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003149.

Mainassara:2015:SSL

[MR15a] Yacouba Boubacar Maïnassara and Hamdi Raïssi. Semi-strong linearity testing in linear models with dependent but uncorrelated errors. Statistics & Probability Letters, 103(??):110– 115, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001145.

Molchanov:2015:MPP

[MR15b] Ilya Molchanov and Kostiantyn Ralchenko. Multifractional Poisson process, multistable subordinator and related limit theorems. Statistics & Probability Letters, 96(??):95–101, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003228.

Mai:2010:PRS

[MS10a] Jan-Frederik Mai and Matthias Scherer. The Pickands representation of survival Marshall–Olkin copulas. Statistics & Probability Letters, 80(5–6):357–360, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004349.

Meitz:2010:NGE

[MS10b] Mika Meitz and Pentti Saikkonen. A note on the geometric ergodicity of a nonlinear AR–ARCH model. Statistics & Probability Letters, 80(7–8):631–638, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004787.

Meschenmoser:2011:FCL

[MS11] D. Meschenmoser and A. Shashkin. Functional Central Limit Theorem for the volume of excursion sets generated by associated random fields. Statistics & Probability Letters, 81(6):642– 646, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000587.

Mercadier:2012:ORC

[MS12a] Cécile Mercadier and Philippe Soulier. Optimal rates of convergence in the Weibull model based on kernel-type estimators. Statistics & Probability Letters, 82(3):548–556, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003762.

Mnatsakanov:2012:VKD

[MS12b] Robert Mnatsakanov and Khachatur Sarkisian. Varying kernel density estimation on R_{+}. Statistics & Probability Letters, 82(7):1337–1345, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001265.

Meerschaert:2013:TFB

[MS13] Mark M. Meerschaert and Farzad Sabzikar. Tempered fractional Brownian motion. Statistics & Probability Letters, 83(10):2269– 2275, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002216.

Matsuda:2016:PCPb

[MS16a] Takeru Matsuda and William E. Strawderman. Pitman closeness properties of Bayes shrinkage procedures in estimation and prediction. Statistics & Probability Letters, 119(??):21–29, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301213.

Matsuda:2016:PCPa

[MS16b] Takeru Matsuda and William E. Strawderman. Pitman closeness properties of point estimators and predictive densities with parametric constraints. Statistics & Probability Letters, 116(??):101–106, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300438.

Mnatsakanov:2018:RQQ

[MS18a] Robert M. Mnatsakanov and Aleksandre Sborshchikovi. Recovery of quantile and quantile density function using the frequency moments. Statistics & Probability Letters, 140(??):53–62, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301792.

Mojirsheibani:2018:CIF

[MS18b] Majid Mojirsheibani and Crystal Shaw. Classification with incomplete functional covariates. Statistics & Probability Letters, 139(??):40–46, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301305.

Mselmi:2018:CIS

[Mse18] Farouk Mselmi. Characterization of the inverse stable subordinator. Statistics & Probability Letters, 140(??):37–43, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301470.

Murillo-Salas:2017:RCS

[MSPSJ17] A. Murillo-Salas, J. L. Pérez, and A. Siri-Jégousse. Refracted continuous-state branching processes: Self-regulating populations. Statistics & Probability Letters, 123(??):34–44, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302796.

Meng:2018:EOC

[MSS18] Qingxin Meng, Yang Shen, and Peng Shi. On the existence of optimal controls for backward stochastic partial differential equations. Statistics & Probability Letters, 137(??):113–123, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300154.

Mitric:2010:MTC

[MST10] Ilie-Radu Mitric, Kristina P. Sendova, and Cary Chi-Liang Tsai. On a multi-threshold compound Poisson process perturbed by diffusion. Statistics & Probability Letters, 80(5–6):366–375, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209004362.

Muller:2015:ESR

[MSW15] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Estimators in step regression models. Statistics & Probability Letters, 100(??):124–129, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000620.

Maejima:2013:DRP

[MT13] Makoto Maejima and Ciprian A. Tudor. On the distribution of the Rosenblatt process. Statistics & Probability Letters, 83(6):1490–1495, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000692.

Milev:2017:ECF

[MT17a] Mariyan Milev and Aldo Tagliani. Entropy convergence of finite moment approximations in Hamburger and Stieltjes problems. Statistics & Probability Letters, 120(??):114–117, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630181X.

Mohammadi:2017:EPS

[MT17b] Zohreh Mohammadi and Mina Towhidi. Estimating the parameters of a selected bivariate normal population. Statistics & Probability Letters, 122(??):205–210, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302450.

Maeyama:2011:PTE

[MTT11] Yusuke Maeyama, Kenichiro Tamaki, and Masanobu Taniguchi. Preliminary test estimation for spectra. Statistics & Probability Letters, 81(11):1580–1587, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002082.

Maejima:2010:NBG

[MU10] Makoto Maejima and Yohei Ueda. A note on a bivariate gamma distribution. Statistics & Probability Letters, 80(23–24):1991–1994, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002543.

Maejima:2013:ESD

[MU13] Makoto Maejima and Yohei Ueda. Examples of α-selfdecomposable distributions. Statistics & Probability Letters, 83(1):286–291, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003689.

Mu:2015:LCM

[Mu15] Xiaosheng Mu. Log-concavity of a mixture of beta distributions. Statistics & Probability Letters, 99(??):125–130, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000176.

Meintanis:2016:NPW

[MU16] Simos G. Meintanis and Nikolai G. Ushakov. Nonparametric probability weighted empirical characteristic function and applications. Statistics & Probability Letters, 108:52–61, January 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003120.

Mukhopadhyay:2010:CIM

[Muk10] Nitis Mukhopadhyay. A convolution identity and more with illustrations. Statistics & Probability Letters, 80(23–24):1980– 1984, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521000252X.

Mukeru:2017:RLT

[Muk17] Safari Mukeru. Representation of local times of fractional Brownian motion. Statistics & Probability Letters, 131(??):1–12, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217302596.

Muller:2019:RCL

[Mül19] Ulrich K. Müller. Refining the central limit theorem approximation via extreme value theory. Statistics & Probability Letters, 155(??):Article 108564, December 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521930210X.

Michael:2010:CTN

[MV10] Skevi Michael and Stanislav Volkov. On a coloured tree with non i.i.d. random labels. Statistics & Probability Letters, 80(23–24):1896–1903, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002403.

Marinucci:2015:NGS

[MV15] Domenico Marinucci and Sreekar Vadlamani. A note on global suprema of band-limited spherical random functions. Statistics & Probability Letters, 96(??):141–148, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003290.

Mertens:2011:UDC

[MvdBV^{+}11] B. J. A. Mertens, Y. E. M. van der Burgt, B. Velstra, W. E. Mesker, R. A. E. M. Tollenaar, and A. M. Deelder. On the use of double cross-validation for the combination of proteomic mass spectral data for enhanced diagnosis and prediction. Statistics & Probability Letters, 81(7):759–766, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000836.

Molenberghs:2011:PLM

[MVI11] Geert Molenberghs, Geert Verbeke, and Samuel Iddi. Pseudolikelihood methodology for partitioned large and complex samples. Statistics & Probability Letters, 81(7):892–901, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000198.

Mallik:2011:CLT

[MW11] Atul Mallik and Michael Woodroofe. A Central Limit Theorem for linear random fields. Statistics & Probability Letters, 81(11):1623–1626, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002094.

Mena:2012:EIS

[MW12a] Ramsés H. Mena and Stephen G. Walker. An EPPF from independent sequences of geometric random variables. Statistics & Probability Letters, 82(6):1059–1066, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000776.

Meyer:2012:IPO

[MW12b] Mary C. Meyer and Jianqiang C. Wang. Improved power of one-sided tests. Statistics & Probability Letters, 82(8):1619– 1622, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001599.

Malham:2014:EAE

[MW14a] Simon J. A. Malham and Anke Wiese. Efficient almost-exact Lévy area sampling. Statistics & Probability Letters, 88(??): 50–55, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000339.

Morey:2014:SRB

[MW14b] Richard D. Morey and Eric-Jan Wagenmakers. Simple relation between Bayesian order-restricted and point-null hypothesis tests. Statistics & Probability Letters, 92(??):121–124, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001862.

Ma:2016:ESN

[MW16] Ni Ma and Ward Whitt. Efficient simulation of non-Poisson non-stationary point processes to study queueing approximations. Statistics & Probability Letters, 109:202–207, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215300079.

Matthews:2017:URS

[MW17] Michael J. Matthews and Douglas A. Wolfe. Unified ranked sampling. Statistics & Probability Letters, 122(??):173–178, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302681.

Ma:2018:NSG

[MW18] Yutao Ma and Xinyu Wang. A note on the spectral gap for general harmonic measures on spheres. Statistics & Probability Letters, 141(??):56–61, October 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302098.

Miao:2016:LTO

[MWA16] Yu Miao, Rujun Wang, and Andre Adler. Limit theorems for order statistics from exponentials. Statistics & Probability Letters, 110(??):51–57, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003892.

Miecznikowski:2017:NEM

[MWGT17] Jeffrey C. Miecznikowski, Jiefei Wang, Daniel P. Gaile, and David L. Tritchler. A novel exact method for significance of higher criticism via Steck’s determinant. Statistics & Probability Letters, 130(??):105–110, November 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302420.

Ming:2017:MED

[MWH17] Ruixing Ming, Wenyuan Wang, and Yijun Hu. On maximizing expected discounted taxation in a risk process with interest. Statistics & Probability Letters, 122(??):128–140, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302449.

Miao:2011:SLB

[MWZ11] Yu Miao, Ke Wang, and Fangfang Zhao. Some limit behaviors for the LS estimator in simple linear EV regression models. Statistics & Probability Letters, 81(1):92–102, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002737.

Ma:2017:MDN

[MX17a] Xiaocui Ma and Fubao Xi. Moderate deviations for neutral stochastic differential delay equations with jumps. Statistics & Probability Letters, 126(??):97–107, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300901.

McRedmond:2017:EDP

[MX17b] James McRedmond and Chang Xu. On the expected diameter of planar Brownian motion. Statistics & Probability Letters, 130(??):1–4, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302341.

Mu:2018:CSF

[MX18] Weiyan Mu and Shifeng Xiong. A class of space-filling designs and their projection properties. Statistics & Probability Letters, 141(??):129–134, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302256.

Mao:2018:CTV

[MXZZ18] Yong-Hua Mao, Liping Xu, Ming Zhang, and Yu-Hui Zhang. Convergence in total variation distance for (in)homogeneous Markov processes. Statistics & Probability Letters, 137(??):54– 62, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300130.

Misagh:2011:WIE

[MY11] F. Misagh and G. H. Yari. On weighted interval entropy. Statistics & Probability Letters, 81(2):188–194, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521000324X.

Ma:2014:SNF

[MY14] Chunhua Ma and Xu Yang. Small noise fluctuations of the CIR model driven by α-stable noises. Statistics & Probability Letters, 94(??):1–11, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002375.

Mudholkar:2015:MCG

[MYA15] Govind S. Mudholkar, Ziji Yu, and Saria S. Awadalla. The modecentric M-Gaussian distribution: A model for right skewed data. Statistics & Probability Letters, 107:1–10, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002497.

Myronyuk:2019:CDI

[Myr19] Margaryta Myronyuk. Characterization of distributions of Q-independent random variables on locally compact Abelian groups. Statistics & Probability Letters, 152(??):82–88, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301130.

Mao:2017:SSD

[MZ17] Yong-Hua Mao and Pan Zhao. Strong stationary duality for discrete time Möbius monotone Markov chains on Z_{+}^{d}. Statistics & Probability Letters, 123(??):183–192, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302942.

Mackey:2019:DBS

[MZ19] Benjamin Mackey and Dapeng Zhan. Decomposition of backward SLE in the capacity parametrization. Statistics & Probability Letters, 146(??):27–35, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303456.

Ma:2015:CEE

[MZZ15] Huijuan Ma, Feipeng Zhang, and Yong Zhou. Composite estimating equation approach for additive risk model with lengthbiased and right-censored data. Statistics & Probability Letters, 96(??):45–53, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003095.

Naik:2013:FDE

[NA13] Shanoja R. Naik and Bovas Abraham. The fractional-diffusion equation and a new distribution to model positively skewed data with heavy tails. Statistics & Probability Letters, 83(7):1759– 1769, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001041.

Nadarajah:2013:DGP

[NAC13] Saralees Nadarajah, Emmanuel Afuecheta, and Stephen Chan. A double generalized Pareto distribution. Statistics & Probability Letters, 83(12):2656–2663, December 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002903.

Nadarajah:2013:EBE

[Nad13] Saralees Nadarajah. Expansions for bivariate extreme value distributions. Statistics & Probability Letters, 83(3):744–752, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004270.

Nadarajah:2015:CAE

[Nad15a] Saralees Nadarajah. Complete asymptotic expansions for normal extremes. Statistics & Probability Letters, 103(??):127– 133, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001388.

Nadarajah:2015:EBC

[Nad15b] Saralees Nadarajah. Expansions for bivariate copulas. Statistics & Probability Letters, 100(??):77–84, May 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000553.

Nadarajah:2016:AEB

[Nad16] Saralees Nadarajah. Asymptotic expansions for bivariate normal extremes. Statistics & Probability Letters, 119(??):124–133, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301456.

Nagaraja:2013:MOS

[Nag13] H. N. Nagaraja. Moments of order statistics and L-moments for the symmetric triangular distribution. Statistics & Probability Letters, 83(10):2357–2363, October 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300240X.

Nagy:2017:IDM

[Nag17a] Stanislav Nagy. Integrated depth for measurable functions and sets. Statistics & Probability Letters, 123(??):165–170, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216303066.

Nagy:2017:MPS

[Nag17b] Stanislav Nagy. Monotonicity properties of spatial depth. Statistics & Probability Letters, 129(??):373–378, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302298.

Nagler:2018:GAN

[Nag18] Thomas Nagler. A generic approach to nonparametric function estimation with mixed data. Statistics & Probability Letters, 137(??):326–330, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300853.

Nagy:2019:SHD

[Nag19] Stanislav Nagy. Scatter halfspace depth for K-symmetric distributions. Statistics & Probability Letters, 149(??):171–177, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300586.

Nakamura:2013:QID

[Nak13a] Takashi Nakamura. A quasi-infinitely divisible characteristic function and its exponentiation. Statistics & Probability Letters, 83(10):2256–2259, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002204.

Nakatsu:2013:ACL

[Nak13b] Tomonori Nakatsu. Absolute continuity of the laws of a multi-dimensional stochastic differential equation with coefficients dependent on the maximum. Statistics & Probability Letters, 83(11):2499–2506, November 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002605.

Nakata:2015:LDG

[Nak15] Toshio Nakata. Limit distributions of generalized St. Petersburg games. Statistics & Probability Letters, 96(??):307–314, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003599.

Nakata:2016:WLL

[Nak16] Toshio Nakata. Weak laws of large numbers for weighted independent random variables with infinite mean. Statistics & Probability Letters, 109:124–129, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003855.

Nakata:2017:NAM

[Nak17] Toshio Nakata. A note on the asymptotics of the maxima for the St. Petersburg game. Statistics & Probability Letters, 129(??):284–287, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302146.

Nanda:2010:CDT

[Nan10] Asok K. Nanda. Characterization of distributions through failure rate and mean residual life functions. Statistics & Probability Letters, 80(9–10):752–755, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000118.

Nasari:2012:SLL

[Nas12] Masoud M. Nasari. Strong Law of Large Numbers for weighted U-statistics: Application to incomplete U-statistics. Statistics & Probability Letters, 82(6):1208–1217, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200082X.

Navarro:2014:CBM

[Nav14] Jorge Navarro. Can the bounds in the multivariate Chebyshev inequality be attained? Statistics & Probability Letters, 91(??): 1–5, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001217.

Navarro:2018:PDI

[Nav18] Jorge Navarro. Preservation of DMRL and IMRL aging classes under the formation of order statistics and coherent systems. Statistics & Probability Letters, 137(??):264–268, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830049X.

Ni:2016:HPC

[NC16] Wenqing Ni and Zhenlong Chen. Hitting probabilities of a class of Gaussian random fields. Statistics & Probability Letters, 118(??):145–155, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216300943.

Nov:2010:MNE

[ND10] Yuval Nov and Ori Davidov. Minimum-norm estimation for a bi-exponential survival model. Statistics & Probability Letters, 80(7–8):648–653, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004805.

Nanda:2012:SOM

[ND12] Asok K. Nanda and Suchismita Das. Stochastic orders of the Marshall–Olkin extended distribution. Statistics & Probability Letters, 82(2):295–302, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003245.

Nilsson:2012:BCI

[NdBC12] William Nilsson and Tomás del Barrio Castro. Bootstrap confidence interval for a correlation curve. Statistics & Probability Letters, 82(1):1–6, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002884.

Neunhauserer:2013:ACR

[Neu13] J. Neunhäuserer. Absolutely continuous random power series in reciprocals of Pisot numbers. Statistics & Probability Letters, 83(2):431–435, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003963.

Neunhauserer:2017:RFR

[Neu17] Jörg Neunhäuserer. Return of Fibonacci random walks. Statistics & Probability Letters, 121(??):51–53, February 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301917.

Nguyen:2016:ALB

[Ngu16] Dao Nguyen. Another look at Bayes map iterated filtering. Statistics & Probability Letters, 118(??):32–36, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300633.

Niu:2019:LST

[NHBG19] Zhenzhen Niu, Jiang Hu, Zhidong Bai, and Wei Gao. On LR simultaneous test of high-dimensional mean vector and covariance matrix under non-normality. Statistics & Probability Letters, 145(??):338–344, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303249.

Nikulin:2011:VFA

[NHN^{+}11] Vladimir Nikulin, Tian-Hsiang Huang, Shu-Kay Ng, Suren I. Rathnayake, and Geoffrey J. McLachlan. A very fast algorithm for matrix factorization. Statistics & Probability Letters, 81(7):

773–782, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000393.

Nicolau:2017:SNM

[Nic17] João Nicolau. A simple nonparametric method to estimate the expected time to cross a threshold. Statistics & Probability Letters, 123(??):146–152, April 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302954.

Nielsen:2016:MCC

[Nie16] Adam Nielsen. The Monte Carlo computation error of transition probabilities. Statistics & Probability Letters, 118(??):163–170, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216300931.

Niemiro:2019:FRP

[Nie19] Wojciech Niemiro. Fixed relative precision estimators of growth rate for compound Poisson and lévy processes. Statistics & Probability Letters, 153(??):151–156, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301671.

Nanda:2011:CTR

[NK11] Asok K. Nanda and Amarjit Kundu. Comparison of two repairable systems. Statistics & Probability Letters, 81(3):446– 450, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210003482.

Nagase:2017:INS

[NK17] Mario Nagase and Yutaka Kano. Identifiability of nonrecursive structural equation models. Statistics & Probability Letters, 122(??):109–117, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302504.

Nagatsuka:2013:EFS

[NKKY13] Hideki Nagatsuka, Hiroshi Kawakami, Toshinari Kamakura, and Hisashi Yamamoto. The exact finite-sample distribution of the median absolute deviation about the median of continuous random variables. Statistics & Probability Letters, 83(4):999– 1005, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004828.

Nkurunziza:2011:SSS

[Nku11] Sévérien Nkurunziza. Shrinkage strategy in stratified random sample subject to measurement error. Statistics & Probability Letters, 81(2):317–325, February 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003068.

Ngo:2019:TEM

[NL19] Hoang Long Ngo and Duc Trong Luong. Tamed Euler– Maruyama approximation for stochastic differential equations with locally Hölder continuous diffusion coefficients. Statistics & Probability Letters, 145(??):133–140, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303055.

Nott:2011:ISV

[NLF11] David J. Nott, Jialiang Li, and Mark Fielding. Importance sampling as a variational approximation. Statistics & Probability Letters, 81(8):1052–1055, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000745.

Nguyen:2018:CAE

[NM18] Hien D. Nguyen and Geoffrey J. McLachlan. Chunkedand-averaged estimators for vector parameters. Statistics & Probability Letters, 137(??):336–342, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300968.

Ng:2019:GRD

[NM19] Tin Lok James Ng and Thomas Brendan Murphy. Generalized random dot product graph. Statistics & Probability Letters, 148(??):143–149, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300173.

Nishimura:2015:MEC

[NMS15] Kazuya Nishimura, Shun Matsuura, and Hideo Suzuki. Multivariate EWMA control chart based on a variable selection using AIC for multivariate statistical process monitoring. Statistics & Probability Letters, 104(??):7–13, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001522.

Nguyen:2016:LMA

[NMUJ16] Hien D. Nguyen, Geoffrey J. McLachlan, Jeremy F. P. Ullmann, and Andrew L. Janke. Laplace mixture autoregressive models. Statistics & Probability Letters, 110(??):18–24, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215300110.

Nogales:2013:ERC

[Nog13] A. G. Nogales. The existence of regular conditional probabilities for Markov kernels. Statistics & Probability Letters, 83(3):891– 897, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004579.

Nolan:2018:TFM

[Nol18] John P. Nolan. Truncated fractional moments of stable laws. Statistics & Probability Letters, 137(??):312–318, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300543.

Noste:2019:UDI

[Nos19] María Eugenia Szretter Noste. Using DAGs to identify the sufficient dimension reduction in the Principal Fitted Components model. Statistics & Probability Letters, 145(??):317–320, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302839.

Novak:2017:LLH

[Nov17] S. Y. Novak. On the length of the longest head run. Statistics & Probability Letters, 130(??):111–114, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302249.

Nowak:2016:MME

[Now16] Piotr Boleslaw Nowak. The MLE of the mean of the exponential distribution based on grouped data is stochastically increasing. Statistics & Probability Letters, 111(??):49–54, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521530064X.

Nowak:2019:MTE

[Now19] Piotr Bolesław Nowak. Moment-type estimation from grouped samples. Statistics & Probability Letters, 149(??):80–85, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300161.

Neves:2010:DFR

[NP10] Cláudia Neves and António Pereira. Detecting finiteness in the right endpoint of light-tailed distributions. Statistics & Probability Letters, 80(5–6):437–444, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004453.

Nedenyi:2016:ISL

[NP16] Fanni Nedényi and Gyula Pap. Iterated scaling limits for aggregation of random coefficient AR(1) and INAR(1) processes. Statistics & Probability Letters, 118(??):16–23, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300840.

Nastic:2012:SGM

[NR12] Aleksandar S. Nastić and Miroslav M. Ristić. Some geometric mixed integer-valued autoregressive (INAR) models. Statistics & Probability Letters, 82(4):805–811, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521200017X.

Navarro:2010:RBC

[NS10] Jorge Navarro and Fabio Spizzichino. On the relationships between copulas of order statistics and marginal distributions. Statistics & Probability Letters, 80(5–6):473–479, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004490.

Nayak:2012:SAM

[NS12] Tapan K. Nayak and Bimal Sinha. Some aspects of minimum variance unbiased estimation in presence of ancillary statistics. Statistics & Probability Letters, 82(6):1129–1135, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000715.

Nakai:2013:MFG

[NS13a] Eiichi Nakai and Gaku Sadasue. Maximal function on generalized martingale Lebesgue spaces with variable exponent. Statistics & Probability Letters, 83(10):2168–2171, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002125.

Novikov:2013:RMI

[NS13b] Alexander Novikov and Albert Shiryaev. Remarks on moment inequalities and identities for martingales. Statistics & Probability Letters, 83(4):1260–1261, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000126.

Nevzorov:2014:RC

[NS14] V. B. Nevzorov and A. Stepanov. Records with confirmation. Statistics & Probability Letters, 95(??):39–47, ????

2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002818.

Nie:2017:UER

[NSH17] Keyu Nie, Bikas K. Sinha, and A. S. Hedayat. Unbiased estimation of reliability function from a mixture of two exponential distributions based on a single observation. Statistics & Probability Letters, 127(??):7–13, August 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301268.

Navarro:2011:CBM

[NSL11] J. Navarro, S. M. Sunoj, and M. N. Linu. Characterizations of bivariate models using dynamic Kullback–Leibler discrimination measures. Statistics & Probability Letters, 81(11):1594–1598, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521100191X.

Nanda:2014:RRE

[NSS14] Asok K. Nanda, P. G. Sankaran, and S. M. Sunoj. Rényi’s residual entropy: a quantile approach. Statistics & Probability Letters, 85(??):114–121, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003970.

Ning:2014:ETP

[NT14] Zijun Ning and Linjun Tang. Estimation and test procedures for composite quantile regression with covariates missing at random. Statistics & Probability Letters, 95(??):15–25, ???? 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521400282X.

Ngo:2017:SCE

[NT17] Hoang-Long Ngo and Dai Taguchi. Strong convergence for the Euler–Maruyama approximation of stochastic differential equations with discontinuous coefficients. Statistics & Probability Letters, 125(??):55–63, June 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300494.

Nane:2018:RRA

[NTT18] Erkan Nane, Nguyen Hoang Tuan, and Nguyen Huy Tuan. A random regularized approximate solution of the inverse problem for Burgers’ equation. Statistics & Probability Letters, 132(??):46–54, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302742.

Nelsen:2012:HCP

[NÚF12] Roger B. Nelsen and Manuel Úbeda-Flores. How close are pairwise and mutual independence? Statistics & Probability Letters, 82(10):1823–1828, October 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002179.

Nair:2011:ACA

[NV11] N. Unnikrishnan Nair and B. Vineshkumar. Ageing concepts: An approach based on quantile function. Statistics & Probability Letters, 81(12):2016–2025, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002720.

Nashimoto:2010:MTS

[NW10a] Kane Nashimoto and F. T. Wright. Marcus–Talpaz simultaneous lower confidence bounds for contrasts between treatment and control means. Statistics & Probability Letters, 80(2):128–133, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771520900368X.

Niu:2010:CLT

[NW10b] Yingxuan Niu and Yi Wang. The Central Limit Theorem and ergodicity. Statistics & Probability Letters, 80(15–16):1180–1184, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000945.

Narayanan:2013:MDA

[NW13] Rajendran Narayanan and Martin T. Wells. On the maximal domain of attraction of Tracy–Widom distribution for Gaussian unitary ensembles. Statistics & Probability Letters, 83(10):2364– 2371, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002411.

Noonan:2019:ASC

[NZ19] Jack Noonan and Anatoly Zhigljavsky. Approximating Shepp’s constants for the Slepian process. Statistics & Probability Letters, 153(??):21–31, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301300.

Ozkale:2015:FOC

[ÖA15] M. Revan Özkale and Engin Arican. First-order r–d class estimator in binary logistic regression model. Statistics & Probability Letters, 106(??):19–29, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002072.

Oz:2013:TPA

[ÖÇ13] Mehmet Öz and Mine Çaglar. Tail probability of avoiding Poisson traps for branching Brownian motion. Statistics & Probability Letters, 83(9):2034–2038, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001776.

Oesting:2015:DMS

[Oes15] Marco Oesting. On the distribution of a max-stable process conditional on max-linear functionals. Statistics & Probability Letters, 100(??):158–163, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000528.

Ogasawara:2017:EPI

[Oga17] Haruhiko Ogasawara. Extensions of Pearson’s inequality between skewness and kurtosis to multivariate cases. Statistics & Probability Letters, 130(??):12–16, November 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302365.

Ogasawara:2018:ISF

[Oga18] Haruhiko Ogasawara. The inverse survival function for multivariate distributions and its application to the product moment. Statistics & Probability Letters, 142(??):71–76, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302530.

Oh:2014:BCM

[Oh14] Man-Suk Oh. Bayesian comparison of models with inequality and equality constraints. Statistics & Probability Letters, 84(??):176–182, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003398.

Ogundimu:2015:ETP

[OH15] Emmanuel O. Ogundimu and Jane L. Hutton. On the extended two-parameter generalized skew-normal distribution. Statistics & Probability Letters, 100(??):142–148, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000668.

Ohyama:2013:PVI

[Ohy13] Tetsuji Ohyama. Prior value incorporated calibration estimator in stratified random sampling. Statistics & Probability Letters, 83(1):46–51, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003288.

Okolewski:2015:SEI

[OK15] A. Okolewski and M. Kaluszka. Stability of expected Lstatistics against weak dependence of observations. Statistics & Probability Letters, 106(??):157–164, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002606.

Otneim:2013:BBL

[OKT13] Håkon Otneim, Hans Arnfinn Karlsen, and Dag Tjøstheim. Bias and bandwidth for local likelihood density estimation. Statistics & Probability Letters, 83(5):1382–1387, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000424.

Olkin:2014:DIC

[Olk14] Ingram Olkin. A determinantal inequality for correlation matrices. Statistics & Probability Letters, 88(??):88–90, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000236.

Onskog:2013:EPU

[Öns13] Thomas Önskog. Existence of pathwise unique Langevin processes on polytopes with perfect reflection at the boundary. Statistics & Probability Letters, 83(10):2211–2219, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002010.

Onzon:2011:MCR

[Onz11] Emmanuel Onzon. Multivariate Cramér-Rao inequality for prediction and efficient predictors. Statistics & Probability Letters, 81(3):429–437, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210003470.

Oral:2011:RAR

[OO11] Evrim Oral and Ece Oral. A robust alternative to the ratio estimator under non-normality. Statistics & Probability Letters, 81(8):930–936, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001313.

Orsingher:2012:SFP

[OP12] Enzo Orsingher and Federico Polito. The space-fractional Poisson process. Statistics & Probability Letters, 82(4):852– 858, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521100410X.

Orsingher:2013:IFP

[OP13] Enzo Orsingher and Federico Polito. On the integral of fractional Poisson processes. Statistics & Probability Letters, 83(4):1006– 1017, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004750.

Oh:2016:BVS

[OPS16] Man-Suk Oh, Eun Sug Park, and Beong-Soo So. Bayesian variable selection in binary quantile regression. Statistics & Probability Letters, 118(??):177–181, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521530242X.

Ou:2010:SAI

[OQ10] Zujun Ou and Hong Qin. Some applications of indicator function in two-level factorial designs. Statistics & Probability Letters, 80(1):19–25, January 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209003484.

Otiniano:2015:IFM

[ORO15] C. E. G. Otiniano, P. N. Rathie, and L. C. S. M. Ozelim. On the identifiability of finite mixture of Skew-Normal and Skew-t distributions. Statistics & Probability Letters, 106(??):103–108, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002485.

Ostrovska:2010:NPP

[OS10]  Sofiya Ostrovska and Jordan Stoyanov.  A new proof

that the product of three or more exponential random variables is moment-indeterminate. Statistics & Probability Letters, 80(9–10):792–796, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000179.

Ohashi:2015:NSC

[OS15] Alberto Ohashi and Alexandre B. Simas. A note on the sharp L^{p}-convergence rate of upcrossings to the Brownian local time. Statistics & Probability Letters, 100(??):137–141, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000504.

Oshime:2018:PIR

[OS18] Takayoshi Oshime and Yasutaka Shimizu. Parametric inference for ruin probability in the classical risk model. Statistics & Probability Letters, 133(??):28–37, February 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303140.

Osekowski:2010:LEN

[Osȩ10a] Adam Osȩkowski. Logarithmic estimates for nonsymmetric martingale transforms. Statistics & Probability Letters, 80 (7–8):678–682, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209004842.

Osekowski:2010:SMB

[Osȩ10b] Adam Osȩkowski. Sharp maximal bound for continuous martingales. Statistics & Probability Letters, 80(17–18):1405– 1408, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210001410.

Osekowski:2010:WTI

[Osȩ10c] Adam Osȩkowski. Weak type inequalities for conditionally symmetric martingales. Statistics & Probability Letters, 80(23–24):2009–2013, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002579.

Osekowski:2011:SMI

[Osȩ11] Adam Osȩkowski. Sharp maximal inequality for nonnegative martingales. Statistics & Probability Letters, 81(12):1945–1952, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002732.

Osekowski:2012:BCW

[Osȩ12a] Adam Osȩkowski. Best constants in the weak type inequalities for a martingale conditional square function. Statistics & Probability Letters, 82(5):885–893, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200034X.

Osekowski:2012:IMT

[Osȩ12b] Adam Osȩkowski. Inequalities for martingale transforms and related characterizations of Hilbert spaces. Statistics & Probability Letters, 82(1):186–190, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100318X.

Osekowski:2012:WNI

[Osȩ12c] Adam Osȩkowski. Weak norm inequalities for martingales and geometry of Banach spaces. Statistics & Probability Letters, 82(3):411–418, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003890.

Osekowski:2013:MWE

[Osȩ13a] Adam Osȩkowski. On martingales whose exponential processes satisfy Muckenhoupt’s condition A_{1}. Statistics & Probability Letters, 83(8):1849–1853, August 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001338.

Osekowski:2013:PIM

[Osȩ13b] Adam Osȩkowski. A prophet inequality for L^{2}-martingales. Statistics & Probability Letters, 83(10):2319–2323, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521300237X.

Osekowski:2014:IMT

[Osȩ14a] Adam Osȩkowski. Inequalities for martingales taking values in 2-convex Banach spaces. Statistics & Probability Letters, 84(??):102–107, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003301.

Osekowski:2014:SIH

[Osȩ14b] Adam Osȩkowski. Sharp L^{2} log L inequalities for the Haar system and martingale transforms. Statistics & Probability Letters, 94(??):91–97, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002429.

Osekowski:2014:WTI

[Osȩ14c] Adam Osȩkowski. A weak-type inequality for the martingale square function. Statistics & Probability Letters, 95(??):139–143, December 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003009.

Osekowski:2015:SES

[Osȩ15a] Adam Osȩkowski. Sharp ^{L}^{1}(ℓ^{q}) estimate for a sequence and its predictable projection. Statistics & Probability Letters, 104(??):82–86, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521500156X.

Osekowski:2015:SMI

[Osȩ15b] Adam Osȩkowski. A sharp maximal inequality for onedimensional Dunkl martingales. Statistics & Probability Letters, 105(??):114–119, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001947.

Osxcxakowski:2017:FSI

[Osȩ17a] Adam Osȩkowski. A Fefferman–Stein inequality for the martingale square and maximal functions. Statistics & Probability Letters, 129(??):81–85, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301839.

Osekowski:2017:WIM

[Osȩ17b] Adam Osȩkowski. Weighted inequalities for the martingale square and maximal functions. Statistics & Probability Letters, 120(??):95–100, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215300146.

Osekowski:2019:AHT

[Ose19] Adam Osekowski. On the action of the Hilbert transform on ℓ_{1}^{N}- valued functions. Statistics & Probability Letters, 146(??):213– 223, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303778.

Ostrovski:2013:SNA

[Ost13] Vladimir Ostrovski. Stability of no-arbitrage property under model uncertainty. Statistics & Probability Letters, 83(1):89– 92, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003318.

Ostrovski:2017:TEM

[Ost17] Vladimir Ostrovski. Testing equivalence of multinomial distributions. Statistics & Probability Letters, 124(??):77–82, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300111.

Ostrovski:2018:TEF

[Ost18] Vladimir Ostrovski. Testing equivalence to families of multinomial distributions with application to the independence model. Statistics & Probability Letters, 139(??):61–66, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301342.

Osuna:2012:CPD

[Osu12] Edgar Elias Osuna. Crossing points of distributions and a theorem that relates them to second order stochastic dominance. Statistics & Probability Letters, 82(4):758–764, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003920.

Ouyang:2018:MIR

[OSW18] Cheng Ouyang, Yinghui Shi, and Dongsheng Wu. Mutual intersection for rough differential systems driven by fractional Brownian motions. Statistics & Probability Letters, 135(??):83– 91, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303668.

Olkin:2015:CBB

[OT15] Ingram Olkin and Thomas A. Trikalinos. Constructions for a bivariate beta distribution. Statistics & Probability Letters, 96(??):54–60, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003241.

Ostrovska:2019:SCS

[OT19] Sofiya Ostrovska and Mehmet Turan. q-Stieltjes classes for some families of q-densities. Statistics & Probability Letters, 146(??):118–123, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303602.

Ouadah:2013:UBN

[Oua13] Sarah Ouadah. Uniform-in-bandwidth nearest-neighbor density estimation. Statistics & Probability Letters, 83(8):1835– 1843, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001351.

Ouyang:2010:GC

[Ouy10] Shun-Xiang Ouyang. Gluing and coupling. Statistics & Probability Letters, 80(15–16):1196–1199, August 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000969.

Omey:2015:IAD

[OV15] Edward Omey and Stefan Van Gulck. Intuitive approximations in discrete renewal theory, part 1: Regularly varying case. Statistics & Probability Letters, 104(??):68–74, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001510.

Olhede:2018:FSD

[OW18] Sofia C. Olhede and Patrick J. Wolfe. The future of statistics and data science. Statistics & Probability Letters, 136(??):46– 50, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300877.

Owo:2015:BDS

[Owo15] Jean-Marc Owo. Backward doubly stochastic differential equations with stochastic Lipschitz condition. Statistics & Probability Letters, 96(??):75–84, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400323X.

Oz:2016:SBB

[Öz16] Mehmet Öz. Survival of branching Brownian motion in a uniform trap field. Statistics & Probability Letters, 110(??):211– 216, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215003624.

Pacheco:2016:RMS

[Pac16] Carlos G. Pacheco. A random matrix from a stochastic heat equation. Statistics & Probability Letters, 113(??):71–78, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000419.

Pacheco:2017:RES

[Pac17] Carlos G. Pacheco. Random eigenvalues from a stochastic heat equation. Statistics & Probability Letters, 123(??):114–121, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302929.

Packham:2018:OCU

[Pac18] N. Packham. Optimal contracts under competition when uncertainty from adverse selection and moral hazard are present. Statistics & Probability Letters, 137(??):99–104, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300166.

Pakes:2013:LLU

[Pak13] Anthony G. Pakes. Limit laws for UGROW random graphs. Statistics & Probability Letters, 83(12):2607–2614, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002836.

Pakes:2018:LFN

[Pak18] Anthony G. Pakes. The Lambert W function, Nuttall’s integral, and the Lambert law. Statistics & Probability Letters, 139(??): 53–60, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301354.

Pamen:2018:SAS

[Pam18] Olivier Menoukeu Pamen. On some applications of Sobolev flows of SDEs with unbounded drift coefficients. Statistics & Probability Letters, 141(??):114–124, October 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302165.

Panov:2017:LTS

[Pan17] Vladimir Panov. Limit theorems for sums of random variables with mixture distribution. Statistics & Probability Letters, 129(??):379–386, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302213.

Panov:2019:SPO

[Pan19] Vladimir Panov. Some properties of the one-dimensional subordinated stable model. Statistics & Probability Letters, 146(??): 80–84, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303493.

Papanicolaou:2016:AL

[Pap16a] Vassilis G. Papanicolaou. An arctangent law. Statistics & Probability Letters, 116(??):62–64, September 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300335.

Papastathopoulos:2016:CIC

[Pap16b] Ioannis Papastathopoulos. Conditional independence and conditioned limit laws. Statistics & Probability Letters, 112(??): 1–4, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521530016X.

Park:2014:SEN

[Par14] Junyong Park. Shrinkage estimator in normal mean vector estimation based on conditional maximum likelihood estimators. Statistics & Probability Letters, 93(??):1–6, October 1–142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002077.

Park:2017:SPQ

[Par17a] Jinho Park. Solution path for quantile regression with epsiloninsensitive loss in a reproducing kernel Hilbert space. Statistics & Probability Letters, 126(??):205–211, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300974.

Park:2017:TIR

[Par17b] Junyong Park. Tolerance intervals from ridge regression in the presence of multicollinearity and high dimension. Statistics & Probability Letters, 121(??):128–135, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302152.

Parvardeh:2013:CRL

[PB13] A. Parvardeh and N. Balakrishnan. Conditional residual lifetimes of coherent systems. Statistics & Probability Letters, 83(12):2664–2672, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521300285X.

Parvardeh:2015:MSM

[PB15a] A. Parvardeh and N. Balakrishnan. On mixed δ-shock models. Statistics & Probability Letters, 102(??):51–60, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001169.

Pati:2015:ABI

[PB15b] Debdeep Pati and Anirban Bhattacharya. Adaptive Bayesian inference in the Gaussian sequence model using exponentialvariance priors. Statistics & Probability Letters, 103(??):100– 104, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001224.

Pal:2016:DNB

[PB16] Suvra Pal and N. Balakrishnan. Destructive negative binomial cure rate model and EM-based likelihood inference under Weibull lifetime. Statistics & Probability Letters, 116(??):9–20, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216300232.

Perez-Casany:2014:RMP

[PCGV14] M. Pérez-Casany, J. Ginebra, and J. Valero. On recovering a mixed Poisson distribution from its left-truncated version. Statistics & Probability Letters, 92(??):89–94, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001849.

Petrovic:2011:ISC

[PD11] Ljiljana Petrović and Sladjana Dimitrijević. Invariance of statistical causality under convergence. Statistics & Probability Letters, 81(9):1445–1448, September 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001611.

Petrovic:2012:CFH

[PD12] Ljiljana Petrović and Sladjana Dimitrijević. Causality with finite horizon of the past in continuous time. Statistics & Probability Letters, 82(7):1219–1223, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001253.

Peng:2010:DPE

[PDW10] Hanxiang Peng, Xin Dang, and Xueqin Wang. The distribution of partially exchangeable random variables. Statistics & Probability Letters, 80(11–12):932–938, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521000043X.

Pei:2015:LSI

[Pei15] Ling Pei. Logarithmic Sobolev inequality on free path space over a compact Riemannian manifold. Statistics & Probability Letters, 98(??):12–19, March 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004076.

Peng:2011:UHE

[Pen11] Qidi Peng. Uniform Hölder exponent of a stationary increments Gaussian process: Estimation starting from average values. Statistics & Probability Letters, 81(8):1326–1335, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001271.

Peng:2014:NFP

[Pen14] Jun Peng. A note on the first passage time of diffusions with holding and jumping boundary. Statistics & Probability Letters, 93(??):58–64, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002144.

Peterson:2011:CDP

[Pet11] Lisa D. Peterson. Convergence in distribution of point processes on Polish spaces to a simple limit. Statistics & Probability Letters, 81(12):1859–1861, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002495.

Peyhardi:2017:CCS

[PF17] Jean Peyhardi and Pierre Fernique. Characterization of convolution splitting graphical models. Statistics & Probability Letters, 126(??):59–64, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300743.

Pawlas:2010:CLI

[PH10a] Zbynek Pawlas and Ondrej Honzl. Comparison of lengthintensity estimators for segment processes. Statistics & Probability Letters, 80(9–10):825–833, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000283.

Peng:2010:RPO

[PH10b] Jiangyan Peng and Jin Huang. Ruin probability in a onesided linear model with constant interest rate. Statistics & Probability Letters, 80(7–8):662–669, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004829.

Pai:2013:ABM

[PH13a] Hui-Ming Pai and Chii-Ruey Hwang. Accelerating Brownian motion on N-torus. Statistics & Probability Letters, 83(5):1443– 1447, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000527.

Preuss:2013:CSD

[PH13b] Philip Preuß and Thimo Hildebrandt. Comparing spectral densities of stationary time series with unequal sample sizes. Statistics & Probability Letters, 83(4):1174–1183, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000163.

Ping:2012:AOS

[Pin12] Sun Ping. The application of order statistics to multiple integration over a canonical simplex. Statistics & Probability Letters, 82(9):1641–1647, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001940.

Pinelis:2013:ERT

[Pin13] Iosif Pinelis. Exact Rosenthal-type inequalities for p = 3, and related results. Statistics & Probability Letters, 83(12):2634–2637, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002873.

Pingel:2014:SAL

[Pin14] Ronnie Pingel. Some approximations of the logistic distribution with application to the covariance matrix of logistic regression. Statistics & Probability Letters, 85(??):63–68, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521300388X.

Pinelis:2015:CFP

[Pin15a] Iosif Pinelis. Characteristic function of the positive part of a random variable and related results, with applications. Statistics & Probability Letters, 106(??):281–286, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002722.

Pinelis:2015:STN

[Pin15b] Iosif Pinelis. On the supremum of the tails of normalized sums of independent Rademacher random variables. Statistics & Probability Letters, 99(??):131–134, April 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000164.

Pinelis:2017:CBP

[Pin17] Iosif Pinelis. Contrast between populations versus spread within populations. Statistics & Probability Letters, 121(??):99–100, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216302103.

Peng:2010:SCT

[PJ10] Shiguo Peng and Baoguo Jia. Some criteria on p-th moment stability of impulsive stochastic functional differential equations. Statistics & Probability Letters, 80(13–14):1085–1092, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000726.

Persing:2013:LCH

[PJ13] Adam Persing and Ajay Jasra. Likelihood computation for hidden Markov models via generalized two-filter smoothing. Statistics & Probability Letters, 83(5):1433–1442, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000485.

Park:2011:QRE

[PK11] Jinho Park and Jeankyung Kim. Quantile regression with an epsilon-insensitive loss in a reproducing kernel Hilbert space. Statistics & Probability Letters, 81(1):62–70, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002695.

Park:2014:CRE

[PK14] Sangun Park and Ilmun Kim. On cumulative residual entropy of order statistics. Statistics & Probability Letters, 94 (??):170–175, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214002582.

Park:2014:IOB

[PKK14] Yousung Park, Daeyoung Kim, and Seongyong Kim. Identification of the occurrence of boundary solutions in a contingency table with nonignorable nonresponse. Statistics & Probability Letters, 93(??):34–40, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002132.

Peng:2015:GMM

[PL15a] Lihua Peng and Junping Li. A generalization of Φ-moment martingale inequalities. Statistics & Probability Letters, 102(??): 61–68, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500098X.

Peng:2015:SOA

[PL15b] Zuoxiang Peng and Xin Liao. Second-order asymptotics for convolution of distributions with light tails. Statistics & Probability Letters, 106(??):199–208, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002680.

Peng:2017:FTP

[PL17] Xiaofan Peng and Li Luo. Finite time Parisian ruin of an integrated Gaussian risk model. Statistics & Probability Letters, 124(??):22–29, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300056.

Palacio:2018:ICC

[PL18] Alan Riva Palacio and Fabrizio Leisen. Integrability conditions for compound random measures. Statistics & Probability Letters, 135(??):32–37, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303516.

Park:2019:NAR

[PL19] Seongoh Park and Johan Lim. Non-asymptotic rate for highdimensional covariance estimation with non-independent missing observations. Statistics & Probability Letters, 153(??):113– 123, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301543.

Pan:2013:SOV

[PLH13] Xiaoqing Pan, Xuan Leng, and Taizhong Hu. The secondorder version of Karamata’s theorem with applications. Statistics & Probability Letters, 83(5):1397–1403, May 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000497.

Peng:2013:DPD

[PLL13] Dan Peng, Donghai Liu, and Zaiming Liu. Dividend problems in the dual risk model with exponentially distributed observation time. Statistics & Probability Letters, 83(3):841–849, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004415.

Peng:2016:EPS

[PLN16] Zuoxiang Peng, Chunqiao Li, and Saralees Nadarajah. Extremal properties of the skew-t distribution. Statistics & Probability Letters, 112(??):10–19, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000134.

Pan:2018:NRL

[PLY18] Donghui Pan, Jia-Bao Liu, and Wenzhi Yang. A new result on lifetime estimation based on skew-Wiener degradation model. Statistics & Probability Letters, 138(??):157–164, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301238.

Park:2012:MIO

[PN12] Sangun Park and Hon Keung Tony Ng. Missing information and an optimal one-step plan in a Type II progressive censoring scheme. Statistics & Probability Letters, 82(2):396–402, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003336.

Pham:2014:SBB

[PN14] Tung-Dinh Pham and Nam-Ky Nguyen. Small Box–Behnken designs with orthogonal blocks. Statistics & Probability Letters, 85(??):84–90, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003854.

Pimentel:2015:AZI

[PNBW15] Ronald S. Pimentel, Magdalena Niewiadomska-Bugaj, and Jung-Chao Wang. Association of zero-inflated continuous variables. Statistics & Probability Letters, 96(??):61–67, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003137.

Park:2015:FID

[PNC15] Sangun Park, Hon Keung Tony Ng, and Ping Shing Chan. On the Fisher information and design of a flexible progressive censored experiment. Statistics & Probability Letters, 97(??):142– 149, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003952.

Patch:2015:CTC

[PNT15] Brendan Patch, Yoni Nazarathy, and Thomas Taimre. A correction term for the covariance of renewal-reward processes with multivariate rewards. Statistics & Probability Letters, 102(??): 1–7, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000942.

Penda:2018:MDP

[PO18] S. Valère Bitseki Penda and Adélaïde Olivier. Moderate deviation principle in nonlinear bifurcating autoregressive models. Statistics & Probability Letters, 138(??):20–26, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300828.

Pogany:2016:IFC

[Pog16] Tibor K. Pogány. Integral form of the COM-Poisson renormalization constant. Statistics & Probability Letters, 119(??):144– 145, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301249.

Poinas:2019:BMC

[Poi19] Arnaud Poinas. A bound of the β-mixing coefficient for point processes in terms of their intensity functions. Statistics & Probability Letters, 148(??):88–93, May 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218304012.

Pollak:2013:SDP

[Pol13]  Moshe Pollak.  A stochastic dominance property com-

mon to the boy-or-girl paradox and the lottery. Statistics & Probability Letters, 83(1):410–413, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200377X.

Poschadel:2010:CVC

[Pos10] Norbert Poschadel. On a characterization of variance and covariance. Statistics & Probability Letters, 80(23–24):1739–1743, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002129.

Potgieter:2015:FDB

[Pot15] Paul Potgieter. The Fourier dimension of Brownian limsup fractals. Statistics & Probability Letters, 106(??):228–238, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002710.

Pfeiffer:2011:VCF

[PP11] R. M. Pfeiffer and E. Petracci. Variance computations for functionals of absolute risk estimates. Statistics & Probability Letters, 81(7):807–812, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521100040X.

Paparoditis:2014:EBL

[PP14a] Efstathios Paparoditis and Philip Preuß. Estimation of the bispectrum for locally stationary processes. Statistics & Probability Letters, 89(??):8–16, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000650.

Perez:2014:CSS

[PP14b] María-Eglée Pérez and Luis Raúl Pericchi. Changing statistical significance with the amount of information: The adaptive α significance level. Statistics & Probability Letters, 85(??):20– 24, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003611.

Park:2015:CRK

[PP15] Sangun Park and Reza Pakyari. Cumulative residual Kullback– Leibler information with the progressively type-II censored data. Statistics & Probability Letters, 106(??):287–294, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002709.

Panga:2018:MMC

[PP18a] Zacarias Panga and Luísa Pereira. On the maxima and minima of complete and incomplete samples from nonstationary random fields. Statistics & Probability Letters, 137(??):124–134, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300300.

Pilipenko:2018:PON

[PP18b] Andrey Pilipenko and Frank Norbert Proske. On perturbations of an ODE with non-Lipschitz coefficients by a small self-similar noise. Statistics & Probability Letters, 132(??):62–73, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302912.

Panga:2019:ASC

[PP19] Zacarias Panga and Luísa Pereira. On the almost sure convergence for the joint version of maxima and minima of stationary sequences. Statistics & Probability Letters, 155(??): Article 108540, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301774.

Perez:2016:NNE

[PPA16] Ana Pérez and Mercedes Prieto-Alaiz. A note on nonparametric estimation of copula-based multivariate extensions of Spearman’s rho. Statistics & Probability Letters, 112(??):41–50, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000110.

Parui:2018:OCR

[PPM18] Shyamsundar Parui, Rajender Parsad, and B. N. Mandal. A-optimal completely randomized designs for incomplete factorial structures with three factors. Statistics & Probability Letters, 137(??):343–348, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300816.

Popovic:2010:MTS

[PPN10] Bozidar V. Popović, Tibor K. Pogány, and Saralees Nadarajah. On mixed AR(1) time series model with approximated beta marginal. Statistics & Probability Letters, 80(19–20):1551–1558, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210001707.

Peng:2014:ELT

[PQW14] Liang Peng, Yongcheng Qi, and Ruodu Wang. Empirical likelihood test for high dimensional linear models. Statistics & Probability Letters, 86(??):85–90, March 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213004185.

Palacios:2012:PSH

[PR12] José Luis Palacios and José M. Renom. On partial sums of hitting times. Statistics & Probability Letters, 82(4):783–785, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003944.

Pai:2015:FAL

[PR15] Jeffrey Pai and Nalini Ravishanker. Fast approximate likelihood evaluation for stable VARFIMA processes. Statistics & Probability Letters, 103(??):160–168, August 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001108.

Pelekis:2016:LBP

[PR16] Christos Pelekis and Jan Ramon. A lower bound on the probability that a binomial random variable is exceeding its mean. Statistics & Probability Letters, 119(??):305–309, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301614.

Petrovich:2017:APP

[PR17] Justin Petrovich and Matthew Reimherr. Asymptotic properties of principal component projections with repeated eigenvalues. Statistics & Probability Letters, 130(??):42–48, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302377.

Pogorui:2013:RMG

[PRD13] Anatoliy A. Pogorui and Ramón M. Rodríguez-Dagnino. Random motion with gamma steps in higher dimensions. Statistics & Probability Letters, 83(7):1638–1643, July 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000941.

Pronzato:2013:DSO

[Pro13] Luc Pronzato. A delimitation of the support of optimal designs for Kiefer’s Φ_{p}-class of criteria. Statistics & Probability Letters, 83(12):2721–2728, December 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003076.

Park:2012:CRK

[PRS12] Sangun Park, Murali Rao, and Dong Wan Shin. On cumulative residual Kullback–Leibler information. Statistics & Probability Letters, 82(11):2025–2032, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200226X.

PintoDaCosta:2015:WRC

[PRS15] Joaquim Pinto Da Costa, Luís A. C. Roque, and Carlos Soares. The weighted rank correlation coefficient r_{W 2} in the case of ties. Statistics & Probability Letters, 99(??):20–26, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000024.

Pausinger:2018:OJS

[PRS18] Florian Pausinger, Manas Rachh, and Stefan Steinerberger. Optimal Jittered Sampling for two points in the unit square. Statistics & Probability Letters, 132(??):55–61, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302961.

Prus:2019:VOC

[Pru19] Maryna Prus. Various optimality criteria for the prediction of individual response curves. Statistics & Probability Letters, 146(??):36–41, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303468.

Polpo:2011:CBN

[PS11a] A. Polpo and D. Sinha. Correction in Bayesian nonparametric estimation in a series system or a competing-risk model. Statistics & Probability Letters, 81(12):1756–1759, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002549.

Proschan:2011:ABD

[PS11b] Michael A. Proschan and Pamela A. Shaw. Asymptotics of Bonferroni for dependent normal test statistics. Statistics & Probability Letters, 81(7):739–748, July 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003329.

Pilipauskaite:2015:JAR

[PS15] Vytaute Pilipauskaite and Donatas Surgailis. Joint aggregation of random-coefficient AR(1) processes with common innovations. Statistics & Probability Letters, 101(??):73–82, June 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000826.

Pakhteev:2016:SGR

[PS16a] A. Pakhteev and A. Stepanov. Simulation of gamma records. Statistics & Probability Letters, 119(??):204–212, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521630147X.

Papastathopoulos:2016:CIA

[PS16b] Ioannis Papastathopoulos and Kirstin Strokorb. Conditional independence among max-stable laws. Statistics & Probability Letters, 108:9–15, January 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002874.

Popivoda:2016:EGF

[PS16c] Goran Popivoda and Sinisa Stamatović. Extremes of Gaussian fields with a smooth random variance. Statistics & Probability Letters, 110(??):185–190, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215004071.

Privault:2017:CSA

[PS17] Nicolas Privault and Qihao She. Conditional Stein approximation for Itô and Skorohod integrals. Statistics & Probability Letters, 128(??):1–7, September 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301323.

Pass:2018:KTI

[PS18a] Brendan Pass and Susanna Spektor. On Khintchine type inequalities for k-wise independent Rademacher random variables. Statistics & Probability Letters, 132(??):35–39, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302882.

Peng:2018:DEC

[PS18b] Qidi Peng and Henry Schellhorn. On the distribution of extended CIR model. Statistics & Probability Letters, 142(??):23– 29, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302347.

Printechapat:2018:FNA

[PS18c] Tanes Printechapat and Songkiat Sumetkijakan. Factorizable non-atomic copulas. Statistics & Probability Letters, 143(??):86– 94, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302803.

Pakhteev:2019:DRL

[PS19a] A. Pakhteev and A. Stepanov. Discrete records: Limit theorems for their spacings and generation methods. Statistics & Probability Letters, 148(??):134–142, May 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300227.

Popivoda:2019:PHE

[PS19b] Goran Popivoda and Sinisa Stamatović. On probability of high extremes of Gaussian fields with a smooth random trend. Statistics & Probability Letters, 147(??):29–35, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303808.

Pacheco:2012:SNG

[PSC12] A. Pacheco, S. K. Samanta, and M. L. Chaudhry. A short note on the GI/Geo/1 queueing system. Statistics & Probability Letters, 82(2):268–273, February 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003191.

Pellerey:2012:CCA

[PSS12] Franco Pellerey, Moshe Shaked, and Salimeh Yasaei Sekeh. Comparisons of concordance in additive models. Statistics & Probability Letters, 82(11):2059–2067, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002659.

Park:2013:TIH

[PSS13] Junyong Park, Bimal K. Sinha, and Arvind K. Shah. Testing interval hypotheses for scale parameters in gamma distributions. Statistics & Probability Letters, 83(10):2172–2178, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001958.

Pagui:2015:FEM

[PSS15] E. C. Kenne Pagui, A. Salvan, and N. Sartori. On full efficiency of the maximum composite likelihood estimator. Statistics & Probability Letters, 97(??):120–124, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003721.

Patra:2016:NNR

[PSS16] Rohit K. Patra, Bodhisattva Sen, and Gábor J. Székely. On a nonparametric notion of residual and its applications. Statistics & Probability Letters, 109:208–213, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301930.

Pipiras:2010:RIR

[PT10] Vladas Pipiras and Murad S. Taqqu. Regularization and integral representations of Hermite processes. Statistics & Probability Letters, 80(23–24):2014–2023, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002580.

Papastathopoulos:2013:GSD

[PT13] Ioannis Papastathopoulos and Jonathan A. Tawn. A generalised Student’s t-distribution. Statistics & Probability Letters, 83(1):70–77, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003392.

Phuong:2019:DDG

[PT19] Cao Xuan Phuong and Le Thi Hong Thuy. Density deconvolution from grouped data with additive errors. Statistics & Probability Letters, 148(??):74–81, May 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300136.

Phong:2019:EDN

[PTT19] Duong Thanh Phong, Pham-Gia Thu, and Dinh Ngoc Thanh. Exact distribution of the non-central Wilks’s statistic of the second kind. Statistics & Probability Letters, 153(??):80–89, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301476.

Pestman:2016:PUP

[PTV16] Wiebe Pestman, Francis Tuerlinckx, and Wolf Vanpaemel. Persistently unbounded probability densities. Statistics & Probability Letters, 118(??):135–138, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521630102X.

Pakdaman:2010:DAB

[PTW10] Khashayar Pakdaman, Michèle Thieullen, and Gilles Wainrib. Diffusion approximation of birth-death processes: Comparison in terms of large deviations and exit points. Statistics & Probability Letters, 80(13–14):1121–1127, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000763.

Peng:2019:EPP

[PTW19] Zuoxiang Peng, Jinjun Tong, and Zhichao Weng. Exceedances point processes in the plane of stationary Gaussian sequences with data missing. Statistics & Probability Letters, 149(??):73– 79, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300288.

Puccetti:2013:SBE

[Puc13] Giovanni Puccetti. Sharp bounds on the expected shortfall for a sum of dependent random variables. Statistics & Probability Letters, 83(4):1227–1232, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000230.

Paindaveine:2017:MHD

[PV17] Davy Paindaveine and Germain Van Bever. On the maximal halfspace depth of permutation-invariant distributions on the simplex. Statistics & Probability Letters, 129(??):335–339, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302237.

Pechersky:2017:SIM

[PVY17] Eugene Pechersky, Guillem Via, and Anatoly Yambartsev. Stochastic Ising model with plastic interactions. Statistics & Probability Letters, 123(??):100–106, April 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302814.

Pourahmadi:2015:DRC

[PW15] Mohsen Pourahmadi and Xiao Wang. Distribution of random correlation matrices: Hyperspherical parameterization of the Cholesky factor. Statistics & Probability Letters, 106(??):5–12, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002011.

Peng:2017:ECM

[PY17] Liang Peng and Qiwei Yao. Estimating conditional means with heavy tails. Statistics & Probability Letters, 127(??):14– 22, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301232.

Perez:2018:AOU

[PY18] José-Luis Pérez and Kazutoshi Yamazaki. American options under periodic exercise opportunities. Statistics & Probability Letters, 135(??):92–101, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303747.

Pan:2015:SCW

[PYK15] Xiaoqing Pan, Min Yuan, and Subhash C. Kochar. Stochastic comparisons of weighted sums of arrangement increasing random variables. Statistics & Probability Letters, 102(??):42–50, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001017.

Pluta:2018:SMC

[PYS^{+}18] Dustin Pluta, Zhaoxia Yu, Tong Shen, Chuansheng Chen, Gui Xue, and Hernando Ombao. Statistical methods and challenges in connectome genetics. Statistics & Probability Letters, 136(??): 83–86, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300932.

Phillips:2014:AIM

[PZ14] T. R. L. Phillips and A. Zhigljavsky. Approximation of inverse moments of discrete distributions. Statistics & Probability Letters, 94(??):135–143, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002454.

Pang:2017:FLL

[PZ17] Guodong Pang and Yi Zheng. On the functional and local limit theorems for Markov modulated compound Poisson processes. Statistics & Probability Letters, 129(??):131–140, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301840.

Qian:2014:ARR

[QB14] Jing Qian and Rebecca A. Betensky. Assumptions regarding right censoring in the presence of left truncation. Statistics & Probability Letters, 87(??):12–17, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300415X.

Qiu:2014:CMC

[QC14] Dehua Qiu and Pingyan Chen. Complete moment convergence for i.i.d. random variables. Statistics & Probability Letters, 91(??):76–82, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214001321.

Qiu:2015:KAI

[QCZ15] Zhiping Qiu, Xiaoping Chen, and Yong Zhou. A kernelassisted imputation estimating method for the additive hazards model with missing censoring indicator. Statistics & Probability Letters, 98(??):89–97, March 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400409X.

Quang:2013:MCS

[QG13] Nguyen Van Quang and Duong Xuan Giap. Mosco convergence of SLLN for triangular arrays of rowwise independent random sets. Statistics & Probability Letters, 83(4):1117–1126, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004890.

Quang:2013:DCC

[QHS13] Nguyen Van Quang, Nguyen Ngoc Huy, and Le Hong Son. The degenerate convergence criterion and Feller’s Weak Law of Large Numbers for double arrays in noncommutative probability. Statistics & Probability Letters, 83(7):1812–1818, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001247.

Qiu:2014:NCV

[Qiu14] Yanqi Qiu. A non-commutative version of Lépingle–Yor martingale inequality. Statistics & Probability Letters, 91(??):52– 54, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001382.

Qiu:2017:EOS

[Qiu17] Guoxin Qiu. The extropy of order statistics and record values. Statistics & Probability Letters, 120(??):52–60, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301808.

Qiu:2018:REO

[QJ18] Guoxin Qiu and Kai Jia. The residual extropy of order statistics. Statistics & Probability Letters, 133(??):15–22, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303085.

Qiu:2011:DFM

[QL11] Peihua Qiu and Zhonghua Li. Distribution-free monitoring of univariate processes. Statistics & Probability Letters, 81 (12):1833–1840, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002355.

Qiu:2016:SSD

[QLH^{+}16] Qinwei Qiu, Wei Liu, Liangjian Hu, Xuerong Mao, and Surong You. Stabilization of stochastic differential equations with Markovian switching by feedback control based on discrete-time state observation with a time delay. Statistics & Probability Letters, 115(??):16–26, August 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300596.

Quang:2015:SSL

[QN15] Nguyen Van Quang and Pham Tri Nguyen. Some strong laws of large number for double array of random upper semicontinuous functions in convex combination spaces. Statistics & Probability Letters, 96(??):85–94, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003186.

Qin:2015:SPG

[QO15] Huaizhen Qin and Weiwei Ouyang. Statistical properties of gene–gene correlations in omics experiments. Statistics & Probability Letters, 97(??):206–211, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004027.

Qin:2016:ARS

[QO16] Huaizhen Qin and Weiwei Ouyang. Asymmetric risk of the Stein variance estimator under a misspecified linear regression model. Statistics & Probability Letters, 116(??):94–100, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302704.

Quang:2017:SLL

[QSS17] Nguyen Van Quang, Do The Son, and Le Hong Son. The strong laws of large numbers for positive measurable operators and applications. Statistics & Probability Letters, 124(??):110–120, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300378.

Quarteroni:2018:RSE

[Qua18] Alfio Quarteroni. The role of statistics in the era of big data: a computational scientist’ perspective. Statistics & Probability Letters, 136(??):63–67, ???? 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300920.

Qiao:2016:CPI

[QW16] Huijie Qiao and Jiang-Lun Wu. Characterizing the pathindependence of the Girsanov transformation for non-Lipschitz SDEs with jumps. Statistics & Probability Letters, 119(??):326– 333, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301663.

Qin:2013:VAA

[QX13] Yan Qin and Ning-Mao Xia. Variational approach for the adapted solution of the general backward stochastic differential equations under the Bihari condition. Statistics & Probability Letters, 83(4):1271–1281, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000266.

Qiao:2015:WCE

[QY15] Gaoxiu Qiao and Qiang Yao. Weak convergence of equity derivatives pricing with default risk. Statistics & Probability Letters, 103(??):46–56, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521500125X.

Roy:2012:WWE

[RA12] Shongkour Roy and Mian Arif Shams Adnan. Wrapped weighted exponential distributions. Statistics & Probability Letters, 82(1):77–83, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002872.

Razmkhah:2013:PCC

[RA13a] M. Razmkhah and Jafar Ahmadi. Pitman closeness of current k-records to population quantiles. Statistics & Probability Letters, 83(1):148–156, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200315X. See corrigendum [RA13b].

Razmkhah:2013:CPC

[RA13b] Mostafa Razmkhah and Jafar Ahmadi. Corrigendum to “Pitman closeness of current k-records to population quantiles [Statist. Probab. Lett. 83 (1) (2013) 148–156]”. Statistics & Probability Letters, 83(10):2446, October 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002496. See [RA13a].

Radchenko:2019:APH

[Rad19] Vadym Radchenko. Averaging principle for the heat equation driven by a general stochastic measure. Statistics & Probability Letters, 146(??):224–230, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303791.

Rahimov:2011:EOM

[Rah11] I. Rahimov. Estimation of the offspring mean in a supercritical branching process with non-stationary immigration. Statistics & Probability Letters, 81(8):907–914, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001295.

Rahimov:2017:AIN

[Rah17] I. Rahimov. Asymptotic inference for non-supercritical partially observed branching processes. Statistics & Probability Letters, 126(??):26–32, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300809.

Rao:2010:CIH

[Rao10] B. L. S. Prakasa Rao. Chebyshev’s inequality for Hilbert-space-valued random elements. Statistics & Probability Letters, 80 (11–12):1039–1042, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000696.

Rao:2012:RMI

[Rao12] B. L. S. Prakasa Rao. Remarks on maximal inequalities for non-negative demisubmartingales. Statistics & Probability Letters, 82(7):1388–1390, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001125.

Rao:2013:FHT

[Rao13] Shravas K. Rao. Finding hitting times in various graphs. Statistics & Probability Letters, 83(9):2067–2072, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001818.

Rao:2018:MDP

[Rao18] B. L. S. Prakasa Rao. Moderate deviation principle for maximum likelihood estimator for Markov processes. Statistics & Probability Letters, 132(??):74–82, January 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730295X.

Raqab:2010:EMR

[Raq10] Mohammad Z. Raqab. Evaluations of the mean residual lifetime of an m-out-of-n system. Statistics & Probability Letters, 80(5–6):333–342, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004313.

Ratanov:2013:DJT

[Rat13] Nikita Ratanov. Damped jump-telegraph processes. Statistics & Probability Letters, 83(10):2282–2290, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300223X.

Ratanov:2014:PLP

[Rat14] Nikita Ratanov. On piecewise linear processes. Statistics & Probability Letters, 90(??):60–67, July 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001084.

Ratanov:2015:HED

[Rat15] Nikita Ratanov. Hypo-exponential distributions and compound Poisson processes with alternating parameters. Statistics & Probability Letters, 107:71–78, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002850.

Ratanov:2017:PLP

[Rat17] Nikita Ratanov. Piecewise linear process with renewal starting points. Statistics & Probability Letters, 131(??):78–86, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302705.

Rau:2013:BCT

[Rau13] Christian Rau. Bayes classifiers of three-dimensional rotations and the sphere with symmetries. Statistics & Probability Letters, 83(3):930–935, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004713.

Rodrigues:2016:BRC

[RBSB16] Josemar Rodrigues, Jorge L. Bazán, Adriano K. Suzuki, and Narayanaswamy Balakrishnan. The Bayesian restricted Conway–Maxwell-binomial model to control dispersion in count data. Statistics & Probability Letters, 119(??):281–288, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301651.

Riabi:2010:EPT

[RBY10] Mehdi Yaghoobi Avval Riabi, G. R. Mohtashami Borzadaran, and G. H. Yari. β-entropy for Pareto-type distributions and related weighted distributions. Statistics & Probability Letters, 80(19–20):1512–1519, October 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001641.

Reddy:2017:PPR

[Red17] Tulasi Ram Reddy. Probability that product of real random matrices have all eigenvalues real tend to 1. Statistics & Probability Letters, 124(??):30–32, May 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730007X.

Reffel:2014:TIP

[Ref14] Fabian P. Reffel. Termination of the iterative proportional fitting procedure. Statistics & Probability Letters, 92(??):59–64, September 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001783.

Reimherr:2015:FRR

[Rei15] Matthew Reimherr. Functional regression with repeated eigenvalues. Statistics & Probability Letters, 107:62–70, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002783.

Reid:2018:SSW

[Rei18] Nancy Reid. Statistical science in the world of big data. Statistics & Probability Letters, 136(??):42–45, ???? 2018.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300944.

Ren:2013:RTS

[Ren13] Liying Ren. On representation theorem of sublinear expectation related to G-Lévy process and paths of G-Lévy process. Statistics & Probability Letters, 83(5):1301–1310, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000394.

Ren:2014:EBU

[Ren14] Xingwei Ren. On the equivalence of the BLUEs under a general linear model and its restricted and stochastically restricted models. Statistics & Probability Letters, 90(??):1–10, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000996. See corrigendum [Ren15].

Ren:2015:CEB

[Ren15] Xingwei Ren. Corrigendum to “On the equivalence of the BLUEs under a general linear model and its restricted and stochastically restricted models” [Statist. Probab. Lett. 90 (2014) 1– 10]. Statistics & Probability Letters, 104(??):181–185, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001558. See [Ren14].

Ressel:2011:RKR

[Res11] Paul Ressel. A revision of Kimberling’s results — with an application to max-infinite divisibility of some Archimedean copulas. Statistics & Probability Letters, 81(2):207–211, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003263.

Reveillac:2012:OCB

[Rév12] Anthony Réveillac. On the orthogonal component of BSDEs in a Markovian setting. Statistics & Probability Letters, 82(1):151– 157, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003099.

Rezapour:2015:CNA

[Rez15] Mohsen Rezapour. On the construction of nested Archimedean copulas for d-monotone generators. Statistics & Probability Letters, 101(??):21–32, June 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000814.

Rugamer:2018:SIA

[RG18] David Rügamer and Sonja Greven. Selective inference after likelihood- or test-based model selection in linear models. Statistics & Probability Letters, 140(??):7–12, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301640.

Ren:2011:NSD

[RH11] Yong Ren and Lanying Hu. A note on the stochastic differential equations driven by G-Brownian motion. Statistics & Probability Letters, 81(5):580–585, May 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000174.

Rigdon:2015:ECI

[RH15] Joseph Rigdon and Michael G. Hudgens. Exact confidence intervals in the presence of interference. Statistics & Probability Letters, 105(??):130–135, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001972.

Rizzo:2016:EDG

[RH16] Maria L. Rizzo and John T. Haman. Expected distances and goodness-of-fit for the asymmetric Laplace distribution. Statistics & Probability Letters, 117(??):158–164, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300591.

Roman:2014:RGS

[RHP14] Jorge Carlos Román, James P. Hobert, and Brett Presnell. On reparametrization and the Gibbs sampler. Statistics & Probability Letters, 91(??):110–116, August 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001175.

Ricci:2010:AST

[Ric10] Lila Ricci. Adjusted R-squared type measure for exponential dispersion models. Statistics & Probability Letters, 80(17–18):1365–1368, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001288.

Rahmani:2016:SCT

[RIK16] Rabi-Allah Rahmani, Muhyiddin Izadi, and Baha-Eldin Khaledi. Stochastic comparisons of total capacity of weighted-k-out-of-n systems. Statistics & Probability Letters, 117(??):216–220, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300864.

Rubio:2014:IJP

[RL14] Francisco Javier Rubio and Brunero Liseo. On the independence Jeffreys prior for skew-symmetric models. Statistics & Probability Letters, 85(??):91–97, February 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003933.

Rubio:2011:SCG

[RM11] Francisco Rubio and Xavier Mestre. Spectral convergence for a general class of random matrices. Statistics & Probability Letters, 81(5):592–602, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000113.

Richter:2013:SMC

[RM13] Scott J. Richter and Melinda H. McCann. Simultaneous multiple comparisons with a control using median differences and permutation tests. Statistics & Probability Letters, 83(4):1167– 1173, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000151.

Ruiz-Medina:2019:NSC

[RMÁL19] M. D. Ruiz-Medina and J. Álvarez-Liébana. A note on strongconsistency of componentwise ARH(1) predictors. Statistics & Probability Letters, 145(??):224–228, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303031.

Robertson:2017:MBA

[RMPB17] B. L. Robertson, T. McDonald, C. J. Price, and J. A. Brown. A modification of balanced acceptance sampling. Statistics & Probability Letters, 129(??):107–112, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301797.

Remillard:2017:CBC

[RNB17] Bruno Rémillard, Bouchra Nasri, and Taoufik Bouezmarni. On copula-based conditional quantile estimators. Statistics & Probability Letters, 128(??):14–20, September 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301530.

Robert:2010:ADM

[Rob10] C. Y. Robert. On asymptotic distribution of maxima of stationary sequences subject to random failure or censoring. Statistics & Probability Letters, 80(2):134–142, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003691.

Robert:2013:SNC

[Rob13] Christian Y. Robert. Some new classes of stationary maxstable random fields. Statistics & Probability Letters, 83(6):1496– 1503, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000606.

Roberts:2015:DFT

[Rob15] Leigh A. Roberts. Distribution free testing of goodness of fit in a one dimensional parameter space. Statistics & Probability Letters, 99(??):215–222, April 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000085.

Roberts:2019:DFG

[Rob19] Leigh A. Roberts. Distribution free goodness of fit testing of grouped Bernoulli trials. Statistics & Probability Letters, 150(??):47–53, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300410.

Rodriguez:2013:JPM

[Rod13] Abel Rodríguez. On the Jeffreys prior for the multivariate Ewens distribution. Statistics & Probability Letters, 83(6):1539– 1546, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000576.

Rohan:2013:TVM

[Roh13] Neelabh Rohan. A time varying GARCH(p, q) model and related statistical inference. Statistics & Probability Letters, 83 (9):1983–1990, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521300151X.

Rohmer:2016:SRC

[Roh16] Tom Rohmer. Some results on change-point detection in cross-sectional dependence of multivariate data with changes in marginal distributions. Statistics & Probability Letters, 119(??):45–54, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301079.

Rokhlin:2015:CLT

[Rok15] Dmitry B. Rokhlin. Central limit theorem under uncertain linear transformations. Statistics & Probability Letters, 107:191–198, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500317X.

Roland:2010:NPE

[Rol10] Christophe Roland. A note on the parameterized EM method. Statistics & Probability Letters, 80(17–18):1354–1357, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210001264.

Rollin:2018:QBM

[Röl18] Adrian Röllin. On quantitative bounds in the mean martingale central limit theorem. Statistics & Probability Letters, 138(??): 171–176, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301184.

Rosati:2013:ERC

[Ros13] Nicoletta Rosati. Efficiency of repeated-cross-section estimators in fixed-effects models. Statistics & Probability Letters, 83(7):1770–1775, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001260.

Ramos:2017:NEI

[ROSL17] Héctor M. Ramos, Jorge Ollero, and Alfonso Suárez-Llorens. A new explanatory index for evaluating the binary logistic regression based on the sensitivity of the estimated model. Statistics & Probability Letters, 120(??):135–140, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302048.

Roy:2012:SAC

[Roy12] Vivekananda Roy. Spectral analytic comparisons for data augmentation. Statistics & Probability Letters, 82(1):103–108, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002963.

Rozovsky:2010:BLL

[Roz10] Leonid Rozovsky. On the behavior of the log Laplace transform of series of weighted non-negative random variables at infinity. Statistics & Probability Letters, 80(9–10):764–770, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000131.

Rozovsky:2012:SDP

[Roz12] Leonid Rozovsky. Superlarge deviation probabilities for sums of independent lattice random variables with exponential decreasing tails. Statistics & Probability Letters, 82(1):72–76, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002938.

Rozovsky:2014:SDP

[Roz14] L. V. Rozovsky. Small deviation probabilities of weighted sums under minimal moment assumptions. Statistics & Probability Letters, 86(??):1–6, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003921.

Rozovsky:2016:SDP

[Roz16] L. V. Rozovsky. Small deviation probabilities for weighted sum of independent random variables with a common distribution that can decrease at zero fast enough. Statistics & Probability Letters, 117(??):192–200, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300657.

Ramosaj:2019:CER

[RP19] Burim Ramosaj and Markus Pauly. Consistent estimation of residual variance with random forest out-of-bag errors. Statistics & Probability Letters, 151(??):49–57, August 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300938.

Romer:2010:MLE

[RR10a] Megan M. Romer and Donald St. P. Richards. Maximum likelihood estimation of the mean of a multivariate normal population with monotone incomplete data. Statistics & Probability Letters, 80(17–18):1284–1288, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001148.

Ruckdeschel:2010:FIS

[RR10b] Peter Ruckdeschel and Helmut Rieder. Fisher information of scale. Statistics & Probability Letters, 80(23–24):1881–1885, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002385.

Robini:2011:SAT

[RR11] Marc C. Robini and Pierre-Jean Reissman. On simulated annealing with temperature-dependent energy and temperaturedependent communication. Statistics & Probability Letters, 81(8):915–920, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001349.

Roberts:2013:NFC

[RR13] Gareth O. Roberts and Jeffrey S. Rosenthal. A note on formal constructions of sequential conditional couplings. Statistics & Probability Letters, 83(9):2073–2076, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001739.

Reza:2014:SRC

[RR14] Sadat Reza and Paul Rilstone. A simple root-N-consistent semiparametric estimator for discrete duration models. Statistics & Probability Letters, 95(??):150–154, December 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002983.

Ruiz:2016:LTM

[RR16] Patricia Alonso Ruiz and Alexander Rakitko. The limit theorem for maximum of partial sums of exchangeable random variables. Statistics & Probability Letters, 119(??):357–362, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301730.

Rao:2010:SCS

[RS10a] B. L. S. Prakasa Rao and Harshinder Singh. Sufficient conditions for stochastic equality of two distributions under some partial orders. Statistics & Probability Letters, 80(5–6):513–518, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209004544.

Rosalsky:2010:SLL

[RS10b] Andrew Rosalsky and George Stoica. On the Strong Law of Large Numbers for identically distributed random variables irrespective of their joint distributions. Statistics & Probability Letters, 80(17–18):1265–1270, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001124.

Ren:2012:NDI

[RS12a] Yaofeng Ren and Jing Shen. A note on the domination inequalities and their applications. Statistics & Probability Letters, 82(6):1160–1168, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000740.

Rifo:2012:NCI

[RS12b] L. R. Rifo and J. D. Santos. A note on conflict of information and subexponential densities. Statistics & Probability Letters, 82(4): 840–842, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000235.

Rupasinghe:2012:APS

[RS12c] Maduka Rupasinghe and V. A. Samaranayake. Asymptotic properties of sieve bootstrap prediction intervals for FARIMA processes. Statistics & Probability Letters, 82(12):2108–2114, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212002830.

Ranjan:2014:SFL

[RS14a] Pritam Ranjan and Neil Spencer. Space-filling Latin hypercube designs based on randomization restrictions in factorial experiments. Statistics & Probability Letters, 94(??):239–247, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521400279X.

Rydlewski:2014:GES

[RS14b] Jerzy P. Rydlewski and Małgorzata Snarska. On geometric ergodicity of skewed-SVCHARME models. Statistics & Probability Letters, 84(??):192–197, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003519.

Roozegar:2015:ABR

[RS15a] Rasool Roozegar and A. R. Soltani. On the asymptotic behavior of randomly weighted averages. Statistics & Probability Letters, 96(??):269–272, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003526.

Rudolf:2015:EBM

[RS15b] Daniel Rudolf and Nikolaus Schweizer. Error bounds of MCMC for functions with unbounded stationary variance. Statistics & Probability Letters, 99(??):6–12, April 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004246.

Rao:2016:RCL

[RS16] B. L. S. Prakasa Rao and M. Sreehari. Random central limit theorem for associated random variables and the order of approximation. Statistics & Probability Letters, 111(??):1–7, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521600002X.

Rocha:2010:SOA

[RSC10] Andréa V. Rocha, Alexandre B. Simas, and Gauss M. Cordeiro. Second-order asymptotic expressions for the covariance matrix of maximum likelihood estimators in dispersion models. Statistics & Probability Letters, 80(7–8):718–725, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000064.

Reijsbergen:2015:SHT

[RSdB15] Daniël Reijsbergen, Werner Scheinhardt, and Pieter-Tjerk de Boer. A sequential hypothesis test based on a generalized Azuma inequality. Statistics & Probability Letters, 97(??):192– 196, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003940.

Revathi:2016:SFD

[RSR16] P. Revathi, R. Sakthivel, and Yong Ren. Stochastic functional differential equations of Sobolev-type with infinite delay. Statistics & Probability Letters, 109:68–77, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003727.

Rosenberg:2010:OAE

[RSV10] Dinah Rosenberg, Eilon Solan, and Nicolas Vieille. On the optimal amount of experimentation in sequential decision problems. Statistics & Probability Letters, 80(5–6):381–385, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004386.

Renault:2015:CS

[RT15] Eric Renault and Umberto Triacca. Causality and separability. Statistics & Probability Letters, 99(??):1–5, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214004210.

Rubio:2015:PPH

[Rub15] F. J. Rubio. On the propriety of the posterior of hierarchical linear mixed models with flexible random effects distributions. Statistics & Probability Letters, 96(??):154–161, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003435.

Rukhin:2012:ECM

[Ruk12] Andrew L. Rukhin. Estimating common mean and heterogeneity variance in two study case meta-analysis. Statistics & Probability Letters, 82(7):1318–1325, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001241.

Rukhin:2016:CRC

[Ruk16] Andrew L. Rukhin. Confidence regions for comparison of two normal samples. Statistics & Probability Letters, 119(??):273– 280, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301560.

Rulloni:2014:UCA

[Rul14] Valeria Rulloni. Uniqueness condition for an auto-logistic model. Statistics & Probability Letters, 87(??):1–6, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213004148.

Ruzankin:2014:CKM

[Ruz14] P. S. Ruzankin. On Cox–Kemperman moment inequalities for independent centered random variables. Statistics & Probability Letters, 86(??):80–84, March 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213004045.

Remillard:2019:DPT

[RV19a] Bruno N. Rémillard and Jean Vaillancourt. Detecting periodicity from the trajectory of a random walk in random environment. Statistics & Probability Letters, 155(??):Article 108568, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302147.

Rodriguez:2019:BPM

[RV19b] Daniela Rodriguez and Marina Valdora. The breakdown point of the median of means tournament. Statistics & Probability Letters, 153(??):108–112, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301415.

Rempala:2016:DAC

[RW16a] Grzegorz A. Rempala and Jacek Wesolowski. Double asymptotics for the chi-square statistic. Statistics & Probability Letters, 119(??):317–325, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301699.

Romano:2016:ECA

[RW16b] Joseph P. Romano and Michael Wolf. Efficient computation of adjusted p-values for resampling-based stepdown multiple testing. Statistics & Probability Letters, 113(??):38–40, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000389.

Rekabdarkolaee:2017:VST

[RW17] Hossein Moradi Rekabdarkolaee and Qin Wang. Variable selection through adaptive MAVE. Statistics & Probability Letters, 128(??):44–51, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301517.

Ren:2010:SSV

[RZ10] Yunwen Ren and Xinsheng Zhang. Subset selection for vector autoregressive processes via adaptive lasso. Statistics & Probability Letters, 80(23–24):1705–1712, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002075.

Radchenko:2012:HEG

[RZ12] Vadym Radchenko and Martina Zähle. Heat equation with a general stochastic measure on nested fractals. Statistics & Probability Letters, 82(3):699–704, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211004056.

Samii:2012:EBD

[SA12a] Cyrus Samii and Peter M. Aronow. On equivalencies between design-based and regression-based variance estimators for randomized experiments. Statistics & Probability Letters, 82(2):365–370, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003452.

Sankaran:2012:AHM

[SA12b] P. G. Sankaran and P. Anisha. Additive hazards models for gap time data with multiple causes. Statistics & Probability Letters, 82(7):1454–1462, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000727.

Swanepoel:2013:SNR

[SA13] J. W. H. Swanepoel and J. S. Allison. Some new results on the empirical copula estimator with applications. Statistics & Probability Letters, 83(7):1731–1739, July 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001107.

Sabelfeld:2017:MFF

[Sab17] Karl K. Sabelfeld. A mesh free floating random walk method for solving diffusion imaging problems. Statistics & Probability Letters, 121(??):6–11, February 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302139.

Sabelfeld:2018:AMF

[Sab18] Karl K. Sabelfeld. Application of the von Mises–Fisher distribution to Random Walk on Spheres method for solving high-dimensional diffusion–advection–reaction equations. Statistics & Probability Letters, 138(??):137–142, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301160.

Sakhanenko:2012:NIE

[Sak12] Lyudmila Sakhanenko. Numerical issues in estimation of integral curves from noisy diffusion tensor data. Statistics & Probability Letters, 82(6):1136–1144, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000867.

Sakhanenko:2015:RAE

[Sak15] Lyudmila Sakhanenko. Rate acceleration for estimators of integral curves from diffusion tensor imaging (DTI) data. Statistics & Probability Letters, 107:286–295, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301000.

Sakhanenko:2017:SOK

[Sak17] Lyudmila Sakhanenko. In search of an optimal kernel for a bias correction method for density estimators. Statistics & Probability Letters, 122(??):42–50, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302279.

Sadooghi-Alvandi:2013:NTH

[SAM13] Soltan Mohammad Sadooghi-Alvandi and Ahad Malekzadeh. A note on testing homogeneity of the scale parameters of several inverse Gaussian distributions. Statistics & Probability Letters, 83(8):1844–1848, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001405.

Sangalli:2018:RSE

[San18] Laura M. Sangalli. The role of statistics in the era of Big Data. Statistics & Probability Letters, 136(??):1–3, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830155X.

Sason:2013:ILB

[Sas13a] Igal Sason. Improved lower bounds on the total variation distance for the Poisson approximation. Statistics & Probability Letters, 83(10):2422–2431, October 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002514.

Sason:2013:TEB

[Sas13b] Igal Sason. Tightened exponential bounds for discrete-time conditionally symmetric martingales with bounded jumps. Statistics & Probability Letters, 83(8):1928–1936, August 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001363.

Saussereau:2012:DPB

[Sau12] Bruno Saussereau. Deviation probability bounds for fractional martingales and related remarks. Statistics & Probability Letters, 82(8):1610–1618, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001708.

Subramanian:2010:DRS

[SB10] Sundarraman Subramanian and Dipankar Bandyopadhyay. Doubly robust semiparametric estimation for the missing censoring indicator model. Statistics & Probability Letters, 80 (7–8):621–630, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715209004775.

Sharp:2016:CFD

[SB16] Julia L. Sharp and John J. Borkowski. A conditional frequency distribution test for analyzing 2 × c tables. Statistics & Probability Letters, 119(??):334–343, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301559.

Seo:2019:CLP

[SB19] Won-Ki Seo and Brendan K. Beare. Cointegrated linear processes in Bayes Hilbert space. Statistics & Probability Letters, 147(??):90–95, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303882.

Salehi:2012:PPH

[SBA12] E. T. Salehi, F. G. Badía, and M. Asadi. Preservation properties of a homogeneous Poisson process stopped at an independent random time. Statistics & Probability Letters, 82(3):574– 585, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003592.

Stepanov:2016:CRL

[SBN16] Alexei Stepanov, Alexandre Berred, and Valery B. Nevzorov. Concomitants of records: Limit results, generation techniques, correlation. Statistics & Probability Letters, 109:184–188, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003879.

Shangari:2012:PME

[SC12a] Dhruv Shangari and Jiahua Chen. Partial monotonicity of entropy measures. Statistics & Probability Letters, 82 (11):1935–1940, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002635.

Shen:2012:SIR

[SC12b] Guangjun Shen and Chao Chen. Stochastic integration with respect to the sub-fractional Brownian motion with H ∈ (0, 1/2). Statistics & Probability Letters, 82(2):240–251, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003233.

Shiu:2015:RCG

[SC15] Shang-Ying Shiu and Ting-Li Chen. On the rate of convergence of the Gibbs sampler for the 1-D Ising model by geometric bound. Statistics & Probability Letters, 105(??):14–19, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521500190X.

Schott:2012:NML

[Sch12] James R. Schott. A note on maximum likelihood estimation for covariance reducing models. Statistics & Probability Letters, 82(9):1629–1631, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200171X.

Schlemm:2013:ENS

[Sch13] Eckhard Schlemm. On the expected number of successes in a sequence of nested Bernoulli trials. Statistics & Probability Letters, 83(7):1619–1623, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001016.

Schlemm:2014:LDM

[Sch14] Eckhard Schlemm. Limiting distribution of the maximal distance between random points on a circle: A moments approach. Statistics & Probability Letters, 92(??):132–136, September 1– 262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001953.

Schott:2016:RPR

[Sch16] James R. Schott. On a robustness property of the Rayleigh and Bingham tests of uniformity. Statistics & Probability Letters, 111(??):55–59, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215303850.

Scott:2018:RSE

[Sco18] E. Marian Scott. The role of statistics in the era of big data: Crucial, critical and under-valued. Statistics & Probability Letters, 136(??):20–24, ???? 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300956.

Shi:2014:REV

[SCS14] Jianhong Shi, Kun Chen, and Weixing Song. Robust errorsin-variables linear regression via Laplace distribution. Statistics & Probability Letters, 84(??):113–120, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003349.

Shi:2015:SRN

[SCZ15] Jianhua Shi, Xiaoping Chen, and Yong Zhou. The strong representation for the nonparametric estimator of length-biased and right-censored data. Statistics & Probability Letters, 104(??):49– 57, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001418.

Schuhmacher:2010:CML

[SD10a] Dominic Schuhmacher and Lutz Dümbgen. Consistency of multivariate log-concave density estimators. Statistics & Probability Letters, 80(5–6):376–380, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004374.

Song:2010:DFL

[SD10b] Weixing Song and Juan Du. Distribution-free lack-of-fit tests in balanced mixed models. Statistics & Probability Letters, 80(17–18):1378–1387, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001380.

Shi:2011:BVS

[SD11] Minghui Shi and David B. Dunson. Bayesian variable selection via particle stochastic search. Statistics & Probability Letters, 81(2):283–291, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002944.

Sreehari:2014:LIR

[SD14] Maddipatla Sreehari and Gooty Divanji. Limit infimum results for subsequences of partial sums and random sums of i.i.d. random variables. Statistics & Probability Letters, 95(??):101–109, December 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002740.

Shen:2015:HTI

[SD15] Gang Shen and Karl D’Silva. A hybrid test for the isotonic change-point problem. Statistics & Probability Letters, 99(??): 36–43, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000036.

Sengupta:2016:SBD

[SD16] Debasis Sengupta and Sudipta Das. Sharp bounds on DMRL and IMRL classes of life distributions with specified mean. Statistics & Probability Letters, 119(??):101–107, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301286.

Soni:2015:TSD

[SDJ15] Pooja Soni, Isha Dewan, and Kanchan Jain. Tests for successive differences of quantiles. Statistics & Probability Letters, 97(??): 1–8, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003629.

Stange:2016:MDA

[SDNS16] Jens Stange, Thorsten Dickhaus, Arcadi Navarro, and Daniel Schunk. Multiplicity- and dependency-adjusted p-values for control of the family-wise error rate. Statistics & Probability Letters, 111(??):32–40, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000043.

Subtil:2012:CDD

[SdOG12] Ana Subtil, M. Rosário de Oliveira, and Luzia Gonçalves. Conditional dependence diagnostic in the latent class model: a simulation study. Statistics & Probability Letters, 82(7):1407– 1412, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521200123X.

Secchi:2018:RSE

[Sec18] Piercesare Secchi. On the role of statistics in the era of big data: a call for a debate. Statistics & Probability Letters, 136(??): 10–14, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300865.

Seol:2015:LTD

[Seo15] Youngsoo Seol. Limit theorems for discrete Hawkes processes. Statistics & Probability Letters, 99(??):223–229, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000292.

Seol:2017:LTC

[Seo17] Youngsoo Seol. Limit theorems for the compensator of Hawkes processes. Statistics & Probability Letters, 127(??):165–172, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301347.

Seol:2019:LTI

[Seo19] Youngsoo Seol. Limit theorems for an inverse Markovian Hawkes process. Statistics & Probability Letters, 155(??):Article 108580, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302263.

Sepehri:2018:ANT

[Sep18]  Amir Sepehri.  Asymptotic normality of the trace for

a class of distributions on orthogonal matrices. Statistics & Probability Letters, 137(??):14–18, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300051.

Serdyukova:2017:CIG

[Ser17] Nora Serdyukova. A concentration inequality for a Gaussian process indexed by matrices. Statistics & Probability Letters, 123(??):23–26, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216302772.

Serdyukova:2018:LBE

[Ser18] Nora Serdyukova. Lower bounds in estimation at a point under multi-index constraint. Statistics & Probability Letters, 138(??): 183–189, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521830107X.

Su:2012:EMG

[SF12] Jianxi Su and Edward Furman. Erratum to “On a multivariate Gamma distribution” by E. Furman [Statist. Probab. Lett. 78 (2008) 2353–2360]. Statistics & Probability Letters, 82(5):1040– 1041, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000399. See [Fur08].

Soltani:2014:SMR

[SG14] A. R. Soltani and H. Ghasemi. Semi-Markov and reward fields. Statistics & Probability Letters, 95(??):71–76, ???? 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002910.

Semrau-Gilka:2015:ASO

[SG15a] Alina Semrau-Gilka. On approximation of solutions of onedimensional reflecting SDEs with discontinuous coefficients. Statistics & Probability Letters, 96(??):315–321, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003605.

Sun:2015:IPF

[SG15b] Xiaoxia Sun and Feng Guo. On integration by parts formula and characterization of fractional Ornstein–Uhlenbeck process. Statistics & Probability Letters, 107:170–177, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003144.

Sun:2018:MEC

[SG18] Xiaoxia Sun and Feng Guo. On moment estimates and continuity for solutions of SDEs driven by fractional Brownian motions under non-Lipschitz conditions. Statistics & Probability Letters, 132(??):116–124, January 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302948.

Singh:2010:NCS

[SGG10] Parminder Singh, Anju Goyal, and A. N. Gill. A note on comparing several variances with a control variance. Statistics & Probability Letters, 80(23–24):1995–2002, December 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002555.

Sun:2015:AAD

[SGV15] Xianming Sun, Siqing Gan, and Michèle Vanmaele. Analytical approximation for distorted expectations. Statistics & Probability Letters, 107:246–252, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003247.

Shen:2011:MSP

[SH11] Dan Shen and Xiaoyu Hu. The multifractal structure of the product of two stable occupation measures. Statistics & Probability Letters, 81(4):478–488, April 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003457.

Shin:2013:SBC

[SH13] Dong Wan Shin and Eunju Hwang. Stationary bootstrapping for cointegrating regressions. Statistics & Probability Letters, 83(2):474–480, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003793.

Sarkar:2017:CPH

[SH17] A. Sarkar and M. Haenggi. Continuum percolation with holes. Statistics & Probability Letters, 126(??):212–218, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301013.

Shakhatreh:2012:TPW

[Sha12a] M. K. Shakhatreh. A two-parameter of weighted exponential distributions. Statistics & Probability Letters, 82(2):252–261, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003294.

Shalit:2012:UOT

[Sha12b] Haim Shalit. Using OLS to test for normality. Statistics & Probability Letters, 82(11):2050–2058, November 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002763.

Shang:2012:LPM

[Sha12c] Zuofeng Shang. On latent process models in multi-dimensional space. Statistics & Probability Letters, 82(7):1259–1266, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001150.

Shan:2013:MEU

[Sha13a] Guogen Shan. More efficient unconditional tests for exchangeable binary data with equal cluster sizes. Statistics & Probability Letters, 83(2):644–649, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004300.

Shashkin:2013:FCL

[Sha13b] A. Shashkin. A functional Central Limit Theorem for the level measure of a Gaussian random field. Statistics & Probability Letters, 83(2):637–643, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004233.

Shan:2014:EAT

[Sha14] Guogen Shan. Exact approaches for testing non-inferiority or superiority of two incidence rates. Statistics & Probability Letters, 85(??):129–134, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521300391X.

Shao:2017:DAR

[Sha17] Hui Shao. Decomposing aggregate risk into marginal risks under partial information: A top-down method. Statistics & Probability Letters, 124(??):97–100, May 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300391.

Sharples:2018:RSE

[Sha18] Linda D. Sharples. The role of statistics in the era of big data: Electronic health records for healthcare research. Statistics & Probability Letters, 136(??):105–110, ???? 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300890.

Spagnoli:2011:HMM

[SHBHD11] Alessandra Spagnoli, Robin Henderson, Richard J. Boys, and Jeanine J. Houwing-Duistermaat. A hidden Markov model for informative dropout in longitudinal response data with crisis states. Statistics & Probability Letters, 81(7):730–738, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000435.

Shen:2011:PIC

[She11] Gang Shen. On periodically isotonic climate change. Statistics & Probability Letters, 81(11):1627–1634, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001891.

Shen:2013:ELI

[She13] Gang Shen. On empirical likelihood inference of a changepoint. Statistics & Probability Letters, 83(7):1662–1668, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000977.

Shimura:2012:LDR

[Shi12] Takaaki Shimura. Limit distribution of a roundoff error. Statistics & Probability Letters, 82(4):713–719, April 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211004135.

Shimizu:2014:BNA

[Shi14] Kenichi Shimizu. Bootstrapping the nonparametric ARCH regression model. Statistics & Probability Letters, 87(??):61– 69, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000066.

Shi:2018:HDS

[Shi18] Jian Qing Shi. How do statisticians analyse big data — our story. Statistics & Probability Letters, 136(??):130–133, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300889.

Shi:2019:GFB

[Shi19a] Qun Shi. Generalized fractional BSDE with jumps and Lipschitz coefficients. Statistics & Probability Letters, 155(??): Article 108549, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301865.

Shi:2019:NLD

[Shi19b] Wanlin Shi. A note on large deviation probabilities for empirical distribution of branching random walks. Statistics & Probability Letters, 147(??):18–28, April 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303845.

Shin:2019:CLB

[Shi19c] Jiyong Shin. Construction of Liouville Brownian motion via Dirichlet form theory. Statistics & Probability Letters, 148(??): 128–132, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300197.

Spade:2014:NRT

[SHK14] David A. Spade, Radu Herbei, and Laura S. Kubatko. A note on the relaxation time of two Markov chains on rooted phylogenetic tree spaces. Statistics & Probability Letters, 84(??):247– 252, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213003155.

Soubeyrand:2015:WCP

[SHL15]  Samuel Soubeyrand and Emilie Haon-Lasportes.  Weak

convergence of posteriors conditional on maximum pseudolikelihood estimates and implications in ABC. Statistics & Probability Letters, 107:84–92, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002825.

Shmerling:2013:RRM

[Shm13] Efraim Shmerling. A range reduction method for generating discrete random variables. Statistics & Probability Letters, 83(4):1094–1099, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000035.

Shushi:2016:PCC

[Shu16] Tomer Shushi. A proof for the conjecture of characteristic function of the generalized skew-elliptical distributions. Statistics & Probability Letters, 119(??):301–304, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301626.

Shushi:2018:GSE

[Shu18a] Tomer Shushi. Generalized skew-elliptical distributions are closed under affine transformations. Statistics & Probability Letters, 134(??):1–4, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303267.

Shushi:2018:PEM

[Shu18b] Tomer Shushi. A proof for the existence of multivariate singular generalized skew-elliptical density functions. Statistics & Probability Letters, 141(??):50–55, October 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302050.

Shushi:2018:SLT

[Shu18c] Tomer Shushi. Stein’s lemma for truncated elliptical random vectors. Statistics & Probability Letters, 137(??):297–303, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300531.

Shushi:2019:MLS

[Shu19a] Tomer Shushi. The Minkowski length of a spherical random vector. Statistics & Probability Letters, 153(??):104–107, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301555.

Shushi:2019:NCE

[Shu19b] Tomer Shushi. A note on the coefficients of elliptical random variables. Statistics & Probability Letters, 152(??):153–155, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219301336.

Szymkowiak:2016:CDW

[SI16] Magdalena Szymkowiak and Maria Iwińska. Characterizations of Discrete Weibull related distributions. Statistics & Probability Letters, 111(??):41–48, April 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216000031.

Smirnova:2018:PGB

[SIBC18] Ekaterina Smirnova, Andrada Ivanescu, Jiawei Bai, and Ciprian M. Crainiceanu. A practical guide to big data. Statistics & Probability Letters, 136(??):25–29, ???? 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300592.

Singh:2019:TLO

[Sin19] Rakhi Singh. On three-level D-optimal paired choice designs. Statistics & Probability Letters, 145(??):127–132, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303043.

Sun:2014:ELB

[SJ14] Haoze Sun and Yuexiang Jiang. Empirical likelihood based confidence intervals for the tail index when γ < −1/2. Statistics & Probability Letters, 84(??):149–157, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003350.

Singh:2011:PEL

[SK11] Sarjinder Singh and Jong-Min Kim. A pseudo-empirical log-likelihood estimator using scrambled responses. Statistics & Probability Letters, 81(3):345–351, March 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003305.

Sen:2013:BHL

[SK13] Pranab K. Sen and Moonsu Kang. Bivariate high-level exceedance and the Chen–Stein theorem in genomics multiple hypothesis testing perspectives. Statistics & Probability Letters, 83(7):1725–1730, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001028.

Seo:2015:PIS

[SK15] Jung-In Seo and Suk-Bok Kang. Pivotal inference for the scaled half logistic distribution based on progressively Type-II censored samples. Statistics & Probability Letters, 104(??):109–116, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001625.

Spanhel:2016:PCP

[SK16] Fabian Spanhel and Malte S. Kurz. The partial copula: Properties and associated dependence measures. Statistics & Probability Letters, 119(??):76–83, December 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215303552.

Soja-Kukiela:2017:AOS

[SK17] Natalia Soja-Kukiela. Asymptotics of the order statistics for a process with a regenerative structure. Statistics & Probability Letters, 131(??):108–115, December 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302869.

Shah:2014:RTD

[SKB14] Imtiyaz A. Shah, A. H. Khan, and H. M. Barakat. Random translation, dilation and contraction of order statistics. Statistics & Probability Letters, 92(??):209–214, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002016.

Shah:2015:TCD

[SKB15] Imtiyaz A. Shah, A. H. Khan, and H. M. Barakat. Translation, contraction and dilation of dual generalized order statistics. Statistics & Probability Letters, 107:131–135, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003065.

Skipper:2010:LTP

[Ski10] Max Skipper. Limit theorems for projections of random walk on a hypersphere. Statistics & Probability Letters, 80 (9–10):771–778, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000143.

Skinner:2015:CCS

[Ski15] C. J. Skinner. Cross-classified sampling: Some estimation theory. Statistics & Probability Letters, 104(??):163–168, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001819.

Sepehrifar:2015:RIM

[SKJ15] Mohammad B. Sepehrifar, Kavoos Khorshidian, and Ahmad R. Jamshidian. On renewal increasing mean residual life distributions: An age replacement model with hypothesis testing application. Statistics & Probability Letters, 96(??):117–122, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003204.

Sun:2010:HBM

[SKK10] Hokeun Sun, Robert W. Keener, and Dong-Yun Kim. Hybrid bootstrap for mapping quantitative trait loci. Statistics & Probability Letters, 80(15–16):1253–1259, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001100.

Skorniakov:2019:ANC

[Sko19] V. Skorniakov. On asymptotic normality of certain linear rank statistics. Statistics & Probability Letters, 145(??):63–73, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302918. See erratum [Sko20].

Skorniakov:2020:EAN

[Sko20] V. Skorniakov. Erratum to “On asymptotic normality of certain linear rank statistics” [statist. probab. lett. 145 (2019) 63–73]. Statistics & Probability Letters, 156(??):Article 108608, January 2020. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302548.

Siriteanu:2016:CSM

[SKRT16] Constantin Siriteanu, Satoshi Kuriki, Donald Richards, and Akimichi Takemura. Chi-square mixture representations for the distribution of the scalar Schur complement in a noncentral Wishart matrix. Statistics & Probability Letters, 115(??):79– 87, August 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215304429.

Stoica:2010:KFL

[SL10] George Stoica and Deli Li. On the Kolmogorov–Feller law for exchangeable random variables. Statistics & Probability Letters, 80(9–10):899–902, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000374.

Shi:2012:NGV

[SL12] Jian-Hong Shi and Jiang-Long Lv. A new generalized p-value for testing equality of inverse Gaussian means under heterogeneity. Statistics & Probability Letters, 82(1):96–102, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002860.

Seo:2013:NUC

[SL13a] Byungtae Seo and Bruce G. Lindsay. Nearly universal consistency of maximum likelihood in discrete models. Statistics & Probability Letters, 83(7):1699–1702, July 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001089.

Song:2013:ELI

[SL13b] Yuping Song and Zhengyan Lin. Empirical likelihood inference for the second-order jump-diffusion model. Statistics & Probability Letters, 83(1):184–195, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003422.

Song:2015:EFP

[SL15] Iickho Song and Seungwon Lee. Explicit formulae for product moments of multivariate Gaussian random variables. Statistics & Probability Letters, 100(??):27–34, May 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521500036X.

Sun:2016:AUF

[SLC16] Tien-Lung Sun, Kuo-Hung Lo, and Juei-Chao Chen. An accurate updating formula to calculate sample variance from weighted successive differences. Statistics & Probability Letters, 114(??):14–19, July 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000559.

Shi:2019:MSU

[SLM19] Guiling Shi, Chae Young Lim, and Tapabrata Maiti. Model selection using mass-nonlocal prior. Statistics & Probability Letters, 147(??):36–44, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303821.

Shi:2013:CPR

[SLZ13] Yafeng Shi, Peng Liu, and Chunsheng Zhang. On the compound Poisson risk model with dependence and a threshold dividend strategy. Statistics & Probability Letters, 83(9):1998–2006, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001697.

Sankaran:2010:WTD

[SM10] P. G. Sankaran and N. N. Midhu. On waiting time distributions for patterns in a sequence of multistate trials. Statistics & Probability Letters, 80(23–24):1798–1805, December 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002270.

Sparks:2013:PTC

[SM13] Joshua Sparks and Hosam M. Mahmoud. Phases in the two-color tenable zero-balanced Pólya process. Statistics & Probability Letters, 83(1):265–271, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003252.

Schmidt:2016:MML

[SM16] Daniel F. Schmidt and Enes Makalic. Minimum message length analysis of multiple short time series. Statistics & Probability Letters, 110(??):318–328, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003399.

Schmidt:2019:ECB

[SM19] Mikkel N. Schmidt and Morten Mørup. Efficient computation for Bayesian comparison of two proportions. Statistics & Probability Letters, 145(??):57–62, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302864.

Smaga:2014:NSC

[Sma14] Lukasz Smaga. Necessary and sufficient conditions in the problem of D-optimal weighing designs with autocorrelated errors. Statistics & Probability Letters, 92(??):12–16, September 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001679.

Smaga:2015:WTS

[Sma15] Lukasz Smaga. Wald-type statistics using {2}-inverses for hypothesis testing in general factorial designs. Statistics & Probability Letters, 107:215–220, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003156.

Smucler:2019:CGD

[Smu19] Ezequiel Smucler. Consistency of generalized dynamic principal components in dynamic factor models. Statistics & Probability Letters, 155(??):Article 108536, December 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301701.

Shi:2018:NQE

[SMZ18] Jianhua Shi, Huijuan Ma, and Yong Zhou. The nonparametric quantile estimation for length-biased and right-censored data. Statistics & Probability Letters, 134(??):150–158, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303425.

Sankaran:2010:QBT

[SNS10] P. G. Sankaran, N. Unnikrishnan Nair, and E. P. Sreedevi. A quantile based test for comparing cumulative incidence functions of competing risks models. Statistics & Probability Letters, 80(9–10):886–891, May 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000350.

Sohl:2010:PSA

[Söh10] Jakob Söhl. Polar sets for anisotropic Gaussian random fields. Statistics & Probability Letters, 80(9–10):840–847, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000301.

Song:2010:LDE

[Son10a] Seongjoo Song. Lévy density estimation via information projection onto wavelet subspaces. Statistics & Probability Letters, 80(21–22):1623–1632, November 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001884.

Song:2010:MDD

[Son10b] Weixing Song. Moderate deviations for deconvolution kernel density estimators with ordinary smooth measurement errors. Statistics & Probability Letters, 80(3–4):169–176, February 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003897.

Song:2012:ABS

[Son12a] Jian Song. Asymptotic behavior of the solution of heat equation driven by fractional white noise. Statistics & Probability Letters, 82(3):614–620, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003713.

Song:2012:SCE

[Son12b] Kyungchul Song. On the smoothness of conditional expectation functionals. Statistics & Probability Letters, 82(5):1028– 1034, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000375.

Song:2014:DFE

[Son14] Yulin Song. Derivative formula and exponential convergence for semilinear SPDEs driven by Lévy processes. Statistics & Probability Letters, 89(??):99–109, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400087X.

Song:2015:NND

[Son15] Yongsheng Song. A note on G-normal distributions. Statistics & Probability Letters, 106(??):142–146, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002473.

Song:2016:AES

[Son16] Yan-Hong Song. Algebraic ergodicity for SDEs driven by Lévy processes. Statistics & Probability Letters, 119(??):108–115, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301201.

Schmitt:2014:FSB

[SÖV14] Eric Schmitt, Viktoria Öllerer, and Kaveh Vakili. The finite sample breakdown point of PCS. Statistics & Probability Letters, 94(??):214–220, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002648.

Spataru:2010:CRP

[Spa10a] Aurel Spataru. A CLT for renewal processes with a finite set of interarrival distributions. Statistics & Probability Letters, 80(21–22):1680–1683, November 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002063.

Spataru:2010:GTK

[Spa10b] Aurel Spataru. Generalizing a theorem of Katz. Statistics & Probability Letters, 80(13–14):1136–1140, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000787.

Spataru:2011:OFT

[Spa11] Aurel Spataru. Only the first term of some series counts. Statistics & Probability Letters, 81(10):1547–1551, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001878.

Su:2018:NNM

[SQY18] Liju Su, Xiaodi Qian, and Ting Yan. A note on a network model with degree heterogeneity and homophily. Statistics & Probability Letters, 138(??):27–30, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300919.

Soltani:2012:DRO

[SR12a] Ahmad Reza Soltani and Rasool Roozegar. On distribution of randomly ordered uniform incremental weighted averages: Divided difference approach. Statistics & Probability Letters, 82(5):1012–1020, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000545.

Szekely:2012:UDC

[SR12b] Gábor J. Székely and Maria L. Rizzo. On the uniqueness of distance covariance. Statistics & Probability Letters, 82 (12):2278–2282, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003124.

Sreehari:2010:ELD

[Sre10] M. Sreehari. On the equivalence of limit distributions of a sum and of a maximum sum of independent random variables. Statistics & Probability Letters, 80(9–10):860–863, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000325.

Sreehari:2012:CBN

[Sre12] M. Sreehari. A characterization of the bivariate negative binomial distribution via α-monotonicity. Statistics & Probability Letters, 82(3):433–437, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003646.

Sriram:2015:SLC

[Sri15] Karthik Sriram. A sandwich likelihood correction for Bayesian quantile regression based on the misspecified asymmetric Laplace density. Statistics & Probability Letters, 107:18–26, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500276X.

Shen:2010:CSR

[sS10a] Pao sheng Shen. A class of semiparametric rank-based tests for right-truncated data. Statistics & Probability Letters, 80(17–18):1459–1466, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001562.

Shen:2010:SES

[sS10b] Pao sheng Shen. Semiparametric estimation of survival function when data are subject to dependent censoring and left truncation. Statistics & Probability Letters, 80(3–4):161–168, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003885.

Stamatovic:2010:CLT

[SS10c] Biljana Stamatovic and Sinisa Stamatovic. Cox limit theorem for large excursions of a norm of a Gaussian vector process. Statistics & Probability Letters, 80(19–20):1479–1485, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001598.

Simon:2011:URT

[SS11a] Richard Simon and Noah Robin Simon. Using randomization tests to preserve type I error with response adaptive and covariate adaptive randomization. Statistics & Probability Letters, 81(7):767–772, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210003688.

Srihera:2011:KAD

[SS11b] Ramidha Srihera and Winfried Stute. Kernel adjusted density estimation. Statistics & Probability Letters, 81(5):571–579, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000204.

Skinner:2012:EFF

[SS12a] C. J. Skinner and N. Shlomo. Estimating frequencies of frequencies in finite populations. Statistics & Probability Letters, 82(12):2206–2212, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002957.

Sunoj:2012:QBE

[SS12b] S. M. Sunoj and P. G. Sankaran. Quantile based entropy function. Statistics & Probability Letters, 82(6):1049–1053, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000521.

Shevchenko:2013:MRS

[SS13a] Georgiy Shevchenko and Taras Shalaiko. Malliavin regularity of solutions to mixed stochastic differential equations. Statistics & Probability Letters, 83(12):2638–2646, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002885.

Singh:2013:SCI

[SS13b] Parminder Singh and Navdeep Singh. Simultaneous confidence intervals for ordered pairwise differences of exponential location parameters under heteroscedasticity. Statistics & Probability Letters, 83(12):2673–2678, December 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002939.

Sen:2014:ISF

[SS14] Arusharka Sen and Winfried Stute. Identification of survival functions through hazard functions in the Claytonfamily. Statistics & Probability Letters, 87(??):94–97, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521400025X.

Shen:2015:CMP

[sS15a] Pao sheng Shen. Conditional MLE for the proportional hazards model with left-truncated and interval-censored data. Statistics & Probability Letters, 100(??):164–171, May 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000656.

Shen:2015:IPW

[sS15b] Pao sheng Shen. The inverse probability weighted estimators for distribution functions of the bivariate recurrent events. Statistics & Probability Letters, 106(??):91–99, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002400.

Sun:2015:IEE

[SS15c] Jing Sun and Qihang Sun. An improved and efficient estimation method for varying-coefficient model with missing covariates. Statistics & Probability Letters, 107:296–303, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003284.

Shokrollahi:2017:HFB

[SS17] Foad Shokrollahi and Tommi Sottinen. Hedging in fractional Black–Scholes model with transaction costs. Statistics & Probability Letters, 130(??):85–91, November 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521730247X.

Sabzikar:2018:TFB

[SS18] Farzad Sabzikar and Donatas Surgailis. Tempered fractional Brownian and stable motions of second kind. Statistics & Probability Letters, 132(??):17–27, January 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302845.

Singer:2012:PME

[SSL^{+}12] Julio M. Singer, Edward J. Stanek III, Viviana B. Lencina, Luz Mery González, Wenjun Li, and Silvina San Martino. Prediction with measurement errors in finite populations. Statistics & Probability Letters, 82(2):332–339, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003348.

Sunoj:2013:QBE

[SSN13] S. M. Sunoj, P. G. Sankaran, and Asok K. Nanda. Quantile based entropy function in past lifetime. Statistics & Probability Letters, 83(1):366–372, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003628.

Sankaran:2016:KLD

[SSN16] P. G. Sankaran, S. M. Sunoj, and N. Unnikrishnan Nair. Kullback–Leibler divergence: A quantile approach. Statistics & Probability Letters, 111(??):72–79, April 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216000067.

Stabile:2010:LDP

[ST10] Gabriele Stabile and Giovanni Luca Torrisi. Large deviations of Poisson shot noise processes under heavy tail semiexponential conditions. Statistics & Probability Letters, 80(15– 16):1200–1209, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210000970.

Soltani:2013:CCK

[ST13a] A. R. Soltani and L. Tafakori. A class of continuous kernels and Cauchy type heavy tail distributions. Statistics & Probability Letters, 83(4):1018–1027, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200483X.

Son:2013:BPS

[ST13b] Ta Cong Son and Dang Hung Thang. The Brunk–Prokhorov Strong Law of Large Numbers for fields of martingale differences taking values in a Banach space. Statistics & Probability Letters, 83(8):1901–1910, August 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001442.

Singh:2014:DRR

[ST14] Housila P. Singh and Tanveer A. Tarray. A dexterous randomized response model for estimating a rare sensitive attribute using Poisson distribution. Statistics & Probability Letters, 90(??): 42–45, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001126.

Schumann:2018:CPW

[ST18] Martin Schumann and Gautam Tripathi. Convexity of probit weights. Statistics & Probability Letters, 143(??):81–85, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302669.

Salehi:2019:SCC

[ST19a] Ebrahim Salehi and Mahdi Tavangar. Stochastic comparisons on conditional residual lifetime and inactivity time of coherent systems with exchangeable components. Statistics & Probability Letters, 145(??):327–337, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303237.

Sirovich:2019:FPN

[ST19b] Roberta Sirovich and Luisa Testa. On the first positive and negative excursion exceeding a given length. Statistics & Probability Letters, 150(??):137–145, July 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300847.

Stadje:2012:EMC

[Sta12] Wolfgang Stadje. Embedded Markov chain analysis of the superposition of renewal processes. Statistics & Probability Letters, 82(8):1497–1503, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001538.

Stark:2016:STI

[Sta16] Dudley Stark. Bin sizes in time-inhomogeneous infinite Polya processes. Statistics & Probability Letters, 113(??):49–53, June 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303588.

Son:2012:RCC

[STD12] Ta Cong Son, Dang Hung Thang, and Le Van Dung. Rate of complete convergence for maximums of moving average sums of martingale difference fields in Banach spaces. Statistics & Probability Letters, 82(11):1978–1985, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002258.

Stepanov:2011:LTR

[Ste11] A. Stepanov. Limit theorems for runs based on ‘small spacings’. Statistics & Probability Letters, 81(1):54–61, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002683.

Stepanov:2014:UBC

[Ste14] Alexei Stepanov. On the use of the Borel–Cantelli lemma in Markov chains. Statistics & Probability Letters, 90(??):149– 154, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001187.

Stepniak:2015:DLT

[Ste15] Czeslaw Stepniak. On distribution of the leadership time in counting votes and predicting winners. Statistics & Probability Letters, 106(??):109–112, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002321.

Steinerberger:2019:SEP

[Ste19] Stefan Steinerberger. A sharp estimate for probability distributions. Statistics & Probability Letters, 155(??):Article 108584, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219302305.

Stojakovic:2012:SVP

[Sto12] Mila Stojaković. Set valued probability and its connection with set valued measure. Statistics & Probability Letters, 82(6):1043– 1048, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000685.

Song:2019:MGF

[STW19] Pingfan Song, Changchun Tan, and Shaochen Wang. On the moment generating function for random vectors via inverse survival function. Statistics & Probability Letters, 145(??):345–350, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303365.

Su:2010:SOC

[Su10] Zhonggen Su. On the second-order correlation of characteristic polynomials of Hermite β ensembles. Statistics & Probability Letters, 80(19–20):1500–1507, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001628.

Subramanian:2012:MBL

[Sub12]  Sundarraman Subramanian.  Model-based likelihood ra-

tio confidence intervals for survival functions. Statistics & Probability Letters, 82(3):626–635, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003828.

Sugrue:2019:IAN

[Sug19] Darren Sugrue. Introducing article numbering to Statistics and Probability Letters. Statistics & Probability Letters, 155(??): Article 108573, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219302196.

Sun:2013:LDT

[Sun13a] Hongyan Sun. A large deviation theorem for a branching Brownian motion with random immigration. Statistics & Probability Letters, 83(6):1559–1566, June 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000771.

Sung:2013:SLL

[Sun13b] Soo Hak Sung. On the Strong Law of Large Numbers for pairwise i.i.d. random variables with general moment conditions. Statistics & Probability Letters, 83(9):1963–1968, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001703.

Sun:2017:ATI

[Sun17] Yan Sun. Asymptotic tests for interval-valued means. Statistics & Probability Letters, 121(??):70–77, February 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300833.

Suzuki:2018:MDI

[Suz18] Ryoichi Suzuki. Malliavin differentiability of indicator functions on canonical Lévy spaces. Statistics & Probability Letters, 137(??):183–190, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521830035X.

See correction [Suz20].

Suzuki:2020:CMD

[Suz20] Ryoichi Suzuki. Correction to “Malliavin differentiability of indicator functions on canonical Lévy spaces” [statist. probab. lett. 137 (2018) 183–190]. Statistics & Probability Letters, 156(??):Article 108614, January 2020. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219302603.

Schwarz:2010:CDD

[SV10] Maik Schwarz and Sébastien Van Bellegem. Consistent density deconvolution under partially known error distribution. Statistics & Probability Letters, 80(3–4):236–241, February 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003988.

Sinova:2015:CST

[SV15] Beatriz Sinova and Stefan Van Aelst. On the consistency of a spatial-type interval-valued median for random intervals. Statistics & Probability Letters, 100(??):130–136, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000644.

Sottinen:2017:PLF

[SV17] Tommi Sottinen and Lauri Viitasaari. Prediction law of fractional Brownian motion. Statistics & Probability Letters, 129(??):155–166, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301815.

Salminen:2019:MIE

[SV19] Paavo Salminen and Lioudmila Vostrikova. On moments of integral exponential functionals of additive processes. Statistics & Probability Letters, 146(??):139–146, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303584.

Shneer:2017:PSO

[SvdV17] Seva Shneer and Peter M. van de Ven. Per-site occupancy in the discrete parking problem. Statistics & Probability Letters, 120(??):141–146, January 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301833.

Scott:2011:GAP

[SW11] Alexandre Scott and François Watier. Goal achieving probabilities of constrained mean-variance strategies. Statistics & Probability Letters, 81(8):1021–1026, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000691.

Schick:2012:EED

[SW12] Anton Schick and Wolfgang Wefelmeyer. On efficient estimation of densities for sums of squared observations. Statistics & Probability Letters, 82(9):1637–1640, September 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001642.

Sharipov:2013:NLN

[SW13] Olimjon Sh. Sharipov and Martin Wendler. Normal limits, nonnormal limits, and the bootstrap for quantiles of dependent data. Statistics & Probability Letters, 83(4):1028–1035, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004762.

Stoev:2019:ERP

[SW19] Stilian Stoev and Yizao Wang. Exchangeable random partitions from max-infinitely-divisible distributions. Statistics & Probability Letters, 146(??):50–56, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303559.

Senn:2011:CSM

[SWH^{+}11] Stephen Senn, James Weir, Tsushung A. Hua, Conny Berlin, Michael Branson, and Ekkehard Glimm. Creating a suite of macros for meta-analysis in SAS(R): a case study in collaboration. Statistics & Probability Letters, 81(7):842–851, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000484.

Su:2012:RMO

[SWH12] Xiaonan Su, Wensheng Wang, and Kyo-Shin Hwang. Riskminimizing option pricing under a Markov-modulated jumpdiffusion model with stochastic volatility. Statistics & Probability Letters, 82(10):1777–1785, October 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200209X.

Shi:2010:NAA

[SWL10] Xiaoping Shi, Yuehua Wu, and Yu Liu. A note on asymptotic approximations of inverse moments of nonnegative random variables. Statistics & Probability Letters, 80(15–16):1260–1264, August 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210001112.

Schick:2011:TNB

[SWW11] Anton Schick, Yishi Wang, and Wolfgang Wefelmeyer. Tests for normality based on density estimators of convolutions. Statistics & Probability Letters, 81(2):337–343, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003081.

Song:2015:FHT

[SWW15] Shiyu Song, Suxin Wang, and Yongjin Wang. First hitting times for doubly skewed Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 96(??):212–222, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400340X.

Shi:2019:ACD

[SWZ19] X. Shi, A. Wong, and S. Zheng. Approximations of the cumulative distribution function for infinite weighted sum of random variables. Statistics & Probability Letters, 155(??): Article 108567, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219302135.

Shen:2013:SCI

[SX13] Jia Shen and Yuan Xie. Strong consistency of the internal estimator of nonparametric regression with dependent data. Statistics & Probability Letters, 83(8):1915–1925, August 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521300148X.

Shen:2016:PLD

[SXM16] Xinmei Shen, Menghao Xu, and Ebenezer Fiifi Emire Atta Mills. Precise large deviation results for sums of subexponential claims in a size-dependent renewal risk model. Statistics & Probability Letters, 114(??):6–13, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216000444.

Shi:2010:SLP

[SY10] Zhiyan Shi and Weiguo Yang. Some limit properties for the m th-order nonhomogeneous Markov chains indexed by an m rooted Cayley tree. Statistics & Probability Letters, 80(15–16):1223–1233, August 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001008.

Song:2011:LFT

[SY11] Weixing Song and Weixin Yao. A lack-of-fit test in Tobit errors-in-variables regression models. Statistics & Probability Letters, 81(12):1792–1801, December 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521100246X.

Shi:2017:FSD

[SY17] Qun Shi and Xianye Yu. Fractional smoothness of derivative of self-intersection local times. Statistics & Probability Letters, 129(??):241–251, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302110.

Shen:2018:EML

[SY18] Yuncong Shen and Jiangang Ying. Excessive measures for linear diffusions. Statistics & Probability Letters, 135(??):132– 139, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217303644.

Salah:2019:NRR

[SY19a] Khardani Salah and Slaoui Yousri. Nonparametric relative regression under random censorship model. Statistics & Probability Letters, 151(??):116–122, August 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300951.

Sun:2019:IMO

[SY19b] Wenjie Sun and Jiangang Ying. Intrinsic metric and onedimensional diffusions. Statistics & Probability Letters, 150(??): 146–151, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300859.

Sendova:2018:DBS

[SYZ18] Kristina P. Sendova, Chen Yang, and Ruixi Zhang. Dividend barrier strategy: Proceed with caution. Statistics & Probability Letters, 137(??):157–164, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300270.

Shen:2012:MDR

[SZ12] Xinmei Shen and Yi Zhang. Moderate deviations for a risk model based on the customer-arrival process. Statistics & Probability Letters, 82(1):116–122, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002975.

Shen:2013:RPT

[SZ13a] Xinmei Shen and Yi Zhang. Ruin probabilities of a twodimensional risk model with dependent risks of heavy tail. Statistics & Probability Letters, 83(7):1787–1799, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001120.

Sun:2013:SAA

[SZ13b] Zhimeng Sun and Zhongzhan Zhang. Semiparametric analysis of additive isotonic errors-in-variables regression models. Statistics & Probability Letters, 83(1):100–114, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003331.

Sun:2018:LDP

[SZ18] Hongyan Sun and Lin Zhang. Large deviation principle for the maximal positions in critical branching random walks with small drifts. Statistics & Probability Letters, 139(??):31–39, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301317.

Szablowski:2015:MNC

[Sza15] Pawel J. Szablowski. Moments of q-normal and conditional q-normal distributions. Statistics & Probability Letters, 106 (??):65–72, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521500231X.

Szablowski:2019:PAC

[Sza19] Pawel J. Szablowski. On probabilistic aspects of Chebyshev polynomials. Statistics & Probability Letters, 145(??):205–215, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521830292X.

Szewczak:2010:LTR

[Sze10] Zbigniew S. Szewczak. A limit theorem for random sums modulo 1. Statistics & Probability Letters, 80(9–10):747–751, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000106.

Szewczak:2011:NML

[Sze11] Zbigniew S. Szewczak. A note on Marcinkiewicz laws for strictly stationary ϕ-mixing sequences. Statistics & Probability Letters, 81(11):1738–1741, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002033.

Szewczak:2012:DI

[Sze12] Zbigniew S. Szewczak. On Dobrushin’s inequality. Statistics & Probability Letters, 82(6):1202–1207, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000661.

Szewczak:2015:MMI

[Sze15] Zbigniew S. Szewczak. A moment maximal inequality for dependent random variables. Statistics & Probability Letters, 106(??):129–133, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002436.

Szewczak:2016:CMS

[Sze16] Zbigniew S. Szewczak. Convergence of moments for strictly stationary sequences. Statistics & Probability Letters, 119(??):200– 203, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301481.

Song:2010:SLS

[SZW10] Lixin Song, Yue Zhao, and Xiaoguang Wang. Sieve least squares estimation for partially nonlinear models. Statistics & Probability Letters, 80(17–18):1271–1283, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001136.

Shang:2012:LWA

[SZW12] Suoping Shang, Changliang Zou, and Zhaojun Wang. Local Walsh-average regression for semiparametric varying-coefficient models. Statistics & Probability Letters, 82(10):1815–1822, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212002131.

Torehzadeh:2014:NSW

[TA14] S. Torehzadeh and M. Arashi. A note on shrinkage wavelet estimation in Bayesian analysis. Statistics & Probability Letters, 84(??):231–234, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213003404.

Tabacu:2018:WCL

[Tab18] Lucia Tabacu. Weak convergence of the linear rank statistics under strong mixing conditions. Statistics & Probability Letters, 132(??):28–34, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302870.

Taheriyoun:2012:TCF

[Tah12] Ali Reza Taheriyoun. Testing the covariance function of stationary Gaussian random fields. Statistics & Probability Letters, 82(3):606–613, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003671.

Tahmasebi:2014:SDS

[Tah14] M. Tahmasebi. Smooth density for the solution of scalar SDEs with locally Lipschitz coefficients under Hörmander condition. Statistics & Probability Letters, 85(??):51–62, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003842.

Takagi:2012:ESP

[Tak12] Yoshiji Takagi. On the estimation of the shape parameter of the gamma distribution in second-order asymptotics. Statistics & Probability Letters, 82(1):15–21, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002896.

Takeuchi:2019:IPF

[Tak19] Atsushi Takeuchi. Integration by parts formulas for marked Hawkes processes. Statistics & Probability Letters, 145(??):229– 237, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303195.

Tan:2013:ASL

[Tan13a] Zhongquan Tan. An almost sure limit theorem for the maxima of smooth stationary Gaussian processes. Statistics & Probability Letters, 83(9):2135–2141, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002022.

Tan:2013:LTE

[Tan13b] Zhongquan Tan. The limit theorems on extremes for Gaussian random fields. Statistics & Probability Letters, 83(2):436–444, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003975.

Tang:2015:EMD

[Tan15a] Yongqiang Tang. An efficient monotone data augmentation algorithm for Bayesian analysis of incomplete longitudinal data. Statistics & Probability Letters, 104(??):146–152, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001716.

Tanguy:2015:SSI

[Tan15b] Kevin Tanguy. Some superconcentration inequalities for extrema of stationary Gaussian processes. Statistics & Probability Letters, 106(??):239–246, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002692.

Tan:2019:DRE

[Tan19] Zhiqiang Tan. On doubly robust estimation for logistic partially linear models. Statistics & Probability Letters, 155(??): Article 108577, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219302238.

Tao:2019:LTM

[Tao19] Yubo Tao. Limit theory for moderate deviation from Integrated GARCH processes. Statistics & Probability Letters, 150(??):126– 136, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521930077X.

Tappe:2010:NSI

[Tap10] Stefan Tappe. A note on stochastic integrals as L^{2}-curves. Statistics & Probability Letters, 80(13–14):1141–1145, July 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000799.

Taufer:2015:EPS

[Tau15] Emanuele Taufer. On the empirical process of strongly dependent stable random variables: asymptotic properties, simulation and applications. Statistics & Probability Letters, 106(??):262–271, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002734.

Terpstra:2012:LLN

[TE12a] Jeff T. Terpstra and Tamer Elbayoumi. A Law of Large Numbers result for a bifurcating process with an infinite moving average representation. Statistics & Probability Letters, 82(1):123– 129, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003002.

Tzavelas:2012:CPL

[TE12b] George Tzavelas and Polychronis Economou. Characterization properties of the log-normal distribution obtained with the help of divergence measures. Statistics & Probability Letters, 82(10):1837–1840, October 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002143.

Tzavelas:2014:CPB

[TE14] George Tzavelas and Polychronis Economou. Characterization properties based on the Fisher information for weighted distributions. Statistics & Probability Letters, 84(??):54–59, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003209.

Tahmasebi:2017:GCE

[TE17] Saeid Tahmasebi and Maryam Eskandarzadeh. Generalized cumulative entropy based on k th lower record values. Statistics & Probability Letters, 126(??):164–172, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300925.

Teran:2015:CCL

[Ter15] Pedro Terán. Counterexamples to a Central Limit Theorem and a Weak Law of Large Numbers for capacities. Statistics & Probability Letters, 96(??):185–189, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002867.

Terhesiu:2017:NTL

[Ter17] Dalia Terhesiu. Non trivial limit distributions for transient renewal chains. Statistics & Probability Letters, 129(??):189–195, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301888.

Tu:2014:ABS

[TH14] Shuheng Tu and Wu Hao. Anticipated backward stochastic differential equations with jumps under the non-Lipschitz condition. Statistics & Probability Letters, 92(??):215–225, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521400203X.

Tanabe:2016:OPZ

[TH16] Ryunosuke Tanabe and Etsuo Hamada. Objective priors for the zero-modified model. Statistics & Probability Letters, 112(??): 92–97, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215300602.

Tu:2017:ASC

[THC17] Shuheng Tu, Wu Hao, and Jing Chen. The adapted solutions and comparison theorem for anticipated backward stochastic differential equations with Poisson jumps under the weak conditions. Statistics & Probability Letters, 126(??):7–17, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300780.

Thulin:2014:SSR

[Thu14] Måns Thulin. On split sample and randomized confidence intervals for binomial proportions. Statistics & Probability Letters, 92(??):65–71, September 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214001813.

Tian:2016:NCR

[Tia16] Tian Tian. A note on continual reassessment method. Statistics & Probability Letters, 113(??):94–102, June 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216000432.

Tien:2013:EPS

[Tie13] Dung Nguyen Tien. The existence of a positive solution for a generalized delay logistic equation with multifractional noise. Statistics & Probability Letters, 83(4):1240–1246, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004889.

Timmermann:2015:SDG

[Tim15] Hella Timmermann. Sequential detection of gradual changes in the location of a general stochastic process. Statistics & Probability Letters, 99(??):85–93, April 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000073.

Tian:2010:ESB

[TJD10] Dejian Tian, Long Jiang, and Matt Davison. On the existence of solutions to BSDEs with generalized uniformly continuous generators. Statistics & Probability Letters, 80(9–10):903–909, May 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000386.

Tahmasebi:2016:SRR

[TJE16] Saeid Tahmasebi, Ali Akbar Jafari, and Maryam Eskandarzadeh. Some results on residual entropy of ranked set samples. Statistics & Probability Letters, 112(??):137–145, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000158.

Tian:2013:SBS

[TJS13] Dejian Tian, Long Jiang, and Xuejun Shi. L^{p} solutions to backward stochastic differential equations with discontinuous generators. Statistics & Probability Letters, 83(2):503–510, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003987.

Tan:2015:SDN

[TJS15] Li Tan, Wei Jin, and Yongqiang Suo. Stability in distribution of neutral stochastic functional differential equations. Statistics & Probability Letters, 107:27–36, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002746.

Tyran-Kaminska:2010:FLT

[TK10] Marta Tyran-Kamińska. Functional limit theorems for linear processes in the domain of attraction of stable laws. Statistics & Probability Letters, 80(11–12):975–981, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000519.

Taskinen:2012:RPC

[TKO12] Sara Taskinen, Inge Koch, and Hannu Oja. Robustifying principal component analysis with spatial sign vectors. Statistics & Probability Letters, 82(4):765–774, April 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000028.

Tang:2012:ELA

[TL12] Cheng Yong Tang and Chenlei Leng. An empirical likelihood approach to quantile regression with auxiliary information. Statistics & Probability Letters, 82(1):29–36, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002902.

Tang:2015:CIC

[TL15] Nian-Sheng Tang and Xian-Gui Luo. Confidence interval construction for sensitivity difference of two continuous-scale diagnostic tests at the fixed level of two specificities. Statistics & Probability Letters, 97(??):32–40, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003654.

Tabacu:2019:CPA

[TL19] Lucia Tabacu and Mark Ledbetter. Change-point analysis using logarithmic quantile estimation. Statistics & Probability Letters, 150(??):94–100, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300665.

Thavaneswaran:2012:IRC

[TLF12] A. Thavaneswaran, You Liang, and Julieta Frank. Inference for random coefficient volatility models. Statistics & Probability Letters, 82(12):2086–2090, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002805.

Tokola:2011:PSS

[TLNO11] K. Tokola, D. Larocque, J. Nevalainen, and H. Oja. Power, sample size and sampling costs for clustered data. Statistics & Probability Letters, 81(7):852–860, July 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000447.

Tong:2012:LTM

[TLZ12] Changqing Tong, Zhengyan Lin, and Jing Zheng. The local time of the Markov processes of Ornstein–Uhlenbeck type. Statistics & Probability Letters, 82(7):1229–1234, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000752.

Tayob:2016:NRM

[TM16] Nabihah Tayob and Susan Murray. Nonparametric restricted mean analysis across multiple follow-up intervals. Statistics & Probability Letters, 109:152–158, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521500382X.

Taskinen:2016:MES

[TMN16] Sara Taskinen, Jari Miettinen, and Klaus Nordhausen. A more efficient second order blind identification method for separation of uncorrelated stationary time series. Statistics & Probability Letters, 116(??):21–26, September 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300256.

Trottini:2011:MTD

[TMS11] Mario Trottini, Krish Muralidhar, and Rathindra Sarathy. Maintaining tail dependence in data shuffling using t copula. Statistics & Probability Letters, 81(3):420–428, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003421.

Talebi:2019:RNB

[TMS19] Ali Talebi, M. S. Moslehian, and Ghadir Sadeghi. A representation of noncommutative BMO spaces. Statistics & Probability Letters, 153(??):65–70, October 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301324.

Tuan:2017:ISP

[TN17] Nguyen Huy Tuan and Erkan Nane. Inverse source problem for time-fractional diffusion with discrete random noise. Statistics & Probability Letters, 120(??):126–134, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301882.

Trong:2017:DCS

[TNP17] Dang Duc Trong, Ton That Quang Nguyen, and Cao Xuan Phuong. Deconvolution of P(X < Y) with compactly supported error densities. Statistics & Probability Letters, 123(??):171– 176, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521630308X.

Tomczak:2016:SPL

[Tom16] Jakub M. Tomczak. On some properties of the low-dimensional Gumbel perturbations in the Perturb-and-MAP model. Statistics & Probability Letters, 115(??):8–15, August 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300189.

Tone:2016:CLT

[Ton16] Cristina Tone. A central limit theorem for quadruplewise independent arrays of random variables. Statistics & Probability Letters, 110(??):58–61, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215300109. See corrigendum [Ton17].

Tone:2017:CCL

[Ton17] Cristina Tone. Corrigendum to “A central limit theorem for quadruple-wise independent arrays of random variables [Statist. Probab. Lett. 110 (2016) 58–61]. Statistics & Probability Letters, 121(??):163, February 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302176. See [Ton16].

Torrisi:2013:FSL

[Tor13] Giovanni Luca Torrisi. Functional Strong Law of Large Numbers for loads in a planar network model. Statistics & Probability Letters, 83(3):718–723, March 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004294.

Tomazella:2019:ZAR

[TPNSN19] Vera Tomazella, Gustavo H. A. Pereira, Juvêncio S. Nobre, and Manoel Santos-Neto. Zero-adjusted reparameterized Birnbaum–Saunders regression model. Statistics & Probability Letters, 149(??):142–145, June 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300252.

Tchetgen:2012:PRS

[TR12] Eric J. Tchetgen Tchetgen and James Robins. On parametrization, robustness and sensitivity analysis in a marginal structural Cox proportional hazards model for point exposure. Statistics & Probability Letters, 82(5):907–915, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000296.

Torrecilla:2018:DLB

[TR18] José L. Torrecilla and Juan Romo. Data learning from big data. Statistics & Probability Letters, 136(??):15–19, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830083X.

Trabs:2014:IDD

[Tra14] Mathias Trabs. On infinitely divisible distributions with polynomially decaying characteristic functions. Statistics & Probability Letters, 94(??):56–62, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002387.

Triacca:2018:GCB

[Tri18] Umberto Triacca. Granger causality between vectors of time series: a puzzling property. Statistics & Probability Letters, 142(??):39–43, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302323.

Truquet:2010:MIM

[Tru10] Lionel Truquet. A moment inequality of the Marcinkiewicz– Zygmund type for some weakly dependent random fields. Statistics & Probability Letters, 80(21–22):1673–1679, November 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002051.

Trutschnig:2013:CCI

[Tru13] Wolfgang Trutschnig. On Cesáro convergence of iterates of the star product of copulas. Statistics & Probability Letters, 83(1):357–365, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003719.

Tsuruta:2017:HOK

[TS17] Yasuhito Tsuruta and Masahiko Sagae. Higher order kernel density estimation on the circle. Statistics & Probability Letters, 131(??):46–50, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302638.

Tsao:2013:EDF

[Tsa13] Min Tsao. An empirical depth function for multivariate data. Statistics & Probability Letters, 83(1):213–218, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003446.

Tsionas:2012:EMH

[Tsi12] Efthymios G. Tsionas. Estimating multivariate heavy tails and principal directions easily, with an application to international exchange rates. Statistics & Probability Letters, 82 (11):1986–1989, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002647.

Tsionas:2017:NIT

[Tsi17] Mike G. Tsionas. A non-iterative (trivial) method for posterior inference in stochastic volatility models. Statistics & Probability Letters, 126(??):83–87, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300913.

Toker:2013:ICR

[TSK13] Selma Toker, Gülesen Üstündag Siray, and Selahattin Kaçiranlar. Inequality constrained ridge regression estimator. Statistics & Probability Letters, 83(10):2391–2398, October 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002356.

Tao:2018:EAC

[TSL18] Yanfang Tao, Biqin Song, and Luoqing Li. Error analysis for coefficient-based regularized regression in additive models. Statistics & Probability Letters, 134(??):22–28, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303152.

Tsukuma:2010:SME

[Tsu10] Hisayuki Tsukuma. Shrinkage minimax estimation and positivepart rule for a mean matrix in an elliptically contoured distribution. Statistics & Probability Letters, 80(3–4):215–220, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003952.

Tang:2015:TET

[TSZ15] Yanlin Tang, Xinyuan Song, and Zhongyi Zhu. Threshold effect test in censored quantile regression. Statistics & Probability Letters, 105(??):149–156, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001807.

Thapliyal:2015:RIO

[TT15] Richa Thapliyal and H. C. Taneja. On residual inaccuracy of order statistics. Statistics & Probability Letters, 97(??):125–131, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003794.

Tan:2017:MSH

[TT17] Zhongquan Tan and Linjun Tang. On the maxima and sums of homogeneous Gaussian random fields. Statistics & Probability Letters, 125(??):44–54, June 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300482.

Turcanu:2017:SAG

[TU17] Teodor Turcanu and Constantin Udriste. Stochastic accessibility on Grushin-type manifolds. Statistics & Probability Letters, 125(??):196–201, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300585.

Tudor:2014:CEA

[Tud14] Ciprian A. Tudor. Chaos expansion and asymptotic behavior of the Pareto distribution. Statistics & Probability Letters, 91(??): 62–68, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001436.

Tumlinson:2015:NEM

[Tum15] Samuel E. Tumlinson. On the non-existence of maximum likelihood estimates for the extended exponential power distribution and its generalizations. Statistics & Probability Letters, 107:111– 114, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002886.

Turchyn:2014:SSS

[Tur14] Ievgen Turchyn. Simulation of a strictly sub-Gaussian random field. Statistics & Probability Letters, 92(??):183–189, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001989.

Ta:2017:SPB

[TV17a] Bao Quoc Ta and Chung Pham Van. Some properties of bivariate Schur-constant distributions. Statistics & Probability Letters, 124(??):69–76, May 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300275.

Tille:2017:RVD

[TV17b] Yves Tillé and Audrey-Anne Vallée. Revisiting variance decomposition when independent samples intersect. Statistics & Probability Letters, 130(??):71–75, November 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302456.

Turner:2015:PSD

[TW15] Amanda Turner and John Whitehead. Partial stochastic dominance for the multivariate Gaussian distribution. Statistics & Probability Letters, 103(??):80–85, August 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001297.

Tian:2017:SPM

[TW17] Yongge Tian and Cheng Wang. On simultaneous prediction in a multivariate general linear model with future observations. Statistics & Probability Letters, 128(??):52–59, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301384.

Tarr:2015:DSQ

[TWM15] G. Tarr, N. C. Weber, and S. Müller. The difference of symmetric quantiles under long range dependence. Statistics & Probability Letters, 98(??):144–150, March 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400426X.

Tang:2018:ANW

[TXWW18] Xufei Tang, Mengmei Xi, Yi Wu, and Xuejun Wang. Asymptotic normality of a wavelet estimator for asymptotically negatively associated errors. Statistics & Probability Letters, 140(??):191–201, September 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301810.

Tian:2016:AVS

[TXX16] Ruiqin Tian, Liugen Xue, and Dengke Xu. Automatic variable selection for varying coefficient models with longitudinal data. Statistics & Probability Letters, 119(??):84–90, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301274.

Thanh:2015:WSS

[TY15] Le Van Thanh and G. Yin. Weighted sums of strongly mixing random variables with an application to nonparametric regression. Statistics & Probability Letters, 105(??):195–202, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001881.

Tykesson:2011:VWB

[Tyk11] Johan Tykesson. On the visibility in well-behaved random sets in Euclidean space. Statistics & Probability Letters, 81(8):1256– 1259, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001088.

Takasu:2018:SRS

[TYK18] Yuya Takasu, Keisuke Yano, and Fumiyasu Komaki. Scoring rules for statistical models on spheres. Statistics & Probability Letters, 138(??):111–115, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300993.

Tyler:2010:NML

[Tyl10] David E. Tyler. A note on multivariate location and scatter statistics for sparse data sets. Statistics & Probability Letters, 80(17–18):1409–1413, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001422.

Thas:2015:PST

[TYNZ15] Olivier Thas, Ao Yuan, Hon Keung Tony Ng, and Gang Zheng. A proportional score test over the nuisance parameter space: Properties and applications. Statistics & Probability Letters, 106(??):295–300, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002631.

Tyurin:2012:SOB

[Tyu12] I. S. Tyurin. Some optimal bounds in the Central Limit Theorem using zero biasing. Statistics & Probability Letters, 82(3):514– 518, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003634.

Tian:2016:CAO

[TZ16] Yongge Tian and Xuan Zhang. On connections among OLSEs and BLUEs of whole and partial parameters under a general linear model. Statistics & Probability Letters, 112(??):105– 112, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215302339.

Tong:2013:SDF

[TZB13] Jinying Tong, Zhenzhong Zhang, and Jianhai Bao. The stationary distribution of the facultative population model with a degenerate noise. Statistics & Probability Letters, 83(2):655– 664, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004105.

Tzioufas:2017:RFH

[Tzi17] Achillefs Tzioufas. The randomly fluctuating hyperrectangles are spatially monotone. Statistics & Probability Letters, 130(??):76– 79, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217302493.

Tian:2016:ELC

[TZT16] Yuzhu Tian, Qianqian Zhu, and Maozai Tian. Estimation of linear composite quantile regression using EM algorithm. Statistics & Probability Letters, 117(??):183–191, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300700.

Tang:2012:WCQ

[TZW12] Linjun Tang, Zhangong Zhou, and Changchun Wu. Weighted composite quantile estimation and variable selection method for censored regression model. Statistics & Probability Letters, 82(3):653–663, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003750.

Tang:2013:TLE

[TZW13] Linjun Tang, Zhangong Zhou, and Changchun Wu. Testing the linear errors-in-variables model with randomly censored data. Statistics & Probability Letters, 83(3):875–884, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004592.

Uematsu:2016:AEO

[Uem16] Yoshimasa Uematsu. Asymptotic efficiency of the OLS estimator with singular limiting sample moment matrices. Statistics & Probability Letters, 114(??):104–110, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300177.

Upadhye:2018:PBA

[UK18] N. S. Upadhye and A. N. Kumar. Pseudo-binomial approximation to (k_{1}, k_{2})-runs. Statistics & Probability Letters, 141(??):19– 30, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302037.

Uno:2013:ATT

[Uno13] Chikara Uno. Asymptotic theory for a two-stage procedure in sequential interval estimation of a normal mean. Statistics & Probability Letters, 83(5):1420–1423, May 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000503.

Urban:2012:MPA

[Urb12] Roman Urban. Markov processes on the Adeles and Dedekind’s zeta function. Statistics & Probability Letters, 82(8):1583–1589, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001617.

Ushakov:2011:OCS

[Ush11a] N. G. Ushakov. One characterization of symmetry. Statistics & Probability Letters, 81(5):614–617, May 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003603.

Ushakov:2011:SIA

[Ush11b] N. G. Ushakov. Some inequalities for absolute moments. Statistics & Probability Letters, 81(12):2011–2015, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002811.

Ushakov:2011:CMG

[UU11] N. G. Ushakov and V. G. Ushakov. On convergence of moment generating functions. Statistics & Probability Letters, 81(4):502– 505, April 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210003536.

Upadhye:2013:IBA

[UV13] N. S. Upadhye and P. Vellaisamy. Improved bounds for approximations to compound distributions. Statistics & Probability Letters, 83(2):467–473, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003938.

Volfovsky:2016:STV

[VA16] Alexander Volfovsky and Edoardo M. Airoldi. Sharp total variation bounds for finitely exchangeable arrays. Statistics & Probability Letters, 114(??):54–59, July 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216000390.

Vardar-Acar:2015:BEV

[VAB15] Ceren Vardar-Acar and Hatice Bulut. Bounds on the expected value of maximum loss of fractional Brownian motion. Statistics & Probability Letters, 104(??):117–122, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001509.

VanBever:2016:SBT

[Van16] Germain Van Bever. Simplicial bivariate tests for randomness. Statistics & Probability Letters, 112(??):20–25, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000092.

Vantini:2018:WNP

[Van18] Simone Vantini. Wishing the non-parametric re-evolution. Statistics & Probability Letters, 136(??):139–141, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300907.

Vasudeva:2018:FLI

[Vas18] R. Vasudeva. Functional law of the iterated logarithm for partial sum processes of non-negative valued iid random variables belonging to the domain of partial attraction of a semistable law. Statistics & Probability Letters, 137(??):34–39, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303838.

Valdes:2010:SSC

[VAZB10] José E. Valdés, Gerardo Arango, Romulo I. Zequeira, and Gerandy Brito. Some stochastic comparisons in series systems with active redundancy. Statistics & Probability Letters, 80(11–12):945–949, June 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000453.

Vidal:2011:BER

[VB11] Ignacio Vidal and Heleno Bolfarini. Bayesian estimation of regression parameters in elliptical measurement error models. Statistics & Probability Letters, 81(9):1398–1406, September 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001544.

Vecera:2017:AAS

[VB17] Jakub Vecera and Viktor Benes. Approaches to asymptotics for U-statistics of Gibbs facet processes. Statistics & Probability Letters, 122(??):51–57, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302231.

vanBentum:2019:SMM

[vBC19] Thomas van Bentum and Erhard Cramer. Stochastic monotonicity of MLEs of the mean for exponentially distributed lifetimes under hybrid censoring. Statistics & Probability Letters, 148 (??):1–8, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218304000.

V:2014:LPM

[VCM14] José M. Martínez V., Reinaldo A. Vallejos C., and Marta Barría M. On the limiting probabilities of the M/E_{r} /1 queueing system. Statistics & Probability Letters, 88(??):56–61, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000479.

Vujacic:2018:CDI

[VD18] Ivan Vujačić and Itai Dattner. Consistency of direct integral estimator for partially observed systems of ordinary differential equations. Statistics & Probability Letters, 132(??):40–45, January 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217302730.

vanDoorn:2019:BEK

[vDLMW19] Johnny van Doorn, Alexander Ly, Maarten Marsman, and Eric-Jan Wagenmakers. Bayesian estimation of Kendall’s τ using a latent normal approach. Statistics & Probability Letters, 145(??):268–272, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218303201.

Velilla:2012:NSQ

[Vel12] Santiago Velilla. A note on the structure of the quadratic subspace in discriminant analysis. Statistics & Probability Letters, 82(4):739–747, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211004123.

Vetter:2014:ILM

[Vet14] Mathias Vetter. Inference on the Lévy measure in case of noisy observations. Statistics & Probability Letters, 87(??):125– 133, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000194.

Villasenor:2015:TFI

[VGE15a] José A. Villaseñor and Elizabeth González-Estrada. Tests of fit for inverse Gaussian distributions. Statistics & Probability Letters, 105(??):189–194, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002035.

Villasenor:2015:VRT

[VGE15b] José A. Villaseñor and Elizabeth González-Estrada. A variance ratio test of fit for Gamma distributions. Statistics & Probability Letters, 96(??):281–286, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003502.

Vidmar:2017:PTS

[Vid17] Matija Vidmar. Proper two-sided exits of a Lévy process. Statistics & Probability Letters, 125(??):160–163, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303655.

Vidmar:2018:CRC

[Vid18] Matija Vidmar. A couple of remarks on the convergence of σ-fields on probability spaces. Statistics & Probability Letters, 134(??):86–92, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303395.

Vidmar:2019:OLP

[Vid19] Matija Vidmar. Observing a Lévy process up to a stopping time. Statistics & Probability Letters, 150(??):121–125, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300781.

Vieu:2018:DRM

[Vie18] Philippe Vieu. On dimension reduction models for functional data. Statistics & Probability Letters, 136(??):134–138, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300774.

Vignat:2012:GIT

[Vig12] C. Vignat. A generalized Isserlis theorem for location mixtures of Gaussian random vectors. Statistics & Probability Letters, 82(1):67–71, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002951.

Viitasaari:2016:RSS

[Vii16] Lauri Viitasaari. Representation of stationary and stationary increment processes via Langevin equation and self-similar processes. Statistics & Probability Letters, 115(??):45–53, August 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300190.

Vila:2011:NMS

[Vil11] Jean-Pierre Vila. Nonparametric multi-step prediction in nonlinear state space dynamic systems. Statistics & Probability Letters, 81(1):71–76, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002701.

Vila:2012:ECR

[Vil12] Jean-Pierre Vila. Enhanced consistency of the Resampled Convolution Particle Filter. Statistics & Probability Letters, 82(4): 786–797, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000041.

Vinogradov:2011:KRR

[Vin11] Vladimir Vinogradov. On Kendall–Ressel and related distributions. Statistics & Probability Letters, 81(10):1493–1501, October 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001842.

Virta:2016:OSE

[Vir16] J. Virta. One-step M-estimates of scatter and the independence property. Statistics & Probability Letters, 110(??):133– 136, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215303977.

vanLieshout:2016:LBI

[vL16] M. N. M. van Lieshout. Likelihood based inference for partially observed renewal processes. Statistics & Probability Letters, 118(??):190–196, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301183.

Vu:2019:SNE

[VL19] Vincent Q. Vu and Jing Lei. Squared-norm empirical processes. Statistics & Probability Letters, 150(??):108–113, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300604.

vanLeeuwaarden:2011:UAE

[vLT11] J. S. H. van Leeuwaarden and N. M. Temme. A uniform asymptotic expansion for weighted sums of exponentials. Statistics & Probability Letters, 81(11):1571–1579, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100188X.

Vrbik:2012:ACE

[VM12] I. Vrbik and P. D. McNicholas. Analytic calculations for the EM algorithm for multivariate skew-t mixture models. Statistics & Probability Letters, 82(6):1169–1174, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000673.

Veronese:2018:SAR

[VM18] Piero Veronese and Eugenio Melilli. Some asymptotic results for fiducial and confidence distributions. Statistics & Probability Letters, 134(??):98–105, March 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303255.

Volodko:2014:SDD

[Vol14] Nadezhda V. Volodko. Small deviations of the determinants of random matrices with Gaussian entries. Statistics & Probability Letters, 84(??):48–53, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003179.

Valjarevic:2012:SCO

[VP12] Dragana Valjarević and Ljiljana Petrović. Statistical causality and orthogonality of local martingales. Statistics & Probability Letters, 82(7):1326–1330, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001290.

Valero:2013:PSS

[VPCG13] Jordi Valero, Marta Pérez-Casany, and Josep Ginebra. On Poisson-stopped-sums that are mixed Poisson. Statistics & Probability Letters, 83(8):1830–1834, August 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001302.

Ventura:2012:IPE

[VR12] Laura Ventura and Walter Racugno. On interval and point estimators based on a penalization of the modified profile likelihood. Statistics & Probability Letters, 82(7):1285–1289, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001186.

Vrins:2016:CFT

[Vri16] Frédéric Vrins. Characteristic function of time-inhomogeneous Lévy-driven Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 116(??):55–61, September 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215303849.

Ventura:2013:NAB

[VRR13] Laura Ventura, Erlis Ruli, and Walter Racugno. A note on approximate Bayesian credible sets based on modified loglikelihood ratios. Statistics & Probability Letters, 83(11):2467–2472, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002563.

VanLieshout:2010:NPL

[VS10] M. N. M. Van Lieshout and R. S. Stoica. A note on pooling of labels in random fields. Statistics & Probability Letters, 80(17–18):1431–1436, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521000146X.

Vasudeva:2013:SLD

[VS13] R. Vasudeva and G. Srilakshminarayana. On strong large deviation results for lightly trimmed sums and some applications. Statistics & Probability Letters, 83(7):1745–1753, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001077.

vanStaden:2015:QBS

[vSK15] Paul J. van Staden and Robert A. R. King. The quantilebased skew logistic distribution. Statistics & Probability Letters, 96(??):109–116, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003125.

Varagnolo:2013:VNP

[VSP13] Damiano Varagnolo, Luca Schenato, and Gianluigi Pillonetto. A variation of the Newton–Pepys problem and its connections to size-estimation problems. Statistics & Probability Letters, 83(5):1472–1478, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213000515.

Veillette:2010:UDE

[VT10] Mark Veillette and Murad S. Taqqu. Using differential equations to obtain joint moments of first-passage times of increasing Lévy processes. Statistics & Probability Letters, 80(7–8):697–705, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210000039.

VanderWeele:2011:EPU

[VT11] Tyler J. VanderWeele and Eric J. Tchetgen Tchetgen. Effect partitioning under interference in two-stage randomized vaccine trials. Statistics & Probability Letters, 81(7):861–869, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000654.

Volny:2014:QCL

[VW14] Dalibor Volný and Michael Woodroofe. Quenched Central Limit Theorems for sums of stationary processes. Statistics & Probability Letters, 85(??):161–167, February 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003313.

vanWieringen:2017:MSE

[vW17] Wessel N. van Wieringen. On the mean squared error of the ridge estimator of the covariance and precision matrix. Statistics & Probability Letters, 123(??):88–92, April 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302863.

vanWaaij:2017:FAS

[vWvZ17] Jan van Waaij and Harry van Zanten. Full adaptation to smoothness using randomly truncated series priors with Gaussian coefficients and inverse gamma scaling. Statistics & Probability Letters, 123(??):93–99, April 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302930.

Vakeroudis:2012:CLT

[VY12] S. Vakeroudis and M. Yor. A Central Limit Theorem for a sequence of Brownian motions in the unit sphere in R^{n}. Statistics & Probability Letters, 82(3):599–605, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003725.

Vexler:2018:ELR

[VZ18] Albert Vexler and Li Zou. Empirical likelihood ratio tests with power one. Statistics & Probability Letters, 140(??):160–166, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301895.

Wang:2017:DSD

[WA17] Yaping Wang and Mingyao Ai. Definitive screening designs with extreme numbers of level changes. Statistics & Probability Letters, 131(??):13–18, December 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302626.

Walk:2010:SCK

[Wal10] Harro Walk. Strong consistency of kernel estimates of regression function under dependence. Statistics & Probability Letters, 80(15–16):1147–1156, August 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000891.

Walker:2015:PPH

[Wal15] Stephen G. Walker. A probabilistic proof of the Hardy inequality. Statistics & Probability Letters, 103(??):6–7, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000966.

Walker:2016:BIE

[Wal16] Stephen G. Walker. Bayesian information in an experiment and the Fisher information distance. Statistics & Probability Letters, 112(??):5–9, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216000109.

Walker:2017:SIC

[Wal17] Stephen G. Walker. A self-improvement to the Cauchy–Schwarz inequality. Statistics & Probability Letters, 122(??):86–89, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302401.

Wang:2011:HIO

[Wan11a] Jian Wang. Harnack inequalities for Ornstein–Uhlenbeck processes driven by Lévy processes. Statistics & Probability Letters, 81(9):1436–1444, September 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001593.

Wang:2011:SCS

[Wan11b] Jiantian Wang. A simple characterization of Student’s distributions and normal distributions. Statistics & Probability Letters, 81(8):903–906, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211001283.

Wan:2012:ELC

[Wan12a] Shuwen Wan. An empirical likelihood confidence interval for the volume under ROC surface. Statistics & Probability Letters, 82(7):1463–1467, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001502.

Wang:2012:NPA

[Wan12b] Antai Wang. On the nonidentifiability property of Archimedean copula models under dependent censoring. Statistics & Probability Letters, 82(3):621–625, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003580.

Wang:2012:IOC

[Wan12c] Weizhen Wang. An inductive order construction for the difference of two dependent proportions. Statistics & Probability Letters, 82(8):1623–1628, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212001289.

Wang:2013:NBC

[Wan13a] Weizhen Wang. A note on bootstrap confidence intervals for proportions. Statistics & Probability Letters, 83(12):2699–2702, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002940.

Wang:2013:MDE

[Wan13b] Yanqing Wang. Moderate deviations for the energy of charged polymer. Statistics & Probability Letters, 83(4):1078–1082, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004725.

Wang:2013:EPS

[Wan13c] Yun Wang. On efficiency properties of an R-square coefficient based on final prediction error. Statistics & Probability Letters, 83(10):2276–2281, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002332.

Wang:2014:APL

[Wan14a] Cheng Wang. Asymptotic power of likelihood ratio tests for high dimensional data. Statistics & Probability Letters, 88(??): 184–189, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000662.

Wang:2014:SPS

[Wan14b] Jian Wang. Smooth properties for semigroups of Lévy processes and their application. Statistics & Probability Letters, 89(??): 23–30, June 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214000741.

Wang:2014:LIL

[Wan14c] Wensheng Wang. Laws of the iterated logarithm of Chovertype for operator stable Lévy processes. Statistics & Probability Letters, 92(??):17–25, September 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001655.

Wang:2014:CMP

[Wan14d] Yizao Wang. Convergence to the maximum process of a fractional Brownian motion with shot noise. Statistics & Probability Letters, 90(??):33–41, July 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001072.

Wang:2015:SCR

[Wan15] Jiantian Wang. A stochastic comparison result about hazard rate ordering of two parallel systems comprising of geometric components. Statistics & Probability Letters, 106(??):86–90, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002394.

Wang:2016:LJI

[Wan16] Yizao Wang. Large jumps of q-Ornstein–Uhlenbeck processes. Statistics & Probability Letters, 118(??):110–116, November 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301067.

Wang:2017:PRL

[Wan17a] Jinxia Wang. Probabilistic representation and local existence for the quasi-linear partial integro-differential equations with Sobolev initial value. Statistics & Probability Letters, 120(??):69–80, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301729.

Wang:2017:EIE

[Wan17b] Weizhen Wang. On exact interval estimation for the odds ratio in subject-specific table. Statistics & Probability Letters, 129(??):360–366, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302201.

Wang:2018:NWC

[Wan18] Jin Wang. A note on weak convergence of general halfspace depth trimmed means. Statistics & Probability Letters, 142(??):50–56, November 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302487.

Wang:2019:ADT

[Wan19]  Wensheng Wang.  Asymptotics for discrete time hedg-

ing errors under fractional Black–Scholes models. Statistics & Probability Letters, 149(??):160–170, June 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300598.

Wu:2014:CLT

[WC14] Xinxing Wu and Guanrong Chen. Central limit theorem and chaoticity. Statistics & Probability Letters, 92(??):137– 142, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521400193X.

Wang:2015:GCL

[WC15] Qing Wang and Shiwen Chen. A general class of linearly extrapolated variance estimators. Statistics & Probability Letters, 98(??):29–38, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214004143.

Wang:2017:AOP

[WC17] Jiantian Wang and Bin Cheng. Answer to an open problem on mean residual life ordering of parallel systems under multiple-outlier exponential models. Statistics & Probability Letters, 130(??):80–84, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302468.

Wu:2018:AVR

[WC18] Chi-Hao Wu and Ting-Li Chen. On the asymptotic variance of reversible Markov chain without cycles. Statistics & Probability Letters, 137(??):224–228, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300312.

Wu:2019:EEA

[WCD19] Hongping Wu, Xiaomin Cao, and Caifeng Du. Estimating equations of additive mean residual life model with censored length-biased data. Statistics & Probability Letters, 155(??): Article 108552, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301890.

Wei:2011:TCN

[WCM11] Jiawei Wei, Raymond J. Carroll, and Arnab Maity. Testing for constant nonparametric effects in general semiparametric regression models with interactions. Statistics & Probability Letters, 81(7):717–723, July 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521000310X.

Wang:2018:FBM

[WCZ18] Xiaodi Wang, Xueping Chen, and Yingshan Zhang. Feasible blocked multi-factor designs of unequal block sizes. Statistics & Probability Letters, 135(??):102–109, April 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303759.

Wei:2012:TSD

[WD12] Fan Wei and Richard M. Dudley. Two-sample Dvoretzky– Kiefer–Wolfowitz inequalities. Statistics & Probability Letters, 82(3):636–644, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003658.

Weiss:2010:IPH

[Wei10] Christian H. Weiß. INARCH(1) processes: Higher-order moments and jumps. Statistics & Probability Letters, 80(23–24): 1771–1780, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715210002245.

Weissman:2017:SSU

[Wei17] Ishay Weissman. Sum of squares of uniform random variables. Statistics & Probability Letters, 129(??):147–154, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302018. See comment [For18].

Wei:2019:NSR

[Wei19] Qingda Wei. Nonzero-sum risk-sensitive finite-horizon continuoustime stochastic games. Statistics & Probability Letters, 147(??): 96–104, April 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303924.

Westgate:2013:BCC

[Wes13] Philip M. Westgate. A bias-corrected covariance estimator for improved inference when using an unstructured correlation with quadratic inference functions. Statistics & Probability Letters, 83(6):1553–1558, June 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000710.

Wesolowski:2015:MYT

[Wes15] Jacek Wesolowski. On the Matsumoto–Yor type regression characterization of the gamma and Kummer distributions. Statistics & Probability Letters, 107:145–149, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215002771.

Wang:2015:CML

[WF15] HaiYing Wang and Nancy Flournoy. On the consistency of the maximum likelihood estimator for the three parameter lognormal distribution. Statistics & Probability Letters, 105(??):57– 64, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001856.

Wang:2011:CLT

[WGL11] Hesong Wang, Zhiqiang Gao, and Quansheng Liu. Central Limit Theorems for a supercritical branching process in a random environment. Statistics & Probability Letters, 81(5):539–547, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000046.

White:2010:FSA

[WH10] W. T. J. White and M. D. Hendy. A fast and simple algorithm for finding the modes of a multinomial distribution. Statistics & Probability Letters, 80(1):63–68, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900354X.

Wu:2011:KTT

[WH11] Fuke Wu and Shigeng Hu. Khasminskii-type theorems for stochastic functional differential equations with infinite delay. Statistics & Probability Letters, 81(11):1690–1694, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211001763.

Wei:2014:CCR

[WH14] Linxiao Wei and Yijun Hu. Coherent and convex risk measures for portfolios with applications. Statistics & Probability Letters, 90(??):114–120, July 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214000984.

Wang:2010:DFT

[WHB10] Haiyan Wang, James Higgins, and Dale Blasi. Distributionfree tests for no effect of treatment in heteroscedastic functional data under both weak and long range dependence. Statistics & Probability Letters, 80(5–6):390–402, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004404.

Whitt:2012:FBD

[Whi12] Ward Whitt. Fitting birth-and-death queueing models to data. Statistics & Probability Letters, 82(5):998–1004, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000570.

Wang:2011:MID

[WHRY11] Xuejun Wang, Shuhe Hu, B. L. S. Prakasa Rao, and Wenzhi Yang. Maximal inequalities for N-demimartingale and Strong Law of Large Numbers. Statistics & Probability Letters, 81(9):1348–1353, September 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001581.

Wang:2010:EII

[WHYL10] Xuejun Wang, Shuhe Hu, Wenzhi Yang, and Nengxiang Ling. Exponential inequalities and inverse moment for NOD sequence.

Statistics & Probability Letters, 80(5–6):452–461, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004477.

Wishart:2011:MLB

[Wis11] Justin Rory Wishart. Minimax lower bound for kink location estimators in a nonparametric regression model with long-range dependence. Statistics & Probability Letters, 81(12):1871–1875, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002501.

Wit:2018:BDB

[Wit18] Ernst C. Wit. Big data and biostatistics: the death of the asymptotic Valhalla. Statistics & Probability Letters, 136(??): 30–33, ???? 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218300841.

Wang:2011:BEB

[WJY11] Baobin Wang, Hui Jiang, and Jinyou Yu. Berry–Esseen bound for parameter estimation in some time inhomogeneous diffusions and applications. Statistics & Probability Letters, 81(8):921– 929, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211001337.

Wu:2010:DAM

[WK10a] Haizhen Wu and Giorgi Kvizhinadze. Diversity analysis in multiple-choice questionnaires. Statistics & Probability Letters, 80(13–14):1103–1110, July 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521000074X.

Wu:2010:NPE

[WK10b] Xiaowei Wu and Marek Kimmel. A note on the path to extinction of critical Markov branching processes. Statistics & Probability Letters, 80(5–6):263–269, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004088.

Williams:2011:LRT

[WK11a] M. R. Williams and D. Kim. Likelihood ratio tests for continuous monotone hazards with an unknown change point. Statistics & Probability Letters, 81(11):1599–1603, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521100215X.

Wu:2011:MLT

[WK11b] Haizhen Wu and Giorgi Kvizhinadze. Martingale limit theorems of divisible statistics in a multinomial scheme with mixed frequencies. Statistics & Probability Letters, 81(8):1128–1135, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000903.

Wang:2011:BAA

[WL11] Lianming Wang and Xiaoyan Lin. A Bayesian approach for analyzing case 2 interval-censored data under the semiparametric proportional odds model. Statistics & Probability Letters, 81(7): 876–883, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000800.

Wang:2013:URD

[WL13] Juan Wang and Junping Li. Uniqueness, recurrence and decay properties of collision branching processes with immigration. Statistics & Probability Letters, 83(7):1603–1612, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000989.

Wang:2014:NEE

[WL14a] Kangning Wang and Lu Lin. New efficient estimation and variable selection in models with single-index structure. Statistics & Probability Letters, 89(??):58–64, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000820.

Woody:2014:LRM

[WL14b] Jonathan Woody and Robert Lund. A linear regression model with persistent level shifts: An alternative to infill asymptotics. Statistics & Probability Letters, 95(??):118–124, December 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002995.

Wang:2015:NAP

[WL15a] Jiantian Wang and Henry Laniado. A note on allocation policy in two-parallel-series and two-series-parallel systems with respect to likelihood ratio order. Statistics & Probability Letters, 102(??): 17–21, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500084X.

Wang:2015:MPS

[WL15b] Min Wang and Tao Lu. A matching prior for the shape parameter of the exponential power distribution. Statistics & Probability Letters, 97(??):150–154, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003927.

Weng:2017:SOE

[WL17] Zhichao Weng and Xin Liao. Second order expansions of distributions of maxima of bivariate Gaussian triangular arrays under power normalization. Statistics & Probability Letters, 125(??): 33–43, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300469.

Wang:2017:CSC

[WLLL17] Yuejiao Wang, Zaiming Liu, Yingqiu Li, and Quansheng Liu. On the concept of subcriticality and criticality and a ratio theorem for a branching process in a random environment. Statistics & Probability Letters, 127(??):97–103, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300792.

Wang:2011:BFP

[WLLZ11] Linglu Wang, Qizhai Li, Zhaohai Li, and Gang Zheng. Bayes factors in the presence of population stratification. Statistics & Probability Letters, 81(7):836–841, July 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000666.

Wang:2017:SOE

[WLP17] Rui Wang, Xin Liao, and Zuoxiang Peng. Second-order expansions for maxima of dynamic bivariate normal copulas. Statistics & Probability Letters, 129(??):275–283, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302158.

Wang:2011:EET

[WLS11] Lichun Wang, Heng Lian, and Radhey S. Singh. On efficient estimators of two seemingly unrelated regressions. Statistics & Probability Letters, 81(5):563–570, May 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000149.

Won:2014:MFD

[WLY^{+}14] Joong-Ho Won, Johan Lim, Donghyeon Yu, Byung Soo Kim, and Kyunga Kim. Monotone false discovery rate. Statistics & Probability Letters, 87(??):86–93, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213004100.

Wang:2018:TSE

[WM18] Wenyuan Wang and Ruixing Ming. Two-side exit problems for taxed Lévy risk process involving the general draw-down time. Statistics & Probability Letters, 138(??):66–74, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830110X.

Wang:2015:LLQ

[WMFW15] Jiang-Feng Wang, Wei-Min Ma, Guo-Liang Fan, and Li-Min Wen. Local linear quantile regression with truncated and dependent data. Statistics & Probability Letters, 96(??):232–240, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214003496.

Wang:2011:EDP

[WMH11] Wenyuan Wang, Ruixing Ming, and Yijun Hu. On the expected discounted penalty function for risk process with tax. Statistics & Probability Letters, 81(4):489–501, April 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003524.

Wang:2015:SCT

[WMM15] Zhen Wang, Jianyong Mu, and Yu Miao. Some convergence theorems for RM algorithm. Statistics & Probability Letters, 99(??): 54–60, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214004271.

Wang:2013:ANL

[WMZW13] Jiang-Feng Wang, Wei-Min Ma, Hui-Zeng Zhang, and Li-Min Wen. Asymptotic normality for a local composite quantile regression estimator of regression function with truncated data. Statistics & Probability Letters, 83(6):1571–1579, June 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000722.

Wang:2010:DEO

[WN10a] Ke Wang and H. N. Nagaraja. Distribution of extremal order statistics from large subsets of concomitants. Statistics & Probability Letters, 80(7–8):534–539, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771520900457X.

Withers:2010:DQF

[WN10b] Christopher S. Withers and Saralees Nadarajah. The distribution and quantiles of functionals of weighted empirical distributions when observations have different distributions. Statistics & Probability Letters, 80(13–14):1093–1102, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000738.

Withers:2010:EEP

[WN10c] Christopher S. Withers and Saralees Nadarajah. Edgeworth expansions for the product of two complex random matrices each with IID components. Statistics & Probability Letters, 80(23–24):1954–1961, December 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210002488.

Withers:2011:ELB

[WN11] Christopher S. Withers and Saralees Nadarajah. Estimates of low bias for the multivariate normal. Statistics & Probability Letters, 81(11):1635–1647, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002161.

Withers:2013:ABM

[WN13a] Christopher S. Withers and Saralees Nadarajah. Asymptotic behavior of the maximum from distributions subject to trends in location and scale. Statistics & Probability Letters, 83(10):2143– 2151, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521300206X.

Withers:2013:BE

[WN13b] Christopher S. Withers and Saralees Nadarajah. Bayesian efficiency. Statistics & Probability Letters, 83(4):1203–1212, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000205.

Withers:2014:CMN

[WN14a] Christopher S. Withers and Saralees Nadarajah. Cumulants of multinomial and negative multinomial distributions. Statistics & Probability Letters, 87(??):18–26, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213004136.

Withers:2014:DMM

[WN14b] Christopher S. Withers and Saralees Nadarajah. The distribution of the maximum of the multivariate AR(p) and multivariate MA(p) processes. Statistics & Probability Letters, 95(??): 48–56, ???? 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002880.

Withers:2014:DMC

[WN14c] Christopher S. Withers and Saralees Nadarajah. The dual multivariate Charlier and Edgeworth expansions. Statistics & Probability Letters, 87(??):76–85, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000145.

Withers:2015:JDM

[WN15] Christopher S. Withers and Saralees Nadarajah. The joint distribution of the maximum and minimum of an AR(1) process. Statistics & Probability Letters, 99(??):77–84, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000139.

Withers:2019:DMP

[WN19a] Christopher S. Withers and Saralees Nadarajah. The distribution of the minimum of a positive sample. Statistics & Probability Letters, 151(??):89–96, August 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301051.

Withers:2019:MCE

[WN19b] Christopher S. Withers and Saralees Nadarajah. Moments and cumulants of the extremes of a sample from a uniform distribution. Statistics & Probability Letters, 145(??):238–247, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303225.

Wojcik:2013:HFI

[Wój13a] Michal Ryszard Wójcik. How fast increasing powers of a continuous random variable converge to Benford’s law. Statistics & Probability Letters, 83(12):2688–2692, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002927.

Wojewodka:2013:ERC

[Woj13b] H. Wojewódka. Exponential rate of convergence for some Markov operators. Statistics & Probability Letters, 83(10):2337–2347, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002034.

Wong:2013:NBF

[Won13] A. C. M. Wong. A note on Bayesian and frequentist parametric inference for a scalar parameter of interest. Statistics & Probability Letters, 83(1):414–421, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003781.

Wong:2018:BDB

[Won18] Limsoon Wong. Big data and a bewildered lay analyst. Statistics & Probability Letters, 136(??):73–77, ???? 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300786.

Woody:2015:TSR

[Woo15] Jonathan Woody. Time series regression with persistent level shifts. Statistics & Probability Letters, 102(??):22–29, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001005.

Wang:2014:NMQ

[WP14] Hui Wang and Jiazhu Pan. Normal mixture quasi maximum likelihood estimation for non-stationary TGARCH(1,1) models. Statistics & Probability Letters, 91(??):117–123, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001205.

Wu:2015:MBT

[WQD15] Jianhong Wu, Jinxu Qin, and Qing Ding. A moment-based test for individual effects in the error component model with incomplete panels. Statistics & Probability Letters, 104(??):153–162, September 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001649.

Wu:2012:ADB

[WQY12] Lan Wu, Yongcheng Qi, and Jingping Yang. Asymptotics for dependent Bernoulli random variables. Statistics & Probability Letters, 82(3):455–463, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003841.

Wu:2010:BPP

[WR10] Changbao Wu and J. N. K. Rao. Bootstrap procedures for the pseudo empirical likelihood method in sample surveys. Statistics & Probability Letters, 80(19–20):1472–1478, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210001586.

Wang:2018:APG

[WR18] Xin Wang and Vivekananda Roy. Analysis of the Pólya–Gamma block Gibbs sampler for Bayesian logistic linear mixed models. Statistics & Probability Letters, 137(??):251–256, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300476.

Wright:2017:EOS

[Wri17] Tommy Wright. Exact optimal sample allocation: More efficient than Neyman. Statistics & Probability Letters, 129(??):50– 57, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301657.

Wang:2011:FQT

[WRvdL11] Hui Wang, Sherri Rose, and Mark J. van der Laan. Finding quantitative trait loci genes with collaborative targeted maximum likelihood learning. Statistics & Probability Letters, 81(7): 792–796, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210003093.

Wang:2010:ASM

[WS10a] Yizao Wang and Stilian A. Stoev. On the association of sum- and max-stable processes. Statistics & Probability Letters, 80(5–6):480–488, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004507.

Wu:2010:EML

[WS10b] Jianhong Wu and Weihua Su. Estimation of moments for linear panel data models with potential existence of time effects. Statistics & Probability Letters, 80(23–24):1933–1939, December 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210002452.

Wang:2012:OBA

[WS12] Haiying Wang and Dongchu Sun. Objective Bayesian analysis for a truncated model. Statistics & Probability Letters, 82(12):2125–2135, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002854.

Wang:2013:ISV

[WS13] Mingqiu Wang and Lixin Song. Identification for semiparametric varying coefficient partially linear models. Statistics & Probability Letters, 83(5):1311–1320, May 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000461.

Wu:2015:HTE

[WS15] Jianghong Wu and Weixing Song. On Hong–Tamer’s estimator in nonlinear errors-in-variable regression models. Statistics & Probability Letters, 97(??):165–175, February 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003976.

Weiss:2016:BCM

[WS16] Christian H. Weiß and Sebastian Schweer. Bias corrections for moment estimators in Poisson INAR(1) and INARCH(1) processes. Statistics & Probability Letters, 112(??):124–130, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521530345X.

Wen:2017:ABS

[WS17] Jiaqiang Wen and Yufeng Shi. Anticipative backward stochastic differential equations driven by fractional Brownian motion. Statistics & Probability Letters, 122(??):118–127, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302516.

Weiss:2018:MSS

[WSMG18] Christian H. Weiß, Manuel G. Scotto, Tobias A. Möller, and Sónia Gouveia. Mean-square stability of delayed stochastic neural networks with impulsive effects driven by G-Brownian motion. Statistics & Probability Letters, 143(??):56–66, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302682.

Wang:2014:MAL

[WST14] Lu Wang, Jincheng Shen, and Peter F. Thall. A modified adaptive Lasso for identifying interactions in the Cox model with the heredity constraint. Statistics & Probability Letters, 93(??):126–133, October 1–142, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002351.

Wolf:2011:URD

[WSU11] Petra Wolf, Georg Schmidt, and Kurt Ulm. The use of ROC for defining the validity of the prognostic index in censored data. Statistics & Probability Letters, 81(7):783–791, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000678.

Wang:2010:BEG

[WSW10] Mingqiu Wang, Lixin Song, and Xiaoguang Wang. Bridge estimation for generalized linear models with a diverging number of parameters. Statistics & Probability Letters, 80(21–22):1584–1596, November 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001732.

Wu:2014:TSE

[WT14] Fan Wu and Min Tsao. Two-sample extended empirical likelihood. Statistics & Probability Letters, 84(??):81–87, January 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003143.

Wei:2018:HCE

[WT18] Fengrong Wei and Weizhong Tian. Heterogeneous connection effects. Statistics & Probability Letters, 133(??):9–14, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303097.

Wu:2011:SPS

[Wu11a] Dongsheng Wu. On the solution process for a stochastic fractional partial differential equation driven by space-time white noise. Statistics & Probability Letters, 81(8):1161–1172, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000952.

Wu:2011:ASL

[Wu11b] Qunying Wu. Almost sure limit theorems for stable distributions. Statistics & Probability Letters, 81(6):662–672, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000411.

Wu:2014:LAD

[Wu14] Rongning Wu. Least absolute deviation estimation for general fractionally integrated autoregressive moving average time series models. Statistics & Probability Letters, 94(??):69–76, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002466.

Wang:2010:PDD

[WW10] Xue Wang and Stephen G. Walker. A penalised data-driven block shrinkage approach to empirical Bayes wavelet estimation. Statistics & Probability Letters, 80(11–12):990–996, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521000060X.

Wang:2011:SOF

[WW11] Bing Wang and Min Wang. Stochastic ordering of folded normal random variables. Statistics & Probability Letters, 81(4):524– 528, April 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210003676.

Wang:2013:SEP

[WW13a] Qin Wang and Rongning Wu. Shrinkage estimation of partially linear single-index models. Statistics & Probability Letters, 83(10):2324–2331, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213002241.

Wang:2013:SIC

[WW13b] Xinghui Wang and Xuejun Wang. Some inequalities for conditional demimartingales and conditional N-demimartingales. Statistics & Probability Letters, 83(3):700–709, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004336.

Wang:2014:ALE

[WW14] Mingqiu Wang and Xiuli Wang. Adaptive Lasso estimators for ultrahigh dimensional generalized linear models. Statistics & Probability Letters, 89(??):41–50, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000789.

Worms:2016:LBI

[WW16] J. Worms and R. Worms. A Lynden–Bell integral estimator for extremes of randomly truncated data. Statistics & Probability Letters, 109:106–117, February 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521500379X.

Wu:2018:CPI

[WW18] Bo Wu and Jiang-Lun Wu. Characterising the path-independent property of the Girsanov density for degenerated stochastic differential equations. Statistics & Probability Letters, 133(??):71– 79, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521730319X.

Witkovsky:2015:LLR

[WWD15] Viktor Witkovský, Gejza Wimmer, and Tomy Duby. Logarithmic Lambert W × F random variables for the family of chi-squared distributions and their applications. Statistics & Probability Letters, 96(??):223–231, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003484.

Wu:2017:CMC

[WWH17] Yi Wu, Xuejun Wang, and Shuhe Hu. Complete moment convergence for weighted sums of weakly dependent random variables and its application in nonparametric regression model. Statistics & Probability Letters, 127(??):56–66, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521730127X.

Wang:2018:NJO

[WWPY18] Wenyuan Wang, Xueyuan Wu, Xingchun Peng, and Kam C. Yuen. A note on joint occupation times of spectrally negative Lévy risk processes with tax. Statistics & Probability Letters, 140(??):13–22, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301706.

Wu:2012:ABS

[WWR12] Hao Wu, Wenyuan Wang, and Jie Ren. Anticipated backward stochastic differential equations with non-Lipschitz coefficients. Statistics & Probability Letters, 82(3):672–682, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003907.

Wang:2019:CLT

[WWS19] Sumin Wang, Dongying Wang, and Fasheng Sun. A central limit theorem for marginally coupled designs. Statistics & Probability Letters, 146(??):168–174, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303717.

Wang:2010:STW

[WWW10] Ran Wang, Xinyu Wang, and Liming Wu. Sanov’s theorem in the Wasserstein distance: a necessary and sufficient condition. Statistics & Probability Letters, 80(5–6):505–512, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004532.

Wang:2014:LTB

[WWW14] Guanying Wang, Xingchun Wang, and Yongjin Wang. Long time behavior for nonlocal stochastic Kuramoto–Sivashinsky equations. Statistics & Probability Letters, 87(??):54–60, April 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214000030.

Wang:2019:CMS

[WWW19] Bin Wang, Ruodu Wang, and Yuming Wang. Compatible matrices of Spearman’s rank correlation. Statistics & Probability Letters, 151(??):67–72, August 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300914.

Wang:2017:LTS

[WWX17] Guanying Wang, Xingchun Wang, and Guangli Xu. Long time stability of nonlocal stochastic Kuramoto–Sivashinsky equations with jump noises. Statistics & Probability Letters, 127(??):23– 32, August 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217301244.

Wei:2012:OSS

[WWY12] Jiaqin Wei, Rongming Wang, and Hailiang Yang. Optimal surrender strategies for equity-indexed annuity investors with partial information. Statistics & Probability Letters, 82(7):1251– 1258, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001149.

Wang:2018:LTB

[WWZ18] Guanying Wang, Xingchun Wang, and Ke Zhou. Long time behavior for stochastic Burgers equations with jump noises. Statistics & Probability Letters, 141(??):41–49, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302104.

Wang:2012:SPC

[WX12] Li Wang and Xingzhong Xu. Step-up procedure controlling generalized family-wise error rate. Statistics & Probability Letters, 82(4):775–782, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200003X.

Wang:2018:RCP

[WX18] Shasha Wang and Wen-Qing Xu. Random cyclic polygons from Dirichlet distributions and approximations of π. Statistics & Probability Letters, 140(??):84–90, September 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301883.

Wang:2019:CRM

[WX19a] Wensheng Wang and Yimin Xiao. The Csörgő–Révész moduli of non-differentiability of fractional Brownian motion. Statistics & Probability Letters, 150(??):81–87, July 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300689.

Wu:2019:PIS

[WX19b] Zhen Wu and Ruimin Xu. Probabilistic interpretation for Sobolev solutions of McKean–Vlasov partial differential equations. Statistics & Probability Letters, 145(??):273–283, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303171.

Wang:2019:CNC

[WXL19] Zhaoliang Wang, Liugen Xue, and Juanfang Liu. Checking nonparametric component for partially nonlinear model with missing response. Statistics & Probability Letters, 149(??):1–8, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300264.

Wang:2010:ARP

[WY10] Yinfeng Wang and Chuancun Yin. Approximation for the ruin probabilities in a discrete time risk model with dependent risks. Statistics & Probability Letters, 80(17–18):1335–1342, September 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210001239.

Wang:2016:PPS

[WY16] Min Wang and Mingan Yang. Posterior property of Studentt linear regression model using objective priors. Statistics & Probability Letters, 113(??):23–29, June 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215302893.

Wong:2018:PPF

[WY18] Tak Kwong Wong and Sheung Chi Phillip Yam. A probabilistic proof for Fourier inversion formula. Statistics & Probability Letters, 141(??):135–142, October 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302153.

Wang:2015:IEP

[WYC15] Bing Xing Wang, Keming Yu, and Frank P. A. Coolen. Interval estimation for proportional reversed hazard family based on lower record values. Statistics & Probability Letters, 98(??):115– 122, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214004222.

Wang:2014:NIN

[WYH14] Shaoli Wang, Weixin Yao, and Mian Huang. A note on the identifiability of nonparametric and semiparametric mixtures of GLMs. Statistics & Probability Letters, 93(??):41–45, October 1– 142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002120.

Wang:2010:CRM

[WYL10] Chunwei Wang, Chuancun Yin, and Erqiang Li. On the classical risk model with credit and debit interests under absolute ruin. Statistics & Probability Letters, 80(5–6):427–436, March 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004441.

Wysocki:2012:CAC

[Wys12] Włodzimierz Wysocki. Constructing Archimedean copulas from diagonal sections. Statistics & Probability Letters, 82(4):818– 826, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000181.

Wysocki:2013:WCA

[Wys13] Włodzimierz Wysocki. When a copula is archimax. Statistics & Probability Letters, 83(1):37–45, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003409.

Wang:2013:EOR

[WYY13] Kaiyong Wang, Yang Yang, and Changjun Yu. Estimates for the overshoot of a random walk with negative drift and non-convolution equivalent increments. Statistics & Probability Letters, 83(6):1504–1512, June 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000588.

Wang:2014:WCF

[WYY14] Zhi Wang, Litan Yan, and Xianye Yu. Weak convergence to the fractional Brownian sheet using martingale differences. Statistics & Probability Letters, 92(??):72–78, September 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521400145X.

Wei:2010:MSU

[WZ10] Jiawei Wei and Lan Zhou. Model selection using modified AIC and BIC in joint modeling of paired functional data. Statistics & Probability Letters, 80(23–24):1918–1924, December 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210002439.

Wang:2013:FDE

[WZ13a] Ling-Di Wang and Yu-Hui Zhang. The first Dirichlet eigenvalue of birth-death process on trees. Statistics & Probability Letters, 83(9):1973–1982, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001521.

Wu:2013:PSM

[WZ13b] Jing Wu and Hua Zhang. Penalization schemes for multi-valued stochastic differential equations. Statistics & Probability Letters, 83(2):481–492, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212003914.

Wan:2015:UPO

[WZ15a] Shuwen Wan and Biao Zhang. Using proportional odds models for semiparametric ROC surface estimation. Statistics & Probability Letters, 105(??):74–79, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001959.

Wang:2015:EDR

[WZ15b] Ruodu Wang and Johanna F. Ziegel. Elicitable distortion risk measures: A concise proof. Statistics & Probability Letters, 100(??):172–175, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000541.

West:2016:RDM

[WZ16] Kenneth D. West and Zifeng Zhao. Regressor and disturbance have moments of all orders, least squares estimator has none. Statistics & Probability Letters, 115(??):54–59, August 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300207.

Wilk:2017:ODR

[WZ17] M. Wilk and A. Zaigraev. DS-optimal designs for random coefficient first-degree regression model with heteroscedastic errors. Statistics & Probability Letters, 128(??):28–34, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301554.

Wang:2019:PUN

[WZ19] Chen Wang and Tusheng Zhang. Pathwise uniqueness and non-explosion of SDEs driven by compensated Poisson random measures. Statistics & Probability Letters, 150(??):61–67, July 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300628.

Wang:2014:MWS

[WZG14] Xiao Wang, Bo Zhao, and Joseph Glaz. A multiple window scan statistic for time series models. Statistics & Probability Letters, 94(??):196–203, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002636.

Wang:2016:JEL

[WZG16] Dongliang Wang, Yichuan Zhao, and Dirk W. Gilmore. Jackknife empirical likelihood confidence interval for the Gini index. Statistics & Probability Letters, 110(??):289–295, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003430.

Wang:2017:LMC

[WZS17] Antai Wang, Yilong Zhang, and Yongzhao Shao. On the likelihood of mixture cure models. Statistics & Probability Letters, 131(??):51–55, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302663.

Wang:2012:LEJ

[WZT12] Yunyan Wang, Lixin Zhang, and Mingtian Tang. Local M-estimation for jump-diffusion processes. Statistics & Probability Letters, 82(7):1273–1284, July 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001174.

Wang:2014:FCD

[WZT14] Xiaodi Wang, Yingshan Zhang, and Yincai Tang. Feasible criterion for designs based on fixed effect ANOVA model. Statistics & Probability Letters, 87(??):134–142, April 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000315.

Wang:2017:RPE

[WZW17] Xiuli Wang, Shengli Zhao, and Mingqiu Wang. Restricted profile estimation for partially linear models with large-dimensional covariates. Statistics & Probability Letters, 128(??):71–76, September 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301529.

Wang:2017:CNO

[WZYL17] Xiao-Lei Wang, Yu-Na Zhao, Jian-Feng Yang, and Min-Qian Liu. Construction of (nearly) orthogonal sliced Latin hypercube designs. Statistics & Probability Letters, 125(??):174–180, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300536.

Xiong:2015:RPE

[XD15] Wenjun Xiong and Juan Ding. Robust procedures for experimental design in group testing considering misclassification. Statistics & Probability Letters, 100(??):35–41, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000279.

Xu:2018:GCN

[XDWY18] Jin Xu, Xiaojun Duan, Zhengming Wang, and Liang Yan. A general construction for nested Latin hypercube designs. Statistics & Probability Letters, 134(??):134–140, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217303449.

Xiao:2017:RTG

[XF17] Lishun Xiao and Shengjun Fan. A representation theorem for generators of BSDEs with general growth generators in y and its applications. Statistics & Probability Letters, 129(??):297– 305, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217302183.

Xu:2017:HTP

[XFC17] Hong-Xia Xu, Guo-Liang Fan, and Zhen-Long Chen. Hypothesis tests in partial linear errors-in-variables models with missing response. Statistics & Probability Letters, 126(??):219–229, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300949.

Xie:2012:RCT

[XH12] Tingfan Xie and Jianjun He. Rate of convergence in a theorem of Heyde. Statistics & Probability Letters, 82(8):1576–1582, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001277.

Xu:2015:FRE

[XH15] Kai Xu and Daojiang He. Further results on estimation of covariance matrix. Statistics & Probability Letters, 101(??):11– 20, June 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215000796.

Xiang:2017:SSD

[Xia17] Jim X. Xiang. Sample size determination of a nonparametric test based on weighted L_{2}-Wasserstein distance. Statistics & Probability Letters, 123(??):160–164, April 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302760.

Xia:2019:NAB

[Xia19a] Dong Xia. Non-asymptotic bounds for percentiles of independent non-identical random variables. Statistics & Probability Letters, 152(??):111–120, September 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521930121X.

Xia:2019:PME

[Xia19b] Wanwan Xia. Partial monotonicity of entropy revisited. Statistics & Probability Letters, 145(??):248–253, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218303146.

Xiang:2019:EMM

[Xia19c] Jim X. Xiang. Estimation of minimum and maximum correlation coefficients. Statistics & Probability Letters, 145(??):81– 88, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302852.

Xie:2013:LSD

[Xie13] Junshan Xie. Limiting spectral distribution of normalized sample covariance matrices with p/n → 0. Statistics & Probability Letters, 83(2):543–550, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003860.

Xing:2012:ASA

[Xin12] Fei Xing. Almost sure asymptotic for Ornstein–Uhlenbeck processes of Poisson potential. Statistics & Probability Letters, 82(12):2091–2102, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212002842.

Xia:2015:SCM

[XJW15] Tian Xia, Xuejun Jiang, and Xueren Wang. Strong consistency of the maximum quasi-likelihood estimator in quasilikelihood nonlinear models with stochastic regression. Statistics & Probability Letters, 103(??):37–45, August 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001261.

Xiao:2015:DFM

[XKBG15] Yuewen Xiao, Yu-Cheng Ku, Peter Bloomfield, and Sujit K. Ghosh. On the degrees of freedom in MCMCbased Wishart models for time series data. Statistics & Probability Letters, 98(??):59–64, March 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214004155.

Xing:2016:NLC

[XL16] Jiamin Xing and Yong Li. Nonlinear Lyapunov criteria for stochastic explosive solutions. Statistics & Probability Letters, 109:63–67, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003752.

Xu:2018:SDE

[XL18] Liping Xu and Jiaowan Luo. Stochastic differential equations driven by fractional Brownian motion. Statistics & Probability Letters, 142(??):102–108, November 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302359.

Xiong:2019:NRS

[XLL19] Xianzhu Xiong, Rui Li, and Heng Lian. On nonparametric randomized sketches for kernels with further smoothness. Statistics & Probability Letters, 153(??):139–142, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301531.

Xie:2017:GAC

[XLY17] Jiehua Xie, Feng Lin, and Jingping Yang. On a generalization of Archimedean copula family. Statistics & Probability Letters, 125(??):121–129, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300512.

Xu:2018:ILE

[XS18] Ancha Xu and Lijuan Shen. Improved on-line estimation for gamma process. Statistics & Probability Letters, 143(??):67–73, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302657.

Xifara:2014:LDM

[XSL^{+}14] T. Xifara, C. Sherlock, S. Livingstone, S. Byrne, and M. Girolami. Langevin diffusions and the Metropolis-adjusted Langevin algorithm. Statistics & Probability Letters, 91(??):14–19, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001333.

Xun:2017:EEQ

[XSZ17] Li Xun, Li Shao, and Yong Zhou. Efficiency of estimators for quantile differences with left truncated and right censored data. Statistics & Probability Letters, 121(??):29–36, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302085.

Xue:2011:FCD

[XT11] Jing-Hao Xue and D. Michael Titterington. The p-folded cumulative distribution function and the mean absolute deviation from the p-quantile. Statistics & Probability Letters, 81(8):1179– 1182, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000976.

Xu:2018:AMS

[XTT18] Zhengchun Xu, Zhongquan Tan, and Linjun Tang. Approximation of the maximum of storage process with fractional Brownian motion as input. Statistics & Probability Letters, 140(??):147–159, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218302025.

Xu:2012:RSG

[Xu12a] Xiaoming Xu. Reflected solutions of generalized anticipated BS-DEs and application to reflected BSDEs with functional barrier. Statistics & Probability Letters, 82(6):1185–1192, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212000648.

Xu:2012:RMB

[Xu12b] Yitian Xu. A rough margin-based linear support ν vector regression. Statistics & Probability Letters, 82(3):528–534, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003609.

Xu:2013:MCA

[Xu13] Fangjun Xu. Markov chain approximations to singular stablelike processes. Statistics & Probability Letters, 83(3):790–796, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004488.

Xu:2014:PET

[Xu14] Wei Xu. Parameter estimation in two-type continuousstate branching processes with immigration. Statistics & Probability Letters, 91(??):124–134, August 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001618.

Xu:2015:RCP

[Xu15] Wen-Qing Xu. Random circumscribing polygons and approximations of π. Statistics & Probability Letters, 106(??):52–57, November 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215002308.

Xu:2018:SIS

[Xu18] Lihu Xu. Singular integrals of stable subordinator. Statistics & Probability Letters, 139(??):115–118, August 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301482.

Xue:2017:MFL

[Xue17] Xiaofeng Xue. Mean field limit for survival probability of the high-dimensional contact process. Statistics & Probability Letters, 127(??):178–184, August 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301499. See priority acknowledge-

ment [Xue19].

Xue:2018:CIR

[Xue18] Xiaofeng Xue. The critical infection rate of the highdimensional two-stage contact process. Statistics & Probability Letters, 140(??):115–125, September 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301871.

Xue:2019:PRM

[Xue19] Xiaofeng Xue. Priority of the result in ‘Mean field limit for survival probability of the high-dimensional contact process’. Statistics & Probability Letters, 148(??):133, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219300239. See [Xue17].

Xia:2010:NPR

[XW10] Tian Xia and Yuebao Wang. A note on the properties of the reproductive dispersion model. Statistics & Probability Letters, 80(17–18):1397–1404, September 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001409.

Xu:2016:SMM

[XW16] Guangli Xu and Yongjin Wang. On stability of the Markovmodulated skew CIR process. Statistics & Probability Letters, 109:139–144, February 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003739.

Xu:2017:ASC

[XW17] Feng Xu and Qunying Wu. Almost sure central limit theorem for self-normalized partial sums of ρ^{−}-mixing sequences. Statistics & Probability Letters, 129(??):17–27, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301621.

Xu:2019:SFS

[XWML19] Jie Xu, Jiaping Wen, Jianyong Mu, and Jicheng Liu. Stochastic flows of SDEs with non-Lipschitz coefficients and singular time. Statistics & Probability Letters, 148(??):118–127, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521930032X.

Xiao:2019:LPP

[XWT19] Jinghong Xiao, Yuechun Wen, and Zhongquan Tan. The limit properties of point processes of upcrossings in nonstationary strongly dependent Gaussian models. Statistics & Probability Letters, 149(??):38–46, June 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300343.

Xie:2013:RRS

[XX13] Jian-Sheng Xie and Yuan-Yuan Xu. Range-renewal structure of transient simple random walk. Statistics & Probability Letters, 83(10):2220–2221, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001995.

Xing:2012:NTD

[XXY12] Xiaoyu Xing, Yongsheng Xing, and Xuewei Yang. A note on transition density for the reflected Ornstein–Uhlenbeck process. Statistics & Probability Letters, 82(3):586–591, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003737.

Xie:2017:LSL

[XXY17] Fang Xie, Lihu Xu, and Youcai Yang. Lasso for sparse linear regression with exponentially β-mixing errors. Statistics & Probability Letters, 125(??):64–70, June 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300457.

Xiao:2012:PAL

[XY12] Xiaoyong Xiao and Hongwei Yin. Precise asymptotics in the law of iterated logarithm for the first moment convergence of i.i.d. random variables. Statistics & Probability Letters, 82(8):1590– 1596, August 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001629.

Xi:2013:SFP

[XY13] Fubao Xi and George Yin. The strong Feller property of switching jump-diffusion processes. Statistics & Probability Letters, 83(3):761–767, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004373.

Xiong:2017:SEU

[XY17] Jie Xiong and Xu Yang. Strong existence and uniqueness to a class of nonlinear SPDEs driven by Gaussian colored noises. Statistics & Probability Letters, 129(??):113–119, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217301876.

Xu:2016:SDL

[XYZ16] Lihu Xu, Wen Yue, and Tusheng Zhang. Smooth densities of the laws of perturbed diffusion processes. Statistics & Probability Letters, 119(??):55–62, December 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301298.

Xu:2013:SBA

[XZ13] Dengke Xu and Zhongzhan Zhang. A semiparametric Bayesian approach to joint mean and variance models. Statistics & Probability Letters, 83(7):1624–1631, July 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000734.

Xu:2016:MPS

[XZ16] Siyan Xu and Mengqi Zheng. A maximum principle for the stochastic variational inequalities. Statistics & Probability Letters, 116(??):116–121, September 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215301449.

Xiao:2013:PRG

[XZY13] Xiao-Yong Xiao, Li-Xin Zhang, and Hong-Wei Yin. Precise rates in the generalized law of the iterated logarithm. Statistics & Probability Letters, 83(2):616–623, February 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200421X.

Xun:2019:GES

[XZZ19] Li Xun, Yangzhi Zhou, and Yong Zhou. A generalization of Expected Shortfall based capital allocation. Statistics & Probability Letters, 146(??):193–199, March 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303389.

Yabe:2017:ADC

[Yab17] Ryota Yabe. Asymptotic distribution of the conditional-sum-of-squares estimator under moderate deviation from a unit root in MA(1). Statistics & Probability Letters, 125(??):220–226, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300731.

Yakubovich:2015:DAM

[Yak15] Yu. Yakubovich. On descents after maximal values in samples of discrete random variables. Statistics & Probability Letters, 105(??):203–208, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215002060.

Yamada:2015:FST

[Yam15] Toshihiro Yamada. A formula of small time expansion for Young SDE driven by fractional Brownian motion. Statistics & Probability Letters, 101(??):64–72, June 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215000619.

Yang:2012:EPL

[Yan12] Xiangfeng Yang. An elementary proof of the lower bound of Cramér’s Theorem in R^{d}. Statistics & Probability Letters, 82(2):291–294, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003282.

Yang:2013:OBS

[Yan13] Po Yang. Optimal blocking and semifoldover plans for 2^{n−p} designs. Statistics & Probability Letters, 83(2):624–630, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003999.

Yan:2014:EAB

[Yan14] Jun Yan. Estimations and asymptotic behaviors of coherent entropic risk measure for sums of random variables. Statistics & Probability Letters, 91(??):171–180, August 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400159X.

Yan:2015:DCC

[Yan15a] Jun Yan. Deviations of convex and coherent entropic risk measures. Statistics & Probability Letters, 100(??):56–66, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000516.

Yan:2015:NAD

[Yan15b] Ting Yan. A note on asymptotic distributions in maximum entropy models for networks. Statistics & Probability Letters, 98(??):1–5, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214004064.

Yang:2015:EUT

[Yan15c] Xiangfeng Yang. Exact upper tail probabilities of random series. Statistics & Probability Letters, 99(??):13–19, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000048.

Yan:2017:DAB

[Yan17a] Jun Yan. Deviations and asymptotic behavior of convex and coherent entropic risk measures for compound Poisson process influenced by jump times. Statistics & Probability Letters, 125(??): 71–79, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300470.

Yang:2017:MLT

[Yan17b] Xu Yang. Maximum likelihood type estimation for discretely observed CIR model with small α-stable noises. Statistics & Probability Letters, 120(??):18–27, January 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521630178X.

Yang:2019:SLL

[Yan19a] Hui Yang. Scaling limit of the local time of the reflected (1, 2)-random walk. Statistics & Probability Letters, 155(??): Article 108578, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521930224X.

Yang:2019:RBS

[Yan19b] Xue Yang. Reflected backward stochastic partial differential equations with jumps in a convex domain. Statistics & Probability Letters, 152(??):126–136, September 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301221.

Yao:2012:NMP

[Yao12a] Jianfeng Yao. A note on a Marcenko–Pastur type theorem for time series. Statistics & Probability Letters, 82(1):22–28, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002756.

Yao:2012:BCN

[Yao12b] Weixin Yao. A bias corrected nonparametric regression estimator. Statistics & Probability Letters, 82(2):274–282, February 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211003270.

Yao:2013:NGS

[Yao13a] Changlong Yao. A note on geodesics for supercritical continuum percolation. Statistics & Probability Letters, 83(3):797– 804, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212004464.

Yao:2013:NEA

[Yao13b] Weixin Yao. A note on EM algorithm for mixture models. Statistics & Probability Letters, 83(2):519–526, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212003896.

Yao:2014:RPA

[Yao14] Yi-Ching Yao. Rate of Poisson approximation for nearest neighbor counts in large-dimensional Poisson point processes. Statistics & Probability Letters, 92(??):143–147, September 1–262, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214001904.

Yao:2018:MDM

[Yao18] Nian Yao. Moderate deviations for multivariate Hawkes processes. Statistics & Probability Letters, 140(??):71–76, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301809.

Yasuda:2013:MPA

[Yas13] Kumi Yasuda. Markov processes on the Adeles and Chebyshev function. Statistics & Probability Letters, 83(1):238–244, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003458.

Yaskov:2019:PRS

[Yas19] Pavel Yaskov. On pathwise Riemann–Stieltjes integrals. Statistics & Probability Letters, 150(??):101–107, July 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300574.

Yatracos:2015:BSS

[Yat15] Yannis G. Yatracos. Balancing scores for simultaneous comparisons of multiple treatments. Statistics & Probability Letters, 107:178–182, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215003090.

Yang:2016:BUP

[YAT16a] Jun Yang, Fady Alajaji, and Glen Takahara. On bounding the union probability using partial weighted information. Statistics & Probability Letters, 116(??):38–44, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300323.

Yatracos:2016:BRV

[Yat16b] Yannis G. Yatracos. Bias reduction with variable percent bias reducing matching. Statistics & Probability Letters, 110(??):181– 184, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215301723.

Yazigi:2015:RSS

[Yaz15] Adil Yazigi. Representation of self-similar Gaussian processes. Statistics & Probability Letters, 99(??):94–100, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000188.

Yanev:2016:CED

[YC16] George P. Yanev and Santanu Chakraborty. A characterization of exponential distribution and the Sukhatme– Rényi decomposition of exponential maxima. Statistics & Probability Letters, 110(??):94–102, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003934.

Yang:2014:ROA

[YCL14] Xue Yang, Hao Chen, and Min-Qian Liu. Resolvable orthogonal array-based uniform sliced Latin hypercube designs. Statistics & Probability Letters, 93(??):108–115, October 1–142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002235.

Yong:2016:DWR

[YCQY16] Zhang Yong, Siyu Chen, Hong Qin, and Ting Yan. Directed weighted random graphs with an increasing bi-degree sequence. Statistics & Probability Letters, 119(??):235–240, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301523.

Yao:2012:EST

[YD12] Fengqi Yao and Feiqi Deng. Exponential stability in terms of two measures of impulsive stochastic functional differential systems via comparison principle. Statistics & Probability Letters, 82(6):1151–1159, June 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000065.

Yu:2013:DSD

[YDG13] Zhou Yu, Yuexiao Dong, and Ranwei Guo. On determining the structural dimension via directional regression. Statistics & Probability Letters, 83(4):987–992, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004816.

Yang:2010:NED

[YE10] Zhaojun Yang and Christian-Oliver Ewald. On the nonequilibrium density of geometric mean reversion. Statistics & Probability Letters, 80(7–8):608–611, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004751.

Ye:2016:CTM

[Ye16] Zi Ye. Cramér type moderate deviations for the number of renewals. Statistics & Probability Letters, 119(??):194–199, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216301341.

Yeh:2011:MSL

[Yeh11] Hsiaw-Chan Yeh. A multivariate semi-logistic autoregressive process and its characterization. Statistics & Probability Letters, 81(9):1370–1379, September 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211001301.

Yan:2013:MEM

[YG13] Jun Yan and Fuqing Gao. The minimal entropy martingale measure of a jump process influenced by jump times. Statistics & Probability Letters, 83(1):83–88, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003380.

Yang:2014:REE

[YGL14] Hu Yang, Chaohui Guo, and Jing Lv. A robust and efficient estimation method for single-index varying-coefficient models. Statistics & Probability Letters, 94(??):119–127, November 1– 272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002491.

Yin:2013:ENR

[YGTT13] Zanhua Yin, Wei Gao, Man-Lai Tang, and Guo-Liang Tian. Estimation of nonparametric regression models with a mixture of Berkson and classical errors. Statistics & Probability Letters, 83(4):1151–1162, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521300014X.

Yang:2014:LDL

[YH14] Wenzhi Yang and Shuhe Hu. Large deviation for a least squares estimator in a nonlinear regression model. Statistics & Probability Letters, 91(??):135–144, August 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521400162X.

Yuan:2015:CVE

[YH15] Demei Yuan and Xuemei Hu. A conditional version of the extended Kolmogorov–Feller weak law of large numbers. Statistics & Probability Letters, 97(??):99–107, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003757.

Yang:2016:NDM

[YH16] Jianping Yang and Taizhong Hu. New developments on the L_{p}-metric between a probability distribution and its distortion. Statistics & Probability Letters, 110(??):236–243, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003582.

Yang:2017:KEM

[YH17] Lifang Yang and Jiang Hu. Kernel estimators for marčenko-Pastur law of quaternion sample covariance matrices. Statistics & Probability Letters, 126(??):230–237, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301153.

You:2015:RES

[YHMM15] Surong You, Liangjian Hu, Wei Mao, and Xuerong Mao. Robustly exponential stabilization of hybrid uncertain systems by feedback controls based on discrete-time observations. Statistics & Probability Letters, 102(??):8–16, July 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000954.

Yang:2011:TPP

[YHW11] Yingying Yang, Shuhe Hu, and Tao Wu. The tail probability of the product of dependent random variables from max-domains of attraction. Statistics & Probability Letters, 81(12):1876–1882, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211002288.

Young:2016:DMS

[YHY16] Phil D. Young, Jane L. Harvill, and Dean M. Young. A derivation of the multivariate singular skew-normal density function. Statistics & Probability Letters, 117(??):40–45, October 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216300463.

Yang:2015:CTE

[YIS15] Yang Yang, Egle Ignataviciute, and Jonas Siaulys. Conditional tail expectation of randomly weighted sums with heavy-tailed distributions. Statistics & Probability Letters, 105(??):20–28, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215001777.

Yao:2012:EHI

[YJ12] Lili Yao and Wenxin Jiang. On extensions of Hoeffding’s inequality for panel data. Statistics & Probability Letters, 82(3):446– 454, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003567.

Yang:2016:SBM

[YJLL16] Aijun Yang, Xuejun Jiang, Pengfei Liu, and Jinguan Lin. Sparse Bayesian multinomial probit regression model with correlation prior for high-dimensional data classification. Statistics & Probability Letters, 119(??):241–247, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301535.

Yu:2011:MTS

[YK11] Jihnhee Yu and James L. Kepner. On the maximum total sample size of a group sequential test about bivariate binomial proportions. Statistics & Probability Letters, 81(7):829–835, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211000629.

Young:2017:ISM

[YKY17] Phil D. Young, David J. Kahle, and Dean M. Young. On the independence of singular multivariate skew-normal subvectors. Statistics & Probability Letters, 122(??):58–62, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302036.

Yang:2010:ELS

[YL10] Hu Yang and Tingting Li. Empirical likelihood for semiparametric varying coefficient partially linear models with longitudinal data. Statistics & Probability Letters, 80(2):111–121, January 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003666.

Yang:2019:AFT

[YL19] Haizhong Yang and Jinzhu Li. On asymptotic finite-time ruin probability of a renewal risk model with subexponential main claims and delayed claims. Statistics & Probability Letters, 149(??):153–159, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219300434.

Yin:2019:CPM

[YLHY19] Cuihong Yin, X. Sheldon Lin, Rongtan Huang, and Haili Yuan. On the consistency of penalized MLEs for Erlang mixtures. Statistics & Probability Letters, 145(??):12–20, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302797.

Yang:2010:LPL

[YLS10] Yang Yang, Remigijus Leipus, and Jonas Siaulys. Local precise large deviations for sums of random variables with O-regularly varying densities. Statistics & Probability Letters, 80(19–20):1559–1567, October 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210001720.

Yang:2012:TPR

[YLS12] Yang Yang, Remigijus Leipus, and Jonas Siaulys. Tail probability of randomly weighted sums of subexponential random variables under a dependence structure. Statistics & Probability Letters, 82(9):1727–1736, September 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212001915.

Yang:2014:CPM

[YLS14] Yang Yang, Remigijus Leipus, and Jonas Siaulys. Closure property and maximum of randomly weighted sums with heavytailed increments. Statistics & Probability Letters, 91(??):162– 170, August 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214001606.

Yoo:2015:TNO

[Yoo15] Jae Keun Yoo. A theoretical note on optimal sufficient dimension reduction with singularity. Statistics & Probability Letters, 99(??):109–113, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000103.

Yordzhev:2014:PTR

[Yor14] Krasimir Yordzhev. On the probability of two randomly generated S-permutation matrices to be disjoint. Statistics & Probability Letters, 91(??):47–51, August 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001370.

Younso:2017:CNK

[You17] Ahmad Younso. On the consistency of a new kernel rule for spatially dependent data. Statistics & Probability Letters, 131(??):64–71, December 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302687.

Youndje:2018:EBM

[You18] É. Youndjé. Equivalence between Mallows and Girone–Cifarelli tests. Statistics & Probability Letters, 141(??):125–128, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521830227X.

Yu:2010:ACM

[YQW10] Qiqing Yu, Hao Qin, and Jiaping Wang. About conditional masking probability models. Statistics & Probability Letters, 80(15–16):1174–1179, August 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000933.

Yang:2011:LTD

[YR11] Ting Yang and Yan-Xia Ren. Limit theorem for derivative martingale at criticality w.r.t branching Brownian motion. Statistics & Probability Letters, 81(2):195–200, February 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003251.

Yan:2010:CLT

[YS10] Litan Yan and Guangjun Shen. On the collision local time of sub-fractional Brownian motions. Statistics & Probability Letters, 80(5–6):296–308, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004210.

Yamada:2012:APC

[YS12] Takayuki Yamada and Tetsuro Sakurai. Asymptotic power comparison of three tests in GMANOVA when the number of observed points is large. Statistics & Probability Letters, 82(3):692– 698, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211003865.

Yang:2013:SPD

[YS13] Haizhong Yang and Suting Sun. Subexponentiality of the product of dependent random variables. Statistics & Probability Letters, 83(9):2039–2044, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001788.

Yao:2019:NUE

[YS19] Xuan Yao and Eric Slud. Nonexistence of an unbiased estimating function for the Cox model. Statistics & Probability Letters, 152(??):21–27, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301099.

Suen:2014:CRM

[ySDM14] Chung yi Suen, Ashish Das, and C. K. Midha. On the construction of restricted minimum aberration designs. Statistics & Probability Letters, 94(??):162–169, November 1–272, 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214002569.

Yang:2010:SDU

[YSLL10] Jianfeng Yang, Fasheng Sun, Dennis K. J. Lin, and Min-Qian Liu. A study on design uniformity under errors in the level values. Statistics & Probability Letters, 80(19–20):1467–1471, October 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715210001574.

Yang:2019:EDD

[YSZ19] Yang Yang, Wen Su, and Zhimin Zhang. Estimating the discounted density of the deficit at ruin by Fourier cosine series expansion. Statistics & Probability Letters, 146(??):147– 155, March 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218303626.

Yu:2011:MTB

[Yu11] Lin Yu. Martingale transforms between Hardy–Orlicz spaces Q_{Φ_{1}} and Q_{Φ_{2}} of martingales. Statistics & Probability Letters, 81(8):1086–1093, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000824.

Yu:2017:NVM

[Yu17] Yaming Yu. On normal variance-mean mixtures. Statistics & Probability Letters, 121(??):45–50, February 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302012.

Yu:2019:NLA

[Yu19] Xianye Yu. Non-Lipschitz anticipated backward stochastic differential equations driven by fractional Brownian motion. Statistics & Probability Letters, 155(??):Article 108582, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302287.

Yun:2018:MDP

[Yun18] Youngyun Yun. The moments of a diffusion process. Statistics & Probability Letters, 138(??):36–41, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521830052X.

Yang:2010:ARP

[YW10] Yang Yang and Yuebao Wang. Asymptotics for ruin probability of some negatively dependent risk models with a constant interest rate and dominatedly-varying-tailed claims. Statistics & Probability Letters, 80(3–4):143–154, February 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209003708.

Yi:2017:BSM

[YW17] Taihe Yi and Zhengming Wang. Bayesian sieve method for piece-wise smooth regression. Statistics & Probability Letters, 130(??):5–11, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217302389.

Yu:2010:LLU

[YWC10] Changjun Yu, Yuebao Wang, and Zhaolei Cui. Lower limits and upper limits for tails of random sums supported on R. Statistics & Probability Letters, 80(13–14):1111–1120, July 1– 15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000751.

Yi:2014:NJN

[YWJ14] Rui Yi, Lian Wu, and Yong Jiao. New John–Nirenberg inequalities for martingales. Statistics & Probability Letters, 86(??): 68–73, March 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213004094.

Yu:2010:CCE

[YWY10] Changjun Yu, Yuebao Wang, and Yang Yang. The closure of the convolution equivalent distribution class under convolution roots with applications to random sums. Statistics & Probability Letters, 80(5–6):462–472, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004489.

Yin:2014:APM

[YWZ14] Changming Yin, Zhanfeng Wang, and Hong Zhang. Asymptotic properties of maximum likelihood estimator for two-step logit models. Statistics & Probability Letters, 85(??):135–143, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213003957.

Yang:2014:SCB

[YXL14] Suigen Yang, Liugen Xue, and Gaorong Li. Simultaneous confidence band for single-index random effects models with longitudinal data. Statistics & Probability Letters, 85(??):6–14, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521300357X.

Yen:2011:TFL

[YY11a] Ju-Yi Yen and Marc Yor. Truncation functions and Laplace transform. Statistics & Probability Letters, 81(3):417–419, March 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210003469.

Yin:2011:OTD

[YY11b] Chuancun Yin and Kam Chuen Yuen. Optimality of the threshold dividend strategy for the compound Poisson model. Statistics & Probability Letters, 81(12):1841–1846, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002537.

Yen:2013:ILB

[YY13] Ju-Yi Yen and Marc Yor. On an identity in law between Brownian quadratic functionals. Statistics & Probability Letters, 83(9):2015–2018, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001740.

Yu:2014:MTW

[YY14] Lin Yu and Huan Yin. Martingale transforms and weak Orlicz–Hardy spaces of predictable martingales. Statistics & Probability Letters, 89(??):89–98, June 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000819.

Yang:2016:RPE

[YY16] Jing Yang and Hu Yang. A robust penalized estimation for identification in semiparametric additive models. Statistics & Probability Letters, 110(??):268–277, March 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003454.

Yan:2018:BFS

[YY18] Litan Yan and Xiuwei Yin. Bismut formula for a stochastic heat equation with fractional noise. Statistics & Probability Letters, 137(??):165–172, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300294.

Yan:2017:DIL

[YYC17] Litan Yan, Xianye Yu, and Ruqing Chen. Derivative of intersection local time of independent symmetric stable motions. Statistics & Probability Letters, 121(??):18–28, February 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303060.

Yan:2016:ABS

[YYS16] Litan Yan, Xianye Yu, and Xichao Sun. Asymptotic behavior of the solution of the fractional heat equation. Statistics & Probability Letters, 117(??):54–61, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300578.

Yang:2018:UMM

[YYS18] Xue Yang, Gui-Jun Yang, and Ya-Juan Su. Uniform minimum moment aberration designs. Statistics & Probability Letters, 137(??):26–33, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218300026.

Yano:2017:WCT

[YYY17] Kouji Yano, Yuko Yano, and Ju-Yi Yen. Weak convergence of h-transforms for one-dimensional diffusions. Statistics & Probability Letters, 122(??):152–156, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302474.

Yuan:2013:CNA

[YZ13] Demei Yuan and Jianhua Zheng. Conditionally negative association resulting from multinomial distribution. Statistics & Probability Letters, 83(10):2222–2227, October 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213002095.

Yang:2014:NLR

[YZ14] Jianping Yang and Weiwei Zhuang. A note on likelihood ratio ordering of order statistics from two samples. Statistics & Probability Letters, 84(??):135–139, January 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003386.

Yan:2016:AST

[YZ16] Ting Yan and Yunpeng Zhao. Asymptotics of score test in the generalized β-model for networks. Statistics & Probability Letters, 119(??):163–169, December 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216301365.

Yi:2018:PUU

[YZ18] Si-Yu Yi and Yong-Dao Zhou. Projection uniformity under mixture discrepancy. Statistics & Probability Letters, 140(??):96–105, September 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218301858.

Zarepour:2012:RSD

[ZA12] Mahmoud Zarepour and Luai Al Labadi. On a rapid simulation of the Dirichlet process. Statistics & Probability Letters, 82(5): 916–924, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212000302.

Zajkowski:2014:NGT

[Zaj14] Krzysztof Zajkowski. A note on the gambling team method. Statistics & Probability Letters, 85(??):45–50, February 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003647.

Zajkowski:2019:NSE

[Zaj19] Krzysztof Zajkowski. Norms of sub-exponential random vectors. Statistics & Probability Letters, 152(??):147–152, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219301312.

Zamanzade:2012:PBT

[ZAV12] Ehsan Zamanzade, Nasser Reza Arghami, and Michael Vock. Permutation-based tests of perfect ranking. Statistics & Probability Letters, 82(12):2213–2220, December 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212002945.

Zhao:2010:OPC

[ZB10] Peng Zhao and N. Balakrishnan. Ordering properties of convolutions of heterogeneous Erlang and Pascal random variables. Statistics & Probability Letters, 80(11–12):969–974, June 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000507.

Zhao:2011:NRC

[ZB11] Peng Zhao and N. Balakrishnan. New results on comparisons of parallel systems with heterogeneous gamma components. Statistics & Probability Letters, 81(1):36–44, January 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521000266X.

Zhu:2015:BSD

[ZB15] Xiaojun Zhu and N. Balakrishnan. Birnbaum–Saunders distribution based on Laplace kernel and some properties and inferential issues. Statistics & Probability Letters, 101(??):1–10, June 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000577.

Zhu:2019:SSN

[ZBL19] Jiahui Zhu, Zdzisław Brzeźniak, and Wei Liu. L_{p}-solutions for stochastic Navier–Stokes equations with jump noise. Statistics & Probability Letters, 155(??):Article 108563, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302093.

Zhao:2011:RDQ

[ZBZ11] Mu Zhao, Fangfang Bai, and Yong Zhou. Relative deficiency of quantile estimators for left truncated and right censored data.

Statistics & Probability Letters, 81(11):1725–1732, November 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002264.

Zhou:2012:SES

[ZC12] Jianjun Zhou and Min Chen. Spline estimators for semifunctional linear model. Statistics & Probability Letters, 82 (3):505–513, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003816.

Zong:2013:HIM

[ZC13] Gaofeng Zong and Zengjing Chen. Harnack inequality for mean-field stochastic differential equations. Statistics & Probability Letters, 83(5):1424–1432, May 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000473.

Zhao:2019:IML

[ZC19] Linjie Zhao and Dayue Chen. The invariant measures and the limiting behaviors of the facilitated TASEP. Statistics & Probability Letters, 155(??):Article 108557, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715219302032.

Zhang:2015:ECM

[ZCM15] Panpan Zhang, Chen Chen, and Hosam Mahmoud. Explicit characterization of moments of balanced triangular Pólya urns by an elementary approach. Statistics & Probability Letters, 96(??):149–153, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003277.

Zhao:2018:ERA

[ZCZ18] Jun Zhao, Yingyu Chen, and Yi Zhang. Expectile regression for analyzing heteroscedasticity in high dimension. Statistics & Probability Letters, 137(??):304–311, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300506.

Zhao:2018:SPL

[ZDD18] Xianghua Zhao, Hua Dong, and Hongshuai Dai. On spectrally positive Lévy risk processes with Parisian implementation delays in dividend payments. Statistics & Probability Letters, 140(??):176–184, September 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302001.

Zhang:2017:FIC

[ZDM17] Qingzhao Zhang, Xiaogang Duan, and Shuangge Ma. Focused information criterion and model averaging with generalized rank regression. Statistics & Probability Letters, 122(??):11–19, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302206.

Zhang:2014:STU

[ZDX14] Liangyong Zhang, Xiaofang Dong, and Xingzhong Xu. Sign tests using ranked set sampling with unequal set sizes. Statistics & Probability Letters, 85(??):69–77, February 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213003878.

Zeng:2014:DEL

[Zen14] Xingyuan Zeng. Distribution of eigenvalues of large Euclidean matrices generated from ℓ_{p} ellipsoid. Statistics & Probability Letters, 91(??):181–191, August 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001576.

Zeng:2015:NLR

[Zen15] Xingyuan Zeng. A note on the large random inner-product kernel matrices. Statistics & Probability Letters, 99(??):192–201, April 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000206.

Zeng:2017:LED

[Zen17] Xingyuan Zeng. Limiting empirical distribution for eigenvalues of products of random rectangular matrices. Statistics & Probability Letters, 126(??):33–40, July 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217300810.

Zeng:2019:SDL

[Zen19] Xingyuan Zeng. Spectral distribution of large generalized random kernel matrices. Statistics & Probability Letters, 152 (??):100–110, September 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219301191.

Zhang:2013:RTG

[ZF13] HengMin Zhang and ShengJun Fan. A representation theorem for generators of BSDEs with finite or infinite time intervals and linear-growth generators. Statistics & Probability Letters, 83(3):724–734, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212004361.

Zhou:2016:CEO

[ZF16] Jie Zhou and Hai-Lin Feng. Consistent estimation of ordinary differential equation when the transformation parameter is unknown. Statistics & Probability Letters, 115(??):60–69, August 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000377.

Zhang:2013:MRC

[ZFZ13] Shuhong Zhang, Xiuying Feng, and Peng Zhao. On a mixture representation of the conditional inactivity time of a coherent system. Statistics & Probability Letters, 83(10):2297–2307, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002228.

Zhao:2010:LDT

[ZG10] Shoujiang Zhao and Fuqing Gao. Large deviations in testing Jacobi model. Statistics & Probability Letters, 80(1):34– 41, January 2010. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715209003502.

Zhao:2016:SSD

[ZG16] Bo Zhao and Joseph Glaz. Scan statistics for detecting a local change in variance for normal data with unknown population variance. Statistics & Probability Letters, 110(??):137– 145, March 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016771521500406X.

Zhou:2017:NUE

[ZG17a] Bu Zhou and Jia Guo. A note on the unbiased estimator of Σ^{2}. Statistics & Probability Letters, 129(??):141–146, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521730189X.

Zhou:2017:NCC

[ZG17b] Qi Zhou and Bing Guo. A note on constructing clear compromise plans. Statistics & Probability Letters, 130(??):17–24, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521730233X.

Zhang:2019:LSP

[ZG19] Xiaoke Zhang and Joseph L. Gastwirth. Large sample properties of a new measure of income inequality. Statistics & Probability Letters, 145(??):50–56, February 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218302840.

Zhigljavsky:2016:DDU

[ZGG16] Anatoly Zhigljavsky, Nina Golyandina, and Svyatoslav Gryaznov. Deconvolution of a discrete uniform distribution. Statistics & Probability Letters, 118(??):37–44, November 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300876.

Zheng:2013:GRS

[ZGK13] Qi Zheng, Colin Gallagher, and K. B. Kulasekera. The growth rate of significant regressors for high dimensional data. Statistics & Probability Letters, 83(9):1969–1972, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213001508.

Zhelonkin:2012:RTS

[ZGR12] Mikhail Zhelonkin, Marc G. Genton, and Elvezio Ronchetti. On the robustness of two-stage estimators. Statistics & Probability Letters, 82(4):726–732, April 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211004068.

Zhou:2018:NEA

[ZGZ18] Qi Zhou, Bing Guo, and Shengli Zhao. A note on extending the alias length pattern. Statistics & Probability Letters, 138(??): 82–89, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218301056.

Zhang:2012:FPT

[ZH12a] Xuan Zhang and Zhenting Hou. The first-passage times of phase semi-Markov processes. Statistics & Probability Letters, 82(1):40–48, January 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211002859.

Zhou:2012:CIB

[ZH12b] Ling Zhou and Ze-Chun Hu. Chebyshev’s inequality for Banachspace-valued random elements. Statistics & Probability Letters, 82(5):925–931, May 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715212000351.

Zhang:2014:NPR

[ZH14] Haimeng Zhang and Chunfeng Huang. A note on processes with random stationary increments. Statistics & Probability Letters, 94(??):153–161, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002545.

Zhang:2010:NPG

[Zha10] Shunpu Zhang. A note on the performance of the gamma kernel estimators at the boundary. Statistics & Probability Letters, 80(7–8):548–557, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004593.

Zhang:2011:PEF

[Zha11a] Chenhua Zhang. Parameter estimation for first-order bifurcating autoregressive processes with Weibull innovations. Statistics & Probability Letters, 81(12):1961–1969, December 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715211002781.

Zhao:2011:ELI

[Zha11b] Yichuan Zhao. Empirical likelihood inference for the accelerated failure time model. Statistics & Probability Letters, 81(5):603– 610, May 2011. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715211000150.

Zhao:2012:SRF

[Zha12] Feng-Zhen Zhao. Some recursive formulas related to inverse moments of the random variables with binomial-type distributions. Statistics & Probability Letters, 82(7):1290–1296, July 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212001198.

Zhang:2013:HIS

[Zha13a] Shao-Qin Zhang. Harnack inequality for semilinear SPDE with multiplicative noise. Statistics & Probability Letters, 83(4):1184– 1192, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213000102.

Zhao:2013:HMS

[Zha13b] James Y. Zhao. The Hopfield model with superlinearly many patterns. Statistics & Probability Letters, 83(1):350–356, January 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003720.

Zhang:2014:UAT

[Zha14a] Chenhua Zhang. Uniform asymptotics for the tail probability of weighted sums with heavy tails. Statistics & Probability Letters, 94(??):221–229, November 1–272, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214002600.

Zhang:2014:SCI

[Zha14b] Guoyi Zhang. Simultaneous confidence intervals for several inverse Gaussian populations. Statistics & Probability Letters, 92(??):125–131, September 1–262, 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001898.

Zhang:2016:BAS

[Zha16] Yazhe Zhang. Binomial approximation for sum of indicators with dependent neighborhoods. Statistics & Probability Letters, 119(??):146–154, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301353.

Zhang:2017:LLN

[Zha17a] Na Zhang. On the law of large numbers for discrete Fourier transform. Statistics & Probability Letters, 120(??):101–107, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301821.

Zhang:2017:EIM

[Zha17b] Tonglin Zhang. An example of inconsistent MLE of spatial covariance parameters under increasing domain asymptotics. Statistics & Probability Letters, 120(??):108–113, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216302097.

Zhang:2017:LLI

[Zha17c] Yong Zhang. The limit law of the iterated logarithm for linear processes. Statistics & Probability Letters, 122(??):147– 151, March 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215303321.

Zhang:2018:OAA

[Zha18] Yiying Zhang. Optimal allocation of active redundancies in weighted k-out-of-n systems. Statistics & Probability Letters, 135(??):110–117, April 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303760.

Zhang:2019:ITL

[Zha19a] Qingyang Zhang. Independence test for large sparse contingency tables based on distance correlation. Statistics & Probability Letters, 148(??):17–22, May 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715218304048.

Zhang:2019:SPI

[Zha19b] Weiwei Zhang. Super Poincaré inequality for a dynamic model of the two-parameter Dirichlet process. Statistics & Probability Letters, 151(??):97–105, August 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300318.

Zheng:2011:SRM

[Zhe11] Yanbing Zheng. Shape restriction of the multi-dimensional Bernstein prior for density functions. Statistics & Probability Letters, 81(6):647–651, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000459.

Zheng:2017:RHU

[Zhe17] Songfeng Zheng. A refined Hoeffding’s upper tail probability bound for sum of independent random variables. Statistics & Probability Letters, 131(??):87–92, December 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217302729.

Zhou:2013:NME

[ZHL13] Jie Zhou, Lu Han, and Sanyang Liu. Nonlinear mixedeffects state space models with applications to HIV dynamics. Statistics & Probability Letters, 83(5):1448–1456, May 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213000448.

Zhou:2010:CMC

[Zho10] Xingcai Zhou. Complete moment convergence of moving average processes under ϕ-mixing assumptions. Statistics & Probability Letters, 80(5–6):285–292, March 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004192.

Zhou:2013:GEE

[Zho13a] Ding-Xuan Zhou. On grouping effect of elastic net. Statistics & Probability Letters, 83(9):2108–2112, September 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213001752.

Zhou:2013:HTD

[Zho13b] Ke Zhou. Hitting time distribution for skip-free Markov chains: a simple proof. Statistics & Probability Letters, 83(7):1782– 1786, July 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213001296.

Zhou:2015:DIM

[Zho15] Jie Zhou. Detection of influential measurement for ordinary differential equation with application to HIV dynamics. Statistics & Probability Letters, 107:324–332, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003260.

Zhu:2013:MDH

[Zhu13a] Lingjiong Zhu. Moderate deviations for Hawkes processes. Statistics & Probability Letters, 83(3):885–890, March 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212004609.

Zhu:2013:NDE

[Zhu13b] Ying Zhu. Nonparametric density estimation based on the truncated mean. Statistics & Probability Letters, 83(2):445–451, February 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212003951.

Zhu:2014:LDP

[Zhu14] Lingjiong Zhu. Large deviations for product of sums of random variables. Statistics & Probability Letters, 89(??):17–22, June 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771521400073X.

Zhu:2015:LDO

[Zhu15] Lingjiong Zhu. Large deviations for one-dimensional random walks on discrete point processes. Statistics & Probability Letters, 97(??):69–75, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003770.

Zhao:2012:NUM

[ZHWH12] Jing Zhao, ChongZhao Han, Bin Wei, and DeQiang Han. A novel Univariate Marginal Distribution Algorithm based discretization algorithm. Statistics & Probability Letters, 82(11):2001–2007, November 2012. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715212001976.

Zintout:2013:TVD

[Zin13] Rola Zintout. The total variation distance between two double Wiener–Itô integrals. Statistics & Probability Letters, 83 (10):2160–2167, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715213001983.

Zhang:2017:ISS

[ZIPP17] Qiang Zhang, Edward H. Ip, Junhao Pan, and Robert Plemmons. Individual-specific, sparse inverse covariance estimation in generalized estimating equations. Statistics & Probability Letters, 122(??):96–103, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215304338.

Zhang:2013:SMO

[ZJ13] Na Zhang and Guangyan Jia. Stochastic monotonicity and order-preservation for a type of nonlinear semigroups. Statistics & Probability Letters, 83(1):422–429, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003744.

Zhao:2015:BEB

[ZJ15] Mu Zhao and Hongmei Jiang. Berry–Esseen bounds for the percentile residual life function estimators. Statistics & Probability Letters, 104(??):133–140, September 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001601.

Zhao:2013:NEM

[ZJL13] Mu Zhao, Hongmei Jiang, and Xu Liu. A note on estimation of the mean residual life function with left-truncated and rightcensored data. Statistics & Probability Letters, 83(10):2332– 2336, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715213002320.

Zhang:2018:SFA

[ZJL18] Tianfang Zhang, Junyu Jiang, and Zhiming Li. Some families of asymmetric nested orthogonal arrays and asymmetric sliced orthogonal arrays. Statistics & Probability Letters, 141(??):109– 113, October 2018. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715218302268.

Zhang:2013:BMS

[ZJS13] Jingsi Zhang, Wenxin Jiang, and Xiaofeng Shao. Bayesian model selection based on parameter estimates from subsamples. Statistics & Probability Letters, 83(4):979–986, April 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715212004798.

Zhao:2013:DSQ

[ZJZ13] Mu Zhao, Hongmei Jiang, and Yong Zhou. L_{1}-deficiency of the sample quantile estimator with respect to a kernel quantile estimator. Statistics & Probability Letters, 83(10):2399–2406, October 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002472.

Zheng:2016:RET

[ZJZ16] Wei Zheng, Yong Jin, and Guoyi Zhang. Recursive estimation of time-average variance constants through prewhitening. Statistics & Probability Letters, 114(??):30–37, July 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216000407.

Zaigraev:2010:EBP

[ZK10] Alexander Zaigraev and Serguei Kaniovski. Exact bounds on the probability of at least k successes in n exchangeable Bernoulli trials as a function of correlation coefficients. Statistics & Probability Letters, 80(13–14):1079–1084, July 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715210000714.

Zeifman:2014:PBC

[ZK14] A. I. Zeifman and V. Yu. Korolev. On perturbation bounds for continuous-time Markov chains. Statistics & Probability Letters, 88(??):66–72, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214000480.

Zeifman:2015:TSB

[ZK15] A. I. Zeifman and V. Yu. Korolev. Two-sided bounds on the rate of convergence for continuous-time finite inhomogeneous Markov chains. Statistics & Probability Letters, 103(??):30–36, August 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215001236.

Zheng:2010:LAS

[ZKG10] Qi Zheng, K. B. Kulasekera, and Colin Gallagher. Local adaptive smoothing in kernel regression estimation. Statistics & Probability Letters, 80(7–8):540–547, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715209004581.

Zeifman:2018:LBR

[ZKSK18] A. I. Zeifman, V. Yu. Korolev, Ya. A. Satin, and K. M. Kiseleva. Lower bounds for the rate of convergence for continuous-time inhomogeneous Markov chains with a finite state space. Statistics & Probability Letters, 137(??):84–90, June 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218300038.

Zhao:2012:MDS

[ZL12] Shoujiang Zhao and Qiaojing Liu. Moderate deviations for some nonparametric estimators with errors in variables. Statistics & Probability Letters, 82(6):1175–1184, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000624.

Zhou:2014:ATL

[ZL14] Zhiyong Zhou and Zhengyan Lin. Asymptotic theory for LAD estimation of moderate deviations from a unit root. Statistics & Probability Letters, 90(??):25–32, July 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214000972.

Zheng:2015:RTG

[ZL15] Shiqiu Zheng and Shoumei Li. Representation theorems for generators of BSDEs with monotonic and convex growth generators. Statistics & Probability Letters, 97(??):197–205, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214004052.

Zhao:2016:LAN

[ZL16a] Weihua Zhao and Heng Lian. Local asymptotics for nonparametric quantile regression with regression splines. Statistics & Probability Letters, 117(??):209–215, October 2016. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216300839.

Zhou:2016:HOE

[ZL16b] Wei Zhou and Chengxiu Ling. Higher-order expansions of powered extremes of normal samples. Statistics & Probability Letters, 111(??):12–17, April 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521530198X.

Zou:2017:ETM

[ZL17] Yuye Zou and Xiangdong Liu. An extension of a theorem of mikosch. Statistics & Probability Letters, 120(??):81–86, January 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715216301791.

Zhang:2018:MSL

[ZL18] Lixin Zhang and Jinghang Lin. Marcinkiewicz’s strong law of large numbers for nonlinear expectations. Statistics & Probability Letters, 137(??):269–276, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300336.

Zhao:2018:LDP

[ZLC18] Shoujiang Zhao, Qiaojing Liu, and Ting Chen. On the large deviation principle for maximum likelihood estimator of α-Brownian bridge. Statistics & Probability Letters, 138(??):143–150, July 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301159.

Zheng:2011:PIS

[ZLT11] Jing Zheng, Zhengyan Lin, and Changqing Tong. The packing indices for some Lévy processes. Statistics & Probability Letters, 81(11):1683–1689, November 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002343.

Zhang:2016:JEL

[ZLZ16] Zhengjia Zhang, Tianqing Liu, and Baoxue Zhang. Jackknife empirical likelihood inferences for the population mean with ranked set samples. Statistics & Probability Letters, 108:16–22, January 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215003351.

Zhang:2016:DCP

[ZM16a] Panpan Zhang and Hosam M. Mahmoud. Distributions in a class of Poissonized urns with an application to Apollonian networks. Statistics & Probability Letters, 115(??):1–7, August 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215302479.

Zhao:2016:RNK

[ZM16b] Ge Zhao and Yanyuan Ma. Robust nonparametric kernel regression estimator. Statistics & Probability Letters, 116(??):72–79, September 2016. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715216300293.

Zamanzade:2017:MEP

[ZM17] Ehsan Zamanzade and M. Mahdizadeh. A more efficient proportion estimator in ranked set sampling. Statistics & Probability Letters, 129(??):28–33, October 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715217301761.

Zhao:2010:AAE

[ZMG10] Huixiu Zhao, Wen-Qing Ma, and Jianhua Guo. The AU algorithm for estimating equations in the presence of missing data. Statistics & Probability Letters, 80(7–8):639–647, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715209004799.

Zhang:2012:ODQ

[ZP12] Chongqi Zhang and Heng Peng. D-optimal designs for quadratic mixture canonical polynomials with spline. Statistics & Probability Letters, 82(6):1095–1101, June 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000600.

Zhou:2015:TSS

[ZP15] Xianbo Zhou and Zhewen Pan. Two-step semiparametric estimation of the Type-3 Tobit model. Statistics & Probability Letters, 105(??):96–105, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215001984.

Zalzadeh:2016:PDN

[ZP16] Saeed Zalzadeh and Franco Pellerey. A positive dependence notion based on componentwise unimodality of copulas. Statistics & Probability Letters, 112(??):51–57, May 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215303400.

Zou:2019:LPB

[ZP19] Nan Zou and Dimitris N. Politis. Linear process bootstrap unit root test. Statistics & Probability Letters, 145(??):74–80, February 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302815.

Zaigraev:2014:MIL

[ZPK14] A. Zaigraev and A. Podraza-Karakulska. Maximum integrated likelihood estimator of the interest parameter when the nuisance parameter is location or scale. Statistics & Probability Letters, 88(??):99–106, May 2014. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214000352.

Zhang:2018:RCN

[ZPZ18] Shucong Zhang, Jing Pan, and Yong Zhou. Robust conditional nonparametric independence screening for ultrahigh-dimensional data. Statistics & Probability Letters, 143(??):95–101, December 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218302785.

Zhou:2012:RBS

[ZR12] Qing Zhou and Yong Ren. Reflected backward stochastic differential equations with time delayed generators. Statistics & Probability Letters, 82(5):979–990, May 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212000594.

Zinodiny:2013:BME

[ZRAN13] S. Zinodiny, S. Rezaei, O. Naghshineh Arjmand, and S. Nadarajah. Bayes minimax estimation of the multivariate normal mean vector under quadratic loss functions. Statistics & Probability Letters, 83(9):2052–2056, September 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521300182X.

Zinodiny:2014:BME

[ZRN14] S. Zinodiny, S. Rezaei, and S. Nadarajah. Bayes minimax estimation of the multivariate normal mean vector under balanced loss function. Statistics & Probability Letters, 93(??):96–101, October 1–142, 2014. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715214002338.

Zinodiny:2017:BME

[ZRN17] S. Zinodiny, S. Rezaei, and S. Nadarajah. Bayes minimax estimation of the mean matrix of matrix-variate normal distribution under balanced loss function. Statistics & Probability Letters, 125(??):110–120, June 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215303254.

Zhang:2019:NBC

[ZS19] Yi-Chen Zhang and Lyudmila Sakhanenko. The naive Bayes classifier for functional data. Statistics & Probability Letters, 152(??):137–146, September 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301208.

Zhu:2019:EUS

[ZT19] Runyu Zhu and Dejian Tian. Existence and uniqueness of solutions for BDSDEs with weak monotonicity coefficients. Statistics & Probability Letters, 153(??):48–55, October 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219301464.

Zamanzade:2015:VER

[ZV15] Ehsan Zamanzade and Michael Vock. Variance estimation in ranked set sampling using a concomitant variable. Statistics & Probability Letters, 105(??):1–5, October 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521500187X.

Zhou:2011:WMW

[ZW11] Weihua Zhou and Jin Wang. On the weighted multivariate Wilcoxon rank regression estimate. Statistics & Probability Letters, 81(6):704–713, June 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521100068X.

Zhang:2012:RPB

[ZW12] Yuanyuan Zhang and Wensheng Wang. Ruin probabilities of a bidimensional risk model with investment. Statistics & Probability Letters, 82(1):130–138, January 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211002987.

Zhang:2013:ODM

[ZW13] Chongqi Zhang and Weng Kee Wong. Optimal designs for mixture models with amount constraints. Statistics & Probability Letters, 83(1):196–202, January 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715212003343.

Zhang:2015:PID

[ZW15a] Lihua Zhang and Yingzhe Wang. L^{1}-Poincaré inequality for discrete time Markov chains. Statistics & Probability Letters, 100(??):93–97, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215000267.

Zhou:2015:OTR

[ZW15b] Jiang Zhou and Lan Wu. Occupation times of refracted double exponential jump diffusion processes. Statistics & Probability Letters, 106(??):218–227, November 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215002643.

Zeng:2018:WDD

[ZW18a] Xiaochen Zeng and Jinru Wang. Wavelet density deconvolution estimations with heteroscedastic measurement errors. Statistics & Probability Letters, 134(??):79–85, March 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217303383.

Zhang:2018:OBI

[ZW18b] Duo Zhang and Min Wang. Objective Bayesian inference for the intraclass correlation coefficient in linear models. Statistics & Probability Letters, 137(??):292–296, June 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218300488.

Zhang:2018:OWS

[ZW18c] Xiaoke Zhang and Jane-Ling Wang. Optimal weighting schemes for longitudinal and functional data. Statistics & Probability Letters, 138(??):165–170, July 2018. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218301214.

Zhou:2017:OTL

[ZWB17] Jiang Zhou, Lan Wu, and Yang Bai. Occupation times of Lévydriven Ornstein–Uhlenbeck processes with two-sided exponential jumps and applications. Statistics & Probability Letters, 125(??): 80–90, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715217300433.

Zhang:2015:NLB

[ZWHQ15] Qionghui Zhang, Zhenghong Wang, Jianwei Hu, and Hong Qin. A new lower bound for wrap-around L_{2}-discrepancy on two and three mixed level factorials. Statistics & Probability Letters, 96(??):133–140, January 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214003113.

Zhou:2017:CEC

[ZWLB17] Sanyu Zhou, Fang Wan, Wei Liu, and Frank Bretz. Computation of an exact confidence set for a maximum point of a univariate polynomial function in a given interval. Statistics & Probability Letters, 122(??):157–161, March 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715216302395.

Zhou:2015:SCP

[ZWT15] Mi Zhou, Huixia Judy Wang, and Yanlin Tang. Sequential change point detection in linear quantile regression models. Statistics & Probability Letters, 100(??):98–103, May 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715215000371.

Zhang:2012:NMT

[ZX12] Jin-Ting Zhang and Shengning Xiao. A note on the modified two-way MANOVA tests. Statistics & Probability Letters, 82(3):519–527, March 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211003877.

Zhao:2017:GCC

[ZX17a] Li Zhao and Xingzhong Xu. Generalized canonical correlation variables improved estimation in high dimensional seemingly unrelated regression models. Statistics & Probability Letters, 126(??):119–126, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300937.

Zheng:2017:PUS

[ZX17b] Jiayu Zheng and Jie Xiong. Pathwise uniqueness for stochastic differential equations driven by pure jump processes. Statistics & Probability Letters, 130(??):100–104, November 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217302481.

Zhao:2018:NTC

[ZX18] Jingxin Zhao and Chuanlong Xie. A nonparametric test for covariate-adjusted models. Statistics & Probability Letters, 133(??):65–70, February 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217303188.

Zhang:2019:NMF

[ZX19] Shuaiqi Zhang and Jie Xiong. A numerical method for forward-backward stochastic equations with delay and anticipated term. Statistics & Probability Letters, 149(??):107– 115, June 2019. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715219300380.

Zhang:2010:OPR

[ZY10a] Zhengcheng Zhang and Yonghong Yang. Ordered properties on the residual life and inactivity time of (n − k + 1)-out-of-n systems under double monitoring. Statistics & Probability Letters, 80(7–8):711–717, April 1–15, 2010. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210000052.

Zhang:2010:GPF

[ZY10b] Zhimin Zhang and Hu Yang. A generalized penalty function in the Sparre–Andersen risk model with two-sided jumps. Statistics & Probability Letters, 80(7–8):597–607, April 1–15, 2010. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S016771520900474X.

Zheng:2011:ELA

[ZY11] Ming Zheng and Wen Yu. An empirical likelihood approach to data analysis under two-stage sampling designs. Statistics & Probability Letters, 81(8):947–956, August 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211000186.

Zhou:2017:PSR

[ZYG17] Jieming Zhou, Xiangqun Yang, and Junyi Guo. Portfolio selection and risk control for an insurer in the Lévy market under mean-variance criterion. Statistics & Probability Letters, 126(??):139–149, July 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217300998.

Zhang:2011:SBB

[ZYL11] Tian-Fang Zhang, Jian-Feng Yang, and Dennis K. J. Lin. Small Box–Behnken design. Statistics & Probability Letters, 81(8):1027–1033, August 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715211000708.

Zhang:2014:SGD

[ZYL14] Wen Zhang, Jun Ye, and Haibo Li. Stability with general decay rates of stochastic differential delay equations with Poisson jumps and Markovian switching. Statistics & Probability Letters, 92(??):1–11, September 2014. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715214001643.

Zhang:2019:NET

[ZYS19] Xuekang Zhang, Haoran Yi, and Huisheng Shu. Nonparametric estimation of the trend for stochastic differential equations driven by small α-stable noises. Statistics & Probability Letters, 151(??):8–16, August 2019. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715219300884.

Zeng:2015:NCM

[ZYW15] Bilin Zeng, Zhou Yu, and Xuerong Meggie Wen. A note on cumulative mean estimation. Statistics & Probability Letters, 96(??):322–327, January 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715214003253.

Zhao:2011:ELC

[ZZ11] Yichuan Zhao and Meng Zhao. Empirical likelihood for the contrast of two hazard functions with right censoring. Statistics & Probability Letters, 81(3):392–401, March 2011. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715210003408.

Zhang:2012:NST

[ZZ12a] Xiang Zhang and Yanbing Zheng. A note on spatial-temporal lattice modeling and maximum likelihood estimation. Statistics & Probability Letters, 82(12):2145–2155, December 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S016771521200291X.

Zhang:2012:ADB

[ZZ12b] Xueying Zhang and Chuanzhou Zhang. Atomic decompositions of Banach lattice-valued martingales. Statistics & Probability Letters, 82(3):664–671, March 2012. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715211003853.

Zhang:2013:SLO

[ZZ13a] Zhongyang Zhang and Lixin Zhang. Scaling limits for onedimensional long-range percolation: Using the corrector method. Statistics & Probability Letters, 83(11):2459–2466, November 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002484.

Zhao:2013:SLD

[ZZ13b] Shoujiang Zhao and Yanping Zhou. Sharp large deviations for the log-likelihood ratio of an α-Brownian bridge. Statistics & Probability Letters, 83(12):2750–2758, December 2013. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715213002083.

Zhang:2015:ESO

[ZZ15a] Fuxi Zhang and Wei Zhang. On the equilibrium state of the one-dimensional near-symmetric simple exclusion process. Statistics & Probability Letters, 98(??):20–28, March 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214004040.

Zhang:2015:ACD

[ZZ15b] Li-Xin Zhang and Yang Zhang. Asymptotics for a class of dependent random variables. Statistics & Probability Letters, 105(??):47–56, October 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715215001790.

Zhang:2015:ASI

[ZZ15c] Yang Zhang and Li-Xin Zhang. On the almost sure invariance principle for dependent Bernoulli random variables. Statistics & Probability Letters, 107:264–271, December 2015. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715215003272.

Zhou:2015:ELC

[ZZ15d] Mai Zhou and Shihong Zhu. Empirical likelihood confidence band for the difference of survival functions under proportional hazards model. Statistics & Probability Letters, 107:228–235, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215003077.

Zhang:2016:SDD

[ZZ16] Mei Zhang and Yong-Dao Zhou. Spherical discrepancy for designs on hyperspheres. Statistics & Probability Letters, 119 (??):226–234, December 2016. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715216301547.

Zhao:2019:MDP

[ZZ19a] Huiyan Zhao and Chongqi Zhang. Minimum distance parameter estimation for SDEs with small α-stable noises. Statistics & Probability Letters, 145(??):301–311, February 2019.

CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715218303250.

Zhao:2019:LDE

[ZZ19b] Shoujiang Zhao and Qianqian Zhou. On large deviation expansion for log-likelihood ratio of non-homogeneous Ornstein– Uhlenbeck processes. Statistics & Probability Letters, 155(??): Article 108560, December 2019. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715219302068.

Zhao:2015:RAE

[ZZLH15] Weihua Zhao, Riquan Zhang, Jicai Liu, and Hongchang Hu. Robust adaptive estimation for semivarying coefficient models. Statistics & Probability Letters, 97(??):132–141, February 2015. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715214003915.

Zhu:2017:VDA

[ZZLL17] Huafeng Zhu, Xingfa Zhang, Xin Liang, and Yuan Li. On a vector double autoregressive model. Statistics & Probability Letters, 129(??):86–95, October 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0167715217301773.

Zhao:2015:ELI

[ZZM15] Meng Zhao, Yichuan Zhao, and Ian W. McKeague. Empirical likelihood inference for the odds ratio of two survival functions under right censoring. Statistics & Probability Letters, 107:304– 312, December 2015. CODEN SPLTDC. ISSN 0167-7152 (print),

1879-2103 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0167715215003259.

Zhang:2017:EEP

[ZZT17] Zhenzhong Zhang, Xuekang Zhang, and Jinying Tong. Exponential ergodicity for population dynamics driven by α-stable processes. Statistics & Probability Letters, 125(??):149–159, June 2017. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715217300597.

Zhu:2013:ORU

[ZZZ13] Yunzhou Zhu, Lixin Zhang, and Yi Zhang. Optimal reinsurance under the Haezendonck risk measure. Statistics & Probability Letters, 83(4):1111–1116, April 2013. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (elec-

tronic).  URL http://www.sciencedirect.com/science/
article/pii/S0167715213000096.

Zhao:2018:OSF

[ZZZ18] Guoqing Zhao, Kun Zhai, and Gaofeng Zong. On optimal stopping and free boundary problems under ambiguity. Statistics & Probability Letters, 139(??):129–134, August 2018. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0167715218301457.